A Bibliography of Publications in Journal of Parallel and Distributed Computing: 1984–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 2.00

Title word cross-reference

(a, b) [DJM94]. (k, 2) [EMMM94]. (κ − κ) [KT91]. 0 [EE05, PMV05, PM96, SM89b]. 1 [EE05, HV09, PMV05, PM96, SM89b]. 2 [Ano93e, BDKM94, BAES92, CS92, CS93b, HSSM07, HHC98, KLC05, LME95, MD01, SS94b, WC91, WS95, Wu02]. 2 × 2 [PD92]. 3 [BAL05, BC94, CW00, CCCM96, GW99, Joh89, PEC95, WC91, Wan07, WS95, YB01]. 4 [MD01]. 4 × 4 [Jia99]. 5 [CCCM96]. ^{∗1} [HCZ04]. ^{∗2} [HCZ04]. ^{+} [OC07]. ^{,} [HCZ04]. _{2} [ASST05]. _{3} [ASST05]. B [YL89]. C^{3} [HK96]. C^{3}I [PAJC97]. d [DFN^{+}94, LSC00, VB94]. ⋄W [MRRT07]. g [BNP98]. h [GS98]. K [ACU08, BE95, DWG03, HHC98, SHL95, BVB02, DW06, DH91a, GP00, KK98a, PD05, PK04a, PRHB06, PK07, RP98, San99, SAOKM03, SGR03, SLP^{+}98, SZ00b, TT98, WS97b, YTH07, YD98]. k(n − k) [Lin03]. κ [XL95]. M [YLB90, WTB^{+}08]. N [AY89, IHM05, NTA96, SHT^{+}95, AKPT99, BVB02, GL90, NS94, PK04a, RP98, SAOKM03, WS97b, XL95, YTH07, YD98]. ∇^{2}G [CL85]. nn [PK07]. n × n [COS^{+}95, NS94]. O(1) [GP94, Wan07]. O(lg 2N) [BNP02]. O(log_{2}n) [JBL02]. O(log log N) [DP98]. O(log N) [GS99]. Ω [MRRT07]. P [BM97, PMV05]. P^{3}E [HSJP87]. P_{4} [ANP07]. ϕ [AK07]. ±2^{b} [Nas94]. q [DP00, Lat98]. QR [ZLRP91].

-ary [BVB02, DP00, Lat98, PK04a, RP98, SAOKM03, TT98, WS97b, XL95, YTH07, YD98, SHL95]. -Bandwidth [BM97]. -banyan [YL89]. -based [AK07]. -Best [BE95]. -Body [SHT^{+}95, IHM05]. -Chain [BNP98]. -connected [DW06]. -Cube [RP98, PK04a]. -Cubes [XL95, BVB02, SAOKM03, WS97b, YTH07, YD98]. -D [Ano93e, BAES92, CS93b, SS94b, CW00, GW99, PEC95, Wu02, YB01]. -delta [YL89]. -Dimensional [AKPT99, CCCM96, DFN^{+}94, VB94, KLC05, LSC00, SGR03]. -dominating [DW06]. -Extra-Stage [SZ00b]. -Item [San99]. -labeling [CP04a]. -Level [GS98, PRHB06]. -limited [WTB^{+}08]. -omega [GL90]. -optimistic [DWG03]. -page [HSSM07]. -Pairwise [GP00]. -Partite [EMMM94, SLP^{+}98]. -queens [AY89]. -reader [HV09]. -Reducing [GS00]. -satisfiability [Joh89]. -sparse [ANP07]. -Terminal [HHC98]. -Track [MD01]. -Trees [DJM94, HHC98, PD05]. -way [KK98a, ACU08]. -width [DH91a]. -writer [HV09].

/compute [KAS07].

0/1 [LSS88].

1 [HV95, MF94]. 1-Writer [HV95]. 10-Gigabit [HcF05]. 16S [ZFWF06]. 1D [PA04].

2 [ACYS08, AAL95, AR97, BLPV95, BSGM90, CDH84, DPSD08, FPD93, GH90, SI91, SMKL93]. 2-D [AR97, BLPV95]. 2000 [Wee01]. 2002 [Sni03]. 2006 [Ros07]. 2007 [Pan09]. 2008 [Rob09]. 2D [DFRCU99, TKHG04].

3 [BFG94, MKY^{+}97]. 3-D [BFG94, MKY^{+}97]. 3D [AB03a, CGW^{+}03, GS03a, MJ03, NPI^{+}96].

4 [BAM93]. 42 [Ano97c]. 46 [Ano97g].

5 [LAD^{+}96, PTC^{+}93]. 53 [Ano00d].

60 [Ano00b, Ano00c]. 66 [Ano93e, CS93b].

80 [Ano97k]. 802.11 [BCD00]. 802.11e [FA07]. 802.11s [VHH08]. 860 [DHR96].

90 [HLJ98]. 90D [BCF^{+}94]. 90D/HPF [BCF^{+}94].

A* [DM94]. a-cyclic [BD05]. A. [Ano92a]. AA1 [GCM95]. Abduction [eW95]. Abduction-Based [eW95]. Abductive [eW95]. Absolute [Wor93]. Abstract [CGSV93]. Abstraction [GDN^{+}98]. Abstractions [KB01]. Accelerated [Eİ07]. Accelerating [SHT^{+}08]. Access [ADS98, Bal90, BP02, Bit92, BR95c, CW93, CH92, DP00, FY96, HP00, OS93, San98, WMG01, ZRC99, BGLA03, BR91b, Che90, DFP06a, FA07, FC90, HC91, Lan09, MM07c, Pad91, SM89a, SR88b, SR90, WTS03]. Accesses [MRRV98, SR97a, SR97b, JZ05]. accumulations [SAF05]. Accuracy [EH01a, PKK91]. Accurate [DD95, KK88, HDT^{+}05]. ACE [PL98]. Achieving [EH01a, KEA95, NPY^{+}97]. Acknowledgment [KL08a]. Actions [WR95]. Activated [NPP^{+}02]. Active [SKH96, DB86, HOE^{+}09, PMV05, PMV06, YT05]. active/active [HOE^{+}09]. Activity [AS00, CW93]. Activity-Based [AS00]. actor [ASM09]. actors [GE85]. ActorSpace [CA94]. Acyclic [GY92, AFM09, BP89, Zim90]. Ad [Ano01e, GS01b, XG03, AP03, ALF03, BFG^{+}03, BGLA03, BOP06, BDF01, BN03, Bou03, CNS03, CW05, CYZ06, CDCD05, DW06, DMB^{+}03, DB08, EBE08, FVCL05, GAGPK03, GS03b, GMS06, GMXA07, HW03, HJ07, KK06, KNS06, LAZC00, LR03a, LPX05a, LW06a, LR03b, LHT08, OSL05, OMSGNSG05, Pat01, SSM^{+}06, SGS08, SKMM04, VA03, WTB^{+}08, WGS08, WBTM09, XHG03, XWC^{+}08, YC04, ZMC06]. ad-hoc [BOP06, CYZ06]. Ada [Lun90]. Adaptable [Zim96]. adaptation [BK08, GBMZ07, LS06]. Adapting [DKRI09, Wei02]. Adaptive [ASH^{+}01, AA93, AMN00, AYIE98, ACFK07, BLPA05, BPR99, BL90, Bou02, CS00, CLT96, DY99, DHB02, DMB97, DM99, FLS^{+}97, ISM07, JK00, KR97, KKGS01, KLLK98, KB01, Lan94, LLL06, LME95, LEB98, ME04, MV88, MD92, MTS90, OB98, OR97, PW96, PRS97, PIB^{+}01, RDS02, SS06, SKK97, SJ95, SB02, SSOB02, SLG06, SHT^{+}95, TC04, Ten90, WCE97, WA02, YIY97, ZM94a, AOSM05, AGMS04, BČFF05, CL03a, CP04b, CDCD05, DMB^{+}03, ESA03, GBA08, HNSA07, KMF^{+}05, KKS08, LY91, LA04, MCdS^{+}06, MSAF04, OPG08, OS04, SHLN09, TKHG04, TT07, WW04, ZWRI07]. adaptively [Mit07]. Adaptivity [OH02]. ADDAP [DHR96]. Addendum [Ano92a]. Adders [NIR86]. Adding [MSZ05]. addition [OB88]. Additional [LP97, CKN07]. Address [KY96, SL97, TR96]. Addressable [Win85]. Addresses [CGL^{+}95]. Addressing [ZLPP01, Ho91, TY90a]. adjusting [MC91]. ADM [Pad93]. admission [RKK06, XYDL06]. ADMs [FSZ07]. Ads [BA01a]. Advanced [BW95a, SD88a, TSD08, PLL^{+}03, SHT^{+}08]. Advancement [Lan09, LVR90]. advantage [CL03b]. advantages [CCLS94]. Adversarial [GBMZ07]. advertisement [WGC09]. advertisement-based [WGC09]. Affecting [DVW94]. Affine [DR95, DRR96, Dja06, DQR^{+}09]. Affine-by-Statement [DR95]. Affinity [TTG95]. after [DRR96]. against [SCC^{+}06, XCH08]. Agent [Ser97, FCC07, GZMC08, SS06]. agent-based [FCC07, SS06]. agents [AK06, CSWD03, KERUM04, MS05]. aggregate [Yan09]. aggregates [Chi95, Chi95]. aggregation [CDR09a, CDR09b]. agreement [GCS06]. Ahead [PL93, TG04]. AI [Ull84]. Aid [DBKF90]. aided [ZMC06]. air [FL86]. Airshed [SS00]. Algebra [CDH84, DVW94, KL01a, WM92, Joh87, RG87]. Algebraic [PL06, Pat01, BAH04, BM08, CM03]. Algorithm [AAP01, AE95, AM97b, AMS94, Als01, AS95, Ano93e, Ano96l, AS96, ABC^{+}09a, ABZ95, Bai94, BCC95, BGR96, BS97, BPST96, BOSW94, BE95, BDDL09, Bou02, BX93, BHRS95, CLZ02, CGKK97, CCM01, CB99, CSW08, CS93b, CP92, CTZ99, CF98, CRFS94, DA97, DM90a, DMB97, DS01, DS84, DH94, DSAUM99, DLP99, DT97, FY96, FT94, GGN93, Ger98, GRR93, GP00, GS99, Haw97, HH01, HBJ98, HO94, HM99, Hwa97, IZ95, JP95, Jia99, JK00, KRSZ02, Kar02, KSA95, KK98b, Kau94, KF95b, KS97b, KW02, KA97, KC99b, LP96a, LO94, LHVW95, LP97, LWP02, MT97a, Mil99, MV94, MSST99, NTA96, NM02, Par98, PE93, Par96, PL94, PB95, PM96, PRS97, PM92, RR95a, RP95, SAOKMA02, SZ00b, SCC92, SR94, Shu95, SM00, TU92, TZ00, WSRM97, WD94]. Algorithm [WA02, WLID02, XWC^{+}08, YZY96, mYyF92, ZB97, AOS^{+}05, AT03, AK07, ATH91, AGMS04, Ara90, BFG^{+}03, Bad04, BC05, BČFF05, BSG90, BH05, BBL04, Cal06, CR91, CM03, CV90, CS92, Che89, Cho90, CZ90, CRC^{+}02, DK08, DDNS06, DB08, DM90b, DB86, Ebn04, EE05, EDÖ05, Fei03, FSZ07, GPX08, GGR89, GT04, Gue86, GB06, HJ90a, HRJ94, HLM^{+}90, HL07, Kal04, KK06, KM03, KA91, Koc91, LVP08, LSS88, LMZ04, LO91, LB89, LP88, MD07, MM07a, Mar88, McD89, MMS09, MM07c, MP08, MMS90, OS04, OT86, PK05a, PHS04, PB09, QJ05, RH05, RGD03, RKS87, SSTP09, SCJ^{+}08, SA08, SKK91, SM08b, TKHG04, WSH^{+}03, WJV07, Wan07, WG08, WGC09, XHY07, XQ07, Yan04, YME06, YZLT09, ZZ90, ZFWF06]. Algorithm-Based [GRR93, mYyF92, BDDL09, LP88]. Algorithm-system [CSW08]. Algorithmic [Gao89, SCB08, LS05]. Algorithms [ANT02, ABJS01, AKP95, ABM^{+}92, BJ96, BJ99, Bah00, BPJG92, BLPV95, BGJDL02, BAES92, BAGS95, BBM^{+}02, BSDE96, BOP06, BPR99, BSS99, BMRC98, BMRC99, Bro96, BA01b, CTD99, CDY97, Cha94, CGO^{+}96, CDH84, COS^{+}95, CN93, CP91, CHR94, CWP98, CA95b, DS95b, DP98, DHB02, DP99, DM92, DMSH90, DFRCU99, DBKF90, DKMV01, EP90, ESMG96, EMMM94, EL97, Fer95, FR96b, FA95, FV97, FTC00, GG94, GP94, GV94, GM96, GHSJ96, GMM00, HHM94, HQPT99, HCWS94, HR92a, HP97b, HTB98, HO94, IK93, IK94, Iqb92, IM00, JW94, JS94, KRC00, KAM94, KLZ97, KG94, KA99, LHS97, LSH96, LHHB^{+}01, LLCC02, MB96a, MMRS98, MS94, MMVR97, Man97, MT96, Mat93, MHC95, MK92, MS98, MS99b, Nak95, Nas94, PAH^{+}98, PAJC97, Pov99, Pra93, QZ94, QOvdG01, RS96a]. Algorithms [RR95b, Raj01, RSS96, Ram92, RDS02, RSW90, SH90, SS96, San95, San99, San02, SZB92, SY01, Sto90, SYG92, Ten90, TVS97, TC96, ÜD96, VB94, VR95, WNA^{+}94, WR97, WA02, WD92, WN94, WT92, WHT00, WHT02, YMR93, dBL95, AL04, ACFK07, BC06, BS87, BAS06, BOS^{+}91, BFG04, BRPR06, BPP05, BM08, CM04, CF88, CNS03, Che86, Che05, CRA^{+}08, CB06, DS04a, DH91a, Dja04, Dja06, DJT03, DM94, Fen90, FBRW03, FGG08, FJSW90, FM85, FVCL05, GP07, Gos90, GH89b, GWH06, GS03a, GC07, Han89, HSSM07, HSW04, IC05, JMS86, JBM91, KJD03, KS08, KAP90, KKB^{+}06, KS91, KMP^{+}06, LW90, LLL06, LW06a, LS88, Lin91, LS91, LS03, LLW07, LA04, LVB07, LGG08, LV88, MM04, MPZ09, Meg91, MCT06, MM07b, MS88, MGG03, MVV91, MAR87, Nik04]. algorithms [PD05, PY09c, PL03a, PA04, PRG88, PS88, RTCG91, SSM89, SS06, SM89b, ST87, SAF05, SZW05, SGS08, SD88b, Sto87, TY90a, TW87, TK08, WC91, WCWH03, Wri91, ZV09b, dVCP06]. Align [BR95c]. aligning [LVB07]. Alignment [BRR01, CGO^{+}96, DRR96, Mil99, MJ01, SS94a, BBM08, BR91b, BMARM07, LC91a, PTZ06, SK09]. alignments [BW09, ST85]. All-Output-Port [ST02, ST06]. all-pairs [KS91]. All-Port [RJMC95, Dim04]. all-reduce [PY09c]. All-to-All [HP95, LHS97, LWP02, Ede91, LR03b]. Alleviating [Tze91]. Allocating [BPRG04, Hag97, SEP96, SCS^{+}08]. Allocation [AM97b, AERBL92, CS00, yCM98, DSST95, DY99, DL99, DL01, Hwa97, KKGS01, KLS90, Moh96, NSS97, OM84, PT01, SM94, SdS97, SP96, YL98, Zhu92, ALH^{+}09, AKSM08, ATZ07, AH06, BMB^{+}08, BG86, Bat05, BSMH08, BPW05, DM90c, ERS90, GNT04, GRDB05, LHF91, LC91b, Li05, MCC04, PM05, PBS08, RLH03, SSM^{+}06, SSM^{+}07, ZI08]. Allocations [BE95, CT96, SSMS08]. Almost [JBP00, SS95]. Alphabetic [LP96a]. alternate [LS03]. Alternating [BC94]. Alternative [GW99, Pad93, CBV08, GB06, Ros85]. Alternatives [BAHP01, NBSD99]. alternator [LW06b]. ALU [KF90b]. Always [BRR01]. Among [OO85, GM94b, KS03, MT93a, NMS93]. AMR [GWH06]. Analyses [KY96]. Analysis [Abr96, Ano92a, BCV94, BCF97, BN94, Bhu87, BDF01, BLG01, Buc92, CK88, CC91, CAB94, DLLX97, ES96, Fra92, GM94a, GSG^{+}93, GCM95, GC01, HLM^{+}90, HC97, HF96, IM94, JV09, KME92, Kop97, LW89, MF94, MT93b, MM93, MS99a, MRR^{+}02, MT96, MDD97, MHBW86, NBM93, NSM98, OD95b, OS93, PD92, Piu01, PAJC97, RPS93, RKS87, SM89a, SLP^{+}98, SWP90, SHC93, ST08a, VSM96, XL92, ABC^{+}88, BČFF05, BFG04, BM08, CK06, CH06b, CWL^{+}07, CPO^{+}03, FC90, FCS91, FD86, FX06, GBA08, HRC09, HA91, IKS87, IC05, JT88, JBM91, KME89, KA08, KG04, KLL87, Li06a, Li06b, LH05, LP88, MM06, McD89, Pak89, PL06, PRHB06, PI90, Pfe90, PL03b, RM90, RGU08, TMM06, WSH^{+}03, WF89, Yan09, YH07, ZFS07]. Analytic [BS96b, BS96c, Har91]. Analytical [DG94, HW03, QY94, SAOKM03, AP91c, Bat05, BFH09]. Analyzing [CDR09a, CMT92, HcF05, KG94, LMCF90, MSH90, MBH^{+}08, WXZ05]. Anatomy [ZBF05]. Anchored [KS03]. AND-parallelism [DeG88]. AND/OR [RP95]. Animate [MBL^{+}92]. Animation [RGS00]. Anisotropic [PSE^{+}01, Eİ07]. Annealing [Bev02, BA92, HB97, RSS96, Soh96, XH91, AH06, BG89]. Annotated [KBC^{+}01]. Announcement [Ano93a, Ano96k, Ano01c, Ano01d, Ano01e, Ano01a, Ano01b, Ano02a, Ano02b, GHS96, Kai92, Ano00a]. Anomalous [MSH90]. anomaly [DFP06b]. anonymous [AFM09, FKK^{+}04, MSJ05, XLG^{+}06]. ANSWER [OEY07]. Ant [RL02]. antenna [CCHC09]. Anticipative [WLID02]. Any [RCY97]. APHID [BS00]. Appear [Ano00e, Ano00f, Ano00g, Ano00h, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Ano01-32, Ano02q, Ano02r, Ano02s, Ano02t, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p]. Application [AS97, AYIE98, BB03, BSS97, CKK00, CCC92, ES96, HMV07, Kop97, PH00, PP92, Ser97, SM92b, SK93, BCM06, CP05, CD95, DBC03, DKRI09, FGM^{+}03, GP91, KME09, LS06, OSL05, PL06, PGS06, SL90, SFT04, SS94b, VD04, dGP06]. Application-aware [HMV07]. Application-based [BB03]. application-level [VD04]. application-sensitive [CP05]. Application-Specific [PP92, SK93, SS94b]. Applications [ABDS02, Ano96i, AFT^{+}00, BOSW94, BMRC98, CCRS92, CA95a, CDF01, DRC90, DS84, EH01a, FR98, FBK98, GCB^{+}00, GT02, HS94b, KR97, LLS93, MHC95, MB92, MBK^{+}92, NB93, NsPPC02, OS96a, RS92c, SSOB02, UZSS96, VH93, WMG01, Wei02, AKSM08, ARM^{+}05, AGMJ06, BBCLL04, BAS06, BM04b, CCC^{+}04, CC08, CP05, CBM^{+}08, CCM^{+}06, Dim91, EDÖ05, ESA03, GTN^{+}06, GST09, GJA08, HC09, HSLL04, HA91, HL07, KJD03, KHK03, KAS07, Kri91, MMAL^{+}06, Oza04, PLL^{+}03, PF04, SV08, SM89a, SCS^{+}08, TDM05, Ull84, VB08, VM03, YH07]. Applied [CB96, BDDL09, EE05, HSLL04, PR06]. Applying [PEC95]. Approach [AAL95, AM93, Bev02, BR02, BST01, CCM92, CY95, CLZ00, DM95, Fer92, FKT96, FKKC97, GG94, GZ97, HC97, HLJ98, KCRB99, KSB94, LS95, LW95, LLCL98, MSSE02, RJY96, RAS96, SL95, SP96, SZ00a, TC92, WSRM97, WA02, Won99, WLID02, AP91c, Ara90, AH06, BAS06, BW09, CvdBL^{+}08, DBC03, DQR^{+}09, FZC^{+}05, FGZ03, GZ08, GWWL94, GBA08, JLM08, Joh89, KMS^{+}06, LH04, LC07, MS05, MSM09, SU87, SCS^{+}08, TM06, TBZB05, VB08, YF09, ZFL89]. Approaches [CHGM01, QM01, KERUM04, KA05, PR06, dGP06]. Approximate [JSS92, KERUM04, MM07b]. Approximating [FMM^{+}08, PBS08]. Approximation [FV97, HP97b, Mat93, LVP08, LW06a, MK08b]. Approximations [Gon98, BFM06]. AQOR [XG03]. Araneola [MK08a]. arbiter [Bhu87]. arbitrarily [ZV06]. Arbitrary [ERL90, KA97, SS95, YZY96, Ara90, SGE91, Wag89, FII04]. arbitration [ASD09, KS03]. Arc [CA95b, Ros89]. architecting [CCC^{+}04]. Architectural [DZDZ01, GSP02, HPT^{+}97, KC99a, MT96, MG93, WSS93, FZC^{+}05, JBY^{+}05]. Architecture [AGW01, ABZ95, BBD^{+}91, BAHP01, DH95, Gao93, Ger98, GBES93, GM95, HP97a, HGCC96, IWM97, KC94, LBL95, MWL00, MS00, MAM05, MKY^{+}97, MO97, MT85, NEG85, OD95b, OY00, Pad93, PS01, STN92, SSYG97, SH98, VS99, ZYH94, Zim96, ACYS08, AC89, BB87, CCQ^{+}06, CTCX08, CCEB03, CDJ^{+}89, FCS91, GHS86, JS86, JXW06, KRL87, KH89, LAD^{+}96, LZSL06, MM07b, MBH^{+}08, MP08, NW88, PK05b, SDTD04, SR88a, SAB^{+}92, SR91, WL92, XJS03, ZV09a]. architecture-based [CTCX08]. Architectures [AGW98, ABDS02, BBR94, CCM92, CCC90, CT93, CS93c, CP01, CBdCD00, DUSH94, DMSH90, DS02, DT01, DRSB01, DT92, EP90, EL97, FY97, GGB93, KS95, KM97, KG94, LB90, LC90b, LR93, LR94, MSd^{+}95, PP96, PA94, PD92, SH90, SS94a, TG99, ZMPE00, ZL93, AP03, ABC^{+}09a, ABC^{+}09b, BS87, CCK88, Che86, CkLCK04, CkLCK05, CPO^{+}03, GSWW04, GS91a, HSW04, Joh87, Joh91, KF90a, LM05, LS88, LVB07, Pad91, PR06, PLD87, RTCG91, SLG06, SS94b, TKHG04, VM03, WJD91, vS91]. Area [BCD00, CLR90, KF95a, NIR86, Wei98, HZY04, HL07, KCD08, KMF^{+}05]. Area-Time [NIR86, CLR90]. Arithmetic [AK93, CL88, DPRW85, Gro85, Irw88, KK88, KM88, SR88a, Sch87, Sil90, SL90, Tay87]. Arithmetic/Logical [AK93]. Arnold [Ano00d]. arrangement [Lin03, NAK04]. Array [AW95, BCF97, BL90, CT93, CWW^{+}95, ER97, GKHS96, GE94, HQPT99, HCS^{+}00, HCZ04, HLJ98, HLJ01, KRW96, KHS96, KC98, KR87, LP96b, LTH97, Mil99, MJ01, MBK^{+}92, MT97b, OM90, RSB96, Ste95, SOG94, Tse90, WSS93, Win85, dR09, BB85b, BPP05, DS04a, GP05, Lee91, MM07b, NAK04, PLD87, SI86, ST87, SCC^{+}06, YTH07]. Arrays [Ann94, BAGS95, BPST96, BP02, BR95c, CGO^{+}96, Cor93, GP93, GW99, Guo94, IPK85, KLS90, KEA95, KL84, KBG92, MM00, MD01, MT93b, MRK93, MFS93, MFS96, RFM94, RCB93, Swa98, TBPV00, TC96, WCF94, WHT00, BBd90, CL03b, DMCFCM03, Deh90, Dja04, Dja06, EL91, GMH^{+}91, KT89, KT91, KLL87, LB89, Lis90, OT86, RIZ90, SSM89, Sch89b, ST89, SKK91, Ume85, WAS88]. Art [KM92, WCO^{+}09]. article [Ano96l, Ano97k, Ano00d, CS93b]. artifacts [LZ08]. Artificial [MT85, NS92, Piu01, TVO92, KH89, VO89, VM95]. ARTS [BNSP99]. ary [BVB02, DP00, Lat98, PK04a, RP98, SAOKM03, SHL95, TT98, WS97b, XL95, YTH07, YD98]. AS008S [Ano04c]. ASAT [SEP96]. ASCEND [Nas94]. Aspect [BZLI04, MO97]. Aspect-oriented [BZLI04]. aspects [Gao89]. Assembling [KESA07]. assembly [ABCM07]. Asserting [ASST05]. assessment [LY08]. Assign [CYZ06]. Assignment [HBCM99, HB97, KLZ97, SS93, Ste95, VWHL96, WW97, Bat05, BPRS04, QGL^{+}09, UAKI06, WZ91]. Assignments [LL98, Sin87]. Assisted [HlLLY95, LVP07, RG06]. Associative [AA93, DM92, NSM98, Par96, PL98, VR94, Kri91, LL90, SR88a, SI89]. Assurance [BK08, WLL08, XHY07]. Asymmetric [BNS00, ZR00]. asymmetry [AP91b]. Asymptotic [GM94a]. asymptotically [Dja04]. Asynchronism [ÜD96]. Asynchronous [Bah00, BSS99, BS00, CS95c, CA95b, ESMG96, MS02, MM93, MR94a, MR94c, OY00, The02, WT92, BB03, CRC^{+}02, DGFGK05, DB86, FKK^{+}04, MM04, RLH03]. Asynchronous/Synchronous [OY00]. asynchrony [WCYR08]. ATExpert [KW93]. ATM [WR97]. Atomic [HV95, JBP00, WR95, van96, GNS09, HV09]. Atomicity [NA02]. attack [JXW06]. attacks [CH06b, KMMZ06, SCC^{+}06, XYG07, XCH08]. attributes [Par05]. Augmented [MKY^{+}97, Lo92]. Auralization [FJ93]. Aurora [Lu01]. authentication [CL09, NC09]. Author [Ano92b, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98l, Ano99a, Ano99b, Ano99c, Ano99h, Ano00b, Ano00c, Ano01f, Ano01g, Ano01h, Ano01i, Ano02c, Ano02d, Ano03a, Ano03b, Ano04b, Ano04a]. Author-Title [Ano98l, Ano99h, Ano00c, Ano01h, Ano01i, Ano02d, Ano03b]. automata [MS86]. Automated [NM95, NC97]. Automatic [ABCM07, CGO^{+}96, DHR96, KBC^{+}01, LC92, MJ01, SEP96, AAD05]. Automatically [DR98, TG99]. automaton [Cap87]. automorphisms [DH91b]. autonomic [ATZ07, CP05]. autonomous [XCH08, ZV09a, OEY07]. autonomy [LFH^{+}03, ML89]. Availability [HJD^{+}01, LS01, DB08, HOE^{+}09, PF08]. Available [NKC^{+}97]. Average [DF95, Li06b, MDD97, NSM98, Li06a, XBK07]. Average-case [Li06b, Li06a]. AVL [MD98]. Avoidance [MJ94, BB85a, BPRS04]. Award [Ros07]. Aware [ALF03, BPA06, EB09, FGZ03, HMV07, HK05, HK04, KDH08, NP09, OS04, SK05a, XCZL03, ZV09b, ZC04]. awareness [LR03b]. Azriel [Ano04r].

B [CWW^{+}95, CY96, GM95, HS94a, Meg91, OC07, PPC04, WW96]. B-Spline [CWW^{+}95, CY96, GM95, Meg91]. B-Trees [HS94a, WW96, PPC04]. back [HPSM91, KMMZ06]. back-end [HPSM91]. back-propagation [KMMZ06]. backbone [HWWH08]. backbones [KERUM04, XHG03]. Backfilling [SF05]. Backplane [SH98]. backpropagation [SM08b]. backup [AOSM04, HOVC09]. Balance [SEP96, CCK88]. Balanced [GJP96, LT94, NFEG97, PB99, SA93, SBAM96, BNP02, LCW05, XYKA08]. Balancing [Ano97j, BEE00, DHB02, DMB97, DLLX97, DSW94, Efe96, FMP98, FLS^{+}97, FM99b, Gil94, GM96, HlLLY95, HTL99, HO94, HC97, JR92, KGV94, LK94, LHVW95, MP96, NSLK99, OB98, QY94, SH92a, SHT^{+}95, SB97, TSHH01, Wan96, WS97b, XL92, XH93, XL95, ZLP97, ZM94b, AGMS04, BCV05, BFH09, BRPR06, BD04, BM08, CSWD03, CBD^{+}09, CRC^{+}02, Cyb89, DM94, EE05, Gao89, GC05, HJ90a, HLM^{+}90, IC05, JL05, JW89, KKS08, KC04, LTB02, LTL06, LHKL03, Mit07, Nik03, PA04, RN04, SX08, ZV06, ZXP09, dG91, vS91]. Banded [Pov99, ORR03]. Bandwidth [BM97, Cha95, PY09a, PY09c, CCHC09, DK04, HJ90b]. Banerjee [PKK91, Psa96]. Banerjee-Wolfe [Psa96]. bank [QGL^{+}09]. banker [MMS90]. banyan [PL06, Kop97, WN94, Yan00, YN92, YL89]. Banyan-Hypercube [YN92]. Bareiss [HM99]. bargaining [GRDB05].

Barnes [SHT^{+}95]. Barrier [Cha95, JLRA97, OD95b, RSS99, XMN92]. Base [DKMV01, RBD08, DDNS06]. Based [AE95, AS00, Ano99g, BCD95, BPJG92, BGJDL02, BMM97, BN02, BR02, BA92, CGKK97, CC91, CRV94, CS95b, CKL99, CGA98, CHGM01, DA97, DR98, FF98, FKKC97, GS01a, GRR93, Gup92, GS01b, HP00, HB97, HK01, HSJP87, KCRB99, KSP^{+}92, KCDZ95, Lat95, LAZC00, LZ02, MSC96, MB93, MG98, NTA96, NB93, NM02, OM84, Pad93, PN97a, PN97b, PA97, PL95, PM96, PAJC97, RL96, RSD94, RMC97, RSBN01, SMR96, SSRV94, WLY01, WSRM97, WSA^{+}94, Won99, WLID02, XH91, mYyF92, YB01, Zia92, eW95, AL04, ASM09, AK07, ARM^{+}05, ABC^{+}09b, ATZ07, BCM06, BJPPM^{+}08, BB03, BDDL09, BM08, CL03a, CK08, CTCX08, Chi95, CL09, CVJ09, CHC05, CGW^{+}03, CZZY09, DE91, DRST02, DRT07, DQR^{+}09, EDÖ05, FCC07, GBA08, GMXA07, HW03, HV09, HC09, HLM^{+}90, JXW06, JP09]. based [JBY^{+}05, KKVI05, KERUM04, KJD03, KA08, KR06, KC04, LHKL03, LLLY08, LL07, LS03, LHT08, LVB07, LS06, LP88, MCC04, MCdS^{+}06, MMS09, Mit07, MM07c, MBH^{+}08, MRRT07, NJ91, NC09, Nic07, NAK04, PRP09, PK05b, PVPM06, PF04, SS06, SP08, SX08, She09, ST85, TR89, TT07, VD04, VB08, WCC02, WGC09, WLWW09, XHY07, XCLR07, XO05, ZCK^{+}02, ZV09a, dGP06, SM92a, WAS95, ZNQ93, HC91, KKS08, PLD87]. bases [GPT06a, SK90]. basic [BM04a, Joh87]. Basis [TR96]. Batch [LL98]. batched [CK06]. Batcher [NT93]. Batching [DSST95]. Bayesian [FBRW03]. be [BNP02]. beacons [TDC05]. Beamforming [BL90]. Behavior [Abr96, BDF92, BN02, BST01, CMT93, FJ93, LZ08, BS92, JZK04]. Behavior-Based [BN02]. Benchmark [PAJC97, GREC91, Num07, Num08, Num09, WRHR91]. Benchmarking [KA99]. Benchmarks [WAS95, JV06]. Bends [OS97]. Bene [CI03]. Benefit [Wei02]. Benefits [FR92, SS99, Wei98, GK04]. Benes [DD96, Qia97]. Best [BE95, Rob09, SP96, Sni03, Bar05, FPP^{+}08, MAM05, WAE03, Ros07]. best-effort [Bar05, MAM05]. Best-Fit [SP96]. better [AM06]. between [BVB02, BJS03, CG86, DB86, FII04, KNS91, LDCZ97, PHS04, RGD03]. Biconnected [Kar02, Hoh90]. bicriteria [BFG04, BFM06]. Bidimensional [BP02]. Bids [BA01a]. Bimodal [KC95]. Binary [AS94, CS95a, DS93, Efe96, HIKM94, HKMU98, HM01, HR92a, Iqb92, JH94, LP96a, Li92, OOW95, SYO94, Wag93, BL89, DH91a, Wag89, HRJ94]. Binding [LBL95]. Binomial [DP00, WFL98]. bioinformatics [TZ06]. bioinspired [MPZ09, MCT06, dVCP06]. biological [AFM03, BBM08, BA06, BW09, BMARM07, SK09]. biology [AB03b, TZ06]. Bipartite [DS84, LPS^{+}98, SM89b]. bipartitioning [ERS90]. bis [Fen90]. bis-sequential [Fen90]. Bisectors [BEE00]. Bit [HPT^{+}97, MO97, MT97b, SI91, CL90, Ede91, GPX08, KM88]. bit-pipelined [KM88]. Bit-Rate [MO97]. Bit-Serial [MT97b, SI91, CL90]. bit-substitution [GPX08]. bitonic [FCS91]. Bits [GH96, HV09]. bivalued [Zep91]. Blackboard [CC91]. BLAS [HWW96]. BLITZEN [BDHF90]. Block [CT96, FBK98, GHSJ96, PT97, WSA^{+}94, BW08, GPX08, KR06, Sch87, SZW05]. Block-Based [WSA^{+}94, KR06]. Block-Structured [FBK98]. Blocking [BHK^{+}94, QS05]. Blocks [CWW96]. Blue [FGM^{+}03]. BlueCube [CCS06]. Bluetooth [CCS06, SLWW05, WTS03]. board [Ano02e, Ano02f, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q]. Body [HP95, SHT^{+}95, IHM05]. Boltzmann [KA89, WCO^{+}09, ZA91]. Boolean [HJ90c, JH92b]. boosting [FGP05]. Border [DRST02, HR90]. Border-based [DRST02]. both [WAE03]. Bottleneck [WW98]. Bound [GZ97, PM96, CH06a, MCC04, SCS^{+}08, SW90, YZLT09]. Boundaries [Wor93]. boundary [Lin91, RBD08, SCC^{+}06]. Bounded [AW95, BBN93, CLT96, GP97, Pra93, SN93, BD05, BPRG04, LMZ04, MRRT07, NP09, TK07]. Bounding [Lun99]. Bounds [ADS01, BBH^{+}98, DL98, JR95, LPS^{+}98, LP95, Lun94, WW97, FT04, FSZ07, ITT04, KMS07, Mat06]. Branch [GZ97, MCC04, PM96, SCS^{+}08, YZLT09]. Branch-and-bound [MCC04, SCS^{+}08, YZLT09]. Branches [ERA95]. Breaking [FJ93]. Breakpoint [dADB96, MT97a]. Bridge [HR00]. Bridged [EAL90, LCM^{+}06]. bridging [BJS03]. Broadcast [DHB02, OS96a, Pel95, RS96a, RS92c, San99, VB94, BG05, FVLB09, KGN89, MRRT07, PYF08, SGS08, TR08, WL05]. broadcast-based [MRRT07]. Broadcast-Efficient [OS96a]. Broadcasting [BNS00, BPvW96, BMMS01, BOS^{+}95, CW00, CCC92, DLP99, Fra92, FV97, GP97, HIKM94, Lat98, ST02, ST06, SCD99, Wu94, dBL95, oPP00, Che05, CMS04, FMR05, HS06, Ho91, KR87, LR03b, SZ03, Wu03, ZA05]. Broadcasts [WD92]. Broker [HR00]. Browsing [SF90]. Bruijn [ANS97, CT96, FT04, HOS94, MVM04, Swa98]. Brzezinski [Ano96l]. BSP [CTZ99, GS98, GLC01, HH01, HM99, KP00, RGD03]. Bubble [DF94, PIB^{+}01]. buddy [LC91b]. budget [dR09]. Buffer [FM99a, HV95, MSSE02, PY09b, WLID02, BPW05, HV09, PBS08, SCC^{+}06]. buffer-based [HV09]. Buffer-Optimal [HV95]. Buffer-Safe [FM99a]. Buffered [AA95, KJ84]. bufferless [BMIM07]. buffers [DW04, HM06, WAS88]. Building [Haw97, IK93, SK93, CZ90]. Bulk [GV94, Lu01, FXW03]. Bulk-Data [Lu01]. Bulk-Synchronous [GV94]. Bus [CKL99, DVZ96, FZVT02, FY96, GK98, LPZ99, TVS97, VB02, dR09, BPP05, CLM90, DS04a, JSWB92, MS88, MHBW86, YB90]. Bus-Based [CKL99]. Bus-Connected [DVZ96]. Buses [CL96, HQPT99, IM00, KC98, LS94, NS94, TVT96, TBPV00, WHT00, ZLPP01, Car90, JW89, KRL87, Mat06]. Busy [SP96]. Busy-List [SP96]. butterflies [CI03]. Butterfly [JH94, TDM05]. Byzantine [CBV08, PP06].

C [CD98, DZDZ01, LS85, ZH99]. C3 [Ano04c]. C3- [Ano04c]. CA [Chi95]. Cache [DS95a, Dah99, GS96, HP97a, LY98, LY01, LF92, NB93, PL95, PY96, RL96, San95, TTG95, Yan93, BW89, CWLD05, DK04, GJG88, GVA^{+}08, HZY04, HC09, HSMB91, SYYU07, YCC05]. Cache-Affinity [TTG95].

Cache-Based [RL96]. cache-coherent [SYYU07]. Caches [DS95a, YAS98, RFPAG08, SD91]. Caching [BS96b, BS96c, KC99a, KE93, MM93, BLPA05, CR96, LVP07, OC07, TC03]. calculation [SL90]. calculations [HT90]. Calculus [PL98, SC95]. calibration [MMAL^{+}06]. Call [Ano95i, Ano95j, Ano96j, Ano96i, Ano97i, Ano97j, Ano98i, Ano98j, Ano99g, Ano99d, Ano99e, Ano99f, Ano00a, Ano01c, Ano01d, Ano01e, GSC96, RKK06]. Calls [Ano98k]. can [BNP02]. Capabilities [Fra92, MMRS98, TVO92]. Capability [Guo94, SP96, YWP00, Ho91, HK04]. capability-aware [HK04]. Capacitance [YB01]. Capacity [ACD^{+}93, MO97]. capture [JXW06]. Capturing [ISAZ07]. cards [KME09]. Cares [BL94]. Carlo [Bro96]. carried [NCT^{+}07]. Carrier [DL01]. Carry [NIR86]. Carry-Lookahead [NIR86]. Cartesian [ANS97, Dim04]. carving [RRS^{+}08]. Cascaded [Wil90]. Case [BA01b, GT02, HPT^{+}97, MS99a, NSM98, SSG93, WNA^{+}94, WLR90, CCK^{+}08, DI91, GRR^{+}05, HA91, Li06a, Li06b]. Categories [Cou93]. Causal [CLZ02, MT97a, PRS97, RS92c, CZZY09, FJC04]. Cayley [BS03, WLD00]. CBIR [BRPR06]. CBT [GS01b]. CBT-FR [GS01b]. cDNA [TMM06]. CEFT [ZJ06]. Celeste [BCK^{+}09]. Cell [CB99, LTKS90]. cells [Lis90]. Cellular [CS00, DL01, DKMV01, Oru87, Tan84, ZR00, FCG04, GMXA07, MAM05]. cellular-based [GMXA07]. center [FP03]. centric [KSI04, XCLR07]. CFD [BAMM05, Kal04, MS99a]. CGM [KP00]. Chain [BNP98, Lun94, GRV08, MVB05]. Chains [NH93]. challenges [SAB^{+}92]. changes [DB08]. Channel [AM95, BNS00, BPRS04, BKT95, CS00, DSST95, GCKM97, HP00, JK00, KKGS01, LM96, PA97, BGLA03, CCHC09, CLL09, DRT07]. channel-based [DRT07]. channels [CK06, KS03, Lee03]. Characteristics [LHVW95, GF89, JV06, LTD^{+}93, LF03, SCK03, VM03]. Characterization [BF01, KS94, MR94b, RJA97, WP02, LJ86, SR90, WH08]. Characterizing [MS96]. Chare [SK91]. Chasing [YZY96]. checking [CM04, SSS07, SCC^{+}06]. Checkpoint [PT01, JLM08, MM04, PGS06]. Checkpointing [PKD97, WF96, AAFV04, JLM08, LM09, MM06, MM07a, QS05]. Checksum [Par92]. CHEMAS [XYG07]. chemical [MMAL^{+}06]. Cheng [Ano93e]. chessboard [Eİ07]. Chip [ASH^{+}01, MT97b]. Choice [SB02, BL05]. Choices [FR96a]. Cholesky [MVV91]. Choosing [HBCM99]. Chordal [Man97, WT09]. chordal planar [PD05]. Cilk [BJK^{+}96]. cipher [GPX08]. Circle [KSB94]. circles [Wri91]. Circuit [CB99, CCR94, CS93c, GGN93, LK96, EB09, YTH07]. Circuit-Partitioned [CB99]. Circuit-Switched [CCR94, CS93c, GGN93, LK96]. Circuits [KM97, BAH04, LH04, LS05, LH09, TT07]. Circular [BP02, CDP95, JT88, RGU08]. circulation [PV07]. clairvoyant [Li06a]. Class [BNP98, BSB^{+}01, CAB94, CN93, HR00, LYL93, MAS^{+}99, Nas94, TL96, WN94, WLD00, FY86, LLS07, Pak89, SP90, Ume85]. Classes [Par98, LLL06]. Classification [DSAUM99, BCM06, Bod89, DH04]. Client [GM99, HC09, ST08a, TC04]. Client-Server [GM99, HC09]. client-side [TC04]. Clients [ALL99, Yan09]. Clique [FTL92, SSTP09]. cliques [CKO04]. Clock [ASB97, PD92, PB95, PB09]. Clock-Regulated [PD92]. Clocks [DKMV01, YH97, AKD06]. Cloning [DDD98]. Clos [HJDH01]. Closed [TR96]. Closure [YMR93]. Closures [AW95]. cloth [GRR^{+}05]. Cloud [CDJL09]. Cluster [AFT^{+}00, BAHP01, GS01a, HS00, JM00, LS01, MKC01, PT01, ARM^{+}05, BMARM07, FW05, HW03, MAR05, MSJ05, MBH^{+}08, OC07, PSPR05, PVPM06, SAOKZ05a, SAOKZ05b, TC04, VM03, ZBF05]. cluster- [SAOKZ05a, SAOKZ05b]. cluster-based [HW03, MBH^{+}08, PVPM06]. Clustered [CP99, MF94, HRC09, Wan06]. Clustering [ASM09, GY92, HJ07, TZ07, WSH^{+}03, WHT00, GWH06, LLW07, MCC04, RIZ90, SX08, WMW09]. clustering-based [MCC04]. Clusters [AYI97, BJ99, BP01, BDH^{+}97, Dek00, KMKD97, KR98, LC97, PN97a, PN97b, WB96, Wei02, BČFF05, BJS03, FMR05, GJA08, KYL05, ME04, PYF08, PY09c, QJ05, QS05, SM04, TC03, WLNL06, WLWW09, YH07, ZB09, ZI08]. CM [BSGM90, LAD^{+}96, PTC^{+}93, Sab94, SI91]. CM-2 [BSGM90, SI91]. CM-5 [LAD^{+}96, PTC^{+}93]. CMPs [DKRI09]. CMV [WDDK09]. CO [ASST05]. Coarse [BR96, BM04b, CDRC99, DFRCU99, HK96, NS97, SR97a, SR97b, TF01, CT94]. Coarse-Grained [BR96, CDRC99, HK96, SR97a, SR97b]. Coarsening [DR98]. Code [Bec96, FK89, JH94, NS97, RNSB96, BCM87, Gao89, LS06, SY04]. code-based [LS06]. Codes [BVB02, Lat98, GRR^{+}05, HR90, LWR^{+}03]. Coevolutionary [Ser97]. Cogenerator [KSP^{+}92]. cognitive [MKC^{+}09]. Cographs [LO94, LO91]. Coherence [ABP92, CKL99, DS95a, DSS95, GS96, HP97a, HF96, KS95, LY98, LY01, PL95, San95, SDS99, FGP05, GVA^{+}08]. Coherence-Miss [SDS99]. Coherency [TJ92]. Coherent [PY96, SYYU07]. Coincident [ZLPP01]. Cointegration [THN^{+}93]. Coir [SG96]. collaboration [ABCM07]. Collaborative [CH06b, WW07, FM07, GCS06, Wan06, XQ04]. Collapsar [JXW06]. Collection [BS90, KS00, RW01, HMV07, JLM08]. Collection-Oriented [BS90]. Collective [DT01, HK01, TSC01, BRP03]. Collision [YB95, SK05b]. Colony [RL02]. color [Ebn04]. Coloring [LSH96, BGM^{+}08, DJT03, GDP08, HLM^{+}90, KJD03]. Colorings [GJP96, Ros89]. colouring [SS03]. column [Mat06]. COMA [CKL99]. Combinations [Kap93]. Combinatorial [Kap93, KA89, CCLS94]. Combine [BLPV95, Van94]. Combined [OY00, CF88]. Combining [LKK94, LK98, LC96, SZ00a, WR95, GWWL94, HDJ08, TY90a]. Comments [Cha94, GRV08, Pan09]. Commercial [DZDZ01, MKC01, NKC^{+}97]. commit [mYA91]. Committee [Ano93a]. Commodity [PVPM06, MC03, ZB09]. Common [MS99b, ALH^{+}09, MS88, FII04]. common-bus [MS88]. communicating [BFTV87, SSM^{+}06]. Communication [BPR99, BKT95, BCR96, CW00, CCRS92, CGL^{+}95, CS95c, DUSH94, DS95b, ESMG96, Fah96, FM99a, FKT96, FGKT97, FA95, FAM96, Fra92, GRV97, GBES93, GM94a, GK98, GPS96, HQPT99, HH01, HP95, HS93, HA92, IM94, ITT04, Joh87, KL01b, KLS90, KS00, KS02, LHS97, LZ02, LR03a, LO96, LWP02, Mck94, MRRV98, MSST99, PP96, PB99, QH96, RWK95, RS92c, RU99, RMC97, SCM99, SS99, SOG94, SSK96, SBAM96, SKH96, TF92, TSHH01, TSC01, VM03, WR97, XKMN94, Xue97, ZH99, BFTV87, BCM87, BOS^{+}91, BRP03, CCS06, CNS03, CHC05, DW04, Ede91, FW05, GPT06a, GP05, HK05, IB04, KYL05, KSG03, Lai86, Lo92, Lun90, LM09, MAM05, NRM^{+}09, PB90, SS89, SPBR91, Sta04, SW90, ZBF05, ZV09b]. communication-aware [ZV09b]. Communication-Computation [QH96]. Communication-Efficient [HQPT99]. Communication-Free [HS93]. communication-induced [LM09]. Communication-Minimal [Xue97]. Communications [AMN00, BD00, CQ95, DRR96, LLJ00a, SC91a, SHC93, TSC01, WA02, YMG01, ZR00, EB09, GMH^{+}91, LHP07]. Communicator [KF90b]. community [Trä09]. Compact [CDF01, CJ99a, CJY04, CI03, NCTT09]. Compact-Port [CDF01]. Compaction [BHR91, Kar95, WD94]. Comparative [AAD02, GS00, QM01, HA91, PL03b]. Comparing [GGW96, YL98]. Comparison [BSB^{+}01, DRSB01, Fre96, GY92, JNW96, KA08, KA99, OP98, SSOB02, SAC^{+}98, Tay02, AFM03, FGZ03, SS94b]. ComPaSS [XKMN94]. compatible [MP08]. compensation [Yan09]. Competition [eW95, TR89]. Competition-Based [eW95, TR89]. Competitive [DLLX97, GS96, Ser97, VM95]. Competitive-Update [GS96]. Compilation [BCR96, CA96, HHKT96, PA96]. Compile [Fah96, HA92, LPU97, PM96]. Compile-Time [Fah96, HA92, LPU97, PM96]. compiled [KYL05]. Compiler [ABDS02, BW95a, CGSV93, HKT94, KRC00, LY98, LY01, RJY96, SDS99, SD00, Tse90, VV90, WB94, DK04, RG06, Sab94]. Compiler-Controlled [SDS99]. Compiler-Directed [LY98, LY01, RJY96]. Compiler-Optimized [ABDS02]. Compiling [BS90, BCF^{+}94, DRR96, GKHS96, KHS96, SSHC00, SB93, DeG88, LC91a]. Complement [YAS98]. Complete [BP02, Efe96, HKMU98, HM01, SP96, SHL95, TT98, Wag94, ZW00, MPZ09]. completion [KSG03]. Complex [GPS96]. Complexity [BH93, CMS92, Dja06, FAGW95, Fra92, GRV97, Gon98, JBL02, Tay02, BPW05, CH06a, DUW86, SSS88, THSS87, WG08]. compliance [AM06]. Component [HHM94, SR94, CT94]. Components [BJ96, Kar02, BBB^{+}06, Hoh90, LWR^{+}03, MHPR05]. Composed [SM92a]. Composing [BA96]. compositing [WGCZ09]. Composition [HLJ98, Tay02]. Comprehensive [DG94]. compressed [WBTM09]. Compression [SYO94, CD95, NRM^{+}09, SR91]. Comput [WTC08a]. Computation [AM97a, AISS97, BCV94, BP95, BA01b, CA95a, GM94a, GM95, HR92b, HR92a, JSS92, KF95a, KS00, LHM95, PB99, QH96, Sch90, Sin87, SA93, TR96, Win85, CR96, CL85, DHK04, DWHL87, JT88, KSG03, Lee90, NCTT09, PK07, SD88a, SZ03, VGAB08, WL04, WT09, WCO^{+}09, YJB91]. Computation-Intensive [CA95a]. Computational [DRC90, JBL02, KRW96, KR97, Num08, Num09, AB03b, AGMJ06, CS06a, DHS06, LBE03, MJ03, TZ06, WW03]. Computations [AGF94, AMN00, AP94, Ano92a, BR95a, BDKM94, BW95a, Cas93, CN93, CQ95, CGA98, DUSH94, DN94, GR96, GK98, HH97, HJ01, HF02, KL01a, KME92, KC99a, KS02, LPZ99, Man94, MR94a, MP93, MNM98, NRS95, Nas94, Nic94, OS96b, OSZ98, OP98, SV00, WB96, ZB97, ZYO02, AAD05, AFM03, DMCFCM03, EL91, FXW03, Joh87, KME89, KHK03, ST89, SC04, SK91, SS94b, TG04]. computations/applications [KHK03]. Compute [ABM^{+}92, CM92, CTZ99]. Compute-Intensive [ABM^{+}92, KAS07]. Computer [BCH95a, BS96b, BS96c, Cha94, CDP95, HHM94, IWM97, Kri91, LLS93, LR94, MKY^{+}97, NSS97, PEC95, VV90, WF93, WHT02, Gai87, GE85, Gos90, GREC91, HR89, HR90, Irw88, JW89, KK86, LV88, MP08, SAB^{+}92, Vel89, WJD91]. Computers [Alu97, ADM^{+}94, AB93, BS90, BR95c, yCM98, CCC92, Chi92, CY96, CJ99b, Fer93, KL01a, KGV94, Li01, MT96, MSC96, MYD95, Moh96, NFEG97, NS92, PE93, Ree84, RW01, SR94, Shu95, Sto90, Tan84, TC92, VSM96, WLR90, Yan93, YP96, Zhu92, ZM94a, ALS91, AP91c, BGM^{+}08, BCF^{+}94, Car90, CT94, GMS06, JL05, KESA07, LR06, ML89, PB90, Raj04, Sab94, Sch87, WRHR91, ZLRP91]. Computing [AW95, AL99, AM97b, ANT02, Ano97k, Ano99g, Ano01e, Bai94, BG90b, Bir94, BD00, BSB^{+}01, BDH^{+}97, BNSP99, BS09, CA94, CEF^{+}95, CDJL09, CP99, Deh90, DAYA02, DBP94, ELS94, ES97, FFK97, FPP^{+}08, FGKT97, GRS97, GS01a, HGCC96, HS00, HHC98, KSA95, KMKD97, Kri92, KC99b, LAS^{+}97, LFA96, LS01, MWL00, MAS^{+}99, MC93, NA06, OY00, PN97a, PN97b, Pat01, PT01, PRS97, SM94, SdS97, SR95, SS97, Szy95, VR94, WR97, WSRM97, Wei98, WF96, WLID02, wXH00, YZS96, ZO97, AC89, Arb89, AM06, BC06, BW09, BDDL09, Bou03, BH05, BSMH08, BAK^{+}03, CCS06, CSW08, CVJ09, DK08, DÖ06, EL88, FCG04, GWWL94, Han89, IB04, KS08, KVHS07, KL05, KBD05, KC04, KMS^{+}06, LTL06, LCC^{+}05, Li05, LY08, LPX05b, LR05, Luk85, LLS07, ME04, MCT06, MMS09]. computing [MSJ05, MC03, NAK04, NRM^{+}09, Oza04, Raj04, RRS^{+}08, SAOKZ05a, SAOKZ05b, SCS^{+}08, SAB^{+}92, SFEF06, SB04, ST08a, TZ07, TRSS06, UAKI06, VD04, WS06, XQ04, YC04, ZV09b, ZB03, ZFWF06, ZHO03, Ano99g, Ano97j, BS09, CDJL09, Gra09, Lan09]. Concentrate [LW95]. Concentration [JL05]. Concepts [TAS^{+}01, ZZ90]. Concerning [IPK85]. Concurrency [Ahu90, KCV99, LZCY09, MS96, NMS93, RM90]. Concurrent [AyJ93, CCM92, FPD93, IM94, Joh94, MM04, RSD94, RS92d, WCF94, WW96, WG93, WT92, Chi95, CMT92, DB08, FJSW90, GV86, KME89, Par89, ST05, TK07, Chi95]. Condition [SJ96]. Conditional [CW09, ERA95, RLS96]. Conditions [DJ98, HM96, MI92]. Condor [HS97]. Condors [BZH06]. configuration [BL05, FVCL05, NP09]. Configurations [LK94]. configured [ZV06]. Conflict [BP02, CH92, DP00, DFP06a, HV09]. Conflict-Free [BP02, CH92, DP00, DFP06a, HV09]. Conformance [CY95]. Congestion [BDF01, XWC^{+}08]. Conjugate [Bas97, McA89]. Connected [Ann94, ADM^{+}94, BJ96, BCH95b, yCM98, CCC92, CWW^{+}95, CT94, CY96, CDP95, DVZ96, Fer93, HHM94, LH92, MD01, Moh96, SR94, Tze93, Zhu92, ZYO02, dBL95, BB85b, BBd90, Car90, DW06, GP07, HJ07, HSW04, HR89, HR90, JT88, JL05, KT91, KF90a, LC90a, LC91b, Li06b, LV88, MHPR05, PB90, Raj04, SI86, ST06, SSM89, SC91a, TR08, YME06, HWW96]. Connecting [FT94]. Connection [AyJ93, GHKS98, ML89, TT07, YSL08, CM93, CRFS94, EHS94, LAD^{+}96, LTD^{+}93, Sab94]. connection-based [TT07]. connection-level [YSL08]. Connectionist [MBK^{+}92, TR89]. Connections [Goe94, TC03]. Connectivity [Wil92, ASM09, DH91a, OMSGNSG05, SK89a]. Conquer [CTZ99, AY89, BW09, Sto87]. conscious [OC07]. consensus [ISM07, MR09, WTC08a, WTC08b, WCYR08, XBK07, DS04b]. Conservation [FLS^{+}97]. Conservative [LA93, BD04]. Considerations [Ger98, VWHL96]. Consistency [Bir94, CA95b, GAG^{+}92, SS08, Fei03, HC09, WDDK09, XO05]. Consistency-driven [SS08]. Consistent [KCDZ95, HK08, JLM08, LFA05]. constancy [Ebn04]. Constant [BGOS95, BPP05, BTZ98, COS^{+}95, DS01, KBG92, RO92, TVS97]. Constant-Time [BGOS95, COS^{+}95, DS01]. Constrained [AZ01, BSDE96, MMVR97, RL95, BKS05, HP06, KSI04, WTB^{+}08, ZV09b]. Constraint [GHH92, LP97, Mon94, CLL09, UAPM07]. Constraints [BA96, KB96b, LTWY95, van96, AP91a, AY89, ACU08, DUW86, FVLB09, Li06b, SSM^{+}07]. Construct [BW96]. Constructing [CCS06, CS06a, Hal05, HS94b, BBL04, DW06, GC07, LMZ04, LH04, OMSGNSG05, WC91, YZLT09]. Construction [BCH95b, DM95, DFN^{+}94, DJM94, BFG^{+}03, LT07, LS05]. Constructions [FA95, HV95, HV09]. constructor [tH90]. Constructs [Ano92a, KME92]. consumption [LCW05]. Contact [PAH^{+}98]. contemporary [VM03]. Content [Li99, Win85, Bar05, Fei03, FM07, SZ09, SCK03]. Content-Addressable [Win85]. Contention [BCD00, LKK94, FA07, JW89, NSTN91, Nik03]. Context [Cou93, Ano04d, BPA06, IB04, YK04]. context-aware [BPA06]. context-sensitive [Ano04d, YK04]. Continuous [NH93, MCdS^{+}06, dGP06]. continuously [AKSM08]. Continuum [MP96]. Contractions [BBN93, Ros89]. contributions [RGU08]. Control [AGW98, AGW01, BJP91, BBM^{+}02, BCLR96, BCD00, BDF01, DSST95, ESA03, FR96a, FT94, KSP^{+}92, LM96, MS96, Nie94, OS93, SG96, THBF97, WLID02, Ahu90, BMF05, CF88, Che89, CLM90, FL86, HCZ04, KNS91, LL90, LZCY09, LCW05, MG09, RKK06, TG04, WJD91, XYDL06, XWC^{+}08]. Control-Memory [BCLR96]. Controlled [CGSV93, Li99, MG91, SDS99, SD00]. controls [YSL08]. convection [CEGS07]. convergecast [KK06]. Convergence [GCM95, ÜD96, YBOY97, Tor89]. conversion [SMH91]. Convex [DS84, DFRCU99, LP97, Wu02, DDNS06, GS03a, RBD08]. Convexity [BOS^{+}95, BGOS95]. convolver [Kep03]. Cooperative [BW95b, GRDB05, NP09, YF07]. Coordinated [VPHML06]. Coordinating [DZ97, CHC05]. Coordination [DRST02, DRT07, MS05]. Coprocessors [SS99]. Copy [Ano93e, CS93b, CS92]. CORBA [CCC^{+}04, LWR^{+}03, MSAF04, RSR04, wXH00]. CORDIC [CL88, HBH93]. Core [BCR96, PL94, KKB^{+}06, LHT08, Raj04, SCB09, Trä09, WJV07]. core-based [LHT08]. Cored [GS01b]. Cored-Based [GS01b]. Correct [JF95]. Correcting [BA01b]. Correction [Lat98]. Correctness [BCC95, GG94, KS94]. corrector [GGR89]. corresponding [BS03]. Corrigendum [WTC08a]. Coscheduled [KCD08]. Coscheduling [ABM^{+}92, NBSD99]. Coset [Oru87]. cosmology [LTL06]. Cost [AZ01, Ano92c, BC01, DT97, FM99a, GPS96, HCS^{+}00, JH92a, JLRA97, KER01, LO96, Nic07, PP96, QM01, SC95, WC91, Wei02, AP91c, BJS03, CL09, GJXZ05, LMZ04, MSM09, SSM^{+}07, Yan09, ZJ06]. Cost-Driven [FM99a]. Cost-Effective [BC01, ZJ06]. Cost-efficient [Nic07]. Cost-Optimal [DT97, WC91]. cost-performance [BJS03]. Costs [Fah96, WF90, PB90]. coterie [SGR03]. Coteries [WRC^{+}02]. Countering [SFEF06]. Counting [KS00, SS96, WW98, WW04]. coupled [AjHcC90, BBB^{+}06, BMF05, SA90]. Coupling [GT02, YWD08]. Cover [Ano04e, ANP07, DDNS06]. Coverability [SP90]. coverage [HWC08, WMW09, ZC04]. coverage-oriented [ZC04]. coverings [Bod89]. Covers [ABCP96]. Covert [BKT95]. Cowichan [ASST05]. crash [BG05, ISM07, MFVP08, MR09]. crash-prone [MFVP08]. crash-recovery [BG05]. Cray [CDH84, SI91]. Cray-2 [SI91]. CRCW [GM94b]. create [AM07, MMAL^{+}06]. Creating [DHS06]. Creation [Ric98]. CREL [KMB91]. CREW [OOW95]. Criteria [BSS99, CCR94, LL07]. Criterion [SS93]. Critical [BLG01, Seb95, GST09, TYH09]. cross [CI03, KPR88]. Crossbar [CP01, KJ84, OK01, PD92, LW89, McA89, Wil90]. crossed [CW09]. crossing [HSSM07]. Crosstalk [Qia97]. CSA [Ebe94]. Cube [BCH95b, JH94, MS85, RP98, Tze93, AP91b, JT88, JL05, KF90a, PK04a, ST06, LH05]. Cube-Connected [BCH95b, Tze93, JT88, JL05, KF90a, ST06]. Cubes [HJ90c, HTHH02, JH92b, Lat98, XL95, BVB02, CW09, FLPJ07, SAOKM03, WS97b, YTH07, YD98]. Cubic [CP98, MP88, YME06]. CUDA [CBM^{+}08, KME09]. Cumulative [Ano98l, Ano99h]. Current [KS95]. curve [Gue86]. Customized [Isl97, ZLP97]. Cut [DRSB01, KLLK98]. Cut-Through [DRSB01, KLLK98]. Cutsets [DH94]. Cycle [Ano00d, KK95, LS97, Ros99, HDT^{+}05]. cycle-accurate [HDT^{+}05]. Cycle-Stealing [Ano00d, Ros99]. Cycles [BCH95b, Tze93, Wan01a, dBL95, JT88, JL05, KF90a, PK04b, ST06]. Cycletrees [VB96]. Cyclic [OP96, PT97, SSG93, BD05, HS03, PK05a, Sch87, ST87]. cyclic-by-rows [ST87]. Cylindrical [WN94].

D [Ano92a, Ano93e, BAES92, CS93b, SS94b, AR97, BLPV95, BFG94, BAL05, BC94, CW00, CS92, DSAUM99, GW99, HHKT96, HKT94, LME95, MKY^{+}97, PEC95, Wan07, WS95, Wu02, YB01]. D-ISODATA [DSAUM99]. DADO [SM86]. Daemon [KY02]. DAG [CJ99a, CJY04, DQR^{+}09]. Dags [BCLR96]. daisy [GRV08, MVB05]. DARPA [WRHR91]. Data [AOS^{+}05, AL04, AAL95, ALS91, Ano96j, Ano00d, ADM^{+}94, BVB02, BCD95, Bal90, BBB^{+}06, BG90b, BHS^{+}94, BR95c, BR02, BS09, CDY97, CK08, CGL^{+}95, CP92, CHR94, CRFS94, DOP98, DRC90, DSAUM99, DRST02, DHR96, DSD^{+}97, DSS95, Fah96, FMP98, FKKC97, FMW^{+}94, GG94, GP93, GC01, GDN^{+}98, GS96, Gup92, HK01, HJD^{+}01, ISZBM99, JW94, JS86, JB93, KR97, KLS90, KRS01, LSCA93, LZ02, LAS^{+}97, LY98, LY01, LO96, LL95, Lu01, MS85, MRRV98, MK92, MRK93, MNB95, MNM98, NBP98, Nic94, OK02, OP98, PHB96, PH91, PL98, PT97, QZ94, QH96, RSW90, Ros99, RW93, SS89, SMH94, SG99, SR97a, SR97b, SAC^{+}98, SSHC00, SHT^{+}95, SS94a, SSYG97, SIR92, Ste95, SC91b, SV00, SG96, TSC01, TR96, VBM90, WB94, WNA^{+}94, WPKK94, WSS93, Wei02, WS97a]. Data [ZRC99, Ara90, BH86, BR91b, CK06, CF88, CKN07, CLL09, CZ90, CTT08, ESTA94, EDÖ05, FP03, Gao89, GE85, GS91a, GJA08, GBA08, HMV07, HLS03, HSMB91, HP06, HA05, JZ05, KA08, KHK03, KAS07, KL05, LHF91, LC91a, LLW07, Lon04, LA04, MCdS^{+}06, ME04, MP08, NS90, NCT^{+}07, NAK04, NTC03, Pad91, PSPR05, PLR07, Psa96, SS08, SC04, SM08a, SK05a, SD88a, SR91, ST08a, TR89, TBHA07, TZH^{+}06, TK07, VB08, WSH^{+}03, WT09, WL05, RAB08, WLL08]. data- [KAS07]. data-/compute-intensive [KAS07]. data-center [FP03]. Data-Driven [JB93, VBM90, WSS93, BH86, KHK03]. Data-Flow [BG90b, GE85]. data-gathering [LLW07]. Data-Intensive [BS09, SC04, VB08]. Data-Parallel [AAL95, Ano00d, BCD95, BHS^{+}94, CGL^{+}95, DSD^{+}97, FKKC97, KR97, OP98, QZ94, QH96, Ros99, RW93, SAC^{+}98, SSHC00, Ste95, WB94, WNA^{+}94]. Data-stream-based [CK08]. Database [DSW94, HlLLY95, HTL99, LLS93, LHM95, MB93, RSD94, YMR93, BH86, CI86, HPSM91, LY91, LZCY09]. Databases [BM95, CS95b, FCF00, MFS93, Ahu90, BA06, CG86, PF08, Ram89]. datacenter [MG09]. Dataflow [BG86, BCF97, BPN90, BJP91, BH93, GGB93, Gao93, HCAA93, LB90, MNB95, NBM93, RSBN01, SA93, SBKB90, VV90, YMR93, Bic90, KLL87]. Dataflow-Based [RSBN01]. dataraces [SSS07]. datasets [ZB09]. DAWGS [CM92]. dBBlue [SLWW05]. DCC [BCD00]. DCT [Jia99]. DDE [WS97b]. DDoS [CH06b]. Deadline [LTWY95]. deadlines [BSMH08, KSS^{+}07, WL05]. Deadlock [Ano96l, BHRS95, CP01, CMS92, KS94, Li92, MJ94, PA97, PA01, SJ96, TT07, ZN01, BB85a]. Deadlock-Free [CMS92, Li92, PA97, PA01, SJ96, ZN01, TT07]. Deadlocks [RP95, WP02, LJ05]. Dealing [BKS05, FP03]. DEAR [ALF03]. debug [BBCLL04]. Debugger [MB96b, BBCLL04]. Debugging [MI92, MLC^{+}90, SG93]. Decaying [GM96]. Decentralized [GMXA07, HS97, Che89, SL06, mYA91]. Decision [ADS01, BF01, LFA96, KC04, PP06]. Decision-Tree [BF01]. Decomposable [KS08]. Decomposition [Bai94, BBCD02, CP92, HJ90c, HBH93, KBG92, LS95, NPY^{+}97, PE93, QZ94, Ara90, ACFK07, CvdBL^{+}08, Luk85, OT86, SK09, TW87, XWC^{+}08, ZWRI07]. Decompositions [ABCP96, KRW96, Oru87]. decoupled [CTCX08, DBC03]. Decreasing [TSHH01]. dedicated [AM07, MAR05, WLNL06, ZV09b]. defense [XCH08]. definite [KK86]. Degenerate [HF96]. Degradable [BBR94, CGA98, LH92, RCB93]. degradation [NSTN91, WCYR08]. Degree [DS96, Pra93, RL95, LVP08]. Degree-Constrained [RL95]. degrees [ZDC06]. Delaunay [ABC^{+}09a, ABC^{+}09b]. Delay [AZ01, HWWH08, LMZ04, MD07, SGR03]. Delay-Constrained [AZ01]. delay-guaranteed [HWWH08]. delay-optimal [MD07]. Delays [GM94a, GK98, KL01b, Sta04]. Deleting [BCK^{+}09, PPC04]. Delivery [CLZ02, CLV95, Bar05, KMF^{+}05, KNS06, SZ09, WGCZ09, XYDL06]. Delta [KJ84, YL89]. Demand [DSST95, HLL^{+}95, JSCB95, BSW07, FVLB09, SFEF06, WL05, XG03]. dendritic [WCKD06]. denial [KMMZ06]. denial-of-service [KMMZ06]. Dense [DVW94]. densities [DHK04]. Dependability [SM92a, WLID02]. Dependable [MAJJ05]. Dependence [GSG^{+}93, KK95, Xue97, CC87, Psa96]. dependences [NCT^{+}07]. Dependencies [KBG92, TC96, BSMH08]. Dependency [GP94]. dependent [AL04, BH05]. deployment [VHH08, ZC04]. depth [BP89, LH04, PV07]. depth-first [PV07]. deques [ST08b]. derivatives [PK04a]. Descriptor [Bal90]. Design [Ano92c, BAHP01, BCD00, CGKK97, Car95, CCC90, CT93, CAB94, CW93, DBKF90, DVW94, ES96, EMP^{+}96, FC90, FR96a, Fer92, GRV08, GFB^{+}92, Ger98, GRS97, GSP02, HP97b, JH92a, LL90, Lee91, LH92, LLS93, MKC01, MVB05, MG09, MML07, NBM93, NJ91, Nie94, NsPPC02, OS93, PA01, PI90, RCB93, RPS93, RKK97, SAOKZ05a, SAOKZ05b, SRK95, SHC93, SOG94, WNA^{+}94, WH97, XKMN94, BZLI04, Che86, Chi95, CC96, DE91, Fer90, FD86, GREC91, HDT^{+}05, HWWH08, Lon04, LVB07, Nap90, PLD87, RGD03, TM06, TB90, VHH08, WSG91, ZV09b, ZFWF06]. Designing [BC01, CB06, DH91b, GP93, GB93, KT89, NS92, Oru87, SRGB90, TC96, KAS07]. Designs [HCS^{+}00, LHM95, MD01, Oru94, Bhu87, CP04b, Sch89b, WAS88]. desktop [CCEB03]. Detect [XCH08]. Detecting [CK97, NCT^{+}07, Tse95]. Detection [Ano96l, BN02, BHRS95, BST01, CW93, CY95, CDP95, dADB96, GCKM97, GS96, HTB98, ISZBM99, KSB94, KS94, LLLY08, MMRS98, Par92, PAH^{+}98, Ram89, RP95, SL97, WCF94, AMK^{+}07, BXA08, CRK^{+}09, CV90, CH06b, DFP06b, Eri88, FM85, Gue86, GH89b, KHK03, Lai86, LJ05, MD07, MFVP08, YF07]. Detections [Yen01]. detector [SLG06]. detectors [DGFGK05, LFA05, MFVP08]. detention [JXW06]. Determinacy [BN94]. determination [MJ03]. Determining [GRR93, LAS^{+}97, DH91a].

Deterministic [AS91, BBCD02, OS96a, SGS08]. Development [BR95b, FSD04, PH00, DBC03]. deviation [XBK07]. Device [DM90a, ALF03]. devices [Ano04d, WL04, YK04, ZV09a, ZV09b]. DEVS [PK05c]. DHT [BJPPM^{+}08, SP08, SX08, ZH07]. DHT-based [BJPPM^{+}08, SP08]. DI-multicomputer [CC96]. Diagnosing [Qia97]. Diagnosis [BW95b, Kav93, KF95b, RFM94, Wan01b, eW95, FY86, FZ90, Yan04]. diagonal [PRHB06]. Diagram [RR95b]. diagrams [SZ03]. Diameter [DF95, LP95, RS96b, RLS96, WIKC97, BBL04, CW09, SLWW05]. Diameters [Als01]. DICE [CKL99]. Dictionaries [MD98]. dictionary [GA90]. difference [HT90]. Differences [LDCZ97]. Different [GAG^{+}92, PD92, Bhu87, GPT06b, MM06, She06]. differential [GGR89]. differentiated [AM07]. differentiation [ZI08]. Diffracting [DLS00, HPT07]. Diffusion [SKK97, BFH09, CEGS07, MMS09, RN04]. diffusion-based [MMS09]. diffusion-type [BFH09]. Digit [BOI91]. Digital [ZRC99, NAK04, PR06]. Digitized [HHM94, Ara90]. Digraphs [BMMS01, TZ00, BP89]. Dilated [Iqb92, Qia97]. Dilation [CCCM96]. Dilation- [CCCM96]. Dimension [HSW04, RS96a, WS97b, XL92, XL95]. Dimension-exchange [HSW04]. Dimensional [AKPT99, CCCM96, DFN^{+}94, FLS^{+}97, Hwa97, KR98, LHS97, LP96b, LP95, NEG85, TC96, VB94, YCY^{+}00, AB05, DMCFCM03, Deh90, FCG04, GSSS03, HT90, KVHS07, KLC05, LSC00, LC91b, NBP98, NAK04, PTA08, PK07, SGR03]. DINO [RSW91]. Direct [GV94, LLCC02, TF01, ACFK07, ACU08]. Directed [GY92, LSC00, LY98, LY01, RJY96, BD05, WCWH03, Wu03]. Direction [BC94, Ebe94]. directional [CCHC09]. Directory [GS00, JSM94, RFPAG08]. disasters [FP03]. Disciplines [MSd^{+}95]. disconnected [LR03a]. Discovery [CHGM01, AOS^{+}05, KOA09, KKS09, MKC^{+}09, She09, TDC05]. Discrete [Ano02v, AB93, BBM^{+}02, Bou02, DMSH90, Lin93b, Lin93c, LLCL98, NC97, Pra93, CVJ09, CRC^{+}02, TKHG04, ZZ90, ZCK^{+}02]. Discrete-Event [DMSH90, Pra93]. Discrete-Time [BBM^{+}02]. Disjoint [BGR96, GT97, GP00, NS90, RSS99, WB01, Lin03, LS03, TDM05]. Disk [CT93, Cor93, ER97, GP93, LP96b, MKC01, MRK93, MFS93, Raj01, RCB93, CL03b]. Diskless [PKD97]. Disks [MT93b, MB93, MFS96, CkLCK04, CkLCK05, OC07, VA07]. Dispersing [Gil94]. displays [Tay05]. dissemination [MCdS^{+}06]. Distance [BVB02, CW00, CDF01, DS01, DF95, ST02, DS04a, Eİ07, Hsi04, MBR08, ST06, WCWH03]. Distance-Hereditary [CDF01, Hsi04]. Distance-Insensitive [ST02, ST06]. DistDLB [LTL06]. DistOpt [CLRW00]. Distrib [WTC08a]. Distribute [LW95]. Distribute- [LW95]. Distributed [AE95, AL99, AM97a, AM97b, AMN00, AFS96, ABJS01, Alu97, AYI97, Ano96j, Ano96l, Ano97j, Ano99g, Ano02v, Ano02u, ABCP96, BR95a, BR96, BFTV87, BGLA03, BCV94, Bas97, BA01a, BCH95a, BAS06, BPR99, Bir94, BCD00, BCR96, Bou02, BSB^{+}01, BHRS95, BNSP99, BS09, CS00, CTD99, CCM01, CC08, CL91a, CS93a, Cha94, Cha96, CKK00, CNS03, CC94, CK97, CDJL09, CB95, CWP98, CM92, CA95b, CLRW00, CJ99b, CP99, DWG03, DY99, DA97, DUSH94, DS95b, DOP98, DMSH90, DN94, DSW94, DSAUM99, DAYA02, DL99, DH95, dADB96, EP90, FR96a, FFK97, FKSW97, FM99b, FY97, FTC00, FBDC99, GDP08, GP07, GCKM97, GM94a, Gra09, Gup92, GKHS96, GHSJ96, HR00, HBCM99, Haw97, HK01, HP97b, HLJ01, JF95, JSM94, JNW96, JRR99, KKGS01, KY02, KRC00, KS97a]. Distributed [KHS96, Kel00, KB96a, KZ96, KCV99, KS00, KC94, KS94, KS02, KC99b, Lan09, LWY97, LTH97, LZ02, LC90b, LHM95, Li99, Li01, Lin93c, LLW07, LHT08, Lon04, Lu01, LS01, MI92, Man97, MS99a, MLC^{+}90, MT97a, Mat93, MSS00, MFS96, MSST99, MK08b, NSS97, NTA96, NBP98, NM02, OK01, PHB96, PAM94, PA96, PB99, PK07, PM92, RSB96, RWK95, RS92c, RDS02, RJY96, RGS00, RAS96, Ros07, RP95, SM94, Sch89a, Seb95, SRGB90, SZW05, Shu95, Sin87, Sin93, SS94a, SM08a, Sni03, Soh96, SIR92, SBAM96, The02, TSC01, TAS^{+}01, TG97, TB90, Tse95, TY95, Wan01b, WCWH03, WW98, Wee01, WRC^{+}02, WMG01, WF96, WLID02, WUG99, Wu02, XBK07, wXH00, XQ04, YH97, YB01, ZV06, ZM94b, van96, AT03, ALH^{+}09, AAFV04, AL04, Ahu90, AGMS04, AFM09]. distributed [AMK^{+}07, AH06, BFG^{+}03, BCV05, BMB^{+}08, BLPA05, BG89, BNP02, Bar05, BB03, BRP03, BK08, BD04, BMF05, BH05, BGM^{+}08, BCF^{+}94, BFKP04, BBL04, CSWD03, Car95, CG86, CV90, CvdBL^{+}08, CTCX08, CS08, CLM90, CkLCK04, CkLCK05, CGG^{+}09, CJA09, CI86, CPO^{+}03, CTT08, CK91, Cyb89, DK08, DM04, DRT07, DHK04, DH04, DJT03, EBE08, ESA03, FCC07, Fer90, FL86, FX06, Gai87, GCS06, Gos90, GWWL94, GC05, GL90, HJ90a, Hoh90, HLM^{+}90, HKW05, HL07, ITT04, IB04, IS06, JLM08, JZ05, Joh91, KUA07, KK06, KMMZ06, KAS07, KCD08, KL05, KBD05, KP05, KC04, Lai86, LTL06, LLL06, LVP08, LL90, LJ05, LY91, LZCY09, LVR90, LC91a, LVP07, LB09, LA04, LCM^{+}06, Lun90, LM09, MD07, MM07a, MSM09, MHPR05, MBR08, MS86, MTS90, MM07c, MFVP08, NP09, OFS03, PKN08]. distributed [PK05b, PRHB06, PGS06, PL03a, Pop91, PF04, Ram89, RLH03, RKS87, SDTD04, SSS88, SU87, SC04, She09, SCS^{+}08, SK90, SXZ06, SKK91, SK89b, SM04, TG04, TBZB05, TZH^{+}06, VB08, WW07, WTC08a, WTC08b, WW04, WL92, XHY07, XQ07, ZCK^{+}02, ZV09a, ZWRI07, ZWL03, dG91]. Distributed-Memory [AMN00, CB95, CJ99b, DY99, Gup92, GKHS96, GHSJ96, KRC00, KHS96, NSS97, PHB96, RGS00, Soh96, BGM^{+}08, CPO^{+}03, GL90, ITT04, LC91a, Pop91]. distributed-Web [KCD08]. distributing [TY90a]. Distribution [BRR01, BR02, CLZ00, DHR96, KL01a, LAS^{+}97, LL98, MNM98, SSYG97, ASM09, Fei03, FM07, GRV08, GBA08, HSW04, LLL06, LT07, MVB05, PV89, SS06, ZWL03]. distributions [Nic07]. Distributively [VR94, FPP^{+}08]. divergence [Tor89]. Divide [AY89, CTZ99, BW09, Sto87]. divide-and-conquer [BW09, Sto87]. Divisible [VB02, CVJ09, DW04, MVB05, ZV06]. Division [HP00, QMCL94, ZLPP01, EL91]. DMON [HP97a]. DNA [GPX08, JV09]. DO [CC87, CCC90]. Doan [Ano92c]. Document [ZWL03, XCZL03].

Documents [ALL99, Fei03]. dollar [SSM^{+}07]. Domain [NPY^{+}97, SK09]. Domains [DR95, BMF05, dGP06]. dominance [EE05]. Dominating [RDL95, DW06, HJ07, WCWH03]. domination [GP07]. Don’t [BL94]. DOOR [Won99]. DOOR/MM [Won99]. Doubly [OOW95, ST08b]. down [Sch89b]. Draw [Mil93]. Drawing [CP98]. Driven [CB99, CP99, FM99a, JB93, The02, TVO92, VBM90, WSS93, BH86, GK04, KHK03, QJ05, SS08, VO89, YCC05]. DSDV [BDF01]. DSM [BJS03, ISZBM99, NPP^{+}02, Nik03]. DSMs [KG04]. DSP [QSL^{+}08]. DSS [FGP05, MKC01]. DTN [VV90]. DTNs [Yan09]. Dual [XWC^{+}08, ZW00, MAJJ05, WCC02, WL05]. dual-Hamiltonian-path-based [WCC02]. Duane [BS96c]. due [BKS91]. Duplex [RS94]. Duplication [BA97, DA97, BKS05, BD05]. duplications [SCJ^{+}08]. during [VWHL96]. Dyn [WLNL06]. Dyn-MPI [WLNL06]. Dynamic [AGF94, ALL99, Ano97j, BR95a, BJPPM^{+}08, BPN90, BR02, CJ99a, Cyb89, DL01, FCC07, Fer95, FMP98, GP94, HM01, HC97, KKGS01, KPC96, KC99a, KS97a, LHKL03, LPS^{+}98, LL98, MAS^{+}99, MSd^{+}95, MSSE02, Moh97, MNM98, NPP^{+}02, NPY^{+}97, PHB96, QMCL94, RDS02, Ric98, RGVB00, RN04, San95, SZ00a, SLP^{+}98, SSB98, SB97, SG96, WCE97, WJD91, WLID02, XL92, XH93, ZLP97, ZA05, ZM94b, Ano04d, BCV05, BGLA03, BNP02, BB03, BK08, CBD^{+}09, CW05, CGG^{+}09, CDCD05, EE05, Fei03, FXW03, FKLB08, GCS06, GBA08, IC05, KMS07, KMS^{+}06, LTB02, LLLY08, LC91b, LPX05a, LS06, MC91, MK08a, Mit07, MML07, NCT^{+}07, PKN08, PM05, PSPR05, QJ05, SS06, SSS07, SZD07, SCK03, SLG06, TZ07, TH08, TT07, YK04]. Dynamically [JB98, KSS^{+}07, SB84, dSR00, Kep03, Lai86, Mat06]. Dynamics [ES96, JBL02, NPY^{+}97, PAH^{+}98, TSA97, AGMJ06, CvdBL^{+}08, GBMZ07, LYL08].

E-ODMRP [OPG08]. E-R [BG90a]. earthquake [KME09]. EB [SM92b]. EB-Equivalence [SM92b]. ECC [CL09, GCS06]. ECC-based [CL09]. EcliPSe [RS92d]. EDAs [MMAL^{+}06, dGP06]. eddy [SM04]. Edge [BGR96, BS97, GT97, LSH96, TDM05, WB01, CL85, DJT03, GDP08, Lin03, SS03]. Edge-Coloring [LSH96, GDP08]. Edge-Disjoint [BGR96, WB01, TDM05, Lin03]. Edges [HHC98, FPP^{+}08]. editing [RS90b]. editor [WW03, AB03b, Ano01l, Ano02g, Cas93, Che92, Cho93, Her92, Kri92, Lin93b, Pan09, Sch90, Sto90]. Editorial [Ano94e, Ano95k, Ano96k, Ano99i, Ano02e, Ano02f, GHS94, GHS95, GHS96, GHS97, Kai92, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q]. editors [XO05, AP93, AL99, Ano01j, Ano01k, Ano02h, Ano02i, BG90b, BD00, DOP98, ES97, GGB93, GC95, JW94, MC93, NT90, OW01, PN97a, PN97b, PA96, SH92a, TY95, WC05]. Effect [ACD^{+}93, IS06, BL05, JZ05]. Effective [Ano97k, BC01, GM96, HH97, LT96, MAR05, QM01, TC92, VH93, WLID02, YZS96, DK04, FII04, NAK04, ZJ06]. Effectiveness [GMM00, HKT^{+}91, KS97a, LKK94, NRS95, TC03]. Effects [AMB95, DZDZ01, KB96b, ÜD96, CK88, HLS03, KG04, SPBR91]. Efficiency [EH01a, GG01, GSWW04, LZSL06, SI86, YF09]. Efficient [AOSM04, AP94, AKP95, AG86, AMK^{+}07, BGH^{+}03, BAGS95, BAH04, BRP03, BJK^{+}96, BDH^{+}97, BMIM07, CM04, CRK^{+}09, CKK00, CCC92, CN93, CS95c, DDNS06, EP90, EL97, FGG08, FBK98, FMR05, GPT06a, Gao93, GR96, GCKM97, GM94b, GRS97, GP00, GKHS96, GNW03, HQPT99, HH01, HSLL04, HHC98, HBH93, HO94, Hwa97, Iqb92, JB93, KPC96, KHS96, KLZ97, KKB^{+}06, KA97, KBG92, LJ05, LY01, MD01, Mat93, MHC95, MS99b, NB93, NT93, NIR86, OS96a, OK01, OP96, Pad91, Par98, PA97, Pra93, RSS99, RSB96, Ric98, RJMC95, San02, SW96, SSHC00, Sin87, TU92, TR96, VB02, VBM90, WRC^{+}02, WHT00, XMN92, YD98, YZLT09, ZB97, Zhu92, ZH07, BR91a, Bic90, CKN07, CGW^{+}03, GT04, GLD06, HS06, HRJ94, Hsi04, Joh87, KL05, KA05, LMZ04, LS91, LR03b, LHP07]. efficient [Lon04, LA06, MD07, Nic07, PVGG06, SX08, SM08b, WJV07, Wan07, WTC08a, WTC08b, WMW09, WGCZ09, ZB03, ZHO03, LM09]. Efficiently [MT95, Coh90, CCM^{+}06, FP03]. effort [Bar05, MAM05]. EGEE [VPHML06]. Eigenvalue [Kau94, LYL08]. eigenvalues [VGAB08, ZB03]. Electing [SK94]. Election [AS96, KB96a, FKK^{+}04, KGN89, Pel90, SS05]. Elections [FM96]. Electric [IWM97]. Electrical [MO97]. electron [FCG04, FGG08]. Electronic [WH97, AA93]. Element [BCV94, CSSY94, KME09]. elementary [FK89]. Elements [GB93, KNS91]. Eleven [BSB^{+}01]. Eliminating [DR98]. Elimination [BPST96, BMM97, CS95b, Cap87, KA91, Vel89]. Elimination-Based [CS95b]. Elliptic [PSE^{+}01, BGH^{+}03]. EM-4 [BAM93]. embed [SKK91]. Embedded [WA02, CkLCK04, CkLCK05, DQR^{+}09, GSWW04, KR06, MBR08, PRHB06]. Embedding [ANS97, Ann94, AM93, BL89, CCCM96, CS95a, Efe91, Efe96, HKMU98, HJ90c, LSC00, LPS^{+}98, Lin03, NPI^{+}96, PM92, QM01, RWY93, SHL95, SLP^{+}98, TT98, TLW94, TL96, Var91, Wag89, Wag93, Wag94, Wan01a, Wu85, WFL98, BG90a, FLPJ07, FT04]. Embeddings [GH93, HM01, HOS94, KC98, MT93a, OS97, OD95a, CL91a, GNW03, YTH07]. Emerging [Ano02v]. Empirical [FTC00, LR93, XZS96]. Employing [AGMJ06]. empty [Deh90]. Emulation [JH94, PRW94]. Emulations [RGD03]. Enabled [MWL00, CCN06, KTF03]. Enabling [SP08, ZPI06]. Encoded [JH94, CLV95]. Encoding [AAL95]. End [Ano08, Ano09, FGP05, GBMZ07, HPSM91]. end-systems [GBMZ07]. endpoint [Hsi04]. Energy [ALF03, DKY01, LR03b, PLR07, QSL^{+}08, FKLB08, GTN^{+}06, HP06, KSI04, LCW05, WMW09]. energy-constrained [KSI04]. energy-efficient [WMW09]. ENF [CK97]. Enforcing [KMF^{+}05]. Engine [KSL85, Ram92, SD88b]. Engineering [LWR^{+}03, Gai87, PRHB06]. Engines [SD00]. Enhance [WLID02]. Enhanced [BOSW94, OPG08, OS96b, OSZ98]. Enhancement [KJ84, TC92, DK04, RH05, RM90]. Enhancing [AYIE98, CRA^{+}08, CCHC09, JBY^{+}05, VA03, WXZ05]. Ensuring [JF95]. enterprise [BJPPM^{+}08, CCEB03]. entities [Ahu90]. Entropia [CCEB03]. Entropy [TVO92, VO89, WMW09]. Entropy-Driven [TVO92]. enumeration [SSTP09, SR90]. envelope [GC07]. Envelopes [BMRC98]. Environment [AT94, AD95, ALL99, AA95, BB93, CP97, CLZ02, CCRS92, CHR94, CB96, DKY01, DRSB01, KZ96, KC99b, LC90b, LAS^{+}97, Li99, MHF93, RS92b, RSD94, SG93, SRGB90, SS00, WH97, ZL93, AOS^{+}05, CK88, CCS06, KVHS07, KSS^{+}07, LLLY08, MAR05, MML07, SSM^{+}06]. Environments [CTD99, CLRW00, CP99, KRW96, KR97, KER01, LTH97, PRS97, PRG88, SSK96, WSRM97, WSA^{+}94, ATZ07, BAL05, BPA06, BH05, BSMH08, CLL09, DBC03, JS86, KAS07, LY91, LWR^{+}03, MK08a, NP09, PP06, TG03, YT05, YCC05]. epistatic [HLS03]. epochs [PBS08]. EPPOD [WH97]. EPSILON [GH90]. EPSILON-2 [GH90]. equal [ST85]. Equation [DM90a, RW01, Gao86, LYL08]. Equations [IK94, MV94, PSE^{+}01, QOvdG01, TH02, CM03, GGR89, GS91b]. Equivalence [OO85, CM04, SM92b]. Ercegovac [Ano92a]. EREW [DL98, HS94a, ZK94]. Erratum [Ano92c, Ano93e, Ano96l, Ano00d, BS96c]. Error [Lat98, Par92, WCF94, PKN08, RIZ90]. Error-Correction [Lat98]. Essay [Mil93]. Essential [DSS95]. Estimating [CCK88, MK92]. Estimation [CP92, Fah96, PKN08, SPVvH03]. Ethernet [HcF05, KYL05, PYF08]. Euclidean [DS01, DS04a]. Eulerian [Kal04]. EUROGRID [LBE03]. European [LBE03]. Evaluating [BL96, BC01, CLRW00, FW05, HCS^{+}00, HKT94, LR94, RS92b, SS99, TTG95, ZYH94]. Evaluation [ATM01, BPJG92, BS92, BCD00, BM95, CT93, CEF^{+}95, CP01, CP04b, CP91, CP92, DT01, FR96a, FTC00, GGD93, GS96, GS00, HJ90b, HN91, yHY97, JB93, KCDZ95, LLS93, LYL93, LP96b, MT95, MS85, MKC01, MB92, MJ01, NBP98, PEC95, PTC^{+}93, RCB93, RNSB96, RKK97, SM92a, SDS99, SOG94, THBF97, TH02, VBM90, Bat05, CkLCK04, CkLCK05, CC96, DM88, GRV08, GE85, GS91a, HW03, LL90, MS88, MVB05, Sch89b, SWP90, WL90, XQ07, XWC^{+}08]. evaluator [MS87, MP88]. Even [NT93]. Event [Ano02v, AB93, Bou02, CK97, DMSH90, Lin93b, Lin93c, Pra93, BXA08, CK08, LVR90, Tay05, ZZ90, ZCK^{+}02]. Events [Yen01]. Eventually [LFA05]. everything [CCM^{+}06]. everything-shared [CCM^{+}06]. Evolution [JM00, Ngo06]. Evolutionary [Ano99g, MSSE02, SdS97, SS97, ZO97, AC89, BH05, GB06, SCS^{+}08]. Evolving [GR96, OH02]. Exact [RS96b, OFS03, Psa96]. examination [FL86, SMH91]. examples [FK89]. Exchange [VB94, WS97b, XL92, XL95, Dim04, HSW04]. Exchanging [GPT06b]. Exclusion [AE95, Cha94, Cha96, FTC00, GBG93, KY02, KUFM02, NTA96, NM02, Sin93, YZY96, AK07, BAS06, CW05, CH06a, CB06, DGFGK05, Gos90, MM07c]. executed [SP90]. executing [AKSM08, CDJ^{+}89, QJ05]. Execution [CCC90, Cou93, DD95, Gup92, GKHS96, HS86, LAS^{+}97, LTIK05, Mah95, MM93, Mer96, Mir91, NBM93, NS97, NDZA99, OKB95, RSD94, RHH96, RSBN01, SCMB90, SA93, Sun02, WB96, ARM^{+}05, Bic90, CC87, DeG88, DKRI09, FCC07, GK04, MSM09, RG06, SS06]. Executions [LMCF90]. expandable [SSB91]. Expanding [Zia92]. expansion [SL89]. Expected [Ros99, CLL09, SSS88, SC91a]. expected-time [CLL09]. Experience [SH92b, Chi95]. Experiences [ARM^{+}05, CDH84]. experiment [PF04]. Experimental [BJ96, BFG04, FCS91, Hag97, HBJ98, MJ01, PTC^{+}93, YMR93, ZYH94, Bad04, CT94]. Experimenting [AD95]. Experiments [RS92d, CF88]. Expert [DSW94]. Explicit [CP90, DS02, Fre96]. exploit [ZPI06]. exploitation [PVGG06]. Exploiting [CKK00, DL99, FKLB08, FY97, HT90, JBY^{+}05, MNB95, NMS93, SH92b, GJXZ05]. exploits [GBMZ07]. exploration [TD07]. Exploring [LR93]. expression [GS91a, WSH^{+}03]. Expressions [GKHS96, Mer96, DeG88, DM90b, JK89, MP88]. Extended [BLG01, LWOG02, Ree84, Eİ07]. Extending [BBCLL04]. Extensibility [MB96b, LFH^{+}03]. extensible [ZWL03]. extensions [DPSD08, Oza04, JM00]. external [DO89, JZK04]. Extra [SZ00b]. Extraction [YB01, HP06, Pla08, Raj08, WJV07]. Extrema [AFS96, RKS87].

Factor [GG01]. Factored [BSGM90]. factorization [MVV91, She06, ZLRP91]. Factors [BP98, EL88]. Faddeeva [CF98]. failed [Trä09]. Failure [FCF00, DGFGK05, HK05, LFA05, MFVP08, YF07]. Failures [ADS01, DT02, VR94, VR95, GPT06a, HRC09, MR09, RLH03]. Fair [ALH^{+}09, KY02, GNT04, KS03, KDH08, XWC^{+}08]. False [HF96, KG04]. family [NS90, ZDC06]. farm [TBZB05]. Fast [ABCP96, BC06, BF97, CK06, Cor93, DP00, DS04a, DPRW85, EM89, FZC^{+}05, FR96b, GM94b, Gil94, GSC96, GZ97, GJXZ05, HN91, IK94, JNW96, KK06, Lat98, LH09, PH91, PA04, PT97, RHH96, SS03, San98, SR94, SHT^{+}95, SGS08, SA08, SDG08, ST05, TF01, YZY96, YD98, YB01, BC05, Cal06, Kep03, KA91, KP05, LLS07, ST85, TS91, Yan04, LLCL98]. Faster [BMM97, GS03a, LS05, CM03]. fat [CI03, CS06b, SK05b]. fat-stack [CS06b]. fat-tree [SK05b]. Fault [AE95, AM97a, AM95, BXA08, BSS97, BMM97, BW95b, BPA06, BCH95b, CRV94, CL93, CKN07, CY95, CC94, CDR09b, CF98, FY86, FM99b, GNS09, GRR93, HGCC96, HTHH02, KP00, Lan94, LBT94, LGG08, LC96, MD01, MMRS98, Pak89, PB95, Piu01, PKD97, PM92, RLS96, SCC92, SS95, UR94, VR95, WIKC97, WW97, Wu94, XCS06, mYyF92, YBOY97, mYA91, ZYO02, AOSM05, BB87, BDDL09, BPP05, CL91a, CW09, CWL^{+}07, CDR09a, CMT92, CMS04, DH91b, EBE08, FLPJ07, FZ90, LCC^{+}05, LH05, LP88, PR06, PL06, ZV09b, ZJ06]. Fault-Detection [CY95]. Fault-Induced [WIKC97]. Fault-Sensitive [VR95]. Fault-Tolerant [AE95, AM97a, AM95, BW95b, BCH95b, CRV94, CL93, CC94, FM99b, HGCC96, HTHH02, KP00, Lan94, LBT94, LC96, MD01, PB95, PKD97, SCC92, WIKC97, Wu94, YBOY97, ZYO02, BPA06, CKN07, GNS09, XCS06, mYA91, AOSM05, CL91a, CMT92, CMS04, DH91b, FLPJ07, PR06, PL06, ZV09b, ZJ06]. Faults [LT96, WFL98, ISM07]. Faulty [GP97, HIKM94, NSLK99, Pel95, RS96a, Tse95, TL96, Wan01a, Wu02, YTR94, oPP00, Che05, DD96, PK04b, SKK91, YTH07]. FDM [ORR03]. FDM/FEM [ORR03]. feature [PFJ04]. federate [CTCX08]. FEM [ORR03]. fetch [AK07]. fetch-and- [AK07]. FFT [ABZ95, HR92a, JMS86, RKK97, Tay87]. FFTs [BH93]. Fibonacci [Alu97]. Field [BA92]. fields [CDR90, Eİ07]. FIFO [BCLR96]. File [FPD93, GL92, KE93, MS96, WDDK09, WMG01, HOE^{+}09, KUA07, LCM^{+}06, SC04, SZ09, Wan06, ZJ06]. file-sharing [KUA07]. Files [BNS00, JSM94, Lin93a, WRC^{+}02, BCK^{+}09, Che89]. Filling [BFG94]. Filter [LWOG02]. Filtering [BTG02, CH06b, Kep03, PVG09, ZCK^{+}02]. find [Hoh90]. Finding [AFS96, BS97, BE95, CCC92, DH94, DWHL87, FTL92, HHC98, KRSZ02, Kar02, MT97a, MHPR05, OMSGNSG05, PGS06, SH92b, RKS87, WCWH03]. Fine [CLZ00, FR92, IBP08, LFA96, NS97, PY96, SA93, WD94, FW05, FSD04, GVA^{+}08, IKS87, PL03b, TKHG04, LM09]. Fine-Grain [FR92, LFA96, FW05, PL03b, TKHG04]. Fine-Grained [PY96, WD94, IBP08, FSD04, GVA^{+}08, IKS87]. Finite [BCV94, CSSY94, HB97, HNM02, WLD00, CDR90, HM06, HT90, KME09, Sil90]. finite-element [KME09]. Finite-State [HNM02]. FIR [GLD06]. Firehose [KM97]. Firing [KM91, Nie94]. first [Lai86, MP87, PV07, TBZB05]. first-order [MP87]. Fit [SP96, HLS03]. Fitting [CY96, MRRV98]. Fixed [GHKS98, HCWS94, ACU08, BCM06, GREC91, Hsi04, ZDC06]. Fixed-Connection [GHKS98]. fixed-time [GREC91]. Flexible [CCR94, ESMG96, HGCC96, RS92c, VB96, MM07b, PR06]. Flexibly [SA90]. Floating [MRK93, Gro85, MP08]. floating-point [Gro85, MP08]. flock [BZH06]. flooding [XCH08]. Flow [AS95, BJP91, BG90b, ESMG96, LLS93, LM96, MK92, BAMM05, Boż09, CF88, Gao89, GE85, LHF91, MG09, Oza04, TR89, TBZB05, TY90b]. flow-time [TBZB05]. flows [SM89b]. flowtime [LZ05]. fluid [AGMJ06]. flush [CK06]. Flux [Ull84]. FM [LC97]. Folded [Wan01a, SGR03]. folding [LYL08]. FORALL [ALS91]. forces [Num08, Num09]. Forecast [RHH96]. forest [BC06]. Formal [AS00, LSCA93, Eri88]. formalism [PK05c, PSPR05]. Formation [Wu02]. Forms [TR96, WNA^{+}94]. Formulation [JBL02]. Forthcoming [Ano00e, Ano00f, Ano00g, Ano00h, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Ano01-32, Ano02q, Ano02r, Ano02s, Ano02t]. FORTRAN [FC95, AH94, BCF^{+}94, HHKT96, HKT94, HLJ98, Sab94]. Forward [NS95]. Forwarding [GS01b, WTB^{+}08, XYG07]. foundation [DHS06]. four [FZ90]. Fourier [LLCL98, DPRW85, HN91, TS91]. FP [WB94]. FPGAs [MSSE02]. FR [GS01b]. Fractal [LS06]. Fraction [GP97]. fractions [CR91]. fragment [CZZY09]. Framework [AGG98, CLRW00, EMP^{+}96, GHSJ96, KZ96, KK95, LAZC00, Sin95, ZM94b, AK06, BA06, BČFF05, BGM^{+}08, CCC^{+}04, DMB^{+}03, FGM^{+}03, GRDB05, HDT^{+}05, KL05, LS06, MJ03, RGD03, RW02, TZH^{+}06, YT05]. Fraud [BST01]. Free [BP02, CMS92, CG02, CH92, DP00, HPT02, HS93, KM97, Li92, PA97, PA01, RP98, SJ96, SH98, ZN01, CB06, DFP06a, HV09, HA06, SD91, ST05, ST08b, TT07]. Free-Space [KM97, RP98, SH98]. Frequent [AAP01]. Frequently [LL95]. FTN [Seb91]. Full [BBN93, SR88b, SR90, HH97]. full-access [SR88b, SR90]. Fully [BNP02, Fer95, KP00, SJ95, CP04b, DM90b, tH90, SI89, TR08, YME06, LM09]. Function [AGG98, HLJ98, MJ94, SB02]. Function-Composition [HLJ98]. Functional [AB84, Mah95, SC95, QSL^{+}08, YJB91]. Functions [TG97, VR94]. Fundamental [GL92]. Funnels [SZ00a]. Further [PMV06]. Fusing [TVT96]. Fusion [AMB95, STN92, QSL^{+}08]. Future [AE88, KS95]. Fuzzy [BCF97, KC04, NC09]. fuzzy-based [NC09]. fuzzy-decision [KC04].

G2 [KTF03]. Galactica [WL92]. Gallop [Wei98]. Game [ABJS01, BS00, Sch89a, Zep91]. Game-Theoretic [ABJS01]. Game-Tree [BS00, Sch89a]. Games [DKY01]. Gamma [VR86]. Gang [FR92, FR96a]. gap [BJS03]. GAPP [KA91]. Garbage [KS00]. GAs [KZ96]. Gate [OM90]. Gate-Array [OM90]. gather [BM04b]. Gathering [Lat98, LLW07]. Gauss [HO94]. Gaussian [BPST96, BMM97, Cap87, DPRW85, KA91, Vel89]. GCD [Psa96]. GCSPNs [Buc92]. GEL [LTIK05]. gene [WSH^{+}03, FGM^{+}03]. Genehunter [CPO^{+}03]. General [Ano96l, BHRS95, CG02, GFB^{+}92, KL08b, Seb95, VA07, BČFF05, CBM^{+}08, CYZ06, FK89, LB09, OFS03, PK05a, Pel90, RGD03]. General-Purpose [GFB^{+}92, KL08b, CBM^{+}08, RGD03]. Generalization [GCM95]. Generalizations [Oru94]. Generalized [AKPT99, Bai94, BETD94, BR91b, DMCFCM03, Fer93, FAM96, JH92b, Lee94, PE93, SSB91, WIKC97, XL92, XL95, YN92, ZLPP01, FK89, KMP^{+}06, Luk85, Nic88, TDM05, YCC05]. generals [CBV08]. Generating [AMS94, Bec96, CGL^{+}95, CJ07, GHSJ96, SS96, SOG94, TH02, Wri91]. Generation [ASR93, AAP01, AS94, CCM01, DT97, Kap93, KHS96, KBC^{+}01, Lin93a, NC97, RGS00, RNSB96, SSHC00, ABC^{+}09a, ABC^{+}09b, AFM09, Arb89, FK89, Gao89, GMXA07, LC92, Meg91, RKK06, SB04, Trä09]. generator [WSG91]. Generators [Alu97, Bro96, PK89]. Generic [PA01, AK07]. Genetic [ANT02, CGKK97, KRSZ02, KA97, PAJC97, WSRM97, WA02, WLID02, AL04, BČFF05, HSSM07, KM03, LA04]. Genetic-Algorithm [WA02]. Genetic-Algorithm-Based [WSRM97]. genomes [KESA07]. genomic [HLS03]. geocast [CL03a]. geographic [LAGK07]. Geographical [PFJ04]. geographically [ZWL03]. Geometric [Abr96, BMRC99, CDRC99, GM96, KV88, WPKK94, AG86, KK06]. Geometric-Decaying [GM96]. Geometry [DRC90]. Geomulticast [AP03]. GF [KA91]. Gigabit [HcF05]. given [DDNS06]. Global [BLPV95, KCRB99, LWY97, LA93, MT95, MI92, Mat93, OK02, Par96, TG97, Van94, WT09, Yen01, AY89, Car90, CK08, DK04, GJG88, JLM08, Lun90, SK89a, VB08]. glueless [RFPAG08]. GMA [ZFS07]. GMAC [GZMC08]. Gnutella [BAL05]. go [PL03a]. Good [BEE00, DP99, SK94]. Goscinski [BCC95]. gossip [FM07]. Gossiping [FV97, GRV97, SGS08]. gossipings [KLC05]. GPU [RRS^{+}08, SA08, SDG08, TSD08]. GPU-sorting [SA08]. GPUs [IBP08, RGU08, SHT^{+}08]. Gracefully [BBR94, CGA98, LH92, RCB93]. Gradient [Bas97, BM08, PB09]. gradients [McA89]. GrADSolve [VD04]. Grain [FR92, LFA96, Mah95, NS97, SA93, CT94, FW05, GSWW04, PL03b, TKHG04]. Grained [BR96, CDRC99, CLZ00, DFRCU99, HK96, PY96, SR97a, SR97b, WD94, BM04b, FSD04, GVA^{+}08, IKS87, IBP08]. Gram [ZLRP91]. grand [SAB^{+}92]. Granularity [CDH84]. Graph [AyJ93, CCM01, CHGM01, GJP96, HJ90c, Kar95, KK98b, KC98, KA99, Lat95, MJ94, OSZ98, RW97, RWY93, RLS96, SAOKMA02, TVS97, TLW94, WCE97, ZW00, BDjQ86, BM08, CM03, DeG88, HLM^{+}90, MPZ09, MMS09, PK07, Ros89, SW91, SGR03, WCC02, ZNQ93]. Graph-Based [CHGM01]. graph-partitioning [SW91]. Graphical [CMT93]. graphics [CBM^{+}08, KL08b, KME09, SCB08]. Graphs [ANS97, AKPT99, AS96, AKP95, BS97, BP98, CP98, CA95a, CDF01, DDD98, DS84, DH94, EMMM94, FA95, GY92, GS98, GSG^{+}93, GS99, HOS94, IZ95, JR95, JSS92, KK98a, KW02, KA97, OS97, PRW94, Par98, RDL95, TL96, VB96, WIKC97, WLD00, ANP07, BC06, BKS05, BD05, CP04a, FT04, Hsi04, HS03, Lin03, Lo92, MHPR05, MSZ05, Nik04, PD05, PK04b, SS03, SP90]. Gray [BVB02, HHM94, HRJ94, JH94]. Gray-Scale [HHM94]. Gray-to-binary [HRJ94]. Great [KF90b]. Greater [Ebe94]. Greedy [KNS06, BGM^{+}08, HDJ08, LLS07, Cho90]. Grid [AKPT99, BR02, BAK^{+}03, SDG08, TF01, CCEB03, CGW^{+}03, Eİ07, FGZ03, LBE03, LFH^{+}03, LB09, MC03, PF04, VD04, ZV09b, AOS^{+}05, ABCM07, BAS06, CS06a, CTT08, CCN06, DBC03, EDÖ05, GBA08, KTF03, KVHS07, KKS08, LCC^{+}05, LLLY08, Li05, LL07, LTIK05, LR05, MCT06, RAB08, SV08, SAOKZ05a, SAOKZ05b, SXZ06, SSM^{+}06, SFEF06, TYH09, TMM06, TD07, VPHML06, WS06, YT05, YWD08]. grid-aware [FGZ03]. Grid-Based [BR02, VD04, KKS08, GBA08, LLLY08]. Grid-computing [BAK^{+}03, SAOKZ05a, SAOKZ05b]. Grid-enabled [KTF03]. GridBench [TD07]. Grids [CCCM96, HKMU98, HOS94, ACFK07, GRDB05, JV09, Mit07, PHS04, ABCM07, GTN^{+}06, GJA08, Ngo06, TZ06, VB08, WW03, WLL08]. grooming [FMM^{+}08, WG08]. ground [BFKP04]. Group [RGVB00, CHC05, Dim91, LHT08, MM07c, XO05]. group-shared [LHT08]. Grouping [CWP98]. Groups [Oru87, WLD00, CHC05, GCS06, MS05, Ros89]. Growing [CRFS94, WLR90, MGG03]. growth [WCKD06]. GSM [TM06]. GSPN [CCM92, CCM01, SM92b]. guaranteed [HWWH08, PY09a]. Guaranteeing [Sch91]. Guarantees [MS00, OY00]. Guessing [DKY01]. Guest [WW03, AP93, AL99, AB03b, Ano01j, Ano01k, Ano01l, Ano02g, Ano02h, Ano02i, BG90b, BD00, Cas93, Che92, Cho93, DOP98, ES97, GGB93, GC95, Her92, JW94, Kri92, Lin93b, MC93, NT90, OW01, PN97a, PN97b, Pan09, PA96, Sch90, SH92a, Sto90, TY95, WC05]. Guidelines [Ano00d, Ros99].

h [CP04a]. HA03094L [Ano04e]. Half [RS94]. Half-Duplex [RS94]. Hamiltonian [DP98, Hsi04, LSC00, Nik04, Wan01a, WCC02, YTH07]. Hamiltonicity [HTHH02]. handheld [WL04]. Handling [BW09, CVJ09, SYG92]. Handoff [SK05a]. Hard [DJ98, BRR01]. Hardware [GS00, MD01, RPS93, SCC^{+}06, TF92, The02, TH08, VH93, ABC^{+}09a, AF06, BJS03, GVA^{+}08, HDJ08, Nik03, NAK04, PVG09]. Hardware-Efficient [MD01]. Hardware-Only [GS00]. Hardware/software [SCC^{+}06]. Hardwired [DM88]. HARNESS [MSS00]. harnessing [VPHML06]. HARP [SSB98]. harvest [WS06]. Hash [SX08, HKW05, TC04]. Hash-based [SX08]. hashed [HSMB91]. Hashing [WPKK94, YB95]. having [BSMH08]. Hawkeye [ZFS07]. Hazards [AGG98]. HCL [Pfe90]. Heap [DP98, ZK94]. heat [LGG08]. Height [LP96a]. Height-Limited [LP96a]. Helary [Ano96l]. helper [DKRI09]. Hereditary [CDF01, Hsi04]. heterogeneity [BKS05, CL03b, XQ07]. Heterogeneous [ANT02, Ano97k, BSS97, BPR99, BSB^{+}01, CP97, CA94, CEF^{+}95, DAYA02, DBP94, HS94b, HC97, KL01a, LAS^{+}97, LHHB^{+}01, MAS^{+}99, MSd^{+}95, MP96, NRS95, NDZA99, PP92, SC91b, WR97, WSRM97, Won99, YZS96, ALF03, BD05, BČFF05, BR08, BRP03, BH05, BSMH08, CSW08, CCK^{+}08, CDR09b, CGW^{+}03, DK08, DÖ06, FMR05, GRV08, GNT04, GWWL94, GMXA07, KUA07, KSS^{+}07, KAS07, KMS^{+}06, LR06, LLL06, LMR05, LPX05b, LLS07, MVB05, MTS90, NP09, PKN08, PTA08, Pla08, QJ05, QGL^{+}09, RN04, SX08, SCS^{+}08, SSM^{+}06, TG03, UAKI06, WSG91, YH07]. HeteroMPI [LR06]. Heuristic [BA92, DDD98, EHMN95, KLZ97, XH93, HS06, KJD03, PM05, SWP90, VB08]. Heuristics [BSB^{+}01, GY92, GJP96, IAS^{+}92, KUA07, TSC01, AKSM08, KA08, LLS07, ZHO03]. heuristics-based [KA08]. hexagonal [GSSS03]. HHN [YP96]. Hiding [HF02, WL92]. Hierarchical [AGF94, Buc92, BM95, CAB94, FR96a, HR92b, HR92a, yHY97, KZ96, LLJ00a, MS00, OM90, SHT^{+}95, TM06, TJ92, Tan84, TW89, VSIR91, WHT00, YP96, AGMS04, BAS06, CKO04, DE91, DM04, LTL06, RH05, RR05, SS05, WLL08, ZZ90]. Hierarchical-Memory [VSIR91]. Hierarchies [VN93, BW89]. hierarchy [Pad91]. High [ABDS02, BJ99, BBH^{+}97, BNSP99, CY99, CD98, DS02, FGKT97, FM99b, GP93, JSCB95, JLRA97, KMKD97, KS95, KRS01, LC97, LS01, MR94b, NKC^{+}97, NTC03, PF08, PVG09, TF92, TMM06, AB03b, BSW07, BDDL09, CCC^{+}04, CBP02, CTCX08, DK08, DB08, FGP05, GTN^{+}06, HOE^{+}09, HCZ04, HT90, JBY^{+}05, KME09, LWR^{+}03, LVB07, LZSL06, MG09, Nap90, NRM^{+}09, PK07, SD91, SC04, SAB^{+}92, SR91]. High-Availability [LS01]. high-dimensional [HT90, PK07]. high-end [FGP05]. High-Level [BBH^{+}97, CCC^{+}04]. high-order [KME09]. High-Performance [BNSP99, CY99, FGKT97, JLRA97, KMKD97, KRS01, NTC03, AB03b, CBP02, HCZ04, JBY^{+}05, LWR^{+}03, LVB07, NRM^{+}09, SD91, SC04]. High-Priority [TF92]. high-radix [MG09]. High-Speed [BBH^{+}97, SR91]. High-Throughput [FM99b, BSW07]. Higher [GSSS03, AM06]. Highly [BDHF90, CAB94, Joh94, MD01, NKC^{+}97, VH93, WIKC97, ATH91, GV86, SM08b]. Histogramming [BJ96]. historical [SFT04]. history [WBTM09]. HLR [FCF00]. Hoang [Ano92c]. Hoc [Ano01e, BDF01, GS01b, LAZC00, Pat01, AP03, ALF03, BFG^{+}03, BGLA03, BOP06, BN03, Bou03, CNS03, CW05, CYZ06, CDCD05, DW06, DMB^{+}03, DB08, EBE08, FVCL05, GAGPK03, GS03b, GMS06, GMXA07, HW03, HJ07, KK06, KNS06, LR03a, LPX05a, LW06a, LR03b, LHT08, OSL05, OMSGNSG05, SSM^{+}06, SGS08, SKMM04, VA03, WTB^{+}08, WGS08, WBTM09, XHG03, XWC^{+}08, XG03, YC04, ZMC06]. Homogeneous [LS97, GHS86, SCJ^{+}08]. honeycomb [BPRS04]. honeyfarm [JXW06]. Honeypot [KMMZ06]. hop [BSW07, MAM05, NC09, YMG01]. Horizons [BP95]. Hot [LKK94, NS95, TY90a]. hot-spot [TY90a]. hotspots [MLG05]. Hough [BA95, CP91, Fer93, GZ97, JS94, SSL04]. HPC [SCB09]. HPF [BCF^{+}94, CA96, HLJ01, KHS96, SS00]. Hull [DFRCU99]. hulls [GS03a]. Hut [SHT^{+}95]. Hybrid [Dah99, FA07, Gao93, NBM93, OS93, AC89, BAMM05, CCQ^{+}06, FX06, SCS^{+}08, SHLN09, SSL04, SA08]. Hydrodynamic [HC97]. Hydrodynamics [PAH^{+}98]. Hyperbolic [SSK96]. hyperconcentrator [CL90]. hypercontexts [LM05]. Hypercube [AGF94, AM93, BKT95, BC94, CS93c, DP98, DMSH90, DRC90, DFN^{+}94, FAM96, FPD93, GGD93, GT97, GBG93, HGCC96, IK93, IK94, JR92, JB98, KB96b, KM91, Lan94, LH92, LLJ00b, LEB98, Man94, MP93, MW95, MYD95, NSLK99, NT93, Nas94, OM90, RS94, Raj96, SYO94, SCC92, SY01, Sto90, TLW94, TL96, TC92, WIKC97, Wag93, Wag94, XMN92, YP96, Zia92, Cap87, CCS06, DE91, Efe91, EAL90, ERS90, Joh87, KAP90, LEN90, LSS88, LS91, MVM04, MAR87, RS90a, RS90b, RIZ90, SW90, TS91, Wag89, Yan04, ZLRP91, YN92]. Hypercube-Based [Zia92, DE91]. Hypercube-Connected [LH92]. Hypercubes [AD95, AERBL92, Ann94, CL93, CCCM96, CS95a, CCR94, Efe96, Fag92, FM96, Fra92, GP00, GH93, HM01, HOS94, Kav93, KF95b, Li92, LBT94, LW95, LT96, Moh97, OD95a, OP96, Pel95, PM92, RS96a, RJMC95, SHL95, SR95, TT98, WW97, Wan01a, Wu94, WFL98, YTR94, BG90a, BM04a, BOS^{+}91, BL89, CL91a, CL91b, Che05, Ede91, FT04, GT04, GNW03, HNSA07, Ho91, HRJ94, LW90, SS89, Var91, Wu85, Wu03, XCS06]. hypergraph [ACU08, CBD^{+}09, DHK04, KJD03, TK08]. Hypermeshes [OK01, Szy95]. Hyperoctrees [DFN^{+}94]. Hyperplane [HS93]. Hyperreconfigurable [LM05]. hyperspectral [PVPM06, Pla08]. Hypersphere [AM93]. Hyperstar [AAD98]. hypertree [LTD^{+}93].

I-Caching [MM93]. I/O [AW95, CkLCK04, CkLCK05, Cho93, CQ95, CD95, DD93, DT01, DJT03, EH01a, GGD93, JSCB95, JSWB92, LTH97, MLG05, NSSS99, NsPPC02, WLWW09]. I/O-Intensive [EH01a, CkLCK04, CkLCK05]. IBM [ASH^{+}01, BAHP01, BR95b]. Id [HCAA93]. Identification [CS95b, EBE08, FCC07]. Identification- [CS95b]. Identify [XYG07]. Identifying [HS03]. Idle [CW93, CM92]. IDOS [BA01a]. IEEE [Ano93a, BCD00, FA07, VHH08]. II [HR92a, SAOKZ05b, SR97b]. ILU [SZW05]. Image [BJ96, BM95, ELS94, HSJP87, HC95, KSL85, KC99b, LWY97, MWL00, MG98, NEG85, OS98, RS90a, RG87, SR94, SD88b, WS95, ZM94a, CDJ^{+}89, CCN06, GSWW04, IKS87, Kep03, KM03, Lee91, MGG03, PI90, Pfe90, Sto87, SA90, UAPM07, Wan07, WRHR91, WJD91, WGCZ09]. Image-Processing [KSL85, SD88b]. imagery [PVPM06, Pla08]. Images [SYO94, Ara90, CL85, DH91a, NAK04]. immersive [MBH^{+}08]. Impact [Buc92, Kel00, Tze91, GSWW04, MLG05, SYYU07]. Implementation [AS95, BAHP01, BHS^{+}94, CP91, CP92, CS95c, DM90a, DBKF90, EP90, HS97, HBH93, KM91, MSS00, NT93, NsPPC02, OS98, OP98, PAJC97, RL02, RW01, Shu95, SM00, Ski96, SOG94, TVO92, VBM90, XMN92, YB01, BFTV87, BG89, CEGS07, CPO^{+}03, FGG08, Gro85, JK89, KTF03, KA91, KP05, ML89, MML07, OO05, PLD87, SM08b, SMKL93, TR89, Tay87, XWC^{+}08]. Implementations [DT01, KL84, SAC^{+}98, WPKK94, BCM06, BRPR06, GNS09]. Implementing [BC94, Coh90, DRC90, GSC96, HK08, MT95, DM90b, OB88]. Implications [AH94, BS96a, GTN^{+}06, MT96, MG93, SH92b, TSA97]. Implicit [BAM93, Fre96]. Implicitly [SAC^{+}98]. imposed [BKS91]. Improve [CB02, DS95a, SKH96, CDR09a]. Improved [AM97b, AS91, CLZ02, Che05, DL98, FT04, GJP96, JR95, KLC05, Mil99, PB95, Tsu07, Wor93, Bad04, TDC05]. Improvement [yCM98, IAS^{+}92]. Improvements [GCB^{+}00, WSS93, DPSD08]. Improving [Dah99, DK04, GT02, GP05, GMM00, Kan05, LTL06, MBR08, WTB^{+}08, CCK88, YF09]. in-order [KMF^{+}05]. inclusion [RFPAG08]. Incomplete [OD95a, PK04a, SCD99, TC92]. Incompletely [BSGM90]. inconsistency [Ram89, TK07]. Incorporating [AISS97, VWHL96]. increasing [RS08]. Incremental [ZN01, LY08]. incrementally [SSB91]. Independent [BSB^{+}01, Ger98, Hag97, MAS^{+}99, NMS93, PS93, AK06, AY09, CL91b, Li06a, LH09, LB09, LLS07, SZW05, SSM^{+}07, YWD08]. independently [XCH08]. Index [Ano92b, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98l, Ano99a, Ano99b, Ano99c, Ano99h, Ano00b, Ano00c, Ano01f, Ano01g, Ano01h, Ano01i, Ano02c, Ano02d, Ano03a, Ano04b, Ano04a, Ano08, Ano09, KHS96, SSHC00, Ano03b]. indexes [OC07]. indexing [FKJG08, GZ08]. indirect [Ho91]. Induced [WIKC97, LM09]. Induction [BF01]. indulgent [WCYR08]. Industrial [MS99a]. Inexpensive [MT93b]. Inference [AyJ93, FBRW03, PTZ06]. inferencing [MK08b]. InfiniBand [ASD09, PK05b]. Influential [TAS^{+}01]. Information [Bal90, BS96a, CY99, LA93, Oza04, CKN07, DB86, LY91, Pla08, Psa96, Raj08, RFPAG08, SSS07, SFT04, XCS06, XQ04, ZFS07]. Informed [LM09]. infostations [BPRG04]. Infrastructure [GC01, LCM^{+}06]. infrastructures [Ano04d, BJPPM^{+}08, TD07, YK04]. Inherent [WW98]. Initial [dGP06]. Initializing [Nak95]. initiation [MM04]. Injured [Wu94, Wu03]. inner [Lis90, ST85]. input [LY08, NAK04, PMV05]. Insensitive [ST02, ST06]. INSIGNIA [LAZC00]. Instance [SM94]. Instruction [AGG98, LPU97, Gro85, Sch89b]. Instructions [dSR00]. Instrumentation [GP91]. Integer [DL98, Fag92, SS96, KKVI05, VM95]. Integral [Ten90]. Integrated [BDHF90, DAYA02, OY00, PW96, WAE03, YSL08, ZR00, ZMC06, HC09, SKMM04, XYDL06, XHY07]. Integrating [Bir94, DRST02, FKT96, Lu01, OK02, PY96, YT05]. Integration [ISZBM99, KL84, LY01, Ano04d, HMV07, YK04]. integrity [LZSL06]. Intel [FPD93, SMKL93]. Intelligence [MT85]. Intelligent [IAS^{+}92, KSP^{+}92, SH98, ZL93, CDJ^{+}89, She09, WJD91]. Intensive [ABM^{+}92, BS09, CA95a, EH01a, SW90, CkLCK04, CkLCK05, KAS07, SC04, VB08]. inter [FKLB08, Kan05]. inter-node [FKLB08]. inter-procedural [Kan05]. Interaction [CCM92, DH95, LLCC02]. interactions [CK08]. Interactive [LHM95, RGS00, MAR05, TSD08, TD07]. Interactive-Rate [RGS00]. Interconnect [HP97a, WLY01, MG09]. Interconnected [DH95, EH01b, Guo94, KM97, QMCL94, GMH^{+}91, McA89, TRSS06]. Interconnection [AAD98, AA95, BETD94, CW01, CJA09, DVZ96, FD86, KRSZ02, KAM94, Lat95, LYL93, MLW^{+}97, MSH90, MC93, MJ94, OM84, OO85, Pad93, PL93, SW96, SZB92, Szy95, TH02, Tze91, VB96, Wan96, Wan01b, Wil92, YWP00, ZMPE00, ZW00, dBL95, BDjQ86, BHR91, BR91a, Bhu87, BR91b, CM04, CKO04, CS06b, DE91, FJC04, Har91, JBM91, KRL87, LK90, MHBW86, Pak89, Par05, SSB91, SL89, SH89, WCC02, Wil90, ZDC06]. Interconnections [LLJ00b, SL97, THN^{+}93, Oza04, YB90]. Interconnectivity [DSD^{+}97]. Interconnects [ES97, HP00, MO97, MG93, PEC95]. Interface [BAHP01, BF97, BDH^{+}97, CD98, IWM97, PS01, RS92c, KTF03]. interference [BPRS04, KDH08]. interference-aware [KDH08]. interleaved [NC09]. interlock [CCK88]. Intermittent [DT02]. Internal [Bal90, JZK04]. International [Ros07, Sni03, Wee01]. Internet [Bar05, KA08, She09, TB90, WLID02, XO05]. Internet-based [She09, XO05]. Interpolation [CWW^{+}95, Goe94, SAOKMA02, Nic07, PHS04, Sch89b, SDG08]. Interpretation [FAGW95]. Interpretive [PH00]. Interprocedural [HHKT96, CK88]. Interval [CI03, PT01, Sch87, Sta04]. Interworking [WH08]. intra [Kan05]. intra-procedural [Kan05]. Introduction [AP93, AL99, AB03b, Ano01j, Ano01k, Ano01l, Ano02g, Ano02h, Ano02i, BG90b, BD00, Cas93, Che92, Cho93, DOP98, ES97, GGB93, Gau06, GC95, Her92, JW94, Kri92, Lin93b, LR05, MC93, MGS^{+}06, NT90, OW01, PN97a, PN97b, PA96, Sch90, SH92a, Sto90, TY95, IB04, WW03, WC05]. intruder [ISAZ07]. Intrusion [BN02, LLLY08]. invalidation [OFS03]. invention [MC03]. Inverse [CTZ99]. Inversion [SW96, mYyF92]. investigation [CD95]. Invited [Ano01m]. invocation [BBB^{+}06]. IP [HZY04, HC09, JP09, JBY^{+}05, KERUM04, LAZC00]. IP-Based [LAZC00, JBY^{+}05]. IPDPS [Pan09, Rob09]. iPSC [DHR96, FPD93, SMKL93]. iPSC/2 [FPD93, SMKL93]. iPSC/860 [DHR96]. IRISGrid [VPHML06]. Irregular [Ano96i, DUSH94, FR98, FBK98, FY97, KK98a, LWP02, MRRV98, Nic94, NsPPC02, RWK95, WP02, CB06, GRR^{+}05, SPBR91]. Irregularly [MNM98]. ISA [SM08b]. Island [CGKK97, GB06]. Island-Based [CGKK97]. islands [dGP06]. islands-based [dGP06]. Iso [KF95a]. Iso-rectangles [KF95a]. ISODATA [DSAUM99]. Isomorphism [GS99, KW02]. isosurface [WJV07, ZB09]. Issue [AP93, AL99, Ano95i, Ano96j, Ano96i, Ano97j, Ano99g, Ano01e, Ano02v, BG90b, BD00, BS09, Chi92, CDJL09, DOP98, Dek00, DT92, ES97, FR98, GC95, GS01a, Gra09, JW94, KRS01, Lan09, Lin93b, Mir91, NT90, Ola01, PN97a, PN97b, PA96, Sch90, SH92a, SB97, Sto90, TY95, Wee01, YW91, ZO97, AB03b, BOP06, Irw88, IB04, KL08a, KL08b, Raj08, SXZ06, WW03, XJS03, dVCP06]. Issues [Ano95j, Ano00e, Ano00f, Ano00g, Ano00h, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Ano01-32, Ano02q, Ano02r, Ano02s, Ano02t, DVW94, MFS93, Nie94, PS01, THBF97, BAK^{+}03, GCY^{+}04, TB90]. Item [AAP01, San99]. Iteration [BW96, CC87, RS92a]. Iteration-level [CC87]. Iterations [AR97]. Iterative [Bah00, BSS99, CTD99, CHR94, ESMG96, IPK85, LPX05b, ÜD96, WB96, CF88, CRC^{+}02, FGG08, KMS^{+}06, VGAB08]. iterator [Lon04]. iTPS [TDC05].

J [WTC08a]. Jacobi [EP90, MVV91, MV94, RS08, ST87, ZB97]. Jacobi-Type [MV94, MVV91]. James [Ano92c]. JAVA [MSS00, AFT^{+}00, CCK^{+}08, Dek00, GCB^{+}00, GLC01, HR00, HS00, JM00, MWL00, SCB09]. Java-Enabled [MWL00]. JBSP [GLC01]. Jean [Ano96l]. Jean-Michel [Ano96l]. Jerzy [Ano96l]. JESSICA [MWL00]. JMX [JM00]. Job [FKSW97, Li05, FCC07, GRDB05, LC90a, MS86]. Jobs [CB02, CL91b, LF03, MLG05, QJ05, SF05]. Join [HTL99, LT94]. joins [CG86]. Joint [AF06]. Josephus [LH05]. Journal [Ano99g, Ano97j, BS09, CDJL09, Gra09, Lan09]. jpdc.1999.1564 [Ano00d]. JPEG [CD95]. Jumping [HIKM94]. Just [FKLB08]. Just-in-time [FKLB08]. juxtaposition [BKS91].

Kalman [LWOG02]. Kapelnikov [Ano92a]. Karhunen [FSD04]. Kautz [CC94]. Keeping [Bal90]. kernel [IBP08, SK91]. key [GCS06, GMXA07, SB04]. keys [PPC04]. Kinetic [RW01]. Knapsack [FR96b, Ten90, EE05, LSS88, LS91, PMV05, GT04]. Knapsack-like [FR96b]. Knowledge [CHGM01, DL99, EHS94]. Kohonen [VM95]. Krylov [BSGM90]. Kutta [KR06].

L [Ano00d, CS93b, CP04a]. L. [Ano93e]. Labeled [FM96]. Labeling [HHM94, KR98, SR94, CP04a, CT94]. LAD [DFP06b]. Lagged [Alu97]. Lagrange [Goe94, SAOKMA02, ZZC92]. Lagrangian [Kal04, Kal04]. lags [LY91]. Lamport [Lo92]. LAN [HWW96]. LAN-Connected [HWW96]. Lanczos [Bas97]. Landmark [XHG03]. Language [BCD95, BBH^{+}97, BN94, BHS^{+}94, CC91, DRST02, FCO90, FC95, FKKC97, FMW^{+}94, LS85, Chi95, ESA03, PLD87, Pfe90, RSW91, ESA03, LTIK05, SBKB90]. Languages [BS90, KBC^{+}01]. Large [ABDS02, Ano92c, BP01, BMCP98, Efe96, Fag92, GK98, GK93, JH92a, LK98, Lin93a, OK01, PTZ06, SR95, SM04, VN93, WRC^{+}02, BMB^{+}08, BA06, BMF05, CS06a, CVJ09, DHK04, HRC09, KESA07, LYL08, Lon04, PK07, RW02, VM03, XHY07, YH07, ZV09a]. Large-eddy [SM04]. Large-Scale [ABDS02, BMCP98, LK98, OK01, VN93, BMB^{+}08, BMF05, KESA07, LYL08, VM03, XHY07, ZV09a]. large-size [CVJ09]. Larger [Mah95]. largest [Deh90]. LARPBS [dR09]. Last [Tay02, RFPAG08]. last-level [RFPAG08]. Latency [GS00, HF02, KUFM02, MR94c, MG91, RJY96, ZYH94, IS06, KS03, PRHB06, WL92]. Latency-Tolerating [GS00]. lattice [GMS06, IBP08, WCO^{+}09]. Laws [FLS^{+}97]. Layer [BNSP99, KNS06]. Layered [DDD98, SSK96, CI03, LHF91]. Layers [ZAW94]. Layout [MB96a, MLG05]. Lazy [GSC96, MYD95, DS04b]. LDU [MVV91]. Leader [AS96, SS05, Pel90]. Leaders [SK94]. Learning [CW92, WT92, AC89, MS86, SM08b, Tor89, VM95]. Leashing [DHS06]. Least [CB95, HLS03, KAP90, ZYO02, BBd90, SMKL93, TBZB05, XBK07]. least-mean-square [XBK07]. Least-Squares [CB95, ZYO02, HLS03, KAP90, BBd90, SMKL93]. LED [MLW^{+}97]. Lee [BVB02]. legacy [LWR^{+}03]. Legion [LFH^{+}03]. Length [BL94, MP08]. Level [BBH^{+}97, BSS97, CD98, GS98, HKT^{+}91, HWW96, Kav93, KOW97, KL84, MR94b, MHC95, Qia97, RP95, SSHC00, SBKB90, AY09, ACU08, CCC^{+}04, CC87, CTCX08, IKS87, MAJJ05, PRHB06, Pfe90, RFPAG08, VD04, WCKD06, YSL08]. Leveled [PRW94, BMIM07]. levels [Wu03]. Lexicographic [AMS94, DT97]. Lexicon [Haw97]. Libraries [KBC^{+}01, ZRC99]. Library [BMCP98, CJ99b, DVW94, FKKC97, GLC01, HWW96, SKH96, LR06, RR05]. Library-Based [FKKC97]. Life [HSJP87]. lifetime [HP06]. Light [RGVB00, Koc91, Wan06]. Light-Weight [RGVB00, Wan06]. Lightweight [HS00, CL09]. like [FR96b, GL90]. Limit [MO97]. Limitations [BKS91, LS97]. Limited [yHY97, LP96a, LK98, BKS05, DW04, WTB^{+}08]. limits [DW04]. Line [BDKM94, BMMS01, LTY96, RR95b, Yen01, BS92, DMCFCM03, DJ98, EL88, GH89b, GC07, KM88, LHK03, SSL04, SL90]. line-sweep [DMCFCM03]. Linear [Bah00, BBM^{+}02, BMM97, BCZ95, CDH84, CCC92, DVW94, IPK85, IK94, KL01a, KF95b, LP97, PM96, Pov99, RFM94, RS92b, ST89, TBPV00, ZZC92, dR09, BGH^{+}03, BAH04, BPP05, Car90, CM03, CEGS07, DS04a, Dja06, GPT06a, GRV08, Gao86, GS91b, HR89, Joh87, KKVI05, KT89, MP88, MP87, MVB05, NCTT09, YTH07]. Linearization [FZVT02]. Linearly [BBd90, PB90]. Lines [HKMU98, Wri91]. Link [GDP08, MLW^{+}97, VR94, VR95, WFL98, RH05, SW90, WTS03]. link-bound [SW90]. link-selection [RH05]. linkage [CPO^{+}03]. linked [Han89, HA05, ST08b]. Links [ABJS01, KJ84, RS94, WW97, Wan01a, KPR88]. Linpack [Num07, Num08]. Linux [BP01]. List [BBH^{+}98, SP96, SGS99, Han89, LPX05b, Vis87]. Lists [BP02, VSIR91, ST08b]. Load [Ano97j, BEE00, BM08, CS93a, CRL04, CLZ00, DHB02, DMB97, DLLX97, DSW94, Efe96, EE05, FMP98, FLS^{+}97, FM99b, GK98, Gil94, GM96, HS97, HlLLY95, HTL99, HO94, HC97, JR92, JW89, KGV94, LK94, LHVW95, LT94, LL98, MDD97, MP96, NSLK99, NFEG97, OB98, PB99, QY94, SH92a, SHT^{+}95, SB97, SBAM96, TSHH01, TT98, Wan96, WS97b, XYKA08, XL92, XH93, XL95, ZLP97, ZXP09, ZM94b, vS91, AGMS04, BCV05, BFH09, BRPR06, BD04, CSWD03, CBD^{+}09, CVJ09, Cho90, CRC^{+}02, Cyb89, DW04, DM94, GRV08, GC05, HJ90a, HLM^{+}90, IC05, IS06, JL05, KKS08, KC04, LTB02, LTL06, LLL06, LHKL03, LY91, MVB05, MTS90, Mit07, MGG03, Nik03, PA04, RN04, SU87, SX08, TBZB05, TKHG04, ZV06, dG91]. load-adaptive [TKHG04]. Load-Balanced [LT94, NFEG97, XYKA08]. Load-Balancing [DHB02, FM99b, HO94, HC97, Wan96, ZXP09]. load-sharing [SU87]. Loads [KC95, VB02, GRV08, MVB05, ZV06]. Local [AD02, BSS99, BCD00, CGL^{+}95, FLS^{+}97, HR00, SR94, AK07, BMARM07, CKN07, GJG88, MS88, MAR05, XCS06]. local-spin [AK07]. localities [GJXZ05]. Locality [BS96a, CL96, FJG06, KCRB99, KRC00, MNB95, SCM99, SHT^{+}95, FPP06, Kan05, KR06, SZD07, WLL08, XCZL03]. locality-aware [XCZL03]. Localization [DFP06b]. localized [Cal06, KNS06, LS03]. locally [AMK^{+}07]. Location [KER01, LS03, LAGK07, MMRS98, XCLR07, CZ90, TZ07, TDC05, ZMC06, ZHO03]. location-aided [ZMC06]. Location-based [LS03]. Location-centric [XCLR07]. Lock [DR98, ST08b, CB06, Dim91, HA06, ST05, XO05]. Lock-free [ST08b, CB06, HA06, ST05]. Locking [MS98, XO05, DM04]. lockless [HMBW07]. Locks [JNW96]. Lockup [SD91]. Lockup-free [SD91]. Loève [FSD04]. Log [NTA96]. Logarithmic [Nas94, OOW95]. Logarithmic-Time [Nas94]. logging [CZZY09, DWG03, JLM08]. LogGP [AISS97]. Logic [AyJ93, CC91, CBdCD00, Mon94, Tan84, DeG88, MV88, MC91, NAK04, SK90, WF89]. logic-oriented [SK90]. Logical [AK93, YMG01]. LogP [AISS97, BHPP05, RGD03]. Long [AISS97, GO95, Lin93a, MBR08, TDC05]. long-distance [MBR08]. long-range [TDC05]. Longest [MS99b, PK04b]. Look [PL93, TG04].

Look-Ahead [PL93, TG04]. Lookahead [NIR86, SF05]. lookup [JP09]. Loop [AMB95, BCH95a, BCZ95, CG02, DR95, DS95b, Nic88, OK02, PB99, QGL^{+}09, AL04, KSG03, MP08, NCT^{+}07, QSL^{+}08]. loop-carried [NCT^{+}07]. Loop-Free [CG02]. Looping [Ano92a, KME92]. Loops [CCC90, CWW96, DRR96, HS93, KK95, KBG92, SCMB90, SG99, Xue97, CC87, SGE91]. Loosely [SKR93, AjHcC90, BMF05]. lossless [PY09b]. Low [AZ01, Ano92c, IKS87, JH92a, JNW96, JLRA97, KS00, MHC95, SD00, CBP02, CL09, GE85, GJXZ05, KS03, Pfe90, SZ09, SLWW05]. Low-Cost [AZ01, Ano92c, JH92a, JLRA97, CL09, GJXZ05]. low-latency [KS03]. Low-Level [MHC95, IKS87, Pfe90]. low-memory [CBP02]. Low-Overhead [SD00, SZ09]. low-resolution [GE85]. Lower [BMRC98, JR95, LPS^{+}98, TC96, WW97, FT04, ITT04]. Lower-Dimensional [TC96]. LPAR [BK95]. LQ [BBM^{+}02]. LR [CB96]. LU [OT86, She06]. Lyapunov [MV94, QOvdG01].

M [Ano92a, FC95, LZSL06, ZBF05]. M-TREE [LZSL06]. M-VIA [ZBF05]. MAC [CCHC09, Los08]. Machine [BG86, BDHF90, CA95b, LWOG02, MB93, SYO94, SR97a, SR97b, TVS97, ZL93, BH86, HS86, HPSM91, KNS91, KA89, Ros85, SM86, WF89, CM93, CRFS94, CGSV93, EHS94, LAD^{+}96, LTD^{+}93, Sab94]. Machines [BR96, BPN90, BCR96, CWP98, ERL90, Gup92, GKHS96, HK96, HB97, HLJ01, KRC00, KHS96, KLS90, LWY97, MK92, PAM94, RS94, RWK95, RGS00, SSG93, SCMB90, San02, TSA97, Zak01, AE88, Fen90, FX06, GA90, IKS87, Koc91, KP05, LC91a, Mar88, MAR87, SW90, Ume85, ZA91]. macroeconomic [BMB^{+}08]. macropipelines [WAS88]. magnetic [CCN06]. Main [DM99]. Maintaining [HS94a, LY98, YC04]. maintenance [CDCD05, WDDK09, XO05]. Major [SSL04]. majority [ZWS09]. makespan [LZ05, SSM^{+}07]. Making [LFA96, VR95]. manage [ASD09]. Management [BR02, CKK00, CY99, HlLLY95, HTL99, JM00, KER01, LZ02, LO96, RDS02, RSBN01, TJ92, WLID02, YD98, ZRC99, GPT06a, GJG88, GBA08, HMV07, HC09, HSLL04, LCC^{+}05, LAGK07, NTC03, PY09b, PF04, SDTD04, SS08, SK05a, SL06, TZ07, TB90, ZMC06, ZHO03]. manager [Gai87]. Managers [AB84]. Managing [FGKT97, SEP96]. MANETs [ZA05]. Manipulation [PH91]. Manipulator [MS85, NS90]. Many [HP95, SR97b]. Many-Body [HP95]. many-core [Trä09]. Many-to-One [SR97b]. map [IB04]. Mapped [BF97]. Mapper [AM93]. Mapping [AGG98, BR08, BSB^{+}01, BA92, CN93, CHR94, CW92, Dja04, GH89a, GW99, IAS^{+}92, KBG92, LW90, LWY97, MM00, MAS^{+}99, NB93, SH90, Ser97, SBAM96, XH91, ZZ90, BS87, DQR^{+}09, HA91, KSS^{+}07, KMS^{+}06, LB89, Lo92, LLS07, ZWRI07]. Mappings [BP02, DP00, Iqb92, SR97a, SR97b, SSHC00]. Marginal [WLID02]. Mark [Ano92c]. Marked [JSS92]. Marker [MG98]. Marker-Based [MG98]. Marking [BBH^{+}98]. Markov [DHK04, NH93, PF91]. Markovian [VM95]. Mass [HLL^{+}95]. Massive [SANY94, ZB09]. Massively [BS90, BDHF90, EHMN95, GGN93, GBES93, JBL02, Kri92, KP05, MM93, MT96, NDZA99, NS92, PE93, Sch90, SRK95, TSA97, WT92, YP96, BB87, BBCLL04, GP91, HS86, Koc91, SPBR91, TS91, WZ91, LTKS90]. master [HSLL04, LZ05, LYL08, YH07]. Matching [BL94, DS84, DAYA02, LO94, Par98, WSRM97, MM07b, RS90a]. matchings [SM89b]. matchmaking [LR05]. Mathematical [HNSA07, ZA91]. Matlab [MJ01]. MatlabMPI [KA04]. Matrices [Bas97, BSGM90, SH97, BW08, ORR03, VGAB08, WF90]. Matrix [CT96, CTZ99, DBKF90, GK98, GE94, KCRB99, KK98b, LPZ99, Li01, Man94, MSC96, NFEG97, Par92, PKD97, SW96, TLW94, UZSS96, WM92, Win85, mYyF92, AAD05, BB85b, CLR90, Dja06, Ede91, EL91, EM89, ITT04, KK86, MS87, NJ91, NCTT09, OT86, SAOKM03, ST89, SM08b, ZB03]. Matrix-Based [KCRB99]. matrix-transpose [SAOKM03]. Matrix-Vector [GK98, MSC96, NFEG97, CLR90]. Matter [FGM^{+}03]. MAWS [AK06]. Max [DP98]. Maxcut [HP97b]. maxima [GS03a]. Maximal [CWW96, GS99, KW02, SSTP09]. maximally [Gao86]. Maximizing [MSC96, Ros99, AH06, KNS06]. Maximum [Als01, AS95, DDD98, FTL92, HP06, KEA95, Par98, mYyF92, SM89b, WMW09]. Maze [EL97]. Mbps [MLW^{+}97]. MDS2 [ZFS07]. Mean [BA92, JBM91, LZ05, XBK07]. Measure [ASR93, Kav93, PS93, SK89a]. Measurement [FPD93, KL01b]. measurements [JZK04]. Measures [GRR93]. Measuring [ZYH94, DI91]. Mechanism [Bal90, BCD00, JSM94, GVA^{+}08, She09, XO05, YF07]. Mechanisms [KPC96, KC99a]. Media [WUG99, HK05, XYDL06, XHY07]. media-based [XHY07]. Median [CCC92]. medical [CCN06, TSD08]. Medium [MSST99, WLNL06]. medium-scale [WLNL06]. membranes [PMV05, PMV06]. Memoriam [Ano04r]. Memories [CH92, PH91, Sin95, Yan93]. Memory [AD95, ACD^{+}93, AMN00, Alu97, ADS98, AS91, BR96, Bas97, BS96a, BCLR96, BF97, Bit92, BCR96, CB95, CP91, CWP98, CA95b, CJ99b, DS95a, DY99, DA97, DUSH94, DP00, DH95, DM99, DT92, EP90, FY97, GAG^{+}92, Gra09, Gup92, GKHS96, GHSJ96, Haw97, HPT02, HA92, HA05, HLJ01, IWM97, JF95, KRC00, KS97a, KHS96, Kel00, KC94, LWY97, LK98, Li01, LA93, MF94, MR94c, MS98, MG91, NSS97, OS98, PHB96, PAM94, PA96, PB99, PL95, PY96, RL96, RSB96, RWK95, RJY96, RGS00, SL95, Shu95, SS94a, SDS99, Soh96, SC91b, SB84, SN93, TJ92, TTG95, TY95, VSIR91, VN93, WW96, WD94, Wil92, YW91, YMR93, YB01, YL98, Zak01, AL04, BC06, BBM08, BJS03, BS92, BGM^{+}08, BCF^{+}94, CBP02, Car95, CJA09, CPO^{+}03, CK91, Cyb89]. memory [DFP06a, DI91, FZC^{+}05, FJC04, GJG88, GL90, HMBW07, HC91, ITT04, Joh91, KP05, LL90, LC91a, NSTN91, Nik03, Pad91, PK05b, PL03a, Pop91, QGL^{+}09, RFPAG08, SYYU07, SZD07, SM04, TW89, WL92, YLB90, ZFL89]. Memory-Access [Bit92]. Memory-Bounded [SN93]. Memory-Electric [IWM97]. Memory-side [HA05]. Merge [NT93, SM00]. Merging [VSIR91, AY09, DO89]. Mesh [AP94, Ann94, ADM^{+}94, yCM98, CCC92, CWW^{+}95, CLT96, CY96, CDP95, EL97, EH01b, FZVT02, Fer93, HHM94, IM00, JP95, JS94, JB98, KB01, LLJ00b, LME95, MD01, MP96, Moh96, Nak95, NSSS99, OS96a, RO92, RR95b, RR95a, SP96, SR94, SM00, Zhu92, ZYO02, ABC^{+}09a, ABC^{+}09b, BB85b, CL03a, Car90, CWL^{+}07, Dja04, Efe91, GH89b, HWWH08, HWC08, HR89, HR90, KDH08, KT91, LZ08, LC90a, LC91b, Li06b, Los08, LVB07, LV88, MLG05, MBR08, PB90, Raj04, SI86, SSM89, SC91a, SS94b, SZ03, VHH08, WH08, XYKA08, YSL08]. mesh-based [CL03a, LVB07]. Mesh-Connected [Ann94, ADM^{+}94, yCM98, CCC92, CWW^{+}95, CY96, CDP95, Fer93, HHM94, MD01, Zhu92, ZYO02, BB85b, Car90, HR89, HR90, KT91, LV88, PB90, SI86, SSM89, SC91a]. Meshes [BLPV95, BPvW96, BA97, BSDE96, BM97, BOSW94, BOS^{+}95, BGOS95, CW00, COS^{+}95, CL96, DS01, FF98, HCWS94, HJ90c, LS95, LSC00, LS94, MT93a, NPI^{+}96, NS94, OS97, OS96b, OSZ98, OB98, RWY93, ST02, SKK97, SJ95, VB94, WCE97, Wu02, YTR94, YCY^{+}00, BM04a, CI03, GLD06, KLC05, Mat06]. Meshing [YIY97]. Message [Ano94e, Ano95k, BB93, BKT95, BDH^{+}97, CW92, CZZY09, CD98, DMSH90, dADB96, GBES93, GHS94, GHS95, GHS97, HNM02, Isl97, Kar92, LK96, Li92, LW95, MD92, PY96, SCMB90, WTC08a, WTC08b, XH93, ZN01, BHR91, BR91a, BPW05, CV90, FM07, GH89a, GK04, Hal05, JLM08, LR06, LR03a, She06, TW87, vS91, KTF03, PS01]. message-driven [GK04]. message-optimal [CV90]. Message-Passing [CW92, dADB96, GBES93, HNM02, MD92, XH93, ZN01, GH89a, LR06]. Messages [AISS97, DLP99, FBDC99, LTWY95, LTY96, SKH96, BD04, CL90, GPT06b, KLC05]. Messengers [FBDC99]. Meta [SWC^{+}91, DÖ06, ZHO03]. meta-heuristics [ZHO03]. Meta-rules [SWC^{+}91]. meta-task [DÖ06]. metacomputers [Li05, LCM^{+}06]. metacomputing [BGH^{+}03]. metadata [HOE^{+}09]. Metalevel [Zim96]. metaphor [SK89b]. Metasystems [GWWL94]. Method [BC94, GHH92, KLLK98, PB99, WS97b, XL92, XL95, ZYH94, ABC^{+}09b, BFH09, BR91a, BBB^{+}06, KP05, Luk85, Mit07, ORR03, SMKL93, WCKD06, XWC^{+}08, ZB03]. Methodological [Bev02]. Methodology [Ano92a, BJ99, KME92, LR93, MB92, NMS93, PA94, PA01, SKR93, SK93, Che86, GL89, KME89, PF91]. Methods [Bas97, BSGM90, BR95c, Cas93, FGKT97, GL92, Kap93, KB01, Par92, SHT^{+}95, Wor93, XH93, BDjQ86, BM08, CEGS07, EE05, KG04, SWP90, UAPM07, VGAB08]. Metric [RJA97, ZYH94, SSMS08]. metrics [BSW07]. Michel [Ano96l]. microarray [WSH^{+}03]. Microelectronic [THN^{+}93]. microscope [FCG04]. Middleware [BNSP99, GJA08, SB04, Ano04d, CTT08, KAS07, MSAF04, PF04, SDTD04, YK04]. middleware-based [PF04]. Migratable [KOW97]. Migration [AMB95, CLZ00, Lat95, NPP^{+}02, SZ00b, CR96, Gai90, HSMB91, TH08, XYKA08]. Migratory [GS96]. Millenium [TAS^{+}01]. MIMD [BCF^{+}94, CJ99b, FAGW95, GGW96, GP91, HPSM91, MSC96, OD95b, PK89, RS90a, Shu95, UR94, VSM96, Vel89]. MIMDIX [MHF93]. Min [DP98, CRV94, ZNQ93]. MIN-Based [ZNQ93, CRV94]. MIN-Graph [ZNQ93]. Min-Max-Pair [DP98]. mincut [ERS90]. Minimal [CLT96, SJ95, SR90, Xue97, ZAW94, OMSGNSG05, SR88b]. Minimization [OKB95, QSL^{+}08, VA07]. Minimize [Als01, SBAM96, KSG03]. minimized [SCJ^{+}08]. Minimizing [KER01, LZ05, LO96, FSZ07]. Minimum [CW00, DH94, Li92, RDL95, WW97, BC06, BBL04, tH90, LVP08, LMZ04, OMSGNSG05, SL89, WCWH03, YZLT09]. minimum-spanning-tree [tH90]. Mining [GC01, HK01, KRS01, Zak01, CTT08, GJA08]. mirrored [BL05]. Miss [SDS99]. Misses [DSS95]. mitigating [KMMZ06]. mix [Ahu90]. Mixed [CDY97, MRR^{+}02, NDZA99, SV00, van96, BKS91, FCS91, Kal04]. Mixed-Mode [NDZA99, BKS91, FCS91]. Mixed-Technology [MRR^{+}02]. MKCE [RW01]. MM [Won99]. MMR [CCQ^{+}06]. Mobile [Ano01e, BD00, BN02, BST01, CS00, CCK^{+}08, DKY01, DL01, GS01b, KER01, LAZC00, MS00, Pat01, PRS97, SMR96, WLID02, ZR00, AP03, Ano04d, AK06, BN03, Bou03, CSWD03, CNS03, CW05, CDCD05, DB08, EBE08, FCC07, GRDB05, GZMC08, HKW05, KERUM04, Lan09, LZCY09, LPX05a, LLW07, LHT08, LS06, MS05, MSJ05, RKK06, SY04, SGS08, TZ07, TM06, VA03, XHG03, XG03, YK04, YC04, YCC05, ZMC06, ZHO03, HC09]. Mobile-Process-Based [SMR96]. Mobility [FCF00, GCB^{+}00, FX06, HC09, RKK06, SK05a]. Mode [NDZA99, WSA^{+}94, BKS91, FCS91]. Model [AGW01, AISS97, Ano97k, BPJG92, BA97, CC91, DL98, DG94, DF94, FTL92, Gao93, GS98, GDN^{+}98, HK96, HR92b, HR92a, JRR99, KSP^{+}92, KCV99, MRRV98, MNB95, NDZA99, OKB95, QY94, SANY94, SAC^{+}98, SSK96, WSA^{+}94, YZS96, eW95, BMB^{+}08, Bic90, BG05, CBD^{+}09, CH06a, CDJ^{+}89, CRC^{+}02, DRT07, Kal04, LFH^{+}03, LTKS90, LA06, MM06, NSTN91, NJ91, OO05, RSR04, SSS07, SL90, SK05b, TR89, VHH08, YJB91, ZA91, dR09, GB06]. Model-Based [KSP^{+}92]. Modeling [ATM01, CR91, CCM92, Chi92, CM93, CLRW00, DI91, FMW^{+}94, JZ05, JZK04, KNS91, LP96b, Pat01, QS05, RP98, SCM99, SCK03, SS00, TK07, AP91c, FX06, Joh91, KME09, LF03, SV08]. modelling [HNSA07, KME89, SAOKM03]. Models [AGW98, Ano96l, ABM^{+}92, BDF92, Bir94, BSS99, BHRS95, CDY97, CDF01, GAG^{+}92, MM00, MLC^{+}90, RHH96, SM92a, SSOB02, SM92b, CkLCK04, CkLCK05, CJA09, DHK04, Har91, HK05, MMAL^{+}06, PL03a, PF91, Pop91, SS06, ZZ90, dG91]. Modes [GGW96, SSG93]. Modifications [PM92]. Modified [WS97b, ZLRP91]. modify [CH06a]. Modular [AM95, DD93, FC95, RAS96, CBP02, Dja06]. modularity [GK04]. Module [AM97b, EL91, MC91, ZFL89]. Modules [DP00]. modulo [YLB90]. Moldability [CB02]. Molecular [ES96, NPY^{+}97, SPVvH03, TSA97, FGM^{+}03]. Monitoring [MLC^{+}90, TG97, ZNQ93, CK08, YT05, ZFS07]. monitors [TH08]. Monotone [HJDH01]. Monsoon [HCAA93, NCA93]. Monte [Bro96]. morphological [SSL04]. Moset [MSJ05]. Most [BS97, HHC98, TAS^{+}01]. mother [MC03]. Motion [CP92, RR95b, OPG08]. MP [CDH84]. MPEG [AAL95, CLV95]. MPEG-2 [AAL95]. MPEG-Encoded [CLV95]. MPI [PS01, ATM01, BA06, BDH^{+}97, CEGS07, DPS05, DPSD08, FKLB08, HcF05, KYL05, LC97, WLNL06]. MPI-2 [DPSD08]. MPI-FM [LC97]. MPICH [KTF03]. MPICH-G2 [KTF03]. MPP [DM90a]. MR [MF94]. MR-1 [MF94]. MRI [SHT^{+}08]. MST [Fer95]. Mukesh [Ano96l]. Multi [ACU08, BG86, BA95, MAM05, NBP98, SR88a, Ser97, SM00, WW96, Wil92, YMG01, AGMJ06, BSW07, CCHC09, CLL09, DMCFCM03, FM07, GCS06, JV06, Kep03, KVHS07, LL07, LS03, NC09, QSL^{+}08, QGL^{+}09, RR05, SCB09, ST05, Trä09]. Multi-Agent [Ser97]. multi-bank [QGL^{+}09]. multi-channel [CCHC09, CLL09]. multi-core [SCB09]. multi-criteria [LL07]. Multi-dimensional [NBP98, DMCFCM03, KVHS07]. multi-functional-unit [QSL^{+}08]. Multi-hop [MAM05, YMG01, BSW07, NC09]. Multi-level [ACU08]. Multi-Mesh [SM00]. multi-message [FM07]. Multi-operand [SR88a]. multi-party [GCS06]. multi-path [LS03]. Multi-processor [Wil92, RR05]. Multi-Ring [BA95, BG86]. multi-thread [ST05]. multi-threaded [Kep03]. Multi-Version [WW96]. multi-zone [AGMJ06, JV06]. multi/many [Trä09]. multi/many-core [Trä09]. Multibody [JBL02]. Multicast [AZ01, ABP92, CLZ02, GK98, LEN90, Lan94, LHHB^{+}01, LME95, Mck94, RJMC95, RMC97, SY01, WB01, Yan00, CS08, GZMC08, GS03b, HL07, KDH08, LMZ04, LHT08, MK08a, PY09a, SKMM04, XLG^{+}06, YF07]. Multicasting [BETD94, FF98, Gon98, GS01b, LBT94, WCC02, XCS06]. Multichannel [HP97a, Mck94]. Multicomponent [RW01]. Multicomputer [ASB97, DG94, GBES93, HlLLY95, JR95, LK96, MLW^{+}97, PA01, RU99, XH93, AP91a, CC96, DB86, Li06b, RS90b, Yan04]. Multicomputers [AGF94, CSSY94, CW92, DY99, DFRCU99, GGD93, Lan94, LME95, LEB98, NSLK99, OK01, PHB96, RS92a, RSB96, SP96, SCC92, SB84, Swa98, TJ92, WN94, XH91, XMN92, YB01, GH89a, HSMB91, RS90a]. multicore [ABC^{+}09b, WCO^{+}09]. Multidimensional [GC01, LS94, RS92a, KT91, LB89, PMV05, QSL^{+}08, SC91a]. Multifaceted [Won99]. Multigauge [LR94]. multigrain [ABC^{+}09b]. Multigrid [MT96, MHC95, PSE^{+}01, IHM05]. multihop [CDCD05, HW03]. Multilevel [BW89, KK98a, KK98b, SKK97, MMS09, SZW05, TK08]. MultiMedia [CCQ^{+}06, ALL99, AZ01, GC95, JSCB95, LBL95, Won99, WUG99, ZR00, LVP07, ZV09a]. Multimedia-on-Demand [JSCB95]. Multimessage [Gon98]. Multinode [VB94]. Multipacket [MS94, RR95a]. multipartitioning [DMCFCM03]. Multipath [LYL93, KPR88, SH89, WGS08]. multiperiodic [TW89]. Multiple [ALL99, ADS98, BOSW94, BOS^{+}95, CCC92, DLP99, FGKT97, GH93, KS97a, KC98, KJ84, KM91, LMCF90, LSC00, Par92, SM94, TVS97, VSIR91, VB02, WNA^{+}94, Wan96, ACU08, BXA08, BSMH08, BFKP04, Car90, CHC05, CCLS94, DMB^{+}03, GRV08, JSWB92, JP09, JW89, KAP90, KSS^{+}07, KR87, LLL06, LPX05a, LVB07, MVB05, MHBW86, PTZ06, PHS04, SK09, SZ03, YB90]. multiple-bus [MHBW86, YB90]. Multiple-Pass [Wan96]. Multiple-Writer [KS97a]. Multiplexed [HP00]. Multiplexing [AM95, PD92, QMCL94, QM01, ZLPP01]. Multiplication [Fag92, Li01, NFEG97, CLR90, EL91, ITT04]. multiplier [MS87]. Multipliers [SRK95, BOI91]. Multipole [SHT^{+}95, YB01, KP05]. Multipole-Based [YB01]. multiprecision [MS87]. multiprefix [Coh90]. Multiprocessing [CDH84, MBK^{+}92, ABC^{+}88, JS86]. Multiprocessor [BW95b, CKL99, CP91, DS96, DRC90, DFN^{+}94, GH90, GMM00, HP00, HC95, HN91, KS97b, LYC02, LF92, Lun94, MF94, MMRS98, MT95, MMVR97, MD92, OM90, PL95, PM96, PP92, QY94, RS92b, SEP96, Soh96, WF93, XZS96, ZNQ93, AOSM05, BHR91, BR91a, BS92, DI91, GL89, HDT^{+}05, HA91, HC91, KA05, Lee90, LHK03, LW89, LVB07, McA89, PK05a, PI90, SK09, SM89a, SYYU07, TS91, YL89, ZZ90]. Multiprocessors [AMB95, AM95, BJ99, Bas97, BS96a, BL96, BC01, BLG01, CB95, DS95a, DJ98, DZDZ01, DT92, GY92, GZ97, HJ01, HA92, KSB94, KB96b, KA97, LK98, LA93, MB92, MS98, MG91, NB93, NS97, NPP^{+}02, PH91, PY96, PT97, RL96, RJY96, SMH94, SCM99, SY01, SDS99, SD00, SC91b, TTG95, VSIR91, YW91, YMR93, YL98, AP91b, BC05, CLM90, Cyb89, FZC^{+}05, FGP05, Gai90, GL90, KA03, LEN90, LE91, NSTN91, Nik03, RFPAG08, SPBR91, SD91, SMH91, SA90, YB90]. Multiprogrammed [MS98, NSS97, NPP^{+}02, YL98]. multiprogramming [DI91]. Multirate [HJDH01]. Multireader [HV95]. Multiresolution [KZ96, ZM94a, CL85]. Multiscalar [VS99]. Multisearch [ADM^{+}94]. Multiset [AFS96]. Multistage [AA95, BETD94, LC96, OM84, PL93, SZB92, TH02, Tze91, UR94, Wan96, Wan01b, YWP00, ATH91, CM04, FZ90, HJ90b, Har91, JBM91, LK90, MVM04, SH89]. Multistage-Network [UR94]. multistep [GGR89]. Multithreaded [BJK^{+}96, BLG01, GGB93, GRS97, KC99a, Lun99, PS01, RNSB96, RSBN01, SAC^{+}98, SSYG97, TG99, YMR93, ABC^{+}09a, SLG06, TKHG04]. Multithreading [BL96, FKT96, KPC96]. multitonic [Sei05]. Multiuser [BAL05, ZRC99]. Multivalued [HV95, HV09]. Multivariate [HK01, MMAL^{+}06]. multiversioned [Ahu90]. Multiway [SM00]. Munin [Car95]. Muntz [Ano92a]. MUPPET [MSS88]. Mutual [AE95, Cha94, Cha96, DGFGK05, FTC00, GBG93, KY02, KUFM02, NTA96, NM02, Sin93, XLG^{+}06, YZY96, AK07, BAS06, CW05, CH06a, CB06, Gos90, MM07c, Ram89]. MVAMIN [JBM91]. Myrinet [KL01b, QS05].

name [TB90]. NAP [KF90b]. NAS [JV06, WAS95]. Natural [LS95, VB96]. NC [LO91, RDL95]. Near [FTL92, HA92, San99, UR94, CCN06]. Near-Maximum [FTL92]. Near-Neighbor [HA92]. Near-Optimal [San99]. Nearest [HH01, OS96b, KS08, NA06, Wan07]. Nearest-Neighbor [OS96b]. Nearly [Nas94, SSM89]. Necessary [SJ96]. Necessity [MC03]. Negotiation [LL98]. Neighbor [HA92, OS96b, UR94, KS08, MKC^{+}09, Wan07]. Neighborhood [LYC02]. neighbors [NA06]. Nested [BHS^{+}94, CWW96, DRR96, HS93, KBG92, Mer96, RSS99, SCB09, AGMJ06, BFTV87, EB09]. Nests [DR95]. Net [BPJG92, BDF92, Chi92, Fer92, SP90, NM95, WL92]. Netfinity [BAHP01].

Nets [BPJG92, CMT92]. Network [AA93, AAD98, ABM^{+}92, ABCP96, BBH^{+}97, BBCD02, BA95, BC01, BF97, BST01, CGKK97, CW01, Cha95, CW92, DLLX97, DSAUM99, DVZ96, DBP94, DKMV01, DH95, ESMG96, FFK97, FAM96, FTL92, GRS97, GS01a, GH93, HH97, HPT^{+}97, KC95, Kop97, LS97, LK94, LC96, MM00, MJ94, MSS88, NBSD99, OM84, PN97a, PN97b, Pat01, RCY97, RJY96, SM00, SBAM96, SS95, TSC01, Tze91, UR94, WMG01, YZY96, ZLP97, ZMPE00, ZW00, dBL95, AP91b, Ano04d, AF06, BXA08, Bat05, BAL05, BPA06, CKO04, CKN07, CDB04, CWL^{+}07, Che89, CVJ09, DE91, FK89, Gai87, GZMC08, HWWH08, HWC08, IS06, JXW06, Joh89, JZK04, KERUM04, KJD03, KCD08, KO90, KPR88, LAD^{+}96, LTD^{+}93, LY08, Nap90, NS90, OO05, PL06, RH05, RD05, SSB91, SS05, SY04, SK89a, SMKL93, TM06]. network [VM95, VHH08, VR86, YK04, ZWS09, ZWRI07, dG91]. Network-Based [GS01a, OM84, PN97a, PN97b, CVJ09, KJD03]. Networked [FGKT97, HS97, LHM95, OEY07, BW09, HP06, SS08]. Networking [Ano01e, GCY^{+}04, Bou03]. Networks [AAD02, AZ01, AS97, ABP92, Ann94, Ano92c, Ano93e, Ano00d, AA95, BSS97, BAES92, BCH95a, BETD94, BCD00, BDF01, BCH95b, CP97, CT96, CS00, CAB94, CS93b, CC94, CS95c, DS95b, DHB02, DP99, DS93, DL01, DF95, DZ97, DC94, FCF00, FT94, GGN93, GK98, GHKS98, GO95, GPS96, GB93, GS01b, HIKM94, yHY97, HLCZ00, HJDH01, HJD^{+}01, JR92, JH92a, JLRA97, JH94, KKGS01, KL01a, KRSZ02, KAM94, KB96a, KL01b, KR98, KJ84, Lat95, LBL95, LYL93, Lee94, LLJ00a, LAZC00, LPS^{+}98, LWOG02, LHHB^{+}01, LP95, MS00, Man94, MLW^{+}97, MSH90, MS85, Mck94, MDD97, NRS95, NSSS99, NS92, OD95a, Ola01, OO85, Oru87, Oru94, OK01, PRW94, PA97, PA01, PL93, Piu01, PKD97, Pra93, QMCL94, Qia97, QM01, RS96b, RP98, RMC97, Ros99, RLS96, SW96, Sei05]. Networks [SZB92, SLP^{+}98, SZ00b, SF90, SCD99, Szy95, TVO92, TH02, VB02, WM92, Wan96, WR97, Wan01b, WB01, WP02, WAS95, Wil92, WT92, YWP00, Yan00, YN92, YMG01, YP96, ZZC92, AP91a, ASM09, AAD03, AB05, AP03, Arb89, ALF03, BFG^{+}03, BCV05, BSW07, BGLA03, BS03, BDjQ86, BHR91, BR91a, BPRS04, BOP06, Bhu87, Bod89, BR91b, BN03, BZLI04, BMIM07, CI03, CM04, CNS03, CKN07, CW05, CS06b, CCK^{+}08, CS92, CDR09a, CDR09b, CYZ06, CGG^{+}09, CDCD05, CM93, CMS04, CT04, DW06, DD96, DMB^{+}03, DB08, Dim04, DFP06b, DH04, EAL90, EBE08, Fei03, FY86, FZ90, FJG06, FKJG08, FMM^{+}08, FVCL05, FD86, GPT06a, GRV08, GDP08, GP07, GCY^{+}04, GSSS03, GH89a, GAGPK03, GJXZ05, GS03b, GMXA07, HW03, HMV07, HJ07, HJ90b, Har91, HS06, HZY04, HT06]. networks [HDT^{+}05, Hoh90, HL07, ISAZ07, IB04, JT88, JBM91, JBY^{+}05, KKVI05, KSI04, KK06, KOA09, KGN89, KMF^{+}05, KKS09, KMS07, KDH08, KH89, KNS06, LZ08, LK90, LR06, LHKL03, LR03a, LCW05, LPX05a, LW06a, LT07, LS03, LC07, LR03b, LLW07, LHT08, LHP07, Los08, MLG05, MM04, MAM05, MSM09, MV88, MSZ05, MS88, MVB05, MBR08, MKC^{+}09, MAJJ05, MVM04, MHBW86, MK08b, NJ91, NC09, OMSGNSG05, Pak89, Par05, PK05a, PL06, Pel90, PV07, Pla08, PLR07, PB09, RKK06, RHL08, SAOKZ05a, SAOKZ05b, SX08, SZ09, SGS08, SKMM04, SK05a, SL89, SR88b, SR90, SK05b, SH89, TBHA07, TDC05, TDM05, TR08, VO89, Var91, VA03, WCC02, WW07, WG08, WTB^{+}08, WGS08, WMW09, WBTM09, Wil90, Wu85, WTS03, WH08, XYKA08, XCLR07, XHG03, XQ04, XWC^{+}08, XG03, YME06, YF09, YL89]. networks [YSL08, ZV06, ZMC06, ZXP09, ZDC06, ZHO03, ZC04, MR03, PRP09]. Neural [AA93, Ano92c, BST01, CW92, FTL92, HPT^{+}97, JH92a, KJD03, Kri92, LWOG02, MM00, Mon94, NS92, Piu01, Ram92, TVO92, WT92, ZZC92, eW95, Arb89, FK89, GH89a, Joh89, KH89, SMKL93, Tor89, VM95]. Neural-Network [CW92]. Neuro [MT97b]. Neuro-Chip [MT97b]. Neurocomputer [GFB^{+}92, Ram92]. Neurocomputing [Ebe94]. neuronal [VO89]. next [RKK06, SB04]. next-generation [RKK06]. Nexus [FKT96]. NIC [JBY^{+}05]. NMC [SANY94]. No [KF90b]. Node [AAD03, KKS09, FKLB08, TR08]. Node-ranking [AAD03]. Nodes [GP97, NSLK99, SS95, CK91, DB86, WGS08, XYG07]. Non [BH05, BGH^{+}03, CLL09, GRDB05, Lai86, Li06a, MM07c, MAR05, QS05, WLNL06]. non-blocking [QS05]. non-clairvoyant [Li06a]. non-cooperative [GRDB05]. non-dedicated [MAR05, WLNL06]. Non-evolutionary [BH05]. non-first-in-first-out [Lai86]. non-linear [BGH^{+}03]. non-uniform [CLL09, MM07c]. Nonatomic [Sin95]. Nonblocking [JSM94, MS98]. Noncooperative [GC05]. Nondedicated [Ano97k, YZS96]. nondense [WF90]. Nondeterministic [CY95]. nonequivalent [NJ91]. Nonexpansive [Bah00]. Nonlinear [AM93, DM90a, ESMG96, MHC95, CEGS07, KKB^{+}06]. Nonloop [Bec96]. Nonoblivious [FY96]. Nonredundant [Wu94]. Nonuniform [AA95, KRW96, KR97, LK90, OP98, WLR90]. normal [ZB03]. Note [Ano01-34, Ano02j, Pel95, Num07, Ano04d]. Notes [THSS87]. Nothing [LT94, PVGG06]. Notification [ABP92]. Noting [HTL99]. notion [LJ86]. Novel [LYC02, LLCL98, OS96a, CWLD05, CCHC09, LTB02, SKMM04, YF09, ZV09a]. NP [BRR01, MPZ09]. NP-Hard [BRR01]. NT [BAHP01]. Null [DMSH90, BD04]. NUMA [LE91, WF93]. Number [Alu97, Ano92a, Ano92c, Ano93e, Ano96l, Ano97k, Ano00d, Bro96, BS96c, CS93b, SS95, ZAW94, DDNS06, FSZ07, HSSM07, IC05, PK89]. Numbers [NS94]. Numerical [BK95, LLCC02, RW01, NAK04]. NUTS [LK90]. NVIDIA [KME09].

O [AW95, Cho93, CQ95, CD95, DD93, DT01, DJT03, GGD93, JSCB95, JSWB92, LTH97, MLG05, NSSS99, NsPPC02, WLWW09]. O-Intensive [EH01a, CkLCK04, CkLCK05]. Object [CSSY94, CS95b, DR98, GCB^{+}00, HS00, JRR99, KC99a, LLS93, LTH97, SG96, WPKK94, WLID02, WH97, ACFK07, Chi95, KC04, LFH^{+}03, SK90, SCK03, YJB91, ZV09a]. Object-Based [DR98, WLID02, ZV09a]. Object-Oriented [CSSY94, CS95b, HS00, SG96, Chi95, YJB91]. object-space-parallel [ACFK07]. Objects [CLZ00, CDP95, HPT02, Kap93, SBAM96, VWHL96, WG93, Won99, van96]. Oblivious [IM00, YME06]. observations [ZHO03]. Obtaining [AFT^{+}00]. Occam [LC92]. Occamflow [GL89]. Ocean [SAC^{+}98, SH92b]. Octree [FLS^{+}97]. Octrees [BFG94]. Odd [DS96, NT93, SL95, ZDC06]. Odd-Even [NT93]. ODEs [Wor93]. ODMRP [OPG08]. Off [BCLR96, GK98, LPU97, BS92, PF08, ZB09]. off-line [BS92]. off-the-shelf [PF08, ZB09]. offer [Trä09]. offloading [WL04]. offs [CLR90]. Olden [CR96]. Omega [Ano93e, CS93b, SZ00b, GL90, CS92]. omega-like [GL90]. on-demand [BSW07, FVLB09, SFEF06, WL05, XG03]. On-Line [BDKM94, LTY96, Yen01, DJ98, EL88, LHK03, KM88, SL90]. One [Ano93e, CS93b, LP95, PTA08, SR97a, SR97b, YAS98, ZB97, Che05, CS92, Deh90, Yan04]. one- [Deh90]. One-Copy [Ano93e, CS93b, CS92]. One-Dimensional [LP95, PTA08]. One-Sided [ZB97]. one-step [Yan04]. one-to-all [Che05]. One-to-Many [SR97b]. One-to-One [SR97a]. Online [HCWS94, QM01, BFG04, Li06a]. Only [GS00]. onto [BR08, BS90, BSB^{+}01, DAYA02, Dja04, DQR^{+}09, ERL90, ERS90, GH89a, GW99, KMS^{+}06, LLS07, MM00, MAS^{+}99, XH91]. Ontology [PRP09]. Ontology-based [PRP09]. Open [CA94]. OpenMP [AGMJ06, CCM^{+}06, HLCZ00, LA06]. operand [SR88a]. Operating [MBL^{+}92, SEP96, CDJ^{+}89]. Operation [HLJ01, Coh90, KNS91]. Operational [RHH96]. Operations [BTZ98, DP98, FAGW95, HTL99, HLJ98, KSA95, PKD97, Van94, ZK94, BM04b, LMR05, JSWB92]. operator [CL85, TG03]. Operators [BDKM94, SR94]. opportunistic [AM07]. opportunity [KS03]. OPS5 [GF89, HS86]. Optical [AK93, Ano93e, BA97, BC01, CS93b, CLM90, DP99, DSD^{+}97, ELS94, ES97, GB93, HP97a, HQPT99, IWM97, LLJ00a, LLJ00b, LPZ99, MR03, MC93, MB93, MG93, OS97, OS93, PEC95, QM01, RP98, SHC93, SL97, Szy95, SH98, THN^{+}93, TBPV00, WLY01, WHT00, YWP00, YMG01, ZMPE00, ZLPP01, CS92, McA89, NAK04, WG08, dR09]. Optically [DH95, EH01b, Guo94, KM97, MKY^{+}97, QMCL94, GMH^{+}91, TRSS06]. Optimal [AMS94, AR97, AKPT99, BNS00, BBM^{+}02, BSDE96, BOS^{+}91, BOSW94, BHK^{+}94, CW00, CS93a, CA95a, CW92, CA96, DS95b, DP00, DLP99, DT97, DF90, Ede91, FLPJ07, FM96, FXW03, FA95, FAM96, FY96, GS91a, HV95, HKMU98, HM01, Ho91, HJD^{+}01, IZ95, JP95, JBP00, KERUM04, KUFM02, KS97b, KW02, LHS97, LSC00, LK94, LCW05, LO94, LO96, LV88, LS01, MS94, Man97, MW95, Nak95, OS96b, OSZ98, OH02, PM05, PP06, PK05a, Pel95, PL94, PV07, PM96, RR95b, San99, San02, SJ95, SZ00b, Sin87, SV00, TR08, WL90, WLY01, WR97, WS95, WS97a, WN94, Wu94, WHT02, Wu03, WLL08, ZWS09, ZWRI07, oPP00, ANP07, BM04a, BS92, CV90, CMS04, CZ90, Dja04, Gue86, HDJ08, LH04, LS05, Lis90, MD07, NW88, PY09c, Pel90, PA04, PLR07, SGR03]. optimal [SSM89, SGE91, WC91, XWC^{+}08]. optimality [HV09]. Optimally [TBPV00, GC07]. optimisation [LL07]. Optimistic [HF02, NH93, PW96, SS93, DWG03, JLM08, QS05]. Optimization [BLG01, CLRW00, FM99a, FCF00, HA92, KCRB99, KZ96, KLS90, LWY97, OK02, RL02, RNSB96, SMH94, TRSS06, VSM96, WCO^{+}09, ATH91, AF06, BCM87, CI86, KA89, KKB^{+}06, KLL87, QS05, RRS^{+}08, SCC^{+}06, SZD07, SK90, WMW09, Wol88, YWD08, ZI08]. Optimizations [BW95a, DUSH94, HKT94, KY96, RSB96, ZH99, ABC^{+}09a]. Optimize [DRR96, HLJ01, SF05]. Optimized [ABDS02, Bar05, BKS91]. Optimizer [HlLLY95]. Optimizer-Assisted [HlLLY95]. Optimizing [CG86, KRC00, KR06, LMR05, NCTT09, Sab94, WMG01, WLWW09, ARM^{+}05, MSM09]. Optimum [LP96a]. Opto [AA93]. Opto-electronic [AA93]. Optoelectronic [HPT^{+}97, MLW^{+}97, MB93, HNSA07]. Order [AMS94, Bit92, CLZ02, DT97, BCM06, BG05, GA90, KMF^{+}05, KME09, MP87]. Ordered [GS98, TS91, JW89, Tay05]. Ordering [KK98b, PRS97, RS96a, ZB97, CHC05]. Orders [SH97, Sta04]. ordinary [GGR89]. Organization [AP94, CT04, Ull84]. organizations [BW89]. organizing [BFKP04, BZH06]. orientations [AFM09]. Oriented [BS90, CSSY94, CS95b, Fer92, HS00, SG96, Bic90, BZLI04, Chi95, CTT08, KBD05, RKK06, SK90, SFEF06, YJB91, ZC04]. Origin2000 [SSOB02]. ORION [PRP09]. Orthogonal [AR97, Wu02, GS91b, HC91, SM89a]. orthogonal-access [HC91]. Orthogonally [CP98]. Other [Kap93]. OTIS [ZMPE00, ZXP09]. Out-of-Core [BCR96, Raj04, KKB^{+}06, WJV07]. Outerplanar [GS99, KW02]. Output [ASR93, GC07, PD92, Ros99, ST02, GS03a, PY09a, ST06]. Output-sensitive [GC07, GS03a]. Overall [LO96, SEP96]. overcome [KG04]. overflow [SCC^{+}06]. Overhead [DR98, JNW96, KS00, SD00, BCM87, BD04, CX05, FGP05, SC91a, SZ09]. overheads [DI91]. Overlap [QH96]. Overlapped [Lin93a, KNS91]. Overlapping [CQ95, Wil92, CHC05, KSG03]. Overlay [PRP09, GZMC08, HK04, LCM^{+}06, XLG^{+}06, YF07]. overlays [ZH07]. overloading [AOSM04]. Overview [EMP^{+}96, KS93, ABC^{+}88].

P [ASST05, dR09, PMV06]. P2MCMD [LC07]. P2P [CWLD05, GJXZ05, RHL08, She09, SZ09, SHLN09]. P2P-based [She09]. PACK [BR96]. PACK/UNPACK [BR96]. Package [HS97, KOW97, XKMN94, CPO^{+}03]. packages [PL03b]. Packet [GHKS98, GO95, JK00, LYL93, LS94, NS95, OY00, PRW94, PV89, RD05, SL97, BMIM07, KMF^{+}05, Nap90, OS04, PY09a, YSL08]. packet-level [YSL08]. packet-size [OS04]. packet-switched [Nap90]. Packets [GRV97]. Packing [Hwa97, LTW^{+}90, SF05]. Page [LE91, NPP^{+}02, HSSM07, TH08]. Pagenumber [KRSZ02]. pages [Ano96l, Ano97k, Ano00d, CS93b]. Paging [DM99]. Pair [DP98]. Pairs [BGR96, TU92, KS91]. Pairwise [GP00, CK08]. Pancake [BS03, KAM94]. panel [Rob09]. Paper [Ano01m, Ros07]. Papers [Ano95i, Ano95j, Ano96j, Ano96i, Ano97i, Ano97j, Ano98k, Ano98i, Ano98j, Ano99g, Ano99d, Ano99e, Ano99f, Ano00a, Ano00e, Ano00f, Ano00g, Ano00h, Ano01c, Ano01d, Ano01e, Ano01n, Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano01-29, Ano01-30, Ano01-31, Ano01-32, Ano02q, Ano02r, Ano02s, Ano02t, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Sni03, Rob09]. Para [CD98]. Paradigm [KBD05, RS92d, BAMM05, CVJ09, MSJ05]. Paradigm-oriented [KBD05]. Paradigms [Ano99g, CEF^{+}95, YMR93, XQ04]. Paragon [CCRS92]. Parallel [ASR93, AGW01, AT94, AGF94, AAL95, ANT02, AISS97, AP94, Als01, ABJS01, Alu97, AFM03, AS97, AS95, AH94, Ano92a, Ano93a, Ano96j, Ano97j, Ano97k, Ano99g, Ano00d, Ano02v, ABZ95, AKP95, ADM^{+}94, AS94, ADS98, AB93, BK95, BJ96, BR96, BCD95, BBD^{+}91, Bai94, BW08, BBH^{+}97, Bal90, BDF92, BGR96, BS97, BCV94, BFG94, BN94, BB93, BBM^{+}02, BL94, Bev02, BBH^{+}98, BP95, BEE00, BS90, BHS^{+}94, BDHF90, BP89, BR95c, BRPR06, BMARM07, BMRC98, BMRC99, BS00, BTZ98, Bro96, BX93, BDH^{+}97, BA01b, BTG02, BMCP98, BM95, BNSP99, BS09, CP97, CMT93, CP98, CGKK97, Cas93, CC91, CDY97, CDRC99, CB99, CKK00, CvdBL^{+}08, CCRS92, CGL^{+}95, CCC90, CS95b, CW93, CA95a, CWW^{+}95, Chi92, CV91, CDJL09, CN93, CP92, Cho93, CHR94]. Parallel [CY96, CWP98, CB96, CQ95, CGA98, CH92, CP94, CA95b, CHGM01, CRFS94, CLZ00, CBdCD00, DM90a, DM95, DOP98, DP00, DM92, DRC90, DH91a, DS84, DO89, DH94, DN94, DJM94, DSW94, DT01, DSD^{+}97, DBKF90, DD95, DZ97, DJT03, ES96, ERL90, ERA95, EMMM94, ELS94, ES97, EHS94, EHMN95, Fah96, FBRW03, Fer95, FR96b, Fer92, FMP98, FLS^{+}97, FC95, FKKC97, FJ93, FMW^{+}94, Fre96, FT94, GG94, GP94, GCB^{+}00, GGN93, GV94, Ger98, GBES93, GGD93, GJP96, GC01, GSC96, GM95, GSP02, Gra09, GL92, GH89b, GHH92, GWH06, GK93, GHSJ96, GS99, GRR^{+}05, Hag97, HHM94, HK96, HH97, HGCC96, Han89, HB97, HBJ98, HP95, HR92b, HR92a, HHC98, HP97b, HN91, HTB98, HR89, IK94, IZ95, IWM97, IHM05, JW94, JBL02, JSM94, Jia99]. Parallel [KR97, KF95a, KME92, Kap93, KSA95, Kar92, KK98b, Kau94, KZ96, KR98, KB01, KKS08, KE93, KS93, Kri92, KW02, KG94, KGV94, KM92, KA97, KC99b, LSCA93, Lan09, LP96a, LMCF90, LWY97, LTH97, LJKS02, LS97, LC90b, LAS^{+}97, LPZ99, Li01, LWOG02, LYL08, LSH96, LS88, Lin91, Lin93b, LA93, LO94, LLCC02, LP97, LFA96, MB96a, MHF93, Mah95, MM93, MS99a, MLC^{+}90, MR94a, MPZ09, MT96, MB96b, MP93, MSH90, MD98, MHC95, MB92, MSd^{+}95, MMAL^{+}06, Mer96, Mil93, Mir91, MB93, MG98, Moh96, MS96, MS99b, NSS97, Nas94, NFEG97, NMS93, NS97, Ngo06, NT90, NKC^{+}97, NH93, Nic94, Nie94, Nik04, NsPPC02, NDZA99, NS92, NPY^{+}97, OO05, OY00, OB98, OP98, ORR03, OR97, PH91, PD05, PP96, PH00, Par98, PE93, Par96, PL03a, PL94, Pla08, PAH^{+}98]. Parallel [PAJC97, PSE^{+}01, QZ94, QH96, QOvdG01, Raj01, RSS96, Ram92, RL02, RS92b, Ree84, RW01, RGS00, RPS93, RSW90, RIZ90, RJA97, Ros99, Ros07, RW93, SSG93, SH90, SS96, San98, SMR96, San02, SAOKMA02, SH97, SG93, Sch90, SM89b, SW96, Sch91, SdS97, SAF05, SR97a, SR97b, SAC^{+}98, She06, SS92, SSHC00, STN92, Shu95, SGS99, Sil90, SM00, SRK95, SSRV94, SB93, SC95, Ski96, Sni03, Soh96, SL97, SIR92, SK93, SMKL93, Ste95, SSK96, SWC^{+}91, SF90, SYG92, SS97, Szy95, TSA97, TW87, Ten90, TAS^{+}01, TR96, THBF97, TVO92, TZ00, TK08, TF01, UAPM07, VSM96, VGAB08, WB94, WCE97, WLY01, WM92, WNA^{+}94, WPKK94, WB96, WTC08a, WMW09, WSA^{+}94, WD94, Wee01, Wei98, WMG01, Wei02, WA02, WAS95, WS95, WS97a, Wor93]. Parallel [Wri91, WT92, WH97, WHT00, WHT02, YZS96, YB95, YIY97, YB01, YP96, Zak01, Zep91, ZYH94, ZK94, ZB97, Zhu92, ZH99, ZM94a, ZO97, ZYO02, ZA91, ACYS08, ALS91, AP91c, ATH91, Ara90, AE88, ANP07, AG86, ACFK07, Bad04, BC05, BCM87, BB87, BBCLL04, BBM08, BA06, BČFF05, BAH04, BFH09, BS87, BSG90, BR91b, BGM^{+}08, Boż09, CK88, CR91, CCS06, CRL04, CEGS07, Che86, CC87, CJ07, CT94, CDJ^{+}89, CL85, CZ90, CB06, CD95, CK91, DFP06a, DRT07, DM90b, DM90c, DQR^{+}09, DUW86, DM94, DWHL87, Ebn04, ESTA94, EE05, Eİ07, FCG04, FGG08, FCS91, FSD04, GGR89, GS91a, GP91, GT04, GWWL94, GS03a, GC07, GB06, HM06, HOE^{+}09, HS86, HA91, Hsi04, JT88, JSWB92, JMS86, JL05, JP09, JZ05, JV06, KESA07, KV88, Kep03, KHK03, KM03, Koc91]. parallel [KK86, KS91, KMP^{+}06, KP05, LTB02, Li06a, Li06b, LT07, LTKS90, LC92, LH04, LS05, LH09, LF03, Luk85, ME04, Mar88, MV88, McD89, MCT06, MP87, MAR05, NW88, Nic07, NCTT09, OS04, PB90, PPC04, PK89, PF91, PVPM06, PHS04, Pop91, PF04, PRG88, QJ05, Raj08, RSR04, RGD03, RG87, Ros89, RSW91, RTCG91, SI86, SS03, SPBR91, SV08, SI89, SC91a, SS06, SSTP09, SC04, SZW05, SF05, SK91, SA08, Sta04, SDG08, SR91, TM06, TW89, TRSS06, TS91, Trä09, VD04, VA07, Vis87, WL90, WC91, WJV07, WBTM09, WRHR91, WJD91, WZ91, WF89, WLWW09, WGCZ09, YJB91, YZLT09, ZFL89, ZJ06, ZFWF06, dVCP06, dGP06, CPO^{+}03, Ree84]. Parallel-depth [BP89]. parallel-processing [Trä09]. Parallel/Distributed [KZ96]. Parallelisation [HSSM07, Kal04]. Parallelism [Bec96, BAM93, DRST02, FM85, FKKC97, FY97, GSG^{+}93, HKT^{+}91, KRC00, MR94b, MK92, SSG93, SW91, SH92b, SV00, SG96, GV86, HS03, Irw88, PVGG06, RS08, SCB09, DeG88]. Parallelization [BPST96, BF01, DHR96, HO94, KR97, NM95, NC97, Pov99, SANY94, UZSS96, WCKD06, AAD05, AGMJ06, CVJ09, IBP08, SGE91]. parallelizations [CCLS94]. Parallelized [DR98, MJ01, SPVvH03]. Parallelizing [HWW96, RHH96, Tse90, DMCFCM03]. Parameter [FCF00, SPVvH03]. Parameterized [dR09, NSTN91, PW96]. Parameterizing [TSHH01]. Parameters [Fer90]. Parametric [DR95]. Parentheses [MW95]. Pareto [BFM06]. Parity [CT93, MRK93]. ParList [FMP98]. Parser [CB96]. PARSIMONY [GC01]. Part [SAOKZ05a, SAOKZ05b]. Partial [FY96, HHC98, SH97, VB94, JL05, Ros89, Vel89]. Partially [FII04, SKK91, Tay05]. Particle [BTG02, PAH^{+}98, SDG08, CvdBL^{+}08, LTKS90]. particle-in-cell [LTKS90]. Particle-To-Grid [SDG08]. Particles [LLCC02]. Particles-Turbulence [LLCC02]. Partite [EMMM94, SLP^{+}98]. partition [LM05].

Partitionability [SZ00b]. Partitionable [NMS93, SB84, CL91b, LC90a, LC91b]. Partitioned [CB99, LJKS02, YI96, CG86, Gai90, Mat06, OT86, SR88a, SM08a, MR03]. Partitioner [SSB98]. Partitioning [Als01, AYIE98, BW96, Bou02, CN93, GK98, HS93, Kar95, KK98a, KK98b, Lee90, Mah95, Moh96, MFS96, Nic94, PHB96, PB99, TG99, WCE97, WF93, ACU08, CP05, DHK04, LVP07, Mit07, PA04, PTA08, SW91, TK08]. Partitions [SS96, MMS09]. partner [FCC07]. party [GCS06]. PARULEL [SWC^{+}91]. Pascal [PLD87, Ree84]. Pascal-based [PLD87]. Pass [Wan96, DD96]. passable [VR86]. Passing [BB93, BDH^{+}97, CW92, CD98, dADB96, GBES93, HNM02, Isl97, Kar92, KTF03, LK96, MD92, PY96, PS01, SCMB90, XH93, ZN01, BPW05, GH89a, Hal05, LR06, She06, vS91]. Passive [MR03, DS04b, YT05]. Past [TAS^{+}01]. patch [Meg91]. Patches [GM95]. Path [BLG01, DP00, FF98, HTB98, IZ95, LK96, NTA96, OC07, RMC97, TU92, TZ00, ATH91, ANP07, DM90b, EDÖ05, Hsi04, KS91, LS03, NS90, Ros89, SYYU07, WCC02, YME06]. Path-Based [FF98, RMC97]. Paths [BGR96, BP02, GT97, GP00, DMB^{+}03, FLPJ07, PK04b]. Pattern [AA93, BMRC99, LW95, Lon04]. Patterns [GSP02, KS02, LL95, BHR91, BR91a, NAK04, RGU08, SPBR91]. PCB [wXH00]. PCG [ORR03]. PCS [FCF00]. PCT [AT03]. PdBCube [CAB94]. PDE [CHR94, GV86]. PDES [PW96]. PEACE [BNSP99]. PEC [LP95, RS96b]. peer [ALH^{+}09, ABCM07, BCK^{+}09, BAL05, FJG06, FKJG08, FVCL05, HK04, LC07, MSZ05, OSL05, WXZ05, WGC09, WDDK09, YF09]. peer-to-peer [ALH^{+}09, ABCM07, BCK^{+}09, BAL05, FJG06, FKJG08, FVCL05, HK04, LC07, MSZ05, OSL05, WXZ05, WGC09, WDDK09, YF09]. Penalties [SDS99]. Perceptron [ZAW94]. Perceptual [CWP98]. Percolation [MSH90]. Perfect [BAES92, AB05]. Perfectly [Lin93a]. perform [EL91]. Performance [AP91a, Abr96, ABDS02, AP93, ACD^{+}93, ATM01, AYIE98, AH94, Ano92a, Ano97k, AA95, BJ99, BBH^{+}97, BPJG92, BCV94, BS96a, BAMM05, BL96, BCD00, BP01, BLG01, BNSP99, CTD99, yCM98, CY99, CB02, CP99, DS95a, Dah99, DPSD08, DY99, DS02, DW04, DF94, ER97, FR92, Fer92, FGKT97, FPD93, GCB^{+}00, GE85, GT02, GM94a, GGD93, GDN^{+}98, GM99, GRR93, GBA08, GK93, GK04, HMBW07, HCS^{+}00, HCAA93, HSMB91, HP97b, HN91, HLL^{+}95, yHY97, HTL99, IC05, JSCB95, JV06, JB93, JLRA97, Joh91, KME92, KMKD97, KC95, KS95, KMS07, KB96b, KG04, KEA95, KJ84, KRS01, KLL87, KMB91, LC97, LLS93, LYL93, LP96b, LPU97, LPX05a, LTD^{+}93, LHVW95, LDCZ97, Lun94, MF94, MT95, MSAF04, MM06, MSC96, MB92, MS96, NBP98, NCA93, NKC^{+}97, OD95b, PH00, PS93]. Performance [PD92, PEC95, PTC^{+}93, PAJC97, PS01, RPS93, RW93, RGU08, SMH94, SSG93, SPBR91, SV08, SKR93, SG93, SB02, SLP^{+}98, SKH96, THBF97, TTG95, TH02, Tze91, VSM96, VHH08, WAS95, WF89, WLID02, XQ07, XZS96, YB90, Yan93, YZS96, YI96, YAS98, Yan00, YB95, YMG01, ZNQ93, AB03b, AP91c, BL05, BW89, Bat05, BČFF05, BKS91, BH86, BJS03, BDDL09, CK06, CF88, CBP02, CBM^{+}08, CCEB03, CkLCK04, CkLCK05, CC96, DK08, FGP05, FJSW90, FD86, GRV08, GST09, HW03, HNSA07, HCZ04, HA91, HcF05, HC91, JBY^{+}05, KKS08, LL90, LWR^{+}03, Li06b, LVB07, MVB05, MG09, MBH^{+}08, NSTN91, Nap90, NTC03, NRM^{+}09, OSL05, Par05, PRHB06, RH05, RM90, RTCG91, SAOKZ05a, SAOKZ05b, SCB08, SD91, SC04, SAB^{+}92, TB90, TMM06, TD07, WTB^{+}08, WS06, WH08]. Performance-Driven [CP99]. performance/cost [AP91c]. Performances [MS99a]. performing [GA90, VM95]. Perimeter [KF95a, KOA09]. Periodic [Abr96, BNP98, BBM^{+}02, RDS02, WCF94, FXW03]. Peripheral [MBK^{+}92]. Permutation [AKP95, CL93, DT97, GT97, IZ95, Oru87, Oru94, QM01, RDL95, TBPV00, WS97a, YWP00, HRJ94, JL05, KO90]. Permutations [AMS94, BP98, CS93c, JH92b, Kap93, RS94, MR03, VR86]. Permuting [Cor93]. persistent [TC03]. Personal [ZR00]. Personalized [LHS97, RWK95, Ede91]. perspective [Los08, Wan07]. perspectives [WH08]. pessimistic [Yan04]. Petersen [SGR03]. Petri [BPJG92, BDF92, Chi92, Fer92, NM95, SP90]. Petri-Net [NM95]. Pfair [HA06]. Pfair-scheduled [HA06]. Phase [AT94, DRR96, LC91a]. Phase-Reconfigurable [AT94]. phases [BKS91, SZD07]. PHOEBUS [MB93]. Phone [BN02, BST01]. Photonic [Qia97, RKK97]. phylogenetic [FBRW03]. phylogenetics [SPVvH03]. phylogeny [PTZ06]. Physical [WH97, BPA06, FD86, KNS06]. physics [GTN^{+}06]. PIC [SDG08]. Picture [HHM94]. PID [WLID02]. pin [AP91a]. pin-out [AP91a]. Ping [LF92]. Ping-Pong [LF92]. PIOUS [MS96]. Pipe [KSL85, SD88b]. Pipeline [DT97, DF94, VSM96, BR08, JS86, ZWRI07]. Pipelined [GMH^{+}91, KSL85, KL84, LPZ99, MP93, PH91, Pov99, RFM94, RS92b, SG99, SV00, TG03, dR09, BPP05, CCK88, DS04a, Gao86, Gao89, tH90, KM88, KSG03, MP08, PYF08, SD88b]. pipelined-loop-compatible [MP08]. pipelines [JP09]. Pipelining [LYC02, MK92, WGCZ09, DF90, JS86, KR06]. Pivoting [mYyF92, Vel89]. Pixel [Tay02]. Pixels [HPT^{+}97]. Placement [CB99, HJD^{+}01, KL05, LE91, PFJ04, RBD08, VA07, WLL08, WLWW09]. placements [AB03a, AB05, ZWS09]. placing [DDNS06]. Planar [SL97, TZ00, CP04a, CZ90, PD05]. Plane [OS98, RR95b]. Planning [RR95b, FL86]. plans [CBV08]. Plants [KSP^{+}92]. plasma [SDG08]. Platform [HS94b, AK06]. platform-independent [AK06]. platforms [AM07, BR08, GZMC08, HK08, LMR05, PVGG06, SK09, She06, WCO^{+}09]. PLUM [OB98]. PMESC [CJ99b]. PODC [KBD05]. Point [Als01, BSG90, CZ90, Gro85, MP08, PK07]. Pointer [Gup92, SCC^{+}06]. Pointer-Based [Gup92]. Points [Tor89]. Poisson [DM90a]. polarizations [PMV06]. Policies [DY99, KE93, ATZ07, FW05, LE91, PM05]. Policy [GM96, HBCM99, GNT04, YCC05]. polling [FA07]. Pollution [SS00]. Polygon [KSA95, RBD08]. Polygons [Wu02]. polynomial [DM88, MS87, Nic07, Sch89b, Sil90]. Pong [LF92]. POPS [RD05]. Popularity [SHLN09]. Port [CDF01, RJMC95, ST02, Dim04, ST06]. Portable [BK95, BHS^{+}94, LWP02, RHH96]. Portal [PLL^{+}03]. portals [BAK^{+}03]. Porting [KME09]. Ports [AW95]. positive [KK86]. Potato [NS95]. Potential [MK92]. Power [Ebe94, EB09, TVT96, BAPRS91, HMV07, WTB^{+}08, dR09]. Power-aware [EB09]. power-constrained [WTB^{+}08]. pp [Ano92a, Ano92c, Ano93e, BS96c]. PRA* [EHMN95]. Practical [Ger98, HCWS94, HR92b, HR92a, KK95, SGS99, YZS96, McD89]. practice [PTA08]. Practicle [Ano97k]. PRAM [AS91, DL98, HS94a, PRW94, Pra93, ZK94]. PRAMS [MR94c, FII04, GM94b]. Pre [VWHL96, RG06]. pre-execution [RG06]. Pre-run-time [VWHL96]. prearranged [SW90]. Precedence [JR95, KB96b, MMVR97, BKS05, DUW86, Li06b, ZV09b]. Precluding [Yen01]. Preconditioned [BSGM90]. preconditioners [SZW05]. Predicate [TG04, Yen01, AMK^{+}07]. Predicates [CK97, GCKM97, RS92b]. Predictable [CKK00]. Predicting [FFK97, Lun99, SSG93, SZD07, SFT04, Wei02]. Prediction [Ano97k, CTD99, KL01b, PH00, YZS96, YI96, CDB04, SM08a, SK05a]. Predictions [DD95, XZS96]. Predictive [DSW94, RKK06]. predictor [GGR89]. predictor-corrector [GGR89]. Preemption [MS98]. Preemption-Safe [MS98]. Preemptive [GAGPK03, Mar88]. Preface [Ano01-33, Ola01]. Prefetch [SD00]. Prefetching [BL96, KS97a, LY98, LY01, MG91, SMH94, SG99, SD00, HA05]. Prefix [HJ01, MP93, San02, AFM03, BS03, Han89, LH04, LS05, LH09]. Preliminaries [NBM93]. preprocessing [FSZ07]. Presence [ADS01, LT96, ISM07, RLH03, SAOKM03]. Preserving [NA02, JP09, OMSGNSG05]. pricing [GRDB05]. primary [AOSM04, BB03]. primary-backup [AOSM04]. prime [YLB90]. Primitives [FAM96]. Principles [KAS07, FK89]. priorities [BSMH08, KSS^{+}07]. prioritized [LW89]. Priority [BM97, BTZ98, Joh94, JNW96, KB96b, San98, TF92, FC90, HM06, MM07c, ST05]. priority-based [MM07c]. prism [Ros85]. privacy [LZSL06]. private [RFPAG08]. Proactive [RLH03, HOVC09, SZ09]. Probabilistic [CWL^{+}07, DM92, MK08b, SU87, ZA05]. probability [KNS06]. probe [ZFWF06]. Problem [AS95, AM93, ASST05, CLRW00, CRFS94, GP00, HH01, HC97, Kau94, KBC^{+}01, KLZ97, LF92, NW88, RDL95, TU92, TZ00, WH97, Zia92, AY89, ANP07, BB85a, BSG90, BFG04, BFM06, Boż09, DM90c, EE05, FMM^{+}08, GT04, HSSM07, Hsi04, IHM05, Joh89, KS91, LM05, LSS88, LWR^{+}03, LYL08, LS91, LH09, MGG03, Ngo06, PMV05, PBS08, SU87, WZ91]. problem-size-independent [LH09]. Problem-Solving [KBC^{+}01, LWR^{+}03]. Problems [Ano96i, Ano99g, ADS01, BK95, BOS^{+}95, BEE00, BGOS95, BMCP98, CB95, DS02, ESMG96, FR96b, FR98, FT94, GL92, KL01a, LSH96, MS94, MP96, MS99b, OR97, RS96b, Ser97, SN93, Ten90, TF01, WM92, WLR90, WHT02, WH08, ATH91, AG86, BGH^{+}03, BS03, BBd90, CEGS07, KJD03, LW06a, Lin91, Los08, LGG08, LV88, MPZ09, Nik04]. procedural [Kan05]. procedures [DWHL87]. Process [CCM92, IAS^{+}92, Kar95, KSP^{+}92, KOW97, Qia97, Ric98, SMR96, SS93, SF90, Ara90, Bic90, Gai87, Gai90, Lo92]. process-and-data-decomposition [Ara90]. process-oriented [Bic90]. process-time [Lo92]. Processes [DZ97, VWHL96, BFTV87, MAR05]. Processing [AyJ93, AK93, CS95b, GC95, HPT^{+}97, HSJP87, HR90, IWM97, KSL85, Kri92, LWY97, LS97, LS85, LT94, MSH90, MT85, NSM98, NMS93, Ros07, SH90, Sni03, SD88b, SSK96, SWC^{+}91, TAS^{+}01, THBF97, VB02, Wee01, WRC^{+}02, WSS93, Wei98, WA02, ZM94a, ZM94b, BB87, CC08, CRL04, CCN06, DW04, GSWW04, GWWL94, HR89, JMS86, KL08b, KNS91, Lee91, MS86, PI90, PVPM06, RG87, RTCG91, SCB08, SK89b, Sto87, SA90, TZH^{+}06, Trä09, WW07, Wan07, WJD91, XHY07, XQ04, Ano93a]. Processor [AW95, AERBL92, Ann94, BG86, CW93, CWW^{+}95, CkLCK04, CkLCK05, DY99, DDD98, GW99, Goe94, Guo94, HO94, Hwa97, JB98, KC98, KF90b, KBG92, LS91, MSd^{+}95, Moh96, MNM98, MBK^{+}92, NSS97, OS98, Par96, PT01, RKK97, SS93, SHC93, SS97, WCF94, YD98, YL98, Zhu92, ZYO02, ACYS08, Bat05, Bod89, CL88, CL85, Deh90, Eİ07, Gro85, HK08, HA05, Kri91, KR87, Lee91, Li05, MM07b, OT86, PLD87, RR05, RLH03, SI86, SI89, SSM89, SKK91, ST85, TR08, Wil92, LTKS90]. Processor-efficient [LS91]. Processor-embedded [CkLCK04, CkLCK05]. processor-in-memory [HA05]. processor-node [TR08]. Processors [CMS92, DBKF90, GR96, Hag97, HQPT99, HBH93, JR95, LPU97, MP96, AjHcC90, BD05, Bat05, BB85b, BR91b, CBM^{+}08, CK91, DPRW85, IC05, KK88, PK89, SC91a]. Product [AAD02, GE94, MSC96, CI03, Dim04, Dja06, ISAZ07, ST85]. Production [BBD^{+}91, HKT^{+}91, KM91, KM92, Nie94, Sch91, DM90c, GF89, HS86, SM86]. Products [ANS97, WLD00]. Professor [Ano04r]. Profiling [BST01]. profit [AM06]. Program [BDF92, BE95, DBP94, DD95, ERL90, Fer92, FJ93, GSG^{+}93, LSCA93, LMCF90, LAS^{+}97, MDD97, Mil93, NBM93, PP96, PS01, RRS^{+}08, SH92b, The02, WF93, YB01, ZYH94, GJG88, Kan05, RM90, ESA03]. programmable [AC89, MM07b]. Programming [AT94, AM93, AB84, BK95, BJ99, BCD95, Bal90, BN94, BB93, CP97, CCRS92, CCC92, CEF^{+}95, CBdCD00, CJ99b, FC95, Fre96, FBDC99, GP94, GGW96, GAG^{+}92, GLC01, HR00, JW94, JRR99, NT90, PA94, PM96, RAS96, SSOB02, Sin95, SC95, ZZC92, AE88, BAMM05, Boż09, CK88, CCC^{+}04, DRT07, EE05, EC89, ESA03, GL89, KKVI05, MSS88, RSR04, RR05, RSW91]. programming-based [KKVI05]. Programs [AH94, BB93, BCR96, BLG01, CMT93, CDY97, CGL^{+}95, CMS92, DR98, dADB96, ERA95, Fah96, Gup92, GHSJ96, HLJ01, Kar92, KY96, LP97, Lun94, Lun99, Mah95, MI92, QZ94, QH96, RJA97, RW93, SKR93, SG93, SSHC00, SK93, TR96, TG97, YI96, ZN01, ZH99, AY09, Bic90, DeG88, FKLB08, HK08, HS03, LC91a, LC92, McD89, NCT^{+}07, Nic07, Pop91, THSS87]. Progressive [RGS00, YIY97]. Project [FCO90]. Projection [AAP01, HSJP87].

Projection-Based [HSJP87]. projections [KM03]. PROLOG [SS97]. promoting [ABCM07]. prone [MFVP08]. Pronto [PF08]. PROOF [YJB91]. Propagation [CDP95, DF94, AAFV04, CKN07, CDB04, KMMZ06, PLR07]. Propagations [WD92]. Properties [BR95a, CW01, DC94, GK93, KAM94, YN92, NS90, PL06]. property [PB09]. Proposal [HPT^{+}97, VO89]. proposals [RFPAG08]. Protecting [SY04, LZSL06]. protection [DHS06]. protein [FGZ03, GZ08, LYL08, LVB07, Ngo06]. Protocol [BMMS01, CKL99, GRS97, GS96, GS01b, HP00, KUFM02, KB96a, LL98, Seb95, The02, ALF03, CL03a, CCHC09, CS08, CL09, CHC05, EBE08, Eri88, GCS06, LS06, Lun90, LM09, MCdS^{+}06, PGS06, ZPI06, ZWS09]. Protocols [AS00, DS95a, Dah99, Dol97, DSS95, GS00, HNM02, KCDZ95, AP03, BW89, BSW07, BPA06, CB06, FW05, GS03b, JBY^{+}05, LPX05a, Los08, MAM05, OSL05, Seb91, VA03, WTC08a, WTC08b, WCYR08, mYA91]. Prototype [CSSY94, KYL05]. Prototyping [DN94, WH97, PRG88]. Provable [KMP^{+}06]. Provably [DP99]. proxies [TC04]. Proximity [OSZ98, SX08]. proxy [HC09, KERUM04]. proxy-based [HC09]. pruning [MCC04]. PSIST [GZ08]. PTASs [LW06a]. PTRAN [ABC^{+}88]. PTW [PW96]. public [AM06]. Publisher [Ano04d]. Pulse [ZLPP01]. Purdue [SAB^{+}92]. Purpose [GFB^{+}92, CBM^{+}08, KL08b, Lo92, RGD03]. Push [AS95]. Push-Relabel [AS95]. puzzling [SPVvH03]. PVM [KOW97, LDCZ97, SKH96, WAS95, ZPI06]. PVM-Based [WAS95]. PVMe [BR95b]. Pyramid [DS93, RL95, Tan84, LW90, Ros85, WW04]. Pyramids [NPI^{+}96]. Python [DPS05, DPSD08].

QAP [BMCP98]. QC [ACYS08]. QC-2 [ACYS08]. QCD [IBP08]. QoS [CS08, DMB^{+}03, DÖ06, LL07, MS00, NP09, OY00, TBHA07, XHY07, XG03, YSL08]. QoS-aware [NP09]. QR [Kau94]. QSM [RGD03]. Quadratic [WNA^{+}94, MP88]. Quadrature [MD92]. Quadtree [IK93, WF90]. quadtrees [HR89]. Qualitative [Buc92]. Quality [LAZC00, SS08]. Quantifying [AAFV04, LDCZ97, Nik03]. Quantitative [Buc92, NBM93, MMAL^{+}06, ZI08]. Quantization [ZCK^{+}02, Nic88]. Quantization-based [ZCK^{+}02]. quartet [SPVvH03]. Quasi [AB05, Nik04]. Quasi-perfect [AB05]. quasi-threshold [Nik04]. Quasirandom [Bro96, CJ07]. queens [AY89]. queries [CI86, RHL08]. Query [AyJ93, CS95b, DM92, SK90, PRP09, KSI04]. Queue [BTZ98, CLT96, Joh94, RO92, Che90, CP04b, ACYS08]. queued [PY09a]. Queueing [dG91, HM06, KS03]. Queues [BM97, BCLR96, Kop97, PD92, San98, FC90, ST05]. Quicksort [BX93, CV91]. quiescent [MRRT07]. Quorum [NM02]. Quorum-Based [NM02]. quorums [BJPPM^{+}08].

R [Ano92a, BG90a, ZFS07]. R-GMA [ZFS07]. Race [HM96, ISZBM99]. radii [OMSGNSG05]. Radio [CGKK97, CDB04, CCS06, GPT06a, KK06, MKC^{+}09]. Radio-wave [CDB04]. Radiosity [SHT^{+}95, YIY97]. Radix [BVB02, BDKM94, LJKS02, MG09]. Radix- [BVB02, BDKM94]. RAID [CT93]. Raking [BCZ95]. Random [Alu97, BA01a, PK89, SR97a, SR97b, SLP^{+}98, SS97, Li06b, SKK91]. randomization [CJ07, FII04]. Randomized [AFM09, BDF01, CDCD05, HBJ98, HT06, LW06b, MVM04, RR95a, Raj96, San98, Vis87, Bad04, DJT03, SK05b]. Randomly [SS96]. Range [SIR92, TDC05]. ranges [CYZ06]. Ranking [SGS99, AAD03, Vis87]. Rapid [PRHB06, CL85]. raster [Wri91]. Rate [MO97, RGS00, ÜD96]. Ratio [MO97]. Rational [GM95, KM88]. Ray [RGS00, CDB04]. Ray-Tracing [RGS00, CDB04]. Raynal [Ano96l]. RC [VV90]. RCC [HH97]. RCC-Full [HH97]. Re [FVCL05, PRHB06]. re-engineering [PRHB06]. Reachability [CCM01]. Reactivation [CW93]. Reactive [DLS00, HPT07]. Reactor [KKS08]. read [CH06a, GNS09]. read-modify-write [CH06a]. read/write [GNS09]. Reader [JBP00, HV09]. Ready [JM00]. Real [AAL95, AK93, Ano92c, BPJG92, BA96, BA01b, CS93a, Cha94, DJ98, EMP^{+}96, GMM00, JH92a, KS97b, Lee03, LTY96, LM96, MMRS98, MMVR97, Moh97, MSST99, OY00, PS93, RDS02, RU99, RAS96, STN92, THBF97, WLID02, Zim96, van96, AOSM04, AOSM05, BW08, CCN06, EDÖ05, Gos90, HOVC09, HA06, HL07, LHK03, LZCY09, MAM05, NA06, QJ05, RLH03, TZH^{+}06, WL05, XO05, ZB03]. Real-Time [AAL95, AK93, Ano92c, BPJG92, BA96, BA01b, CS93a, Cha94, DJ98, EMP^{+}96, GMM00, JH92a, KS97b, LTY96, LM96, MMRS98, MMVR97, Moh97, MSST99, OY00, PS93, RDS02, RU99, RAS96, STN92, THBF97, WLID02, Zim96, van96, Lee03, AOSM04, AOSM05, CCN06, EDÖ05, Gos90, HOVC09, HA06, HL07, LHK03, LZCY09, MAM05, QJ05, RLH03, TZH^{+}06, WL05, XO05]. realistic [KNS06]. RealTimeTalk [EMP^{+}96]. rearrangeability [DD96]. Rearrangeable [CS93c, HJDH01, FY86, Pak89]. Rearrangement [BVB02, GL92]. Reasoning [PS88, Ste95, eW95]. recipients [Ros07]. reciprocal [SL90]. Reclaiming [GMM00, MMVR97]. reclamation [HMBW07]. Recognition [BMRC99, RGU08, SP96, WPKK94, LO91, PD05]. Reconfigurable [AT94, BAGS95, BSDE96, BBR94, BM97, BA95, BGOS95, COS^{+}95, CGG^{+}09, DS01, EL97, EH01b, FZVT02, HQPT99, HCWS94, JP95, JS94, JB98, KF90a, LS95, LPZ99, LR93, MD01, MG93, MT97b, Nak95, NS94, OS96a, TVS97, TBPV00, WHT00, dR09, BM04a, BPP05, CDJ^{+}89, DS04a, FX06, HPSM91, Mat06, MP08, PVG09, SI89, SL89, TRSS06, WJD91]. Reconfiguration [CGA98, QMCL94, UR94, YTR94, BAPRS91, PSPR05]. Reconstructing [OOW95]. reconstruction [FCG04, FGG08, KM03]. reconstructions [SHT^{+}08]. Recovery [CP01, FCF00, JF95, LS01, MFS93, BG05, DWG03, MM04, MM06, MS02, PGS06]. rectangle [Deh90, LV88]. rectangles [KF95a]. Rectangular [CWW96, Dja04]. Rectilinear [Nic94]. Recurrence [CP94, Car90, MP87]. Recurrences [BCZ95, GP94, NCTT09].

Recurrent [WT92]. Recursive [CW01, CB95, CTZ99, GHSJ96, KC99b, Lee94, LT07, RS92b, SCD99, ZYO02, ERS90, SMKL93, DC94]. Redaction [SWC^{+}91]. Redistribution [PT97, RSB96, BBB^{+}06, GP05]. Reduce [KLS90, SDS99, LMR05, Lun90, MP08, PY09c]. Reduced [AP94, CC87, Gro85, HJ90b]. reduced-instruction-set [Gro85]. Reducible [DH94]. Reducing [BCM87, BD04, FGP05, GS00, PB90, SS93, CX05]. Reduction [PA97, RJY96, SSG93, SM92b, LS88, Sch87, ST08a]. Redundant [MT93b, MFS93, MFS96]. Reference [KS00, CH06a, FPP06, WL92]. references [SYYU07]. refillable [ALH^{+}09]. refined [Mit07]. Refinement [FLS^{+}97, NA02, Mit07]. refinement-tree [Mit07]. Reflections [Zim96]. reformation [LHT08]. refresh [OPG08]. Region [CRFS94, WLR90, DDNS06, MGG03, TYH09]. Regions [GS01b]. Register [JBP00]. registers [GNS09]. Regression [HK01]. Regular [KBG92, NIR86, SZB92, SS92, SS95, TC96, TL96, Eİ07, Hal05, Lee90, WG08]. Regulated [PD92]. reindexing [DQR^{+}09]. Relabel [AS95]. Related [Ann94, Dol97, JR92, Man94, MS99b, OD95a]. Relations [OO85, CG86, CC87]. Relationship [MDD97]. Relaxation [MHC95, Tor89, RS08]. relay [LR03a]. relaying [TBHA07]. Release [KCDZ95, LTWY95]. Reliability [GP93, GST09, HHC98, MT93b, AH06, MSM09, QJ05, TYH09]. reliability-driven [QJ05]. Reliable [BG05, DM99, GS01b, KGN89, LHP07, Tze93, HOVC09, KSI04, KL05, MK08a, MRRT07]. Relocation [YCY^{+}00]. Remapping [OR97, ACFK07, FXW03]. Remappings [CA96]. remark [PMV06]. Remote [DM99, KS00, WMG01, BBB^{+}06, CH06a, Lon04, MSJ05, WGCZ09, ZWRI07]. Remotely [DSAUM99]. Removal [KK95, SSL04]. Renaming [Gil94]. Rendering [Tay02, WS97a, ACFK07, WJV07]. rendezvous [PHS04]. Rent [Oza04]. Repartitioning [MNM98, PP96, SKK97, CBD^{+}09]. Replacement [CKL99, YCC05]. Replay [ZN01, NRM^{+}09]. Replica [RAB08, WLL08]. Replicas [HJD^{+}01, ZWS09]. Replicated [JSM94, LO96]. Replication [CA95a, JRR99, Li99, DS04b, KA08, LA04, SZ09, ZWL03]. Report [FCO90, SAB^{+}92]. repositories [KUA07]. Representation [CJ99a, TLW94, CJY04, EHS94, JZ05, VO89, WF90, Wri91]. Representational [Ebe94]. representations [BHR91, NCTT09]. Representative [BW96]. representing [BR91a, NAK04]. reproducible [PK05c]. reputation [SL06]. request [XHY07]. requesting [XO05]. Requests [TSC01, BPRG04]. Requirement [DDD98]. requirements [DÖ06, MVM04]. Research [Ano01-34, Lan09]. reservation [RKK06]. Resettable [AKD06]. Resetting [YH97]. Residual [DRR96, SR95]. residue [DPRW85, Tay87]. resilience [WXZ05]. resilient [LAGK07]. Resolution [YB95, GE85, LJ05]. Resolving [LKK94]. resonance [CCN06]. Resource [AB84, BMF05, CKK00, GMM00, MMVR97, NSTN91, OM84, RDS02, RSBN01, SM94, YT05, ZI08, ALH^{+}09, AB03a, AB05, AKSM08, ATZ07, BMB^{+}08, BSMH08, HSLL04, JK89, LCC^{+}05, LC91b, RKK06, RLH03, She09, SSMS08]. Resources [HS94b, AM06, AM07, NAK04, SSM^{+}06, SSM^{+}07]. response [DHK04, VA07]. restoration [UAPM07]. Restricted [Fra92, MSSE02, BS03, BBM08, DeG88]. Restrictions [Li92]. result [Lon04]. Results [IPK85, Sch91, SH92b, BR95b, SZ03]. Retargetability [MB96b]. Retrieval [AA93, CLV95, KV88, Lon04]. reuse [CCHC09, DK04]. Reversal [NTA96, Ede91]. reversals [BS03]. Reverse [LP97, JXW06]. Reviewer [Ano08, Ano09]. reviewers [KL08a]. revisit [LLS07]. revisited [GDP08]. Revisiting [MR09]. Reward [SM92a, CMT92]. RFID [CRK^{+}09, CL09]. rhombic [Wil90]. Riccati [MV94]. Rigid [JBL02, LF03]. Ring [BA95, CMS92, FFK97, Goe94, GH96, HJD^{+}01, MBK^{+}92, ZB97, BG86, MM04, PV89, RKS87, YC04, ZWS09]. Ringed [DVZ96]. Rings [FKSW97, GR96, KY02, KUFM02, LHS97, LSC00, MS94, Man97, YTR94, CL91a, FKK^{+}04, LC92, LW06b, Sil90, Tsu07, WT09]. RISC [HC91, LPU97, MSC96]. RISC-based [HC91]. rising [ORR03]. RMF [YT05]. RMI [CCK^{+}08]. road [IB04]. roadway [XCLR07]. Robin [CMS04]. Robust [SSM^{+}07, AKSM08, GA90, TZH^{+}06]. robustness [Par05, SSMS08]. Role [Cha95, Won99]. Role-Based [Won99]. Rollback [JF95, AAFV04]. Rollbacks [SS93]. root [EL91]. Rosenberg [Ano00d]. Rosenfeld [Ano04r]. ROSS [CBP02]. Rotation [HC95, HBH93, Ara90, EL88]. Round [CMS04]. route [CDCD05, LPX05a]. Routed [FF98, NSSS99, RJMC95, RMC97, XMN92, MVM04, SAOKM03, WCC02]. Router [DRSB01, PIB^{+}01, MBR08, XYKA08, CCQ^{+}06]. Routers [CP01, CP04b]. routine [IBP08]. Routing [ASH^{+}01, AZ01, ABJS01, BLPV95, BPvW96, BP98, BA97, BA01a, BW95b, BDF01, BN03, CRV94, CL93, CW01, CL96, CC94, CLT96, CCR94, CS93c, CDF01, CG02, Dol97, DG94, EL97, GG01, GHKS98, GO95, GT97, HCWS94, HJDH01, IM00, JR92, KLLK98, LS94, LTWY95, LTY96, Li92, LME95, LW95, LEB98, MS00, MS94, MW95, MR03, MJ94, NSSS99, NS95, OM90, PRW94, Par96, PA97, PA01, PL93, RS94, RS96b, RH05, RO92, RR95a, RW97, SJ95, SJ96, SB02, SZB92, TBPV00, WLY01, Wan96, WN94, WLD00, YBOY97, PRP09, BSW07, BR91b, BPA06, CI03, CL03a, CS06b, CS08, CDCD05, CL90, DMB^{+}03, EB09, GAGPK03, GLD06, HNSA07, HL07, JL05, KSI04, KMF^{+}05, KO90, KT91, KNS06, LPX05a, LS03, LAGK07, LH05, MCdS^{+}06, MBR08, MVM04, NJ91, OS04]. routing [OSL05, RD05, RHL08, SLWW05, SK05b, TC04, TT07, VA03, WTB^{+}08, WGS08, XHG03, XG03, YME06, ZV06, ZMC06, ALF03]. Routings [WIKC97]. row [Mat06]. row/column [Mat06]. rows [ST87]. RPC [BF97, VD04, WSG91]. rRNA [ZFWF06]. Rule [KM91, Mir91, Nie94, SWC^{+}91, XH91, MCdS^{+}06, Oza04]. Rule-Based [XH91, MCdS^{+}06]. Rule-Firing [Nie94]. Rules [RSD94, SM92b, SWC^{+}91]. Run [FBK98, FY97, LPU97, LFA96, MDD97, PM92, SCMB90, GF89, SFT04, VWHL96]. Run-Time [FBK98, FY97, LPU97, LFA96, PM92, SCMB90, GF89]. Runge [KR06].

Running [CCM^{+}06, FGP05]. Runs [Lin93a]. Runtime [Bir94, BJK^{+}96, KR97, KPC96, NS97, NsPPC02, PT97, SK91].

S [ASST05, BPJG92]. S-Nets [BPJG92]. Safe [FM99a, MS98, HV09]. safety [Wu03, XCS06, XCLR07]. SAGE [Num09]. Sampling [OS96a, SS92]. SAMR [CP05, LTL06]. SAN [SM92a]. SAN-Based [SM92a]. sandboxing [SFEF06]. satellite [TZH^{+}06]. Satisfaction [GHH92]. Satisfiability [Soh96, Joh89]. Saturation [Tze91]. save [FKLB08]. Saving [DKY01]. Sawchuk [Ano93e]. Scalability [AFT^{+}00, BCV94, BP01, DVW94, KS91, KG94, MR94a, QZ94, SSRV94, Sun02, ZYH94, ZFS07, CSW08, KR06, YH07]. Scalable [AYI97, BMRC99, CSWD03, CSSY94, CAB94, CLV95, CBdCD00, Cou93, DA97, DD93, DM04, DSW94, DFRCU99, DSD^{+}97, DT92, DM94, FR96b, GH02, HA92, KA03, KP00, KC94, KGV94, LZ02, Li01, LWP02, NKC^{+}97, NRM^{+}09, NPY^{+}97, PA94, Pra93, SMH94, SN93, Sun02, WPKK94, WW96, XKMN94, ZMPE00, ZB09, BGM^{+}08, CS08, CD95, DS04a, GZ08, GREC91, HWC08, LHK03, LC07, LB09, MK08a, SSTP09, WJV07, XCZL03, XJS03]. Scalar [VH93]. ScalaTrace [NRM^{+}09]. Scale [ABDS02, BMCP98, FZVT02, GK93, HHM94, KL84, LK98, OK01, RFM94, VN93, BMB^{+}08, BMF05, KESA07, LYL08, PTZ06, RW02, VM03, WLNL06, XHY07, ZV09a]. Scaleable [BMRC98]. Scaling [SSS07, TBPV00, FKLB08, Num07]. Scan [KB96b]. scanners [CCN06]. scatter [BM04b, LMR05]. scattering [DB86]. scatternet [SLWW05]. schedule [KSG03]. Scheduled [LB90, HA06]. Scheduler [NPP^{+}02, HDJ08, KS03, LB09]. Scheduler-Activated [NPP^{+}02]. schedules [Dja06, DQR^{+}09]. Scheduling [AGF94, ALL99, AMN00, AGG98, AS97, AYIE98, AKPT99, AjHcC90, BPJG92, BD05, BPN90, Bec96, BCLR96, CDY97, CL91b, CLL09, CJ99a, DA97, DR95, DDD98, DP99, DS84, DAYA02, DÖ06, DJ98, ERL90, ERA95, FAGW95, FVLB09, FR92, FR96a, FKSW97, Gai90, GR96, GY92, GM99, HO94, JSCB95, JSWB92, JR95, KS97b, KB96b, KA97, KA99, LPU97, LYC02, Lun94, MMRS98, Mah95, MSd^{+}95, MSSE02, MYD95, Moh97, MSST99, NSSS99, OH02, PKN08, PAM94, PS93, PM96, QM01, RU99, SCMB90, Ser97, SH92a, Sta04, SD88b, SYG92, TSC01, TTG95, VB02, VWHL96, WCF94, WSRM97, WA02, WUG99, YI96, YWD08, dSR00, AL04, AOSM04, AOSM05, BKS05, BGLA03, BFG04, BFM06, BH05, Cal06, CGW^{+}03, CRA^{+}08, CJY04, DBC03, DK08, DUW86, DJT03, FA07, FW05, GDP08, GP05]. scheduling [HDJ08, KA03, KUA07, KVHS07, KDH08, KMP^{+}06, KA05, LHK03, LC90a, Li05, Li06a, Li06b, LL07, LPX05b, Lo92, Mar88, MS86, MAR05, ORR03, PY09a, PK05a, QJ05, QSL^{+}08, QGL^{+}09, SV08, SLG06, SCJ^{+}08, SWP90, TG03, Tsu07, VD04, VB08, WJD91, WAE03, WL05, XQ07, ZV06, ZV09b]. Schemas [Arb89, BG90a]. Scheme [BDF01, FY96, JB93, KK98a, LO96, MYD95, OS96a, Wu94, YD98, AOSM05, CWLD05, EL88, GMXA07, HC09, HOVC09, KVHS07, KRL87, LTB02, LHF91, NC09, RS08, SZ09, SKMM04, TDC05, WL04, WW04, XYDL06].

Schemes [yCM98, FM99b, GG01, LL95, LS01, SKK97, WRC^{+}02, ZLPP01, AAD03, BLPA05, BR91b, CI03, GJXZ05, HSMB91, MM06, TW89]. Schmidt [ZLRP91]. Scientific [CCRS92, DUSH94, FMW^{+}94, GT02, HS94b, KBC^{+}01, AOS^{+}05, AE88, NTC03, VM03]. SCO [WTS03]. SCP [VB08]. SCP-based [VB08]. screening [AT03]. Scrolling [Tay05]. SCSI [HZY04]. SCSI-to-IP [HZY04]. SCTP [ZPI06]. SDEF [EC89]. SDSM [CCM^{+}06]. Seamless [HR00]. Search [BOSW94, BS00, BMCP98, CDRC99, DM95, DM92, EHMN95, Fen90, LYC02, SIR92, BNP02, BP89, CCLS94, GJXZ05, KA05, MSM09, PRHB06, Par89, PVGG06, RHL08, SP08, SHLN09, WGC09, YF09, Zep91, ZH07]. Searching [NBP98, NSM98, SH97, BA06, Sch89a]. Section [Seb95]. Sections [BW96]. Secure [BKT95]. Securing [SL06]. Security [SXZ06, BAK^{+}03, LZSL06, LCM^{+}06, SFEF06, VA03, XQ07]. sediment [CvdBL^{+}08]. Segmentation [KC99b, MG98, MGG03]. Segmenting [TVT96]. Segments [RR95b, GC07]. Seidel [HO94]. Selecting [SSG93, KERUM04]. Selection [JK00, LK96, PT01, Raj96, RW97, RCY97, Raj01, SH97, SB02, VS99, WSA^{+}94, WRC^{+}02, Bad04, EDÖ05, RH05, RAB08, RD05, SSS88]. selection-based [EDÖ05]. selections [JW89]. selective [XYG07]. Self [Ano02u, AS96, ABZ95, BGJDL02, Bec96, BBCD02, BAGS95, CW05, CT04, DB08, Dol97, GH02, GS03b, HPT07, HPT02, HNM02, KY02, MM07a, NM02, PK05c, SZB92, SEP96, BFG^{+}03, BR91b, BFKP04, BZH06, CRA^{+}08, KO90, MC91, NJ91, SLWW05]. self-adjusting [MC91]. Self-Allocating [SEP96]. Self-organization [CT04]. self-organizing [BFKP04, BZH06]. Self-reproducible [PK05c]. Self-Routing [SZB92, BR91b, KO90, NJ91, SLWW05]. Self-Scheduling [Bec96, CRA^{+}08]. Self-Simulation [BAGS95]. Self-Sorting [ABZ95]. Self-Stabilization [GH02, HPT02]. Self-Stabilizing [Ano02u, AS96, BGJDL02, BBCD02, Dol97, HNM02, KY02, NM02, CW05, DB08, GS03b, MM07a, BFG^{+}03]. Self-tuning [HPT07]. selfish [WGS08]. Semantic [FKJG08, RHL08, CM93, LR05, ZH07]. Semantics [JK89, HK05]. Semi [DS04b, XZS96, KMS^{+}06]. Semi-empirical [XZS96]. Semi-passive [DS04b]. semi-static [KMS^{+}06]. Semiconductor [DM90a]. Semidirect [WLD00]. semifast [GNS09]. Sensed [DSAUM99]. sensing [HP06]. Sensitive [VR95, Ano04d, CP05, GS03a, GC07, PFJ04, YK04]. Sensitivity [HJ90a]. Sensor [KSI04, STN92, ASM09, BXA08, BPA06, BZLI04, CKN07, CDR09a, CDR09b, CT04, DW06, DFP06b, DH04, GDP08, GCY^{+}04, HMV07, HP06, IB04, KKVI05, KOA09, LLW07, LHP07, MSM09, MK08b, NC09, OMSGNSG05, PFJ04, PLR07, PB09, SS08, TBHA07, TDC05, WW07, WMW09, XCLR07, XQ04, ZC04, OEY07]. Sensor-centric [KSI04]. sensorial [VO89]. sensors [BFKP04, Cal06]. Sequence [JP09, Zak01, AFM03, BBM08, BW09, BMARM07, FCS91, JV09, PTZ06, TMM06]. Sequence-preserving [JP09]. sequencer [BCM06]. sequencer-based [BCM06]. sequencers [CHC05]. Sequences [Swa98, TR96, CJ07, LVB07, SK09, Sei05]. sequencing [CRL04]. Sequential [KF95b, BFTV87, Fen90]. sequentially [HK08]. Serial [EMMM94, MT97b, BOI91, CR91, CL90, SD88a, SI91]. serial-data [SD88a]. Serializable [Sch91]. Series [CA95a]. Series-Parallel [CA95a]. Server [ALL99, AYI97, CM92, GM99, HBCM99, JSCB95, RU99, HC09, OS04, PM05, TBZB05, WLWW09, ZI08]. Servers [FM99b, Bar05, BPRG04, CSWD03, KCD08, PSPR05, Wan06, WDDK09, ZWL03]. Service [CTT08, JRR99, LAZC00, RGVB00, DB08, HOE^{+}09, KMMZ06, LC07, NP09, PY09b, SFEF06, WS06, Yan09, ZI08]. service-aggregate [Yan09]. Service-oriented [CTT08, SFEF06]. Services [ZR00, AK06, AM07, LCC^{+}05, XJS03, YWD08]. sessions [TK07]. Set [Als01, BCD95, DM92, KF95a, KSA95, KHS96, RDL95, BD05, CC87, DW06, Gro85, HJ07, Lon04, Nic07, SZW05, WCWH03, WCKD06, ASST05]. Set-Based [BCD95]. set-distributions [Nic07]. Sets [AAP01, CGL^{+}95, EP90, GT97, Pov99, Lon04, MP08, PK07]. Several [CP92]. ShadowObjects [JRR99]. shallow [CvdBL^{+}08]. share [PVGG06]. share-nothing [PVGG06]. Shared [AGW98, AGW01, AD95, BS96a, BJS03, CP91, DS95a, DH95, GDN^{+}98, HV95, HS00, HPT02, HTL99, HA92, JF95, KRC00, KS97a, Kel00, KC94, KY96, LK98, LA93, LT94, Lu01, MF94, MS98, MG91, MSST99, PY96, RL96, RJY96, SDS99, SC91b, TJ92, TTG95, TY95, Wil92, YW91, YMR93, YL98, Zak01, AL04, BC06, Car95, CCM^{+}06, DI91, FZC^{+}05, LHT08, NSTN91, OC07, Pad91, PY09b, PK05b, RFPAG08, SM04, TK07, WL92]. Shared-Memory [BS96a, CP91, DS95a, HA92, KS97a, LK98, MF94, MG91, SDS99, TTG95, YW91, YL98, Zak01, BC06, DI91, FZC^{+}05, NSTN91, PK05b, RFPAG08]. Shared-Nothing [LT94]. Shared-Noting [HTL99]. Sharing [CS93a, DY99, HS97, HF96, Cho90, IS06, KUA07, KG04, LY91, MTS90, SU87, TBZB05, WTS03, XCZL03]. Shear [SSM89]. Shear-sort [SSM89]. shelf [PF08, ZB09]. Shifts [OP96]. shop [Boż09]. Short [ESTA94, KLC05]. Shortest [BGR96, HTB98, IZ95, KC99b, TU92, TZ00, ATH91, KS91, YME06]. shortest-path [KS91, YME06]. shrew [CH06b]. SHRIMP [BF97]. Shuffle [BAES92, JH92b, Pad93, PA97, JT88, Var91]. Shuffle-Based [Pad93, PA97]. Shuffles [Ano93e, CS93b, CS92]. side [CK88, HA05, TC04, XCH08]. Sided [ZB97]. SIEVE [SG93]. Signal [RTCG91, SH90, THBF97, WW07, XQ04]. Signal-processing [RTCG91]. Silence [DKY01, FJ93]. Silicon [THN^{+}93]. SIMD [AB93, BAES92, Che05, CP94, CD95, FAGW95, GGW96, GSWW04, HCS^{+}00, HCZ04, Ho91, IK93, IKS87, JMS86, KNS91, KLS90, LWOG02, ML89, NT93, Nas94, RS96a, RS90b, SI91, Ume85, WSA^{+}94, WLR90, ZLRP91]. SIMD/SPMD [WSA^{+}94]. Simple [BW96, GPS96, GB93, GS99, KW02, LW06a, PL94, TZ00, Koc91, MRRT07, MC03, BJ99]. Simplex [Shu95]. Simulated [Bev02, BH86, HB97, HC91, RSS96, Soh96, XH91, AH06, BG89, GE85, Ume85]. Simulating [DS02, DN94, LC90b, eW95, WCKD06]. Simulation [ABDS02, Ano92c, Ano02v, AS91, AB93, BAGS95, Bou02, Cha96, DMSH90, DS93, EH01a, GGN93, JH92a, KZ96, LZ02, Lin93b, Lin93c, LA93, LLCC02, MHF93, MRR^{+}02, NH93, Pra93, RSD94, RS92d, SMR96, SH92b, SSRV94, SS93, The02, ZL93, BM04a, BD04, BAL05, BMF05, CvdBL^{+}08, CTCX08, FGM^{+}03, GRR^{+}05, HDT^{+}05, Koc91, LVR90, Mat06, QS05, RW02, WBTM09, WF89, ZZ90, ZCK^{+}02]. Simulation-Based [RSD94, SSRV94]. Simulations [ASR93, Ger98, GM94b, HP95, KP00, LHM95, NM95, PAH^{+}98, RPS93, FII04, LTL06, SDG08, SM04]. simulative [HW03]. Simultaneous [CW93, ABC^{+}09a, BPRG04, Che90, FC90, MR09, PTZ06, SLG06]. Singhal [Ano96l]. Single [ALL99, JBP00, MWL00, TZ00]. Single-Source [TZ00]. Single-System-Image [MWL00]. Singular [Bai94, HBH93, PE93, Luk85]. Sisal [FCO90, PAM94]. Site [MFS96, LFH^{+}03]. situation [LR03b]. sixth [Arb89]. Size [COS^{+}95, CLT96, CVJ09, GSWW04, LH09, NW88, OS04]. sizes [GPT06b]. Skeletons [GSP02, Ski96, BR08]. Skew [SYG92]. skewing [TW89]. SLA [ATZ07, AM06]. slack [FKLB08]. Slant [ESTA94]. slave [LZ05, YH07]. sliced [KRL87]. Sliding [OS98]. Slotted [HQPT99, MSST99]. Small [CDH84, LH04, MSZ05]. small-world [MSZ05]. Smaller [HH01]. Smallest [Wu02]. Smart [HPT^{+}97, MKC01, CkLCK04, CkLCK05]. Smoothed [JK00, PAH^{+}98]. smoothing [HT06]. SMP [Bev02, FGP05, KA03]. SMPs [BJ99, BC05, BJS03, FW05, HLCZ00]. SMT [ABC^{+}09b]. SMT-based [ABC^{+}09b]. snap [PV07, MT85]. snap-stabilizing [PV07]. Snapshots [Mat93]. Snooping [Dah99]. social [SK89b]. socket [MAJJ05]. soft [AOSM05]. Software [AL99, CR96, CHR94, CLRW00, GS00, Gro85, HS94b, KCDZ95, Kel00, KB01, KS95, MLC^{+}90, MG91, NT90, SG99, San95, SZ00a, TY90a, VSM96, XKMN94, ABC^{+}09a, CMT92, DHS06, KG04, LZSL06, SCC^{+}06, SMH91]. Software-Based [KCDZ95]. Software-Controlled [MG91]. Software-Only [GS00]. Solaris [Lun99]. Solution [DM90a, FLS^{+}97, LF92, OH02, PW96, RW01, AY89, ANP07, Bat05, GS91b, LYL08, WZ91]. Solutions [Ano99g, CLRW00, RS96b, AG86, BAH04, LZ08]. Solver [BMM97, CSSY94, BAMM05, CK91, GV86, Gao86, KKB^{+}06, MP87]. Solvers [CHR94, CP94, MS99a, TF01, KR06]. Solving [BCZ95, Boż09, BMCP98, Car90, CRFS94, GL92, IK94, KL01a, KBC^{+}01, Mon94, PMV05, QOvdG01, WM92, WLR90, WH97, CM03, GGR89, GT04, LWR^{+}03, PF91]. Some [BDKM94, DKMV01, IPK85, KAM94, Oru94, Par98, SI86, SZ03, ZHO03, AG86, BS03, BDjQ86]. SoMR [CS08]. Song [Ano97k]. Sort [LJKS02, Tay02, SSM89]. Sort-Last [Tay02]. Sorted [SH97]. Sorters [BNP98]. Sorting [ABZ95, CQ95, DL98, FKK^{+}04, FY96, HQPT99, HBJ98, JP95, Lee94, Lin93a, MP93, NS94, OS96a, RW97, SCC92, SS92, SM00, VN93, WRC^{+}02, Che89, FCS91, MS88, PB90, SSM89, Sei05, SA08, Ull84, ZFL89]. Sorts [ZAW94, SI86]. SOS [PP92]. Sound [DKY01]. Source [AY09, TZ00, LPX05a]. sources [Lon04]. SP [ASH^{+}01]. SP1 [BR95b].

Space [BW96, BH93, DY99, GG01, GW99, GRS97, KM97, KY96, LZ02, NC97, RP98, SH98, WA02, WS97a, ACFK07, BBM08, Dja04, HV09, KA05, MSM09, MSS00]. Space-Based [LZ02]. Space-Efficiency [GG01]. space-optimal [Dja04]. space-optimality [HV09]. Space-Time [WA02]. Spaces [RS92a]. Spanners [RL95]. Spanning [FA95, KC98, KC99b, WB01, BFG^{+}03, BC05, BC06, BBL04, tH90, LVP08, Lin03, OMSGNSG05, TDM05]. Sparse [Bas97, BW95a, KK98b, Man94, MSC96, NFEG97, Shu95, UZSS96, Win85, AAD05, ANP07, BC06, She06]. Spatial [GSG^{+}93, NPY^{+}97, CCHC09, MLG05, NAK04]. Spatial-Temporal [GSG^{+}93]. Spatially [DS02]. SPEAR [RG06]. Special [AP93, AL99, AB03b, Ano95i, Ano95j, Ano96j, Ano96i, Ano97j, Ano99g, Ano01e, Ano02v, BOP06, BG90b, BD00, BS09, Chi92, CDJL09, DOP98, Dek00, DT92, ES97, FR98, GC95, GS01a, Gra09, Irw88, IB04, JW94, KL08b, KRS01, Lan09, Lin93b, Mir91, NT90, Ola01, PN97a, PN97b, PA96, Raj08, Sch90, SXZ06, SH92a, SB97, Sto90, TY95, Wee01, XJS03, YW91, ZO97, dVCP06, KL08a, WW03]. Specialized [QOvdG01]. Speciating [GB06]. Specific [PP92, SK93, SS94b]. Specification [AS00, BR95a, BN94, RSW90]. Specifications [LSCA93, BCM06]. Spectral [SANY94, SSB98, AT03, CH06b]. spectral-screening [AT03]. Speculative [RG06, MG09]. Speed [BBH^{+}97, Fer95, PVG09, SR91, WCYR08, HP97a]. Speedup [AMB95, DBP94, FFK97, Lun99, SN93, YH07, NW88, SC91a]. speedups [Vis87]. spikes [ST08a]. spin [AK07]. Spinning [BHK^{+}94]. Spite [VR94, DB08]. Spline [CWW^{+}95, CY96, GM95, Meg91]. split [WCWH03]. split-stars [WCWH03]. splitting [PVGG06, WSH^{+}03]. SPMD [Gup92, OKB95, RW93, WSA^{+}94]. spoofing [KMMZ06]. Spot [LKK94, TY90a]. spots [LK90]. square [BB85b, EL91, LTW^{+}90, XBK07]. squared [RIZ90]. Squares [CB95, ZYO02, BBd90, HLS03, KAP90, LTW^{+}90, SMKL93]. Squashed [BG90a]. Squid [SP08]. SRAM [JP09]. SRAM-based [JP09]. Stability [Wor93, KMS07]. Stabilization [CG02, GH02, HPT02, NA02]. Stabilization-Preserving [NA02]. Stabilizer [AD02]. Stabilizing [Ano02u, AS96, BGJDL02, BBCD02, Dol97, GH96, HNM02, KY02, Kar02, NM02, BFG^{+}03, CW05, DB08, GS03b, MM07a, PV07]. stable [AMK^{+}07]. Stack [PVGG06, CS06b]. Stage [FT94, SZ00b, HDJ08]. staging [EDÖ05]. Staircase [Mck94]. stalling [BHPP05]. Standard [CB99, PF08]. Star [FA95, KAM94, Lat95, LK94, MJ94, OS97, OS93, PRW94, RW97, RWY93, RLS96, SAOKMA02, dBL95, AAD03, CM03, DFP06a, FMM^{+}08, PK04b, SS05, WCC02]. star-access [DFP06a]. Star-Connected [dBL95]. Stardust [CP97]. Stars [MR03, WCWH03]. State [HB97, HNM02, KM92, NC97, CWLD05, KA05, LMR05, LW06b, MSM09, WCO^{+}09]. state-of-the-art [WCO^{+}09]. State-Space [NC97, MSM09]. Statement [AMB95, DR95, ALS91]. Statements [KHS96, SOG94]. States [Kop97, TG97, FZ90]. Static [AKSM08, BPN90, BSB^{+}01, BSMH08, CC91, ERA95, GF89, LC90a, LK94, LA04, MSd^{+}95, OD95b, SSM^{+}06, DK08, KA08, KMS^{+}06, McD89, SSMS08, SWP90, SSM^{+}07]. Statically [LB90, Mat06]. station [GPT06a, RBD08]. Stations [DKMV01, DDNS06]. statistics [GA90]. steady [LMR05]. steady-state [LMR05]. Stealing [Ano00d, LS97, Ros99]. Stein [QOvdG01]. Step [CW00, Yan04]. steroids [Bar05]. sticker [GPX08]. Sticky [Kop97]. STICS [HZY04]. Stigmergic [PR06]. Stochastic [CTD99, FX06, HPT^{+}97, JSS92, QZ94, RS92d, SSMS08, CMT92, MM06, MS86]. Stopping [BSS99]. Storage [CLV95, HLL^{+}95, LL95, BL05, BCK^{+}09, CGG^{+}09, HZY04, HK04, XCLR07, ZV09a]. Store [CP90, NS95, VA07]. Store-and-Forward [NS95]. straight [GC07, Wri91]. Strategies [AM07, BDjQ86, BHK^{+}94, BCR96, CP92, CGA98, DL01, FF98, GJG88, GM99, LK98, LHM95, Lun94, MS99a, OP98, SMH94, VB02, VA03, YB95, YL98, Zhu92, ZM94b, BMARM07, DM94, GRV08, MVB05, PP06, RAB08]. Strategy [CS00, GMM00, HHC98, KBC^{+}01, PAM94, RS92b, ASD09, BBM08, GRDB05, GLD06, Hsi04, LY91, LL07, LVP07, Ngo06, SK09, WCC02, ZV06]. Stream [HPT^{+}97, ARM^{+}05, CK08, Eİ07]. stream-based [ARM^{+}05]. streaming [HK05, XYDL06]. Streams [MM93, WUG99, LVP07, LY08]. Stretch [GG01]. String [BL94, RS90b, Kri91, MM07b]. strings [SCS^{+}08]. Striping [CT93]. Strongly [SZB92, MHPR05]. Structural [AGG98, SM92b]. Structure [DL99, FMP98, MNB95, PL98, Tze93, BB85a, CZ90, FGZ03, GV86, HK05, Lis90, MJ03, MSZ05, Par89, YC04]. structure-aware [HK05]. Structured [BE95, FBK98, KB01, Lun94, MRRV98, MNM98, WM92, CWLD05, FJG06, FKJG08, GA90, GWH06, IKS87, SZ09, WXZ05]. Structures [Ano96j, ADM^{+}94, CCRS92, DOP98, DRC90, Gup92, SIR92, ZM94a, FCG04, GZ08, HA05, LJ86, NCT^{+}07]. stub [WSG91]. Studies [GT02, HCAA93, SCB08]. Study [AAD02, BJ96, BA01b, BS96b, BS96c, Cha96, GK98, Hag97, HPT^{+}97, HBJ98, MS99a, NBP98, Oru94, QM01, RSD94, SSG93, SSRV94, WNA^{+}94, WLR90, YMR93, AP91b, Bad04, CBM^{+}08, CCK^{+}08, CT94, DI91, GRR^{+}05, HJ90a, HA91, LPX05a, TB90]. Style [SS00]. subclasses [CP04a]. Subcubes [SR95]. subdomain [CEGS07]. submachine [FPP06]. Submesh [SP96]. Subsequence [MS99b]. substitution [GPX08]. Substrate [KMKD97]. Substring [CB96]. Subsystem [GGD93]. subtasks [SSM^{+}06]. Subtree [DP00]. succinct [BHR91]. Sufficient [SJ96, AjHcC90]. Suffix [DP98, CS06a, GZ08]. suitable [PGS06]. Suited [PRS97, GS91b]. summary [Rob09]. summation [IHM05]. Summing [San02]. sums [HLS03]. Super [WLY01, SAOKZ05a, SAOKZ05b]. super- [SAOKZ05a, SAOKZ05b]. Supercomputer [CB02, GHS86, Ull84]. Supercomputers [AP93, CRV94, CP94, LF03]. Supercomputing [Ano96i, FR98, HRC09]. superconcentration [JL05]. Superconcurrent [NRS95]. supercube [SSB91]. SuperNode [AT94]. Superscalar [LPU97]. Superstabilizing [KUFM02]. Supertoroidal [DF95]. supervision [BPA06].

Support [AL99, AH94, CP99, FBK98, KR97, KC99a, LTH97, LFH^{+}03, MBL^{+}92, NS97, PL95, RPS93, TF92, BAL05, CCQ^{+}06, CCC^{+}04, CCK^{+}08, JBY^{+}05, RR05, SK91, SAB^{+}92, TYH09, ZWRI07]. Supporting [HA06, Sto87, WLNL06, BSW07, SKMM04]. Surface [CWW^{+}95, CY96]. surrogate [UAPM07]. Survey [BCH95a, GHKS98]. Survivable [HWWH08]. suspect [XYG07]. SVD [CL88, RS08, ZB97]. swap [FPP^{+}08]. Swapped [Par05, ZXP09]. Sweep [GGN93, DMCFCM03, KMP^{+}06]. Switch [ASH^{+}01, OK01, PD92, CL90, LHKL03, WLWW09]. switch-based [LHKL03, WLWW09]. Switchable [SB84]. Switched [CCR94, CS93c, GGN93, LK96, WB01, EB09, KYL05, Nap90, PYF08]. Switches [KJ84, PL93, TF92, MG09, PY09a, PY09b]. Switching [DRSB01, GB93, Guo94, LYL93, OY00, ST02, BMIM07, PL06, ST06]. Symbolic [YI96, CJY04]. Symmetric [BJ99, DHB02, DZDZ01, HOE^{+}09, HJ01, Kau94, Oru87, BC05, BW08, BB85b, EM89, KA03, VGAB08]. Symmetrical [IM94, QY94]. Symmetry [Kel00, HT90, MJ03]. Symposium [Wee01, Ros07, Sni03]. SYN [XCH08]. Synapse [Ram92]. Synchronization [ASB97, AGW98, ABP92, AH94, BA96, Cha95, FR92, GVA^{+}08, JLRA97, MRRV98, OKB95, PB95, RL96, RSS99, The02, WUG99, XMN92, CRA^{+}08, FZC^{+}05, HMBW07, HA06, LA06, PB09, TG04]. synchronized [JS86]. Synchronizing [DKMV01]. Synchronous [BCV05, CS95c, GV94, NSLK99, OY00, SKR93, Sch91, Soh96, FXW03, PK05a, WTC08a, WTC08b]. synchronously [SP90]. Synthesis [HLJ01, Lis90, PP92, HDT^{+}05, KKB^{+}06]. Synthesize [HLJ98]. Synthesizes [Ram92]. Synthesizing [SL89, Che86]. Synthetic [Pop91]. Sysplex [NKC^{+}97]. System [BK95, BBD^{+}91, BA01a, Bev02, BMM97, BJK^{+}96, CP92, CP99, DHR96, DSD^{+}97, DH95, DT92, FPD93, GH90, HBCM99, HCS^{+}00, HLL^{+}95, Kav93, KMB91, LP96b, Lu01, MWL00, MKY^{+}97, MBL^{+}92, MO97, MS96, NKC^{+}97, NsPPC02, SEP96, SG96, Tse95, UR94, wXH00, ZMPE00, dR09, ABC^{+}88, AMK^{+}07, BL05, BCK^{+}09, BMF05, BPP05, CBP02, Car95, CSW08, CCEB03, CDJ^{+}89, CK91, DS04a, DI91, DB86, EC89, Fer90, HJ90a, HM06, JW89, KCD08, KMF^{+}05, KC04, LFH^{+}03, LC91b, MM07a, MK08a, MC03, NAK04, NTC03, OEY07, PKN08, PK05b, SC04, SK91, SSL04, SM86, VD04, Wan06, WS06, YLB90, ZV09a, ZMC06, ZJ06, HCAA93]. System-Level [Kav93]. Systematic [IAS^{+}92, KK95, LB89, WAS88]. Systems [ASH^{+}01, AM97a, AM97b, AMN00, Ano92c, Ano02u, ADS98, Bah00, BBM^{+}02, BBR94, BPR99, BW95b, Bou02, BN02, BSB^{+}01, BS96b, BS96c, Cas93, CS93a, Cha94, CKK00, CY95, CK97, Cho93, CBdCD00, DSST95, DA97, DS96, DSW94, DAYA02, DG94, EMP^{+}96, FGKT97, FTC00, GCKM97, GM99, GRR93, GK93, GMM00, HKT^{+}91, HNM02, HlLLY95, HTL99, HM99, IM94, IK94, ISZBM99, JR95, JH92a, JF95, JSM94, JRR99, KS97a, KBC^{+}01, KCV99, KE93, KS93, KM91, KM92, LH92, LF92, LT94, MMRS98, MAS^{+}99, MT95, MMVR97, MM93, MRR^{+}02, MC93, Mir91, NSS97, NMS93, Nie94, NDZA99, OM84, PA96, PB99, PT01, Pov99, PP92, QY94, Raj01, RDS02, RAS96, SM94, Sch91, Ser97, SL95, SRGB90, SSRV94, Sun02, THN^{+}93, TH02, TY95, Wil92, WF93, WF96, WUG99, XH91, YH97, ZR00]. Systems [Zia92, ZM94b, van96, AL04, AOSM04, AOSM05, AFM09, AF06, AH06, BMB^{+}08, BB03, BW09, BRP03, BJS03, BK08, BS92, BD04, BPW05, CWLD05, CRK^{+}09, CF88, Car90, CCS06, CVJ09, CGW^{+}03, CI86, DK08, DFP06a, DGFGK05, DM04, DM90c, DQR^{+}09, DÖ06, DW04, DH91b, FJC04, GL89, GNT04, Gos90, GS91b, GWWL94, GC05, GBMZ07, GF89, HRC09, Hal05, HC09, HOE^{+}09, HCZ04, HS86, HA06, HP06, HA91, HA05, IS06, JSWB92, JMS86, JLM08, KME89, KUA07, KAS07, KL05, KMS^{+}06, Lai86, LTB02, LTL06, Lan09, LLL06, Lee90, LHF91, LHK03, LJ05, LZCY09, LZ05, LC90a, Li06b, LVP07, LW89, LCM^{+}06, LLS07, LM09, MS86, MTS90, MFVP08, MBH^{+}08, OS04, PMV05, PMV06, PRHB06, PTA08, PF91, RLH03, RN04, SDTD04, SP08, SYYU07, SS08, SCB09, SU87, She09, SCS^{+}08, SXZ06]. systems [SHLN09, SY04, SCJ^{+}08, ST05, TW89, TRSS06, TB90, UAKI06, WXZ05, WTC08a, WTC08b, WDDK09, WSG91, XHY07, XQ07, Yan04, YL89, ZZ90, ZFS07, ZV09b, Zim90, dG91]. Systolic [AMS94, BPST96, BMM97, BL90, CDR90, GE94, IPK85, KL84, LJ86, MM00, Meg91, MV94, MT97b, Ram92, TY90b, Tse90, Win85, WD92, CL85, Dja06, EL91, KT89, KH89, LB89, Lis90, MP88, PS88, Sch89b, ST87, ST89, THSS87, Ume85, WAS88, Zim90].

tables [ASD09, HKW05]. tag [CRK^{+}09]. Tagging [GHH92]. Taking [CL03b]. TAM [CGSV93]. Target [ERL90, WW07, YCC05]. target-driven [YCC05]. targets [BFKP04]. Task [AKPT99, AH06, CDY97, DA97, DDD98, DAYA02, DL99, DRST02, ERS90, FKKC97, FY97, HBCM99, HKT^{+}91, KLZ97, KA97, KA99, LL98, MSSE02, Moh97, SdS97, SZ00b, SCJ^{+}08, SS94a, SV00, SBKB90, SYG92, UAKI06, UR94, VS99, WSRM97, YCY^{+}00, BKS05, BD05, Bat05, DK08, DÖ06, KA05, LLL06, MCC04, PK05a, SWP90]. Task-Level [HKT^{+}91, SBKB90]. tasking [Lun90]. Tasks [ABM^{+}92, BSB^{+}01, DJ98, ERL90, Hag97, Lat95, LWY97, MAS^{+}99, MMVR97, NMS93, PS93, RDS02, Sin87, AOSM05, BH05, BSMH08, CDJ^{+}89, IKS87, KUA07, KSS^{+}07, KMS^{+}06, LHK03, Li06a, Li06b, LB09, LLS07, PK05a, RR05, SSM^{+}07, ZV09b]. Taxonomy [HM96, Sin93]. TDFL [SBKB90]. TDM [LLJ00b]. Technical [Ano93a]. Technique [BN94, CLV95, DAYA02, Fer95, KBG92, PM96, ZLPP01, CX05, DeG88, EE05, Nic88, PVGG06]. Techniques [ADM^{+}94, CS95b, Dah99, ELS94, FY97, Gil94, GS00, HlLLY95, HTL99, JSCB95, KGV94, NPY^{+}97, PA96, PYF08, RSS99, Tay02, UZSS96, AOSM04, CDB04, CDR09a, CD95, FM85, Gao89, GRR^{+}05, KA08, LP88, Pla08, Raj08, RG87, SFEF06, TZ07]. Technology [Ano02v, ER97, GC95, MKY^{+}97, MRR^{+}02, OB88, TMM06]. Telegraphos [KMKD97]. Telemedicine [CY99]. Telescience [PLL^{+}03]. Telescoping [KBC^{+}01]. template [RS90a]. Templates [ADS98, DP00]. Temporal [GSG^{+}93, Lo92, RJA97, VWHL96, BKS91]. temporary [Wan06]. Ten [TAS^{+}01, KA08]. Terabit [SH98]. Terminal [HHC98]. Terminating [Lin93c]. Termination [ASR93, CW93, HTB98, KHK03, Lai86, Ric98, Tse95, BFTV87, CV90, Eri88, MD07, MFVP08]. ternary [GNW03]. Test [GRS97, PKK91, Soh96, WW97, DWHL87]. test-and-treatment [DWHL87]. testbed [LBE03]. testbeds [VPHML06]. Testing [CY95, GFB^{+}92, GS99, KW02, WG93]. tests [Psa96]. Their [Kop97, BM08, SI86, TDM05]. Themes [RCY97]. Theoretic [ABJS01]. Theoretical [HC97]. Theory [CC08, DM90a, PTA08, VBM90, BDjQ86, BM08, GRDB05, Zim90]. thin [ST08a]. thread [SLG06, ST05]. Threaded [NS97, Kep03, CGSV93]. threading [Ngo06]. Threads [GSC96, LFA96, SEP96, TG99, DKRI09, PL03b]. threats [SFEF06]. Three [FCG04, FLS^{+}97, FT94, GG01, GH96, KR98, NEG85, PD92, SSG93, SSOB02, YMR93, LW06b, ZFS07]. Three-Dimensional [FLS^{+}97, KR98, NEG85, FCG04]. Three-Stage [FT94]. three-state [LW06b]. Threshold [CGA98, PAM94, Nik04]. Threshold-Based [CGA98]. throttle [XCH08]. Through-Wafer [MLW^{+}97]. Throughput [FM99b, HWC08, JSS92, BSW07, LMR05]. Throughput-coverage [HWC08]. Throwing [Tse95]. Tight [BBH^{+}98, FSZ07, Mat06, CH06a]. Tiling [AR97, CWW96, RS92a, Xue97, KSG03]. Time [AAL95, AK93, Ano92c, ADS01, BPJG92, BBM^{+}02, BA96, BM04a, BOSW94, BH93, BGOS95, BTZ98, BA01b, CW00, CS93a, Cha94, COS^{+}95, DP98, DS01, DJ98, DD95, EL97, EMP^{+}96, Fah96, FBK98, FY97, GS99, GMM00, HA92, JR95, JH92a, KF95b, KS97b, KEA95, LTWY95, LTY96, LPU97, LVR90, LM96, LAS^{+}97, LFA96, MMRS98, MT95, MMVR97, Mat93, MDD97, Moh97, MSST99, MS99b, Nas94, NIR86, NH93, NP09, OY00, OOW95, OS96b, OSZ98, PW96, Pel90, Pel95, PS93, PM96, PM92, QMCL94, RDS02, RU99, RAS96, Ric98, SCMB90, STN92, Sun02, THBF97, TVS97, WBTM09, WA02, WS97a, WLID02, ZLPP01, Zim96, van96, AOSM04, AOSM05, BNP02, BPP05, CH06a, CLL09, CLR90, CCN06, DHK04, EDÖ05, FKLB08, Gos90, GF89, GREC91, HOVC09, HA06, HL07, KRL87, KSG03]. time [LHK03, Lee03, LZCY09, Lis90, Lo92, MAM05, NA06, QJ05, RLH03, SI86, TBZB05, TZH^{+}06, VWHL96, VA07, Wan07, WTC08a, WTC08b, WL05, XO05, ACD^{+}93, CBP02, CX05]. Time-bounded [NP09]. Time-Division [QMCL94, ZLPP01]. Time-Efficient [EL97, MS99b]. Time-Optimal [BOSW94, OS96b, OSZ98, Pel90, Lis90]. Time-parallel [WBTM09]. time-sliced [KRL87]. Time-Step [CW00]. Time-Varying [KEA95]. Timed [NM95]. timeliness [ISM07]. times [SFT04]. timestamps [MS02]. Timing [ADS01, BSS99, CB99, Kar92, FVLB09, ISM07]. Timing-Driven [CB99]. TInMANN [VM95]. Title [Ano98l, Ano99h, Ano00c, Ano01h, Ano01i, Ano02d, Ano03b, Ano04a]. Tlib [RR05]. Toeplitz [BBd90, HM99, VGAB08]. Together [WLID02]. Token [AE95, BGJDL02, CP90, FFK97, GH96, HP00, YZY96, HSW04, PV07]. Token-Based [AE95, BGJDL02, HP00]. Token-Chasing [YZY96]. Tokens [SA93]. Tolerance [BSS97, Piu01, PM92, mYyF92, BDDL09, CWL^{+}07, CDR09a, LCC^{+}05, LH05, LP88, Pak89]. Tolerant [AE95, AM97a, AM95, BMM97, BW95b, BCH95b, CRV94, CL93, CC94, CF98, FM99b, GRR93, HGCC96, HTHH02, KP00, Lan94, LBT94, LC96, MD01, PB95, PKD97, SCC92, SS95, WIKC97, Wu94, YBOY97, ZYO02, AOSM05, BB87, BXA08, BPA06, BPP05, CL91a, CKN07, CDR09b, CMT92, CMS04, DH91b, FLPJ07, GNS09, LGG08, PR06, PL06, XCS06, mYA91, ZV09b, ZJ06]. Tolerate [VR95]. Tolerating [DT02, GS00, MG91]. tomography [FCG04, FGG08, PLL^{+}03]. Tool [BN94, DBKF90, ZNQ93, KKVI05, PF04, TD07]. Tools [Bal90, Cas93, MLC^{+}90, MSH90, NT90, MC03, YT05]. Top [Sch89b, TAS^{+}01]. Top-down [Sch89b]. topography [SK05a]. topography-aware [SK05a]. Topological [DC94, Par05, YN92, PL06]. Topologies [YZY96, YMG01, SL89]. Topology [CCM92, DS96, Seb95, WLY01, AP91b, DB08, GL90, LCW05, Seb91]. TOPSYS [BB93]. Tori [LHS97, MT93a, Man97, AB03a, GLD06]. Tornado [HK04]. toroidal [AB05]. Torus [CT96, RMC97, WB01, YMG01, RH05]. Total [CW00, CHC05, BCM06, BG05, Dim04, SL89]. TPC [DZDZ01]. TPC-C [DZDZ01]. traces [NRM^{+}09]. Tracing [RGS00, CDB04]. Track [MD01]. Tracking [BFKP04, WW07]. Trade [BCLR96, GK98, LPU97, CLR90]. Trade-Off [BCLR96, GK98, LPU97]. trade-offs [CLR90]. Tradeoff [TSHH01, HWC08]. traditional [BBCLL04]. Traffic [AA95, DSS95, FT94, KC95, LK94, OY00, TF92, FL86, FMM^{+}08, LK90, SAOKM03, SKMM04, WG08]. Training [LWOG02, SMKL93]. transaction [YWD08, Yan09]. Transactional [Gra09]. Transceiver [DKMV01]. Transfer [Lu01, CK06, LGG08]. Transfers [NSSS99]. Transform [BA95, CP91, DS01, Fer93, GZ97, HN91, JS94, CVJ09, DS04a, DPRW85, ESTA94, FSD04, SSL04, TKHG04, LLCL98]. Transformation [MG98, SC91b, WD92, FM85, GJG88, MRRT07]. Transformations [HBH93, OK02, JV09, Kan05]. transforms [TS91]. Transient [DT02, PAH^{+}98, GPT06a]. Transitive [AW95, YMR93]. Translating [FPP06]. translators [YLB90]. Transmission [DP99, JK00, HOVC09, OS04, OMSGNSG05]. transmitting [BR91a]. transparent [GVA^{+}08]. Transparently [AFT^{+}00]. Transport [GRS97, MSH90]. transportation [OO05]. Transpose [CT96, ZMPE00, SAOKM03]. Transposing [Swa98]. transposition [Ede91]. transputer [LC92]. TRAP [GRS97]. Traps [SD00]. Traversals [OOW95, Eİ07]. TreadMarks [LDCZ97]. treatment [DWHL87]. Tree [AAP01, AS96, BBR94, BM97, BCLR96, BE95, BF01, BS00, COS^{+}95, DVZ96, FA95, Goe94, GS01b, HR92a, KC99b, LPS^{+}98, OD95a, OOW95, PL94, SLP^{+}98, Ski96, Tze91, Wag94, BFG^{+}03, BC05, BBL04, EB09, FMM^{+}08, FJSW90, GA90, HSW04, tH90, IKS87, Mit07, OC07, PV07, Sch89a, SAF05, SK05b, TG03, Wu85, LZSL06]. tree-connected [HSW04]. Tree-Dags [BCLR96]. Tree-Related [OD95a]. tree-structured [GA90, IKS87]. Trees [AP94, AS94, ADS98, BBN93, BP02, CS95a, DM95, DP00, DLS00, DJM94, DLP99, DS93, Efe96, HKMU98, HM01, HS94a, HHC98, Iqb92, LP96a, MD98, PM92, ST02, SHL95, TT98, Wag93, WW96, WB01, WFL98, oPP00, BNP02, BL89, BMIM07, CI03, CS06a, CDR09a, Efe91, GZ08, GNW03, HPT07, LVP08, LMZ04, Lin03, LHT08, OMSGNSG05, PD05, PPC04, SKK91, TDM05, Wag89, WL90, WC91, YZLT09, Zep91]. Trellis [LCM^{+}06]. Trends [ER97]. Triangular [IK94]. Triangularization [KK86, CDR90, EM89]. Triangularizations [Par92]. Triangulation [DFRCU99, LS95]. Tridiagonal [CTZ99, Kau94, CK91, EM89, Gao86]. Tridiagonalization [BB85b, BW08]. trigger [FMR05]. trigger-broadcasting [FMR05]. true [CP04b]. trust [LAGK07]. trusted [SFEF06]. TrustGuard [SL06]. Truthful [WGS08]. Tunability [CKK00]. Tuning [SB02, HPT07, MML07]. tuple [DRT07]. Turbulence [LLCC02]. TWDM [LLJ00b]. Twisted [HTHH02, AP91b, FLPJ07]. Two [ABJS01, BNS00, BP01, Cha94, CCC92, CEF^{+}95, DD96, Gos90, GT97, Hwa97, KLZ97, KL84, LHS97, LP96b, LK94, LLCC02, NAK04, Qia97, RFPAG08, RP95, SSM89, SSHC00, YCY^{+}00, AB05, ARM^{+}05, CF88, CG86, Deh90, HDJ08, Hsi04, LC91b, PMV06, SS94b]. Two- [Hwa97]. Two-Dimensional [LP96b, YCY^{+}00, NAK04, AB05, Deh90, LC91b]. two-fixed-endpoint [Hsi04]. Two-Level [KL84, Qia97, RP95, SSHC00]. Two-pass [DD96]. two-stage [HDJ08]. Two-Variable [CCC92]. Two-Way [LK94, LLCC02]. Type [HO94, SC91b, BFH09, QGL^{+}09, MV94, MVV91]. TYPHOON [HKW05].

UC [BCD95]. UCT [AKPT99]. UDP [ZBF05]. UET [AKPT99]. UET-UCT [AKPT99]. UET/UET [AKPT99]. UET/UET-UCT [AKPT99]. ultra [RW02]. ultra-large-scale [RW02]. ultrametric [YZLT09]. unbalancing [MGG03]. unbounded [SP90]. Uncertainty [ADS01, ZC04]. Uncertainty-aware [ZC04]. understand [BČFF05]. Understanding [BDF92, DBKF90, NEG85, CDJ^{+}89, WRHR91]. Unfair [KY02]. unicast [SKMM04]. Unidirectional [KY02, KUFM02, RMC97]. unification [RM90]. Unified [AGG98, BL90, DM95, JBL02, KH89]. Uniform [AS94, BGJDL02, DR95, GM95, KY02, SR88b, TT98, TC96, VN93, Xue97, ZM94b, CLL09, LW06b, Mar88, MM07c]. Uniformization [DHK04, NH93]. Unifying [RCY97]. Union [KF95a]. unique [WCWH03]. Unit [AGW98, KNS91, KM88, QSL^{+}08]. Units [AM97a, AGG98, KL08b, SCB08]. universal [CS06b]. universality [SH89]. universioned [Ahu90]. unknown [MJ03]. Unlabeled [Man97]. UNPACK [BR96]. Unreliable [KB96a, AM06]. Unstructured [OB98, WCE97, ACFK07, MSZ05, YF09]. Unsupervised [BST01, DSAUM99]. untraceability [CL09]. unwinding [Nic88]. Update [GS96, LSH96, RTCG91]. Updating [JSM94, SDS99, KAP90]. upon [AFM09]. upper [GC07]. Usage [BS96a]. Use [BW96, BST01, Kar92, SV00, MSZ05, NAK04, SSMS08]. Used [LL95]. Useful [Bal90, GSG^{+}93, FM85]. Useless [Yen01]. User [GRS97, KOW97, RKK06, MAJJ05]. User-Level [KOW97, MAJJ05].

User-Space [GRS97]. Users [BST01, ZR00, SY04]. Using [AyJ93, BA97, BCLR96, BLG01, CCRS92, CP92, CB02, DS95a, DHB02, DMSH90, FR96a, FZVT02, FA95, HPT^{+}97, HK01, HS97, HC97, Hwa97, KJ84, KA97, Lat98, LMCF90, LPZ99, LFA96, LL98, MD98, MP96, MS86, Moh96, MFS93, NH93, NS92, NPY^{+}97, OS93, PH91, Par92, Par96, PKD97, SSG93, SM92a, SEP96, SP96, SM00, SD00, SL97, SIR92, SWC^{+}91, SKH96, Swa98, TSC01, TR96, WPKK94, WW96, WSRM97, WB01, WRC^{+}02, WS97a, WCYR08, XH91, YMG01, ZMPE00, AFM03, ASST05, Ara90, AK06, Bar05, BD05, BAMM05, BS92, CSWD03, CF88, CK08, CvdBL^{+}08, CKN07, CBM^{+}08, CDB04, CH06b, CMT92, CL85, DPRW85, DKRI09, DJT03, DH91b, DWHL87, EE05, Eİ07, FM07, FCS91, GZ08, GRDB05, GCS06, HC91, JP09, JZK04, KL08b, Kan05, KME09, KR06, KKB^{+}06, KA05]. using [LR03a, LA04, MM06, MS02, MK08b, MC03, NCTT09, PKN08, PBS08, PVG09, Pla08, SSM89, SSS07, SCB09, SCJ^{+}08, SA08, SK05a, SFEF06, SM08b, SL06, TMM06, UAPM07, Wu03, XCS06, ZV06, ZV09a, ZHO03]. utilities [AM06]. Utility [LL07, QH96, ASST05, CRL04]. Utility-based [LL07]. Utilization [AS91, LT96, CCHC09]. Utilizing [AM06, CM92, LA93].

V THR [DMB97]. valid [FZ90]. Validation [KM03]. Value [Bai94, HBH93, PE93, JBM91, Lin91, Luk85]. Values [HH01]. Variable [BL94, CCC92, IC05, MP08]. variable-length [MP08]. Variables [HV95, HS00, Hal05, HV09]. Variants [XL95]. Variation [YI96]. variational [MK08b]. Variety [WM92]. various [LW06a]. Varying [KEA95, PP96]. VCR [DSST95]. Vector [AMB95, CP94, GK98, GE94, LS85, MSC96, NFEG97, Ric98, Wol88, Yan93, YH97, AKD06, CLR90, CK91, ESTA94, GS03a, GHS86, KK88, LS88, MS02, Sch87]. Vectorizable [VH93]. vectorization [Wol88]. Vectors [TR96]. Vehicle [DH04]. VERDI [SRGB90]. Verification [AS00, BR95a, MB96a, Eri88, LAGK07]. Verifying [WG93]. Versatile [DVZ96]. Version [WW96]. versions [BSMH08]. versus [FBDC99, GST09, LPU97, Sun02, TSHH01]. vertically [LHF91, SM08a]. vertices [ACU08]. Very [OP96, DHK04]. via [BP98, CJ07, CVJ09, CRA^{+}08, MBR08, NS95, PRHB06, ZV06, ZBF05]. Viable [KLLK98]. victim [XCH08]. Video [AAL95, CLV95, DSST95, HLL^{+}95, JK00, RU99, ZRC99, Bar05, LVP07]. Video-on-Demand [DSST95, HLL^{+}95]. View [Buc92]. Views [CMT93, LMCF90, Won99, BB03]. Virtual [AD95, BAHP01, BF97, DRSB01, KS97a, KLLK98, KKS08, LM96, Mat93, PA97, PL95, TJ92, BJS03, BAL05, FX06, KS03, PY09a, PK05b, TT07, WDDK09, ZV06, ZJ06, DHS06]. Virtual-Channel [PA97]. virtualization [GTN^{+}06]. viruses [MJ03]. visibility [BSG90]. Vision [LR94, MBL^{+}92, MHC95, MAR87, WHT02, Kri91, WJD91]. vision/image [WJD91]. Visual [BN94, SRGB90]. Visualization [BB93, Cas93, Cou93, KS93, Mil93, NT90, MBH^{+}08, NCA93, TSD08, WGCZ09, ZB09, ZWRI07]. Visualizations [LSCA93, SK93]. Visualizing [RW93, SKR93, ZNQ93].

Vital [BS97, HHC98]. VLIW [dSR00]. VLSI [BB85a, BBR94, CCC90, FM85, GS91b, Gue86, KM97, KLL87, MB96a, MS87, ML89, MRR^{+}02, MT85, MT97b, NEG85, OB88, OT86, PR06, TU92, TF92, WSS93]. VLSI-suited [GS91b]. VM [JXW06]. VM-based [JXW06]. VoD [Bar05, LC07]. voice [WTS03]. voltage [FKLB08]. Volume [Ano92a, Ano92c, Ano93e, Ano96l, Ano97k, Ano00d, Ano01g, Ano01h, Ano01i, Ano02d, Ano03b, Ano04a, Ano08, Ano09, BS96c, CS93b, WS97a, ACFK07, Ano92b, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano99a, Ano99b, Ano99c, Ano00b, Ano00c]. Volumes [Ano98l, Ano99h]. Voronoi [RR95b, SZ03]. Voting [LO96, ZWS09]. vs [Wol88].

WAdL [GMS06]. Wafer [KL84, MLW^{+}97, RFM94]. Wafer-Scale [KL84, RFM94]. Wait [HPT02]. Wait-Free [HPT02]. Walk [SLP^{+}98]. Walks [BA01a]. warning [XCLR07]. Warp [ACD^{+}93, CBP02, CX05, PW96]. Warping [WS95, WS97a]. water [CvdBL^{+}08]. Watershed [MG98]. Wave [CDP95, CDB04]. wavefront [OT86]. Wavelength [HP00, MVM04]. wavelength-routed [MVM04]. Wavelet [HK01, CVJ09, TKHG04]. Wavelet-Based [HK01]. Way [LK94, LLCC02, ACU08, KK98a, VPHML06]. WDM [DP99, MVM04, OS93, WG08]. Weakest [Bit92]. weakly [HJ07]. Weather [RHH96]. Web [KCD08, AK06, BLPA05, CSWD03, SCK03, TC03, TC04, TK07, Wan06, XCZL03, XJS03, ZWL03]. Weight [RDL95, RGVB00, Tse95, YI96, LVP08, Wan06]. Weight-Throwing [Tse95]. Weighted [BS97]. weighting [CRA^{+}08]. well [EB09]. well-nested [EB09]. whole [Kan05]. whole-program [Kan05]. Wide [WM92, Wei98, HL07]. Wide-Area [Wei98]. width [DH91a]. window [LVP07]. window-assisted [LVP07]. winners [PL03a]. Wire [yHY97]. Wire-Limited [yHY97]. Wireless [BCD00, BD00, BDF01, Bou03, GMS06, JK00, KKGS01, MS00, Ola01, WL05, ASM09, AP03, BFG^{+}03, BSW07, BXA08, BPRS04, BOP06, BN03, BPA06, CKN07, CCK^{+}08, CLL09, CMS04, DW06, DMB^{+}03, DFP06b, EBE08, GDP08, GP07, GCY^{+}04, GMXA07, HMV07, HJ07, HWWH08, HWC08, KKVI05, KSI04, KOA09, KDH08, KNS06, LZ08, Lan09, LCW05, LW06a, LS03, LR03b, LLW07, Los08, MBR08, NC09, PLR07, SKMM04, SK05a, TBHA07, VHH08, WW07, WTB^{+}08, WMW09, WBTM09, WH08, XYKA08, YSL08]. Wireless/Mobile [MS00]. Wires [GO95]. within [THN^{+}93]. without [FSZ07, HP95, Ho91, MS02, OS97, SA93, XO05]. WK [DC94, SCD99]. WK-Recursive [DC94, SCD99]. WLANs [CCHC09, FA07]. Wolfe [Psa96]. work [BM04a]. worker [HSLL04]. workflow [FCC07]. workflows [TYH09]. Workload [DZDZ01, IM94, SSYG97, FGP05, GNT04, LLLY08, LF03].

Workloads [MKC01, CCQ^{+}06, CkLCK04, CkLCK05]. workshop [SAB^{+}92]. Workstation [AYI97, HN91, KMKD97, LC97, PN97a, PN97b, WB96, ME04]. Workstations [AS97, Ano00d, ABM^{+}92, BSS97, BDH^{+}97, CP97, CM92, DSAUM99, DZ97, HS97, HWW96, JLRA97, KR98, LS97, LHHB^{+}01, MDD97, NBSD99, PKD97, Ros99, ZLP97, BMARM07, CDB04, PY09c, Pla08]. world [FL86, MSZ05, MMS90]. worlds [WAE03]. Worm [NS95]. Wormhole [BLPV95, BPvW96, DG94, DRSB01, FF98, LME95, LEB98, NSSS99, PA97, RP98, RJMC95, RMC97, SJ95, SJ96, SB02, WB01, XMN92, HNSA07, Lee03, SAOKM03, WCC02]. Wormhole-Routed [FF98, NSSS99, RJMC95, RMC97, XMN92, SAOKM03, WCC02]. Wormhole-Switched [WB01]. Write [DS95a, CH06a, GNS09]. Writeback [KE93]. Writer [JBP00, KS97a, HV09, HV95]. wrong [SYYU07]. wrong-path [SYYU07]. WSNs [MCdS^{+}06]. WWW [AYI97, AYIE98].

X [CDH84, DS93]. X-MP-2 [CDH84]. X-Trees [DS93]. Xiaodong [Ano97k]. XML [CC08]. XOR [LL95].

Yan [Ano97k]. Yields [FY96]. Yong [Ano97k]. Yongsheng [Ano97k].

ZENTURIO [PF04]. zero [WCYR08]. Zhang [Ano97k]. zone [AGMJ06, JV06]. ZRAM [BMCP98].

References

Ahmed:1993:AOE

[AA93] Farid Ahmed and Abdul Ahad S. Awwal. An adaptive optoelectronic neural network for associative pattern retrieval. Journal of Parallel and Distributed Computing, 17(3):245–250, March 1993. CODEN JPDCER. ISSN 0743-7315 (print),

              1096-0848 (electronic). URL http://www.idealibrary.com/
              links/doi/10.1006/jpdc.1993.1023/production;   http:
              //www.idealibrary.com/links/doi/10.1006/jpdc.1993.1023/
              production/pdf.

Atiquzzaman:1995:PBM

[AA95] M. Atiquzzaman and M. S. Akhtar. Performance of buffered multistage interconnection networks in a nonuniform traffic environment. Journal of Parallel and Distributed Computing, 30(1):52–63, October 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

              URL http://www.idealibrary.com/links/doi/10.1006/
              jpdc.1995.1125/production;   http://www.idealibrary.
              com/links/doi/10.1006/jpdc.1995.1125/production/pdf.

Al-Ayyoub:1998:HIN

[AAD98] Abdel-Elah Al-Ayyoub and Khaled Day. The hyperstar interconnection network. Journal of Parallel and Distributed Computing, 48(2):175–199, February 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1414/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1414/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1414/production/ref.

Al-Ayyoub:2002:CSP

[AAD02] Abdel-Elah Al-Ayyoub and Khaled Day. Comparative study of product networks. Journal of Parallel and Distributed Computing, 62(1):1–18, January 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1770;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1770/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1770/ref.

Al-Ayyoub:2003:NRS

[AAD03] Abdel-Elah Al-Ayyoub and Khaled Day. Node-ranking schemes for the star networks. Journal of Parallel and Distributed Computing, 63(3):239–250, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Adle:2005:TAP

[AAD05] Roxane Adle, Marc Aiguier, and Franck Delaplace. Toward an automatic parallelization of sparse matrix computations. Journal of Parallel and Distributed Computing, 65(3):313–330, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agbaria:2004:QRP

[AAFV04] Adnan Agbaria, Hagit Attiya, Roy Friedman, and Roman Vitenberg. Quantifying rollback propagation in distributed checkpointing. Journal of Parallel and Distributed Computing, 64(3):370–384, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Akramullah:1995:DPA

[AAL95] Shahriar M. Akramullah, Ishfaq Ahmad, and Ming L. Liou. A data-parallel approach for real-time MPEG-2 video encoding. Journal of Parallel and Distributed Computing, 30(2):129–146, November 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1133/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1133/production/pdf.

Agarwal:2001:TPA

[AAP01] Ramesh C. Agarwal, Charu C. Aggarwal, and V. V. V. Prasad. A tree projection algorithm for generation of frequent item sets. Journal of Parallel and Distributed Computing, 61(3):350–371, March 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1693;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1693/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1693/ref.

Arvind:1984:RMF

[AB84] Arvind and Dean J. Brock. Resource managers in functional programming. Journal of Parallel and Distributed Computing, 1(1):5–21, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ayani:1993:PDE

[AB93] Rassul Ayani and Boris Berkman. Parallel discrete event simulation on SIMD computers. Journal of Parallel and Distributed Computing, 18(4):501–508, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1080/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1080/production/pdf.

AlBdaiwi:2003:RPT

[AB03a] Bader F. AlBdaiwi and Bella Bose. On resource placements in 3D tori. Journal of Parallel and Distributed Computing, 63(9): 838–845, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aluru:2003:GEI

[AB03b] Srinivas Aluru and David A. Bader. Guest Editor’s introduction: Special issue on high-performance computational biology. Journal of Parallel and Distributed Computing, 63(7–8):671– 673, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

AlBdaiwi:2005:QPR

[AB05] Bader F. AlBdaiwi and Bella Bose. Quasi-perfect resource placements for two-dimensional toroidal networks. Journal of Parallel and Distributed Computing, 65(7):815–831, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Allen:1988:OPA

[ABC^{+}88] F. Allen, M. Burke, P. Charles, R. Cytron, and J. Ferrante. An overview of the PTRAN analysis system for multiprocessing. Journal of Parallel and Distributed Computing, 5(5):617–640, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Antonopoulos:2009:ASH

[ABC^{+}09a] Christos D. Antonopoulos, Filip Blagojevic, Andrey N. Chernikov, Nikos P. Chrisochoides, and Dimitrios S. Nikolopoulos. Algorithm, software, and hardware optimizations for Delaunay mesh generation on simultaneous multithreaded architectures. Journal of Parallel and Distributed Computing, 69 (7):601–612, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Antonopoulos:2009:MDM

[ABC^{+}09b] Christos D. Antonopoulos, Filip Blagojevic, Andrey N. Chernikov, Nikos P. Chrisochoides, and Dimitrios S. Nikolopoulos. A multigrain Delaunay mesh generation method for multicore SMT-based architectures. Journal of Parallel and Distributed Computing, 69(7):589–600, July 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andrade:2007:AGA

[ABCM07] Nazareno Andrade, Francisco Brasileiro, Walfredo Cirne, and Miranda Mowbray. Automatic Grid assembly by promoting collaboration in peer-to-peer Grids. Journal of Parallel and Distributed Computing, 67(8):957–966, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Awerbuch:1996:FDN

[ABCP96] Baruch Awerbuch, Bonnie Berger, Lenore Cowen, and David Peleg. Fast distributed network decompositions and covers. Journal of Parallel and Distributed Computing, 39(2):105–114, December 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0159/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0159/production/pdf.

Adve:2002:COS

[ABDS02] Vikram S. Adve, Rajive Bagrodia, Ewa Deelman, and Rizos Sakellariou. Compiler-optimized simulation of large-scale applications on high performance architectures. Journal of Parallel and Distributed Computing, 62(3):393–426, March 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1800;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1800/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1800/ref.

Altman:2001:RTP

[ABJS01] Eitan Altman, Tamer Başar, Tania Jiménez, and Nahum Shimkin. Routing into two parallel links: Game-theoretic distributed algorithms. Journal of Parallel and Distributed Computing, 61(9):1367–1381, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1754;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1754/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1754/ref.

Atallah:1992:MAC

[ABM^{+}92] Mikhail J. Atallah, Christina Lock Black, Dan C. Marinescu, Howard Jay Siegel, and Thomas L. Casavant. Models and algorithms for coscheduling compute-intensive tasks on a network of workstations. Journal of Parallel and Distributed Computing, 16(4):319–327, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andrews:1992:NMN

[ABP92] John B. Andrews, Carl J. Beckmann, and David K. Poulsen. Notification and multicast networks for synchronization and coherence. Journal of Parallel and Distributed Computing, 15(4): 332–350, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abrams:1996:GPA

[Abr96] Marc Abrams. Geometric performance analysis of periodic behavior. Journal of Parallel and Distributed Computing, 33(2):131–144, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0032/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0032/production/pdf.

Arguello:1995:PAS

[ABZ95] Francisco Arguello, Javier D. Bruguera, and Emilio L. Zapata. A parallel architecture for the self-sorting FFT algorithm. Journal of Parallel and Distributed Computing, 31(1):88–97, November 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1147/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1147/production/pdf.

Akingbehin:1989:HAP

[AC89] Kiumi Akingbehin and Michael Conrad. A hybrid architecture for programmable computing and evolutionary learning. Journal of Parallel and Distributed Computing, 6(2):245–263, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Akyildiz:1993:EMC

[ACD^{+}93] Ian F. Akyildiz, Liang Chen, Samir Ranjan Das, Richard M. Fujimoto, and Richard F. Serfozo. The effect of memory capacity on Time Warp performance. Journal of Parallel and Distributed Computing, 18(4):411–422, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1074/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1074/production/pdf.

Aykanat:2007:ADR

[ACFK07] Cevdet Aykanat, B. Barla Cambazoglu, Ferit Findik, and Tahsin Kurc. Adaptive decomposition and remapping algorithms for object-space-parallel direct volume rendering of unstructured grids. Journal of Parallel and Distributed Computing, 67(1):77–99, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aykanat:2008:MLD

[ACU08] Cevdet Aykanat, B. Barla Cambazoglu, and Bora Uçar. Multilevel direct K-way hypergraph partitioning with multiple constraints and fixed vertices. Journal of Parallel and Distributed Computing, 68(5):609–625, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abderazek:2008:QPQ

[ACYS08] Ben A. Abderazek, Arquimedes Canedo, Tsutomu Yoshinaga, and Masahiro Sowa. The QC-2 parallel Queue processor architecture. Journal of Parallel and Distributed Computing, 68(2): 235–245, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agarwala:1995:ESV

[AD95] Amit Agarwala and Chita R. Das. Experimenting with a shared virtual memory environment for hypercubes. Journal of Parallel and Distributed Computing, 29(2):228–235, September 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1120/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1120/production/
pdf.

Afek:2002:LS

[AD02] Yehuda Afek and Shlomi Dolev. Local stabilizer. Journal of Parallel and Distributed Computing, 62(5):745–765, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1823;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1823/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1823/ref.

Atallah:1994:MTP

[ADM^{+}94] Mikhail J. Atallah, Frank Dehne, Russ Miller, Andrew Rau-Chaplin, and Jyh-Jong Tsay. Multisearch techniques: Parallel data structures on mesh-connected computers. Journal of Parallel and Distributed Computing, 20(1):1–13, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1001/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1001/production/
pdf.

Auletta:1998:MTA

[ADS98] Vincenzo Auletta, Amelia De Vivo, and Vittorio Scarano. Multiple templates access of trees in parallel memory systems. Journal of Parallel and Distributed Computing, 49(1):22–39, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1426/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1426/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1426/production/ref.

Attiya:2001:TBD

[ADS01] Hagit Attiya and Taly Djerassi-Shintel. Time bounds for decision problems in the presence of timing uncertainty and failures. Journal of Parallel and Distributed Computing, 61(8):1096–1109, August 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1730;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1730/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1730/ref.

Arvind:1988:FSP

[AE88] Arvind and K. Ekanadham. Future scientific programming on parallel machines. Journal of Parallel and Distributed Computing, 5(5):460–493, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agrawal:1995:TBF

[AE95] D. Agrawal and A. El Abbadi. A token-based fault-tolerant distributed mutual exclusion algorithm. Journal of Parallel and Distributed Computing, 24(2):164–176, February 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1016/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1016/production/pdf.

Al-Bassam:1992:PAH

[AERBL92] Sulaiman Al-Bassam, Hesham El-Rewini, Bella Bose, and T. G. Lewis. Processor allocation for hypercubes. Journal of Parallel and Distributed Computing, 16(4):394–401, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ardagna:2006:JOH

[AF06] Danilo Ardagna and Chiara Francalanci. Joint optimization of hardware and network systems. Journal of Parallel and Distributed Computing, 66(3):469–485, March 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aluru:2003:PBS

[AFM03] Srinivas Aluru, Natsuhiko Futamura, and Kishan Mehrotra. Parallel biological sequence comparison using prefix computations. Journal of Parallel and Distributed Computing, 63(3): 264–272, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arantes:2009:RGA

[AFM09] Gladstone M. Arantes, Jr., Felipe M. G. França, and Carlos A. Martinhon. Randomized generation of acyclic orientations upon anonymous distributed systems. Journal of Parallel and Distributed Computing, 69(3):239–246, March 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Alimonti:1996:FED

[AFS96] Paola Alimonti, Paola Flocchini, and Nicola Santoro. Finding the extrema of a distributed multiset. Journal of Parallel and Distributed Computing, 37(2):123–133, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0113/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0113/production/
pdf.

Aridor:2000:TOS

[AFT^{+}00] Yariv Aridor, Michael Factor, Avi Teperman, Tamar Eilam, and Assaf Schuster. Transparently obtaining scalability for Java applications on a cluster. Journal of Parallel and Distributed Computing, 60(10):1159–1193, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1649;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1649/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1649/ref.

Atallah:1986:EPS

[AG86] M. J. Atallah and M. T. Goodrich. Efficient parallel solutions to some geometric problems. Journal of Parallel and Distributed Computing, 3(4):492–507, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ahmad:1994:HSD

[AGF94] Ishfaq Ahmad, Arif Ghafoor, and Geoffrey C. Fox. Hierarchical scheduling of dynamic parallel computations on hypercube multicomputers. Journal of Parallel and Distributed Computing, 20(3):317–329, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1030/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1030/production/pdf.

Altman:1998:UFI

[AGG98] Erik R. Altman, R. Govindarajan, and Guang R. Gao. A unified framework for instruction scheduling and mapping for function units with structural hazards. Journal of Parallel and Distributed Computing, 49(2):259–293, March 15, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1442/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1442/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1442/production/ref.

Ayguade:2006:ENO

[AGMJ06] Eduard Ayguade, Marc Gonzalez, Xavier Martorell, and Gabriele Jost. Employing nested OpenMP for the parallelization of multi-zone computational fluid dynamics applications.

Journal of Parallel and Distributed Computing, 66(5):686–697, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Antonis:2004:HAD

[AGMS04] K. Antonis, J. Garofalakis, I. Mourtos, and P. Spirakis. A hierarchical adaptive distributed algorithm for load balancing. Journal of Parallel and Distributed Computing, 64(1):151–162, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abu-Ghazaleh:1998:MCU

[AGW98] Nael B. Abu-Ghazaleh and Philip A. Wilsey. Models for control unit synchronization on shared control architectures. Journal of Parallel and Distributed Computing, 52(1):69–81, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1468/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1468/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1468/production/ref.

Abu-Ghazaleh:2001:SCP

[AGW01] Nael B. Abu-Ghazaleh and Philip A. Wilsey. The shared control parallel architecture model. Journal of Parallel and Distributed Computing, 61(6):767–783, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1712;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1712/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1712/ref.

Anik:1994:PIS

[AH94] S. Anik and W. M. W. Hwu. Performance implications of synchronization support for parallel Fortran programs. Journal of Parallel and Distributed Computing, 22(2):202–215, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1081/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1081/production/
pdf.

Attiya:2006:TAM

[AH06] Gamal Attiya and Yskandar Hamam. Task allocation for maximizing reliability of distributed systems: a simulated annealing approach. Journal of Parallel and Distributed Computing, 66 (10):1259–1266, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ahuja:1990:CCM

[Ahu90] Mohan Ahuja. Concurrency control for a mix of universioned and multiversioned entities in distributed databases. Journal of Parallel and Distributed Computing, 10(3):233–252, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Alexandrov:1997:LIL

[AISS97] Albert Alexandrov, Mihai F. Ionescu, Klaus E. Schauser, and Chris Scheiman. LogGP: Incorporating long messages into the LogP model for parallel computation. Journal of Parallel and Distributed Computing, 44(1):71–79, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1346/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1346/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1346/production/ref.

Anger:1990:SSL

[AjHcC90] F. D. Anger, Jing jang Hwang, and Yuan chieh Chow. Scheduling with sufficient loosely coupled processors. Journal of Parallel and Distributed Computing, 9(1):87–92, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Alam:1993:RTO

[AK93] M. S. Alam and M. A. Karim. Real-time optical arithmetic/logical processing. Journal of Parallel and Distributed Computing, 17(3):251–258, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1024/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1024/production/pdf.

Artail:2006:MPI

[AK06] Hassan Artail and Elie Kahale. MAWS: a platform-independent framework for mobile agents using Web services. Journal of Parallel and Distributed Computing, 66(3):428–443, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anderson:2007:GLS

[AK07] James H. Anderson and Yong-Jik Kim. A generic local-spin fetch-and-ϕ-based mutual exclusion algorithm. Journal of Parallel and Distributed Computing, 67(5):551–580, May 2007. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arora:2006:RVC

[AKD06] Anish Arora, Sandeep S. Kulkarni, and Murat Demirbas. Resettable vector clocks. Journal of Parallel and Distributed Computing, 66(2):221–237, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arvind:1995:EPA

[AKP95] K. Arvind, V. Kamakoti, and C. Pandu Rangan. Efficient parallel algorithms for permutation graphs. Journal of Parallel and Distributed Computing, 26(1):116–124, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1052/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1052/production/pdf.

Andronikos:1999:OSU

[AKPT99] Theodore Andronikos, Nectarios Koziris, George Papakonstantinou, and Panayiotis Tsanakas. Optimal scheduling for UET/UET-UCT generalized n-dimensional grid task graphs. Journal of Parallel and Distributed Computing, 57(2):140– 165, May 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1530/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1530/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1530/production/ref.

Ali:2008:SHR

[AKSM08] Shoukat Ali, Jong-Kook Kim, Howard Jay Siegel, and Anthony A. Maciejewski. Static heuristics for robust resource allocation of continuously executing applications. Journal of Parallel and Distributed Computing, 68(8):1070–1080, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ahmad:1999:SIS

[AL99] Ishfaq Ahmad and Francis Lau. Special issue on software support for distributed computing: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 59(2):101–106, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1579/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1579/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1579/production/ref.

Aguilar:2004:DDL

[AL04] Jose L. Aguilar and E. L. Ernst L. Leiss. Data dependent loop scheduling based on genetic algorithms for distributed and shared memory systems. Journal of Parallel and Distributed Computing, 64(5):578–590, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Avudainayagam:2003:DDE

[ALF03] Arun Avudainayagam, Wenjing Lou, and Yuguang Fang. DEAR: a Device and Energy Aware Routing protocol for heterogeneous ad hoc networks. Journal of Parallel and Distributed Computing, 63(2):228–236, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agarwal:2009:FDP

[ALH^{+}09] Sachin Agarwal, Moshe Laifenfeld, Andrew Hagedorn, Ari Trachtenberg, and Murat Alanyali. Fair and distributed peer-to-peer allocation of a common, refillable resource. Journal of Parallel and Distributed Computing, 69(12):974–988, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ahmad:1999:DSM

[ALL99] Ishfaq Ahmad, William Y. M. Lai, and Bo Li. Dynamic scheduling of multimedia documents in a single server multiple clients environment. Journal of Parallel and Distributed Computing, 57(1):91–120, April 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1521/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1521/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1521/production/ref.

Albert:1991:DPC

[ALS91] Eugene Albert, Joan D. Lukas, and Guy L. Steele, Jr. Data parallel computers and the FORALL statement. Journal of Parallel and Distributed Computing, 13(2):185–192, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Alsuwaiyel:2001:PAP

[Als01] Muhammad H. Alsuwaiyel. A parallel algorithm for partitioning a point set to minimize the maximum of diameters. Journal of Parallel and Distributed Computing, 61(5):662–666, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1706;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1706/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1706/ref.

Aluru:1997:LFR

[Alu97] Srinivas Aluru. Lagged Fibonacci random number generators for distributed memory parallel computers. Journal of Parallel and Distributed Computing, 45(1):1–12, August 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1363/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1363/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1363/production/ref.

Antonio:1993:HMN

[AM93] John K. Antonio and Richard C. Metzger. Hypersphere mapper: a nonlinear programming approach to the hypercube embedding problem. Journal of Parallel and Distributed Computing, 19(3):262–270, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1110/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1110/production/pdf.

Alam:1995:CMF

[AM95] M. Sultan Alam and R. G. Melhem. Channel multiplexing in fault-tolerant modular multiprocessors. Journal of Parallel and Distributed Computing, 24(2):115–131, February 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1012/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1012/production/pdf.

Ahuja:1997:UCF

[AM97a] Mohan Ahuja and Shivakant Mishra. Units of computation in fault-tolerant distributed systems. Journal of Parallel and Distributed Computing, 40(2):194–209, February 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1277/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1277/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1277/production/ref.

Ajith:1997:IAM

[AM97b] Tom P. Ajith and C. Siva Ram Murthy. An improved algorithm for module allocation in distributed computing systems. Journal of Parallel and Distributed Computing, 42(1):82–90, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1302/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1302/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1302/production/ref.

Asaduzzaman:2006:UUP

[AM06] Shah Asaduzzaman and Muthucumaru Maheswaran. Utilizing unreliable public resources for higher profit and better SLA compliance in computing utilities. Journal of Parallel and Distributed Computing, 66(6):796–806, June 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Asaduzzaman:2007:SCP

[AM07] Shah Asaduzzaman and Muthucumaru Maheswaran. Strategies to create platforms for differentiated services from dedicated and opportunistic resources. Journal of Parallel and Distributed Computing, 67(10):1119–1134, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Mouhamed:1995:ELF

[AMB95] Mayez Al-Mouhamed and Lubomir Bic. Effects of loop fusion and statement migration on the speedup of vector multiprocessors. Journal of Parallel and Distributed Computing, 31(1):56–64, November 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1144/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1144/production/pdf.

Atreya:2007:EDL

[AMK^{+}07] Ranganath Atreya, Neeraj Mittal, Ajay D. Kshemkalyani, Vijay K. Garg, and Mukesh Singhal. Efficient detection of a locally stable predicate in a distributed system. Journal of Parallel and Distributed Computing, 67(4):369–385, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Mouhamed:2000:ASC

[AMN00] Mayez Al-Mouhamed and Homam Najjari. Adaptive scheduling of computations and communications on distributedmemory systems. Journal of Parallel and Distributed Computing, 60(6):716–740, June 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1628;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1628/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1628/ref.

Akl:1994:OSA

[AMS94] Selim G. Akl, Henk Meijer, and Ivan Stojmenovič. An optimal systolic algorithm for generating permutations in lexicographic order. Journal of Parallel and Distributed Computing, 20(1):84–91, January 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1008/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1008/production/pdf.

Annexstein:1994:EHR

[Ann94] F. Annexstein. Embedding hypercubes and related networks into mesh-connected processor arrays. Journal of Parallel and Distributed Computing, 23(1):72–79, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1120/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1120/production/pdf.

Anonymous:1992:AVN

[Ano92a] Anonymous. Addendum: Volume 14, number 2 (1992), pp. 105–120: A. Kapelnikov, R. R. Muntz, and M. D. Ercegovac, “A Methodology for Performance Analysis of Parallel Computations with Looping Constructs”. Journal of Parallel and Distributed Computing, 15(2):167, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1992:AIV

[Ano92b] Anonymous. Author index for Volume 16. Journal of Parallel and Distributed Computing, 16(4):402, December 1992. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1992:EVN

[Ano92c] Anonymous. Erratum: Volume 14, number 3 (1992), pp. 221– 235: Mark James and Doan Hoang, “Design of Low-Cost, Real-Time Simulation Systems for Large Neural Networks”. Journal of Parallel and Distributed Computing, 15(1):85, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [JH92a].

Anonymous:1993:AIT

[Ano93a] Anonymous. Announcement: IEEE Technical Committee on Parallel Processing. Journal of Parallel and Distributed Computing, 17(1–2):184, January/February 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1993:AIVa

[Ano93b] Anonymous. Author index for Volume 17. Journal of Parallel and Distributed Computing, 17(4):381, April 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1038/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1038/production/pdf.

Anonymous:1993:AIVb

[Ano93c] Anonymous. Author index for Volume 18. Journal of Parallel and Distributed Computing, 18(4):531, August 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1084/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1084/production/pdf.

Anonymous:1993:AIVc

[Ano93d] Anonymous. Author index for Volume 19. Journal of Parallel and Distributed Computing, 19(4):379, December 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1120/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1120/production/pdf.

Anonymous:1993:EVN

[Ano93e] Anonymous. Erratum: Volume 16, number 1 (1992): L. Cheng and A. A. Sawchuk, “A One-Copy Algorithm for 2-D Shuffles for Optical Omega Networks,” pp. 54–66. Journal of Parallel and Distributed Computing, 17(1–2):183, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [CS92].

Anonymous:1994:AIVa

[Ano94a] Anonymous. Author index for Volume 20. Journal of Parallel and Distributed Computing, 20(3):393, March 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1036/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1036/production/pdf.

Anonymous:1994:AIVb

[Ano94b] Anonymous. Author index for Volume 21. Journal of Parallel and Distributed Computing, 21(3):334, June 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1994.1064/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1064/production/pdf.

Anonymous:1994:AIVc

[Ano94c] Anonymous. Author index for Volume 22. Journal of Parallel and Distributed Computing, 22(3):571, September 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1113/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1113/production/pdf.

Anonymous:1994:AIVd

[Ano94d] Anonymous. Author index for Volume 23. Journal of Parallel and Distributed Computing, 23(3):484, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1160/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1160/production/pdf.

Anonymous:1994:EM

[Ano94e] Anonymous. Editorial message. Journal of Parallel and Distributed Computing, 23(3):263, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1138/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1138/production/pdf.

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index for Volume 24. Journal of Parallel and Distributed Computing, 24(2):230, February 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1023/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1023/production/pdf.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index for Volume 25. Journal of Parallel and Distributed Computing, 25(2):223, March 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1042/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1042/production/pdf.

Anonymous:1995:AIVc

[Ano95c] Anonymous. Author index for Volume 26. Journal of Parallel and Distributed Computing, 26(2):270, April 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1066/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1066/production/pdf.

Anonymous:1995:AIVd

[Ano95d] Anonymous. Author index for Volume 27. Journal of Parallel and Distributed Computing, 27(2):213, June 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1084/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1084/production/pdf.

Anonymous:1995:AIVe

[Ano95e] Anonymous. Author index for Volume 28. Journal of Parallel and Distributed Computing, 28(2):215, August 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1101/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1101/production/pdf.

Anonymous:1995:AIVf

[Ano95f] Anonymous. Author index for Volume 29. Journal of Parallel and Distributed Computing, 29(2):239, September 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1121/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1121/production/pdf.

Anonymous:1995:AIVg

[Ano95g] Anonymous. Author index for Volume 30. Journal of Parallel and Distributed Computing, 30(2):217, November 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1139/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1139/production/pdf.

Anonymous:1995:AIVh DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1159/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1159/production/pdf.

Anonymous:1995:CPSa

[Ano95i] Anonymous. Call for papers: Special issue. Journal of Parallel and Distributed Computing, 29(1):104, August 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1995:CPSb

[Ano95j] Anonymous. Call for papers: Special issues. Journal of Parallel and Distributed Computing, 29(2):236, September 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1995:EM

[Ano95k] Anonymous. Editorial message. Journal of Parallel and Distributed Computing, 24(1):1, January 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1001/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1001/production/pdf.

Anonymous:1996:AIVa

[Ano96a] Anonymous. Author index for Volume 32. Journal of Parallel and Distributed Computing, 32(2):233, February 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0018/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0018/production/pdf.

Anonymous:1996:AIVb

[Ano96b] Anonymous. Author index for Volume 33. Journal of Parallel and Distributed Computing, 33(2):214, March 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0040/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0040/production/pdf.

Anonymous:1996:AIVc

[Ano96c] Anonymous. Author index for Volume 34. Journal of Parallel and Distributed Computing, 34(2):234, May 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0061/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0061/production/pdf.

Anonymous:1996:AIVd

[Ano96d] Anonymous. Author index for Volume 35. Journal of Parallel and Distributed Computing, 35(2):219, June 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0084/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0084/production/pdf.

Anonymous:1996:AIVe

[Ano96e] Anonymous. Author index for Volume 36. Journal of Parallel and Distributed Computing, 36(2):206, August 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0102/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0102/production/pdf.

Anonymous:1996:AIVf

[Ano96f] Anonymous. Author index for Volume 37. Journal of Parallel and Distributed Computing, 37(2):221, September 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0123/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0123/production/pdf.

Anonymous:1996:AIVg

[Ano96g] Anonymous. Author index for Volume 38. Journal of Parallel and Distributed Computing, 38(2):289, November 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0149/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0149/production/pdf.

Anonymous:1996:AIVh DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0168/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0168/production/pdf.

Anonymous:1996:CPSb

[Ano96i] Anonymous. Call for papers: Special issue on irregular problems in supercomputing applications. Journal of Parallel and Distributed Computing, 39(2):192, December 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0167/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0167/production/pdf.

Anonymous:1996:CPSa

[Ano96j] Anonymous. Call for papers: Special issue on parallel and distributed data structures. Journal of Parallel and Distributed Computing, 39(2):191, December 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0166/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0166/production/pdf.

Anonymous:1996:EA

[Ano96k] Anonymous. Editorial announcement. Journal of Parallel and Distributed Computing, 35(1):1, May 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0062/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0062/production/pdf.

Anonymous:1996:EVN

[Ano96l] Anonymous. Erratum: Volume 31, number 2 (1995), pages 112–125, in the article “Deadlock Models and a General Algorithm for Distributed Deadlock Detection,” by Jerzy Brzezinski, Jean-Michel Helary, Michel Raynal, and Mukesh Singhal. Journal of Parallel and Distributed Computing, 32(2):232, February 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0017/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0017/production/pdf.

See [BHRS95].

Anonymous:1997:AIVa

[Ano97a] Anonymous. Author index for Volume 40. Journal of Parallel and Distributed Computing, 40(2):265, February 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1308/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1308/production/pdf.

Anonymous:1997:AIVb

[Ano97b] Anonymous. Author index for Volume 41. Journal of Parallel and Distributed Computing, 41(2):235–??, March 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1324/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1324/production/pdf.

Anonymous:1997:AIVc

[Ano97c] Anonymous. Author index for Volume 42. Journal of Parallel and Distributed Computing, 42(2):184, May 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1350/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1350/production/pdf.

Anonymous:1997:AIVd

[Ano97d] Anonymous. Author index for Volume 43. Journal of Parallel and Distributed Computing, 43(2):190, June 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1368/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1368/production/pdf.

Anonymous:1997:AIVe

[Ano97e] Anonymous. Author index for Volume 44. Journal of Parallel and Distributed Computing, 44(2):161, August 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1375/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1375/production/pdf.

Anonymous:1997:AIVf

[Ano97f] Anonymous. Author index for Volume 45. Journal of Parallel and Distributed Computing, 45(2):176, September 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1400/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1400/production/pdf.

Anonymous:1997:AIVg

[Ano97g] Anonymous. Author index for Volume 46. Journal of Parallel and Distributed Computing, 46(2):201, November 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1419/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1419/production/pdf.

Anonymous:1997:AIVh

[Ano97h] Anonymous. Author index for Volume 47. Journal of Parallel and Distributed Computing, 47(2):198, December 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1422/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1422/production/pdf.

Anonymous:1997:CP

[Ano97i] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 44(2):160, August 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1997:CPS

[Ano97j] Anonymous. Call for papers: Special issue of the Journal of Parallel and Distributed Computing on dynamic load balancing. Journal of Parallel and Distributed Computing, 41(1):137, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.9999/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.9999/production/pdf.

Anonymous:1997:VNA

[Ano97k] Anonymous. Volume 38, number 1 (1996), in the article “An Effective and Practicle Performance Prediction Model for Parallel Computing on Nondedicated Heterogeneous NOW,” by Yong Yan, Xiaodong Zhang, and Yongsheng Song, pages 63–80. Journal of Parallel and Distributed Computing, 41(2):234, March 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1312/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1312/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1312/production/ref.

Anonymous:1998:AIVa

[Ano98a] Anonymous. Author index for Volume 48. Journal of Parallel and Distributed Computing, 48(2):215, February 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1453/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1453/production/pdf.

Anonymous:1998:AIVb

[Ano98b] Anonymous. Author index for Volume 49. Journal of Parallel and Distributed Computing, 49(2):294, March 15, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1455/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1455/production/pdf.

Anonymous:1998:AIVc

[Ano98c] Anonymous. Author index for Volume 50. Journal of Parallel and Distributed Computing, 50(1–2):170, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1465/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1465/production/pdf.

Anonymous:1998:AIVd

[Ano98d] Anonymous. Author index for Volume 51. Journal of Parallel and Distributed Computing, 51(2):153, June 15, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1471/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1471/production/pdf.

Anonymous:1998:AIVe

[Ano98e] Anonymous. Author index for Volume 52. Journal of Parallel and Distributed Computing, 52(2):178, August 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1486/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1486/production/pdf.

Anonymous:1998:AIVf

[Ano98f] Anonymous. Author index for Volume 53. Journal of Parallel and Distributed Computing, 53(2):174, September 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1496/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1496/production/pdf.

Anonymous:1998:AIVg

[Ano98g] Anonymous. Author index for Volume 54. Journal of Parallel and Distributed Computing, 54(2):223, November 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1504/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1504/production/pdf.

Anonymous:1998:AIVh

[Ano98h] Anonymous. Author index for Volume 55. Journal of Parallel and Distributed Computing, 55(2):249, December 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1515/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1515/production/pdf.

Anonymous:1998:CPb

[Ano98i] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 51(2):151–152, June 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1998:CPc

[Ano98j] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 54(1):75–76, October 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1497/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1497/production/pdf.

Anonymous:1998:CPa

[Ano98k] Anonymous. Calls for papers. Journal of Parallel and Distributed Computing, 48(2):212–214, February 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:1998:CAT

[Ano98l] Anonymous. Cumulative author-title index for volumes 48–55. Journal of Parallel and Distributed Computing, 55(2):250–254, December 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1516/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1516/
production/pdf.

Anonymous:1999:AIVa

[Ano99a] Anonymous. Author index for Volume 55. Journal of Parallel and Distributed Computing, 56(3):296, March 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1535/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1535/production/pdf.

Anonymous:1999:AIVb

[Ano99b] Anonymous. Author index for Volume 58. Journal of Parallel and Distributed Computing, 58(3):548, September 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1588/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1588/production/pdf.

Anonymous:1999:AIVc

[Ano99c] Anonymous. Author index for Volume 59. Journal of Parallel and Distributed Computing, 59(3):489, December 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1609/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1609/production/pdf.

Anonymous:1999:CPa

[Ano99d] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 58(1):126–127, July 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1562/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1562/production/pdf.

Anonymous:1999:CPb

[Ano99e] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 59(1):98, October 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1590/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1590/production/pdf.

Anonymous:1999:CPc

[Ano99f] Anonymous. Call for papers. Journal of Parallel and Distributed Computing, 59(1):99–100, October 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1589/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1589/production/pdf.

Anonymous:1999:CPS

[Ano99g] Anonymous. Call for papers: Special issue of the Journal of Parallel and Distributed Computing on problems in parallel and distributed computing: Solutions based on evolutionary paradigms. Journal of Parallel and Distributed Computing, 58(3):546–547, September 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1580/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1580/production/pdf.

Anonymous:1999:CAT

[Ano99h] Anonymous. Cumulative author-title index for volumes 56–59. Journal of Parallel and Distributed Computing, 59(3):490–494, December 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1610/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1610/
production/pdf.

Anonymous:1999:E

[Ano99i] Anonymous. Editorial. Journal of Parallel and Distributed Computing, 56(1):1, January 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1517/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1517/production/pdf.

Anonymous:2000:ACP

[Ano00a] Anonymous. ANNOUNCEMENT: Call for papers. Journal of Parallel and Distributed Computing, 60(1):135–136, January 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1619/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1619/production/
pdf.

Anonymous:2000:AIV

[Ano00b] Anonymous. Author index for Volume 60. Journal of Parallel and Distributed Computing, 60(12):1561–1562, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1687;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1687/pdf.

Anonymous:2000:ATI

[Ano00c] Anonymous. Author-title index for Volume 60. Journal of Parallel and Distributed Computing, 60(12):1563–1567, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1688;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1688/pdf.

Anonymous:2000:EVN

[Ano00d] Anonymous. Erratum: Volume 59, number 1 (1999), in article jpdc.1999.1564 “Guidelines for Data-Parallel Cycle-Stealing in Networks of Workstations,” by Arnold L. Rosenberg, pages 31–53. Journal of Parallel and Distributed Computing, 60(1): 134, January 2000. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1608/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1608/
production/pdf. See [Ros99].

Anonymous:2000:PAFa

[Ano00e] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 60(9):1154, September 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1662;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1662/pdf.

Anonymous:2000:PAFb

[Ano00f] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 60(10):1354, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1670;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1670/pdf.

Anonymous:2000:PAFc

[Ano00g] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 60(11):1446, November 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1671;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1671/pdf.

Anonymous:2000:PAFd

[Ano00h] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 60(12):1568, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1689;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1689/pdf.

Anonymous:2001:Aa

[Ano01a] Anonymous. Announcement. Journal of Parallel and Distributed Computing, 61(6):??, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:Ab

[Ano01b] Anonymous. Announcement. Journal of Parallel and Distributed Computing, 61(7):??, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:ACPa

[Ano01c] Anonymous. Announcement: Call for papers. Journal of Parallel and Distributed Computing, 61(5):710–711, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1738;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1738/pdf.

Anonymous:2001:ACPb

[Ano01d] Anonymous. Announcement: Call for papers. Journal of Parallel and Distributed Computing, 61(6):851–852, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1763;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1763/pdf.

Anonymous:2001:ACPc

[Ano01e] Anonymous. Announcement: Call for papers: Special issue on mobile ad hoc networking and computing. Journal of Parallel and Distributed Computing, 61(7):997–998, July 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1766;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1766/pdf.

Anonymous:2001:AI

[Ano01f] Anonymous. Author index. Journal of Parallel and Distributed Computing, 61(12):??, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:AIV

[Ano01g] Anonymous. Author index for volume 61. Journal of Parallel and Distributed Computing, 61(12):1842–1843, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1795;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1795/pdf.

Anonymous:2001:ATIa

[Ano01h] Anonymous. Author-title index for volume 61. Journal of Parallel and Distributed Computing, 61(12):1844–1849, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1796;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1796/pdf.

Anonymous:2001:ATIb

[Ano01i] Anonymous. Author-title index for volume 61. Journal of Parallel and Distributed Computing, 61(12):??, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:GEIa

[Ano01j] Anonymous. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 61(9):??, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:GEIb

[Ano01k] Anonymous. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 61(11):??, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:GEIc

[Ano01l] Anonymous. Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 61(12):??, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:IP

[Ano01m] Anonymous. Invited paper. Journal of Parallel and Distributed Computing, 61(12):??, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFa

[Ano01n] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(1):150, January 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1697;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1697/pdf.

Anonymous:2001:PAFb

[Ano01o] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(2):279, February 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1699;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1699/pdf.

Anonymous:2001:PAFc

[Ano01p] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(3):447, March 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1709;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1709/pdf.

Anonymous:2001:PAFd

[Ano01q] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(4):565, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1731;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1731/pdf.

Anonymous:2001:PAFe

[Ano01r] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(5):709, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1736;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1736/pdf.

Anonymous:2001:PAFf

[Ano01s] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(6):850, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1764;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1764/pdf.

Anonymous:2001:PAFg

[Ano01t] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(6):??, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFh

[Ano01u] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(7):996, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1765;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1765/pdf.

Anonymous:2001:PAFi

[Ano01v] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(7):??, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFj

[Ano01w] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(8):1143, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1775;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1775/pdf.

Anonymous:2001:PAFk

[Ano01x] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(8):??, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFl

[Ano01y] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(9):1389, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1783;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1783/pdf.

Anonymous:2001:PAFm

[Ano01z] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(9):??, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFn

[Ano01-27] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(10):1505, October 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1786;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1786/pdf.

Anonymous:2001:PAFo

[Ano01-28] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(10):??, October 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFp

[Ano01-29] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(11):1705, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1793;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1793/pdf.

Anonymous:2001:PAFq

[Ano01-30] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(11):??, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:PAFr

[Ano01-31] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(12):1850, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1797;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1797/pdf.

Anonymous:2001:PAFs

[Ano01-32] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 61(12):??, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:P

[Ano01-33] Anonymous. Preface. Journal of Parallel and Distributed Computing, 61(7):??, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2001:RN

[Ano01-34] Anonymous. Research note. Journal of Parallel and Distributed Computing, 61(9):??, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:Aa

[Ano02a] Anonymous. Announcement. Journal of Parallel and Distributed Computing, 62(7):??, July 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1858;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1858/pdf.

Anonymous:2002:Ab

[Ano02b] Anonymous. Announcement. Journal of Parallel and Distributed Computing, 62(10):??, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1873;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1873/pdf.

Anonymous:2002:AI

[Ano02c] Anonymous. Author index. Journal of Parallel and Distributed Computing, 62(12):1783–1784, December 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:ATI

[Ano02d] Anonymous. Author-title index for volume. Journal of Parallel and Distributed Computing, 62(12):1785–1790, December 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:EBa

[Ano02e] Anonymous. Editorial board. Journal of Parallel and Distributed Computing, 62(11):CO2, November 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:EBb

[Ano02f] Anonymous. Editorial board. Journal of Parallel and Distributed Computing, 62(12):CO2, December 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:GEIa

[Ano02g] Anonymous. Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 62(3):??, March 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1794;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1794/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1794/ref.

Anonymous:2002:GEIb

[Ano02h] Anonymous. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 62(5):??, May 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1835;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1835/pdf.

Anonymous:2002:GEIc

[Ano02i] Anonymous. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 62(9):??, September 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1849;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1849/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1849/ref.

Anonymous:2002:N

[Ano02j] Anonymous. Note. Journal of Parallel and Distributed Computing, 62(10):??, October 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1847;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2002.1847/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1847/ref.

Anonymous:2002:PAa

[Ano02k] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(5):??, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1845;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1845/pdf.

Anonymous:2002:PAb

[Ano02l] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(6):??, June 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1846;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1846/pdf.

Anonymous:2002:PAc

[Ano02m] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(7):??, July 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1859;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1859/pdf.

Anonymous:2002:PAd

[Ano02n] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(8):??, August 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1865;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1865/pdf.

Anonymous:2002:PAe

[Ano02o] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(9):??, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1867;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1867/pdf.

Anonymous:2002:PAg

[Ano02p] Anonymous. Papers to appear. Journal of Parallel and Distributed Computing, 62(10):??, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1874;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1874/pdf.

Anonymous:2002:PAFa

[Ano02q] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 62(1):172, January 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1803;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1803/pdf.

Anonymous:2002:PAFb

[Ano02r] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 62(2):326, February 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1816;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1816/pdf.

Anonymous:2002:PAFc

[Ano02s] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 62(3):??, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1822;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1822/pdf.

Anonymous:2002:PAFd

[Ano02t] Anonymous. Papers to appear in forthcoming issues. Journal of Parallel and Distributed Computing, 62(4):??, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1836;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1836/pdf.

Anonymous:2002:SSD

[Ano02u] Anonymous. Self-stabilizing distributed systems. Journal of Parallel and Distributed Computing, 62(5):??, May 1, 2002.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2002:SIP

[Ano02v] Anonymous. Special issue on parallel and distributed discrete event simulation — an emerging technology. Journal of Parallel and Distributed Computing, 62(3):??, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. Journal of Parallel and Distributed Computing, 63(12):1326–1327, December 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:ATI

[Ano03b] Anonymous. Author-title index for volume. Journal of Parallel and Distributed Computing, 63(12):1328–1332, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBa

[Ano03c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(1):CO2, January 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBb

[Ano03d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(2):CO2, February 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBc

[Ano03e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(3):CO2, March 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBd

[Ano03f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(4):CO2, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBe

[Ano03g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(5):CO2, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBf

[Ano03h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(6):CO2, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBg

[Ano03i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(7–8):CO2, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBh

[Ano03j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(9):CO2, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBi

[Ano03k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(10):CO2, October 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBj

[Ano03l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(11):CO2, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2003:EBk

[Ano03m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 63(12):CO2, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:ATI

[Ano04a] Anonymous. Author–title index for volume. Journal of Parallel and Distributed Computing, 64(12):1416–1420, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:AI

[Ano04b] Anonymous. Author index. Journal of Parallel and Distributed Computing, 64(12):1414–1415, December 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:CA

[Ano04c] Anonymous. C3- AS008S. Journal of Parallel and Distributed Computing, 64(6):CO3, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:CSM

[Ano04d] Anonymous. A context-sensitive middleware for dynamic integration of mobile devices with network infrastructures– Publisher’s Note. Journal of Parallel and Distributed Computing, 64(6):692, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:CH

[Ano04e] Anonymous. Cover 3: HA03094L. Journal of Parallel and Distributed Computing, 64(11):CO3, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBa

[Ano04f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(1):CO2, January 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBb

[Ano04g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(2):CO2, February 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBc

[Ano04h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(3):CO2, March 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBd

[Ano04i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(4):CO2, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBe

[Ano04j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(5):CO2, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBf

[Ano04k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(6):CO2, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBg

[Ano04l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(7):CO2, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBh

[Ano04m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(8):CO2, August 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBi

[Ano04n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(9):CO2, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBj

[Ano04o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(10):CO2, October 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBk

[Ano04p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(11):CO2, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:EBl

[Ano04q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 64(12):CO2, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2004:MPA

[Ano04r] Anonymous. In memoriam: Professor Azriel Rosenfeld. Journal of Parallel and Distributed Computing, 64(4):569, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2008:EVR

[Ano08] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 68(12):1603–1608, December 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2009:EVR

[Ano09] Anonymous. End of volume reviewer index. Journal of Parallel and Distributed Computing, 69(12):996–999, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Asdre:2007:OPS

[ANP07] Katerina Asdre, Stavros D. Nikolopoulos, and Charis Papadopoulos. An optimal parallel solution for the path cover problem on P_{4}-sparse graphs. Journal of Parallel and Distributed Computing, 67(1):63–76, January 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andreae:1997:ECP

[ANS97] Thomas Andreae, Michael Nolle, and Gerald Schreiber. Embedding Cartesian products of graphs into de Bruijn graphs. Journal of Parallel and Distributed Computing, 46(2):194–200, November 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1384/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1384/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1384/production/ref.

Alba:2002:HCP

[ANT02] Enrique Alba, Antonio J. Nebro, and José M. Troya. Heterogeneous computing and parallel genetic algorithms. Journal of Parallel and Distributed Computing, 62(9):1362–1385, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1851;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1851/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1851/ref.

Abdullah:2005:DDA

[AOS^{+}05] Azizol Abdullah, Mohamed Othman, Md. Nasir Sulaiman, Hamidah Ibrahim, and Abu Talib Othman. Data discovery algorithm for scientific data Grid environment. Journal of Parallel and Distributed Computing, 65(11):1429–1434, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Omari:2004:EOT

[AOSM04] R. Al-Omari, A. K. Arun K. Somani, and G. Manimaran. Efficient overloading techniques for primary-backup scheduling in real-time systems. Journal of Parallel and Distributed Computing, 64(5):629–648, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Omari:2005:ASF

[AOSM05] R. Al-Omari, A. K. Somani, and G. Manimaran. An adaptive scheme for fault-tolerant scheduling of soft real-time tasks in multiprocessor systems. Journal of Parallel and Distributed Computing, 65(5):595–608, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abraham:1991:PMN

[AP91a] Seth Abraham and Krishnan Padmanabhan. Performance of multicomputer networks under pin-out constraints. Journal of Parallel and Distributed Computing, 12(3):237–248, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abraham:1991:TCT

[AP91b] Seth Abraham and Krishnan Padmanabhan. The twisted cube topology for multiprocessors: a study in network asymmetry. Journal of Parallel and Distributed Computing, 13(1):104–110, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andrews:1991:AAP

[AP91c] John B. Andrews and Constantine D. Polychronopoulos. An analytical approach to performance/cost modeling of parallel computers. Journal of Parallel and Distributed Computing, 12 (4):343–356, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Agrawal:1993:SIP

[AP93] Dharma P. Agrawal and Lalit M. Patnaik. Special issue on performance of supercomputers: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 19(3):143–147, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Alnuweiri:1994:EPC

[AP94] Hussein M. Alnuweiri and Viktor K. Prasanna. Efficient parallel computations on the reduced mesh of trees organization. Journal of Parallel and Distributed Computing, 20(2):121–135, February 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1013/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1013/production/pdf.

An:2003:GAP

[AP03] Beongku An and Symeon Papavassiliou. Geomulticast: architectures and protocols for mobile ad hoc wireless networks. Journal of Parallel and Distributed Computing, 63(2):182–195, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andonov:1997:OOT

[AR97] Rumen Andonov and Sanjay Rajopadhye. Optimal orthogonal tiling of 2-D iterations. Journal of Parallel and Distributed Computing, 45(2):159–165, September 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1371/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1371/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1371/production/ref.

Arabnia:1990:PAA

[Ara90] Hamid R. Arabnia. A parallel algorithm for the arbitrary rotation of digitized images using process-and-data-decomposition approach. Journal of Parallel and Distributed Computing, 10 (2):188–192, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arbib:1989:SNN

[Arb89] Michael A. Arbib. Schemas and neural networks for sixth generation computing. Journal of Parallel and Distributed Computing, 6(2):185–216, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Angelov:2005:EOT

[ARM^{+}05] Yavor Angelov, Umakishore Ramachandran, Kenneth Mackenzie, James Matthew Rehg, and Irfan Essa. Experiences with optimizing two stream-based applications for cluster execution. Journal of Parallel and Distributed Computing, 65(6):678–691, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aumann:1991:IMU

[AS91] Yonatan Aumann and Assaf Schuster. Improved memory utilization in deterministic PRAM simulation. Journal of Parallel and Distributed Computing, 12(2):146–151, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Atkinson:1994:UGB

[AS94] M. D. Atkinson and J. R. Sack. Uniform generation of binary trees in parallel. Journal of Parallel and Distributed Computing, 23(1):101–103, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1123/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1123/production/pdf.

Anderson:1995:PIP

[AS95] Richard Anderson and Joao C. Setubal. A parallel implementation of the Push-Relabel algorithm for the maximum flow problem. Journal of Parallel and Distributed Computing, 29(1):17–26, August 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1103/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1103/production/pdf.

Antonoiu:1996:SSL

[AS96] Gheorghe Antonoiu and Pradip K. Srimani. A self-stabilizing leader election algorithm for tree graphs. Journal of Parallel and Distributed Computing, 34(2):227–232, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0059/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0059/production/pdf.

Anastasiadis:1997:PAS

[AS97] Stergios V. Anastasiadis and Kenneth C. Sevcik. Parallel application scheduling on networks of workstations. Journal of Parallel and Distributed Computing, 43(2):109–124, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1335/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1335/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1335/production/ref.

Anand:2000:FVA

[AS00] K. C. Anand and R. K. Shyamasundar. Formal verification of activity-based specification of protocols. Journal of Parallel and Distributed Computing, 60(5):639–676, May 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1622;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1622/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1622/ref.

Abali:1997:CSM

[ASB97] Bulent Abali, Craig B. Stunkel, and Caroline Benveniste. Clock synchronization on a multicomputer. Journal of Parallel and Distributed Computing, 40(1):118–130, January 10, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1268/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1268/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1268/production/ref.

Alfaro:2009:NSM

[ASD09] Francisco J. Alfaro, José L. Sánchez, and José Duato. A new strategy to manage the InfiniBand arbitration tables. Journal of Parallel and Distributed Computing, 69(6):508–520, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abali:2001:ARN

[ASH^{+}01] Bulent Abali, Craig B. Stunkel, Jay Herring, Mohammad Banikazemi, Dhabaleswar K. Panda, Cevdet Aykanat, and Yucel Aydogan. Adaptive routing on the new switch chip for IBM SP systems. Journal of Parallel and Distributed Computing, 61(9):1148–1179, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1747;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1747/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1747/ref.

Akkaya:2009:CWS

[ASM09] Kemal Akkaya, Fatih Senel, and Brian McLaughlan. Clustering of wireless sensor and actor networks based on sensor distribution and connectivity. Journal of Parallel and Distributed Computing, 69(6):573–587, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abrams:1993:TOM

[ASR93] Marc Abrams, Vasant Sanjeevan, and Debra S. Richardson. Termination and output measure generation in parallel simulations. Journal of Parallel and Distributed Computing, 18(4):454–472, August 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1077/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1077/production/pdf.

Anvik:2005:AUC

[ASST05] John Anvik, Jonathan Schaeffer, Duane Szafron, and Kai Tan. Asserting the utility of CO_{2}P_{3}S using the Cowichan Problem Set. Journal of Parallel and Distributed Computing, 65(12):

1542–1557, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Adamo:1994:PEP

[AT94] J. M. Adamo and L. Trejo. Programming environment for phase-reconfigurable parallel programming on SuperNode. Journal of Parallel and Distributed Computing, 23(3):278–292, December 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1140/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1140/
production/pdf.

Achalakul:2003:DSS

[AT03] Tiranee Achalakul and Stephen Taylor. A distributed spectralscreening PCT algorithm. Journal of Parallel and Distributed Computing, 63(3):373–384, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Antonio:1991:HPA

[ATH91] John K. Antonio, Wei K. Tsai, and G. M. Huang. A highly parallel algorithm for multistage optimization problems and shortest path problems. Journal of Parallel and Distributed Computing, 12(3):213–222, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Al-Tawil:2001:PME

[ATM01] Khalid Al-Tawil and Csaba Andras Moritz. Performance modeling and evaluation of MPI. Journal of Parallel and Distributed Computing, 61(2):202–223, February 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1677;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1677/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1677/ref.

Ardagna:2007:SBR

[ATZ07] Danilo Ardagna, Marco Trubian, and Li Zhang. SLA based resource allocation policies in autonomic environments. Journal of Parallel and Distributed Computing, 67(3):259–270, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Aboelaze:1995:PAB

[AW95] Mokhtar A. Aboelaze and Benjamin W. Wah. A processor array with bounded I/O ports for computing transitive closures. Journal of Parallel and Distributed Computing, 29(1):84–90, August 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1108/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1108/production/pdf.

Abramson:1989:DCU

[AY89] Bruce Abramson and Moti Yung. Divide and conquer under global constraints: a solution to the N-queens problem. Journal of Parallel and Distributed Computing, 6(3):649–662, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Asher:2009:SLM

[AY09] Yosi Ben Asher and Moshe Yuda. Source level merging of independent programs. Journal of Parallel and Distributed Computing, 69(6):521–531, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Andresen:1997:TSD

[AYI97] Daniel Andresen, Tao Yang, and Oscar H. Ibarra. Toward a scalable distributed WWW server on workstation clusters. Journal of Parallel and Distributed Computing, 42(1):91–100, April 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1305/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1305/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1305/production/ref.

Andresen:1998:APS

[AYIE98] Daniel Andresen, Tao Yang, Oscar H. Ibarra, and Ömer Eğecioğlu. Adaptive partitioning and scheduling for enhancing WWW application performance. Journal of Parallel and Distributed Computing, 49(1):57–85, February 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1423/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1423/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1423/production/ref.

Abdullah:1993:CQP

[AyJ93] Naim Abdullah and Jie yong Juang. Concurrent query processing for logic inference using the connection graph. Journal of Parallel and Distributed Computing, 17(4):282–297, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1028/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1028/production/
pdf.

Alrabiah:2001:DCL

[AZ01] Tawfig Alrabiah and Taieb Znati. Delay-constrained, low-cost multicast routing in multimedia networks. Journal of Parallel and Distributed Computing, 61(9):1307–1336, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1752;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1752/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1752/ref.

Bultan:1992:NMH

[BA92] Tevfik Bultan and Cevdet Aykanat. A new mapping heuristic based on mean field annealing. Journal of Parallel and Distributed Computing, 16(4):292–305, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhandarkar:1995:HTR

[BA95] Suchendra M. Bhandarkar and Hamid R. Arabnia. The Hough transform on a reconfigurable multi-ring network. Journal of Parallel and Distributed Computing, 24(1):107–114, January 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1011/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1011/production/
pdf.

Bergmans:1996:CSR

[BA96] Lodewijk Bergmans and Mehmet Akšit. Composing synchronization and real-time constraints. Journal of Parallel and Distributed Computing, 36(1):32–52, July 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0088/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0088/production/pdf.

Ben-Asher:1997:ORM

[BA97] Yosi Ben-Asher. Optical routing in meshes using the duplication model. Journal of Parallel and Distributed Computing, 41(1):36–41, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1288/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1288/production/pdf.

Ben-Asher:2001:DRA

[BA01a] Yosi Ben-Asher. Distributed routing of ads and bids through random walks in the IDOS system. Journal of Parallel and Distributed Computing, 61(10):1456–1471, October 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1727;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1727/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1727/ref.

Bruda:2001:CSR

[BA01b] Stefan D. Bruda and Selim G. Akl. A case study in realtime parallel computation: Correcting algorithms. Journal of Parallel and Distributed Computing, 61(5):688–708, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1707;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1707/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1707/ref.

Battre:2006:MFP

[BA06] Dominic Battré and David Sigfredo Angulo. MPI framework for parallel searching in large biological databases. Journal of Parallel and Distributed Computing, 66(12):1503–1511, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bader:2004:IRA

[Bad04] David A. Bader. An improved, randomized algorithm for parallel selection with an experimental study. Journal of Parallel and Distributed Computing, 64(9):1051–1059, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ben-Asher:1992:DSA

[BAES92] Yosi Ben-Asher, David Egozi, and Assaf Schuster. 2-D SIMD algorithms for perfect shuffle networks. Journal of Parallel and Distributed Computing, 16(3):250–257, November 1992. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ben-Asher:1995:ESS

[BAGS95] Yosi Ben-Asher, Dan Gordon, and Assaf Schuster. Efficient self-simulation algorithms for reconfigurable arrays. Journal of Parallel and Distributed Computing, 30(1):1–22, October 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1122/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1122/production/
pdf.

Bahi:2000:AIA

[Bah00] Jacques M. Bahi. Asynchronous iterative algorithms for nonexpansive linear systems. Journal of Parallel and Distributed Computing, 60(1):92–112, January 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1587/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1587/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1587/production/ref.

Ben-Asher:2004:EPS

[BAH04] Yosi Ben-Asher and Gadi Haber. Efficient parallel solutions of linear algebraic circuits. Journal of Parallel and Distributed Computing, 64(1):163–172, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Banikazemi:2001:DAV

[BAHP01] Mohammad Banikazemi, Bulent Abali, Lorraine Herger, and Dhabaleswar K. Panda. Design alternatives for virtual interface architecture and an implementation on IBM Netfinity NT cluster. Journal of Parallel and Distributed Computing, 61(11):1512–1545, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1745;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1745/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1745/ref.

Bai:1994:PAC

[Bai94] Zhaojun J. Bai. A parallel algorithm for computing the generalized singular value decomposition. Journal of Parallel and Distributed Computing, 20(3):280–288, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1027/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1027/production/pdf.

Butt:2003:GCP

[BAK^{+}03] Ali Raza Butt, Sumalatha Adabala, Nirav H. Kapadia, Renato J. Figueiredo, and José A. B. Fortes. Grid-computing portals and security issues. Journal of Parallel and Distributed Computing, 63(10):1006–1014, October 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Balasundaram:1990:MKU

[Bal90] V. Balasundaram. A mechanism for keeping useful internal information in parallel programming tools: The data access descriptor. Journal of Parallel and Distributed Computing, 9 (2):154–170, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boukerche:2005:MVS

[BAL05] Azzedine Boukerche, Regina B. Araujo, and Marcelo Laffranchi. Multiuser 3D virtual simulation environments support in the Gnutella peer-to-peer network. Journal of Parallel and Distributed Computing, 65(11):1462–1469, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bic:1993:EUI

[BAM93] Lubomir Bic and Mayez Al-Mouhamed. The EM-4 under implicit parallelism. Journal of Parallel and Distributed Computing, 19(3):255–261, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1109/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1109/production/pdf.

Berger:2005:PNC

[BAMM05] M. J. Berger, M. J. Aftosmis, D. D. Marshall, and S. M. Murman. Performance of a new CFD flow solver using a hybrid programming paradigm. Journal of Parallel and Distributed Computing, 65(4):414–423, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ben-Asher:1991:PR

[BAPRS91] Y. Ben-Asher, D. Peleg, R. Ramaswami, and A. Schuster. The power of reconfiguration. Journal of Parallel and Distributed Computing, 13(2):139–153, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barlas:2005:VSO

[Bar05] Gerassimos Barlas. VoD on steroids: Optimized content delivery using distributed video servers over best-effort Internet. Journal of Parallel and Distributed Computing, 65(9):1057– 1071, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Basermann:1997:CGL

[Bas97] A. Basermann. Conjugate gradient and Lanczos methods for sparse matrices on distributed memory multiprocessors. Journal of Parallel and Distributed Computing, 45(1):46–52, August 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1364/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1364/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1364/production/ref.

Bertier:2006:DME

[BAS06] Marin Bertier, Luciana Arantes, and Pierre Sens. Distributed mutual exclusion algorithms for Grid applications: a hierarchical approach. Journal of Parallel and Distributed Computing, 66(1):128–144, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bataineh:2005:TAS

[Bat05] Sameer M. Bataineh. Toward an analytical solution to task allocation, processor assignment, and performance evaluation of network processors. Journal of Parallel and Distributed Computing, 65(1):29–47, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertolazzi:1985:VSD

[BB85a] P. Bertolazzi and G. Bongiovanni. A VLSI structure for the deadlock avoidance problem. Journal of Parallel and Distributed Computing, 2(4):352–361, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bojanczyk:1985:TSM

[BB85b] A. Bojanczyk and R. P. Brent. Tridiagonalization of a symmetric matrix on a square array of mesh-connected processors. Journal of Parallel and Distributed Computing, 2(3):261–276, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Balasubramanian:1987:FTM

[BB87] Vijay Balasubramanian and Prithviraj Banerjee. A fault tolerant massively parallel processing architecture. Journal of Parallel and Distributed Computing, 4(4):363–383, August 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bemmerl:1993:VMP

[BB93] Thomas Bemmerl and Peter Braum. Visualization of message passing parallel programs with the TOPSYS parallel programming environment. Journal of Parallel and Distributed Computing, 18(2):118–128, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1993.1051/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1051/production/pdf.

Bartoli:2003:ABD

[BB03] Alberto Bartoli and Ozalp Babaoglu. Application-based dynamic primary views in asynchronous distributed systems. Journal of Parallel and Distributed Computing, 63(4):410–433, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertrand:2006:DRR

[BBB^{+}06] Felipe Bertrand, Randall Bramley, David E. Bernholdt, James A. Kohl, Alan Sussman, Jay W. Larson, and Kostadin B. Damevski. Data redistribution and remote method invocation for coupled components. Journal of Parallel and Distributed Computing, 66(7):931–946, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Belkouch:2002:SSD

[BBCD02] Fatima Belkouch, Marc Bui, Liming Chen, and Ajoy K. Datta. Self-stabilizing deterministic network decomposition. Journal of Parallel and Distributed Computing, 62(4):696–714, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1811;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1811/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1811/ref.

Balle:2004:ETD

[BBCLL04] Susanne M. Balle, Bevin R. Brett, C.-P. Chih-Ping Chen, and David LaFrance-Linden. Extending a traditional debugger to debug massively parallel applications. Journal of Parallel and Distributed Computing, 64(5):617–628, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bojanczyk:1990:LCA

[BBd90] A. W. Bojanczyk, R. P. Brent, and F. R. de Hoog. Linearly connected arrays for Toeplitz least-squares problems. Journal of Parallel and Distributed Computing, 9(3):261–270, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bahr:1991:PPS

[BBD^{+}91] E. Bahr, F. Barachini, J. Doppelbauer, H. Gräbner, F. Kasparec, T. Mandl, and H. Mistelberger. A parallel production system architecture. Journal of Parallel and Distributed Computing, 13(4):456–462, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bal:1997:PHL

[BBH^{+}97] Henry Bal, Raoul Bhoedjang, Rutger Hofman, Ceriel Jacobs, Koen Langendoen, Tim Rühl, and Kees Verstoep. Performance of a high-level parallel language on a highspeed network. Journal of Parallel and Distributed Computing, 40(1):49–64, January 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1265/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1265/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1265/production/ref.

Bhatt:1998:TBP

[BBH^{+}98] Sandeep N. Bhatt, Gianfranco Bilardi, Kieran T. Herley, Geppino Pucci, and Abhiram Ranade. Tight bounds on parallel list marking. Journal of Parallel and Distributed Computing, 51(2):75–88, June 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1447/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1447/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1447/production/ref.

Bui:2004:DAC

[BBL04] Marc Bui, Franck Butelle, and Christian Lavault. A distributed algorithm for constructing a minimum diameter spanning tree. Journal of Parallel and Distributed Computing, 64(5):571–577, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Benner:2002:PAL

[BBM^{+}02] Peter Benner, Ralph Byers, Rafael Mayo, Enrique S. Quintana-Ortí, and Vicente Hernández. Parallel algorithms for LQ optimal control of discrete-time periodic linear systems. Journal of Parallel and Distributed Computing, 62(2):306–325, February 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1790;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1790/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1790/ref.

Batista:2008:PSB

[BBM08] Rodolfo Bezerra Batista, Azzedine Boukerche, and Alba Cristina Magalhaes Alves de Melo. A parallel strategy for biological sequence alignment in restricted memory space. Journal of Parallel and Distributed Computing, 68(4):548–561, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barak:1993:BCF

[BBN93] Amnon Barak and Ron Ben-Natan. Bounded contractions of full trees. Journal of Parallel and Distributed Computing, 17(4):363–369, April 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1035/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1035/production/pdf.

Bertossi:1994:RTA

[BBR94] A. A. Bertossi, M. A. Bonuccelli, and M. Roccetti. Reconfigurable tree architectures for gracefully degradable VLSI systems. Journal of Parallel and Distributed Computing, 23(3):264–277, December 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1139/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1139/production/pdf.

Bruno:1994:IAD

[BC94] J. Bruno and P. Cappello. Implementing the 3D alternating direction method on the hypercube. Journal of Parallel and Distributed Computing, 23(3):411–417, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1994.1150/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1150/production/pdf.

Bianchini:2001:DEC

[BC01] Ricardo Bianchini and Enrique V. Carrera. Designing and evaluating a cost-effective optical network for multiprocessors. Journal of Parallel and Distributed Computing, 61(2):178–201, February 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1673;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1673/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1673/ref.

Bader:2005:FPS

[BC05] David A. Bader and Guojing Cong. A fast, parallel spanning tree algorithm for symmetric multiprocessors (SMPs). Journal of Parallel and Distributed Computing, 65(9):994–1006, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bader:2006:FSM

[BC06] David A. Bader and Guojing Cong. Fast shared-memory algorithms for computing the minimum spanning forest of sparse graphs. Journal of Parallel and Distributed Computing, 66(11): 1366–1378, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baldoni:1995:CGA

[BCC95] Roberto Baldoni, Bruno Ciciani, and Giacomo Cioffi. On the correctness of Goscinski’s algorithm. Journal of Parallel and Distributed Computing, 27(2):201–204, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1082/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1082/production/pdf.

Bagrodia:1995:USB

[BCD95] Rajive Bagrodia, Mani Chandy, and Maneesh Dhagat. UC: a set-based language for data-parallel programming. Journal of Parallel and Distributed Computing, 28(2):186–201, August 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1099/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1099/production/
pdf.

Bononi:2000:DPE

[BCD00] Luciano Bononi, Marco Conti, and Lorenzo Donatiello. Design and performance evaluation of a distributed contention control (DCC) mechanism for IEEE 802.11 wireless local area networks. Journal of Parallel and Distributed Computing, 60(4):407–430, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1612;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1612/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1612/ref.

Bozkus:1994:CFH

[BCF^{+}94] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, Sanjay Ranka, and Min-You Wu. Compiling Fortran 90D/HPF for distributed memory MIMD computers. Journal of Parallel and Distributed Computing, 21(1):15–26, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1039/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1039/production/pdf.

Barthou:1997:FAD

[BCF97] Denis Barthou, Jean-François Collard, and Paul Feautrier. Fuzzy array dataflow analysis. Journal of Parallel and Distributed Computing, 40(2):210–226, February 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1261/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1261/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1261/production/ref.

Bazterra:2005:GFU

[BČFF05] Victor E. Bazterra, Martin Čuma, Marta B. Ferraro, and Julio C. Facelli. A general framework to understand parallel performance in heterogeneous clusters: analysis of a new adaptive parallel genetic algorithm. Journal of Parallel and Distributed Computing, 65(1):48–57, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bermond:1995:DLC

[BCH95a] J. C. Bermond, F. Comellas, and D. F. Hsu. Distributed loop computer networks: a survey. Journal of Parallel and Distributed Computing, 24(1):2–10, January 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1002/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1002/production/pdf.

Bruck:1995:CFT

[BCH95b] Jehoshua Bruck, Robert Cypher, and Ching-Tien T. Ho. On the construction of fault-tolerant cube-connected cycles networks. Journal of Parallel and Distributed Computing, 25(1):98–106, February 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1032/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1032/production/pdf.

Badishi:2009:DFC

[BCK^{+}09] Gal Badishi, Germano Caronni, Idit Keidar, Raphael Rom, and Glenn Scott. Deleting files in the Celeste peer-to-peer storage system. Journal of Parallel and Distributed Computing, 69 (7):613–622, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhatt:1996:STD

[BCLR96] Sandeep N. Bhatt, Fan R. K. Chung, F. Thomas Leighton, and Arnold L. Rosenberg. Scheduling tree-dags using FIFO queues: a control-memory trade-off. Journal of Parallel and Distributed Computing, 33(1):55–68, February 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0024/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0024/production/pdf.

Bailey:1987:RCO

[BCM87] Duane A. Bailey, Janice E. Cuny, and Bruce B. MacLeod. Reducing communication overhead: a parallel code optimization. Journal of Parallel and Distributed Computing, 4(5):505–520, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baldoni:2006:CTO

[BCM06] Roberto Baldoni, Stefano Cimmino, and Carlo Marchetti. A classification of total order specifications and its application to fixed sequencer-based implementations. Journal of Parallel and Distributed Computing, 66(1):108–127, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bordawekar:1996:CCS

[BCR96] Rajesh Bordawekar, Alok Choudhary, and J. Ramanujam. Compilation and communication strategies for out-of-core programs on distributed memory machines. Journal of Parallel and Distributed Computing, 38(2):277–288, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0148/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0148/production/pdf.

Barragy:1994:PSF

[BCV94] E. Barragy, G. F. Carey, and R. Van de Geijn. Performance and scalability of finite element analysis for distributed parallel computation. Journal of Parallel and Distributed Computing, 21(2):202–212, May 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1052/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1052/production/pdf.

Bahi:2005:SDL

[BCV05] Jacques Bahi, Raphaël Couturier, and Flavien Vernier. Synchronous distributed load balancing on dynamic networks. Journal of Parallel and Distributed Computing, 65(11):1397– 1405, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Blelloch:1995:SLR

[BCZ95] Guy E. Blelloch, Siddhartha Chatterjee, and Marco Zagha. Solving linear recurrences with loop raking. Journal of Parallel and Distributed Computing, 25(1):91–97, February 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1031/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1031/production/
pdf.

Boukerche:2000:GEI

[BD00] Azzedine Boukerche and Sajal K. Das. Guest Editors’ introduction: Special issue on wireless and mobile computing and communications. Journal of Parallel and Distributed Computing, 60(4):349–352, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1620;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1620/pdf.

Boukerche:2004:RNM

[BD04] Azzedine Boukerche and Sajal K. Das. Reducing null messages overhead through load balancing in conservative distributed simulation systems. Journal of Parallel and Distributed Computing, 64(3):330–344, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baskiyar:2005:SDC

[BD05] Sanjeev Baskiyar and Christopher Dickinson. Scheduling directed a-cyclic task graphs on a bounded set of heterogeneous processors using task duplication. Journal of Parallel and Distributed Computing, 65(8):911–921, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bosilca:2009:ABF

[BDDL09] George Bosilca, Rémi Delmas, Jack Dongarra, and Julien Langou. Algorithm-based fault tolerance applied to high performance computing. Journal of Parallel and Distributed Computing, 69(4):410–416, April 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Balbo:1992:UPP

[BDF92] Gianfranco Balbo, Susanna Donatelli, and Giuliana Franceschinis. Understanding parallel program behavior through Petri net models. Journal of Parallel and Distributed Computing, 15 (3):171–187, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boukerche:2001:ARC

[BDF01] Azzedine Boukerche, Sajal K. Das, and Alessandro Fabbri. Analysis of a randomized congestion control scheme with DSDV routing in ad hoc wireless networks. Journal of Parallel and Distributed Computing, 61(7):967–995, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1723;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1723/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1723/ref.

Bruck:1997:EMP

[BDH^{+}97] Jehoshua Bruck, Danny Dolev, Ching-Tien Ho, Marcel-Cătălin Roşu, and Ray Strong. Efficient message passing interface (MPI) for parallel computing on clusters of workstations. Journal of Parallel and Distributed Computing, 40(1):19–34, January 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1267/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1267/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1267/production/ref.

Blevins:1990:BHI

[BDHF90] D. W. Blevins, E. W. Davis, R. A. Heaton, and J. H. Feif. BLITZEN: a highly integrated massively parallel machine. Journal of Parallel and Distributed Computing, 8(2):150–160, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bermond:1986:SIN

[BDjQ86] J.-C. Bermond, C. Delorme, and J. j. Quisquater. Strategies for interconnection networks: some methods from graph theory. Journal of Parallel and Distributed Computing, 3(4):433–449, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bajard:1994:SOL

[BDKM94] J.-C. Bajard, J. Duprat, S. Kla, and J.-M. Muller. Some operators for on-line radix-2 computations. Journal of Parallel and Distributed Computing, 22(2):336–345, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1093/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1093/production/pdf.

Billionnet:1995:AFB

[BE95] Alain Billionnet and Sourour Elloumi. An algorithm for finding the K-best allocations of a tree structured program. Journal of Parallel and Distributed Computing, 26(2):225–232, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1061/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1061/production/
pdf.

Beckmann:1996:GSS

[Bec96] Carl J. Beckmann. Generating self-scheduling code for nonloop parallelism. Journal of Parallel and Distributed Computing, 39(2):126–145, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0161/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0161/production/pdf.

Bischof:2000:PLB

[BEE00] Stefan Bischof, Ralf Ebner, and Thomas Erlebach. Parallel load balancing for problems with good bisectors. Journal of Parallel and Distributed Computing, 60(9):1047–1073, September 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1647;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1647/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1647/ref.

Bhattacharya:1994:MGM

[BETD94] S. Bhattacharya, G. Elsesser, W.-T. Tsai, and D.-Z. Du. Multicasting in generalized multistage interconnection networks. Journal of Parallel and Distributed Computing, 22(1):80–95, July 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1071/production;   http:

//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1071/
production/pdf.

Bevilacqua:2002:MAP

[Bev02] Alessandro Bevilacqua. A methodological approach to parallel simulated annealing on an SMP system. Journal of Parallel and Distributed Computing, 62(10):1548–1570, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1863;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1863/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1863/ref.

Bilas:1997:FRS

[BF97] Angelos Bilas and Edward W. Felten. Fast RPC on the SHRIMP virtual memory mapped network interface. Journal of Parallel and Distributed Computing, 40(1):138–146, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.1272/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.1272/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1272/production/ref.

Bradford:2001:CPD

[BF01] Jeffrey P. Bradford and José A. B. Fortes. Characterization and parallelization of decision-tree induction. Journal of Parallel and Distributed Computing, 61(3):322–349, March 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1692;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1692/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1692/ref.

Bauer:1994:PDF

[BFG94] M. A. Bauer, S. T. Feeney, and I. Gargantini. Parallel 3-D filling with octrees. Journal of Parallel and Distributed Computing, 22(1):121–128, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1075/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1075/production/pdf.

Baala:2003:SSD

[BFG^{+}03] H. Baala, O. Flauzac, J. Gaber, M. Bui, and T. El-Ghazawi. A self-stabilizing distributed algorithm for spanning tree construction in wireless ad hoc networks. Journal of Parallel and Distributed Computing, 63(1):97–104, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bilo:2004:EAO

[BFG04] Vittorio Bilò, Michele Flammini, and Roberto Giovannelli. Experimental analysis of online algorithms for the bicriteria scheduling problem. Journal of Parallel and Distributed Computing, 64(9):1086–1100, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Berenbrink:2009:NAM

[BFH09] Petra Berenbrink, Tom Friedetzky, and Zengjian Hu. A new analytical method for parallel, diffusion-type load balancing. Journal of Parallel and Distributed Computing, 69(1):54–61, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brooks:2004:TMT

[BFKP04] Richard Brooks, David Friedlander, John Koch, and Shashi Phoha. Tracking multiple targets with self-organizing distributed ground sensors. Journal of Parallel and Distributed Computing, 64(7):874–884, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bilo:2006:PAB

[BFM06] Vittorio Bilò, Michele Flammini, and Luca Moscardelli. Pareto approximations for the bicriteria scheduling problem. Journal of Parallel and Distributed Computing, 66(3):393–402, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baiardi:1987:DIN

[BFTV87] F. Baiardi, A. Fantechi, A. Tomasi, and M. Vanneschi. Distributed implementation of nested communicating sequential processes: communication and termination. Journal of Parallel and Distributed Computing, 4(6):531–545, December 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barahona:1986:PAM

[BG86] P. M. C. C. Barahona and J. R. Gurd. Processor allocation in a multi-ring dataflow machine. Journal of Parallel and Distributed Computing, 3(3):305–327, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barbosa:1989:DIS

[BG89] Valmir C. Barbosa and Eli Gafni. A distributed implementation of simulated annealing. Journal of Parallel and Distributed Computing, 6(2):411–434, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baru:1990:SER

[BG90a] Chaitanya K. Baru and Piyush Goel. Squashed embedding of E-R schemas in hypercubes. Journal of Parallel and Distributed Computing, 8(4):340–348, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bic:1990:SID

[BG90b] Lubomir Bic and Jean-Luc Gaudiot. Special issue on dataflow computing: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 10(4):277–278, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boichat:2005:RTO

[BG05] Romain Boichat and Rachid Guerraoui. Reliable and total order broadcast in the crash-recovery model. Journal of Parallel and Distributed Computing, 65(4):397–413, April 2005. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barberou:2003:EME

[BGH^{+}03] Nicolas Barberou, Marc Garbey, Matthias Hess, Michael M. Resch, Tuomo Rossi, Jari Toivanen, and Damien Tromeur-Dervout. Efficient metacomputing of elliptic linear and nonlinear problems. Journal of Parallel and Distributed Computing, 63(5):564–577, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Beauquier:2002:TBS

[BGJDL02] Joffroy Beauquier, Maria Gradinariu, Colette Johnen, and Jérôme Durand-Lose. Token-based self-stabilizing uniform algorithms. Journal of Parallel and Distributed Computing, 62(5):899–921, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2001.1832;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2001.
1832/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2001.1832/ref.

Bao:2003:DDC

[BGLA03] Lichun Bao and J. J. Garcia-Luna-Aceves. Distributed dynamic channel access scheduling for ad hoc networks. Journal of Parallel and Distributed Computing, 63(1):3–14, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bozdag:2008:FSG

[BGM^{+}08] Doruk Bozdaḡ, Assefaw H. Gebremedhin, Fredrik Manne, Erik G. Boman, and Umit V. Catalyurek. A framework for scalable greedy coloring on distributed-memory parallel computers. Journal of Parallel and Distributed Computing, 68(4):515–535, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bokka:1995:CTC

[BGOS95] V. Bokka, H. Gurla, S. Olariu, and J. L. Schwing. Constanttime convexity problems on reconfigurable meshes. Journal of Parallel and Distributed Computing, 27(1):86–99, May 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1074/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1074/production/
pdf.

Banerjee:1996:PAS

[BGR96] S. Banerjee, R. K. Ghosh, and A. P. K. Reddy. Parallel algorithm for shortest pairs of edge-disjoint paths. Journal of Parallel and Distributed Computing, 33(2):165–171, March 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0035/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0035/production/
pdf.

Bic:1986:SPD

[BH86] Lubomir Bic and Robert L. Hartmann. Simulated performance of a data-driven database machine. Journal of Parallel and Distributed Computing, 3(1):1–22, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bohm:1993:DTS

[BH93] A. P. Willem Bohm and Robert E. Hiromoto. The dataflow time and space complexity of FFTs. Journal of Parallel and Distributed Computing, 18(3):301–313, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1066/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1066/production/pdf.

Boyer:2005:NEA

[BH05] Wayne F. Boyer and Gurdeep S. Hura. Non-evolutionary algorithm for scheduling dependent tasks in distributed heterogeneous computing environments. Journal of Parallel and Distributed Computing, 65(9):1035–1046, September 2005. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boguslavsky:1994:OSS

[BHK^{+}94] L. Boguslavsky, K. Harzallah, A. Kreinen, K. Sevcik, and A. Vainshtein. Optimal strategies for spinning and blocking. Journal of Parallel and Distributed Computing, 21(2):246– 254, May 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1056/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1056/
production/pdf.

Bilardi:2005:SL

[BHPP05] Gianfranco Bilardi, Kieran Herley, Andrea Pietracaprina, and Geppino Pucci. On stalling in LogP. Journal of Parallel and Distributed Computing, 65(3):307–312, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bernhard:1991:CMP

[BHR91] P. J. Bernhard, H. B. Hunt, III, and D. J. Rosenkrantz. Compaction of message patterns into succinct representations for multiprocessor interconnection networks. Journal of Parallel and Distributed Computing, 12(1):39–49, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brzezinski:1995:DMG

[BHRS95] Jerzy Brzezinski, Jean-Michel M. Helary, Michel Raynal, and Mukesh Singhal. Deadlock models and a general algorithm for distributed deadlock detection. Journal of Parallel and Distributed Computing, 31(2):112–125, December 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1150/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1150/production/pdf.

See erratum [Ano96l].

Blelloch:1994:IPN

[BHS^{+}94] Guy E. Blelloch, Jonathan C. Hardwick, Jay Sipelstein, Marco Zagha, and Siddhartha Chatterjee. Implementation of a portable nested data-parallel language. Journal of Parallel and Distributed Computing, 21(1):4–14, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1037/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1037/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1038/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1038/production/pdf.

Bhuyan:1987:AIN

[Bhu87] Laxmi N. Bhuyan. Analysis of interconnection networks with different arbiter designs. Journal of Parallel and Distributed Computing, 4(4):384–403, August 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bic:1990:POM

[Bic90] Lubomir Bic. A process-oriented model for efficient execution of dataflow programs. Journal of Parallel and Distributed Computing, 8(1):42–51, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Birman:1994:IRC

[Bir94] K. P. Birman. Integrating runtime consistency models for distributed computing. Journal of Parallel and Distributed Computing, 23(2):158–176, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1129/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1129/production/pdf.

Bitar:1992:WMA

[Bit92] Philip Bitar. The weakest memory-access order. Journal of Parallel and Distributed Computing, 15(4):305–331, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bader:1996:PAI

[BJ96] David A. Bader and Joseph JáJá. Parallel algorithms for image histogramming and connected components with an experimental study. Journal of Parallel and Distributed Computing, 35(2):173–190, June 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0079/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0079/production/pdf.

Bader:1999:SMP

[BJ99] David A. Bader and Joseph JáJá. Simple: a methodology for programming high performance algorithms on clusters of symmetric multiprocessors (SMPs). Journal of Parallel and Distributed Computing, 58(1):92–108, July 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1541/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1541/production/pdf.

Blumofe:1996:CEM

[BJK^{+}96] Robert D. Blumofe, Christopher F. Joerg, Bradley C. Kuszmaul, Charles E. Leiserson, Keith H. Randall, and Yuli Zhou. Cilk: An efficient multithreaded runtime system. Journal of Parallel and Distributed Computing, 37(1):55–69, August 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0107/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0107/production/
pdf.

Beck:1991:CFD

[BJP91] Micah Beck, Richard Johnson, and Keshav Pingali. From control flow to dataflow. Journal of Parallel and Distributed Computing, 12(2):118–129, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baldoni:2008:DQD

[BJPPM^{+}08] Roberto Baldoni, Ricardo Jiménez-Peris, Marta Patiño-Martínez, Leonardo Querzoni, and Antonino Virgillito. Dynamic quorums for DHT-based enterprise infrastructures. Journal of Parallel and Distributed Computing, 68(9):1235–1249, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bilas:2003:SVM

[BJS03] Angelos Bilas, Dongming Jiang, and Jaswinder Pal Singh. Shared virtual memory clusters: bridging the cost-performance gap between SMPs and hardware DSM systems. Journal of Parallel and Distributed Computing, 63(12):1257–1276, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Baden:1995:PPP

[BK95] Scott B. Baden and Scott R. Kohn. Portable parallel programming of numerical problems under the LPAR system. Journal of Parallel and Distributed Computing, 27(1):38–55, May 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1070/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1070/production/
pdf.

Biyani:2008:ADA

[BK08] Karun N. Biyani and Sandeep S. Kulkarni. Assurance of dynamic adaptation in distributed systems. Journal of Parallel and Distributed Computing, 68(8):1097–1112, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Berg:1991:LIM

[BKS91] Thomas B. Berg, Shin-Dug Kim, and Howard Jay Siegel. Limitations imposed on mixed-mode performance of optimized phases due to temporal juxtaposition. Journal of Parallel and Distributed Computing, 13(2):154–169, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bansal:2005:DHT

[BKS05] Savina Bansal, Padam Kumar, and Kuldip Singh. Dealing with heterogeneity through limited duplication for scheduling precedence constrained task graphs. Journal of Parallel and Distributed Computing, 65(4):479–491, April 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhattacharya:1995:CCS

[BKT95] Sourav Bhattacharya, Thomas F. Keefe, and Wei-Tek T. Tsai. Covert channel secure hypercube message communication. Journal of Parallel and Distributed Computing, 26(2):233–247, April 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1062/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1062/production/pdf.

Bier:1989:EBT

[BL89] Thomas Bier and Kia-Fock Loe. Embedding of binary trees into hypercubes. Journal of Parallel and Distributed Computing, 6 (3):679–691, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bojanczyk:1990:USA

[BL90] Adam W. Bojanczyk and Franklin T. Luk. A unified systolic array for adaptive beamforming. Journal of Parallel and Distributed Computing, 8(4):388–392, April 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertossi:1994:PSM

[BL94] A. A. Bertossi and F. Logi. Parallel string matching with variable length don’t cares. Journal of Parallel and Distributed Computing, 22(2):229–234, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1083/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1083/production/pdf.

Bianchini:1996:EPM

[BL96] Ricardo Bianchini and Beng-Hong Lim. Evaluating the performance of multithreading and prefetching in multiprocessors. Journal of Parallel and Distributed Computing, 37(1):83–97, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0109/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0109/production/pdf.

Bachmat:2005:ECC

[BL05] Eitan Bachmat and Tao Kai Lam. On the effect of a configuration choice on the performance of a mirrored storage system. Journal of Parallel and Distributed Computing, 65(3):382–395, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Broberg:2001:POU

[BLG01] Magnus Broberg, Lars Lundberg, and Håkan Grahn. Performance optimization using extended critical path analysis in multithreaded programs on multiprocessors. Journal of Parallel and Distributed Computing, 61(1):115–136, January 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1667;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1667/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1667/ref.

Bakiras:2005:ASD

[BLPA05] Spiridon Bakiras, Thanasis Loukopoulos, Dimitris Papadias, and Ishfaq Ahmad. Adaptive schemes for distributed Web caching. Journal of Parallel and Distributed Computing, 65(12): 1483–1496, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Barnett:1995:GCA

[BLPV95] M. Barnett, R. Littlefield, D. G. Payne, and R. Van de Geijn. Global combine algorithms for 2-D meshes with wormhole routing. Journal of Parallel and Distributed Computing, 24(2):191–201, February 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1018/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1018/production/pdf.

Buker:1995:PEH

[BM95] Ulrich Buker and Barbel Mertsching. Parallel evaluation of hierarchical image databases. Journal of Parallel and Distributed Computing, 31(2):141–152, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1152/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1152/production/pdf.

Bertossi:1997:BPQ

[BM97] Alan A. Bertossi and Alessandro Mei. P -bandwidth priority queues on reconfigurable tree of meshes. Journal of Parallel and Distributed Computing, 40(2):248–255, February 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1259/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1259/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1259/production/ref.

Bertossi:2004:TWO

[BM04a] Alan A. Bertossi and Alessandro Mei. Time and work optimal simulation of basic reconfigurable meshes on hypercubes. Journal of Parallel and Distributed Computing, 64(1):173–180, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boxer:2004:CGG

[BM04b] Laurence Boxer and Russ Miller. Coarse grained gather and scatter operations with applications. Journal of Parallel and Distributed Computing, 64(11):1297–1310, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bronevich:2008:LBA

[BM08] Andrey G. Bronevich and Wolfgang Meyer. Load balancing algorithms based on gradient methods and their analysis through algebraic graph theory. Journal of Parallel and Distributed Computing, 68(2):209–220, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boukerche:2007:PSL

[BMARM07] Azzedine Boukerche, Alba Cristina Magalhaes Alves de Melo, Mauricio Ayala-Rincón, and Maria Emilia Machado Telles Walter. Parallel strategies for the local biological sequence alignment in a cluster of workstations. Journal of Parallel and Distributed Computing, 67(2):170–185, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bai:2008:MMR

[BMB^{+}08] Xin Bai, Dan C. Marinescu, Ladislau Bölöni, Howard Jay Siegel, Rose A. Daley, and I-Jeng Wang. A macroeconomic model for resource allocation in large-scale distributed systems. Journal of Parallel and Distributed Computing, 68(2):182–199, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brungger:1998:SLS

[BMCP98] Adrian Brüngger, Ambros Marzetta, Jens Clausen, and Michael Perregaard. Solving large-scale QAP problems in parallel with the search library ZRAM. Journal of Parallel and Distributed Computing, 50(1–2):157–169, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1434/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1434/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1434/production/ref.

Boukerche:2005:RCL

[BMF05] Azzedine Boukerche, Armin Mikkler, and Alessandro Fabri. Resource control for large-scale distributed simulation system over loosely coupled domains. Journal of Parallel and Distributed Computing, 65(10):1171–1189, October 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Busch:2007:EBP

[BMIM07] Costas Busch, Malik Magdon-Ismail, and Marios Mavronicolas. Efficient bufferless packet switching on trees and leveled networks. Journal of Parallel and Distributed Computing, 67(11):

1168–1186, November 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhuvaneswari:1997:NFG

[BMM97] K. Bhuvaneswari, K. N. Balasubramanya Murthy, and C. Siva Ram Murthy. A new and faster Gaussian elimination based fault tolerant systolic linear system solver. Journal of Parallel and Distributed Computing, 44(2):107–122, August 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1352/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1352/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1352/production/ref.

Bermond:2001:BPL

[BMMS01] Jean-Claude Bermond, Xavier Muñoz, and Alberto Marchetti-Spaccamela. A broadcasting protocol in line digraphs. Journal of Parallel and Distributed Computing, 61(8):1013–1032, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1737;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1737/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1737/ref.

Boxer:1998:SPA

[BMRC98] Laurence Boxer, Russ Miller, and Andrew Rau-Chaplin. Scaleable parallel algorithms for lower envelopes with applications. Journal of Parallel and Distributed Computing, 53(2):91–118, September 15, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1478/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1478/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1478/production/ref.

Boxer:1999:SPA

[BMRC99] Laurence Boxer, Russ Miller, and Andrew Rau-Chaplin. Scalable parallel algorithms for geometric pattern recognition. Journal of Parallel and Distributed Computing, 58(3):466–486, September 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1565/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1565/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1565/production/ref.

Beguelin:1994:VPP

[BN94] A. Beguelin and G. Nutt. Visual parallel programming and determinacy: a language specification, an analysis technique, and a programming tool. Journal of Parallel and Distributed Computing, 22(2):235–250, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1084/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1084/production/pdf.

Boukerche:2002:BBI

[BN02] Azzedine Boukerche and Mirela Sechi M. Annoni Notare. Behavior-based intrusion detection in mobile phone systems. Journal of Parallel and Distributed Computing, 62(9):1476–1490, September 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1857;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1857/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1857/ref.

Boukerche:2003:RMW

[BN03] Azzedine Boukerche and Mirela Sechi M. Annoni Notare. Routing in mobile and wireless ad hoc networks. Journal of Parallel and Distributed Computing, 63(2):107–109, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Becker:1998:NCC

[BNP98] Ronald I. Becker, David Nassimi, and Yehoshua Perl. The new class of g-chain periodic sorters. Journal of Parallel and Distributed Computing, 54(2):206–222, November 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1485/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1485/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1485/production/ref.

Barillari:2002:FDD

[BNP02] Fabio Barillari, Enrico Nardelli, and Massimo Pepe. Fully dynamic distributed search trees can be balanced in O(lg 2N) time. Journal of Parallel and Distributed Computing, 62(11): 1617–1628, November 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bar-Noy:2000:OBT

[BNS00] Amotz Bar-Noy and Yaron Shilo. Optimal broadcasting of two files over an asymmetric channel. Journal of Parallel and Distributed Computing, 60(4):474–493, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1618;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1618/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1618/ref.

Buttner:1999:APH

[BNSP99] Lars Büttner, Jörg Nolte, and Wolfgang Schröder-Preikschat. ARTS of PEACE — a high-performance middleware layer for parallel distributed computing. Journal of Parallel and Distributed Computing, 59(2):155–179, November 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1570/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1570/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1570/production/ref.

Bodlaender:1989:CCP

[Bod89] H. L. Bodlaender. The classification of coverings of processor networks. Journal of Parallel and Distributed Computing, 6(1): 166–182, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Balsara:1991:DSM

[BOI91] Poras T. Balsara, Robert M. Owens, and Mary Jane Irwin. Digit serial multipliers. Journal of Parallel and Distributed Computing, 11(2):156–162, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertossi:2006:SIA

[BOP06] Alan A. Bertossi, Stephan Olariu, and Cristina M. Pinotti. Special issue: Algorithms for wireless and ad-hoc networks. Journal of Parallel and Distributed Computing, 66(4):487–488, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertsekas:1991:OCA

[BOS^{+}91] D. P. Bertsekas, C. Ozveren, G. D. Stamoulis, P. Tseng, and J. N. Tsitsiklis. Optimal communication algorithms for hypercubes. Journal of Parallel and Distributed Computing, 11 (4):263–275, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhagavathi:1995:CPM

[BOS^{+}95] D. Bhagavathi, S. Olariu, J. L. Schwing, W. Shen, L. Wilson, and J. Zhang. Convexity problems on meshes with multiple broadcasting. Journal of Parallel and Distributed Computing, 27(2):142–156, June 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1078/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1078/production/pdf.

Bhagavathi:1994:TOM

[BOSW94] D. Bhagavathi, S. Olariu, W. Shen, and L. Wilson. A time-optimal multiple search algorithm on enhanced meshes, with applications. Journal of Parallel and Distributed Computing, 22(1):113–120, July 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1074/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1074/production/pdf.

Boukerche:2002:APA

[Bou02] Azzedine Boukerche. An adaptive partitioning algorithm for distributed discrete event simulation systems. Journal of Parallel and Distributed Computing, 62(9):1454–1475, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1856;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1856/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1856/ref.

Boukerche:2003:WMA

[Bou03] Azzedine Boukerche. Wireless and mobile ad hoc networking and computing. Journal of Parallel and Distributed Computing, 63(1):1–2, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bozejko:2009:SFS

[Boż09] Wojciech Bożejko. Solving the flow shop problem by parallel programming. Journal of Parallel and Distributed Computing, 69(5):470–481, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bongiovanni:1989:PDS

[BP89] G. Bongiovanni and R. Petreschi. Parallel-depth search for acyclic digraphs. Journal of Parallel and Distributed Computing, 7(2):383–390, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bilardi:1995:HPC

[BP95] Gianfranco Bilardi and Franco P. Preparata. Horizons of parallel computation. Journal of Parallel and Distributed Computing, 27(2):172–182, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1080/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1080/production/pdf.

Barth:1998:RPG

[BP98] Dominique Barth and Petrişor Panaite. Routing permutations on graphs via factors. Journal of Parallel and Distributed Computing, 54(2):162–182, November 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1488/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1488/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1488/production/ref.

Brightwell:2001:SPT

[BP01] Ron Brightwell and Steve Plimpton. Scalability and performance of two large Linux clusters. Journal of Parallel and Distributed Computing, 61(11):1546–1569, November 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1756;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1756/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1756/ref.

Bertossi:2002:MCF

[BP02] Alan A. Bertossi and Cristina M. Pinotti. Mappings for conflict-free access of paths in bidimensional arrays, circular lists, and complete trees. Journal of Parallel and Distributed Computing, 62(8):1314–1333, August 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1860;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1860/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1860/ref.

Boukerche:2006:FTW

[BPA06] Azzedine Boukerche, Richard Werner Nelem Pazzi, and Regina Borges Araujo. Fault-tolerant wireless sensor network routing protocols for the supervision of context-aware physical environments. Journal of Parallel and Distributed Computing, 66(4):586–599, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Balaji:1992:NPN

[BPJG92] S. Balaji, L. M. Patnaik, Lawrence Jenkins, and P. S. Goel. S-nets: A Petri net based model for performance evaluation of real-time scheduling algorithms. Journal of Parallel and Distributed Computing, 15(3):225–237, July 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Beck:1990:SSD

[BPN90] M. Beck, K. K. Pingali, and A. Nicolau. Static scheduling for dynamic dataflow machines. Journal of Parallel and Distributed Computing, 10(4):279–288, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bourgeois:2005:CTF

[BPP05] Anu G. Bourgeois, Yi Pan, and Sushil K. Prasad. Constant time fault tolerant algorithms for a linear array with a reconfigurable pipelined bus system. Journal of Parallel and Distributed Computing, 65(3):374–381, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bhat:1999:ACA

[BPR99] Prashanth B. Bhat, Viktor K. Prasanna, and C. S. Raghavendra. Adaptive communication algorithms for distributed heterogeneous systems. Journal of Parallel and Distributed Computing, 59(2):252–279, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1571/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1571/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1571/production/ref.

Bertossi:2004:ASI

[BPRG04] A. A. Bertossi, M. C. Pinotti, R. Rizzi, and P. Gupta. Allocating servers in infostations for bounded simultaneous requests. Journal of Parallel and Distributed Computing, 64(10): 1113–1126, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bertossi:2004:CAI

[BPRS04] Alan A. Bertossi, Cristina M. Pinotti, Romeo Rizzi, and Anil M. Shende. Channel assignment for interference avoidance in honeycomb wireless networks. Journal of Parallel and Distributed Computing, 64(12):1329–1344, December 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bermond:1996:PGE

[BPST96] J.-C. Bermond, C. Peyrat, I. Sakho, and M. Tchuente. Parallelization of the Gaussian elimination algorithm on systolic arrays. Journal of Parallel and Distributed Computing, 33(1):69–75, February 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0025/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0025/production/pdf.

Barnett:1996:BMW

[BPvW96] Mike Barnett, David G. Payne, Robert A. van de Geijn, and Jerrell Watts. Broadcasting on meshes with wormhole routing. Journal of Parallel and Distributed Computing, 35(2):111–122, June 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0074/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0074/production/pdf.

Brodsky:2005:CBA

[BPW05] Alex Brodsky, Jan Bækgaard Pedersen, and Alan Wagner. On the complexity of buffer allocation in message passing systems. Journal of Parallel and Distributed Computing, 65(6):692–713, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bernhard:1991:EMR

[BR91a] P. J. Bernhard and D. J. Rosenkrantz. An efficient method for representing and transmitting message patterns on multiprocessor interconnection networks. Journal of Parallel and Distributed Computing, 11(1):72–85, January 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boppana:1991:GSA

[BR91b] Rajendra V. Boppana and C. S. Raghavendra. Generalized schemes for access and alignment of data in parallel processors with self-routing interconnection networks. Journal of Parallel and Distributed Computing, 11(2):97–111, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Babaoglu:1995:SVD

[BR95a] Ozalp Babaoglu and Michel Raynal. Specification and verification of dynamic properties in distributed computations. Journal of Parallel and Distributed Computing, 28(2):173–185, August 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1098/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1098/production/
pdf.

Bernaschi:1995:DRP

[BR95b] Massimo Bernaschi and Giorgio Richelli. Development and results of PVMe on the IBM 9076 SP1. Journal of Parallel and Distributed Computing, 29(1):75–83, August 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1107/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1107/production/pdf.

Boppana:1995:MAA

[BR95c] Rajendra V. Boppana and C. S. Raghavendra. On methods to align and access data arrays in parallel computers. Journal of Parallel and Distributed Computing, 26(2):261–269, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1065/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1065/production/
pdf.

Bae:1996:PUC

[BR96] Seungjo Bae and Sanjay Ranka. PACK/UNPACK on coarsegrained distributed memory parallel machines. Journal of Parallel and Distributed Computing, 38(2):204–216, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0141/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0141/production/
pdf.

Boukerche:2002:DGB

[BR02] Azzedine Boukerche and Amber Roy. Dynamic grid-based approach to data distribution management. Journal of Parallel and Distributed Computing, 62(3):366–392, March 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1799;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1799/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1799/ref.

Benoit:2008:MPS

[BR08] Anne Benoit and Yves Robert. Mapping pipeline skeletons onto heterogeneous platforms. Journal of Parallel and Distributed Computing, 68(6):790–808, June 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bromley:1996:QNG

[Bro96] B. C. Bromley. Quasirandom number generators for parallel Monte Carlo algorithms. Journal of Parallel and Distributed Computing, 38(1):101–104, October 10, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0132/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0132/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.0133/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0133/production/pdf.

Bhat:2003:ECC

[BRP03] Prashanth B. Bhat, C. S. Raghavendra, and Viktor K. Prasanna. Efficient collective communication in distributed heterogeneous systems. Journal of Parallel and Distributed Computing, 63(3):251–263, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bosque:2006:PCI

[BRPR06] José L. Bosque, Oscar D. Robles, Luis Pastor, and Angel Rodríguez. Parallel CBIR implementations with load balancing algorithms. Journal of Parallel and Distributed Computing, 66(8):1062–1075, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Boudet:2001:ADA

[BRR01] Vincent Boudet, Fabrice Rastello, and Yves Robert. Alignment and distribution is not (always) NP-Hard. Journal of Parallel and Distributed Computing, 61(4):501–519, April 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1683;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1683/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1683/ref.

Berman:1987:MPA

[BS87] Francine Berman and Lawrence Snyder. On mapping parallel algorithms into parallel architectures. Journal of Parallel and Distributed Computing, 4(5):439–458, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Blelloch:1990:CCO

[BS90] G. E. Blelloch and G. W. Sabot. Compiling collection-oriented languages onto massively parallel computers. Journal of Parallel and Distributed Computing, 8(2):119–134, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bolosky:1992:EMM

[BS92] William J. Bolosky and Michael L. Scott. Evaluation of multiprocessor memory systems using off-line optimal behavior. Journal of Parallel and Distributed Computing, 15(4):382–398, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bellosa:1996:PIL

[BS96a] Frank Bellosa and Martin Steckermeier. The performance implications of locality information usage in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 37(1):113–121, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0112/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0112/production/pdf.

Buck:1996:ASC

[BS96b] Duane Buck and Mukesh Singhal. An analytic study of caching in computer systems. Journal of Parallel and Distributed Computing, 32(2):205–214, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0014/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0014/production/pdf.

See erratum [BS96c].

Buck:1996:EVN

[BS96c] Duane Buck and Mukesh Singhal. Erratum: Volume 32, number 2 (1996), pp. 205–214: Duane Buck and Mukesh Singhal, “An Analytic Study of Caching in Computer Systems”. Journal of Parallel and Distributed Computing, 34(2):233, May 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0060/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0060/production/

pdf. See [BS96b].

Banerjee:1997:PAF

[BS97] Sudarshan Banerjee and Sanjeev Saxena. Parallel algorithm for finding the most vital edge in weighted graphs. Journal of Parallel and Distributed Computing, 46(1):101–104, October 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1383/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1383/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1383/production/ref.

Brockington:2000:AAP

[BS00] Mark G. Brockington and Jonathan Schaeffer. APHID: Asynchronous parallel game-tree search. Journal of Parallel and Distributed Computing, 60(2):247–273, February 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1600/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1600/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1600/production/ref.

Bass:2003:PPR

[BS03] Douglas W. Bass and I. Hal Sudborough. Pancake problems with restricted prefix reversals and some corresponding Cayley networks. Journal of Parallel and Distributed Computing, 63 (3):327–336, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Byna:2009:SIJ

[BS09] Surendra Byna and Xian-He Sun. Special issue of the Journal of Parallel and Distributed Computing: Data-intensive computing. Journal of Parallel and Distributed Computing, 69(11):937, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Braun:2001:CES

[BSB^{+}01] Tracy D. Braun, Howard Jay Siegel, Noah Beck, Ladislau L. Bölöni, Muthucumaru Maheswaran, Albert I. Reuther, James P. Robertson, Mitchell D. Theys, Bin Yao, Debra Hensgen, and Richard F. Freund. A comparison of eleven static heuristics for mapping a class of independent tasks onto heterogeneous distributed computing systems. Journal of Parallel and Distributed Computing, 61(6):810–837, June 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1714;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1714/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1714/ref.

Beresford-Smith:1996:OAC

[BSDE96] Bryan Beresford-Smith, Oliver Diessel, and Hossam El-Gindy. Optimal algorithms for constrained reconfigurable meshes. Journal of Parallel and Distributed Computing, 39(1):74–78, November 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0155/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0155/production/pdf.

Bertolazzi:1990:PAV

[BSG90] Paola Bertolazzi, Silvio Salza, and Concettina Guerra. A parallel algorithm for the visibility problem from a point. Journal of Parallel and Distributed Computing, 9(1):11–14, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Berryman:1990:KMP

[BSGM90] H. Berryman, J. Saltz, W. Gropp, and R. Mirchandaney. Krylov methods preconditioned with incompletely factored matrices on the CM-2. Journal of Parallel and Distributed Computing, 8(2):186–190, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Braun:2008:SRA

[BSMH08] Tracy D. Braun, Howard Jay Siegel, Anthony A. Maciejewski, and Ye Hong. Static resource allocation for heterogeneous computing environments with tasks having dependencies, priorities, deadlines, and multiple versions. Journal of Parallel and Distributed Computing, 68(11):1504–1516, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Beguelin:1997:ALF

[BSS97] Adam Beguelin, Erik Seligman, and Peter Stephan. Application level fault tolerance in heterogeneous networks of workstations. Journal of Parallel and Distributed Computing, 43(2):147–155, June 15, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1338/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1338/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1338/production/ref.

Blathras:1999:TML

[BSS99] Kostas Blathras, Daniel B. Szyld, and Yuan Shi. Timing models and local stopping criteria for asynchronous iterative algorithms. Journal of Parallel and Distributed Computing, 58(3):446–465, September 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1549/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1549/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1549/production/ref.

Burge:2001:UNN

[BST01] Peter Burge and John Shawe-Taylor. An unsupervised neural network approach to profiling the behavior of mobile phone users for use in fraud detection. Journal of Parallel and Distributed Computing, 61(7):915–925, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1720;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1720/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1720/ref.

Bai:2007:SHT

[BSW07] Rendong Bai, Mukesh Singhal, and Yongwei Wang. On supporting high-throughput routing metrics in on-demand routing protocols for multi-hop wireless networks. Journal of Parallel and Distributed Computing, 67(10):1108–1118, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brun:2002:PPF

[BTG02] Olivier Brun, Vincent Teuliere, and Jean-Marie Garcia. Parallel particle filtering. Journal of Parallel and Distributed Computing, 62(7):1186–1202, July 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1843;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1843/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1843/ref.

Brodal:1998:PPQ

[BTZ98] Gerth Stolting Brodal, Jesper Larsson Traff, and Christos D. Zaroliagis. A parallel priority queue with constant time operations. Journal of Parallel and Distributed Computing, 49(1):4–21, February 25, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1425/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1425/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1425/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1435/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1435/
production/pdf.

Buchholz:1992:HVG

[Buc92] Peter Buchholz. A hierarchical view of GCSPNs and its impact on qualitative and quantitative analysis. Journal of Parallel and Distributed Computing, 15(3):207–224, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bae:2002:DRB

[BVB02] Myung M. Bae, R. Venkatesan, and Bella Bose. Data rearrangement between radix-k and Lee distance gray codes in k-ary n-cubes. Journal of Parallel and Distributed Computing, 62(1):19–37, January 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1772;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1772/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1772/ref.

Baer:1989:MCH

[BW89] Jean-Loup Baer and Wen-Hann Wang. Multilevel cache hierarchies: organizations, protocols, and performance. Journal of Parallel and Distributed Computing, 6(3):451–476, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bik:1995:ACO

[BW95a] Aart J. C. Bik and Harry A. G. Wijshoff. Advanced compiler optimizations for sparse computations. Journal of Parallel and Distributed Computing, 31(1):14–24, November 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1141/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1141/production/pdf.

Blough:1995:CDR

[BW95b] Douglas M. Blough and Hongying Y. Wang. Cooperative diagnosis and routing in fault-tolerant multiprocessor systems. Journal of Parallel and Distributed Computing, 27(2):205– 211, June 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1083/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1083/
production/pdf.

Bik:1996:UIS

[BW96] Aart J. C. Bik and Harry A. G. Wijshoff. The use of iteration space partitioning to construct representative simple sections. Journal of Parallel and Distributed Computing, 34(1):95–110, April 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0048/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0048/production/pdf.

Bai:2008:PBT

[BW08] Yihua Bai and Robert C. Ward. Parallel block tridiagonalization of real symmetric matrices. Journal of Parallel and Distributed Computing, 68(5):703–715, May 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bharadwaj:2009:HBS

[BW09] Veeravalli Bharadwaj and Han Min Wong. Handling biological sequence alignments on networked computing systems: a divide-and-conquer approach. Journal of Parallel and Distributed Computing, 69(10):854–865, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brown:1993:PQA

[BX93]  Theodore Brown and Renbing B. Xiong.  A parallel

quicksort algorithm. Journal of Parallel and Distributed Computing, 19(2):83–89, October 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1093/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1093/production/pdf.

Banerjee:2008:FTM

[BXA08] Torsha Banerjee, Bin Xie, and Dharma P. Agrawal. Fault tolerant multiple event detection in a wireless sensor network. Journal of Parallel and Distributed Computing, 68(9):1222–1234, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Butt:2006:SOF

[BZH06] Ali R. Butt, Rongmei Zhang, and Y. Charlie Hu. A selforganizing flock of Condors. Journal of Parallel and Distributed Computing, 66(1):145–161, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Brooks:2004:AOD

[BZLI04] R. R. Brooks, Mengxia Zhu, Jacob Lamb, and S. S. Iyengar. Aspect-oriented design of sensor networks. Journal of Parallel and Distributed Computing, 64(7):853–865, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Callsen:1994:OHC

[CA94] C. J. Callsen and G. Agha. Open heterogeneous computing in ActorSpace. Journal of Parallel and Distributed Computing, 21(3):289–300, June 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1060/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1060/production/pdf.

Cheng:1995:ORS

[CA95a] Sheng-Tzong T. Cheng and Ashok K. Agrawala. Optimal replication of series-parallel graphs for computationintensive applications. Journal of Parallel and Distributed Computing, 28(2):113–129, August 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1094/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1094/production/pdf.

Conrad:1995:APA

[CA95b] J. M. Conrad and D. P. Agrawal. Asynchronous parallel arc consistency algorithms on a distributed memory machine. Journal of Parallel and Distributed Computing, 24(1):27–40, January 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1004/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1004/
production/pdf.

Coelho:1996:OCH

[CA96] Fabien Coelho and Corinne Ancourt. Optimal compilation of HPF remappings. Journal of Parallel and Distributed Computing, 38(2):229–236, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0143/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0143/production/pdf.

Chen:1994:DAC

[CAB94] Chienhua H. Chen, D. P. Agrawal, and J. R. Burke. Design and analysis of a class of highly scalable hierarchical networks: PdBCube. Journal of Parallel and Distributed Computing, 22(3):555–564, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1111/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1111/production/pdf.

Calinescu:2006:FLA

[Cal06] Gruia Calinescu. A fast localized algorithm for scheduling sensors. Journal of Parallel and Distributed Computing, 66(4):507– 514, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cappello:1987:GEH

[Cap87] Peter R. Cappello. Gaussian elimination on a hypercube automaton. Journal of Parallel and Distributed Computing, 4 (3):288–308, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carlson:1990:SLR

[Car90] David A. Carlson. Solving linear recurrence systems on meshconnected computers with multiple global buses. Journal of Parallel and Distributed Computing, 8(1):89–95, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carter:1995:DMD

[Car95] John B. Carter. Design of the Munin distributed shared memory system. Journal of Parallel and Distributed Computing, 29(2):219–227, September 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1119/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1119/production/pdf.

Casavant:1993:TMV

[Cas93] Thomas L. Casavant. Tools and methods for visualization of parallel systems and computations: Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 18(2):103– 104, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choi:1995:RLS

[CB95] Jaeyoung Y. Choi and A. W. Bojanczyk. Recursive leastsquares problems on distributed-memory multiprocessors.

Journal of Parallel and Distributed Computing, 24(1):11–26, January 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1003/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1003/
production/pdf.

Clarke:1996:LSP

[CB96] Gwen Clarke and David T. Barnard. An LR substring parser applied in a parallel environment. Journal of Parallel and Distributed Computing, 35(1):2–17, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0063/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0063/production/pdf.

Chandy:1999:PCP

[CB99] John A. Chandy and Prithviraj Banerjee. A parallel circuitpartitioned algorithm for timing-driven standard cell placement. Journal of Parallel and Distributed Computing, 57(1): 64–90, April 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1523/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1523/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1523/production/ref.

Cirne:2002:UMI

[CB02] Walfredo Cirne and Francine Berman. Using moldability to improve the performance of supercomputer jobs. Journal of Parallel and Distributed Computing, 62(10):1571–1601, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1869;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1869/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1869/ref.

Cong:2006:DIP

[CB06] Guojing Cong and David A. Bader. Designing irregular parallel algorithms with mutual exclusion and lock-free protocols. Journal of Parallel and Distributed Computing, 66(6):854–866, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Catalyurek:2009:RHM

[CBD^{+}09] Umit V. Catalyurek, Erik G. Boman, Karen D. Devine, Doruk Bozdaḡ, Robert T. Heaphy, and Lee Ann Riesen. A repartitioning hypergraph model for dynamic load balancing. Journal of Parallel and Distributed Computing, 69(8):711–724, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Costa:2000:PLP

[CBdCD00] Vítor Santos Costa, Ricardo Bianchini, and Inês de Castro Dutra. Parallel logic programming systems on scalable architectures. Journal of Parallel and Distributed Computing, 60(7):835–852, July 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1629;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1629/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1629/ref.

Che:2008:PSG

[CBM^{+}08] Shuai Che, Michael Boyer, Jiayuan Meng, David Tarjan, Jeremy W. Sheaffer, and Kevin Skadron. A performance study of general-purpose applications on graphics processors using CUDA. Journal of Parallel and Distributed Computing, 68(10): 1370–1380, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carothers:2002:RHP

[CBP02] Christopher D. Carothers, David Bauer, and Shawn Pearce. ROSS: a high-performance, low-memory, modular Time Warp system. Journal of Parallel and Distributed Computing, 62(11): 1648–1669, November 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Correia:2008:BGA

[CBV08] Miguel Correia, Alysson Neves Bessani, and Paulo Veríssimo. On Byzantine generals with alternative plans. Journal of Parallel and Distributed Computing, 68(9):1291–1296, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1987:ILP

[CC87] Zen Chen and Chih-Chi Chang. Iteration-level parallel execution of DO loops with a reduced set of dependence relations. Journal of Parallel and Distributed Computing, 4(5):488–504, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Castagnetti:1991:SAP

[CC91] T. Castagnetti and P. Ciancarini. Static analysis of a parallel logic language based on the blackboard model. Journal of Parallel and Distributed Computing, 13(4):412–423, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiang:1994:DFT

[CC94] Wei-Kuo K. Chiang and Rong-Jaye J. Chen. Distributed faulttolerant routing in Kautz networks. Journal of Parallel and Distributed Computing, 20(1):99–106, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1010/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1010/production/pdf.

Choi:1996:DPE

[CC96] Lynn Choi and Andrew A. Chien. The design and performance evaluation of the DI-multicomputer. Journal of Parallel and Distributed Computing, 36(2):119–143, August 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0094/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0094/production/pdf.

Cavendish:2008:DXP

[CC08] Dirceu Cavendish and K. Selçuk Candan. Distributed XML processing: Theory and applications. Journal of Parallel and Distributed Computing, 68(8):1054–1069, August 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1990:DVA

[CCC90] Zen Chen, Chih-Chi Chang, and Tsorng-Lin Chia. On the design of VLSI architectures for parallel execution of DO loops. Journal of Parallel and Distributed Computing, 8(4):393–399, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1992:EMF

[CCC92] Yen-Cheng Chen, Wen-Tsuen Chen, and Gen-Huey Chen. Efficient median finding and its application to two-variable linear programming on mesh-connected computers with multiple broadcasting. Journal of Parallel and Distributed Computing, 15(1):79–84, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cao:2004:FAH

[CCC^{+}04] Jiannong Cao, Min Cao, Alvin S. T. Chan, Gengfeng Wu, and S. K. Sajal K. Das. A framework for architecting and highlevel programming support of CORBA applications. Journal of Parallel and Distributed Computing, 64(6):725–739, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chan:1996:DED

[CCCM96] M. Y. Chan, F. Chin, C. N. Chu, and W. K. Mak. Dilation-5 embedding of 3-dimensional grids into hypercubes. Journal of Parallel and Distributed Computing, 33(1):98–106, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0029/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0029/production/
pdf.

Chien:2003:EAP

[CCEB03] Andrew Chien, Brad Calder, Stephen Elbert, and Karan Bhatia. Entropia: architecture and performance of an enterprise desktop grid system. Journal of Parallel and Distributed Computing, 63(5):597–610, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2009:NMC

[CCHC09] Chih-Yung Chang, Yu-Chieh Chen, Li-Ling Hung, and Sheng-Wen Chang. A novel multi-channel MAC protocol with directional antenna for enhancing spatial reuse and bandwidth utilization in WLANs. Journal of Parallel and Distributed Computing, 69(10):824–837, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Callahan:1988:EII

[CCK88] D. Callahan, J. Cocke, and K. Kennedy. Estimating interlock and improving balance for pipelined architectures. Journal of Parallel and Distributed Computing, 5(4):334–358, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2008:MJR

[CCK^{+}08] Chung-Kai Chen, Cheng-Wei Chen, Chien-Tan Ko, Jenq-Kuen Lee, and Jyh-Cheng Chen. Mobile Java RMI support over heterogeneous wireless networks: a case study. Journal of Parallel and Distributed Computing, 68(11):1425–1436, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Crowl:1994:AMP

[CCLS94] Lawrence A. Crowl, Mark E. Crovella, Thomas J. LeBlanc, and Michael L. Scott. The advantages of multiple parallelizations in combinatorial search. Journal of Parallel and Distributed Computing, 21(1):110–123, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1045/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1045/production/pdf.

Caselli:1992:TPI

[CCM92] Stefano Caselli, Gianni Conte, and Ugo Malavolta. Topology and process interaction in concurrent architectures: A GSPN modeling approach. Journal of Parallel and Distributed Computing, 15(3):270–281, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Caselli:2001:DAG

[CCM01] S. Caselli, G. Conte, and P. Marenzoni. A distributed algorithm for GSPN reachability graph generation. Journal of Parallel and Distributed Computing, 61(1):79–95, January 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1664;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1664/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1664/ref.

Costa:2006:ROA

[CCM^{+}06] J. J. Costa, T. Cortes, X. Martorell, E. Ayguade, and J. Labarta. Running OpenMP applications efficiently on an everything-shared SDSM. Journal of Parallel and Distributed Computing, 66(5):647–658, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Crane:2006:GEM

[CCN06] Jason C. Crane, Forrest W. Crawford, and Sarah J. Nelson. Grid enabled magnetic resonance scanners for near real-time medical image processing. Journal of Parallel and Distributed Computing, 66(12):1524–1533, December 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Caminero:2006:MMR

[CCQ^{+}06] B. Caminero, C. Carrión, F. J. Quiles, J. Duato, and S. Yalamanchili. MMR: A MultiMedia Router architecture to support hybrid workloads. Journal of Parallel and Distributed Computing, 66(2):307–321, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiu:1994:FRC

[CCR94] Ge-Ming M. Chiu, S. Chalasani, and C. S. Raghavendra. Flexible routing criteria for circuit-switched hypercubes. Journal of Parallel and Distributed Computing, 22(2):279–294, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1087/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1087/production/
pdf.

Chase:1992:PPP

[CCRS92] Craig M. Chase, Alex L. Cheung, Anthony P. Reeves, and Mark R. Smith. Paragon: a parallel programming environment for scientific applications using communication structures. Journal of Parallel and Distributed Computing, 16(2):79–91, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2006:BCH

[CCS06] Chao-Tsun Chang, Chih-Yung Chang, and Jang-Ping Sheu. BlueCube: Constructing a hypercube parallel computing and communication environment over Bluetooth radio systems. Journal of Parallel and Distributed Computing, 66(10):1243– 1258, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cook:1995:ISS

[CD95] Gregory W. Cook and Edward J. Delp. An investigation of scalable SIMD I/O techniques with application to parallel JPEG compression. Journal of Parallel and Distributed Computing, 30(2):111–128, November 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1132/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1132/production/pdf.

Coulaud:1998:HLC

[CD98] O. Coulaud and E. Dillon. Para++: a high level C++ interface for message passing. Journal of Parallel and Distributed Computing, 51(1):46–62, May 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1450/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1450/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1450/production/ref.

Chen:2004:RWP

[CDB04] Zhongqiang Chen, Alex Delis, and Henry L. Bertoni. Radiowave propagation prediction using ray-tracing techniques on a network of workstations (NOW). Journal of Parallel and Distributed Computing, 64(10):1127–1156, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choi:2005:RDR

[CDCD05] Wook Choi, Sajal K. Das, Jiannong Cao, and Ajoy K. Datta. Randomized dynamic route maintenance for adaptive routing in multihop mobile ad hoc networks. Journal of Parallel and Distributed Computing, 65(2):107–123, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cicerone:2001:CPR

[CDF01] Serafino Cicerone, Gabriele Di Stefano, and Michele Flammini. Compact-port routing models and applications to distance-hereditary graphs. Journal of Parallel and Distributed Computing, 61(10):1472–1488, October 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1728;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1728/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1728/ref.

Chen:1984:MLA

[CDH84] Steven S. Chen, Jack J. Dongarra, and Christopher C. Hsiung. Multiprocessing linear algebra algorithms on the Cray X-MP-2: Experiences with small granularity. Journal of Parallel and Distributed Computing, 1(1):22–31, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chu:1989:MIO

[CDJ^{+}89] C. Henry Chu, Edward J. Delp, Leah H. Jamieson, Howard Jay Siegel, Francis J. Weil, and Andrew B. Whinston. A model for an intelligent operating system for executing image understanding tasks on a reconfigurable parallel architecture. Journal of Parallel and Distributed Computing, 6(3):598–622, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chockler:2009:SIJ

[CDJL09] Gregory Chockler, Eliezer Dekel, Joseph JaJa, and Jimmy Lin. Special issue of the Journal of Parallel and Distributed Computing: Cloud computing. Journal of Parallel and Distributed Computing, 69(9):813, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cucchiara:1995:DCO

[CDP95] Rita Cucchiara, Luigi Di Stefano, and Massimo Piccardi. Detection of circular objects by wave propagation on a mesh-connected computer. Journal of Parallel and Distributed Computing, 31(1):77–87, November 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1146/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1146/production/pdf.

Cosnard:1990:STF

[CDR90] Michel Cosnard, Jean Duprat, and Yves Robert. Systolic triangularization over finite fields. Journal of Parallel and Distributed Computing, 9(3):252–260, July 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chitnis:2009:ATI

[CDR09a] Laukik Chitnis, Alin Dobra, and Sanjay Ranka. Analyzing the techniques that improve fault tolerance of aggregation trees in sensor networks. Journal of Parallel and Distributed Computing, 69(12):950–960, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chitnis:2009:FTA

[CDR09b] Laukik Chitnis, Alin Dobra, and Sanjay Ranka. Fault tolerant aggregation in heterogeneous sensor networks. Journal of Parallel and Distributed Computing, 69(2):210–219, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chan:1999:CGP

[CDRC99] Albert Chan, Frank Dehne, and Andrew Rau-Chaplin. Coarsegrained parallel geometric search. Journal of Parallel and Distributed Computing, 57(2):224–235, May 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1527/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1527/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1527/production/ref.

Chakrabarti:1997:MSA

[CDY97] Soumen Chakrabarti, James Demmel, and Katherine Yelick. Models and scheduling algorithms for mixed data and task parallel programs. Journal of Parallel and Distributed Computing, 47(2):168–184, December 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1413/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1413/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1413/production/ref.

Chen:1995:ETP

[CEF^{+}95] Song Chen, Mary M. Eshaghian, Richard F. Freund, Jerry L. Potter, and Ying-Chieh Wu. Evaluation of two programming paradigms for heterogeneous computing. Journal of Parallel and Distributed Computing, 31(1):41–55, November 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1143/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1143/production/
pdf.

Chau:2007:MIP

[CEGS07] Ming Chau, Didier El Baz, Ronan Guivarch, and Pierre Spiteri. MPI implementation of parallel subdomain methods for linear and nonlinear convection–diffusion problems. Journal of Parallel and Distributed Computing, 67(5):581–591, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carlson:1988:PCD

[CF88] W. W. Carlson and J. A. B. Fortes. On the performance of combined data flow and control flow systems: experiments using two iterative algorithms. Journal of Parallel and Distributed Computing, 5(4):359–382, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Connolly:1998:FTF

[CF98] Maura P. Connolly and Patrick Fitzpatrick. Fault tolerant Faddeeva algorithm. Journal of Parallel and Distributed Computing, 53(1):78–89, August 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1474/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1474/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1474/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1491/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1491/
production/pdf.

Ceri:1986:OJB

[CG86] S. Ceri and G. Gottlob. Optimizing joins between two partitioned relations in distributed databases. Journal of Parallel and Distributed Computing, 3(2):183–205, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cobb:2002:SGL

[CG02] Jorge A. Cobb and Mohamed G. Gouda. Stabilization of general loop-free routing. Journal of Parallel and Distributed Computing, 62(5):922–944, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1830;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1830/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1830/ref.

Colajanni:1998:TBR

[CGA98] M. Colajanni, V. Grassi, and M. Angelaccio. Thresholdbased reconfiguration strategies for gracefully degradable parallel computations. Journal of Parallel and Distributed Computing, 55(1):138–151, November 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1492/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1492/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1492/production/ref.

Chockler:2009:RDS

[CGG^{+}09] Gregory Chockler, Seth Gilbert, Vincent Gramoli, Peter M. Musial, and Alex A. Shvartsman. Reconfigurable distributed storage for dynamic networks. Journal of Parallel and Distributed Computing, 69(1):100–116, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Calegari:1997:PIB

[CGKK97] Patrice Calégari, Frédéric Guidec, Pierre Kuonen, and Daniel Kobler. Parallel island-based genetic algorithm for radio network design. Journal of Parallel and Distributed Computing, 47(1):86–90, November 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1397/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1397/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1397/production/ref.

Chatterjee:1995:GLA

[CGL^{+}95] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua H. Teng. Generating local addresses and communication sets for data-parallel programs. Journal of Parallel and Distributed Computing, 26(1):72–84, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1049/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1049/
production/pdf.

Chatterjee:1996:AAA

[CGO^{+}96] Siddhartha Chatterjee, John R. Gilbert, Leonid Oliker, Robert Schreiber, and Thomas J. Sheffler. Algorithms for automatic alignment of arrays. Journal of Parallel and Distributed Computing, 38(2):145–157, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0137/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0137/production/pdf.

Culler:1993:TCC

[CGSV93] David E. Culler, Seth Copen Goldstein, Klaus Erik Schauser, and Thorsten Von Eicken. TAM — a compiler controlled Threaded Abstract Machine. Journal of Parallel and Distributed Computing, 18(3):347–370, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1070/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1070/production/pdf.

Chronopoulos:2003:EGB

[CGW^{+}03] Anthony T. Chronopoulos, Daniel Grosu, Andrew M. Wissink, Manuel Benche, and Jingyu Liu. An efficient 3D grid based scheduling for heterogeneous systems. Journal of Parallel and Distributed Computing, 63(9):827–837, September 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Colbourn:1992:CFA

[CH92] Charles J. Colbourn and Katherine Heinrich. Conflict-free access to parallel memories. Journal of Parallel and Distributed Computing, 14(2):193–200, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2006:TBR

[CH06a] Sheng-Hsiung Chen and Ting-Lu Huang. A tight bound on remote reference time complexity of mutual exclusion in the read-modify-write model. Journal of Parallel and Distributed Computing, 66(11):1455–1471, November 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2006:CDF

[CH06b] Yu Chen and Kai Hwang. Collaborative detection and filtering of shrew DDoS attacks using spectral analysis. Journal of Parallel and Distributed Computing, 66(9):1137–1151, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:1994:CTA

[Cha94] Ye-In I. Chang. Comments on “Two Algorithms for Mutual Exclusion in Real-Time Distributed Computer Systems”. Journal of Parallel and Distributed Computing, 23(3):449–454, December 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1155/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1155/
production/pdf.

Charney:1995:RNB

[Cha95] Mark Charney. The role of network bandwidth in barrier synchronization. Journal of Parallel and Distributed Computing, 28(2):202–212, August 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1100/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1100/production/pdf.

Chang:1996:SSD

[Cha96] Ye-In Chang. A simulation study on distributed mutual exclusion. Journal of Parallel and Distributed Computing, 33(2):107–121, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1996.0030/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0030/production/pdf.

Chiu:2005:TOG

[CHC05] Ge-Ming Chiu, Chih-Ming Hsiao, and Wen-Ray Chang. Total ordering group communication protocol based on coordinating sequencers for multiple overlapping groups. Journal of Parallel and Distributed Computing, 65(4):437–447, April 2005. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1986:DMS

[Che86] M. C. Chen. A design methodology for synthesizing parallel algorithms and architectures. Journal of Parallel and Distributed Computing, 3(4):461–491, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cheung:1989:ADC

[Che89] To-Yat Cheung. An algorithm with decentralized control for sorting files in a network. Journal of Parallel and Distributed Computing, 7(3):464–481, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cheng:1990:SAQ

[Che90] Kam Hoi Cheng. A simultaneous access queue. Journal of Parallel and Distributed Computing, 9(1):83–86, May 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1992:GEI

[Che92] Wen-Tsuen Chen. Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 16(2):77–78, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2005:IOA

[Che05] Yu-Wei Chen. Improved one-to-all broadcasting algorithms on faulty SIMD hypercubes. Journal of Parallel and Distributed Computing, 65(12):1596–1600, December 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cook:2001:APG

[CHGM01] Diane J. Cook, Lawrence B. Holder, Gehad Galal, and Ron Maglothin. Approaches to parallel graph-based knowledge discovery. Journal of Parallel and Distributed Computing, 61(3):427–446, March 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1696;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1696/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1696/ref.

Chiola:1992:SIP

[Chi92] Giovanni Chiola. Special issue on Petri net modeling of parallel computers. Journal of Parallel and Distributed Computing, 15 (3):169–170, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chien:1995:CAC

[Chi95]  Andrew A. Chien.  Concurrent Aggregates (CA): de-

sign and experience with a concurrent object-oriented language based on aggregates. Journal of Parallel and Distributed Computing, 25(2):174–196, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1039/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1039/production/pdf.

Chowdhury:1990:GLS

[Cho90] S. Chowdhury. The Greedy load sharing algorithm. Journal of Parallel and Distributed Computing, 9(1):93–99, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choudhary:1993:PSG

[Cho93] Alok Choudhary. Parallel I/O systems: Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 17 (1–2):1–3, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chrisochoides:1994:MAS

[CHR94] Nikos Chrisochoides, Elias Houstis, and John Rice. Mapping algorithms and software environment for data parallel PDE iterative solvers. Journal of Parallel and Distributed Computing, 21(1):75–95, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1043/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1043/production/pdf.

Chung:1986:OQD

[CI86] Chin-Wan Chung and Keki B. Irani. An optimization of queries in distributed database systems. Journal of Parallel and Distributed Computing, 3(2):137–157, June 1986. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Calamoneri:2003:IRL

[CI03] Tiziana Calamoneri and Miriam Di Ianni. Interval routing & layered cross product: compact routing schemes for butterflies, meshes of trees, fat trees and Bene networks. Journal of Parallel and Distributed Computing, 63(11):1017–1025, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cosnard:1999:CDR

[CJ99a] Michel Cosnard and Emmanuel Jeannot. Compact DAG representation and its dynamic scheduling. Journal of Parallel and Distributed Computing, 58(3):487–514, September 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1566/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1566/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1566/production/ref.

Crivelli:1999:PPL

[CJ99b] S. Crivelli and E. R. Jessup. The PMESC programming library for distributed-memory MIMD computers. Journal of Parallel and Distributed Computing, 57(3):295–321, June 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1539/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1539/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1539/production/ref.

Chi:2007:GPQ

[CJ07] Hongmei Chi and Edward L. Jones. Generating parallel quasirandom sequences via randomization. Journal of Parallel and Distributed Computing, 67(7):876–881, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cholvi:2009:IDM

[CJA09] Vicent Cholvi, Ernesto Jiménez, and Antonio Fernández Anta. Interconnection of distributed memory models. Journal of Parallel and Distributed Computing, 69(3):295–306, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cosnard:2004:CDR

[CJY04] Michel Cosnard, Emmanuel Jeannot, and Tao Yang. Compact DAG representation and its symbolic scheduling. Journal of Parallel and Distributed Computing, 64(8):921–935, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Callahan:1988:AIS

[CK88] D. Callahan and K. Kennedy. Analysis of interprocedural side effects in a parallel programming environment. Journal of Parallel and Distributed Computing, 5(5):517–550, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cox:1991:TSS

[CK91] Christopher L. Cox and James A. Knisely. A tridiagonal system solver for distributed memory parallel processors with vector nodes. Journal of Parallel and Distributed Computing, 13(3): 325–331, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiou:1997:DEE

[CK97] Hsien-Kuang Chiou and Willard Korfhage. Detecting ENF event predicates in distributed systems. Journal of Parallel and Distributed Computing, 44(1):1–13, July 10, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1348/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1348/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1348/production/ref.

Camp:2006:FBD

[CK06] Tracy Camp and Phil Kearns. Fast batched data transfer with flush channels: a performance analysis. Journal of Parallel and Distributed Computing, 66(6):780–795, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chandra:2008:DSB

[CK08] Punit Chandra and Ajay D. Kshemkalyani. Data-stream-based global event monitoring using pairwise interactions. Journal of Parallel and Distributed Computing, 68(6):729–751, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2000:EAT

[CKK00] Fangzhe Chang, Vijay Karamcheti, and Zvi Kedem. Exploiting application tunability for efficient, predictable resource management in parallel and distributed systems. Journal of Parallel and Distributed Computing, 60(11):1420–1445, November 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1660;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1660/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1660/ref.

Cho:1999:CRP

[CKL99] Sangyeun Cho, Jinseok Kong, and Gyungho Lee. Coherence and replacement protocol of DICE — a busbased COMA multiprocessor. Journal of Parallel and Distributed Computing, 57(1):14–32, April 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1524/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1524/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1524/production/ref.

Chiu:2004:PED

[CkLCK04] Steve C. Chiu, Wei keng Liao, Alok N. Choudhary, and Mahmut T. Kandemir. Processor-embedded distributed smart disks for I/O-intensive workloads: architectures, performance models and evaluation. Journal of Parallel and Distributed Computing, 64(3):427–446, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiu:2005:PED

[CkLCK05] Steve C. Chiu, Wei keng Liao, Alok N. Choudhary, and Mahmut T. Kandemir. Processor-embedded distributed smart disks for I/O-intensive workloads: architectures, performance models and evaluation. Journal of Parallel and Distributed Computing, 65(4):532–551, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chatzigiannakis:2007:FTE

[CKN07] Ioannis Chatzigiannakis, Athanasios Kinalis, and Sotiris Nikoletseas. Fault-tolerant and efficient data propagation in wireless sensor networks using local, additional network information. Journal of Parallel and Distributed Computing, 67(4):456–473, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Campbell:2004:HCI

[CKO04] Stuart Campbell, Mohan Kumar, and Stephan Olariu. The hierarchical cliques interconnection network. Journal of Parallel and Distributed Computing, 64(1):16–28, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Clark:1985:SPP

[CL85] James J. Clark and Peter D. Lawrence. A systolic parallel processor for the rapid computation of multiresolution edge images using the ∇^{2}G operator. Journal of Parallel and Distributed Computing, 2(3):302–319, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cavallaro:1988:CAS

[CL88] Joseph R. Cavallaro and Franklin T. Luk. CORDIC arithmetic for an SVD processor. Journal of Parallel and Distributed Computing, 5(3):271–290, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cormen:1990:HSR

[CL90] Thomas H. Cormen and Charles E. Leiserson. A hyperconcentrator switch for routing bit-serial messages. Journal of Parallel and Distributed Computing, 10(3):193–204, November 1990.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chan:1991:DFT

[CL91a] Mee Yee Chan and Shiang-Jen Lee. Distributed fault-tolerant embeddings of rings in hypercubes. Journal of Parallel and Distributed Computing, 11(1):63–71, January 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1991:SIJ

[CL91b] Guan-Ing Chen and Ten-Hwang Lai. Scheduling independent jobs on partitionable hypercubes. Journal of Parallel and Distributed Computing, 12(1):74–78, May 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chan:1993:FTP

[CL93] Mee Yee Chan and Shiang-Jen Lee. Fault-tolerant permutation routing in hypercubes. Journal of Parallel and Distributed Computing, 17(4):277–281, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1027/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1027/production/pdf.

Cheung:1996:RLM

[CL96] Steven Cheung and Francis C. M. Lau. Routing with locality on meshes with buses. Journal of Parallel and Distributed Computing, 33(1):84–90, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0027/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0027/production/pdf.

Camp:2003:AMB

[CL03a] Tracy Camp and Yu Liu. An adaptive mesh-based protocol for geocast routing. Journal of Parallel and Distributed Computing, 63(2):196–213, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cortes:2003:TAH

[CL03b] T. Cortes and J. Labarta. Taking advantage of heterogeneity in disk arrays. Journal of Parallel and Distributed Computing, 63(4):448–464, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chien:2009:EBL

[CL09] Hung-Yu Chien and Chi-Sung Laih. ECC-based lightweight authentication protocol with untraceability for low-cost RFID. Journal of Parallel and Distributed Computing, 69(10):848–853, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chung:2009:SNU

[CLL09] Yu-Chi Chung, Lanturn Lin, and Chiang Lee. Scheduling nonuniform data with expected-time constraint in wireless multichannel environments. Journal of Parallel and Distributed Computing, 69(3):247–260, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chiarulli:1990:OBC

[CLM90] D. M. Chiarulli, S. P. Levitan, and R. G. Melhem. Optical bus control for distributed multiprocessors. Journal of Parallel and Distributed Computing, 10(1):45–54, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Codenotti:1990:ATT

[CLR90] B. Codenotti, G. Lotti, and F. Romani. Area-time trade-offs for matrix-vector multiplication. Journal of Parallel and Distributed Computing, 8(1):52–59, January 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Contreras:2000:DSF

[CLRW00] Javier Contreras, Arturo Losi, Mario Russo, and Felix F. Wu. DistOpt: a software framework for modeling and evaluating optimization problem solutions in distributed environments. Journal of Parallel and Distributed Computing, 60(6):741– 763, June 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1626;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1626/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1626/ref.

Chinn:1996:MAR

[CLT96] Donald D. Chinn, Tom Leighton, and Martin Tompa. Minimal adaptive routing on the mesh with bounded queue size. Journal of Parallel and Distributed Computing, 34(2):154–170, May 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0052/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0052/production/
pdf.

Chen:1995:SRT

[CLV95] H. J. Chen, T. D. C. Little, and D. Venkatesh. A storage and retrieval technique for scalable delivery of MPEG-Encoded video. Journal of Parallel and Distributed Computing, 30(2):180–189, November 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1136/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1136/production/pdf.

Corradi:2000:POM

[CLZ00] Antonio Corradi, Letizia Leonardi, and Franco Zambonelli. Parallel objects migration: a fine grained approach to load distribution. Journal of Parallel and Distributed Computing, 60(1):48–71, January 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1593/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1593/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1593/production/ref.

Cai:2002:COD

[CLZ02] Wentong Cai, Bu-Sung Lee, and Junlan Zhou. Causal order delivery in a multicast environment: An improved algorithm. Journal of Parallel and Distributed Computing, 62(1):111–131, January 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1774;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1774/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1774/ref.

Clark:1992:DDC

[CM92] Henry Clark and Bruce McMillin. DAWGS — a distributed compute server utilizing idle workstations. Journal of Parallel and Distributed Computing, 14(2):175–186, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chung:1993:MSN

[CM93] Sang-Hwa H. Chung and Dan I. Moldovan. Modeling semantic networks on the Connection Machine. Journal of Parallel and Distributed Computing, 17(1–2):152–163, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1014/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1014/production/
pdf.

Chandra:2003:FAS

[CM03] Ramesh Chandra and C. Siva Ram Murthy. A faster algorithm for solving linear algebraic equations on the star graph. Journal of Parallel and Distributed Computing, 63(4):465–480, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Calamoneri:2004:EAC

[CM04] Tiziana Calamoneri and Annalisa Massini. Efficient algorithms for checking the equivalence of multistage interconnection networks. Journal of Parallel and Distributed Computing, 64(1): 135–150, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Clark:1992:CDF

[CMS92] W. Edwin Clark, Gregory L. McColm, and W. Richard Stark. On the complexity of deadlock-free programs on a ring of processors. Journal of Parallel and Distributed Computing, 16(1): 67–71, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Clementi:2004:RRO

[CMS04] Andrea E. F. Clementi, Angelo Monti, and Riccardo Silvestri. Round Robin is optimal for fault-tolerant broadcasting on wireless networks. Journal of Parallel and Distributed Computing, 64(1):89–96, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ciardo:1992:ACF

[CMT92] Gianfranco Ciardo, Jogesh K. Muppala, and Kishor S. Trivedi. Analyzing concurrent and fault-tolerant software using stochastic reward nets. Journal of Parallel and Distributed Computing, 15(3):255–269, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cai:1993:GVB

[CMT93] Wentong T. Cai, Wendy J. Milne, and Stephen J. Turner. Graphical views of the behavior of parallel programs. Journal of Parallel and Distributed Computing, 18(2):223–230, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1058/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1058/production/
pdf.

Choi:1993:EAM

[CN93] Hyeong-Ah A. Choi and B. Narahari. Efficient algorithms for mapping and partitioning a class of parallel computations. Journal of Parallel and Distributed Computing, 19(4):349–363, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1117/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1117/production/pdf.

Chatzigiannakis:2003:DCA

[CNS03] Ioannis Chatzigiannakis, Sotiris Nikoletseas, and Paul Spirakis. Distributed communication algorithms for ad hoc mobile networks. Journal of Parallel and Distributed Computing, 63(1): 58–74, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cohn:1990:IMO

[Coh90] E. R. Cohn. Implementing the multiprefix operation efficiently. Journal of Parallel and Distributed Computing, 10(1):29–34, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cormen:1993:FPD

[Cor93] Thomas H. Cormen. Fast permuting on disk arrays. Journal of Parallel and Distributed Computing, 17(1–2):41–57, January/ February 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1004/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1004/
production/pdf.

Chen:1995:CTT

[COS^{+}95] G-H Chen, S. Olariu, J. L. Schwing, B-F Wang, and J. Zhang. Constant-time tree algorithms on reconfigurable meshes on size n × n. Journal of Parallel and Distributed Computing, 26(2):137–150, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1055/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1055/production/pdf.

Couch:1993:CCS

[Cou93] Alva L. Couch. Categories and context in scalable execution visualization. Journal of Parallel and Distributed Computing, 18(2):195–204, June 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1056/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1056/production/pdf.

Culler:1990:ETS

[CP90] D. E. Culler and G. M. Papadopoulos. The explicit token store. Journal of Parallel and Distributed Computing, 10(4):289–308, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choudhary:1991:IEH

[CP91] Alok N. Choudhary and Ravi Ponnusamy. Implementation and evaluation of Hough transform algorithms on a shared-memory multiprocessor. Journal of Parallel and Distributed Computing, 12(2):178–188, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choudhary:1992:PIE

[CP92] Alok N. Choudhary and Ravi Ponnusamy. Parallel implementation and evaluation of a motion estimation system algorithm using several data decomposition strategies. Journal of Parallel and Distributed Computing, 14(1):50–65, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Conn:1994:PRS

[CP94] H. E. Conn and L. J. Podrazik. Parallel recurrence solvers for vector and SIMD supercomputers. Journal of Parallel and Distributed Computing, 23(3):435–441, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1153/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1153/production/pdf.

Cabillic:1997:SEP

[CP97] Gilbert Cabillic and Isabelle Puaut. Stardust: An environment for parallel programming on networks of heterogeneous workstations. Journal of Parallel and Distributed Computing, 40(1):65–80, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1271/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1271/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1271/production/ref.

Calamoneri:1998:ODC

[CP98] Tiziana Calamoneri and Rossella Petreschi. Orthogonally drawing cubic graphs in parallel. Journal of Parallel and Distributed Computing, 55(1):94–108, November 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1495/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1495/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1495/production/ref.

Cruz:1999:TPD

[CP99] John Cruz and Kihong Park. Toward performance-driven system support for distributed computing in clustered environments. Journal of Parallel and Distributed Computing, 59(2):132–154, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1575/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1575/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1575/production/ref.

Choi:2001:ECA

[CP01] Yungho Choi and Timothy Mark Pinkston. Evaluation of crossbar architectures for deadlock recovery routers. Journal of Parallel and Distributed Computing, 61(1):49–78, January 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1669;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1669/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1669/ref.

Calamoneri:2004:HLS

[CP04a] Tiziana Calamoneri and Rossella Petreschi. L (h,1)-labeling subclasses of planar graphs. Journal of Parallel and Distributed Computing, 64(3):414–426, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Choi:2004:EQD

[CP04b] Yungho Choi and T. Timothy Mark Pinkston. Evaluation of queue designs for true fully adaptive routers. Journal of Parallel and Distributed Computing, 64(5):606–616, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chandra:2005:TAA

[CP05] Sumir Chandra and Manish Parashar. Towards autonomic application-sensitive partitioning for SAMR applications. Journal of Parallel and Distributed Computing, 65(4):519–531, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Conant:2003:PGI

[CPO^{+}03] Gavin C. Conant, Steven J. Plimpton, William Old, Andreas Wagner, Pamela R. Fain, Theresa R. Pacheco, and Grant Heffelfinger. Parallel Genehunter: implementation of a linkage analysis package for distributed-memory architectures. Journal of Parallel and Distributed Computing, 63(7–8):674–682, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Clement:1995:OCC

[CQ95] Mark J. Clement and Michael J. Quinn. Overlapping computations, communications and I/O in parallel sorting. Journal of Parallel and Distributed Computing, 28(2):162–172, August 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1097/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1097/production/
pdf.

Carmona:1991:MSP

[CR91] Edward A. Carmona and Michael D. Rice. Modeling the serial and parallel fractions of a parallel algorithm. Journal of Parallel and Distributed Computing, 13(3):286–298, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carlisle:1996:SCC

[CR96] Martin C. Carlisle and Anne Rogers. Software caching and computation migration in Olden. Journal of Parallel and Distributed Computing, 38(2):248–255, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0145/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0145/production/pdf.

Ciorba:2008:ESS

[CRA^{+}08] Florina M. Ciorba, Ioannis Riakiotakis, Theodore Andronikos, George Papakonstantinou, and Anthony T. Chronopoulos. Enhancing self-scheduling algorithms via synchronization and weighting. Journal of Parallel and Distributed Computing, 68 (2):246–264, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cortes:2002:AIL

[CRC^{+}02] A. Cortés, A. Ripoll, F. Cedó, M. A. Senar, and E. Luque. An asynchronous and iterative load balancing algorithm for discrete load model. Journal of Parallel and Distributed Computing, 62(12):1729–1746, December 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Copty:1994:DPA

[CRFS94] Nawal Copty, Sanjay Ranka, Geoffrey Fox, and Ravi V. Shankar. A data parallel algorithm for solving the region growing problem on the Connection Machine. Journal of Parallel and Distributed Computing, 21(1):160–168, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1049/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1049/production/pdf.

Carbunar:2009:ETD

[CRK^{+}09] Bogdan Carbunar, Murali Krishna Ramanathan, Mehmet Koyutürk, Suresh Jagannathan, and Ananth Grama. Efficient tag detection in RFID systems. Journal of Parallel and Distributed Computing, 69(2):180–196, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Charcranoon:2004:LSP

[CRL04] Saravut Charcranoon, Thomas G. Robertazzi, and Serge Luryi. Load sequencing for a parallel processing utility. Journal of Parallel and Distributed Computing, 64(1):29–35, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chalasani:1994:FTR

[CRV94] S. Chalasani, C. S. Raghavendra, and A. Varma. Fault-tolerant routing in MIN-based supercomputers. Journal of Parallel and Distributed Computing, 22(2):154–167, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1078/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1078/production/pdf.

Cheng:1992:OCA

[CS92] L. Cheng and A. A. Sawchuk. A one-copy algorithm for 2D shuffles for optical Omega networks. Journal of Parallel and Distributed Computing, 16(1):54–66, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See erratum [Ano93e].

Chang:1993:OLS

[CS93a] Yi-Chieh C. Chang and Kang G. Shin. Optimal load sharing in distributed real-time systems. Journal of Parallel and Distributed Computing, 19(1):38–50, September 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1088/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1088/production/pdf.

Cheng:1993:VNA

[CS93b] L. Cheng and A. A. Sawchuk. Volume 16, number 1 (1992), in the article “A One-Copy Algorithm for 2-D Shuffles for Optical Omega Networks,” by L. Cheng and A. A. Sawchuk, pages 54–66. Journal of Parallel and Distributed Computing, 17(1–2):183, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1016/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1016/production/pdf.

Choi:1993:RCS

[CS93c] Sang Bang Choi and Arun K. Somani. Rearrangeable circuitswitched hypercube architectures for routing permutations. Journal of Parallel and Distributed Computing, 19(2):125–130, October 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1097/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1097/
production/pdf.

Chen:1995:EBT

[CS95a] Woei-Kae K. Chen and M. F. M. Stallmann. On embedding binary trees into hypercubes. Journal of Parallel and Distributed Computing, 24(2):132–138, February 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1013/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1013/production/pdf.

Chen:1995:IEB

[CS95b] Yaw-Huei H. Chen and Stanley Y. W. Su. Identificationand elimination-based parallel query processing techniques for object-oriented databases. Journal of Parallel and Distributed Computing, 28(2):130–148, August 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1095/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1095/production/pdf.

Choy:1995:EIS

[CS95c] Manhoi Choy and Ambuj K. Singh. Efficient implementation of synchronous communication over asynchronous networks. Journal of Parallel and Distributed Computing, 26(2):166–180, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1057/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1057/
production/pdf.

Cao:2000:ADC

[CS00] Guohong Cao and Mukesh Singhal. An adaptive distributed channel allocation strategy for mobile cellular networks. Journal of Parallel and Distributed Computing, 60(4):451–473, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1999.1614;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1614/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1614/ref.

Chen:2006:CLS

[CS06a] Chunxi Chen and Bertil Schmidt. Constructing large suffix trees on a computational Grid. Journal of Parallel and Distributed Computing, 66(12):1512–1523, December 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2006:FSU

[CS06b] Kevin F. Chen and Edwin H.-M. Sha. The fat-stack and universal routing in interconnection networks. Journal of Parallel and Distributed Computing, 66(5):705–715, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2008:SSD

[CS08] Shigang Chen and Yuval Shavitt. SoMR: a scalable distributed QoS multicast routing protocol. Journal of Parallel and Distributed Computing, 68(2):137–149, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Carey:1994:PSO

[CSSY94] Graham Carey, Joe Schmidt, Vineet Singh, and Dennis Yelton. A prototype scalable, object-oriented finite element solver on multicomputers. Journal of Parallel and Distributed Computing, 20(3):357–379, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1033/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1033/production/pdf.

Chen:2008:ASS

[CSW08] Yong Chen, Xian-He Sun, and Ming Wu. Algorithm-system scalability of heterogeneous computing. Journal of Parallel and Distributed Computing, 68(11):1403–1412, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cao:2003:SLB

[CSWD03] Jiannong Cao, Yudong Sun, Xianbin Wang, and Sajal K. Das. Scalable load balancing on distributed Web servers using mobile agents. Journal of Parallel and Distributed Computing, 63(10): 996–1005, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:1993:DER

[CT93] Shenze Z. Chen and Don Towsley. The design and evaluation of RAID 5 and parity striping disk array architectures. Journal of Parallel and Distributed Computing, 17(1–2):58–74, January/February 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1005/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1005/production/pdf.

Choudhary:1994:CCL

[CT94] Akok Choudhary and Rajeev Thakur. Connected component labeling on coarse grain parallel computers: an experimental study. Journal of Parallel and Distributed Computing, 20(1):78–83, January 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1007/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1007/production/pdf.

Calvin:1996:MTB

[CT96] Christophe Calvin and Denis Trystram. Matrix transpose for block allocations on torus and de Bruijn networks. Journal of Parallel and Distributed Computing, 34(1):36–49, April 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0044/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0044/production/
pdf.

Collier:2004:SOS

[CT04] Travis C. Collier and Charles Taylor. Self-organization in sensor networks. Journal of Parallel and Distributed Computing, 64 (7):866–873, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2008:DFA

[CTCX08] Dan Chen, Stephen John Turner, Wentong Cai, and Muzhou Xiong. A decoupled federate architecture for high level architecture-based distributed simulation. Journal of Parallel and Distributed Computing, 68(11):1487–1503, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Casanova:1999:SPP

[CTD99] Henri Casanova, Michael G. Thomason, and Jack J. Dongarra. Stochastic performance prediction for iterative algorithms in distributed environments. Journal of Parallel and Distributed Computing, 58(1):68–91, July 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1541/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1541/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1543/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1543/production/pdf.

Congiusta:2008:SOM

[CTT08] Antonio Congiusta, Domenico Talia, and Paolo Trunfio. Service-oriented middleware for distributed data mining on the Grid. Journal of Parallel and Distributed Computing, 68(1):3– 15, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Climent:1999:BRD

[CTZ99] Joan-Josep Climent, Leandro Tortosa, and Antonio Zamora. A BSP recursive divide and conquer algorithm to compute the inverse of a tridiagonal matrix. Journal of Parallel and Distributed Computing, 59(3):423–444, December 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1586/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1586/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1586/production/ref.

Chandrasekaran:1990:MOA

[CV90] S. Chandrasekaran and S. Venkatesan. A message-optimal algorithm for distributed termination detection. Journal of Parallel and Distributed Computing, 8(3):245–252, March 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chlebus:1991:PQ

[CV91] Bogdan S. Chlebus and Imrich Vrto. Parallel quicksort. Journal of Parallel and Distributed Computing, 11(4):332–337, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Charles:2008:PDS

[CvdBL^{+}08] W. M. Charles, E. van den Berg, H. X. Lin, A. W. Heemink, and M. Verlaan. Parallel and distributed simulation of sediment dynamics in shallow water using particle decomposition approach. Journal of Parallel and Distributed Computing, 68 (6):717–728, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chin:2009:HLS

[CVJ09] Teo Tse Chin, Bharadwaj Veeravalli, and Jingxi Jia. Handling large-size discrete wavelet transform on network-based computing systems — parallelization via divisible load paradigm. Journal of Parallel and Distributed Computing, 69(2):143–152, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chu:1992:OMN

[CW92] Lon-Chan Chu and Benjamin W. Wah. Optimal mapping of neural-network learning on message-passing multicomputers. Journal of Parallel and Distributed Computing, 14(3):319–339, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cheng:1993:SAD

[CW93] Kam Hoi Cheng and Qingzhou Wang. A simultaneous access design for idle processor reactivation and the detection of the termination of a parallel activity. Journal of Parallel and Distributed Computing, 17(4):370–373, April 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1036/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1036/production/pdf.

Cang:2000:TSO

[CW00] Songluan Cang and Jie Wu. Time-step optimal broadcasting in 3-D meshes with minimum total communication distance. Journal of Parallel and Distributed Computing, 60(8):966–997, August 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1642;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1642/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1642/ref.

Chang:2001:RPR

[CW01] Rocky K. C. Chang and Hong Y. Wang. Routing properties of a recursive interconnection network. Journal of Parallel and Distributed Computing, 61(6):838–849, June 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1729;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1729/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1729/ref.

Chen:2005:SSD

[CW05] Yu Chen and Jennifer L. Welch. Self-stabilizing dynamic mutual exclusion for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 65(9):1072–1089, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:2009:CFD

[CW09] Chien-Ping Chang and Chia-Ching Wu. Conditional fault diameter of crossed cubes. Journal of Parallel and Distributed Computing, 69(1):91–99, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chen:2007:PAM

[CWL^{+}07] Jianer Chen, GaoCai Wang, Chuang Lin, Tao Wang, and Guo-Jun Wang. Probabilistic analysis on mesh network fault tolerance. Journal of Parallel and Distributed Computing, 67(1): 100–110, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cai:2005:NSC

[CWLD05] Hailong Cai, Jun Wang, Dong Li, and Jitender S. Deogun. A novel state cache scheme in structured P2P systems. Journal of Parallel and Distributed Computing, 65(2):154–168, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chung:1998:PAP

[CWP98] Yongwha Chung, Cho-Li Wang, and Viktor K. Prasanna. Parallel algorithms for perceptual grouping on distributed memory machines. Journal of Parallel and Distributed Computing, 50(1–2):123–143, April/May 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1438/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1438/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1438/production/ref.

Cheng:1995:PBS

[CWW^{+}95] Fuhua Cheng, G. W. Wasilkowski, Jiaye Wang, Caiming Zhang, and Wenping Wang. Parallel B-spline surface interpolation on a mesh-connected processor array. Journal of Parallel and Distributed Computing, 24(2):224–229, February 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1022/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1022/production/pdf.

Chen:1996:TNL

[CWW96] Yeong-Sheng Chen, Sheng-De Wang, and Chien-Min Wang. Tiling nested loops into maximal rectangular blocks. Journal of Parallel and Distributed Computing, 35(2):123–132, June 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0075/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0075/production/
pdf.

Chung:2005:ORT

[CX05] Moon Jung Chung and Jinsheng Xu. An overhead reducing technique for Time Warp. Journal of Parallel and Distributed Computing, 65(1):65–73, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cheung:1995:FDA

[CY95] To-Yat Y. Cheung and Xinming M. Ye. A fault-detection approach to the conformance testing of nondeterministic systems. Journal of Parallel and Distributed Computing, 28(1): 94–100, July 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1091/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1091/
production/pdf.

Chung:1996:PBS

[CY96] Kuo-Liang Chung and Wen-Ming Yan. Parallel B-spline surface fitting on mesh-connected computers. Journal of Parallel and Distributed Computing, 35(2):205–210, June 15, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0082/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0082/production/pdf.

Chen:1999:HPT

[CY99] Hong-Mei Chen and David Y. Y. Yun. High-performance telemedicine information management. Journal of Parallel and Distributed Computing, 56(3):235–250, March 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1520/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1520/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1520/production/ref.

Cybenko:1989:DLB

[Cyb89] George Cybenko. Dynamic load balancing for distributed memory multiprocessors. Journal of Parallel and Distributed Computing, 7(2):279–301, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chlebikova:2006:ARG

[CYZ06] Janka Chlebíková, Deshi Ye, and Hu Zhang. Assign ranges in general ad-hoc networks. Journal of Parallel and Distributed Computing, 66(4):489–498, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cole:1990:OPA

[CZ90] Richard Cole and Ofer Zajicek. An optimal parallel algorithm for building a data structure for planar point location. Journal of Parallel and Distributed Computing, 8(3):280–285, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ci:2009:MFB

[CZZY09] Yi-Wei Ci, Zhan Zhang, De-Cheng Zuo, and Xiao-Zong Yang. Message fragment based causal message logging. Journal of Parallel and Distributed Computing, 69(11):915–921, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Darbha:1997:TDB

[DA97] Sekhar Darbha and Dharma P. Agrawal. A task duplication based scalable scheduling algorithm for distributed memory systems. Journal of Parallel and Distributed Computing, 46(1):15–27, October 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1376/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1376/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1376/production/ref.

Drummond:1996:DBD

[dADB96] Lúcia Maria de A. Drummond and Valmir C. Barbosa. Distributed breakpoint detection in message-passing programs. Journal of Parallel and Distributed Computing, 39(2):153– 167, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1996.0163/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.0163/production/pdf.

Dahlgren:1999:TIP

[Dah99] Fredrik Dahlgren. Techniques for improving performance of hybrid snooping cache protocols. Journal of Parallel and Distributed Computing, 59(3):329–359, December 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1558/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1558/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1558/production/ref.

Dhodhi:2002:ITT

[DAYA02] Muhammad K. Dhodhi, Imtiaz Ahmad, Anwar Yatama, and Ishfaq Ahmad. An integrated technique for task matching and scheduling onto distributed heterogeneous computing systems. Journal of Parallel and Distributed Computing, 62(9):1338–1361, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1850;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1850/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1850/ref.

Drezner:1986:AAS

[DB86] Zvi Drezner and Amnon Barak. An asynchronous algorithm for scattering information between the active nodes of a multicomputer system. Journal of Parallel and Distributed Computing, 3(3):344–351, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Derhab:2008:SSA

[DB08] Abdelouahid Derhab and Nadjib Badache. Self-stabilizing algorithm for high service availability in spite of concurrent topology changes in ad hoc mobile networks. Journal of Parallel and Distributed Computing, 68(6):752–768, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dail:2003:DSA

[DBC03] Holly Dail, Fran Berman, and Henri Casanova. A decoupled scheduling approach for Grid application development environments. Journal of Parallel and Distributed Computing, 63 (5):505–524, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dongarra:1990:TAD

[DBKF90] J. Dongarra, O. Brewer, J. A. Kohl, and S. Fineberg. A tool to aid in the design, implementation, and understanding of matrix algorithms for parallel processors. Journal of Parallel and Distributed Computing, 9(2):185–202, June 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

deAzevedo:1995:BAS

[dBL95] Marcelo Moraes de Azevedo, Nader Bagherzadeh, and Shahram Latifi. Broadcasting algorithms for the star-connected cycles interconnection network. Journal of Parallel and Distributed Computing, 25(2):209–222, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1041/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1041/production/pdf.

Donaldson:1994:PSH

[DBP94] V. Donaldson, F. Berman, and R. Paturi. Program speedup in a heterogeneous computing network. Journal of Parallel and Distributed Computing, 21(3):316–322, June 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1062/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1062/production/pdf.

Duh:1994:TPW

[DC94] Dyi-Rong R. Duh and Gen-Huey H. Chen. Topological properties of WK-Recursive networks. Journal of Parallel and Distributed Computing, 23(3):468–474, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1158/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1158/production/pdf.

DeBenedictis:1993:MS

[DD93] Erik P. DeBenedictis and Juan Miguel Del Rosario. Modular scalable I/O. Journal of Parallel and Distributed Computing, 17(1–2):122–128, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1010/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1010/production/pdf.

Driscoll:1995:APP

[DD95] Michael A. Driscoll and W. Robert Daasch. Accurate predictions of parallel program execution time. Journal of Parallel and Distributed Computing, 25(1):16–30, February 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1026/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1026/production/
pdf.

Das:1996:TPR

[DD96] Nabanita Das and Jayasree Dattagupta. Two-pass rearrangeability in faulty Benes networks. Journal of Parallel and Distributed Computing, 35(2):191–198, June 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0080/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0080/production/pdf.

Das:1998:HMP

[DDD98] Dibyendu Das, Pallab Dasgupta, and P. P. Das. A heuristic for the maximum processor requirement for scheduling layered task graphs with cloning. Journal of Parallel and Distributed Computing, 49(2):169–181, March 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1416/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1416/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1416/production/ref.

Das:2006:EAP

[DDNS06] Gautam K. Das, Sandip Das, Subhas C. Nandy, and Bhabani P. Sinha. Efficient algorithm for placing a given number of base stations to cover a convex region. Journal of Parallel and Distributed Computing, 66(11):1353–1358, November 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dandamudi:1991:HBH

[DE91] Sivarama P. Dandamudi and Derek L. Eager. On hypercubebased hierarchical interconnection network design. Journal of Parallel and Distributed Computing, 12(3):283–289, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

DeGroot:1988:TCE

[DeG88] D. DeGroot. A technique for compiling execution graph expressions for restricted AND-parallelism in logic programs. Journal of Parallel and Distributed Computing, 5(5):494–516, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dehne:1990:CLE

[Deh90] Frank Dehne. Computing the largest empty rectangle on oneand two-dimensional processor arrays. Journal of Parallel and Distributed Computing, 9(1):63–68, May 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dekel:2000:SIJ

[Dek00] Eliezer Dekel. Special issue on Java on clusters. Journal of Parallel and Distributed Computing, 60(10):1155–1158, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1648;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1648/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1648/ref.

Dubey:1990:OP

[DF90] Pradeep K. Dubey and Michael J. Flynn. Optimal pipelining. Journal of Parallel and Distributed Computing, 8(1):10– 19, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dubey:1994:BPM

[DF94] P. K. Dubey and M. J. Flynn. A bubble propagation model for pipeline performance. Journal of Parallel and Distributed Computing, 23(3):330–337, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1144/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1144/production/pdf.

Draper:1995:DAD

[DF95] Richard N. Draper and Vance Faber. The diameter and average distance of supertoroidal networks. Journal of Parallel and Distributed Computing, 31(1):1–13, November 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1140/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1140/production/pdf.

Dehne:1994:CDH

[DFN^{+}94] F. Dehne, A. Fabri, M. Nassar, A. Rau-Chaplin, and R. Valiveti. Construction of d-dimensional hyperoctrees on a hypercube multiprocessor. Journal of Parallel and Distributed Computing, 23(2):256–261, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1137/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1137/production/pdf.

Das:2006:CFS

[DFP06a] Sajal K. Das, Irene Finocchi, and Rossella Petreschi. Conflictfree star-access in parallel memory systems. Journal of Parallel and Distributed Computing, 66(11):1431–1441, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Du:2006:LLA

[DFP06b] Wenliang Du, Lei Fang, and Ning Peng. LAD: Localization anomaly detection for wireless sensor networks. Journal of Parallel and Distributed Computing, 66(7):874–886, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Diallo:1999:SCH

[DFRCU99] Mohamadou Diallo, Afonso Ferreira, Andrew Rau-Chaplin, and Stéphane Ubéda. Scalable 2D convex hull and triangulation algorithms for coarse grained multicomputers. Journal of Parallel and Distributed Computing, 56(1):47–70, January 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1503/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1503/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1503/production/ref.

deSouzaeSilva:1991:QNM

[dG91] E. de Souza e Silva and M. Gerla. Queueing network models for load balancing in distributed systems. Journal of Parallel and Distributed Computing, 12(1):24–38, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Draper:1994:CAM

[DG94] J. T. Draper and J. Ghosh. A comprehensive analytical model for wormhole routing in multicomputer systems. Journal of Parallel and Distributed Computing, 23(2):202–214, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1132/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1132/production/
pdf.

Delporte-Gallet:2005:MEA

[DGFGK05] Carole Delporte-Gallet, Hugues Fauconnier, Rachid Guerraoui, and Petr Kouznetsov. Mutual exclusion in asynchronous systems with failure detectors. Journal of Parallel and Distributed Computing, 65(4):492–505, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

delaOssa:2006:IAA

[dGP06] Luis delaOssa, José A. Gámez, and José M. Puerta. Initial approaches to the application of islands-based parallel EDAs in continuous domains. Journal of Parallel and Distributed Computing, 66(8):991–1001, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dehne:1991:PAD

[DH91a] Frank Dehne and Susanne E. Hambrusch. Parallel algorithms for determining k-width connectivity in binary images. Journal of Parallel and Distributed Computing, 12(1):12–23, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dutt:1991:DFT

[DH91b] Shantanu Dutt and John P. Hayes. Designing fault-tolerant systems using automorphisms. Journal of Parallel and Distributed Computing, 12(3):249–268, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dekel:1994:PAF

[DH94] Eliezer Dekel and Jie Hu. A parallel algorithm for finding minimum cutsets in reducible graphs. Journal of Parallel and Distributed Computing, 20(1):43–55, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1004/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1004/production/pdf.

Dowd:1995:MNA

[DH95] Patrick W. Dowd and I-Shyan S. Hwang. Memory and network architecture interaction in an optically interconnected distributed shared memory system. Journal of Parallel and Distributed Computing, 25(2):144–161, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1037/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1037/production/pdf.

Duarte:2004:VCD

[DH04] Marco F. Duarte and Yu Hen Hu. Vehicle classification in distributed sensor networks. Journal of Parallel and Distributed Computing, 64(7):826–838, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Das:2002:ALB

[DHB02] Sajal K. Das, Daniel J. Harvey, and Rupak Biswas. Adaptive load-balancing algorithms using symmetric broadcast networks. Journal of Parallel and Distributed Computing, 62(6):1042– 1068, June 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1819;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1819/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1819/ref.

Dingle:2004:UHP

[DHK04] Nicholas J. Dingle, Peter G. Harrison, and William J. Knottenbelt. Uniformization and hypergraph partitioning for the distributed computation of response time densities in very large Markov models. Journal of Parallel and Distributed Computing, 64(8):908–920, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dierstein:1996:ASI

[DHR96] Anne Dierstein, Roman Hayer, and Thomas Rauber. The AD-DAP system on the iPSC/860: Automatic data distribution and parallelization. Journal of Parallel and Distributed Computing, 32(1):1–10, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dvir:2006:VLC

[DHS06] Ori Dvir, Maurice Herlihy, and Nir N. Shavit. Virtual Leashing: Creating a computational foundation for software protection. Journal of Parallel and Distributed Computing, 66(9): 1233–1240, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dimpsey:1991:MMM

[DI91] R. T. Dimpsey and R. K. Iyer. Modeling and measuring multiprogramming and system overheads on a shared-memory multiprocessor: case study. Journal of Parallel and Distributed Computing, 12(4):402–414, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dimitrovsky:1991:GLA

[Dim91] Isaac A. Dimitrovsky. The group lock and its applications. Journal of Parallel and Distributed Computing, 11(4):291–302, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dimakopoulos:2004:APT

[Dim04] Vassilios V. Dimakopoulos. All-port total exchange in Cartesian product networks. Journal of Parallel and Distributed Computing, 64(8):936–944, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dominic:1998:CLS

[DJ98] M. Dominic and Bijendra N. Jain. Conditions for on-line scheduling of hard real-time tasks on multiprocessors. Journal of Parallel and Distributed Computing, 55(1):121–137, November 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1484/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1484/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1484/production/ref.

Djamegni:2004:MRM

[Dja04] Clémentin Tayou Djamegni. Mapping rectangular mesh algorithms onto asymptotically space-optimal arrays. Journal of Parallel and Distributed Computing, 64(3):345–359, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Djamegni:2006:CMP

[Dja06] Clémentin Tayou Djamegni. Complexity of matrix product on modular linear systolic arrays for algorithms with affine schedules. Journal of Parallel and Distributed Computing, 66(3):323– 333, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Deo:1994:PCT

[DJM94] N. Deo, A. Jain, and M. Medidi. Parallel construction of (a, b)-trees. Journal of Parallel and Distributed Computing, 23(3):442–448, December 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1154/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1154/production/pdf.

Durand:2003:PSU

[DJT03] Dannie Durand, Ravi Jain, and David Tseytlin. Parallel I/O scheduling using randomized, distributed edge coloring algorithms. Journal of Parallel and Distributed Computing, 63 (6):611–618, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ding:2004:IEB

[DK04] Chen Ding and Ken Kennedy. Improving effective bandwidth through compiler enhancement of global cache reuse. Journal of Parallel and Distributed Computing, 64(1):108–134, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Daoud:2008:HPA

[DK08] Mohammad I. Daoud and Nawwaf Kharma. A high performance algorithm for static task scheduling in heterogeneous distributed computing systems. Journal of Parallel and Distributed Computing, 68(4):399–409, April 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dornstetter:2001:SAS

[DKMV01] Jean-Louis Dornstetter, Daniel Krob, Michel Morvan, and Laurent Viennot. Some algorithms for synchronizing clocks of base transceiver stations in a cellular network. Journal of Parallel and Distributed Computing, 61(7):855–867, July 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1716;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1716/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1716/ref.

Ding:2009:AAE

[DKRI09] Yang Ding, Mahmut Kandemir, Padma Raghavan, and Mary Jane Irwin. Adapting application execution in CMPs using helper threads. Journal of Parallel and Distributed Computing, 69(9):790–806, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dolev:2001:SSG

[DKY01] Shlomi Dolev, Ephraim Korach, and Dmitry Yukelson. The sound of silence: Guessing games for saving energy in a mobile environment. Journal of Parallel and Distributed Computing, 61(7):868–883, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1717;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1717/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1717/ref.

Dessmark:1998:IBI

[DL98] Anders Dessmark and Andrzej Lingas. Improved bounds for integer sorting in the EREW PRAM model. Journal of Parallel and Distributed Computing, 48(1):64–70, January 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1401/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1401/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1401/production/ref.

DiStefano:1999:EKT

[DL99] A. Di Stefano and L. Lo Bello. Exploiting the knowledge of task structure for distributed allocation. Journal of Parallel and Distributed Computing, 59(1):54–67, October 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1563/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1563/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1563/production/ref.

Dong:2001:DCA

[DL01] Xuefeng Dong and Ten H. Lai. Dynamic carrier allocation strategies for mobile cellular networks. Journal of Parallel and Distributed Computing, 61(7):926–949, July 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1721;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1721/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1721/ref.

Deng:1997:CAN

[DLLX97] Xiaotie Deng, Hai-Ning Liu, JunSheng Long, and Bing Xiao. Competitive analysis of network load balancing. Journal of Parallel and Distributed Computing, 40(2):162–172, February 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1257/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1257/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1257/production/ref.

Diks:1999:OAB

[DLP99] Krzysztof Diks, Andrzej Lingas, and Andrzej Pelc. An optimal algorithm for broadcasting multiple messages in trees. Journal of Parallel and Distributed Computing, 59(3):465–474, December 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1584/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1584/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1584/production/ref.

Della-Libera:2000:RDT

[DLS00] Giovanni Della-Libera and Nir Shavit. Reactive diffracting trees. Journal of Parallel and Distributed Computing, 60 (7):853–890, July 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2000.1630;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2000.
1630/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2000.1630/ref.

Duprat:1988:HPE

[DM88] Jean Duprat and Jean-Michel Muller. Hardwired polynomial evaluation. Journal of Parallel and Distributed Computing, 5 (3):291–309, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Darling:1990:PAS

[DM90a] J. P. Darling and I. D. Mayergoyz. Parallel algorithm for the solution of nonlinear Poisson equation of semiconductor device theory and its implementation on the MPP. Journal of Parallel and Distributed Computing, 8(2):161–168, February 1990.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dinning:1990:FPA

[DM90b] Anne Dinning and B. Mishra. A fully parallel algorithm for implementing path expressions. Journal of Parallel and Distributed Computing, 10(3):205–221, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dixit:1990:APP

[DM90c] V. V. Dixit and D. I. Moldovan. The allocation problem in parallel production systems. Journal of Parallel and Distributed Computing, 8(1):20–29, January 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Davarakis:1992:PPA

[DM92] C. T. Davarakis and D. G. Maritsas. A probabilistic parallel associative search and query set of algorithms. Journal of Parallel and Distributed Computing, 14(1):37–49, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dutt:1994:SLB

[DM94] S. Dutt and N. R. Mahapatra. Scalable load balancing strategies for parallel A* algorithms. Journal of Parallel and Distributed Computing, 22(3):488–505, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1106/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1106/production/pdf.

Das:1995:UAP

[DM95] Sajal K. Das and Kwang Bae Min. A unified approach to the parallel construction of search trees. Journal of Parallel and Distributed Computing, 27(1):71–78, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1072/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1072/production/pdf.

Dramitinos:1999:ARP

[DM99] George Dramitinos and Evangelos P. Markatos. Adaptive and reliable paging to remote main memory. Journal of Parallel and Distributed Computing, 58(3):357–388, September 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1546/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1546/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1546/production/ref.

Desai:2004:SHL

[DM04] Nirmit Desai and Frank Mueller. Scalable hierarchical locking for distributed systems. Journal of Parallel and Distributed Computing, 64(6):708–724, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dasgupta:1997:VAL

[DMB97] Pallab Dasgupta, A. K. Majumder, and P. Bhattacharya. V THR: An adaptive load balancing algorithm. Journal of Parallel and Distributed Computing, 42(2):101–108, May 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1321/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1321/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1321/production/ref.

Das:2003:AFQ

[DMB^{+}03] S. K. Das, A. Mukherjee, S. Bandyopadhyay, D. Saha, and K. Paul. An adaptive framework for QoS routing through multiple paths in ad hoc wireless networks. Journal of Parallel and Distributed Computing, 63(2):141–153, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Darte:2003:GMM

[DMCFCM03] Alain Darte, John Mellor-Crummey, Robert Fowler, and Daniel Chavarría-Miranda. Generalized multipartitioning of multidimensional arrays for parallelizing line-sweep computations. Journal of Parallel and Distributed Computing, 63(9):887–911, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Davis:1990:DDE

[DMSH90] Nathaniel J. Davis, IV, David L. Mannix, Wade H. Shaw, and Thomas C. Hartrum. Distributed discrete-event simulation using null message algorithms on hypercube architectures. Journal of Parallel and Distributed Computing, 8(4):349–357, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Demeure:1994:PSP

[DN94] I. M. Demeure and G. J. Nutt. Prototyping and simulating parallel, distributed computations. Journal of Parallel and Distributed Computing, 23(1):1–9, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1114/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1114/production/pdf.

Dekel:1989:PEM

[DO89] Eliezer Dekel and Istvan Ozsvath. Parallel external merging. Journal of Parallel and Distributed Computing, 6(3):623–635, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dogan:2006:SMT

[DÖ06] Atakan Doḡan and Füsun Özgüner. Scheduling of a meta-task with QoS requirements in heterogeneous computing systems. Journal of Parallel and Distributed Computing, 66(2):181–196, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dolev:1997:SSR

[Dol97] Shlomi Dolev. Self-stabilizing routing and related protocols. Journal of Parallel and Distributed Computing, 42(2):122–127, May 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1317/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1317/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1317/production/ref.

Das:1998:SIP

[DOP98] Sajal K. Das, Stephan Olariu, and Sushil K. Prasad. Special issue on parallel and distributed data structures: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 49(1):1–3, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Das:1998:TAH

[DP98] Sajal K. Das and M. Cristina Pinotti. O(log log N) time algorithms for Hamiltonian suffix and min-max-pair heap operations on the hypercube. Journal of Parallel and Distributed Computing, 48(2):200–211, February 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1418/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1418/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1418/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1452/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1452/
production/pdf.

Dasgupta:1999:PGA

[DP99] Bhaskar Dasgupta and Michael A. Palis. Provably good algorithms for transmission scheduling in WDM optical networks. Journal of Parallel and Distributed Computing, 57(3):345– 357, June 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1542/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1542/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1542/production/ref.

Das:2000:OMA

[DP00] Sajal K. Das and M. Cristina Pinotti. Optimal mappings of q-ary and binomial trees into parallel memory modules for fast and conflict-free access to path and subtree templates. Journal of Parallel and Distributed Computing, 60(8):998–1027, August 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1638;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1638/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1638/ref.

Despain:1985:FFT

[DPRW85] Alvin M. Despain, Allen M. Peterson, Oscar S. Rothaus, and Erling H. Wold. Fast Fourier transform processors using Gaussian residue arithmetic. Journal of Parallel and Distributed Computing, 2(3):219–237, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dalcin:2005:MP

[DPS05] Lisandro Dalcín, Rodrigo Paz, and Mario Storti. MPI for Python. Journal of Parallel and Distributed Computing, 65(9): 1108–1115, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dalcin:2008:MPP

[DPSD08] Lisandro Dalcín, Rodrigo Paz, Mario Storti, and Jorge D’Elía. MPI for Python: Performance improvements and MPI-2 extensions. Journal of Parallel and Distributed Computing, 68 (5):655–662, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Djamegni:2009:RBA

[DQR^{+}09] Clémentin Tayou Djamegni, Patrice Quinton, Sanjay Rajopadhye, Tanguy Risset, and Maurice Tchuenté. A reindexing based approach towards mapping of DAG with affine schedules onto parallel embedded systems. Journal of Parallel and Distributed Computing, 69(1):1–11, January 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Darte:1995:ASS

[DR95] Alain Darte and Yves Robert. Affine-by-statement scheduling of uniform and affine loop nests over parametric domains. Journal of Parallel and Distributed Computing, 29(1):43–59, August 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1105/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1105/
production/pdf.

Diniz:1998:LCE

[DR98] Pedro C. Diniz and Martin C. Rinard. Lock coarsening: Eliminating lock overhead in automatically parallelized object-based programs. Journal of Parallel and Distributed Computing, 49(2):218–244, March 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1441/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1441/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1441/production/ref.

dAuriol:2009:OPB

[dR09] Brian J. d’Auriol and Juan René Roldán. An optical power budget model for the parameterized linear array with a reconfigurable pipelined bus system (LARPBS(p)) model. Journal of Parallel and Distributed Computing, 69(10):815–823, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dehne:1990:IDS

[DRC90] Frank Dehne and Andrew Rau-Chaplin. Implementing data structures on a hypercube multiprocessor, and applications in parallel computational geometry. Journal of Parallel and Distributed Computing, 8(4):367–375, April 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dion:1996:CAN

[DRR96] Michèle Dion, Cyril Randriamaro, and Yves Robert. Compiling affine nested loops: How to optimize the residual communications after the alignment phase. Journal of Parallel and Distributed Computing, 38(2):176–187, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0139/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0139/production/pdf.

Duato:2001:CRA

[DRSB01] J. Duato, A. Robles, F. Silla, and R. Beivide. A comparison of router architectures for virtual cut-through and wormhole switching in a NOW environment. Journal of Parallel and Distributed Computing, 61(2):224–253, February 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1679;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1679/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1679/ref.

Diaz:2002:BBC

[DRST02] Manuel Díaz, Bartolomé Rubio, Enrique Soler, and José M. Troya. A border-based coordination language for integrating task and data parallelism. Journal of Parallel and Distributed Computing, 62(4):715–740, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1814;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1814/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1814/ref.

Diaz:2007:TCB

[DRT07] Manuel Díaz, Bartolomé Rubio, and José M. Troya. A tuple channel-based coordination model for parallel and distributed programming. Journal of Parallel and Distributed Computing, 67(10):1092–1107, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dekel:1984:PMA

[DS84] Eliezer Dekel and Sartaj Sahni. Parallel matching algorithm for convex bipartite graphs and applications to scheduling. Journal of Parallel and Distributed Computing, 1(2):185–205, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dingle:1993:SBT

[DS93] Adair Dingle and I. Hal Sudborough. Simulation of binary trees and X-trees on pyramid networks. Journal of Parallel and Distributed Computing, 19(2):119–124, October 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1096/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1096/production/pdf.

Dahlgren:1995:UWC

[DS95a] Fredrik Dahlgren and Per Stenstrom. Using write caches to improve performance of cache coherence protocols in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 26(2):193–210, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1059/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1059/production/pdf.

Das:1995:OCA

[DS95b] Rajib K. Das and Bhabani P. Sinha. Optimal communication algorithms in distributed loop networks. Journal of Parallel and Distributed Computing, 30(1):85–90, October 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1128/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1128/production/pdf.

Das:1996:NTO

[DS96] Rajib K. Das and Bhabani P. Sinha. A new topology with odd degree for multiprocessor systems. Journal of Parallel and Distributed Computing, 39(1):87–94, November 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0157/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0157/production/pdf.

Datta:2001:CTA

[DS01] Amitava Datta and Subbiah Soundaralakshmi. Constanttime algorithm for the Euclidean distance transform on reconfigurable meshes. Journal of Parallel and Distributed Computing, 61(10):1439–1455, October 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1684;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1684/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1684/ref.

Deelman:2002:SSE

[DS02] Ewa Deelman and Boleslaw K. Szymanski. Simulating spatially explicit problems on high performance architectures. Journal of Parallel and Distributed Computing, 62(3):446–467, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1802;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1802/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1802/ref.

Datta:2004:FSA

[DS04a] Amitava Datta and Subbiah Soundaralakshmi. Fast and scalable algorithms for the Euclidean distance transform on a linear array with a reconfigurable pipelined bus system. Journal of Parallel and Distributed Computing, 64(3):360–369, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Defago:2004:SPR

[DS04b] Xavier Défago and André Schiper. Semi-passive replication and Lazy Consensus. Journal of Parallel and Distributed Computing, 64(12):1380–1398, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dhodhi:1999:DID

[DSAUM99] Muhammad K. Dhodhi, John A. Saghri, Imtiaz Ahmad, and Raza Ul-Mustafa. D-ISODATA: a distributed algorithm for unsupervised classification of remotely sensed data on network of workstations. Journal of Parallel and Distributed Computing, 59(2):280–301, November 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1573/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1573/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1573/production/ref.

Dines:1997:OIS

[DSD^{+}97] J. A. B. Dines, J. F. Snowdon, M. P. Y. Desmulliez, D. B. Barsky, A. V. Shafarenko, and C. R. Jesshope. Optical interconnectivity in a scalable data-parallel system. Journal of Parallel and Distributed Computing, 41(1):120–130, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1290/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1290/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1290/production/ref.

deSouza:2000:DSV

[dSR00] A. F. de Souza and P. Rounce. Dynamically scheduling VLIW instructions. Journal of Parallel and Distributed Computing, 60(12):1480–1511, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1661;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1661/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1661/ref.

Dubois:1995:EMD

[DSS95] Michel Dubois, Jonas Skeppstedt, and Per Strenstrom. Essential misses and data traffic in coherence protocols. Journal of Parallel and Distributed Computing, 29(2):108–125, September 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1112/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1112/production/
pdf.

Dan:1995:CAU

[DSST95] Asit Dan, Perwez Shahabuddin, Dinkar Sitaram, and Don Towsley. Channel allocation under batching and VCR control in video-on-demand systems. Journal of Parallel and Distributed Computing, 30(2):168–179, November 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1135/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1135/production/pdf.

Dewan:1994:SPD

[DSW94] H. M. Dewan, S. J. Stolfo, and L. Woodbury. Scalable parallel and distributed expert database systems with predictive load balancing. Journal of Parallel and Distributed Computing, 22(3):506–522, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1107/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1107/production/pdf.

Dubois:1992:SIM

[DT92] Michel Dubois and Shreekant Thakkar. Special issue on memory system architectures for scalable multiprocessors. Journal of Parallel and Distributed Computing, 15(4):303–304, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Djamegni:1997:COP

[DT97] Clémentin Tayou Djamegni and Maurice Tchuenté. A costoptimal pipeline algorithm for permutation generation in lexicographic order. Journal of Parallel and Distributed Computing, 44(2):153–159, August 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1347/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1347/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1347/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1369/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1369/
production/pdf.

Dickens:2001:ECI

[DT01] Phillip M. Dickens and Rajeev Thakur. Evaluation of collective I/O implementations on parallel architectures. Journal of Parallel and Distributed Computing, 61(8):1052–1076, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1733;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1733/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1733/ref.

Delaet:2002:TTI

[DT02] Sylvie Delaët and Sébastien Tixeuil. Tolerating transient and intermittent failures. Journal of Parallel and Distributed Computing, 62(5):961–981, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1827;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1827/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1827/ref.

Das:1994:COI

[DUSH94] R. Das, M. Uysal, J. Saltz, and Yuan-Shin S. Hwang. Communication optimizations for irregular scientific computations on distributed memory architectures. Journal of Parallel and Distributed Computing, 22(3):462–478, September 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1104/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1104/production/pdf.

Dolev:1986:PCS

[DUW86] D. Dolev, E. Upfal, and M. K. Warmuth. The parallel complexity of scheduling with precedence constraints. Journal of Parallel and Distributed Computing, 3(4):553–576, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

deVega:2006:SIP

[dVCP06] Francisco Fernández de Vega and Erick Cantú-Paz. Special issue on parallel bioinspired algorithms. Journal of Parallel and Distributed Computing, 66(8):989–990, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dongarra:1994:SIA

[DVW94] J. J. Dongarra, R. A. Van de Geijn, and D. W. Walker. Scalability issues affecting the design of a dense linear algebra library. Journal of Parallel and Distributed Computing, 22(3):523–537, September 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1108/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1108/production/pdf.

Dighe:1996:BCR

[DVZ96] Omkar M. Dighe, Ramachandran Vaidyanathan, and S. Q. Zheng. The bus-connected ringed tree: a versatile interconnection network. Journal of Parallel and Distributed Computing, 33(2):189–196, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0037/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0037/production/pdf.

Drozdowski:2004:PLD

[DW04] Maciej Drozdowski and Paweł Wolniewicz. Performance limits of divisible load processing in systems with limited communication buffers. Journal of Parallel and Distributed Computing, 64 (8):960–973, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dai:2006:CCD

[DW06] Fei Dai and Jie Wu. On constructing k-connected k-dominating set in wireless ad hoc and sensor networks. Journal of Parallel and Distributed Computing, 66(7):947–958, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Damani:2003:DRO

[DWG03] Om P. Damani, Yi-Min Wang, and Vijay K. Garg. Distributed recovery with K-optimistic logging. Journal of Parallel and Distributed Computing, 63(12):1193–1218, December 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Duval:1987:FTT

[DWHL87] Louis D. Duval, Robert A. Wagner, Yijie Han, and Donald W. Loveland. Finding test-and-treatment procedures using parallel computation. Journal of Parallel and Distributed Computing, 4(3):309–318, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dandamudi:1999:PAS

[DY99] Sivarama P. Dandamudi and Hai Yu. Performance of adaptive space sharing processor allocation policies for distributedmemory multicomputers. Journal of Parallel and Distributed Computing, 58(1):109–125, July 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1544/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1544/production/pdf.

Du:1997:CPP

[DZ97] Xing Du and Xiaodong Zhang. Coordinating parallel processes on networks of workstations. Journal of Parallel and Distributed Computing, 46(2):125–135, November 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1378/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1378/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1378/production/ref.

Du:2001:AES

[DZDZ01] Xing Du, Xiaodong Zhang, Yingfei Dong, and Lin Zhang. Architectural effects of symmetric multiprocessors on TPC-C commercial workload. Journal of Parallel and Distributed Computing, 61(5):609–640, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1710;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2000.1710/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1710/ref.

El-Amawy:1990:BHN

[EAL90] A. El-Amawy and S. Latifi. Bridged hypercube networks. Journal of Parallel and Distributed Computing, 10(1):90–95, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

El-Boghdadi:2009:PAR

[EB09] Hatem M. El-Boghdadi. Power-aware routing for well-nested communications on the circuit switched tree. Journal of Parallel and Distributed Computing, 69(2):135–142, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Eberbach:1994:CDG

[Ebe94] E. Eberbach. CSA: In the direction of greater representational power for neurocomputing. Journal of Parallel and Distributed Computing, 22(1):107–112, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1073/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1073/production/pdf.

Elhadef:2008:DFI

[EBE08] Mourad Elhadef, Azzedine Boukerche, and Hisham Elkadiki. A distributed fault identification protocol for wireless and mobile ad hoc networks. Journal of Parallel and Distributed Computing, 68(3):321–335, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ebner:2004:PAC

[Ebn04] Marc Ebner. A parallel algorithm for color constancy. Journal of Parallel and Distributed Computing, 64(1):79–88, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Engstrom:1989:SPS

[EC89] Bradley R. Engstrom and Peter R. Cappello. The SDEF programming system. Journal of Parallel and Distributed Computing, 7(2):201–231, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Edelman:1991:OMT

[Ede91] Alan Edelman. Optimal matrix transposition and bit reversal on hypercubes: all-to-all personalized communication. Journal of Parallel and Distributed Computing, 11(4):328–331, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Eltayeb:2005:PSB

[EDÖ05] Mohammed Eltayeb, Atakan Doḡan, and Füsun Özgüner. A path selection-based algorithm for real-time data staging in Grid applications. Journal of Parallel and Distributed Computing, 65(11):1318–1328, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

ElBaz:2005:LBM

[EE05] D. El Baz and M. Elkihel. Load balancing methods and parallel dynamic programming algorithm using dominance technique applied to the 0–1 knapsack problem. Journal of Parallel and Distributed Computing, 65(1):74–84, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Efe:1991:EMT

[Efe91] Kemal Efe. Embedding mesh of trees in the hypercube. Journal of Parallel and Distributed Computing, 11(3):222–230, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Efe:1996:ELC

[Efe96] Kemal Efe. Embedding large complete binary trees in hypercubes with load balancing. Journal of Parallel and Distributed Computing, 35(1):104–109, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0073/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0073/production/pdf.

Eom:2001:AEA

[EH01a] Hyeonsang Eom and Jeffrey K. Hollingsworth. Achieving efficiency and accuracy in simulation for I/O-intensive applications. Journal of Parallel and Distributed Computing, 61(12):1732–1750, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1769;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1769/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1769/ref.

Eshaghian:2001:OIR

[EH01b] Mary M. Eshaghian and Lili Hai. An optically interconnected reconfigurable mesh. Journal of Parallel and Distributed Computing, 61(6):737–747, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1704;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1704/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1704/ref.

Evett:1995:PMP

[EHMN95] Matthew Evett, James Hendler, Ambuj Mahanti, and Dana Nau. PRA*: Massively parallel heuristic search. Journal of Parallel and Distributed Computing, 25(2):133–143, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1036/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1036/production/
pdf.

Evett:1994:PKR

[EHS94] M. P. Evett, J. A. Hendler, and L. Spector. Parallel knowledge representation on the Connection Machine. Journal of Parallel and Distributed Computing, 22(2):168–184, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1079/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1079/production/pdf.

Es:2007:ARG

[Eİ07] Alphan Es and Veysi İşler. Accelerated regular grid traversals using extended anisotropic chessboard distance fields on a parallel stream processor. Journal of Parallel and Distributed Computing, 67(11):1201–1217, November 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ercegovac:1988:LSC

[EL88] Milos D. Ercegovac and Tomas Lang. On-line scheme for computing rotation factors. Journal of Parallel and Distributed Computing, 5(3):209–227, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ercegovac:1991:MPM

[EL91] Milos D. Ercegovac and Tomas Lang. Module to perform multiplication, division, and square root in systolic arrays for matrix computations. Journal of Parallel and Distributed Computing, 11(3):212–221, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ercal:1997:TEM

[EL97] F. Ercal and H. C. Lee. Time-efficient maze routing algorithms on reconfigurable mesh architectures. Journal of Parallel and Distributed Computing, 44(2):133–140, August 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1353/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1353/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1353/production/ref.

Eshaghian:1994:OTP

[ELS94] M. M. Eshaghian, S. H. Lee, and M. E. Shaaban. Optical techniques for parallel image computing. Journal of Parallel and Distributed Computing, 23(2):190–201, November 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1131/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1131/production/pdf.

Evans:1989:FTS

[EM89] D. J. Evans and G. M. Megson. Fast triangularization of a symmetric tridiagonal matrix. Journal of Parallel and Distributed Computing, 6(3):663–678, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ellis:1994:SPA

[EMMM94] J. A. Ellis, M. Mata-Montero, and H. Müller. Serial and parallel algorithms for (k, 2)-partite graphs. Journal of Parallel and Distributed Computing, 22(2):129–137, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1076/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1076/production/pdf.

Eriksson:1996:ORD

[EMP^{+}96] Christer Eriksson, Jukka Maki-Turja, Kjell Post, Mikael Gustafsson, Jan Gustafsson, Kristian Sandstrom, and Ellus Brorsson. An overview of RealTimeTalk, a design framework for real-time systems. Journal of Parallel and Distributed Computing, 36(1):66–80, July 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0090/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0090/production/pdf.

Eberlein:1990:EIJ

[EP90] P. J. Eberlein and Haesun Park. Efficient implementation of Jacobi algorithms and Jacobi sets on distributed memory architectures. Journal of Parallel and Distributed Computing, 8 (4):358–366, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Elford:1997:TTD

[ER97] Chris L. Elford and Daniel A. Reed. Technology trends and disk array performance. Journal of Parallel and Distributed Computing, 46(2):136–147, November 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1405/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1405/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1405/production/ref.

El-Rewini:1995:SSC

[ERA95] H. El-Rewini and H. H. Ali. Static scheduling of conditional branches in parallel programs. Journal of Parallel and Distributed Computing, 24(1):41–54, January 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1005/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1005/production/pdf.

Eriksen:1988:TDP

[Eri88] Ole Eriksen. A termination detection protocol and its formal verification. Journal of Parallel and Distributed Computing, 5 (1):82–91, February 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

El-Rewini:1990:SPP

[ERL90] H. El-Rewini and T. G. Lewis. Scheduling parallel program tasks onto arbitrary target machines. Journal of Parallel and Distributed Computing, 9(2):138–153, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ercal:1990:TAH

[ERS90] F. Ercal, J. Ramanujam, and P. Sadayappan. Task allocation onto a hypercube by recursive mincut bipartitioning. Journal of Parallel and Distributed Computing, 10(1):35–44, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Eisenhauer:1996:DAP

[ES96] Greg Eisenhauer and Karsten Schwan. Design and analysis of a parallel molecular dynamics application. Journal of Parallel and Distributed Computing, 35(1):76–90, May 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0070/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0070/production/pdf.

Eshaghian:1997:SIP

[ES97] Mary M. Eshaghian and Eugen Schenfeld. Special issue on parallel computing with optical interconnects: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 41(1):1–2, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1280/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1280/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1280/production/ref.

Ensink:2003:PCL

[ESA03] Brian Ensink, Joel Stanley, and Vikram Adve. Program Control Language: a programming language for adaptive distributed applications. Journal of Parallel and Distributed Computing, 63(11):1082–1104, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

ElBaz:1996:AIA

[ESMG96] Didier El Baz, Pierre Spiteri, Jean Claude Miellou, and Didier Gazen. Asynchronous iterative algorithms with flexible communication for nonlinear network flow problems. Journal of Parallel and Distributed Computing, 38(1):1–15, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0124/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0124/production/
pdf.

El-Sharkawy:1994:SDP

[ESTA94] M. El-Sharkawy, Wenlong Tsang, and M. Aburdene. Short data parallel vector Slant transform. Journal of Parallel and Distributed Computing, 23(1):27–36, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1116/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1116/production/pdf.

elAyeb:1995:ABD

[eW95] Béchir el Ayeb and Shengrui Wang. Abduction-based diagnosis: a competition-based neural model simulating abductive reasoning. Journal of Parallel and Distributed Computing, 24(2):202–212, February 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1019/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1019/production/pdf.

Fragopoulou:1995:OCA

[FA95] P. Fragopoulou and S. G. Akl. Optimal communication algorithms on star graphs using spanning tree constructions. Journal of Parallel and Distributed Computing, 24(1):55–71, January 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1006/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1006/
production/pdf.

Fallah:2007:HPC

[FA07] Yaser Pourmohammadi Fallah and Hussein Alnuweiri. Hybrid polling and contention access scheduling in IEEE 802.11e WLANs. Journal of Parallel and Distributed Computing, 67(2): 242–256, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fagin:1992:LIM

[Fag92] Barry S. Fagin. Large integer multiplication on hypercubes. Journal of Parallel and Distributed Computing, 14(4):426–430, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fan:1995:CSM

[FAGW95] Xianzhi Fan, Nael B. Abu-Ghazaleh, and Philip A. Wilsey. On the complexity of scheduling MIMD operations for SIMD interpretation. Journal of Parallel and Distributed Computing, 29(1):91–95, August 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1109/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1109/production/pdf.

Fahringer:1996:CTE

[Fah96] Thomas Fahringer. Compile-time estimation of communication costs for data parallel programs. Journal of Parallel and Distributed Computing, 39(1):46–65, November 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0153/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0153/production/pdf.

Fragopoulou:1996:OCP

[FAM96] Paraskevi Fragopoulou, Selim G. Akl, and Henk Meijer. Optimal communication primitives on the generalized hypercube network. Journal of Parallel and Distributed Computing, 32(2):173–187, February 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0012/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0012/production/pdf.

Fukuda:1999:MVM

[FBDC99] Munehiro Fukuda, Lubomir F. Bic, Michael B. Dillencourt, and Jason M. Cahill. Messages versus messengers in distributed programming. Journal of Parallel and Distributed Computing, 57(2):188–211, May 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1533/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1533/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1533/production/ref.

Fink:1998:ERT

[FBK98] Stephen J. Fink, Scott B. Baden, and Scott R. Kohn. Efficient run-time support for irregular block-structured applications. Journal of Parallel and Distributed Computing, 50(1–2):61–82, April/May 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1437/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1437/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1437/production/ref.

Feng:2003:PAB

[FBRW03] Xizhou Feng, Duncan A. Buell, John R. Rose, and Peter J. Waddell. Parallel algorithms for Bayesian phylogenetic inference. Journal of Parallel and Distributed Computing, 63(7–8): 707–718, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fan:1990:DAS

[FC90] Z. Fan and K. H. Cheng. Design and analysis of simultaneous access priority queues. Journal of Parallel and Distributed Computing, 9(4):387–397, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Foster:1995:FML

[FC95] Ian T. Foster and K. Mani Chandy. FORTRAN M: a language for modular parallel programming. Journal of Parallel and Distributed Computing, 26(1):24–35, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1044/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1044/production/pdf.

Feng:2007:DPI

[FCC07] Yuhong Feng, Wentong Cai, and Jiannong Cao. Dynamic partner identification in mobile agent-based distributed job workflow execution. Journal of Parallel and Distributed Computing, 67(11):1137–1154, November 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fang:2000:FRH

[FCF00] Yugang Fang, Imrich Chlamtac, and Hong-Bing Fei. Failure recovery of HLR mobility databases and parameter optimization for PCS networks. Journal of Parallel and Distributed Computing, 60(4):431–450, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1615;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1615/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1615/ref.

Fernandez:2004:TDR

[FCG04] José-Jesús Fernández, José-María Carazo, and Inmaculada García. Three-dimensional reconstruction of cellular structures by electron microscope tomography and parallel computing. Journal of Parallel and Distributed Computing, 64(2):285–300, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feo:1990:RSL

[FCO90] J. T. Feo, D. C. Cann, and R. R. Oldehoeft. A report on the Sisal language project. Journal of Parallel and Distributed Computing, 10(4):349–366, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fineberg:1991:EAM

[FCS91] Samuel A. Fineberg, Thomas L. Casavant, and Howard Jay Siegel. Experimental analysis of a mixed-mode parallel architecture using bitonic sequence sorting. Journal of Parallel and Distributed Computing, 11(3):239–251, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Franklin:1986:INP

[FD86] Mark A. Franklin and Sanjay Dhar. Interconnection networks: physical design and performance analysis. Journal of Parallel and Distributed Computing, 3(3):352–372, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fei:2003:NCA

[Fei03] Zongming Fei. A new consistency algorithm for dynamic documents in content distribution networks. Journal of Parallel and Distributed Computing, 63(10):916–926, October 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feng:1990:SAB

[Fen90] Tse-Yun Feng. Search algorithms for bis-sequential machines. Journal of Parallel and Distributed Computing, 8(1):1–9, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferrante:1990:PDS

[Fer90] Richard D. Ferrante. Parameters for distributed system design. Journal of Parallel and Distributed Computing, 9(3):323–329, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferscha:1992:PNA

[Fer92] A. Ferscha. A Petri net approach for performance oriented parallel program design. Journal of Parallel and Distributed Computing, 15(3):188–206, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferretti:1993:GHT

[Fer93] Marco Ferretti. The generalized Hough transform on meshconnected computers. Journal of Parallel and Distributed Computing, 19(1):51–57, September 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1089/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1089/production/pdf.

Ferragina:1995:TSP

[Fer95] Paolo Ferragina. A technique to speed up parallel fully dynamic algorithms for MST. Journal of Parallel and Distributed Computing, 31(2):181–189, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1157/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1157/production/pdf.

Fleury:1998:SPB

[FF98] Eric Fleury and Pierre Fraigniaud. Strategies for path-based multicasting in wormhole-routed meshes. Journal of Parallel and Distributed Computing, 53(1):26–62, August 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1473/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1473/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1473/production/ref.

Feldman:1997:PSD

[FFK97] Phillip M. Feldman, Raisa E. Feldman, and David B. Kim. Predicting speedup for distributed computing on a token ring network. Journal of Parallel and Distributed Computing, 45(1):53–62, August 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1365/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1365/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1365/production/ref.

Fernandez:2008:EPI

[FGG08] José-Jesús Fernández, Dan Gordon, and Rachel Gordon. Efficient parallel implementation of iterative reconstruction algorithms for electron tomography. Journal of Parallel and Distributed Computing, 68(5):626–640, May 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Foster:1997:MMC

[FGKT97] Ian Foster, Jonathan Geisler, Carl Kesselman, and Steven Tuecke. Managing multiple communication methods in highperformance networked computing systems. Journal of Parallel and Distributed Computing, 40(1):35–48, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1266/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1266/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1266/production/ref.

Fitch:2003:BMA

[FGM^{+}03] B. G. Fitch, R. S. Germain, M. Mendell, J. Pitera, M. Pitman, A. Rayshubskiy, Y. Sham, F. Suits, W. Swope, T. J. C. Ward, et al. Blue Matter, an application framework for molecular simulation on Blue Gene. Journal of Parallel and Distributed Computing, 63(7–8):759–773, July/August 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Foglia:2005:RCO

[FGP05] Pierfrancesco Foglia, Roberto Giorgi, and Cosimo Antonio Prete. Reducing coherence overhead and boosting performance of high-end SMP multiprocessors running a DSS workload. Journal of Parallel and Distributed Computing, 65(3):289–306, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferrari:2003:GAA

[FGZ03] Carlo Ferrari, Concettina Guerra, and Giuseppe Zanotti. A grid-aware approach to protein structure comparison. Journal of Parallel and Distributed Computing, 63(7–8):728–737, July/ August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fujiwara:2004:PER

[FII04] Toshiyuki Fujiwara, Kazuo Iwama, and Chuzo Iwamoto. Partially effective randomization in simulations between ARBI-TRARY and COMMON PRAMs. Journal of Parallel and Distributed Computing, 64(3):319–326, March 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Francioni:1993:BSA

[FJ93] Joan M. Francioni and Jay Alan Jackson. Breaking the silence: Auralization of parallel program behavior. Journal of Parallel and Distributed Computing, 18(2):181–194, June 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1055/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1055/production/pdf.

Fernandez:2004:ICM

[FJC04] Antonio Fernández, Ernesto Jiménez, and V. Vicent Cholvi. On the interconnection of causal memory systems. Journal of Parallel and Distributed Computing, 64(4):498–506, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferreira:2006:LSP

[FJG06] Ronaldo A. Ferreira, Suresh Jagannathan, and Ananth Grama. Locality in structured peer-to-peer networks. Journal of Parallel and Distributed Computing, 66(2):257–273, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ford:1990:PCT

[FJSW90] Ray Ford, Michael J. Jipping, Roger Shultz, and Brion Wenhardt. On the performance of concurrent tree algorithms. Journal of Parallel and Distributed Computing, 8(3):253–266, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fox:1989:CGG

[FK89] G. C. Fox and J. G. Koller. Code generation by a generalized neural network: general principles and elementary examples. Journal of Parallel and Distributed Computing, 6(2):388–410, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ferreira:2008:SIS

[FKJG08] Ronaldo A. Ferreira, Mehmet Koyutürk, Suresh Jagannathan, and Ananth Grama. Semantic indexing in structured peer-to-peer networks. Journal of Parallel and Distributed Computing, 68(1):64–77, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flocchini:2004:SEA

[FKK^{+}04] Paola Flocchini, Evangelos Kranakis, Danny Krizanc, Flaminia L. Luccio, and Nicola Santoro. Sorting and election in anonymous asynchronous rings. Journal of Parallel and Distributed Computing, 64(2):254–265, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Foster:1997:LBA

[FKKC97] Ian Foster, David R. Kohr, Jr., Rakesh Krishnaiyer, and Alok Choudhary. A library-based approach to task parallelism in a data-parallel language. Journal of Parallel and Distributed Computing, 45(2):148–158, September 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1367/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1367/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1367/production/ref.

Freeh:2008:JTD

[FKLB08] Vincent W. Freeh, Nandini Kappiah, David K. Lowenthal, and Tyler K. Bletsch. Just-in-time dynamic voltage scaling: Exploiting inter-node slack to save energy in MPI programs. Journal of Parallel and Distributed Computing, 68(9):1175–1185, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fizzano:1997:DJS

[FKSW97] Perry Fizzano, David Karger, Clifford Stein, and Joel Wein. Distributed job scheduling in rings. Journal of Parallel and Distributed Computing, 45(2):122–133, September 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1373/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1373/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1373/production/ref.

Foster:1996:NAI

[FKT96] Ian Foster, Carl Kesselman, and Steven Tuecke. The Nexus approach to integrating multithreading and communication. Journal of Parallel and Distributed Computing, 37(1):70–82, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1996.0108/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0108/production/pdf.

Findler:1986:EDP

[FL86] Nicholas V. Findler and Ron Lo. An examination of distributed planning in the world of air traffic control. Journal of Parallel and Distributed Computing, 3(3):411–431, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fan:2007:OFT

[FLPJ07] Jianxi Fan, Xiaola Lin, Yi Pan, and Xiaohua Jia. Optimal fault-tolerant embedding of paths in twisted cubes. Journal of Parallel and Distributed Computing, 67(2):205–214, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flaherty:1997:ALR

[FLS^{+}97] J. E. Flaherty, R. M. Loy, M. S. Shephard, B. K. Szymanski, J. D. Teresco, and L. H. Ziantz. Adaptive local refinement with octree load balancing for the parallel solution of threedimensional conservation laws. Journal of Parallel and Distributed Computing, 47(2):139–152, December 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1412/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1412/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1412/production/ref.

Fortes:1985:PDT

[FM85] J. A. B. Fortes and D. I. Moldovan. Parallelism detection and transformation techniques useful for VLSI algorithms. Journal of Parallel and Distributed Computing, 2(3):277–301, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flocchini:1996:OEL

[FM96] Paola Flocchini and Bernard Mans. Optimal elections in labeled hypercubes. Journal of Parallel and Distributed Computing, 33(1):76–83, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0026/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0026/production/pdf.

Fahringer:1999:BSC

[FM99a] Thomas Fahringer and Eduard Mehofer. Buffer-safe and costdriven communication optimization. Journal of Parallel and Distributed Computing, 57(1):33–63, April 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1512/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1512/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1512/production/ref.

Friedman:1999:LBS

[FM99b] Roy Friedman and Daniel Mosse. Load-balancing schemes for high-throughput distributed fault-tolerant servers. Journal of Parallel and Distributed Computing, 59(3):475–488, December 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1582/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1582/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1582/production/ref.

Fernandess:2007:CCD

[FM07] Yaacov Fernandess and Dahlia Malkhi. On collaborative content distribution using multi-message gossip. Journal of Parallel and Distributed Computing, 67(12):1232–1239, December 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flammini:2008:ATG

[FMM^{+}08] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Mordechai Shalom, and Shmuel Zaks. Approximating the traffic grooming problem in tree and star networks. Journal of Parallel and Distributed Computing, 68(7):939–948, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feschet:1998:PPD

[FMP98] Fabien Feschet, Serge Miguet, and Laurent Perroton. ParList: a parallel data structure for dynamic load balancing. Journal of Parallel and Distributed Computing, 51(2):114–135, June 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1457/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1457/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1457/production/ref.

Fraigniaud:2005:ETB

[FMR05] Pierre Fraigniaud, Bernard Mans, and Arnold L. Rosenberg. Efficient trigger-broadcasting in heterogeneous clusters. Journal of Parallel and Distributed Computing, 65(5):628–642, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Francis:1994:DPS

[FMW^{+}94] Rhys S. Francis, Ian D. Mathieson, Paul G. Whiting, Martin R. Dix, Harvey L. Davies, and Leon D. Rotstayn. A data parallel scientific modeling language. Journal of Parallel and Distributed Computing, 21(1):46–60, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1041/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1041/production/pdf.

Frolund:2003:DED

[FP03] Svend Frølund and Fernando Pedone. Dealing efficiently with data-center disasters. Journal of Parallel and Distributed Computing, 63(11):1064–1081, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

French:1993:PMC

[FPD93] James C. French, Terrence W. Pratt, and Mriganka Das. Performance measurement of the concurrent file system of the Intel iPSC/2 hypercube. Journal of Parallel and Distributed Computing, 17(1–2):115–121, January/February 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1009/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1009/production/pdf.

Fantozzi:2006:TSL

[FPP06] Carlo Fantozzi, Andrea Pietracaprina, and Geppino Pucci. Translating submachine locality into locality of reference. Journal of Parallel and Distributed Computing, 66(5):633–646, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flocchini:2008:CAB

[FPP^{+}08] P. Flocchini, L. Pagli, G. Prencipe, N. Santoro, and P. Widmayer. Computing all the best swap edges distributively. Journal of Parallel and Distributed Computing, 68(7):976–983, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feitelson:1992:GSP

[FR92] Dror G. Feitelson and Larry Rudolph. Gang scheduling performance benefits for fine-grain synchronization. Journal of Parallel and Distributed Computing, 16(4):306–318, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feitelson:1996:EDC

[FR96a] Dror G. Feitelson and Larry Rudolph. Evaluation of design choices for gang scheduling using distributed hierarchical control. Journal of Parallel and Distributed Computing, 35(1):18–34, May 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0064/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0064/production/pdf.

Ferreira:1996:FSP

[FR96b] Afonso Ferreira and John Michael Robson. Fast and scalable parallel algorithms for knapsack-like problems. Journal of Parallel and Distributed Computing, 39(1):1–13, November 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0150/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0150/production/
pdf.

Ferreira:1998:SII

[FR98] Afonso Ferreira and José Rolim. Special issue on irregular problems in supercomputing applications. Journal of Parallel and Distributed Computing, 50(1–2):1–2, April/May 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fraigniaud:1992:CAB

[Fra92] Pierre Fraigniaud. Complexity analysis of broadcasting in hypercubes with restricted communication capabilities. Journal of Parallel and Distributed Computing, 16(1):15–26, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Freeh:1996:CIE

[Fre96] Vincent W. Freeh. A comparison of implicit and explicit parallel programming. Journal of Parallel and Distributed Computing, 34(1):50–65, April 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0045/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0045/production/pdf.

Fleury:2004:DFG

[FSD04] M. Fleury, R. P. Self, and A. C. Downton. Development of a fine-grained parallel Karhunen–Loève transform. Journal of Parallel and Distributed Computing, 64(4):520–535, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Flammini:2007:MNA

[FSZ07] Michele Flammini, Mordechai Shalom, and Shmuel Zaks. On minimizing the number of ADMs—tight bounds for an algorithm without preprocessing. Journal of Parallel and Distributed Computing, 67(4):448–455, April 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Funabiki:1994:PAT

[FT94] Nobuo Funabiki and Yoshiyasu Takefuji. A parallel algorithm for traffic control problems in three-stage connecting networks. Journal of Parallel and Distributed Computing, 20(2):236–240, February 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1022/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1022/production/pdf.

Fidanova:2004:ILB

[FT04] Stefka Fidanova and Denis R. Trystram. Improved lower bounds for embedding hypercubes on de Bruijn graphs. Journal of Parallel and Distributed Computing, 64(3):327–329, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fu:2000:EEM

[FTC00] Shiwa S. Fu, Nian-Feng Tzeng, and Jen-Yao Chung. Empirical evaluation of mutual exclusion algorithms for distributed systems. Journal of Parallel and Distributed Computing, 60(7):785–806, July 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1635;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1635/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1635/ref.

Funabiki:1992:NNM

[FTL92] Nubuo Funabiki, Yoshiyasu Takefuji, and Kuo-Chun Lee. A neural network model for finding a near-maximum clique. Journal of Parallel and Distributed Computing, 14(3):340–344, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fraigniaud:1997:AAB

[FV97] Pierre Fraigniaud and Sandrine Vial. Approximation algorithms for broadcasting and gossiping. Journal of Parallel and Distributed Computing, 43(1):47–55, May 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1318/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1318/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1318/production/ref.

Franciscani:2005:RCA

[FVCL05] Fernanda P. Franciscani, Marisa A. Vasconcelos, Rainer P. Couto, and Antonio A. F. Loureiro. (re)configuration algorithms for peer-to-peer over ad hoc networks. Journal of Parallel and Distributed Computing, 65(2):234–245, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fang:2009:SDB

[FVLB09] Qiu Fang, Susan V. Vrbsky, Ming Lei, and Richard Borie. Scheduling on-demand broadcast with timing constraints. Journal of Parallel and Distributed Computing, 69(8):737–747, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Falsafi:2005:ESP

[FW05] Babak Falsafi and David A. Wood. Evaluating scheduling policies for fine-grain communication protocols on a cluster of SMPs. Journal of Parallel and Distributed Computing, 65 (4):464–478, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fu:2006:SMA

[FX06] Song Fu and Cheng-Zhong Xu. Stochastic modeling and analysis of hybrid mobility in reconfigurable distributed virtual machines. Journal of Parallel and Distributed Computing, 66(11): 1442–1454, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fong:2003:OPR

[FXW03] Ngo-Tai Fong, Cheng-Zhong Xu, and Le Yi Wang. Optimal periodic remapping of dynamic bulk synchronous computations. Journal of Parallel and Distributed Computing, 63(11):1036– 1049, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Feng:1986:FDC

[FY86] Tse-Yun Feng and Wei Young. Fault diagnosis for a class of rearrangeable networks. Journal of Parallel and Distributed Computing, 3(1):23–47, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fujita:1996:NBA

[FY96] Satoshi Fujita and Masafumi Yamashita. A nonoblivious bus access scheme yields an optimal partial sorting algorithm. Journal of Parallel and Distributed Computing, 34(1):111–116, April 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0049/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0049/production/pdf.

Fu:1997:RTT

[FY97] Cong Fu and Tao Yang. Run-time techniques for exploiting irregular task parallelism on distributed memory architectures. Journal of Parallel and Distributed Computing, 42(2):143–156, May 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1319/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1319/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1319/production/ref.

Feng:1990:FDM

[FZ90] Tse-Yun Feng and Qirui Zhang. On fault diagnosis of multistage networks with four valid states. Journal of Parallel and Distributed Computing, 9(3):237–251, July 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fang:2005:FSS

[FZC^{+}05] Zhen Fang, Lixin Zhang, John B. Carter, Liqun Cheng, and Michael Parker. Fast synchronization on shared-memory multiprocessors: An architectural approach. Journal of Parallel and Distributed Computing, 65(10):1158–1170, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fernandez-Zepeda:2002:UBL

[FZVT02] José Alberto Fernández-Zepeda, Ramachandran Vaidyanathan, and Jerry L. Trahan. Using bus linearization to scale the reconfigurable mesh. Journal of Parallel and Distributed Computing, 62(4):495–516, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1810;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1810/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1810/ref.

Goodrich:1990:PRO

[GA90] Michael T. Goodrich and Mikhail J. Atallah. On performing robust order statistics in tree-structured dictionary machines. Journal of Parallel and Distributed Computing, 9(1): 69–76, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gharachorloo:1992:PDM

[GAG^{+}92] Kourosh Gharachorloo, Sarita V. Adve, Anoop Gupta, John L. Hennessey, and Mark D. Hill. Programming for different memory consistency models. Journal of Parallel and Distributed Computing, 15(4):399–407, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Goff:2003:PRA

[GAGPK03] Tom Goff, Nael Abu-Ghazaleh, Dhananjay Phatak, and Ridvan Kahvecioglu. Preemptive routing in ad hoc networks. Journal of Parallel and Distributed Computing, 63(2):123–140, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gait:1987:DPM

[Gai87] Jason Gait. A distributed process manager for an engineering network computer. Journal of Parallel and Distributed Computing, 4(4):423–437, August 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gait:1990:SPM

[Gai90] Jason Gait. Scheduling and process migration in partitioned multiprocessors. Journal of Parallel and Distributed Computing, 8(3):274–279, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gao:1986:MPT

[Gao86] Guang R. Gao. A maximally pipelined tridiagonal linear equation solver. Journal of Parallel and Distributed Computing, 3 (2):215–235, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gao:1989:AAB

[Gao89] G. R. Gao. Algorithmic aspects of balancing techniques for pipelined data flow code generation. Journal of Parallel and Distributed Computing, 6(1):39–61, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gao:1993:EHD

[Gao93] G. R. Gao. An efficient hybrid dataflow architecture model. Journal of Parallel and Distributed Computing, 19(4):293–307, December 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1113/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1113/
production/pdf.

Gaudiot:2006:I

[Gau06] Jean-Luc Gaudiot. Introduction. Journal of Parallel and Distributed Computing, 66(7):873, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guha:1993:DON

[GB93] Aloke Guha and Julian Bristow. Designing optical networks from simple switching elements. Journal of Parallel and Distributed Computing, 17(3):212–221, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1020/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1020/production/pdf.

Gustafson:2006:SIM

[GB06] Steven Gustafson and Edmund K. Burke. The Speciating Island Model: An alternative parallel evolutionary algorithm. Journal of Parallel and Distributed Computing, 66(8):1025–1036, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gu:2008:PAA

[GBA08] Yunfeng Gu, Azzedine Boukerche, and Regina B. Araujo. Performance analysis of an adaptive dynamic Grid-based approach to data distribution management. Journal of Parallel and Distributed Computing, 68(4):536–547, April 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Germain:1993:CAM

[GBES93] Cecile Germain, Jean-Luc L. Bechennec, Daniel Etiemble, and Jean-Paul P. Sansonnet. A communication architecture for a massively parallel message-passing multicomputer. Journal of Parallel and Distributed Computing, 19(4):338–348, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1116/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1116/production/pdf.

Gupta:1993:MEH

[GBG93] Ajay Gupta, Steven C. Bruell, and Sukumar Ghosh. Mutual exclusion on a hypercube. Journal of Parallel and Distributed Computing, 17(4):327–336, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1031/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1031/production/pdf.

Guirguis:2007:AEE

[GBMZ07] Mina Guirguis, Azer Bestavros, Ibrahim Matta, and Yuting Zhang. Adversarial exploits of end-systems adaptation dynamics. Journal of Parallel and Distributed Computing, 67(3):318– 335, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ghafoor:1995:SIM

[GC95] Arif Ghafoor and C-Y Roger Chen. Special issue on multimedia processing and technology: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 30(2):107–110, November 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1131/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1131/production/pdf.

Goil:2001:PIP

[GC01] Sanjay Goil and Alok Choudhary. PARSIMONY: An infrastructure for parallel multidimensional analysis and data mining. Journal of Parallel and Distributed Computing, 61(3):285–321, March 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1691;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2000.1691/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1691/ref.

Grosu:2005:NLB

[GC05] Daniel Grosu and Anthony T. Chronopoulos. Noncooperative load balancing in distributed systems. Journal of Parallel and Distributed Computing, 65(9):1022–1034, September 2005. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:2007:OSA

[GC07] N. Gupta and S. Chopra. Output-sensitive algorithms for optimally constructing the upper envelope of straight line segments in parallel. Journal of Parallel and Distributed Computing, 67 (7):772–782, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Garti:2000:OMP

[GCB^{+}00] Dror Garti, Shem-Tov Cohen, Amnon Barak, Arie Keren, and Ricardo Szmit. Object mobility for performance improvements of parallel Java applications. Journal of Parallel and Distributed Computing, 60(10):1311–1324, October 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1655;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1655/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1655/ref.

Garg:1997:EDC

[GCKM97] V. K. Garg, C. M. Chase, Richard Kilgore, and J. Roger Mitchell. Efficient detection of channel predicates in distributed systems. Journal of Parallel and Distributed Computing, 45(2):134–147, September 15, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1374/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1374/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1374/production/ref.

Giraud-Carrier:1995:ACG

[GCM95] Christophe Giraud-Carrier and Tony Martinez. Analysis of the convergence and generalization of AA1. Journal of Parallel and Distributed Computing, 26(1):125–131, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1053/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1053/production/pdf.

Giruka:2006:DMP

[GCS06] Venkata C. Giruka, Saikat Chakrabarti, and Mukesh Singhal. A distributed multi-party key agreement protocol for dynamic collaborative groups using ECC. Journal of Parallel and Distributed Computing, 66(7):959–970, July 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ganesan:2004:NIW

[GCY^{+}04] Deepak Ganesan, Alberto Cerpa, Wei Ye, Yan Yu, Jerry Zhao, and Deborah Estrin. Networking issues in wireless sensor networks. Journal of Parallel and Distributed Computing, 64 (7):799–814, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Goodeve:1998:TMS

[GDN^{+}98] D. M. Goodeve, S. A. Dobson, J. M. Nash, J. R. Davy, P. M. Dew, M. Kara, and C. P. Wadsworth. Toward a model for shared data abstraction with performance. Journal of Parallel and Distributed Computing, 49(1):156–167, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1427/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1427/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1427/production/ref.

Gandham:2008:LSW

[GDP08] Shashidhar Gandham, Milind Dawande, and Ravi Prakash. Link scheduling in wireless sensor networks: Distributed edgecoloring revisited. Journal of Parallel and Distributed Computing, 68(8):1122–1134, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gaudiot:1985:PES

[GE85] J. L. Gaudiot and M. D. Ercegovac. Performance evaluation of a simulated data-flow computer with low-resolution actors. Journal of Parallel and Distributed Computing, 2(4):321–351, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gusev:1994:NMV

[GE94] M. Gusev and D. J. Evans. A new matrix vector product systolic array. Journal of Parallel and Distributed Computing, 22(2):346–349, August 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1094/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1094/production/pdf.

Gerbessiotis:1998:PCP

[Ger98] Alexandros V. Gerbessiotis. Practical considerations of parallel simulations and architecture independent parallel algorithm design. Journal of Parallel and Distributed Computing, 53(1):1–25, August 25, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1477/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1477/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1477/production/ref.

Gupta:1989:SRT

[GF89] Anoop Gupta and Charles L. Forgy. Static and run-time characteristics of OPS5 production systems. Journal of Parallel and Distributed Computing, 7(1):64–95, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Garzon:1992:DTG

[GFB^{+}92] Max H. Garzon, Stanley P. Franklin, William Baggett, William S. Boyd, Jr., and Dinah Dickerson. Design and testing of a general-purpose neurocomputer. Journal of Parallel and Distributed Computing, 14(3):203–220, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gabarro:1994:ACD

[GG94] Joaquim Gabarró and Ricard Gavaldà. An approach to correctness of data parallel algorithms. Journal of Parallel and Distributed Computing, 22(2):185–201, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1080/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1080/production/pdf.

Gavoille:2001:SER

[GG01] Cyril Gavoille and Marc Gengler. Space-efficiency for routing schemes of stretch factor three. Journal of Parallel and Distributed Computing, 61(5):679–687, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1705;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1705/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1705/ref.

Gao:1993:DMA

[GGB93] Guang Gao, Jean-Luc Gaudiot, and Lubomir Bic. Dataflow and multithreaded architectures: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 18(3):271–272, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ghosh:1993:PEP

[GGD93] Joydeep Ghosh, Kelvin D. Goveas, and Jeffrey T. Draper. Performance evaluation of a parallel I/O subsystem for hypercube multicomputers. Journal of Parallel and Distributed Computing, 17(1–2):90–106, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1007/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1007/production/pdf.

Gaujal:1993:SAM

[GGN93] Bruno Gaujal, Albert G. Greenberg, and David M. Nicol. A sweep algorithm for massively parallel simulation of circuit-switched networks. Journal of Parallel and Distributed Computing, 18(4):484–500, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1079/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1079/production/pdf.

Ghoshal:1989:PMP

[GGR89] S. K. Ghoshal, M. Gupta, and V. Rajaraman. A parallel multistep predictor-corrector algorithm for solving ordinary differential equations. Journal of Parallel and Distributed Computing, 6(3):636–648, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ganesan:1996:CSM

[GGW96] Ravikanth Ganesan, Kannan Govindarajan, and Min-You Wu. Comparing SIMD and MIMD programming modes. Journal of Parallel and Distributed Computing, 35(1):91–96, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0071/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0071/production/
pdf.

Ghosh:1989:MNN

[GH89a] Joydeep Ghosh and Kai Hwang. Mapping neural networks onto message-passing multicomputers. Journal of Parallel and Distributed Computing, 6(2):291–330, April 1989. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guerra:1989:PAL

[GH89b] C. Guerra and S. Hambrusch. Parallel algorithms for line detection on a mesh. Journal of Parallel and Distributed Computing, 6(1):1–19, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Grafe:1990:EMS

[GH90] V. G. Grafe and J. E. Hoch. The EPSILON-2 multiprocessor system. Journal of Parallel and Distributed Computing, 10(4): 309–318, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:1993:MNE

[GH93] Ajay K. Gupta and Susanne E. Hambrusch. Multiple network embeddings into hypercubes. Journal of Parallel and Distributed Computing, 19(2):73–82, October 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1993.1092/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1092/production/pdf.

Gouda:1996:STR

[GH96] Mohamed G. Gouda and F. Furman Haddix. The stabilizing token ring in three bits. Journal of Parallel and Distributed Computing, 35(1):43–48, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0066/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0066/production/pdf.

Ghosh:2002:SSS

[GH02] Sukumar Ghosh and Xin He. Scalable self-stabilization. Journal of Parallel and Distributed Computing, 62(5):945–960, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1824;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1824/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1824/ref.

Guesgen:1992:TMP

[GHH92] Hans W. Guesgen, Kinson Ho, and Paul N. Hilfinger. A tagging method for parallel constraint satisfaction. Journal of Parallel and Distributed Computing, 16(1):72–75, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Grammatikakis:1998:PRF

[GHKS98] Miltos D. Grammatikakis, D. Frank Hsu, Miro Kraetzl, and Jop F. Sibeyn. Packet routing in fixed-connection networks: a survey. Journal of Parallel and Distributed Computing, 54(2):77–132, November 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1483/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1483/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1483/production/ref.

Gustafson:1986:AHV

[GHS86] John L. Gustafson, Stuart Hawkinson, and Ken Scott. The architecture of a homogeneous vector supercomputer. Journal of Parallel and Distributed Computing, 3(3):297–304, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gottlieb:1994:EM

[GHS94] Allan Gottlieb, Kai Hwang, and Sartaj Sahni. Editorial message. Journal of Parallel and Distributed Computing, 20(1):v– vii, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gottlieb:1995:EM

[GHS95] Allan Gottlieb, Kai Hwang, and Sartaj Sahni. Editorial message. Journal of Parallel and Distributed Computing, 31(2):99–100, December 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1148/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1148/production/pdf.

Gottlieb:1996:EA

[GHS96] Allan Gottlieb, Kai Hwang, and Sartaj Sahni. Editorial announcement. Journal of Parallel and Distributed Computing, 35(1):1, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gottlieb:1997:EM

[GHS97] Allan Gottlieb, Kai Hwang, and Sartaj Sahni. Editorial message. Journal of Parallel and Distributed Computing, 43(1):1– 2, May 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1359/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1359/
production/pdf.

Gupta:1996:FGD

[GHSJ96] S. K. S. Gupta, C.-H. Huang, P. Sadayappan, and R. W. Johnson. A framework for generating distributed-memory parallel programs for block recursive algorithms. Journal of Parallel and Distributed Computing, 34(2):137–153, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0051/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0051/production/pdf.

Gil:1994:RDT

[Gil94] J. Gil. Renaming and dispersing: Techniques for fast load balancing. Journal of Parallel and Distributed Computing, 23(2):149–157, November 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1128/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1128/production/pdf.

Glimcher:2008:MDM

[GJA08] Leonid Glimcher, Ruoming Jin, and Gagan Agrawal. Middleware for data mining applications on clusters and Grids. Journal of Parallel and Distributed Computing, 68(1):37–53, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gannon:1988:SCL

[GJG88] D. Gannon, W. Jalby, and K. Gallivan. Strategies for cache and local memory management by global program transformation. Journal of Parallel and Distributed Computing, 5(5):587–616, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gjertsen:1996:PHI

[GJP96] Robert K. Gjertsen, Jr., Mark T. Jones, and Paul E. Plassmann. Parallel heuristics for improved, balanced graph colorings. Journal of Parallel and Distributed Computing, 37(2):171–186, September 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0117/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0117/production/pdf.

Guo:2005:FLC

[GJXZ05] Lei Guo, Song Jiang, Li Xiao, and Xiaodong Zhang. Fast and low-cost search schemes by exploiting localities in P2P networks. Journal of Parallel and Distributed Computing, 65 (6):729–742, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:1993:PPL

[GK93] Anshul Gupta and Vipin Kumar. Performance properties of large scale parallel systems. Journal of Parallel and Distributed Computing, 19(3):234–244, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1107/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1107/production/pdf.

Ghose:1998:LPT

[GK98] Debasish Ghose and Hyoung Joong Kim. Load partitioning and trade-off study for large matrix-vector computations in multicast bus networks with communication delays. Journal of Parallel and Distributed Computing, 55(1):32–59, November 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1494/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1494/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1494/production/ref.

Gursoy:2004:PMB

[GK04] Attila Gürsoy and L. V. Laxmikant V. Kale. Performance and modularity benefits of message-driven execution. Journal of Parallel and Distributed Computing, 64(4):461–480, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:1996:CAE

[GKHS96] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Compiling array expressions for efficient execution on distributed-memory machines. Journal of Parallel and Distributed Computing, 32(2):155–172, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0011/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0011/production/pdf.

Gaudiot:1989:OMP

[GL89] Jean-Luc Gaudiot and Liang-Teh Lee. Occamflow: a methodology for programming multiprocessor systems. Journal of Parallel and Distributed Computing, 7(1):96–124, August 1989. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:1990:OOL

[GL90] M. Gupta and K. Y. Lee. The n-omega: an omega-like topology for distributed-memory multiprocessors. Journal of Parallel and Distributed Computing, 9(4):398–410, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guan:1992:PMS

[GL92] Xiaojun Guan and Michael A. Langston. Parallel methods for solving fundamental file rearrangement problems. Journal of Parallel and Distributed Computing, 14(4):436–439, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gu:2001:JBP

[GLC01] Yan Gu, Bu-Sung Lee, and Wentong Cai. JBSP: a BSP programming library in Java. Journal of Parallel and Distributed Computing, 61(8):1126–1142, August 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1735;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1735/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1735/ref.

Gomez:2006:FER

[GLD06] M. E. Gómez, P. López, and J. Duato. FIR: An efficient routing strategy for tori and meshes. Journal of Parallel and Distributed Computing, 66(7):907–921, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ghose:1994:DCC

[GM94a] D. Ghose and V. Mani. Distributed computation with communication delays: Asymptotic performance analysis. Journal of Parallel and Distributed Computing, 23(3):293–305, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1141/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1141/production/
pdf.

Gil:1994:FES

[GM94b] J. Gil and Y. Matias. Fast and efficient simulations among CRCW PRAMs. Journal of Parallel and Distributed Computing, 23(2):135–148, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1127/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1127/production/pdf.

Gopi:1995:PAC

[GM95] Meenakshi Sundaram Gopi and Swami Manohar. A parallel architecture for the computation of uniform rational B-spline patches. Journal of Parallel and Distributed Computing, 30(1):91–98, October 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1129/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1129/production/pdf.

Gil:1996:ELB

[GM96] Joseph Gil and Yossi Matias. An effective load balancing policy for geometric-decaying algorithms. Journal of Parallel and Distributed Computing, 36(2):185–188, August 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0098/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0098/production/pdf.

Graham:1999:PSS

[GM99] William C. Graham and Shikharesh Majumdar. Performance of scheduling strategies for client-server systems. Journal of Parallel and Distributed Computing, 58(3):389–424, September 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1561/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1561/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1561/production/ref.

Guo:1991:PCO

[GMH^{+}91] Zicheng Guo, Rami G. Melhem, Richard W. Hall, Donald M. Chiarulli, and Steven P. Levitan. Pipelined communications in optically interconnected arrays. Journal of Parallel and Distributed Computing, 12(3):269–282, July 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:2000:NSI

[GMM00] Indranil Gupta, G. Manimaran, and C. Siva Ram Murthy. A new strategy for improving the effectiveness of resource reclaiming algorithms in multiprocessor real-time systems. Journal of Parallel and Distributed Computing, 60(1):113–133, January 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1594/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1594/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1594/production/ref.

Gupta:2006:WAH

[GMS06] Vishakha Gupta, Gaurav Mathur, and Anil M. Shende. Wireless ad hoc lattice computers (WAdL). Journal of Parallel and Distributed Computing, 66(4):531–541, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:2007:DKG

[GMXA07] Ananya Gupta, Anindo Mukherjee, Bin Xie, and Dharma P. Agrawal. Decentralized key generation scheme for cellularbased heterogeneous wireless ad hoc networks. Journal of Parallel and Distributed Computing, 67(9):981–991, September 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Georgiou:2009:FTS

[GNS09] Chryssis Georgiou, Nicolas C. Nicolaou, and Alexander A. Shvartsman. Fault-tolerant semifast implementations of atomic read/write registers. Journal of Parallel and Distributed Computing, 69(1):62–79, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Georgiadis:2004:FWA

[GNT04] Leonidas Georgiadis, Christos Nikolaou, and Alexander Thomasian. A fair workload allocation policy for heterogeneous systems. Journal of Parallel and Distributed Computing, 64(4):507–519, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:2003:EET

[GNW03] Ajay K. Gupta, Donald Nelson, and Hong Wang. Efficient embeddings of ternary trees into hypercubes. Journal of Parallel and Distributed Computing, 63(6):619–629, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Greenberg:1995:PRN

[GO95] Ronald I. Greenberg and H-C C. Oh. Packet routing in networks with long wires. Journal of Parallel and Distributed Computing, 31(2):153–158, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1153/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1153/production/pdf.

Goertzel:1994:LIP

[Goe94] B. Goertzel. Lagrange interpolation on a processor tree with ring connections. Journal of Parallel and Distributed Computing, 22(2):321–323, August 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1090/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1090/production/pdf.

Gonzalez:1998:CAM

[Gon98] Teofilo F. Gonzalez. Complexity and approximations for multimessage multicasting. Journal of Parallel and Distributed Computing, 55(2):215–235, December 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1500/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1500/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1500/production/ref.

Goscinski:1990:TAM

[Gos90] A. Goscinski. Two algorithms for mutual exclusion in real-time distributed computer systems. Journal of Parallel and Distributed Computing, 9(1):77–82, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Glenn:1991:IMP

[GP91] Raymond R. Glenn and Daniel V. Pryor. Instrumentation for a massively parallel MIMD application. Journal of Parallel and Distributed Computing, 12(3):223–236, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gibson:1993:DDA

[GP93] Garth A. Gibson and David A. Patterson. Designing disk arrays for high data reliability. Journal of Parallel and Distributed Computing, 17(1–2):4–27, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1001/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1001/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1993.1002/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1002/production/pdf.

Galil:1994:PAD

[GP94] Zvi Galil and Kunsoo Park. Parallel algorithms for dynamic programming recurrences with more than O(1) dependency. Journal of Parallel and Distributed Computing, 21(2):213– 222, May 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1053/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1053/
production/pdf.

Gasieniec:1997:BBF

[GP97] Leszek Gasieniec and Andrzej Pelc. Broadcasting with a bounded fraction of faulty nodes. Journal of Parallel and Distributed Computing, 42(1):11–20, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1303/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1303/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1303/production/ref.

Gu:2000:EAP

[GP00] Qian-Ping Gu and Shietung Peng. An efficient algorithm for the k-pairwise disjoint paths problem in hypercubes. Journal of Parallel and Distributed Computing, 60(6):764–774, June 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1632;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1632/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1632/ref.

Guo:2005:ICS

[GP05] Minyi Guo and Yi Pan. Improving communication scheduling for array redistribution. Journal of Parallel and Distributed Computing, 65(5):553–563, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gandhi:2007:DAC

[GP07] Rajiv Gandhi and Srinivasan Parthasarathy. Distributed algorithms for connected domination in wireless networks. Journal of Parallel and Distributed Computing, 67(7):848–862, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Greenberg:1996:CCC

[GPS96] David S. Greenberg, James K. Park, and Eric J. Schwabe. The cost of complex communication on simple networks. Journal of Parallel and Distributed Computing, 35(2):133–141, June 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0076/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0076/production/
pdf.

Gaibisso:2006:EMT

[GPT06a] Carlo Gaibisso, Guido Proietti, and Richard B. Tan. Efficient management of transient station failures in linear radio communication networks with bases. Journal of Parallel and Distributed Computing, 66(4):556–565, April 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Goldman:2006:EMD

[GPT06b] Alfredo Goldman, Joseph G. Peters, and Denis Trystram. Exchanging messages of different sizes. Journal of Parallel and Distributed Computing, 66(1):1–18, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Geng:2008:DSA

[GPX08] Xiutang Geng, Linqiang Pan, and Jin Xu. A DNA sticker algorithm for bit-substitution in a block cipher. Journal of Parallel and Distributed Computing, 68(9):1201–1206, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gao:1996:TES

[GR96] Li-Xin Gao and Arnold L. Rosenberg. Toward efficient scheduling of evolving computations on rings of processors. Journal of Parallel and Distributed Computing, 38(1):92–100, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0131/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0131/production/
pdf.

Grahn:2009:SIJ

[Gra09] Håkan Grahn. Special issue of the Journal of Parallel and Distributed Computing: Transactional memory. Journal of Parallel and Distributed Computing, 69(9):814, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ghosh:2005:PSJ

[GRDB05] Preetam Ghosh, Nirmalya Roy, Sajal K. Das, and Kalyan Basu. A pricing strategy for job allocation in mobile grids using a non-cooperative bargaining theory framework. Journal of Parallel and Distributed Computing, 65(11):1366–1383, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gustafson:1991:DSF

[GREC91] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. The design of a scalable, fixed-time computer benchmark. Journal of Parallel and Distributed Computing, 12(4):

388–401, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gross:1985:SIF

[Gro85] Thomas Gross. Software implementation of floating-point arithmetic on a reduced-instruction-set processor. Journal of Parallel and Distributed Computing, 2(4):362–375, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gu:1993:DPM

[GRR93] Dechang C. Gu, D. J. Rosenkrantz, and S. S. Ravi. Determining performance measures of algorithm-based fault tolerant systems. Journal of Parallel and Distributed Computing, 18(1): 56–70, May 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1044/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1044/
production/pdf.

Gutierrez:2005:PTI

[GRR^{+}05] Eladio Gutiérrez, Sergio Romero, Luis F. Romero, Oscar Plata, and Emilio L. Zapata. Parallel techniques in irregular codes: cloth simulation as case of study. Journal of Parallel and Distributed Computing, 65(4):424–436, April 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gomez:1997:EMU

[GRS97] Juan Carlos Gomez, Vernon Rego, and V. S. Sunderam. Efficient multithreaded user-space transport for network computing: Design and test of the TRAP protocol. Journal of Parallel and Distributed Computing, 40(1):103–117, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1269/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1269/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1269/production/ref.

Gargano:1997:CCG

[GRV97] Luisa Gargano, Adele A. Rescigno, and Ugo Vaccaro. Communication complexity of gossiping by packets. Journal of Parallel and Distributed Computing, 45(1):73–81, August 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1357/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1357/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1357/production/ref.

Gallet:2008:CDP

[GRV08] Matthieu Gallet, Yves Robert, and Frédéric Vivien. Comments on “Design and performance evaluation of load distribution strategies for multiple loads on heterogeneous linear daisy chain networks”. Journal of Parallel and Distributed Computing, 68 (7):1021–1031, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [MVB05].

Gilbert:1991:OEE

[GS91a] John R. Gilbert and Robert Schreiber. Optimal expression evaluation for data parallel architectures. Journal of Parallel and Distributed Computing, 13(1):58–64, September 1, 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gotze:1991:VSO

[GS91b] J. Gotze and U. Schwiegelshohn. VLSI-suited orthogonal solution of systems of linear equations. Journal of Parallel and Distributed Computing, 11(4):276–283, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Grahn:1996:ECU

[GS96] Håkan Grahn and Per Stenström. Evaluation of a competitiveupdate cache coherence protocol with migratory data detection. Journal of Parallel and Distributed Computing, 39(2):168–180, December 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0164/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0164/production/pdf.

Gerbessiotis:1998:OLG

[GS98] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Ordered h-level graphs on the BSP model. Journal of Parallel and Distributed Computing, 49(1):98–110, February 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1430/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1430/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1430/production/ref.

Gupta:1999:STP

[GS99] Srabani Sen Gupta and Bhabani P. Sinha. A simple O(log N) time parallel algorithm for testing isomorphism of maximal outerplanar graphs. Journal of Parallel and Distributed Computing, 56(2):144–155, February 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1514/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1514/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1514/production/ref.

Grahn:2000:CEL

[GS00] Håkan Grahn and Per Stenström. Comparative evaluation of latency-tolerating and -reducing techniques for hardware-only and software-only directory protocols. Journal of Parallel and Distributed Computing, 60(7):807–834, July 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1606;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1606/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1606/ref.

Govindarajan:2001:SIC

[GS01a] R. Govindarajan and Anand Sivasubramaniam. Special issue on cluster and network-based computing. Journal of Parallel and Distributed Computing, 61(11):1507–1511, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1742;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1742/pdf.

Gupta:2001:CBT

[GS01b] Sandeep K. S. Gupta and Pradip K. Srimani. Cored-based tree with forwarding regions (CBT-FR); a protocol for reliable multicasting in mobile ad hoc networks. Journal of Parallel and Distributed Computing, 61(9):1249–1277, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1749;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1749/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1749/ref.

Gupta:2003:FOS

[GS03a] Neelima Gupta and Sandeep Sen. Faster output-sensitive parallel algorithms for 3D convex hulls and vector maxima. Journal of Parallel and Distributed Computing, 63(4):488–500, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gupta:2003:SSM

[GS03b] Sandeep K. S. Gupta and Pradip K. Srimani. Self-stabilizing multicast protocols for ad hoc networks. Journal of Parallel and Distributed Computing, 63(1):87–96, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Goldstein:1996:LTI

[GSC96] Seth Copen Goldstein, Klaus Erik Schauser, and David E. Culler. Lazy threads: Implementing a fast parallel call. Journal of Parallel and Distributed Computing, 37(1):5–20, August 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0103/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0103/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0104/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0104/production/pdf.

Gill:1993:STA

[GSG^{+}93] D. H. Gill, T. J. Smith, T. E. Gerasch, J. V. Warren, C. L. McCreary, and R. E. K. Stirewalt. Spatial-temporal analysis of program dependence graphs for useful parallelism. Journal of Parallel and Distributed Computing, 19(2):103–118, October 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1095/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1095/production/
pdf.

Goswami:2002:DPP

[GSP02] Dhrubajyoti Goswami, Ajit Singh, and Bruno R. Preiss. From design patterns to parallel architectural skeletons. Journal of Parallel and Distributed Computing, 62(4):669–695, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1809;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1809/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1809/ref.

Garcia:2003:HDH

[GSSS03] Fabián García, Julio Solano, Ivan Stojmenovic, and Milos Stojmenovic. Higher dimensional hexagonal networks. Journal of Parallel and Distributed Computing, 63(11):1164–1172, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Girault:2009:RVP

[GST09] Alain Girault, Érik Saule, and Denis Trystram. Reliability versus performance for critical applications. Journal of Parallel and Distributed Computing, 69(3):326–336, March 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gentile:2004:IGS

[GSWW04] Antonio Gentile, Sam Sander, Linda Wills, and Scott Wills. The impact of grain size on the efficiency of embedded SIMD image processing architectures. Journal of Parallel and Distributed Computing, 64(11):1318–1327, November 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gu:1997:RPH

[GT97] Qian-Ping Gu and Hisao Tamaki. Routing a permutation in the hypercube by two sets of edge disjoint paths. Journal of Parallel and Distributed Computing, 44(2):147–152, August 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1358/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1358/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1358/production/ref.

Geisler:2002:PCC

[GT02] Jonathan Geisler and Valerie Taylor. Performance coupling: Case studies for improving the performance of scientific applications. Journal of Parallel and Distributed Computing, 62(8):1227–1247, August 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1839;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1839/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1839/ref.

Goldman:2004:EPA

[GT04] A. Goldman and D. Trystram. An efficient parallel algorithm for solving the Knapsack problem on hypercubes. Journal of Parallel and Distributed Computing, 64(11):1213–1222, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gilbert:2006:IVG

[GTN^{+}06] Laura Gilbert, Jeff Tseng, Rhys Newman, Saeed Iqbal, Ronald Pepper, Onur Celebioglu, Jenwei Hsieh, Victor Mashayekhi, and Mark Cobban. Implications of virtualization on Grids for high energy physics applications. Journal of Parallel and Distributed Computing, 66(7):922–930, July 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guerra:1986:VAO

[Gue86] Concettina Guerra. A VLSI algorithm for the optimal detection of curve. Journal of Parallel and Distributed Computing, 3 (2):206–214, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guo:1994:OIP

[Guo94] Zicheng C. Guo. Optically interconnected processor arrays with switching capability. Journal of Parallel and Distributed Computing, 23(3):314–329, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1143/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1143/production/pdf.

Gupta:1992:SEP

[Gup92] Rajiv Gupta. SPMD execution of programs with pointerbased data structures on distributed-memory machines. Journal of Parallel and Distributed Computing, 16(2):92–107, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gannon:1986:SPH

[GV86] Dennis Gannon and John Van Rosendale. On the structure of parallelism in a highly concurrent PDE solver. Journal of Parallel and Distributed Computing, 3(1):106–135, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gerbessiotis:1994:DBS

[GV94] A. V. Gerbessiotis and L. G. Valiant. Direct bulk-synchronous parallel algorithms. Journal of Parallel and Distributed Computing, 22(2):251–267, August 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1085/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1085/production/pdf.

Guo:2008:SCT

[GVA^{+}08] Yao Guo, Vladimir Vlassov, Raksit Ashok, Richard Weiss, and Csaba Andras Moritz. Synchronization coherence: a transparent hardware mechanism for cache coherence and fine-grained synchronization. Journal of Parallel and Distributed Computing, 68(2):165–181, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gil:1999:AMD

[GW99] Joseph Gil and Alan Wagner. An alternative mapping of 3-D space onto processor arrays. Journal of Parallel and Distributed Computing, 59(3):360–380, December 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1585/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1585/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1585/production/ref.

Gunney:2006:PCA

[GWH06] Brian T. N. Gunney, Andrew M. Wissink, and David A. Hysom. Parallel clustering algorithms for structured AMR. Journal of Parallel and Distributed Computing, 66(11):1419–1430, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Grimshaw:1994:MAC

[GWWL94] Andrew S. Grimshaw, Jon B. Weissman, Emily A. West, and Ed C. Loyot, Jr. Metasystems: an approach combining parallel processing and heterogeneous distributed computing systems. Journal of Parallel and Distributed Computing, 21(3):257–270, June 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1057/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1057/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1058/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1058/production/pdf.

Gerasoulis:1992:CCH

[GY92] Apostolos Gerasoulis and Tao Yang. A comparison of clustering heuristics for scheduling directed acyclic graphs on multiprocessors. Journal of Parallel and Distributed Computing, 16(4): 276–291, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Guil:1997:FHT

[GZ97] N. Guil and E. L. Zapata. Fast Hough transform on multiprocessors: a branch and bound approach. Journal of Parallel and Distributed Computing, 45(1):82–89, August 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1362/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1362/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1362/production/ref.

Gao:2008:PSA

[GZ08] Feng Gao and Mohammed J. Zaki. PSIST: a scalable approach to indexing protein structures using suffix trees. Journal of Parallel and Distributed Computing, 68(1):54–63, January 2008.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gotthelf:2008:GOM

[GZMC08] Pablo Gotthelf, Alejandro Zunino, Cristian Mateos, and Marcelo Campo. GMAC: An overlay multicast network for mobile agent platforms. Journal of Parallel and Distributed Computing, 68(8):1081–1096, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Houstis:1991:CPA

[HA91] Catherine E. Houstis and Mokhtar Aboelaze. A comparative performance analysis of mapping applications to parallel multiprocessor systems: a case study. Journal of Parallel and Distributed Computing, 13(1):17–29, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hudak:1992:CTO

[HA92] David E. Hudak and Santosh G. Abraham. Compile-time optimization of near-neighbor communication for scalable sharedmemory multiprocessors. Journal of Parallel and Distributed Computing, 15(4):368–381, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hughes:2005:MSP

[HA05] Christopher J. Hughes and Sarita V. Adve. Memory-side prefetching for linked data structures for processor-in-memory systems. Journal of Parallel and Distributed Computing, 65 (4):448–463, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Holman:2006:SLF

[HA06] Philip Holman and James H. Anderson. Supporting lock-free synchronization in Pfair-scheduled real-time systems. Journal of Parallel and Distributed Computing, 66(1):47–67, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hagerup:1997:AIT

[Hag97] Torben Hagerup. Allocating independent tasks to parallel processors: An experimental study. Journal of Parallel and Distributed Computing, 47(2):185–197, December 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1411/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1411/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1411/production/ref.

Haldar:2005:CRV

[Hal05] Sibsankar Haldar. Constructing regular variables in message passing systems. Journal of Parallel and Distributed Computing, 65(1):15–28, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Han:1989:PAC

[Han89] Yijie Han. Parallel algorithms for computing linked list prefix. Journal of Parallel and Distributed Computing, 6(3):537–557, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Harrison:1991:AMM

[Har91] Peter G. Harrison. Analytic models for multistage interconnection networks. Journal of Parallel and Distributed Computing, 12(4):357–369, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hawking:1997:DMA

[Haw97] David Hawking. A distributed memory algorithm for lexicon building. Journal of Parallel and Distributed Computing, 44(1):80–87, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1997.1344/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1344/production/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1344/production/ref.

Hasteer:1997:SAB

[HB97] Gagan Hasteer and Prithviraj Banerjee. Simulated annealing based parallel state assignment of finite state machines. Journal of Parallel and Distributed Computing, 43(1):21–35, May 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1325/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1325/production/

pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1325/production/ref.

Harchol-Balter:1999:CTA

[HBCM99] Mor Harchol-Balter, Mark E. Crovella, and Cristina D. Murta. On choosing a task assignment policy for a distributed server system. Journal of Parallel and Distributed Computing, 59(2):204–228, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1577/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1577/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1577/production/ref.

Hu:1993:EIS

[HBH93] X. Hu, S. C. Bass, and R. G. Harber. An efficient implementation of singular value decomposition rotation transformations with CORDIC processors. Journal of Parallel and Distributed Computing, 17(4):360–362, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1034/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1034/production/pdf.

Helman:1998:RPS

[HBJ98] David R. Helman, David A. Bader, and Joseph JáJá. A randomized parallel sorting algorithm with an experimental study. Journal of Parallel and Distributed Computing, 52(1):1–23, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1462/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1462/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1462/production/ref.

Hwang:1991:SPR

[HC91] Kai Hwang and Chien-Ming Cheng. Simulated performance of a RISC-based multiprocessor using orthogonal-access memory. Journal of Parallel and Distributed Computing, 13(1):43– 57, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hollis:1995:MIR

[HC95] J. E. L. Hollis and T. E. Cronk. Multiprocessor image rotation. Journal of Parallel and Distributed Computing, 30(1):99–106, October 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1130/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1130/
production/pdf.

Hui:1997:TAH

[HC97] Chi-Chung Hui and Samuel T. Chanson. Theoretical analysis of the heterogeneous dynamic load-balancing problem using a hydrodynamic approach. Journal of Parallel and Distributed Computing, 43(2):139–146, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1337/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1337/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1337/production/ref.

He:2009:PBI

[HC09] Weiping He and Ing-Ray Chen. A proxy-based integrated cache consistency and mobility management scheme for client-server applications in Mobile IP systems. Journal of Parallel and Distributed Computing, 69(6):559–572, June 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hicks:1993:PSI

[HCAA93] James Hicks, Derek Chiou, Boon Seong Ang, and Arvind. Performance studies of Id on the Monsoon Dataflow System. Journal of Parallel and Distributed Computing, 18(3):273–300, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1064/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1064/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1065/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1065/production/pdf.

Hurwitz:2005:AMP

[HcF05] Justin (Gus) Hurwitz and Wu chun Feng. Analyzing MPI performance over 10-gigabit Ethernet. Journal of Parallel and Distributed Computing, 65(10):1253–1260, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Herbordt:2000:SEP

[HCS^{+}00] Martin C. Herbordt, Jade Cravy, Renoy Sam, Owais Kidwai, and Calvin Lin. A system for evaluating performance and cost of SIMD array designs. Journal of Parallel and Distributed Computing, 60(2):217–246, February 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1602/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1602/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1602/production/ref.

Herbordt:1994:PAO

[HCWS94] Martin C. Herbordt, James C. Corbett, Charles C. Weems, and John Spalding. Practical algorithms for online routing on fixed and reconfigurable meshes. Journal of Parallel and Distributed Computing, 20(3):341–356, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1032/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1032/production/pdf.

Herbordt:2004:ACH

[HCZ04] Martin C. Herbordt, Jade Cravy, and Honghai Zhang. Array control for high-performance SIMD systems ^{∗1} ^{,} ^{∗2}. Journal of Parallel and Distributed Computing, 64(3):400–413, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hoare:2008:TSH

[HDJ08] Raymond R. Hoare, Zhu Ding, and Alex K. Jones. A two-stage hardware scheduler combining greedy and optimal scheduling. Journal of Parallel and Distributed Computing, 68(11):1437– 1451, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hoare:2005:FDS

[HDT^{+}05] Raymond R. Hoare, Zhu Ding, Shenchih Tung, Rami Melhem, and Alex K. Jones. A framework for the design, synthesis and cycle-accurate simulation of multiprocessor networks. Journal of Parallel and Distributed Computing, 65(10):1237–1252, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Herath:1992:GEI

[Her92] Jayantha Herath. Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 15(2):87–89, June 1992. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hyde:1996:ADS

[HF96] Randall L. Hyde and Brett D. Fleisch. An analysis of degenerate sharing and false coherence. Journal of Parallel and Distributed Computing, 34(2):183–195, May 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0054/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0054/production/pdf.

Hybinette:2002:LHO

[HF02] Maria Hybinette and Richard M. Fujimoto. Latency hiding with optimistic computations. Journal of Parallel and Distributed Computing, 62(3):427–445, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1801;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1801/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1801/ref.

Hameenanttila:1996:FHN

[HGCC96] Tom Hameenanttila, Xin-Li Guan, Jo Dale Carothers, and Jian-Xin Chen. The flexible hypercube: a new fault-tolerant architecture for parallel computing. Journal of Parallel and Distributed Computing, 37(2):213–220, September 15, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0122/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0122/production/pdf.

Hamdi:1997:RFE

[HH97] Mounir Hamdi and Richard W. Hall. RCC-Full: An effective network for parallel computations. Journal of Parallel and Distributed Computing, 41(2):139–155, March 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1295/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1295/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1295/production/ref.

He:2001:CEB

[HH01] Xin He and Chun-Hsi Huang. Communication efficient BSP algorithm for all nearest smaller values problem. Journal of Parallel and Distributed Computing, 61(10):1425–1438, October 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1741;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1741/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1741/ref.

Ho:1998:EPS

[HHC98] Chin-Wen Ho, Sun-Yuan Hsieh, and Gen-Huey Chen. An efficient parallel strategy for computing K-terminal reliability and finding most vital edges in 2-trees and partial 2-trees. Journal of Parallel and Distributed Computing, 51(2):89–113, June 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1454/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1454/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1454/production/ref.

Hall:1996:ICF

[HHKT96] Mary W. Hall, Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Interprocedural compilation of Fortran D. Journal of Parallel and Distributed Computing, 38(2):114–129, November 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0135/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0135/production/pdf.

Hambrusch:1994:PAG

[HHM94] Susanne Hambrusch, Xin He, and Russ Miller. Parallel algorithms for gray-scale digitized picture component labeling on a mesh-connected computer. Journal of Parallel and Distributed Computing, 20(1):56–68, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1005/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1005/production/pdf.

Han:1994:BFB

[HIKM94] Y. J. Han, Y. Igarashi, K. Kanai, and K. Miura. Broadcasting in faulty binary jumping networks. Journal of Parallel and Distributed Computing, 23(3):462–467, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1157/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1157/production/pdf.

Hac:1990:SSL

[HJ90a] A. Hac and T. J. Johnson. Sensitivity study of the load balancing algorithm in a distributed system. Journal of Parallel and Distributed Computing, 10(1):85–89, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Harper:1990:ERB

[HJ90b] D. T. Harper, III and J. R. Jump. Evaluation of reduced bandwidth multistage networks. Journal of Parallel and Distributed Computing, 9(3):304–311, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ho:1990:EMB

[HJ90c] Ching-Tien Ho and S. Lennart Johnsson. Embedding meshes in Boolean cubes by graph decomposition. Journal of Parallel and Distributed Computing, 8(4):325–339, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Helman:2001:PCS

[HJ01] David R. Helman and Joseph JáJá. Prefix computations on symmetric multiprocessors. Journal of Parallel and Distributed Computing, 61(2):265–278, February 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1678;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1678/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1678/ref.

Han:2007:CWA

[HJ07] Bo Han and Weijia Jia. Clustering wireless ad hoc networks with weakly connected dominating set. Journal of Parallel and Distributed Computing, 67(6):727–737, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hu:2001:PDR

[HJD^{+}01] X.-D. Hu, X.-H. Jia, D.-Z. Du, D.-Y. Li, and H.-J. Huang. Placement of data replicas for optimal data availability in ring networks. Journal of Parallel and Distributed Computing, 61(10):1412–1424, October 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1713;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1713/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1713/ref.

Hu:2001:MRM

[HJDH01] Xiao-Dong Hu, Xiao-Hua Jia, Ding-Zhu Du, and Frank K. Hwang. Monotone routing in multirate rearrangeable Clos networks. Journal of Parallel and Distributed Computing, 61(9):1382–1388, September 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1755;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1755/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1755/ref.

Hambrusch:1996:PMC

[HK96] Susanne E. Hambrusch and Ashfaq A. Khokhar. C^{3}: a parallel model for coarse-grained machines. Journal of Parallel and Distributed Computing, 32(2):139–154, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0010/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0010/production/pdf.

Hershberger:2001:DMR

[HK01] Daryl E. Hershberger and Hillol Kargupta. Distributed multivariate regression using wavelet-based collective data mining. Journal of Parallel and Distributed Computing, 61(3):372–400, March 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1694;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1694/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1694/ref.

Hsiao:2004:TCA

[HK04] Hung-Chang Hsiao and C.-T. Chung-Ta King. Tornado: a capability-aware peer-to-peer storage overlay. Journal of Parallel and Distributed Computing, 64(6):747–758, June 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hoffmann:2005:TSA

[HK05] Martin Hoffmann and Winfried E. Kühnhauser. Towards a structure-aware failure semantics for streaming media communication models. Journal of Parallel and Distributed Computing, 65(9):1047–1056, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Higham:2008:ISC

[HK08] Lisa Higham and Jalal Kawash. Implementing sequentially consistent programs on processor consistent platforms. Journal of Parallel and Distributed Computing, 68(4):488–500, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Heckmann:1998:OEC

[HKMU98] Ralf Heckmann, Ralf Klasing, Burkhard Monien, and Walter Unger. Optimal embedding of complete binary trees into lines and grids. Journal of Parallel and Distributed Computing, 49(1):40–56, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1424/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1424/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1424/production/ref.

Harvey:1991:ETL

[HKT^{+}91] Wilson Harvey, Dirk Kalp, Milind Tambe, David McKeown, and Allen Newell. The effectiveness of task-level parallelism for production systems. Journal of Parallel and Distributed Computing, 13(4):395–411, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hiranandani:1994:ECO

[HKT94] Seema Hiranandani, Ken Kennedy, and Chau-Wen W. Tseng. Evaluating compiler optimizations for Fortran D. Journal of Parallel and Distributed Computing, 21(1):27–45, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1040/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1040/production/
pdf.

Hsiao:2005:TMD

[HKW05] Hung-Chang Hsiao, Chung-Ta King, and Chia-Wei Wang. TY-PHOON: mobile distributed hash tables. Journal of Parallel and Distributed Computing, 65(2):191–206, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:2007:DMR

[HL07] Tzu-Lun Huang and D. T. Lee. A distributed multicast routing algorithm for real-time applications in wide area networks. Journal of Parallel and Distributed Computing, 67(5):516–530, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hu:2000:ONS

[HLCZ00] Y. Charlie Hu, Honghui Lu, Alan L. Cox, and Willy Zwaenepoel. OpenMP for networks of SMPs. Journal of Parallel and Distributed Computing, 60(12):1512–1530, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1658;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1658/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1658/ref.

Hwang:1998:FCA

[HLJ98] Gwan-Hwan Hwang, Jenq Kuen Lee, and Roy Dz-Ching Ju. A function-composition approach to synthesize Fortran 90 array operations. Journal of Parallel and Distributed Computing, 54(1):1–47, October 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1481/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1481/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1481/production/ref.

Hwang:2001:AOS

[HLJ01] Gwan-Hwan Hwang, Jenq Kuen Lee, and Roy Dz-Ching Ju. Array operation synthesis to optimize HPF programs on distributed memory machines. Journal of Parallel and Distributed Computing, 61(4):467–500, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1680;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1680/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1680/ref.

Hsieh:1995:PMS

[HLL^{+}95] Jenwei Hsieh, Mengjou Lin, Jonathan C. L. Liu, David H. C. Du, and Thomas M. Ruwart. Performance of a mass storage system for video-on-demand. Journal of Parallel and Distributed Computing, 30(2):147–167, November 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1134/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1134/production/pdf.

Hua:1995:OAL

[HlLLY95] Kien A. Hua, Yu lung L. Lo, and Honesty C. Young. Optimizer-assisted load balancing techniques for multicomputer database management systems. Journal of Parallel and Distributed Computing, 25(1):42–57, February 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1028/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1028/production/pdf.

Hosseini:1990:AGC

[HLM^{+}90] S. H. Hosseini, B. Litow, M. Malkawi, J. McPherson, and K. Vairavan. Analysis of a graph coloring based distributed load balancing algorithm. Journal of Parallel and Distributed Computing, 10(2):160–166, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hanlon:2003:LSF

[HLS03] Philip Hanlon, William Andrew Lorenz, and Dave Strenski. Least-squares fit of genomic data by sums of epistatic effects. Journal of Parallel and Distributed Computing, 63(7–8):683– 691, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Helmbold:1996:TRC

[HM96] D. P. Helmbold and C. E. McDowell. A taxonomy of race conditions. Journal of Parallel and Distributed Computing, 33(2):159–164, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0034/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0034/production/pdf.

Huang:1999:BBA

[HM99] Yuguang Huang and W. F. McColl. A BSP Bareiss algorithm for Toeplitz systems. Journal of Parallel and Distributed Computing, 56(2):99–121, February 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1509/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1509/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1509/production/ref.

Heun:2001:ODE

[HM01] Volker Heun and Ernst W. Mayr. Optimal dynamic embeddings of complete binary trees into hypercubes. Journal of Parallel and Distributed Computing, 61(8):1110–1125, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1734;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1734/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1734/ref.

Haghighi:2006:PPQ

[HM06] Aliakbar Montazer Haghighi and Dimitar P. Mishev. A parallel priority queueing system with finite buffers. Journal of Parallel and Distributed Computing, 66(3):379–392, March 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hart:2007:PMR

[HMBW07] Thomas E. Hart, Paul E. McKenney, Angela Demke Brown, and Jonathan Walpole. Performance of memory reclamation for lockless synchronization. Journal of Parallel and Distributed Computing, 67(12):1270–1285, December 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Han:2007:AAI

[HMV07] Qi Han, Sharad Mehrotra, and Nalini Venkatasubramanian. Application-aware integration of data collection and power management in wireless sensor networks. Journal of Parallel and Distributed Computing, 67(9):992–1006, September 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Horiguchi:1991:PEP

[HN91] Susumu Horiguchi and Takeo Nakada. Performance evaluation of parallel fast Fourier transform on a multiprocessor workstation. Journal of Parallel and Distributed Computing, 12(2):158– 163, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Howell:2002:FSS

[HNM02] Rodney R. Howell, Mikhail Nesterenko, and Masaaki Mizuno. Finite-state self-stabilizing protocols in message-passing systems. Journal of Parallel and Distributed Computing, 62 (5):792–817, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2001.1825;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2001.
1825/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2001.1825/ref.

Hashemi-Najafabadi:2007:MPM

[HNSA07] H. Hashemi-Najafabadi and H. Sarbazi-Azad. Mathematical performance modelling of adaptive wormhole routing in optoelectronic hypercubes. Journal of Parallel and Distributed Computing, 67(9):967–980, September 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ho:1991:OBS

[Ho91] Ching-Tien Ho. Optimal broadcasting on SIMD hypercubes without indirect addressing capability. Journal of Parallel and Distributed Computing, 13(2):246–255, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:1994:ELB

[HO94] G. Huang and W. Ongsakul. An efficient load-balancing processor scheduling algorithm for parallelization of Gauss– Seidel type algorithms. Journal of Parallel and Distributed Computing, 22(2):350–358, August 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1095/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1095/production/pdf.

He:2009:SAA

[HOE^{+}09] Xubin He, Li Ou, Christian Engelmann, Xin Chen, and Stephen L. Scott. Symmetric active/active metadata service for high availability parallel file systems. Journal of Parallel and Distributed Computing, 69(12):961–973, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hohberg:1990:HFB

[Hoh90] W. Hohberg. How to find biconnected components in distributed networks. Journal of Parallel and Distributed Computing, 9(4):374–386, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Heydemann:1994:EHG

[HOS94] M. C. Heydemann, J. Opatrny, and D. Sotteau. Embeddings of hypercubes and grids into de Bruijn graphs. Journal of Parallel and Distributed Computing, 23(1):104–111, October 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1124/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1124/production/pdf.

Hernandez-Orallo:2009:PBS

[HOVC09] Enrique Hernández-Orallo and Joan Vila-Carbó. A proactive backup scheme for reliable real-time transmission. Journal of Parallel and Distributed Computing, 69(5):482–490, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hendrickson:1995:PMB

[HP95] Bruce Hendrickson and Steve Plimpton. Parallel many-body simulations without all-to-all communication. Journal of Parallel and Distributed Computing, 27(1):15–25, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1068/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1068/production/pdf.

Ha:1997:SDC

[HP97a] Joon-Ho Ha and Timothy Mark Pinkston. SPEED DMON: Cache coherence on an optical multichannel interconnect architecture. Journal of Parallel and Distributed Computing, 41(1):78–91, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1287/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1287/production/pdf.

Homer:1997:DPP

[HP97b] Steven Homer and Marcus Peinado. Design and performance of parallel and distributed approximation algorithms for Maxcut. Journal of Parallel and Distributed Computing, 46(1):48–61, October 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1381/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1381/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1381/production/ref.

Ha:2000:NTB

[HP00] Joon-Ho Ha and Timothy Mark Pinkston. A new token-based channel access protocol for wavelength division multiplexed multiprocessor interconnects. Journal of Parallel and Distributed Computing, 60(2):169–188, February 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1599/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1599/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1599/production/ref.

Hong:2006:MLD

[HP06] Bo Hong and Viktor K. Prasanna. Maximum lifetime data sensing and extraction in energy constrained networked sensor systems. Journal of Parallel and Distributed Computing, 66 (4):566–577, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hurson:1991:RMB

[HPSM91] A. R. Hurson, S. H. Pakzad, D. B. Shin, and L. L. Miller. A reconfigurable MIMD back-end database machine. Journal of Parallel and Distributed Computing, 11(1):37–50, January 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hands:1997:PSB

[HPT^{+}97] M. A. Hands, W. Peiffer, H. Thienpont, A. Kirk, T. J. Hall, D. Pignon, and P. Parmiter. Proposal for stochastic bit stream processing using optoelectronic smart pixels: a neural network architectural case study. Journal of Parallel and Distributed Computing, 41(1):92–108, February 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1283/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1283/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1283/production/ref.

Hoepman:2002:SSW

[HPT02] Jaap-Henk Hoepman, Marina Papatriantafilou, and Philippas Tsigas. Self-stabilization of wait-free shared memory objects. Journal of Parallel and Distributed Computing, 62(5):818–842, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1829;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1829/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.

2001.1829/ref;   http://www.idealibrary.com/links/doi/
10.1006/jpdc.2001.1834;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.2001.1834/pdf;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1834/ref.

Ha:2007:STR

[HPT07] Phuong Hoai Ha, Marina Papatriantafilou, and Philippas Tsigas. Self-tuning reactive diffracting trees. Journal of Parallel and Distributed Computing, 67(6):674–694, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hamdi:1999:CES

[HQPT99] Mounir Hamdi, Chunming Qiao, Yi Pan, and J. Tong. Communication-efficient sorting algorithms on reconfigurable array of processors with slotted optical buses. Journal of Parallel and Distributed Computing, 57(2):166–187, May 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1525/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1525/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1525/production/ref.

Hung:1989:PPL

[HR89] Yubin Hung and Azriel Rosenfeld. Parallel processing of linear quadtrees on a mesh-connected computer. Journal of Parallel and Distributed Computing, 7(1):1–27, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hung:1990:PBC

[HR90] Yubin Hung and A. Rosenfeld. Processing border codes on a mesh-connected computer. Journal of Parallel and Distributed Computing, 10(1):18–28, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Heywood:1992:PHMb

[HR92a] T. Heywood and S. Ranka. A practical hierarchical model of parallel computation. II. binary tree and FFT algorithms. Journal of Parallel and Distributed Computing, 16(3):233–249, November 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Heywood:1992:PHMa

[HR92b] Todd Heywood and Sanjay Ranka. A practical hierarchical model of parallel computation. I. the model. Journal of Parallel and Distributed Computing, 16(3):212–232, November 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

HarEl:2000:JCB

[HR00] Zvi Har’El and Zvi Rosberg. Java class broker — a seamless bridge from local to distributed programming. Journal of Parallel and Distributed Computing, 60(10):1223–1237, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1651;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1651/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1651/ref.

Hacker:2009:ACF

[HRC09] Thomas J. Hacker, Fabian Romero, and Christopher D. Carothers. An analysis of clustered failures on large supercomputing systems. Journal of Parallel and Distributed Computing, 69(7):652–665, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ho:1994:EAG

[HRJ94] Ching-Tien T. Ho, M. T. Raghunath, and S. Lennart Johnsson. An efficient algorithm for Gray-to-binary permutation on hypercubes. Journal of Parallel and Distributed Computing, 20(1):114–120, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1012/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1012/production/pdf.

Hillyer:1986:EOP

[HS86] Bruce K. Hillyer and David Elliot Shaw. Execution of OPS5 production systems on a massively parallel machine. Journal of Parallel and Distributed Computing, 3(2):236–268, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:1993:CFH

[HS93] C. H. Huang and P. Sadayappan. Communication-free hyperplane partitioning of nested loops. Journal of Parallel and Distributed Computing, 19(2):90–102, October 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1094/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1094/production/pdf.

Higham:1994:MBT

[HS94a] L. Higham and E. Schenk. Maintaining B-trees on an EREW PRAM. Journal of Parallel and Distributed Computing, 22 (2):329–335, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1994.1092/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1092/production/pdf.

Homer:1994:SPC

[HS94b] P. T. Homer and R. D. Schlichting. A software platform for constructing scientific applications from heterogeneous resources. Journal of Parallel and Distributed Computing, 21(3):301–315, June 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1061/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1061/production/pdf.

Hou:1997:IDL

[HS97] Chao-Ju Hou and Kang G. Shin. Implementation of decentralized load sharing in networked workstations using the Condor package. Journal of Parallel and Distributed Computing, 40(2):173–184, February 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1244/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1244/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1244/production/ref.

Harris:2000:LOO

[HS00] Jacob Harris and Vivek Sarkar. Lightweight object-oriented shared variables for cluster computing in Java. Journal of Parallel and Distributed Computing, 60(10):1238–1259, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1652;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1652/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1652/ref.

Hwang:2003:IPP

[HS03] Yuan-Shin Hwang and Joel H. Saltz. Identifying parallelism in programs with cyclic graphs. Journal of Parallel and Distributed Computing, 63(3):337–355, March 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Harutyunyan:2006:EHB

[HS06] Hovhannes A. Harutyunyan and Bin Shao. An efficient heuristic for broadcasting in networks. Journal of Parallel and Distributed Computing, 66(1):68–76, January 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hsieh:2004:EPS

[Hsi04] Sun-Yuan Hsieh. An efficient parallel strategy for the two-fixed-endpoint Hamiltonian path problem on distance-hereditary graphs. Journal of Parallel and Distributed Computing, 64 (5):662–685, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hinkle:1987:NLP

[HSJP87] Eric B. Hinkle, Jorge L. C. Sanz, Anil K. Jain, and Dragutin Petkovic. P^{3}E: New life for projection-based image processing. Journal of Parallel and Distributed Computing, 4(1):45–78, February 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Heymann:2004:ERM

[HSLL04] E. Heymann, M. A. Senar, E. Luque, and M. Livny. Efficient resource management applied to master–worker applications. Journal of Parallel and Distributed Computing, 64(6):767–773, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hiranandani:1991:PHC

[HSMB91] Seema Hiranandani, Joel Saltz, Piyush Mehrotra, and Harry Berryman. Performance of hashed cache data migration schemes on multicomputers. Journal of Parallel and Distributed Computing, 12(4):415–422, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

He:2007:PGA

[HSSM07] Hongmei He, Ondrej Sýkora, Ana Salagean, and Erkki Mäkinen. Parallelisation of genetic algorithms for the 2-page crossing number problem. Journal of Parallel and Distributed Computing, 67(2):229–241, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Houle:2004:DEA

[HSW04] Michael E. Houle, Antonios Symvonis, and D. R. David R. Wood. Dimension-exchange algorithms for token distribution on tree-connected architectures. Journal of Parallel and Distributed Computing, 64(5):591–605, May 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hillis:1990:ESH

[HT90] W. Daniel Hillis and Washington Taylor, IV. Exploiting symmetry in high-dimensional finite difference calculations. Journal of Parallel and Distributed Computing, 8(1):77–79, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Herlihy:2006:RSN

[HT06] Maurice Herlihy and Srikanta Tirthapura. Randomized smoothing networks. Journal of Parallel and Distributed Computing, 66(5):626–632, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hribar:1998:TDP

[HTB98] Michelle R. Hribar, Valerie E. Taylor, and David E. Boyce. Termination detection for parallel shortest path algorithms. Journal of Parallel and Distributed Computing, 55(2):153–165, December 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1502/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1502/

production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1502/production/ref.

Huang:2002:FTH

[HTHH02] Wen-Tzeng Huang, Jimmy J. M. Tan, Chun-Nan Hung, and Lih-Hsing Hsu. Fault-tolerant Hamiltonicity of twisted cubes. Journal of Parallel and Distributed Computing, 62(4):591–604, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1813;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1813/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1813/ref.

Hua:1999:PLB

[HTL99] Kien A. Hua, Wallapak Tavanapong, and Yu-Lung Lo. Performance of load balancing techniques for join operations in shared-noting database management systems. Journal of Parallel and Distributed Computing, 56(1):17–46, January 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1507/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1507/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1507/production/ref.

Haldar:1995:BOC

[HV95] S. Haldar and K. Vidyasankar. Buffer-optimal constructions of 1-Writer multireader multivalued atomic shared variables. Journal of Parallel and Distributed Computing, 31(2):174–180, December 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1156/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1156/
production/pdf.

Haldar:2009:SOB

[HV09] Sibsankar Haldar and K. Vidyasankar. On space-optimality of buffer-based conflict-free constructions of 1-writer 1-reader multivalued atomic variables from safe bits. Journal of Parallel and Distributed Computing, 69(1):20–27, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Habetha:2003:ASP

[HW03] Jörg Habetha and Jens Wiegert. Analytical and simulative performance evaluation of cluster-based multihop ad hoc networks. Journal of Parallel and Distributed Computing, 63(2):166–181, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hwang:1997:EPA

[Hwa97] Injae Hwang. An efficient processor allocation algorithm using two- dimensional packing. Journal of Parallel and Distributed Computing, 42(1):75–81, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1306/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1306/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1306/production/ref.

Huang:2008:TCT

[HWC08] Jane-Hwa Huang, Li-Chun Wang, and Chung-Ju Chang. Throughput-coverage tradeoff in a scalable wireless mesh network. Journal of Parallel and Distributed Computing, 68(3): 278–290, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:1996:PLB

[HWW96] Kuo-Chan Huang, Feng-Jian Wang, and Pei-Chi Wu. Parallelizing a level 3 BLAS library for LAN-Connected workstations. Journal of Parallel and Distributed Computing, 38(1):28–36, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1996.0126/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.0126/production/pdf.

Hsu:2008:SDG

[HWWH08] Chun-Yen Hsu, Jean-Lien C. Wu, Shun-Te Wang, and Chi-Yao Hong. Survivable and delay-guaranteed backbone wireless mesh network design. Journal of Parallel and Distributed Computing, 68(3):306–320, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

He:2004:SSI

[HZY04] Xubin He, Ming Zhang, and Qing (Ken) Yang. STICS: SCSIto-IP cache for storage area networks. Journal of Parallel and Distributed Computing, 64(9):1069–1085, September 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ieumwananonthachai:1992:IPM

[IAS^{+}92] Arthur Ieumwananonthachai, Akiko N. Aizawa, Steven R. Schwartz, Benjamin W. Wah, and Jerry C. Yan. Intelligent process mapping through systematic improvement of heuristics. Journal of Parallel and Distributed Computing, 15(2):118–142, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iyengar:2004:SII

[IB04] S. Sitharama Iyengar and R. R. Brooks. Special issue introduction — the road map for distributed sensor networks in the context of computing and communication. Journal of Parallel and Distributed Computing, 64(7):785–787, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ibrahim:2008:FGP

[IBP08] Khaled Z. Ibrahim, François Bodin, and Olivier Pène. Finegrained parallelization of lattice QCD kernel routine on GPUs. Journal of Parallel and Distributed Computing, 68(10):1350– 1359, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iqbal:2005:PAD

[IC05] Saeed Iqbal and Graham F. Carey. Performance analysis of dynamic load balancing algorithms with variable number of processors. Journal of Parallel and Distributed Computing, 65(8): 934–948, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Izaguirre:2005:PMS

[IHM05] Jesús A. Izaguirre, Scott S. Hampton, and Thierry Matthey. Parallel multigrid summation for the N-body problem. Journal of Parallel and Distributed Computing, 65(8):949–962, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ibarra:1993:QBA

[IK93] Oscar H. Ibarra and Myung Hee Kim. Quadtree building algorithms on an SIMD hypercube. Journal of Parallel and Distributed Computing, 18(1):71–76, May 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1045/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1045/production/pdf.

Ibarra:1994:FPA

[IK94] Oscar H. Ibarra and Myung Hee Kim. Fast parallel algorithms for solving triangular systems of linear equations on the hypercube. Journal of Parallel and Distributed Computing, 20(3):303–316, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1029/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1029/production/pdf.

Ibrahim:1987:LLI

[IKS87] Hussein A. H. Ibrahim, John R. Kender, and David Elliot Shaw. Low-level image analysis tasks on fine-grained tree-structured SIMD machines. Journal of Parallel and Distributed Computing, 4(6):546–574, December 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iannello:1994:CWA

[IM94] Giulio Iannello and Nicola Mazzocca. Communication workload analysis for symmetrical concurrent systems. Journal of Parallel and Distributed Computing, 20(2):224–235, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1021/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1021/production/
pdf.

Iwama:2000:ORA

[IM00] Kazuo Iwama and Eiji Miyano. Oblivious routing algorithms on the mesh of buses. Journal of Parallel and Distributed Computing, 60(2):137–149, February 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1597/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1597/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1597/production/ref.

Ibarra:1985:SRC

[IPK85] Oscar H. Ibarra, Michael A. Palis, and Sam M. Kim. Some results concerning linear iterative (systolic) arrays. Journal of Parallel and Distributed Computing, 2(2):182–218, May 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iqbal:1992:EAD

[Iqb92] Mohammad Ashraf Iqbal. Efficient algorithms for dilated mappings of binary trees. Journal of Parallel and Distributed Computing, 15(1):75–78, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Irwin:1988:SIP

[Irw88] M. J. Irwin. Special issue on parallelism in computer arithmetic. Journal of Parallel and Distributed Computing, 5(3):205–208, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Iyengar:2006:ENL

[IS06] M. Sriram Iyengar and Mukesh Singhal. Effect of network latency on load sharing in distributed systems. Journal of Parallel and Distributed Computing, 66(6):839–853, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Imani:2007:CIP

[ISAZ07] N. Imani, H. Sarbazi-Azad, and A. Y. Zomaya. Capturing an intruder in product networks. Journal of Parallel and Distributed Computing, 67(9):1018–1028, September 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Islam:1997:CMP

[Isl97] Nayeem Islam. Customized message passing. Journal of Parallel and Distributed Computing, 41(2):205–224, March 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1294/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1294/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1294/production/ref.

Izumi:2007:ATC

[ISM07] Taisuke Izumi, Akinori Saitoh, and Toshimitsu Masuzawa. Adaptive timeliness of consensus in presence of crash and timing faults. Journal of Parallel and Distributed Computing, 67 (6):648–658, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Itzkovitz:1999:TID

[ISZBM99] Ayal Itzkovitz, Assaf Schuster, and Oren Zeev-Ben-Mordehai. Toward integration of data race detection in DSM systems. Journal of Parallel and Distributed Computing, 59(2):180–203, November 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1574/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1574/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1574/production/ref.

Irony:2004:CLB

[ITT04] Dror Irony, Sivan Toledo, and Alexander Tiskin. Communication lower bounds for distributed-memory matrix multiplication. Journal of Parallel and Distributed Computing, 64(9): 1017–1026, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Irakliotis:1997:OME

[IWM97] Leo J. Irakliotis, Carl W. Wilmsen, and Pericles A. Mitkas. The optical memory-electric computer interface as a parallel processing architecture. Journal of Parallel and Distributed Computing, 41(1):67–77, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1286/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1286/production/pdf.

Ibarra:1995:OSP

[IZ95] O. H. Ibarra and Qi Zheng. An optimal shortest path parallel algorithm for permutation graphs. Journal of Parallel and Distributed Computing, 24(1):94–99, January 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1009/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1009/production/pdf.

Johnson:1993:PED

[JB93] Darin Johnson and Francine Berman. Performance of the efficient data-driven evaluation scheme. Journal of Parallel and Distributed Computing, 18(3):340–346, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1069/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1069/production/pdf.

Joy:1998:HDR

[JB98] Joseph M. Joy and R. Daniel Bergeron. The hypercube as a dynamically reconfigurable processor mesh. Journal of Parallel and Distributed Computing, 48(1):130–142, January 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1402/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1402/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1402/production/ref.

Jaramillo-Botero:2002:UFM

[JBL02] Andrés Jaramillo-Botero and Alfons Crespo I. Lorente. A unified formulation for massively parallel rigid multibody dynamics of O(log_{2}n) computational complexity. Journal of Parallel and Distributed Computing, 62(6):1001–1020, June 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1820;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1820/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1820/ref.

Jiang:1991:MMV

[JBM91] Hong Jiang, Laxmi N. Bhuyan, and Jogesh K. Muppala. MVAMIN: mean value analysis algorithms for multistage interconnection networks. Journal of Parallel and Distributed Computing, 12(3):189–201, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jayanti:2000:AOS

[JBP00] Prasad Jayanti, James Burns, and Gary Peterson. Almost optimal single reader, single writer atomic register. Journal of Parallel and Distributed Computing, 60(2):150–168, February 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1505/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1505/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1505/production/ref.

Jin:2005:ENA

[JBY^{+}05] Hyun-Wook Jin, Pavan Balaji, Chuck Yoo, Jin-Young Choi, and Dhabaleswar K. Panda. Exploiting NIC architectural support for enhancing IP-based protocols on high-performance networks. Journal of Parallel and Distributed Computing, 65(11): 1348–1365, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Janssens:1995:ECR

[JF95] Bob Janssens and W. Kent Fuchs. Ensuring correct rollback recovery in distributed shared memory systems. Journal of Parallel and Distributed Computing, 29(2):211–218, September 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1118/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1118/production/
pdf.

James:1992:DLC

[JH92a] Mark James and Doan Hoang. Design of low-cost, real-time simulation systems for large neural networks. Journal of Parallel and Distributed Computing, 14(3):221–235, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See erratum [Ano92c].

Johnsson:1992:GSP

[JH92b] S. Lennart Johnsson and Ching-Tien Ho. Generalized shuffle permutations on Boolean cubes. Journal of Parallel and Distributed Computing, 16(1):1–14, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnsson:1994:BCE

[JH94] S. Lennart Johnsson and Ching-Tien T. Ho. Binary cube emulation of butterfly networks encoded by Gray code. Journal of Parallel and Distributed Computing, 20(3):261–279, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1026/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1026/production/
pdf.

Jiang:1999:PAD

[Jia99] J. Jiang. A parallel algorithm for 4 × 4 DCT. Journal of Parallel and Distributed Computing, 57(2):257–269, May 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1532/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1532/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1532/production/ref.

Jayaraman:1989:SIR

[JK89] Bharat Jayaraman and Robert M. Keller. Semantics and implementation of resource expressions. Journal of Parallel and Distributed Computing, 7(1):40–63, August 1989. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jiang:2000:PSA

[JK00] Zhimei Jiang and Leonard Kleinrock. A packet selection algorithm for adaptive transmission of smoothed video over a wireless channel. Journal of Parallel and Distributed Computing, 60(4):494–509, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1616;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1616/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1616/ref.

Jan:2005:CLB

[JL05] Gene Eu Jan and Ming-Bo Lin. Concentration, load balancing, partial permutation routing, and superconcentration on cube-connected cycles parallel computers. Journal of Parallel and Distributed Computing, 65(12):1471–1482, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jiang:2008:OCM

[JLM08] Qiangfeng Jiang, Yi Luo, and D. Manivannan. An optimistic checkpointing and message logging approach for consistent global checkpoint collection in distributed systems. Journal of Parallel and Distributed Computing, 68(12):1575–1589, December 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:1997:LCH

[JLRA97] Donald Johnson, David Lilja, John Riedl, and James Anderson. Low-cost, high-performance barrier synchronization on networks of workstations. Journal of Parallel and Distributed Computing, 40(1):131–137, January 10, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1273/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1273/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1273/production/ref.

Jaen-Martinez:2000:JME

[JM00] Javier Jaen-Martinez. The Java Management Extensions (JMX): Is your cluster ready for evolution? Journal of Parallel and Distributed Computing, 60(10):1341–1353, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1657;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1657/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1657/ref.

Jamieson:1986:FAS

[JMS86] Leah H. Jamieson, Philip T. Mueller, Jr., and Howard Jay Siegel. FFT algorithms for SIMD parallel processing systems. Journal of Parallel and Distributed Computing, 3(1):48– 71, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:1996:CFL

[JNW96] Theodore Johnson and Richard Newman-Wolfe. A comparison of fast and low overhead distributed priority locks. Journal of Parallel and Distributed Computing, 32(1):74–89, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0006/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0006/production/
pdf.

Johnsson:1987:CEB

[Joh87] S. Lennart Johnsson. Communication efficient basic linear algebra computations on hypercube architectures. Journal of Parallel and Distributed Computing, 4(2):133–172, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:1989:NNA

[Joh89] James L. Johnson. A neural network approach to the 3satisfiability problem. Journal of Parallel and Distributed Computing, 6(2):435–449, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnsson:1991:PMD

[Joh91] S. Lennart Johnsson. Performance modeling of distributed memory architectures. Journal of Parallel and Distributed Computing, 12(4):300–312, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:1994:HCP

[Joh94] T. Johnson. A highly concurrent priority queue. Journal of Parallel and Distributed Computing, 22(2):367–373, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1097/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1097/production/
pdf.

Jang:1995:OSA

[JP95] Ju-Wook Jang and Viktor K. Prasanna. An optimal sorting algorithm on reconfigurable mesh. Journal of Parallel and Distributed Computing, 25(1):31–41, February 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1027/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1027/production/pdf.

Jiang:2009:SPP

[JP09] Weirong Jiang and Viktor K. Prasanna. Sequence-preserving parallel IP lookup using multiple SRAM-based pipelines. Journal of Parallel and Distributed Computing, 69(9):778–789, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

JaJa:1992:LBR

[JR92] Joseph JáJá and Kwan Woo Ryu. Load balancing and routing on the hypercube and related networks. Journal of Parallel and Distributed Computing, 14(4):431–435, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jain:1995:ILB

[JR95] Kamal Kumar Jain and V. Rajaraman. Improved lower bounds on time and processors for scheduling precedence graphs on multicomputer systems. Journal of Parallel and Distributed Computing, 28(1):101–108, July 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1092/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1092/production/pdf.

Joshi:1999:SPM

[JRR99] Rushikesh K. Joshi, O. Ramakrishna, and D. Janaki Ram. ShadowObjects: a programming model for service replication in distributed object systems. Journal of Parallel and Distributed Computing, 59(1):1–12, October 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1548/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1548/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1548/production/ref.

Jegou:1986:DSP

[JS86] Y. Jegou and A. Seznec. Data synchronized pipeline architecture: pipelining in multiprocessing environments. Journal of Parallel and Distributed Computing, 3(4):508–526, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jenq:1994:RMA

[JS94] Jing-Fu F. Jenq and Sartaj Sahni. Reconfigurable mesh algorithms for the Hough transform. Journal of Parallel and Distributed Computing, 20(1):69–77, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1006/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1006/production/pdf.

Jadav:1995:TSH

[JSCB95] Divyesh Jadav, Chutimet Srinilta, Alok Choudhary, and P. Bruce Berra. Techniques for scheduling I/O in a high performance multimedia-on-demand server. Journal of Parallel and Distributed Computing, 30(2):190–203, November 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1137/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1137/production/pdf.

Jia:1994:PNU

[JSM94] Xiaohua H. Jia, Kentaro Shimizu, and Mamoru Maekawa. A parallel and nonblocking updating mechanism for replicated directory files in distributed systems. Journal of Parallel and Distributed Computing, 20(3):330–340, March 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1031/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1031/production/pdf.

Jungnitz:1992:ATC

[JSS92] Hauke Jungnitz, Beatriz Sanchez, and Manuel Silva. Approximate throughput computation of stochastic marked graphs. Journal of Parallel and Distributed Computing, 15(3):282–295, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jain:1992:SPO

[JSWB92] Ravi Jain, Kiran Somalwar, John Werth, and J. C. Browne. Scheduling parallel I/O operations in multiple bus systems. Journal of Parallel and Distributed Computing, 16(4):352–362, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jain:1988:ACC

[JT88] Bijendra N. Jain and Satish K. Tripathi. An analysis of cubeconnected cycles and circular shuffle networks for parallel computation. Journal of Parallel and Distributed Computing, 5(6): 741–754, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jin:2006:PCM

[JV06] Haoqiang Jin and Rob F. Van der Wijngaart. Performance characteristics of the multi-zone NAS parallel benchmarks. Journal of Parallel and Distributed Computing, 66(5):674–685, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Joshi:2009:ADS

[JV09] Yadnyesh Joshi and Sathish Vadhiyar. Analysis of DNA sequence transformations on grids. Journal of Parallel and Distributed Computing, 69(1):80–90, January 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Juang:1989:LBO

[JW89] Jie-Yong Juang and Benjamin W. Wah. Load balancing and ordered selections in a computer system with multiple contention buses. Journal of Parallel and Distributed Computing, 7(3): 391–415, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

JaJa:1994:SID

[JW94] Joseph JáJá and Pearl Y. Wang. Special issue on data parallel algorithms and programming: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 21(1):1–3, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jiang:2006:CVB

[JXW06] Xuxian Jiang, Dongyan Xu, and Yi-Min Wang. Collapsar: a VM-based honeyfarm and reverse honeyfarm architecture for network attack capture and detention. Journal of Parallel and Distributed Computing, 66(9):1165–1180, September 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Jin:2005:MDD

[JZ05] Dejiang Jin and Sotirios G. Ziavras. Modeling distributed data representation and its effect on parallel data accesses. Journal of Parallel and Distributed Computing, 65(10):1281–1289, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:2004:MEN

[JZK04] Michael Johnson, Albert Y. Zomaya, and Miro Kraetzl. Modeling external network behavior by using internal measurements. Journal of Parallel and Distributed Computing, 64(12):1345– 1359, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Korst:1989:COB

[KA89] Jan H. M. Korst and Emile H. L. Aarts. Combinatorial optimization on a Boltzmann machine. Journal of Parallel and Distributed Computing, 6(2):331–357, April 1989. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koc:1991:FAG

[KA91] Cetin K. Koc and Sarath N. Arachchige. A fast algorithm for Gaussian elimination over GF(2) and its implementation on the GAPP. Journal of Parallel and Distributed Computing, 13(1): 118–122, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kwok:1997:ESA

[KA97] Yu-Kwong Kwok and Ishfaq Ahmad. Efficient scheduling of arbitrary task graphs to multiprocessors using a parallel genetic algorithm. Journal of Parallel and Distributed Computing, 47(1):58–77, November 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1395/production;   http://www.idealibrary.

com/links/doi/10.1006/jpdc.1997.1395/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1395/production/ref.

Kwok:1999:BCT

[KA99] Yu-Kwong Kwok and Ishfaq Ahmad. Benchmarking and comparison of the task graph scheduling algorithms. Journal of Parallel and Distributed Computing, 59(3):381–422, December 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1578/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1578/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1578/production/ref.

Kang:2003:SSS

[KA03] Oh-Han Kang and Dharma P. Agrawal. Scalable scheduling for symmetric multiprocessors (SMP). Journal of Parallel and Distributed Computing, 63(3):273–285, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kepner:2004:M

[KA04] Jeremy Kepner and Stan Ahalt. MatlabMPI. Journal of Parallel and Distributed Computing, 64(8):997–1005, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kwok:2005:MTS

[KA05] Yu-Kwong Kwok and Ishfaq Ahmad. On multiprocessor task scheduling using efficient state space search approaches. Journal of Parallel and Distributed Computing, 65(12):1515–1532, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khan:2008:CAT

[KA08] Samee Ullah Khan and Ishfaq Ahmad. Comparison and analysis of ten static heuristics-based Internet data replication techniques. Journal of Parallel and Distributed Computing, 68(2): 113–136, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kai:1992:EA

[Kai92] Hwang Kai. Editorial announcement. Journal of Parallel and Distributed Computing, 15(1):v–vi, May 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaludercic:2004:PLM

[Kal04] Boris Kaludercic. Parallelisation of the Lagrangian model in a mixed Eulerian–Lagrangian CFD algorithm. Journal of Parallel and Distributed Computing, 64(2):277–284, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ke:1994:SPA

[KAM94] Q. Ke, S. G. Akl, and H. Meijer. On some properties and algorithms for the star and pancake interconnection networks. Journal of Parallel and Distributed Computing, 22(1):16–25, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1066/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1066/production/
pdf.

Kandemir:2005:IWP

[Kan05] Mahmut Taylan Kandemir. Improving whole-program locality using intra-procedural and inter-procedural transformations. Journal of Parallel and Distributed Computing, 65(5):564–582, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:1990:LSM

[KAP90] Sukil Kim, Dharma P. Agrawal, and Robert J. Plemmons. Least-squares multiple updating algorithms on a hypercube. Journal of Parallel and Distributed Computing, 8(1):80–88, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kapralski:1993:NMG

[Kap93] Adam Kapralski. New methods for the generation of permutations, combinations, and other combinatorial objects in parallel. Journal of Parallel and Distributed Computing, 17(4):315–326, April 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1030/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1030/production/pdf.

Karonis:1992:TPP

[Kar92] Nicholas T. Karonis. Timing parallel programs that use message passing. Journal of Parallel and Distributed Computing, 14(1):29–36, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karabeg:1995:PPT

[Kar95] Dino Karabeg. Process partitioning through graph compaction. Journal of Parallel and Distributed Computing, 25(2):115–125, March 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1034/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1034/
production/pdf.

Karaata:2002:SAF

[Kar02] Mehmet Hakan Karaata. A stabilizing algorithm for finding biconnected components. Journal of Parallel and Distributed Computing, 62(5):982–999, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1833;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1833/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1833/ref.

Kim:2007:PDD

[KAS07] Jik-Soo Kim, Henrique Andrade, and Alan Sussman. Principles for designing data-/compute-intensive distributed applications and middleware systems for heterogeneous environments. Journal of Parallel and Distributed Computing, 67(7):755–771, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaufman:1994:PQA

[Kau94] L. Kaufman. A parallel QR algorithm for the symmetric tridiagonal eigenvalue problem. Journal of Parallel and Distributed Computing, 23(3):429–434, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1152/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1152/production/pdf.

Kavianpour:1993:NMS

[Kav93] A. Kavianpour. A new measure in system-level diagnosis of hypercubes. Journal of Parallel and Distributed Computing, 19(4):372–378, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1119/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1119/production/pdf.

Kim:1996:DEP

[KB96a] Junguk L. Kim and Geneva G. Belford. A distributed election protocol for unreliable networks. Journal of Parallel and Distributed Computing, 35(1):35–42, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0065/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0065/production/pdf.

Krueger:1996:EPP

[KB96b] Phillip Krueger and Davender Babbar. The effects of precedence and priority constraints on the performance of scan scheduling for hypercube multiprocessors. Journal of Parallel and Distributed Computing, 39(2):95–104, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0158/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0158/production/
pdf.

Kohn:2001:PSA

[KB01] Scott R. Kohn and Scott B. Baden. Parallel software abstractions for structured adaptive mesh methods. Journal of Parallel and Distributed Computing, 61(6):713–736, June 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1700;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1700/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1700/ref.

Kennedy:2001:TLS

[KBC^{+}01] Ken Kennedy, Bradley Broom, Keith Cooper, Jack Dongarra, Rob Fowler, Dennis Gannon, Lennart Johnsson, John Mellor-Crummey, and Linda Torczon. Telescoping languages: a strategy for automatic generation of scientific problem-solving systems from annotated libraries. Journal of Parallel and Distributed Computing, 61(12):1803–1826, December 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1724;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1724/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1724/ref;   http://
www.netlib.org/netlib/utk/people/JackDongarra/PAPERS/
Telescope.pdf.

Kuang:2005:PPO

[KBD05] Hairong Kuang, Lubomir F. Bic, and Michael B. Dillencourt. PODC: Paradigm-oriented distributed computing. Journal of Parallel and Distributed Computing, 65(4):506–518, April 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kyriakis-Bitzaros:1992:EDT

[KBG92] E. D. Kyriakis-Bitzaros and C. E. Goutis. An efficient decomposition technique for mapping nested loops with constant dependencies into regular processor arrays. Journal of Parallel and Distributed Computing, 16(3):258–264, November 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Krishnamoorthy:1994:SDS

[KC94] S. Krishnamoorthy and A. Choudhary. A scalable distributed shared memory architecture. Journal of Parallel and Distributed Computing, 22(3):547–554, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1110/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1110/production/pdf.

Kim:1995:NPU

[KC95] Jae H. Kim and Andrew A. Chien. Network performance under bimodal traffic loads. Journal of Parallel and Distributed Computing, 28(1):43–64, July 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1088/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1088/production/pdf.

Kim:1998:MGE

[KC98] Sook-Yeon Kim and Kyung-Yong Chwa. Multiple graph embeddings into a processor array with spanning buses. Journal of Parallel and Distributed Computing, 52(1):24–39, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1467/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1467/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1467/production/ref.

Karamcheti:1999:ASM

[KC99a] Vijay Karamcheti and Andrew A. Chien. Architectural support and mechanisms for object caching in dynamic multithreaded computations. Journal of Parallel and Distributed Computing, 58(2):260–300, August 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1555/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1555/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1555/production/ref.

Kwok:1999:PRS

[KC99b] S. H. Kwok and A. G. Constantinides. A parallel recursive shortest spanning tree algorithm for image segmentation in distributed computing environment. Journal of Parallel and Distributed Computing, 56(3):181–207, March 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1518/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1518/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1518/production/ref.

Kwok:2004:NFD

[KC04] Yu-Kwong Kwok and Lap-Sun Cheung. A new fuzzy-decision based load balancing system for distributed object computing. Journal of Parallel and Distributed Computing, 64(2):238–253, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2008:CDW

[KCD08] Jin-Ha Kim, Gyu Sang Choi, and Chita R. Das. Coscheduled distributed-web servers on system area network. Journal of Parallel and Distributed Computing, 68(8):1033–1043, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Keleher:1995:ESB

[KCDZ95] Pete Keleher, Alan L. Cox, Sandhya Dwarkadas, and Willy Zwaenepoel. An evaluation of software-based release consistent protocols. Journal of Parallel and Distributed Computing, 29(2):126–141, September 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1113/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1113/production/pdf.

Kandemir:1999:MBA

[KCRB99] Mahmut Kandemir, Alok Choudhary, J. Ramanujam, and Prith Banerjee. A matrix-based approach to global locality optimization. Journal of Parallel and Distributed Computing, 58(2):190–235, August 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1552/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1552/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1552/production/ref.

Kim:1999:CMD

[KCV99] Myungchul Kim, Samuel T. Chanson, and Son T. Vuong. Concurrency model for distributed systems. Journal of Parallel and Distributed Computing, 59(3):445–464, December 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1583/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1583/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1583/production/ref.

Koutsonikolas:2008:IAF

[KDH08] Dimitrios Koutsonikolas, Saumitra M. Das, and Y. Charlie Hu. An interference-aware fair scheduling for multicast in wireless mesh networks. Journal of Parallel and Distributed Computing, 68(3):372–386, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kotz:1993:CWP

[KE93] David Kotz and Carla Schlatter Ellis. Caching and writeback policies in parallel file systems. Journal of Parallel and Distributed Computing, 17(1–2):140–145, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1012/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1012/production/
pdf.

Kulasinghe:1995:AMP

[KEA95] Priyalal Kulasinghe and Ahmed El-Amawy. On achieving maximum performance in time-varying arrays. Journal of Parallel and Distributed Computing, 31(2):101–111, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1149/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1149/production/
pdf.

Keleher:2000:ISS

[Kel00] Peter J. Keleher. The impact of symmetry on software distributed shared memory. Journal of Parallel and Distributed Computing, 60(11):1388–1419, November 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1637;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1637/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1637/ref.

Kepner:2003:MTF

[Kep03] Jeremy Kepner. A multi-threaded fast convolver for dynamically parallel image filtering. Journal of Parallel and Distributed Computing, 63(3):360–372, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karaata:2001:MCL

[KER01] Mehmet Hakan Karaata and Hesham El-Rewini. On minimizing the cost of location management in mobile environments. Journal of Parallel and Distributed Computing, 61(7):950–966, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1722;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1722/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1722/ref.

Kamal:2004:OAA

[KERUM04] Ahmed Kamal, Hesham El-Rewini, and RazaRaza Ul-Mustafa. Optimal and approximate approaches for selecting proxy agents in mobile IP based network backbones. Journal of Parallel and Distributed Computing, 64(4):554–568, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kalyanaraman:2007:AGL

[KESA07] A. Kalyanaraman, S. J. Emrich, P. S. Schnable, and S. Aluru. Assembling genomes on large-scale parallel computers. Journal of Parallel and Distributed Computing, 67(12):1240–1255, December 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuo:1990:RCC

[KF90a] Sy-Yen Kuo and W. Kent Fuchs. Reconfigurable cubeconnected cycles architectures. Journal of Parallel and Distributed Computing, 9(1):1–10, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuszmaul:1990:NNA

[KF90b] B. C. Kuszmaul and J. Fried. NAP (no ALU processor): The great communicator. Journal of Parallel and Distributed Computing, 8(2):169–179, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kankanhalli:1995:APC

[KF95a] Mohan S. Kankanhalli and Wm. Randolph Franklin. Area and perimeter computation of the union of a set of isorectangles in parallel. Journal of Parallel and Distributed Computing, 27(2):107–117, June 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1076/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1076/production/pdf.

Khanna:1995:LTA

[KF95b] Sanjeev Khanna and W. Kent Fuchs. A linear time algorithm for sequential diagnosis in hypercubes. Journal of Parallel and Distributed Computing, 26(1):48–53, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1046/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1046/production/pdf.

Kumar:1994:ASP

[KG94] V. P. Kumar and A. Gupta. Analyzing scalability of parallel algorithms and architectures. Journal of Parallel and Distributed Computing, 22(3):379–391, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1098/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1098/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1099/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1099/production/pdf.

Kudlur:2004:PAM

[KG04] Manjunath Kudlur and R. Govindarajan. Performance analysis of methods that overcome false sharing effects in software DSMs. Journal of Parallel and Distributed Computing, 64(8): 887–907, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

King:1989:REB

[KGN89] Chung-Ta King, Thomas B. Gendreau, and Lionel M. Ni. Reliable election in broadcast networks. Journal of Parallel and Distributed Computing, 7(3):521–540, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:1994:SLB

[KGV94] V. Kumar, A. Y. Grama, and Nageshwara Rao Vempaty. Scalable load balancing techniques for parallel computers. Journal of Parallel and Distributed Computing, 22(1):60–79, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1070/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1070/production/
pdf.

Kung:1989:USA

[KH89] S. Y. Kung and J. N. Hwang. A unified systolic architecture for artificial neural networks. Journal of Parallel and Distributed Computing, 6(2):358–387, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khokhar:2003:TDD

[KHK03] Ashfaq A. Khokhar, Susanne E. Hambrusch, and Erturk Kocalar. Termination detection in data-driven parallel computations/applications. Journal of Parallel and Distributed Computing, 63(3):312–326, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaushik:1996:EIS

[KHS96] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Efficient index set generation for compiling HPF array statements on distributed-memory machines. Journal of Parallel and Distributed Computing, 38(2):237–247, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0144/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0144/production/pdf.

Kumar:1984:PEB

[KJ84] Manoj Kumar and J. R. Jump. Performance enhancement in buffered delta networks using crossbar switches and multiple links. Journal of Parallel and Distributed Computing, 1(1):81– 103, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaznachey:2003:NNB

[KJD03] Dmitri Kaznachey, Arun Jagota, and Sajal Das. Neural network-based heuristic algorithms for hypergraph coloring problems with applications. Journal of Parallel and Distributed Computing, 63(9):786–800, September 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:1986:TPD

[KK86] S. P. Kumar and J. S. Kowalik. Triangularization of a positive definite matrix on a parallel computer. Journal of Parallel and Distributed Computing, 3(4):450–460, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kirchner:1988:AAV

[KK88] R. Kirchner and U. Kulisch. Accurate arithmetic for vector processors. Journal of Parallel and Distributed Computing, 5 (3):250–270, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Konda:1995:SFD

[KK95] Venkat Konda and Anup Kumar. A systematic framework for the dependence cycle removal in practical loops. Journal of Parallel and Distributed Computing, 27(2):157–171, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1079/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1079/production/
pdf.

Karypis:1998:MWP

[KK98a] George Karypis and Vipin Kumar. Multilevel k-way partitioning scheme for irregular graphs. Journal of Parallel and Distributed Computing, 48(1):96–129, January 10, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1404/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1404/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1404/production/ref.

Karypis:1998:PAM

[KK98b] George Karypis and Vipin Kumar. A parallel algorithm for multilevel graph partitioning and sparse matrix ordering. Journal of Parallel and Distributed Computing, 48(1):71–95, January 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1403/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1403/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1403/production/ref.

Kesselman:2006:FDA

[KK06] Alex Kesselman and Dariusz R. Kowalski. Fast distributed algorithm for convergecast in ad hoc geometric radio networks. Journal of Parallel and Distributed Computing, 66(4):578–585, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Krishnan:2006:ESC

[KKB^{+}06] Sandhya Krishnan, Sriram Krishnamoorthy, Gerald Baumgartner, Chi-Chung Lam, J. Ramanujam, P. Sadayappan, and Venkatesh Choppella. Efficient synthesis of out-of-core algorithms using a nonlinear optimization solver. Journal of Parallel and Distributed Computing, 66(5):659–673, May 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kahol:2001:ADD

[KKGS01] Anurag Kahol, Sumit Khurana, Sandeep K. S. Gupta, and Pradip K. Srimani. Adaptive distributed dynamic channel allocation for wireless networks. Journal of Parallel and Distributed Computing, 61(7):898–914, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1719;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1719/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1719/ref.

Korkhov:2008:GBV

[KKS08] Vladimir V. Korkhov, Valeria V. Krzhizhanovskaya, and P. M. A. Sloot. A Grid-based Virtual Reactor: Parallel performance and adaptive load balancing. Journal of Parallel and Distributed Computing, 68(5):596–608, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Konwar:2009:NDN

[KKS09] Kishori M. Konwar, Dariusz Kowalski, and Alexander A. Shvartsman. Node discovery in networks. Journal of Parallel and Distributed Computing, 69(4):337–348, April 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kadayif:2005:ILP

[KKVI05] I. Kadayif, M. Kandemir, N. Vijaykrishnan, and M. J. Irwin. An integer linear programming-based tool for wireless sensor networks. Journal of Parallel and Distributed Computing, 65 (3):247–260, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kung:1984:WSI

[KL84] H. T. Kung and Monica S. Lam. Wafer-scale integration and two-level pipelined implementations of systolic arrays. Journal of Parallel and Distributed Computing, 1(1):32–63, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kalinov:2001:HDC

[KL01a] Alexey Kalinov and Alexey Lastovetsky. Heterogeneous distribution of computations solving linear algebra problems on networks of heterogeneous computers. Journal of Parallel and Distributed Computing, 61(4):520–535, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1686;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1686/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1686/ref.

Kim:2001:MPC

[KL01b] Sang Cheol Kim and Sunggu Lee. Measurement and prediction of communication delays in Myrinet networks. Journal of Parallel and Distributed Computing, 61(11):1692–1704, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1761;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1761/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1761/ref.

Kosar:2005:FRE

[KL05] Tevfik Kosar and Miron Livny. A framework for reliable and efficient data placement in distributed computing systems. Journal of Parallel and Distributed Computing, 65(10):1146–1157, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaeli:2008:ASI

[KL08a] D. R. Kaeli and M. Leeser. Acknowledgment to special issue reviewers. Journal of Parallel and Distributed Computing, 68 (10):1402, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaeli:2008:SIG

[KL08b] David R. Kaeli and Miriam Leeser. Special issue: Generalpurpose processing using graphics processing units. Journal of Parallel and Distributed Computing, 68(10):1305–1306, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2005:IGS

[KLC05] Jae-Hoon Kim, Jae-Ha Lee, and Kyung-Yong Chwa. Improved gossipings by short messages in 2-dimensional meshes. Journal of Parallel and Distributed Computing, 65(7):793–800, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kung:1987:PAO

[KLL87] S. Y. Kung, P. S. Lewis, and S. C. Lo. Performance analysis and optimization of VLSI dataflow arrays. Journal of Parallel and Distributed Computing, 4(6):592–618, December 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:1998:AVC

[KLLK98] Ho Won Kim, Hyun Suk Lee, Sunggu Lee, and Jong Kim. Adaptive virtual cut-through as a viable routing method. Journal of Parallel and Distributed Computing, 52(1):82–95, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1466/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1466/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1466/production/ref.

Knobe:1990:DOA

[KLS90] K. Knobe, J. D. Lukas, and G. L. Steele, Jr. Data optimization: Allocation of arrays to reduce communication on SIMD machines. Journal of Parallel and Distributed Computing, 8(2): 102–118, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kopidakis:1997:TAP

[KLZ97] Y. Kopidakis, M. Lamari, and V. Zissimopoulos. On the task assignment problem: Two new efficient heuristic algorithms. Journal of Parallel and Distributed Computing, 42(1):21–29, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1311/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1311/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1311/production/ref.

Kornerup:1988:LAU

[KM88] Peter Kornerup and David W. Matula. An on-line arithmetic unit for bit-pipelined rational arithmetic. Journal of Parallel and Distributed Computing, 5(3):310–330, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuo:1991:IMR

[KM91] Steve Kuo and Dan Moldovan. Implementation of multiple rule firing production systems on hypercube. Journal of Parallel and Distributed Computing, 13(4):383–394, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuo:1992:SAP

[KM92] Steve Kuo and Dan Moldovan. The state of the art in parallel production systems. Journal of Parallel and Distributed Computing, 15(1):1–26, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Krishnamoorthy:1997:FAF

[KM97] Ashok V. Krishnamoorthy and David A. B. Miller. Firehose architectures for free-space optically interconnected VLSI circuits. Journal of Parallel and Distributed Computing, 41(1):109–114, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1292/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1292/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1292/production/ref.

Knoll:2003:VPG

[KM03] Peter Knoll and S. Mirzaei. Validation of a parallel genetic algorithm for image reconstruction from projections. Journal of Parallel and Distributed Computing, 63(3):356–359, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuo:1991:PCS

[KMB91] Chin-Ming Kuo, Daniel P. Miranker, and James C. Browne. On the performance of the CREL system. Journal of Parallel and Distributed Computing, 13(4):424–441, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kapelnikov:1989:MMA

[KME89] Alex Kapelnikov, Richard R. Muntz, and Milos D. Ercegovac. A modelling methodology for the analysis of concurrent systems and computations. Journal of Parallel and Distributed Computing, 6(3):568–597, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kapelnikov:1992:MPA

[KME92] A. Kapelnikov, R. R. Muntz, and M. D. Ercegovac. A methodology for performance analysis of parallel computations with looping constructs. Journal of Parallel and Distributed Computing, 14(2):105–120, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Komatitsch:2009:PHO

[KME09] Dimitri Komatitsch, David Michéa, and Gordon Erlebacher. Porting a high-order finite-element earthquake modeling application to NVIDIA graphics cards using CUDA. Journal of Parallel and Distributed Computing, 69(5):451–460, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koibuchi:2005:EOP

[KMF^{+}05] Michihiro Koibuchi, Juan C. Martinez, Jose Flich, Antonio Robles, Pedro Lopez, and Jose Duato. Enforcing in-order packet delivery in system area networks with adaptive routing. Journal of Parallel and Distributed Computing, 65(10): 1223–1236, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Katevenis:1997:TSH

[KMKD97] Manolis G. H. Katevenis, Evangelos P. Markatos, George Kalokerinos, and Apostolos Dollas. Telegraphos: a substrate for high-performance computing on workstation clusters. Journal of Parallel and Distributed Computing, 43(2):94–108, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1334/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1334/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1334/production/ref.

Khattab:2006:HBP

[KMMZ06] Sherif Khattab, Rami Melhem, Daniel Mossé, and Taieb Znati. Honeypot back-propagation for mitigating spoofing distributed denial-of-service attacks. Journal of Parallel and Distributed Computing, 66(9):1152–1164, September 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:2006:PAP

[KMP^{+}06] V. S. Anil Kumar, Madhav V. Marathe, Srinivasan Parthasarathy, Aravind Srinivasan, and Sibylle Zust. Provable algorithms for parallel generalized sweep scheduling. Journal of Parallel and Distributed Computing, 66(6):807–821, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kwok:2006:SSA

[KMS^{+}06] Yu-Kwong Kwok, Anthony A. Maciejewski, Howard Jay Siegel, Ishfaq Ahmad, and Arif Ghafoor. A semi-static approach to mapping dynamic iterative tasks onto heterogeneous computing systems. Journal of Parallel and Distributed Computing, 66 (1):77–98, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koukopoulos:2007:PSB

[KMS07] Dimitrios Koukopoulos, Marios Mavronicolas, and Paul Spirakis. Performance and stability bounds for dynamic networks. Journal of Parallel and Distributed Computing, 67(4):386–399, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:1991:MOO

[KNS91] Shin-Dug Kim, Mark A. Nichols, and Howard Jay Siegel. Modeling overlapped operation between the control unit and processing elements in an SIMD machine. Journal of Parallel and Distributed Computing, 12(4):329–342, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kuruvila:2006:GLR

[KNS06] Johnson Kuruvila, Amiya Nayak, and Ivan Stojmenovic. Greedy localized routing for maximizing probability of delivery in wireless ad hoc networks with a realistic physical layer. Journal of Parallel and Distributed Computing, 66(4):499–506, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koppelman:1990:SRP

[KO90] David M. Koppelman and A. Yavuz Oruc. A self-routing permutation network. Journal of Parallel and Distributed Computing, 10(2):140–151, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khedr:2009:PDW

[KOA09] Ahmed M. Khedr, Walid Osamy, and Dharma P. Agrawal. Perimeter discovery in wireless sensor networks. Journal of Parallel and Distributed Computing, 69(11):922–929, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kochevar:1991:SLS

[Koc91] Peter Kochevar. A simple light simulation algorithm for massively parallel machines. Journal of Parallel and Distributed Computing, 13(2):193–201, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Koppelman:1997:SSB

[Kop97] David M. Koppelman. Sticky states in Banyan network queues and their application to analysis. Journal of Parallel and Distributed Computing, 40(2):147–161, February 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1293/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1293/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1293/production/ref.

Konuru:1997:MUL

[KOW97] Ravi B. Konuru, Steve W. Otto, and Jonathan Walpole. A migratable user-level process package for PVM. Journal of Parallel and Distributed Computing, 40(1):81–102, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1270/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1270/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1270/production/ref.

Kim:2000:FSF

[KP00] Sung-Ryul Kim and Kunsoo Park. Fully scalable fault-tolerant simulations for BSP and CGM. Journal of Parallel and Distributed Computing, 60(12):1531–1560, December 1, 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1659;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1659/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1659/ref.

Kurzak:2005:MPI

[KP05] Jakub Kurzak and B. Montgomery Pettitt. Massively parallel implementation of a fast multipole method for distributed memory machines. Journal of Parallel and Distributed Computing, 65(7):870–881, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karamcheti:1996:RME

[KPC96] Vijay Karamcheti, John Plevyak, and Andrew A. Chien. Runtime mechanisms for efficient dynamic multithreading. Journal of Parallel and Distributed Computing, 37(1):21–40, August 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0105/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0105/production/
pdf.

Kothari:1988:MNC

[KPR88] S. C. Kothari, G. M. Prabhu, and Robert Roberts. A multipath network with cross links. Journal of Parallel and Distributed Computing, 5(2):185–193, April 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:1987:APM

[KR87] V. K. Prasanna Kumar and C. S. Raghavendra. Array processor with multiple broadcasting. Journal of Parallel and Distributed Computing, 4(2):173–190, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaddoura:1997:RSP

[KR97] Maher Kaddoura and Sanjay Ranka. Runtime support for parallelization of data-parallel applications on adaptive and nonuniform computational environments. Journal of Parallel and Distributed Computing, 43(2):163–168, June 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1340/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1340/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1340/production/ref.

Knop:1998:PLT

[KR98] Felipe Knop and Vernon Rego. Parallel labeling of threedimensional clusters on networks of workstations. Journal of Parallel and Distributed Computing, 49(2):182–203, March 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1420/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1420/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1420/production/ref.

Korch:2006:OLS

[KR06] Matthias Korch and Thomas Rauber. Optimizing locality and scalability of embedded Runge–Kutta solvers using block-based pipelining. Journal of Parallel and Distributed Computing, 66 (3):444–468, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kandemir:2000:CAO

[KRC00] M. Kandemir, J. Ramanujam, and A. Choudhary. Compiler algorithms for optimizing locality and parallelism on shared and distributed-memory machines. Journal of Parallel and Distributed Computing, 60(8):924–965, August 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1639;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1639/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1639/ref.

Krikelis:1991:CVA

[Kri91] Anargyros Krikelis. Computer vision applications with the associative string processor. Journal of Parallel and Distributed Computing, 13(2):170–184, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Krikelis:1992:NCM

[Kri92] Anargyros (Argy) Krikelis. Neural computing on massively parallel processing. Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 14(3):201–202, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kovaleski:1987:AIS

[KRL87] A. Kovaleski, S. Ratheal, and F. Lombardi. An architecture and an interconnection scheme for time-sliced buses. Journal of Parallel and Distributed Computing, 4(2):209–229, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:2001:SIH

[KRS01] Vipin Kumar, Sanjay Ranka, and Vineet Singh. Special issue on high-performance data mining. Journal of Parallel and Distributed Computing, 61(3):281–284, March 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1690;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1690/pdf.

Kapoor:2002:GAF

[KRSZ02] Nidhi Kapoor, Mark Russell, Ivan Stojmenovic, and Albert Y. Zomaya. A genetic algorithm for finding the pagenumber of interconnection networks. Journal of Parallel and Distributed Computing, 62(2):267–283, February 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1789;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1789/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1789/ref.

Kaddoura:1996:ADN

[KRW96] Maher Kaddoura, Sanjay Ranka, and Albert Wang. Array decompositions for nonuniform computational environments. Journal of Parallel and Distributed Computing, 36(2):91–105, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0092/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0092/
production/pdf.

Kumar:1991:SPA

[KS91] Vipin Kumar and Vineet Singh. Scalability of parallel algorithms for the all-pairs shortest-path problem. Journal of Parallel and Distributed Computing, 13(2):124–138, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kraemer:1993:VPS

[KS93] Eileen Kraemer and John T. Stasko. The visualization of parallel systems: An overview. Journal of Parallel and Distributed Computing, 18(2):105–117, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1049/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1049/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1993.1050/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1050/production/pdf.

Kshemkalyani:1994:CCD

[KS94] A. D. Kshemkalyani and M. Singhal. On characterization and correctness of distributed deadlock detection. Journal of Parallel and Distributed Computing, 22(1):44–59, July 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1069/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1069/production/pdf.

Kontothanassis:1995:HPS

[KS95] Leonidas I. Kontothanassis and Michael L. Scott. High performance software coherence for current and future architectures. Journal of Parallel and Distributed Computing, 29(2):179–195, September 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1116/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1116/production/pdf.

Karlsson:1997:EDP

[KS97a] Magnus Karlsson and Per Stenström. Effectiveness of dynamic prefetching in multiple-writer distributed virtual shared-memory systems. Journal of Parallel and Distributed Computing, 43(2):79–93, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1333/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1333/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1333/production/ref.

Khemka:1997:OMR

[KS97b] Ashok Khemka and R. K. Shyamasundar. An optimal multiprocessor real-time scheduling algorithm. Journal of Parallel and Distributed Computing, 43(1):37–45, May 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1327/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1327/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1327/production/ref.

Kogan:2000:RRC

[KS00] Dmitry Kogan and Assaf Schuster. Remote reference counting: Distributed garbage collection with low communication and computation overhead. Journal of Parallel and Distributed Computing, 60(10):1260–1292, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1653;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1653/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1653/ref.

Kshemkalyani:2002:CPD

[KS02] Ajay D. Kshemkalyani and Mukesh Singhal. Communication patterns in distributed computations. Journal of Parallel and Distributed Computing, 62(6):1104–1119, June 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1818;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1818/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1818/ref.

Kanhere:2003:AOQ

[KS03] Salil S. Kanhere and Harish Sethu. Anchored opportunity queueing: a low-latency scheduler for fair arbitration among virtual channels. Journal of Parallel and Distributed Computing, 63(12):1288–1299, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khedr:2008:DAN

[KS08] Ahmed M. Khedr and Ahmed Salim. Decomposable algorithms for nearest neighbor computing. Journal of Parallel and Distributed Computing, 68(7):902–912, July 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karinthi:1995:PAC

[KSA95] Raghu Karinthi, Kankanahalli Srinivas, and George Almasi. A parallel algorithm for computing polygon set operations. Journal of Parallel and Distributed Computing, 26(1):85–98, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1050/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1050/production/
pdf.

Kavianpour:1994:NAC

[KSB94] A. Kavianpour, S. Shoari, and N. Bagherzadeh. A new approach for circle detection on multiprocessors. Journal of Parallel and Distributed Computing, 20(2):256–260, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1025/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1025/production/
pdf.

Koziris:2003:PSM

[KSG03] N. Koziris, A. Sotiropoulos, and G. Goumas. A pipelined schedule to minimize completion time for loop tiling with computation and communication overlapping. Journal of Parallel and Distributed Computing, 63(11):1138–1151, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kannan:2004:SCE

[KSI04] Rajgopal Kannan, Sudipta Sarangi, and S. Sitharama Iyengar. Sensor-centric energy-constrained reliable query routing for wireless sensor networks. Journal of Parallel and Distributed Computing, 64(7):839–852, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kent:1985:PPI

[KSL85] Ernest W. Kent, Michael O. Shneier, and Ronald Lumia. Pipe. (pipelined image-processing engine). Journal of Parallel and Distributed Computing, 2(1):50–78, February 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karsai:1992:MBI

[KSP^{+}92] Gabor Karsai, Janos Sztipanovits, Samir Padalkar, Csaba Biegl, Nobuji Miyasaka, and Koji Okuda. Model-based intelligent process control for cogenerator plants. Journal of Parallel and Distributed Computing, 15(2):90–102, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:2007:DMT

[KSS^{+}07] Jong-Kook Kim, Sameer Shivle, Howard Jay Siegel, Anthony A. Maciejewski, Tracy D. Braun, Myron Schneider, Sonja Tideman, Ramakrishna Chitta, Raheleh B. Dilmaghani, Rohit Joshi, Aditya Kaul, Ashish Sharma, Siddhartha Sripada, Praveen Vangari, and Siva Sankar Yellampalli. Dynamically mapping tasks with priorities and multiple deadlines in a heterogeneous environment. Journal of Parallel and Distributed Computing, 67(2):154–169, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kumar:1989:DLS

[KT89] V. K. Prasanna Kumar and Yu-Chen Tsai. Designing linear systolic arrays. Journal of Parallel and Distributed Computing, 7(3):441–463, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kunde:1991:RMM

[KT91] Manfred Kunde and Thomas Tensi. (κ − κ) routing on multidimensional mesh-connected arrays. Journal of Parallel and Distributed Computing, 11(2):146–155, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Karonis:2003:MGG

[KTF03] Nicholas T. Karonis, Brian Toonen, and Ian Foster. MPICH-G2: a Grid-enabled implementation of the Message Passing Interface. Journal of Parallel and Distributed Computing, 63 (5):551–563, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kaya:2007:HSF

[KUA07] Kamer Kaya, Bora Uçar, and Cevdet Aykanat. Heuristics for scheduling file-sharing tasks on heterogeneous systems with distributed repositories. Journal of Parallel and Distributed Computing, 67(3):271–285, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Katayama:2002:LOS

[KUFM02] Yoshiaki Katayama, Eiichiro Ueda, Hideo Fujiwara, and Toshimitsu Masuzawa. A latency optimal superstabilizing mutual exclusion protocol in unidirectional rings. Journal of Parallel and Distributed Computing, 62(5):865–884, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1826;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1826/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1826/ref.

Katz:1988:GRP

[KV88] Martin David Katz and Dennis J. Volper. Geometric retrieval in parallel. Journal of Parallel and Distributed Computing, 5 (1):92–102, February 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Khoo:2007:MDS

[KVHS07] B. T. Benjamin Khoo, Bharadwaj Veeravalli, Terence Hung, and C. W. Simon See. A multi-dimensional scheduling scheme in a Grid computing environment. Journal of Parallel and Distributed Computing, 67(6):659–673, June 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kohn:1993:A

[KW93] James Kohn and Winifred Williams. ATExpert. Journal of Parallel and Distributed Computing, 18(2):205–222, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1057/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1057/production/
pdf.

Ku:2002:OSP

[KW02] Shan-Chyun Ku and Biing-Feng Wang. An optimal simple parallel algorithm for testing isomorphism of maximal outerplanar graphs. Journal of Parallel and Distributed Computing, 62(2):221–227, February 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1787;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1787/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1787/ref.

Krishnamurthy:1996:AOS

[KY96] Arvind Krishnamurthy and Katherine Yelick. Analyses and optimizations for shared address space programs. Journal of Parallel and Distributed Computing, 38(2):130–144, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0136/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0136/production/
pdf.

Kakugawa:2002:USS

[KY02] Hirotsugu Kakugawa and Masafumi Yamashita. Uniform and self-stabilizing fair mutual exclusion on unidirectional rings under unfair distributed daemon. Journal of Parallel and Distributed Computing, 62(5):885–898, May 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1831;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1831/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1831/ref.

Karwande:2005:MPC

[KYL05] Amit Karwande, Xin Yuan, and David K. Lowenthal. An MPI prototype for compiled communication on Ethernet switched clusters. Journal of Parallel and Distributed Computing, 65 (10):1123–1133, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Kim:1996:FMO

[KZ96] Jinwoo Kim and Bernard P. Zeigler. A framework for multiresolution optimization in a parallel/distributed environment: Simulation of hierarchical GAs. Journal of Parallel and Distributed Computing, 32(1):90–102, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0007/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0007/production/pdf.

Lin:1993:UGS

[LA93] Jiajen M. Lin and Santosh G. Abraham. Utilizing global simulation information in conservative parallel simulation on shared memory multiprocessors. Journal of Parallel and Distributed Computing, 18(4):516–523, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1082/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1082/production/pdf.

Loukopoulos:2004:SAD

[LA04] Thanasis Loukopoulos and Ishfaq Ahmad. Static and adaptive distributed data replication using genetic algorithms. Journal of Parallel and Distributed Computing, 64(11):1270–1285, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lopez:2006:ESM

[LA06] F. C. García López and N. L. Frías Arrocha. An efficient synchronization model for OpenMP. Journal of Parallel and Distributed Computing, 66(11):1359–1365, November 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leiserson:1996:NAC

[LAD^{+}96] Charles E. Leiserson, Zahi S. Abuhamdeh, David C. Douglas, Carl R. Feynman, Mahesh N. Ganmukhi, Jeffrey V. Hill, W. Daniel Hillis, Bradley C. Kuszmaul, Margaret A. St Pierre, David S. Wells, Monica C. Wong-Chan, Shaw-Wen Yang, and Robert Zak. The network architecture of the Connection Machine CM-5. Journal of Parallel and Distributed Computing, 33(2):145–158, March 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0033/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0033/production/pdf.

Liu:2007:LVT

[LAGK07] Ke Liu, Nael Abu-Ghazaleh, and Kyoung-Don Kang. Location verification and trust management for resilient geographic routing. Journal of Parallel and Distributed Computing, 67(2): 215–228, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lai:1986:TDD

[Lai86] Ten-Hwang Lai. Termination detection for dynamically distributed systems with non-first-in-first-out communication. Journal of Parallel and Distributed Computing, 3(4):577–599, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lan:1994:AFT

[Lan94] Y. R. Lan. Adaptive fault-tolerant multicast in hypercube multicomputers. Journal of Parallel and Distributed Computing, 23(1):80–93, October 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1121/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1121/production/pdf.

Landfeldt:2009:SIJ

[Lan09] B. Landfeldt. Special issue of the Journal of Parallel and Distributed Computing: Advancement of research in wireless access and mobile systems. Journal of Parallel and Distributed Computing, 69(6):588, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1997:DET

[LAS^{+}97] Yan Alexander Li, John K. Antonio, Howard Jay Siegel, Min Tan, and Daniel W. Watson. Determining the execution time distribution for a data parallel program in a heterogeneous computing environment. Journal of Parallel and Distributed Computing, 44(1):35–52, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1349/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1349/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1349/production/ref.

Latifi:1995:MTI

[Lat95] Shahram Latifi. Migration of tasks in interconnection networks based on the star graph. Journal of Parallel and Distributed Computing, 31(2):166–173, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1155/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1155/production/pdf.

Latifi:1998:FBG

[Lat98] Shahram Latifi. Fast broadcasting and gathering in q-ary cubes using error-correction codes. Journal of Parallel and Distributed Computing, 48(1):52–63, January 10, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1399/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1399/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1399/production/ref.

Lee:2000:IIB

[LAZC00] Seoung-Bum Lee, Gahng-Seop Ahn, Xiaowei Zhang, and Andrew T. Campbell. INSIGNIA: An IP-based quality of service framework for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 60(4):374–406, April 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1613;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1613/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1613/ref.

Ling:1989:SAM

[LB89] Nam Ling and Magdy A. Bayoumi. Systematic algorithm mapping for multidimensional systolic arrays. Journal of Parallel and Distributed Computing, 7(2):368–382, October 1989. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1990:ASS

[LB90] E. A. Lee and J. C. Bier. Architectures for statically scheduled dataflow. Journal of Parallel and Distributed Computing, 10(4): 333–348, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2009:GDS

[LB09] Cong Liu and Sanjeev Baskiyar. A general distributed scalable grid scheduler for independent tasks. Journal of Parallel and Distributed Computing, 69(3):307–314, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lesyng:2003:EEC

[LBE03] Bogdan Lesyng, Piotr Bała, and Dietmar Erwin. EUROGRID — European computational grid testbed. Journal of Parallel and Distributed Computing, 63(5):590–596, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lazar:1995:BAM

[LBL95] Aurel A. Lazar, Shailendra K. Bhonsle, and Koon-Seng S. Lim. A binding architecture for multimedia networks. Journal of Parallel and Distributed Computing, 30(2):204–216, November 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1138/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1138/production/
pdf.

Liang:1994:FTM

[LBT94] A. C. Liang, S. Bhattacharya, and Wei-Tek T. Tsai. Faulttolerant multicasting on hypercubes. Journal of Parallel and Distributed Computing, 23(3):418–428, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1151/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1151/production/pdf.

Li:1990:SJS

[LC90a] Keqin Li and Kam Hoi Cheng. Static job scheduling in partitionable mesh connected systems. Journal of Parallel and Distributed Computing, 10(2):152–159, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1990:SPA

[LC90b] Xiaobo Li and Yian-Leng Chang. Simulating parallel architectures in a distributed environment. Journal of Parallel and Distributed Computing, 9(2):218–223, June 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1991:DAP

[LC91a] Jingke Li and Marina Chen. The data alignment phase in compiling programs for distributed-memory machines. Journal of Parallel and Distributed Computing, 13(2):213–221, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1991:TDB

[LC91b] Keqin Li and Kam-Hoi Cheng. A two-dimensional buddy system for dynamic resource allocation in a partitionable mesh connected system. Journal of Parallel and Distributed Computing, 12(1):79–83, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1992:AGP

[LC92] Yen-Chun Lin and Yu-Ho Cheng. Automatic generation of parallel Occam programs for transputer rings. Journal of Parallel and Distributed Computing, 16(2):121–133, October 1992. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:1996:FTM

[LC96] Neng-Pin Lu and Chung-Ping Chung. A fault-tolerant multistage combining network. Journal of Parallel and Distributed Computing, 34(1):14–28, April 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0042/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0042/production/pdf.

Lauria:1997:MFH

[LC97] Mario Lauria and Andrew Chien. MPI-FM: High performance MPI on workstation clusters. Journal of Parallel and Distributed Computing, 40(1):4–18, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1264/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1264/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1264/production/ref.

Lin:2007:PSA

[LC07] Chow-Sing Lin and Yi-Chi Cheng. P2MCMD: a scalable approach to VoD service over peer-to-peer networks. Journal of Parallel and Distributed Computing, 67(8):903–921, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2005:RMF

[LCC^{+}05] HwaMin Lee, KwangSik Chung, SungHo Chin, JongHyuk Lee, DaeWon Lee, Seongbin Park, and HeonChang Yu. A resource management and fault tolerance services in Grid computing. Journal of Parallel and Distributed Computing, 65(11):1305– 1317, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:2006:TSI

[LCM^{+}06] Paul Lu, Michael Closson, Cam Macdonell, Paul Nalos, Danny Ngo, Morgan Kan, and Mark Lee. The Trellis security infrastructure for overlay metacomputers and bridged distributed file systems. Journal of Parallel and Distributed Computing, 66(9): 1181–1188, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2005:OTC

[LCW05] Yingshu Li, Xiaoyan Cheng, and Weili Wu. Optimal topology control for balanced energy consumption in wireless networks. Journal of Parallel and Distributed Computing, 65(2):124–131, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:1997:QPD

[LDCZ97] Honghui Lu, Sandhya Dwarkadas, Alan L. Cox, and Willy Zwaenepoel. Quantifying the performance differences between PVM and TreadMarks. Journal of Parallel and Distributed Computing, 43(2):65–78, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1332/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1332/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1332/production/ref.

Larowe:1991:PPP

[LE91] Richard P. Larowe, Jr. and Carla Schlatter Ellis. Page placement policies for NUMA multiprocessors. Journal of Parallel and Distributed Computing, 11(2):112–129, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1998:AWR

[LEB98] X. Lin, A-H. Esfahanian, and A. Burago. Adaptive wormhole routing in hypercube multicomputers. Journal of Parallel and Distributed Computing, 48(2):165–174, February 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1417/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1417/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1417/production/ref.

Lee:1990:PRC

[Lee90] Fung F. Lee. Partitioning of regular computation on multiprocessor systems. Journal of Parallel and Distributed Computing, 9(3):312–317, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1991:DAP

[Lee91] De-Lei Lee. Design of an array processor for image processing. Journal of Parallel and Distributed Computing, 11(2):163–169, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1994:GRS

[Lee94] Tony T. Lee. Generalized recursive sorting networks. Journal of Parallel and Distributed Computing, 21(2):237–245, May 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1055/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1055/production/
pdf.

Lee:2003:RTW

[Lee03] Sunggu Lee. Real-time wormhole channels. Journal of Parallel and Distributed Computing, 63(3):299–311, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lan:1990:MHM

[LEN90] Youran Lan, Abdol-Hossein Esfahanian, and Lionel M. Ni. Multicast in hypercube multiprocessors. Journal of Parallel and Distributed Computing, 8(1):30–41, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:1992:SCP

[LF92] Mi Lu and Jesse Zhixi Fang. A solution of the cache ping-pong problem in multiprocessor systems. Journal of Parallel and Distributed Computing, 16(2):158–171, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lublin:2003:WPS

[LF03] Uri Lublin and Dror G. Feitelson. The workload on parallel supercomputers: modeling the characteristics of rigid jobs. Journal of Parallel and Distributed Computing, 63(11):1105–1122, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lowenthal:1996:UFG

[LFA96] David K. Lowenthal, Vincent W. Freeh, and Gregory R. Andrews. Using fine-grain threads and run-time decision making in parallel computing. Journal of Parallel and Distributed Computing, 37(1):41–54, August 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0106/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0106/production/pdf.

Larrea:2005:ECF

[LFA05] Mikel Larrea, Antonio Fernández, and Sergio Arévalo. Eventually consistent failure detectors. Journal of Parallel and Distributed Computing, 65(3):361–373, March 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lewis:2003:SES

[LFH^{+}03] Michael J. Lewis, Adam J. Ferrari, Marty A. Humphrey, John F. Karpovich, Mark M. Morgan, Anand Natrajan, Anh Nguyen-Tuong, Glenn S. Wasson, and Andrew S. Grimshaw. Support for extensibility and site autonomy in the Legion grid system object model. Journal of Parallel and Distributed Computing, 63(5):525–538, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ltaief:2008:FTA

[LGG08] Hatem Ltaief, Edgar Gabriel, and Marc Garbey. Fault tolerant algorithms for heat transfer problems. Journal of Parallel and Distributed Computing, 68(5):663–677, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1992:DGD

[LH92] Tze Chiang Lee and John P. Hayes. Design of gracefully degradable hypercube-connected systems. Journal of Parallel and Distributed Computing, 14(4):390–401, April 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:2004:NAC

[LH04] Yen-Chun Lin and Jun-Wei Hsiao. A new approach to constructing optimal parallel prefix circuits with small depth. Journal of Parallel and Distributed Computing, 64(1):97–107, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Loh:2005:JCA

[LH05] Peter K. K. Loh and Wen Jing Hsu. The Josephus Cube: analysis of routing and fault tolerance. Journal of Parallel and Distributed Computing, 65(1):58–64, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:2009:FPS

[LH09]  Yen-Chun Lin and Li-Ling Hung.  Fast problem-size-

independent parallel prefix circuits. Journal of Parallel and Distributed Computing, 69(4):382–388, April 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1991:VLA

[LHF91] B. Lee, A. R. Hurson, and T. Y. Feng. A vertically layered allocation scheme for data flow systems. Journal of Parallel and Distributed Computing, 11(3):175–187, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Libeskind-Hadas:2001:MAH

[LHHB^{+}01] R. Libeskind-Hadas, J. R. K. Hartline, P. Boothe, G. Rae, and J. Swisher. On multicast algorithms for heterogeneous networks of workstations. Journal of Parallel and Distributed Computing, 61(11):1665–1679, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1759;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1759/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1759/ref.

Lee:2003:LSS

[LHK03] Wan Yeon Lee, Sung Je Hong, and Jong Kim. On-line scheduling of scalable real-time tasks on multiprocessor systems. Journal of Parallel and Distributed Computing, 63(12):1315–1324, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2003:DLB

[LHKL03] Wan Yeon Lee, Sung Je Hong, Jong Kim, and Sunggu Lee. Dynamic load balancing for switch-based networks. Journal of Parallel and Distributed Computing, 63(3):286–298, March 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1995:DDD

[LHM95] Xin Li, Kien A. Hua, and J. Michael Moshell. Distributed database designs and computation strategies for networked interactive simulations. Journal of Parallel and Distributed Computing, 25(1):72–90, February 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1030/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1030/production/pdf.

Loh:2007:REC

[LHP07] Peter Kok Keong Loh, Wen Jing Hsu, and Yi Pan. Reliable and efficient communications in sensor networks. Journal of Parallel and Distributed Computing, 67(8):922–934, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lam:1997:OAA

[LHS97] Chi Chung Lam, C.-H. Huang, and P. Sadayappan. Optimal algorithms for all-to-all personalized communication on rings and two dimensional tori. Journal of Parallel and Distributed Computing, 43(1):3–13, May 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1328/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1328/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1328/production/ref.

Liu:2008:DRC

[LHT08] Bing-Hong Liu, Ping-Chin Huang, and Ming-Jer Tsai. Distributed reformation of core-based group-shared multicast trees in mobile ad hoc networks. Journal of Parallel and Distributed Computing, 68(5):582–595, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Litow:1995:PCL

[LHVW95] Bruce Litow, S. Hossein Hosseini, K. Vairavan, and Gregory S. Wolffe. Performance characteristics of a load balancing algorithm. Journal of Parallel and Distributed Computing, 31(2):159–165, December 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1154/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1154/production/pdf.

Li:1992:MDF

[Li92] Qiang Li. Minimum deadlock-free message routing restrictions in binary hypercubes. Journal of Parallel and Distributed Computing, 15(2):153–159, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1999:CRD

[Li99] Bo Li. Content replication in a distributed and controlled environment. Journal of Parallel and Distributed Computing, 59(2):229–251, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1572/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1572/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1572/production/ref.

Li:2001:SPM

[Li01] Keqin Li. Scalable parallel matrix multiplication on distributed memory parallel computers. Journal of Parallel and Distributed Computing, 61(12):1709–1731, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1768;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1768/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1768/ref.

Li:2005:JSP

[Li05] Keqin Li. Job scheduling and processor allocation for Grid computing on metacomputers. Journal of Parallel and Distributed Computing, 65(11):1406–1418, November 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2006:ACA

[Li06a] Keqin Li. An average-case analysis of online non-clairvoyant scheduling of independent parallel tasks. Journal of Parallel and Distributed Computing, 66(5):617–625, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2006:ACP

[Li06b] Keqin Li. Average-case performance analysis of scheduling random parallel tasks with precedence constraints on mesh connected multicomputer systems. Journal of Parallel and Distributed Computing, 66(8):1090–1102, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1991:PAB

[Lin91] Avi Lin. Parallel algorithms for boundary value problems. Journal of Parallel and Distributed Computing, 11(4):284–290, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1993:POG

[Lin93a] Yen-Chun C. Lin. Perfectly overlapped generation of long runs for sorting large files. Journal of Parallel and Distributed Computing, 19(2):136–142, October 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1099/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1099/production/pdf.

Lin:1993:SIP

[Lin93b] Yi-Bing Lin. Special issue on parallel discrete event simulation: Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 18(4):391–394, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1993:TDD

[Lin93c] Yi-Bing B. Lin. On terminating a distributed discrete event simulation. Journal of Parallel and Distributed Computing, 19(4):364–371, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1118/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1118/production/pdf.

Lin:2003:EED

[Lin03] Chin-Tsai Lin. Embedding k(n−k) edge-disjoint spanning trees in arrangement graphs. Journal of Parallel and Distributed Computing, 63(12):1277–1287, December 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lisper:1990:STO

[Lis90] Bjorn Lisper. Synthesis of time-optimal systolic arrays with cells with inner structure. Journal of Parallel and Distributed Computing, 10(2):182–187, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1986:SSN

[LJ86] H. F. Li and R. Jayakumar. Systolic structures: a notion and characterization. Journal of Parallel and Distributed Computing, 3(3):373–397, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2005:EDR

[LJ05] Soojung Lee and Kil Hong Joo. Efficient detection and resolution of OR deadlocks in distributed systems. Journal of Parallel and Distributed Computing, 65(9):985–993, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2002:PPR

[LJKS02] Shin-Jae Lee, Minsoo Jeon, Dongseung Kim, and Andrew Sohn. Partitioned parallel radix sort. Journal of Parallel and Distributed Computing, 62(4):656–668, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1808;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1808/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1808/ref.

Lang:1990:NTS

[LK90] T. Lang and L. Kurisaki. Nonuniform traffic spots (NUTS) in multistage interconnection networks. Journal of Parallel and Distributed Computing, 10(1):55–67, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1994:OSL

[LK94] Jie Li and H. Kameda. Optimal static load balancing in star network configurations with two-way traffic. Journal of Parallel and Distributed Computing, 23(3):364–375, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1146/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1146/production/
pdf.

Lee:1996:PSM

[LK96] Sunggu Lee and Jong Kim. Path selection for message passing in a circuit-switched multicomputer. Journal of Parallel and Distributed Computing, 35(2):211–218, June 15, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0083/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0083/production/pdf.

Lee:1998:LCS

[LK98] Gyungho Lee and Byung-Chang Kang. Limited combining strategies for large-scale shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 52(2):109–119, August 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1460/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1460/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1460/production/ref.

Lee:1994:ECR

[LKK94] Gyungho G. Lee, Clyde P. Kruskal, and David J. Kuck. On the effectiveness of combining in resolving ‘hot spot’ contention. Journal of Parallel and Distributed Computing, 20(2):136–144, February 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1014/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1014/production/pdf.

Lee:1990:DPE

[LL90] Dik Lun Lee and Chun-Wu Leng. Design and performance evaluation of an associative memory with distributed control. Journal of Parallel and Distributed Computing, 10(1):68–78, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:1995:XSS

[LL95] Zhiyong Y. Liu and Xiaobo B. Li. XOR storage schemes for frequently used data patterns. Journal of Parallel and Distributed Computing, 25(2):162–173, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1038/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1038/production/pdf.

Lu:1998:NPD

[LL98] Qin Lu and Sau-Ming Lau. A negotiation protocol for dynamic load distribution using batch task assignments. Journal of Parallel and Distributed Computing, 55(2):166–191, December 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1501/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1501/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1501/production/ref.

Li:2007:UBQ

[LL07] Chunlin Li and Layuan Li. Utility-based QoS optimisation strategy for multi-criteria scheduling on the Grid. Journal of Parallel and Distributed Computing, 67(2):142–153, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ling:2002:PAP

[LLCC02] W. Ling, J. Liu, J. N. Chung, and C. T. Crowe. Parallel algorithms for particles-turbulence two-way interaction direct numerical simulation. Journal of Parallel and Distributed Computing, 62(1):38–60, January 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1781;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1781/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1781/ref.

Liu:1998:NAF

[LLCL98] J. G. Liu, H. F. Li, F. H. Y. Chan, and F. K. Lam. A novel approach to Fast Discrete Fourier Transform. Journal of Parallel and Distributed Computing, 54(1):48–58, October 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1472/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1472/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1472/production/ref.

Lee:2000:HNO

[LLJ00a] Kyungsook Y. Lee, Guoping Liu, and Harry F. Jordan. Hierarchical networks for optical communications. Journal of Parallel and Distributed Computing, 60(1):1–16, January 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1591/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1591/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1591/production/ref.

Lee:2000:THT

[LLJ00b] Kyungsook Y. Lee, Guoping Liu, and Harry F. Jordan. TDM hypercube and TWDM mesh optical interconnections. Journal of Parallel and Distributed Computing, 60(3):320–333, March 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1999.1605;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1605/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1605/ref.

Lau:2006:ALD

[LLL06] Sau-Ming Lau, Qin Lu, and Kwong-Sak Leung. Adaptive load distribution algorithms for heterogeneous distributed systems with multiple task classes. Journal of Parallel and Distributed Computing, 66(2):163–180, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leu:2008:DWD

[LLLY08] Fang-Yie Leu, Ming-Chang Li, Jia-Chun Lin, and Chao-Tung Yang. Detection workload in a dynamic Grid-based intrusion detection environment. Journal of Parallel and Distributed Computing, 68(4):427–442, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1993:OFC

[LLS93] C. Lee, H. Lam, and S. Y. W. Su. An object flow computer for database applications: Design and performance evaluation. Journal of Parallel and Distributed Computing, 17(4):298–314, April 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1029/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1029/production/pdf.

Luo:2007:RFG

[LLS07] Ping Luo, Kevin Lü, and Zhongzhi Shi. A revisit of fast greedy heuristics for mapping a class of independent tasks onto heterogeneous computing systems. Journal of Parallel and Distributed Computing, 67(6):695–714, June 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2007:DCA

[LLW07] Chuan-Ming Liu, Chuan-Hsiu Lee, and Li-Chun Wang. Distributed clustering algorithms for data-gathering in wireless mobile sensor networks. Journal of Parallel and Distributed Computing, 67(11):1187–1200, November 2007. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1996:RTV

[LM96] Jong-Pyng Li and Matt W. Mutka. Real-time virtual channel flow control. Journal of Parallel and Distributed Computing, 32(1):49–65, January 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0004/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0004/production/pdf.

Lange:2005:HAP

[LM05] Sebastian Lange and Martin Middendorf. Hyperreconfigurable architectures and the partition into hypercontexts problem. Journal of Parallel and Distributed Computing, 65(6):743–754, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Luo:2009:FFI

[LM09] Yi Luo and D. Manivannan. FINE: a Fully Informed aNd Efficient communication-induced checkpointing protocol for distributed systems. Journal of Parallel and Distributed Computing, 69(2):153–167, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

LeBlanc:1990:APP

[LMCF90] Thomas J. LeBlanc, John M. Mellor-Crummey, and Robert J. Fowler. Analyzing parallel program executions using multiple views. Journal of Parallel and Distributed Computing, 9 (2):203–217, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1995:AMW

[LME95] Xiaola Lin, Philip K. McKinley, and Abdol-Hossein Esfahanian. Adaptive multicast wormhole routing in 2D mesh multicomputers. Journal of Parallel and Distributed Computing, 28(1): 19–31, July 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1086/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1086/
production/pdf.

Legrand:2005:OSS

[LMR05] A. Legrand, L. Marchal, and Y. Robert. Optimizing the steadystate throughput of scatter and reduce operations on heterogeneous platforms. Journal of Parallel and Distributed Computing, 65(12):1497–1514, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2004:EAC

[LMZ04] Shu Li, Rami Melhem, and Taieb Znati. An efficient algorithm for constructing delay bounded minimum cost multicast trees. Journal of Parallel and Distributed Computing, 64(12):1399– 1413, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1991:NRA

[LO91] R. Lin and S. Olariu. An NC recognition algorithm for cographs. Journal of Parallel and Distributed Computing, 13 (1):76–90, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lo:1992:TCG

[Lo92] Virginia M. Lo. Temporal communication graphs: Lamport’s process-time graphs augmented for the purpose of mapping and scheduling. Journal of Parallel and Distributed Computing, 16 (4):378–384, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1994:OPM

[LO94] R. Lin and S. Olariu. An optimal parallel matching algorithm for cographs. Journal of Parallel and Distributed Computing, 22(1):26–36, July 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1067/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1067/production/pdf.

Lin:1996:OVS

[LO96] Xuemin Lin and Maria E. Orlowska. An optimal voting scheme for minimizing the overall communication cost in replicated data management. Journal of Parallel and Distributed Computing, 35(1):97–103, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0072/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0072/production/pdf.

Long:2004:DRS

[Lon04] Brad Long. Distributed result set iterator: a design pattern for efficient retrieval of large result sets from remote data sources. Journal of Parallel and Distributed Computing, 64(4):536–545, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Loscri:2008:MPW

[Los08] V. Loscrı‘. MAC protocols over wireless mesh networks: problems and perspective. Journal of Parallel and Distributed Computing, 68(3):387–397, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Luk:1988:AAB

[LP88] Franklin T. Luk and Haesun Park. An analysis of algorithmbased fault tolerance techniques. Journal of Parallel and Distributed Computing, 5(2):172–184, April 1988. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1995:BDO

[LP95] Cho-Chin C. Lin and Viktor K. Prasanna. Bounds on the diameter of one-dimensional PEC networks. Journal of Parallel and Distributed Computing, 29(1):1–16, August 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1102/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1102/production/pdf.

Larmore:1996:PAO

[LP96a] Lawrence L. Larmore and Teresa M. Przytycka. A parallel algorithm for optimum height-limited alphabetic binary trees. Journal of Parallel and Distributed Computing, 35(1):49–56, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0067/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0067/
production/pdf.

Lee:1996:PME

[LP96b] Chiung-San Lee and Tai-Ming Parng. Performance modeling and evaluation of a two-dimensional disk array system. Journal of Parallel and Distributed Computing, 38(1):16–27, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0125/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0125/
production/pdf.

Liu:1997:PAL

[LP97] Shih-Mim Liu and G. P. Papavassilopoulos. A parallel algorithm for linear programs with an additional reverse convex constraint. Journal of Parallel and Distributed Computing, 45(2):91–103, September 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1366/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1366/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1366/production/ref.

Li:1998:LBD

[LPS^{+}98] Keqin Li, Yi Pan, Hong Shen, Gilbert H. Young, and Si Qing Zheng. Lower bounds for dynamic tree embedding in bipartite networks. Journal of Parallel and Distributed Computing, 53(2):119–143, September 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1475/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1475/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1475/production/ref.

Leung:1997:RTV

[LPU97] Allen Leung, Krishna V. Palem, and Cristian Ungureanu. Runtime versus compile-time instruction scheduling in superscalar (RISC) processors: Performance and trade-off. Journal of Parallel and Distributed Computing, 45(1):13–28, August 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1329/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1329/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1329/production/ref.

Li:2005:PSM

[LPX05a] Jie Li, Yi Pan, and Yang Xiao. Performance study of multiple route dynamic source routing protocols for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 65(2): 169–177, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2005:ILS

[LPX05b] G. Q. Liu, K. L. Poh, and M. Xie. Iterative list scheduling for heterogeneous computing. Journal of Parallel and Distributed Computing, 65(5):654–665, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1999:PMC

[LPZ99] Keqin Li, Yi Pan, and Si Qing Zheng. Parallel matrix computations using a reconfigurable pipelined optical bus. Journal of Parallel and Distributed Computing, 59(1):13–30, October 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1569/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1569/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1569/production/ref.

Ligon:1993:EME

[LR93] W. B. Ligon, III and U. Ramachandran. An empirical methodology for exploring reconfigurable architectures. Journal of Parallel and Distributed Computing, 19(4):323–337, December 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1115/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1115/production/
pdf.

Ligon:1994:EMA

[LR94] W. B. Ligon, III and U. Ramachandran. Evaluating multigauge architectures for computer vision. Journal of Parallel and Distributed Computing, 21(3):323–333, June 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1063/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1063/production/pdf.

Li:2003:CDA

[LR03a] Qun Li and Daniela Rus. Communication in disconnected ad hoc networks using message relay. Journal of Parallel and Distributed Computing, 63(1):75–86, January 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lindsey:2003:EEA

[LR03b] Stephanie Lindsey and Cauligi S. Raghavendra. Energy efficient all-to-all broadcasting for situation awareness in wireless ad hoc networks. Journal of Parallel and Distributed Computing, 63 (1):15–21, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ludwig:2005:ISM

[LR05] Simone A. Ludwig and S. M. S. Reyhani. Introduction of semantic matchmaking to Grid computing. Journal of Parallel and Distributed Computing, 65(12):1533–1541, December 2005.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lastovetsky:2006:HTM

[LR06] Alexey Lastovetsky and Ravi Reddy. HeteroMPI: Towards a message-passing library for heterogeneous networks of computers. Journal of Parallel and Distributed Computing, 66(2): 197–220, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1985:VCV

[LS85] Kuo-Cheng Li and Herb Schwetman. Vector C: a vector processing language. Journal of Parallel and Distributed Computing, 2(2):132–169, May 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1988:PVR

[LS88] H. X. Lin and H. J. Sips. Parallel vector reduction algorithms and architectures. Journal of Parallel and Distributed Computing, 5(2):103–130, April 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1991:PEH

[LS91] Jianhua Lin and James A. Storer. Processor-efficient hypercube algorithms for the knapsack problem. Journal of Parallel and Distributed Computing, 13(3):332–337, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leung:1994:MPR

[LS94] Joseph Y.-T. T. Leung and Sunil M. Shende. On multidimensional packet routing for meshes with buses. Journal of Parallel and Distributed Computing, 20(2):187–197, February 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1018/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1018/production/pdf.

Lai:1995:TRM

[LS95] Ten-Hwang H. Lai and Ming-Jye J. Sheng. Triangulation on reconfigurable meshes: a natural decomposition approach. Journal of Parallel and Distributed Computing, 30(1):38–51, October 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1124/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1124/
production/pdf.

Leutenegger:1997:LCS

[LS97] Scott T. Leutenegger and Xian-He Sun. Limitations of cycle stealing for parallel processing on a network of homogeneous workstations. Journal of Parallel and Distributed Computing, 43(2):169–178, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1341/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1341/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1341/production/ref.

Lundberg:2001:ORS

[LS01] Lars Lundberg and Charlie Svahnberg. Optimal recovery schemes for high-availability cluster and distributed computing. Journal of Parallel and Distributed Computing, 61(11):1680–1691, November 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1760;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1760/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1760/ref.

Lin:2003:LBL

[LS03] Xu Lin and Ivan Stojmenovic. Location-based localized alternate, disjoint and multi-path routing algorithms for wireless networks. Journal of Parallel and Distributed Computing, 63 (1):22–32, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:2005:FOP

[LS05] Yen-Chun Lin and Chin-Yu Su. Faster optimal parallel prefix circuits: New algorithmic construction. Journal of Parallel and Distributed Computing, 65(12):1585–1595, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lufei:2006:FMC

[LS06] H. Lufei and W. Shi. Fractal: a mobile code-based framework for dynamic application protocol adaptation. Journal of Parallel and Distributed Computing, 66(7):887–906, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2000:OEM

[LSC00] Jae-Ha Lee, Chan-Su Shin, and Kyung-Yong Chwa. Optimal embedding of multiple directed Hamiltonian rings into d-dimensional meshes. Journal of Parallel and Distributed Computing, 60(6):775–783, June 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1631;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1631/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1631/ref.

LaPolla:1993:DPP

[LSCA93] Mark Vincent LaPolla, Joseph L. Sharnowski, Betty H. C. Cheng, and Kevin Anderson. Data parallel program visualizations from formal specifications. Journal of Parallel and Distributed Computing, 18(2):252–257, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1061/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1061/production/pdf.

Liang:1996:PAE

[LSH96] Weifa Liang, Xiaojun Shen, and Qing Hu. Parallel algorithms for the edge-coloring and edge-coloring update problems. Journal of Parallel and Distributed Computing, 32(1):66–73, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0005/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0005/
production/pdf.

Lee:1988:HAK

[LSS88] J. Lee, E. Shragowitz, and S. Sahni. A hypercube algorithm for the 0/1 knapsack problem. Journal of Parallel and Distributed Computing, 5(4):438–456, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:1994:LBJ

[LT94] Hongjun J. Lu and Kian-Lee L. Tan. Load-balanced join processing in shared-nothing systems. Journal of Parallel and Distributed Computing, 23(3):382–398, December 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1148/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1148/production/pdf.

Lin:1996:EUH

[LT96] Guanghua Lin and Nian-Feng Tzeng. Effective utilization of hypercubes in the presence of faults. Journal of Parallel and Distributed Computing, 32(2):223–231, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0016/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0016/production/pdf.

Li:2007:RCP

[LT07] Shuo-Yen Robert Li and Xuesong Jonathan Tan. Recursive construction of parallel distribution networks. Journal of Parallel and Distributed Computing, 67(6):617–634, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lan:2002:NDL

[LTB02] Zhiling Lan, Valerie E. Taylor, and Greg Bryan. A novel dynamic load balancing scheme for parallel systems. Journal of Parallel and Distributed Computing, 62(12):1763–1781, December 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1993:PCC

[LTD^{+}93] Mengjou Lin, Rose P. Tsang, David H. C. Du, Alan E. Klietz, and Stephen Saroff. Performance characteristics of the Connection Machine hypertree network. Journal of Parallel and Distributed Computing, 19(3):245–254, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1108/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1108/production/pdf.

Lee:1997:PAO

[LTH97] Jeng Kuen Lee, Ing-Kuen Tsaur, and San-Yih Hwang. Parallel array object I/O support on distributed environments. Journal of Parallel and Distributed Computing, 40(2):227–241, February 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1278/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1278/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1278/production/ref.

Lian:2005:GGE

[LTIK05] Chua Ching Lian, Francis Tang, Praveen Issac, and Arun Krishnan. GEL: Grid Execution Language. Journal of Parallel and Distributed Computing, 65(7):857–869, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1990:PPC

[LTKS90] C. S. Lin, A. L. Thring, J. Koga, and E. J. Seiler. A parallel particle-in-cell model for the Massively Parallel Processor. Journal of Parallel and Distributed Computing, 8(2):196–199, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lan:2006:DIC

[LTL06] Zhiling Lan, Valerie E. Taylor, and Yawei Li. DistDLB: Improving cosmology SAMR simulations on distributed computing systems through hierarchical load balancing. Journal of Parallel and Distributed Computing, 66(5):716–731, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leung:1990:PSS

[LTW^{+}90] Joseph Y.-T. Leung, Tommy W. Tam, C. S. Wong, G. H. Young, and F. Y. L. Chin. Packing squares into a square. Journal of Parallel and Distributed Computing, 10(3):271–275, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leung:1995:RMR

[LTWY95] Joseph Y-T Leung, Tommy W. Tam, C. S. Wong, and Gilbert H. Young. Routing messages with release time and deadline constraints. Journal of Parallel and Distributed Computing, 31(1):65–76, November 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1145/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1145/production/pdf.

Leung:1996:LRR

[LTY96] Joseph Y-T Leung, Tommy W. Tam, and Gilbert H. Young. On-line routing of real-time messages. Journal of Parallel and Distributed Computing, 34(2):211–217, May 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0057/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0057/production/pdf.

Lu:2001:IBD

[Lu01] Paul Lu. Integrating bulk-data transfer into the Aurora distributed shared data system. Journal of Parallel and Distributed Computing, 61(11):1609–1632, November 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1758;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1758/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1758/ref.

Luk:1985:PMC

[Luk85] Franklin T. Luk. A parallel method for computing the generalized singular value decomposition. Journal of Parallel and Distributed Computing, 2(3):250–260, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lundberg:1990:PRG

[Lun90] Lars Lundberg. A protocol to reduce global communication in distributed Ada tasking. Journal of Parallel and Distributed Computing, 10(3):261–264, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lundberg:1994:PBM

[Lun94] L. Lundberg. Performance bounds on multiprocessor scheduling strategies for chain structured programs. Journal of Parallel and Distributed Computing, 23(1):112–118, October 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1125/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1125/production/pdf.

Lundberg:1999:PBS

[Lun99] Lars Lundberg. Predicting and bounding the speedup of multithreaded Solaris programs. Journal of Parallel and Distributed Computing, 57(3):322–333, June 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1536/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1536/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1536/production/ref.

Lu:1988:OAR

[LV88] Mi Lu and Peter Varman. Optimal algorithms for rectangle problems on a mesh-connected computer. Journal of Parallel and Distributed Computing, 5(2):154–171, April 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Low:2007:DHP

[LVB07] Diana H. P. Low, Bharadwaj Veeravalli, and David A. Bader. On the design of high-performance algorithms for aligning multiple protein sequences on mesh-based multiprocessor architectures. Journal of Parallel and Distributed Computing, 67(9): 1007–1017, September 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2007:WAV

[LVP07] Xiaorong Li, Bharadwaj Veeravalli, and Viktor K. Prasanna. A window-assisted video partitioning strategy for partitioning and caching video streams in distributed multimedia systems. Journal of Parallel and Distributed Computing, 67(6):738–754, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lavault:2008:DAA

[LVP08] Christian Lavault and Mario Valencia-Pabon. A distributed approximation algorithm for the minimum degree minimum weight spanning trees. Journal of Parallel and Distributed Computing, 68(2):200–208, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:1990:TAD

[LVR90] H. F. Li, K. Venkatesh, and T. Radhakrishnan. Time advancement in distributed event simulation. Journal of Parallel and Distributed Computing, 9(1):15–25, May 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:1989:APC

[LW89] Yu-Cheng Liu and Chungching Wang. Analysis of prioritized crossbar multiprocessor systems. Journal of Parallel and Distributed Computing, 7(2):321–334, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lai:1990:MPA

[LW90] Ten-Hwang Lai and William White. Mapping pyramid algorithms into hypercubes. Journal of Parallel and Distributed Computing, 9(1):42–54, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lin:1995:MPR

[LW95] Ferng-Ching C. Lin and Feng-Hsu H. Wang. Message pattern routing in hypercubes: a distribute- concentrate approach. Journal of Parallel and Distributed Computing, 29(1):27–42, August 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1104/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1104/
production/pdf.

Li:2006:SAA

[LW06a] Xiang-Yang Li and Yu Wang. Simple approximation algorithms and PTASs for various problems in wireless ad hoc networks. Journal of Parallel and Distributed Computing, 66(4):515–530, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Liu:2006:RTS

[LW06b] Tzong-Jye Liu and Lih-Chyau Wuu. Randomized three-state alternator for uniform rings. Journal of Parallel and Distributed Computing, 66(10):1347–1351, October 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2002:EKF

[LWOG02] Shuhui Li, Donald C. Wunsch, Edgar O’Hair, and Michael G. Giesselmann. Extended Kalman filter training of neural networks on a SIMD parallel machine. Journal of Parallel and Distributed Computing, 62(4):544–562, April 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1807;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1807/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1807/ref.

Liu:2002:PSA

[LWP02] Wenheng Liu, Cho-Li Wang, and Viktor K. Prasanna. Portable and scalable algorithm for irregular all-to-all communication. Journal of Parallel and Distributed Computing, 62(10):1493–1526, October 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1862;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1862/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1862/ref.

Li:2003:EHP

[LWR^{+}03] M. Li, D. W. Walker, O. F. Rana, Y. Huang, P. T. Williams, and R. C. Ward. Engineering high-performance legacy codes as CORBA components for problem-solving environments. Journal of Parallel and Distributed Computing, 63(11):1152–1163, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:1997:GOM

[LWY97] Cheolwhan Lee, Yuan-Fang Wang, and Tao Yang. Global optimization for mapping parallel image processing tasks on distributed memory machines. Journal of Parallel and Distributed Computing, 45(1):29–45, August 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1360/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1360/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1360/production/ref.

Leff:1991:ASL

[LY91] Avraham Leff and Philip S. Yu. An adaptive strategy for load sharing in distributed database environments with information lags. Journal of Parallel and Distributed Computing, 13(1):91– 103, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lim:1998:MCC

[LY98] Hock-Beng Lim and Pen-Chung Yew. Maintaining cache coherence through compiler-directed data prefetching. Journal of Parallel and Distributed Computing, 53(2):144–173, September 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1480/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1480/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1480/production/ref.

Lim:2001:EIC

[LY01] Hock-Beng Lim and Pen-Chung Yew. Efficient integration of compiler-directed cache coherence and data prefetching. Journal of Parallel and Distributed Computing, 61(12):1775–1802, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1784;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1784/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1784/ref.

Lin:2008:AIS

[LY08] Cho-Chin Lin and Hao-Yun Yin. On the assessment of input streams for incremental network computing. Journal of Parallel and Distributed Computing, 68(9):1297–1304, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Leung:2002:NNS

[LYC02] K. K. Leung, N. H. C. Yung, and P. Y. S. Cheung. Novel neighborhood search for multiprocessor scheduling with pipelining. Journal of Parallel and Distributed Computing, 62(1):85–110, January 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1779;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1779/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1779/ref.

Lee:1993:PEC

[LYL93] Kyungsook Y. Lee, Hyunsoo Yoon, and Ming T. Liu. Performance evaluation of a class of multipath packet switching interconnection networks. Journal of Parallel and Distributed Computing, 17(4):353–359, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1033/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1033/production/pdf.

Li:2008:PSL

[LYL08] Yiming Li, Shao-Ming Yu, and Yih-Lang Li. Parallel solution of large-scale eigenvalue problem for master equation in protein folding dynamics. Journal of Parallel and Distributed Computing, 68(5):678–685, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lee:2002:SBC

[LZ02] Jong S. Lee and Bernard P. Zeigler. Space-based communication data management in scalable distributed simulation. Journal of Parallel and Distributed Computing, 62(3):336–365, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1798;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1798/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1798/ref.

Leung:2005:MMF

[LZ05] Joseph Y-T. Leung and Hairong Zhao. Minimizing mean flowtime and makespan on master–slave systems. Journal of Parallel and Distributed Computing, 65(7):843–856, July 2005. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Landfeldt:2008:BAS

[LZ08] Bjorn Landfeldt and Albert Y. Zomaya. Behavior, artifacts, and solutions in wireless mesh networks. Journal of Parallel and Distributed Computing, 68(3):275–277, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lei:2009:CCM

[LZCY09] Xiangdong Lei, Yuelong Zhao, Songqiao Chen, and Xiaoli Yuan. Concurrency control in mobile distributed real-time database systems. Journal of Parallel and Distributed Computing, 69(10):866–876, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lu:2006:MTH

[LZSL06] Chenghuai Lu, Tao Zhang, Weidong Shi, and Hsien-Hsin S. Lee. M-TREE: a high efficiency security architecture for protecting integrity and privacy of software. Journal of Parallel and Distributed Computing, 66(9):1116–1128, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Maheshwari:1995:PSP

[Mah95] Piyush Maheshwari. Partitioning and scheduling of parallel functional programs for larger grain execution. Journal of Parallel and Distributed Computing, 26(2):151–165, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1056/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1056/production/
pdf.

Mohamed:2005:DUL

[MAJJ05] Nader Mohamed, Jameela Al-Jaroodi, and Hong Jiang. Dependable user-level socket over dual networks. Journal of Parallel and Distributed Computing, 65(10):1261–1270, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Manoj:2005:MHC

[MAM05] B. S. Manoj, R. Ananthapadmanabha, and C. Siva Ram Murthy. Multi-hop cellular networks: Architecture and protocols for best-effort and real-time communication. Journal of Parallel and Distributed Computing, 65(6):767–791, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Manzini:1994:SMC

[Man94] Giovanni Manzini. Sparse matrix computations on the hypercube and related networks. Journal of Parallel and Distributed Computing, 21(2):169–183, May 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1050/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1050/production/pdf.

Mans:1997:ODA

[Man97] Bernard Mans. Optimal distributed algorithms in unlabeled tori and chordal rings. Journal of Parallel and Distributed Computing, 46(1):80–90, October 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1389/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1389/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1389/production/ref.

Mudge:1987:VAH

[MAR87] T. N. Mudge and T. S. Abdel-Rahman. Vision algorithms for hypercube machines. Journal of Parallel and Distributed Computing, 4(1):79–94, February 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Martel:1988:PAP

[Mar88] Charles U. Martel. A parallel algorithm for preemptive scheduling of uniform machines. Journal of Parallel and Distributed Computing, 5(6):700–715, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mnaouer:2005:ESL

[MAR05] Adel Ben Mnaouer and Badriya Al-Riyami. Effective scheduling of local interactive processes and parallel processes in a nondedicated cluster environment. Journal of Parallel and Distributed Computing, 65(6):755–766, June 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Maheswaran:1999:DMC

[MAS^{+}99] Muthucumaru Maheswaran, Shoukat Ali, Howard Jay Siegel, Debra Hensgen, and Richard F. Freund. Dynamic mapping of a class of independent tasks onto heterogeneous computing systems. Journal of Parallel and Distributed Computing, 59(2):107–131, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1579/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1579/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1579/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1581/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1581/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1581/production/ref.

Mattern:1993:EAD

[Mat93] Friedemann Mattern. Efficient algorithms for distributed snapshots and global virtual time approximation. Journal of Parallel and Distributed Computing, 18(4):423–434, August 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1075/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1075/production/pdf.

Matsumae:2006:TBS

[Mat06] Susumu Matsumae. Tight bounds on the simulation of meshes with dynamically reconfigurable row/column buses by meshes with statically partitioned buses. Journal of Parallel and Distributed Computing, 66(10):1338–1346, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Menasce:1992:MPE

[MB92] Daniel A. Menasce and Luiz Andre Barroso. A methodology for performance evaluation of parallel applications on multiprocessors. Journal of Parallel and Distributed Computing, 14 (1):1–14, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mitkas:1993:POD

[MB93] Pericles A. Mitkas and P. Bruce Berra. PHOEBUS: An optoelectronic database machine based on parallel optical disks. Journal of Parallel and Distributed Computing, 17(3):230–244, March 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1022/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1022/
production/pdf.

MacPherson:1996:PAV

[MB96a] Ky MacPherson and Prithviraj Banerjee. Parallel algorithms for VLSI layout verification. Journal of Parallel and Distributed Computing, 36(2):156–172, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0096/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0096/production/pdf.

May:1996:REP

[MB96b] John May and Francine Berman. Retargetability and extensibility in a parallel debugger. Journal of Parallel and Distributed Computing, 35(2):142–155, June 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0077/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0077/production/pdf.

Morillo:2008:APC

[MBH^{+}08] P. Morillo, A. Bierbaum, P. Hartling, M. Fernández, and C. Cruz-Neira. Analyzing the performance of a cluster-based architecture for immersive visualization systems. Journal of Parallel and Distributed Computing, 68(2):221–234, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Morgan:1992:RAP

[MBK^{+}92] Nelson Morgan, James Beck, Phil Kohn, Jeff Bilmes, Eric Allman, and Joachim Beer. The ring array processor: a multiprocessing peripheral for connectionist applications. Journal of Parallel and Distributed Computing, 14(3):248–259, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Marsh:1992:OSS

[MBL^{+}92] B. Marsh, C. Brown, T. LeBlanc, M. Scott, T. Becker, P. Das, J. Karlsson, and C. Quiroz. Operating system support for animate vision. Journal of Parallel and Distributed Computing, 15(2):103–117, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mirabella:2008:IRL

[MBR08] O. Mirabella, Lo Bello, and A. Raucea. Improving routing in long-distance wireless mesh networks via a distributed embedded router. Journal of Parallel and Distributed Computing, 68 (3):361–371, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Martinez:1991:SAD

[MC91] Tony R. Martinez and Douglas M. Campbell. A self-adjusting dynamic logic module. Journal of Parallel and Distributed Computing, 11(4):303–313, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Melhem:1993:OCI

[MC93] Rami Melhem and Donald Chiarulli. Optical computing and interconnection systems: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 17(3):185–187, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Myers:2003:NMI

[MC03] Daniel S. Myers and Michael P. Cummings. Necessity is the mother of invention: a simple grid computing system using commodity tools. Journal of Parallel and Distributed Computing, 63(5):578–589, May 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

McAulay:1989:CGO

[McA89] A. D. McAulay. Conjugate gradients on optical crossbar interconnected multiprocessor. Journal of Parallel and Distributed Computing, 6(1):136–150, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ma:2004:BBT

[MCC04] Yung-Cheng Ma, Tien-Fu Chen, and Chung-Ping Chung. Branch-and-bound task allocation with task clustering-based pruning. Journal of Parallel and Distributed Computing, 64 (11):1223–1240, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

McDowell:1989:PAS

[McD89] Charles E. McDowell. A practical algorithm for static analysis of parallel programs. Journal of Parallel and Distributed Computing, 6(3):515–536, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Macedo:2006:RBA

[MCdS^{+}06] Daniel F. Macedo, Luiz H. A. Correia, Aldri L. dos Santos, Antonio A. F. Loureiro, and José Marcos S. Nogueira. A rulebased adaptive routing protocol for continuous data dissemination in WSNs. Journal of Parallel and Distributed Computing, 66(4):542–555, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mckinley:1994:MCS

[Mck94] P. K. Mckinley. Multicast communication in staircase multichannel networks. Journal of Parallel and Distributed Computing, 23(1):60–71, October 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1119/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1119/production/pdf.

Melab:2006:GCP

[MCT06] N. Melab, S. Cahon, and E-G. Talbi. Grid computing for parallel bioinspired algorithms. Journal of Parallel and Distributed Computing, 66(8):1052–1061, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Miller:1992:AQM

[MD92] Valerie A. Miller and George J. Davis. Adaptive quadrature on a message-passing multiprocessor. Journal of Parallel and Distributed Computing, 14(4):417–425, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Medidi:1998:PDU

[MD98] Muralidhar Medidi and Narsingh Deo. Parallel dictionaries using AVL trees. Journal of Parallel and Distributed Computing, 49(1):146–155, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1432/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1432/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1432/production/ref.

Mahapatra:2001:HEH

[MD01] Nihar R. Mahapatra and Shantanu Dutt. Hardware-efficient and highly reconfigurable 4- and 2-track fault-tolerant designs for mesh-connected arrays. Journal of Parallel and Distributed Computing, 61(10):1391–1411, October 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1702;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1702/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1702/ref.

Mahapatra:2007:EDO

[MD07] Nihar R. Mahapatra and Shantanu Dutt. An efficient delayoptimal distributed termination detection algorithm. Journal of Parallel and Distributed Computing, 67(10):1047–1066, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Meyer:1997:ALA

[MDD97] Trevor E. Meyer, James A. Davis, and Jennifer L. Davidson. Analysis of load average and its relationship to program run time on networks of workstations. Journal of Parallel and Distributed Computing, 44(2):141–146, August 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1355/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1355/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1355/production/ref.

Mahanti:2004:ADP

[ME04] Anirban Mahanti and Derek L. Eager. Adaptive data parallel computing on workstation clusters. Journal of Parallel and Distributed Computing, 64(11):1241–1255, November 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Megson:1991:SAB

[Meg91] G. M. Megson. Systolic algorithms for B-spline patch generation. Journal of Parallel and Distributed Computing, 11(3): 231–238, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Merrall:1996:PEN

[Mer96] Simon C. Merrall. Parallel execution of nested parallel expressions. Journal of Parallel and Distributed Computing, 33(2):122–130, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0031/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0031/production/pdf.

Mabbs:1994:PAM

[MF94] S. A. Mabbs and K. E. Forward. Performance analysis of MR-1, a clustered shared-memory multiprocessor. Journal of Parallel and Distributed Computing, 20(2):158–175, February 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1016/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1016/production/pdf.

Mourad:1993:RID

[MFS93] Antoine N. Mourad, Kent W. Fuchs, and Daniel G. Saab. Recovery issues in databases using redundant disk arrays. Journal of Parallel and Distributed Computing, 17(1–2):75–89, January/February 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1006/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1006/production/pdf.

Mourad:1996:SPR

[MFS96] Antoine N. Mourad, W. Kent Fuchs, and Daniel G. Saab. Site partitioning for redundant arrays of distributed disks. Journal of Parallel and Distributed Computing, 33(1):1–11, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0019/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0019/
production/pdf.

Mittal:2008:TDC

[MFVP08] Neeraj Mittal, Felix C. Freiling, S. Venkatesan, and Lucia Draque Penso. On termination detection in crash-prone distributed systems with failure detectors. Journal of Parallel and Distributed Computing, 68(6):855–875, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mowry:1991:TLT

[MG91] Todd Mowry and Anoop Gupta. Tolerating latency through software-controlled prefetching in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 12(2):87– 106, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Murdocca:1993:AIR

[MG93] Miles Murdocca and Vipul Gupta. Architectural implications of reconfigurable optical interconnects. Journal of Parallel and Distributed Computing, 17(3):200–211, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1019/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1019/production/pdf.

Moga:1998:PMB

[MG98] Alina N. Moga and Moncef Gabbouj. Parallel marker-based image segmentation with watershed transformation. Journal of Parallel and Distributed Computing, 51(1):27–45, May 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1448/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1448/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1448/production/ref.

Minkenberg:2009:DPS

[MG09] Cyriel Minkenberg and Mitchell Gusat. Design and performance of speculative flow control for high-radix datacenter interconnect switches. Journal of Parallel and Distributed Computing, 69(8):680–695, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Montoya:2003:LUP

[MGG03] Maria Dolores Gil Montoya, C. Gil, and I. García. The load unbalancing problem for region growing image segmentation algorithms. Journal of Parallel and Distributed Computing, 63 (4):387–395, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Monien:2006:I

[MGS^{+}06] Burkhard Monien, Guang Gao, Horst Simon, Paul Spirakis, and Per Stenström. Introduction. Journal of Parallel and Distributed Computing, 66(5):615–616, May 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mudge:1986:AMB

[MHBW86] T. N. Mudge, J. P. Hayes, G. D. Buzzard, and D. C. Winsor. Analysis of multiple-bus interconnection networks. Journal of Parallel and Distributed Computing, 3(3):328–343, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Memin:1995:EPN

[MHC95] E. Memin, F. Heitz, and F. Charot. Efficient parallel nonlinear multigrid relaxation algorithms for low-level vision applications. Journal of Parallel and Distributed Computing, 29(1):96–103, August 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1110/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1110/production/pdf.

Madisetti:1993:MEP

[MHF93] Vijay K. Madisetti, David A. Hardaker, and Richard M. Fujimoto. The MIMDIX environment for parallel simulation. Journal of Parallel and Distributed Computing, 18(4):473–483, August 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1078/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1078/
production/pdf.

McLendon:2005:FSC

[MHPR05] William McLendon III, Bruce Hendrickson, Steven J. Plimpton, and Lawrence Rauchwerger. Finding strongly connected components in distributed graphs. Journal of Parallel and Distributed Computing, 65(8):901–910, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Manabe:1992:GCD

[MI92] Yoshifumi Manabe and Makoto Imase. Global conditions in debugging distributed programs. Journal of Parallel and Distributed Computing, 15(1):62–69, May 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Miller:1993:WDW

[Mil93] Barton P. Miller. What to draw? when to draw? an essay on parallel program visualization. Journal of Parallel and Distributed Computing, 18(2):265–269, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1063/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1063/production/pdf.

Milosavljevi:1999:IAA

[Mil99] Igor Milosavljevi. An improved algorithm for array alignment. Journal of Parallel and Distributed Computing, 58(1): 44–67, July 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1531/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1531/
production/pdf.

Miranker:1991:SIP

[Mir91] Daniel P. Miranker. Special issue on the parallel execution of rule systems. Journal of Parallel and Distributed Computing, 13(4):345–347, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mitchell:2007:RTB

[Mit07] William F. Mitchell. A refinement-tree based partitioning method for dynamic load balancing with adaptively refined grids. Journal of Parallel and Distributed Computing, 67(4): 417–429, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Misic:1994:RFD

[MJ94] J. Misic and Z. Jovanovic. Routing function and deadlock avoidance in a star graph interconnection network. Journal of Parallel and Distributed Computing, 22(2):216–228, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1082/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1082/production/
pdf.

Milosavljevi:2001:EEA

[MJ01] Igor Z. Milosavljevi and Marwan A. Jabri. Experimental evaluation of automatic array alignment in parallelized Matlab. Journal of Parallel and Distributed Computing, 61(6):784–809, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1665;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1665/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1665/ref.

Marinescu:2003:CFS

[MJ03] Dan C. Marinescu and Yongchang Ji. A computational framework for the 3D structure determination of viruses with unknown symmetry. Journal of Parallel and Distributed Computing, 63(7–8):738–758, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mendelson:1992:EPP

[MK92] Bilha Mendelson and Israel Koren. Estimating the potential parallelism and pipelining of algorithms for data flow machines. Journal of Parallel and Distributed Computing, 14(1):15–28, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Melamed:2008:ASR

[MK08a] Roie Melamed and Idit Keidar. Araneola: a scalable reliable multicast system for dynamic environments. Journal of Parallel and Distributed Computing, 68(12):1539–1560, December 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mukherjee:2008:DPI

[MK08b] Sourav Mukherjee and Hillol Kargupta. Distributed probabilistic inferencing in sensor networks using variational approximation. Journal of Parallel and Distributed Computing, 68(1):78– 92, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Memik:2001:DES

[MKC01] Gokhan Memik, Mahmut T. Kandemir, and Alok Choudhary. Design and evaluation of a smart disk cluster for DSS commercial workloads. Journal of Parallel and Distributed Computing, 61(11):1633–1664, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1743;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1743/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1743/ref.

Mittal:2009:NDC

[MKC^{+}09] Neeraj Mittal, Srinivasan Krishnamurthy, R. Chandrasekaran, S. Venkatesan, and Yanyan Zeng. On neighbor discovery in cognitive radio networks. Journal of Parallel and Distributed Computing, 69(7):623–637, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Marchand:1997:OAD

[MKY^{+}97] Philippe J. Marchand, Ashok V. Krishnamoorthy, Gökçe I. Yayla, Sadik C. Esener, and Uzi Efron. Optically augmented 3-D computer: System technology and architecture. Journal of Parallel and Distributed Computing, 41(1):20–35, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1282/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1282/production/

pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1282/production/ref.

Maresca:1989:CAS

[ML89] Massimo Maresca and Hungwen Li. Connection autonomy in SIMD computers: a VLSI implementation. Journal of Parallel and Distributed Computing, 7(2):302–320, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Marinescu:1990:MMD

[MLC^{+}90] D. C. Marinescu, J. E. Lumpp, Jr., T. L. Casavant, Siegel, and H. J. Models for monitoring and debugging tools for parallel and distributed software. Journal of Parallel and Distributed Computing, 9(2):171–184, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mache:2005:ISL

[MLG05] Jens Mache, Virginia Lo, and Sharad Garg. The impact of spatial layout of jobs on I/O hotspots in mesh networks. Journal of Parallel and Distributed Computing, 65(10):1190–1203, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

May:1997:MLT

[MLW^{+}97] Phil May, Myunghee Lee, Scott T. Wilkinson, Olivier Vendier, Zhuang Ho, Steven W. Bond, D. Scott Wills, Martin Brooke, Nan M. Jokerst, and April Brown. A 100 mbps, LED through-wafer optoelectronic link for multicomputer interconnection networks. Journal of Parallel and Distributed Computing, 41(1):3–19, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1281/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1281/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1281/production/ref.

Manning:1993:AAE

[MM93] Serge M. Manning and David G. Meyer. Analysis of asynchronous execution streams with I-caching in massively parallel systems. Journal of Parallel and Distributed Computing, 19(3):279–291, November 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1112/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1112/production/pdf.

Mahapatra:2000:MNN

[MM00] Sudipta Mahapatra and Rabi N. Mahapatra. Mapping of neural network models onto systolic arrays. Journal of Parallel and Distributed Computing, 60(6):677–689, June 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1634;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1634/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1634/ref.

Mandal:2004:CCI

[MM04] P. S. Partha Sarathi Mandal and Krishnendu Mukhopadhyaya. Concurrent checkpoint initiation and recovery algorithms on asynchronous ring networks. Journal of Parallel and Distributed Computing, 64(5):649–661, May 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mandal:2006:PAD

[MM06] Partha Sarathi Mandal and Krishnendu Mukhopadhyaya. Performance analysis of different checkpointing and recovery schemes using stochastic model. Journal of Parallel and Distributed Computing, 66(1):99–107, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mandal:2007:SSA

[MM07a] Partha Sarathi Mandal and Krishnendu Mukhopadhyaya. Selfstabilizing algorithm for checkpointing in a distributed system. Journal of Parallel and Distributed Computing, 67(7):816–829, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Michailidis:2007:PAP

[MM07b] Panagiotis D. Michailidis and Konstantinos G. Margaritis. A programmable array processor architecture for flexible approximate string matching algorithms. Journal of Parallel and Distributed Computing, 67(2):131–141, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mittal:2007:PBD

[MM07c] Neeraj Mittal and Prajwal K. Mohan. A priority-based distributed group mutual exclusion algorithm when group access is non-uniform. Journal of Parallel and Distributed Computing, 67(7):797–815, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mendiburu:2006:PEC

[MMAL^{+}06] A. Mendiburu, J. Miguel-Alonso, J. A. Lozano, M. Ostra, and C. Ubide. Parallel EDAs to create multivariate calibration models for quantitative chemical applications. Journal of Parallel and Distributed Computing, 66(8):1002–1013, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Morajko:2007:DID

[MML07] Anna Morajko, Tomàs Margalef, and Emilio Luque. Design and implementation of a dynamic tuning environment. Journal of Parallel and Distributed Computing, 67(4):474–490, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mahesh:1998:SAF

[MMRS98] K. Mahesh, G. Manimaran, C. Siva Ram Murthy, and Arun K. Somani. Scheduling algorithms with fault detection and location capabilities for real-time multiprocessor systems. Journal of Parallel and Distributed Computing, 51(2):136–150, June 15, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1458/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1458/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1458/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1463/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1463/
production/pdf.

Moser:1990:WBA

[MMS90] L. E. Moser and P. M. Melliar-Smith. The world banker’s algorithm. Journal of Parallel and Distributed Computing, 9(4): 369–373, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Meyerhenke:2009:NDB

[MMS09] Henning Meyerhenke, Burkhard Monien, and Thomas Sauerwald. A new diffusion-based multilevel algorithm for computing graph partitions. Journal of Parallel and Distributed Computing, 69(9):750–761, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Manimaran:1997:NAR

[MMVR97] G. Manimaran, C. Siva Ram Murthy, Machiraju Vijay, and Krithi Ramamritham. New algorithms for resource reclaiming from precedence constrained tasks in multiprocessor real-time systems. Journal of Parallel and Distributed Computing, 44(2):123–132, August 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1354/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1354/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1354/production/ref.

Miller:1995:EDS

[MNB95] W. Marcus Miller, Walid A. Najjar, and A. P. Wim Bohm. Exploiting data structure locality in the dataflow model. Journal of Parallel and Distributed Computing, 27(2):183–200, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1081/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1081/production/
pdf.

Moreira:1998:DDD

[MNM98] José E. Moreira, Vijay K. Naik, and Samuel P. Midkiff. Dynamic data distribution and processor repartitioning for irregularly structured computations. Journal of Parallel and Distributed Computing, 50(1–2):28–60, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1436/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1436/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1436/production/ref.

Miller:1997:LBR

[MO97] D. A. B. Miller and H. M. Ozaktas. Limit to the bitrate capacity of electrical interconnects from the aspect ratio of the system architecture. Journal of Parallel and Distributed Computing, 41(1):42–52, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1285/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1285/production/pdf.

Mohapatra:1996:PAU

[Moh96] Prasant Mohapatra. Processor allocation using partitioning in mesh connected parallel computers. Journal of Parallel and Distributed Computing, 39(2):181–190, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0165/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0165/production/
pdf.

Mohapatra:1997:DRT

[Moh97] Prasant Mohapatra. Dynamic real-time task scheduling on hypercubes. Journal of Parallel and Distributed Computing, 46(1):91–100, October 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1387/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1387/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1387/production/ref.

Monfroglio:1994:NLC

[Mon94] Angelo Monfroglio. Neural logic constraint solving. Journal of Parallel and Distributed Computing, 20(1):92–98, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1009/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1009/production/
pdf.

Meyer:1987:PFO

[MP87] Gerard G. L. Meyer and Louis J. Podrazik. A parallel first-order linear recurrence solver. Journal of Parallel and Distributed Computing, 4(2):117–132, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mathias:1988:SEL

[MP88] P. C. Mathias and L. M. Patnaik. A systolic evaluator for linear, quadratic, and cubic expressions. Journal of Parallel and Distributed Computing, 5(6):729–740, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mayr:1993:PPP

[MP93] Ernst W. Mayr and C. Greg Plaxton. Pipelined parallel prefix computations, and sorting on a pipelined hypercube. Journal of Parallel and Distributed Computing, 17(4):374–380, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1037/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1037/production/
pdf.

Mickle:1996:LBU

[MP96] Marlin H. Mickle and JoAnn M. Paul. Load balancing using heterogeneous processors for continuum problems on a mesh. Journal of Parallel and Distributed Computing, 39(1):66–73, November 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0154/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0154/production/pdf.

Morris:2008:PLC

[MP08] Gerald R. Morris and Viktor K. Prasanna. A pipelinedloop-compatible architecture and algorithm to reduce variablelength sets of floating-point data on a reconfigurable computer. Journal of Parallel and Distributed Computing, 68(7):913–921, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Martinez-Perez:2009:PBA

[MPZ09] Israel Marck Martínez-Pérez and Karl-Heinz Zimmermann. Parallel bioinspired algorithms for NP complete graph problems. Journal of Parallel and Distributed Computing, 69(3): 221–229, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Marinescu:1994:SAP

[MR94a] D. C. Marinescu and J. R. Rice. On the scalability of asynchronous parallel computations. Journal of Parallel and Distributed Computing, 22(3):538–546, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1109/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1109/production/pdf.

Marinescu:1994:HLC

[MR94b] Dan C. Marinescu and John R. Rice. On high level characterization of parallelism. Journal of Parallel and Distributed Computing, 20(1):107–113, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1011/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1011/production/pdf.

Martel:1994:APM

[MR94c] C. Martel and A. Raghunathan. Asynchronous PRAMS with memory latency. Journal of Parallel and Distributed Computing, 23(1):10–26, October 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1115/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1115/production/pdf.

Mei:2003:RPP

[MR03] Alessandro Mei and Romeo Rizzi. Routing permutations in Partitioned Optical Passive Stars Networks. Journal of Parallel and Distributed Computing, 63(9):847–852, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Moses:2009:RSC

[MR09] Yoram Moses and Michel Raynal. Revisiting simultaneous consensus with crash failures. Journal of Parallel and Distributed Computing, 69(4):400–409, April 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Menon:1993:FPD

[MRK93] Jai Menon, James Roche, and Jim Kasson. Floating parity and data disk arrays. Journal of Parallel and Distributed Computing, 17(1–2):129–139, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1011/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1011/production/pdf.

Martin:2002:ASM

[MRR^{+}02] Dale E. Martin, Radharamanan Radhakrishnan, Dhananjai M. Rao, Malolan Chetlur, Krishnan Subramani, and Philip A. Wilsey. Analysis and simulation of mixedtechnology VLSI systems. Journal of Parallel and Distributed Computing, 62(3):468–493, March 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1805;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1805/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1805/ref.

Mostefaoui:2007:SBQ

[MRRT07] Achour Mostefaoui, Sergio Rajsbaum, Michel Raynal, and Corentin Travers. From ⋄W to Ω: a simple bounded quiescent reliable broadcast-based transformation. Journal of Parallel and Distributed Computing, 67(1):125–129, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Melin:1998:SSC

[MRRV98] Emmanuel Melin, Bruno Raffin, Xavier Rebeuf, and Bernard Virot. A structured synchronization and communication model fitting irregular data accesses. Journal of Parallel and Distributed Computing, 50(1–2):3–27, April/May 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1439/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1439/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1439/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1444/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1444/
production/pdf.

McMillen:1985:ECD

[MS85] Robert J. McMillen and Howard Jay Siegel. Evaluation of cube and data manipulator networks. Journal of Parallel and Distributed Computing, 2(1):79–107, February 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mirchandaney:1986:USL

[MS86] R. Mirchandaney and J. A. Stankovic. Using stochastic learning automata for job scheduling in distributed processing systems. Journal of Parallel and Distributed Computing, 3(4):527–552, December 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Makarenko:1987:VMM

[MS87] Darrell Makarenko and Jonathan Schaeffer. A VLSI multiprecision matrix multiplier and polynomial evaluator. Journal of Parallel and Distributed Computing, 4(6):619–628, December 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mikkilineni:1988:ESA

[MS88] Krishna P. Mikkilineni and Stanley Y. W. Su. An evaluation of sorting algorithms for common-bus local networks. Journal of Parallel and Distributed Computing, 5(1):59–81, February 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Makedon:1994:OAM

[MS94] F. Makedon and A. Symvonis. Optimal algorithms for multipacket routing problems on rings. Journal of Parallel and Distributed Computing, 22(1):37–43, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1068/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1068/production/pdf.

Moyer:1996:CCC

[MS96] Steven A. Moyer and V. S. Sunderam. Characterizing concurrency control performance for the PIOUS parallel file system. Journal of Parallel and Distributed Computing, 38(1):81–91, October 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0130/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0130/production/pdf.

Michael:1998:NAP

[MS98] Maged M. Michael and Michael L. Scott. Nonblocking algorithms and preemption-safe locking on multiprogrammed shared — memory multiprocessors. Journal of Parallel and Distributed Computing, 51(1):1–26, May 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1446/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1446/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1446/production/ref.

Manzini:1999:DPS

[MS99a] Gianmarco Manzini and Luca Stolcis. Distributed parallel strategies for industrial CFD solvers: a case study and analysis of performances. Journal of Parallel and Distributed Computing, 57(3):334–344, June 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1537/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1537/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1537/production/ref.

Myoupo:1999:TEP

[MS99b] Jean-Frédéric Myoupo and David Semé. Time-efficient parallel algorithms for the longest common subsequence and related problems. Journal of Parallel and Distributed Computing, 57(2):212–223, May 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1534/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1534/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1534/production/ref.

Mahadevan:2000:HAQ

[MS00] Indu Mahadevan and Krishna M. Sivalingam. A hierarchical architecture for QoS guarantees and routing in wireless/mobile networks. Journal of Parallel and Distributed Computing, 60(4):510–520, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1617;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1617/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1617/ref.

Manivannan:2002:ARU

[MS02] D. Manivannan and M. Singhal. Asynchronous recovery without using vector timestamps. Journal of Parallel and Distributed Computing, 62(12):1695–1728, December 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Macedo:2005:MGA

[MS05] Raimundo J. Araújo Macêdo and Flávio M. Assis Silva. The mobile groups approach for the coordination of mobile agents. Journal of Parallel and Distributed Computing, 65(3):275–288, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Majumdar:2004:PAC

[MSAF04] Shikharesh Majumdar, E-Kai Shen, and Istabrak Abdul-Fatah. Performance of adaptive CORBA middleware. Journal of Parallel and Distributed Computing, 64(2):201–218, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

McLay:1996:MSM

[MSC96] R. T. McLay, S. Swift, and G. F. Carey. Maximizing sparse matrix-vector product performance on RISC based MIMD computers. Journal of Parallel and Distributed Computing, 37(2):146–158, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0115/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0115/production/pdf.

Menasce:1995:SDP

[MSd^{+}95] Daniel A. Menasce, Debanjan Saha, Stella C. da Silva Porto, Virgilio A. F. Almeida, and Satish K. Tripathi. Static and dynamic processor scheduling disciplines in heterogeneous parallel architectures. Journal of Parallel and Distributed Computing, 28(1):1–18, July 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1085/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1085/production/pdf.

McLeod:1990:PAT

[MSH90] R. D. McLeod, J. J. Schellenberg, and P. D. Hortensius. Percolation and anomalous transport as tools in analyzing parallel processing interconnection networks. Journal of Parallel and Distributed Computing, 8(4):376–387, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mohamed:2005:MAR

[MSJ05] M. A. Maluk Mohamed, A. Vijay Srinivas, and D. Janakiram. Moset: An anonymous remote mobile cluster computing paradigm. Journal of Parallel and Distributed Computing, 65 (10):1212–1222, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Manoj:2009:SSS

[MSM09] B. S. Manoj, Archana Sekhar, and C. Siva Ram Murthy. A state-space search approach for optimizing reliability and cost of execution in distributed sensor networks. Journal of Parallel and Distributed Computing, 69(1):12–19, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Muhlenbein:1988:NPM

[MSS88] H. Muhlenbein, T. Schneider, and S. Streitz. Network programming with MUPPET. Journal of Parallel and Distributed Computing, 5(5):641–653, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Migliardi:2000:DJS

[MSS00] Mauro Migliardi, Simon Schubiger, and Vaidy Sunderam. A distributed JAVA SPACE implementation for HARNESS. Journal of Parallel and Distributed Computing, 60(10):1325– 1340, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2000.1656;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2000.

1656/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2000.1656/ref.

Middendorf:2002:EAD

[MSSE02] Martin Middendorf, Bernd Scheuermann, Hartmut Schmeck, and Hossam ElGindy. An evolutionary approach to dynamic task scheduling on FPGAs with restricted buffer. Journal of Parallel and Distributed Computing, 62(9):1407–1420, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1853;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1853/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1853/ref.

Mukherjee:1999:DSA

[MSST99] Sarit Mukherjee, Debanjan Saha, Manas C. Saksena, and Satish K. Tripathi. A distributed scheduling algorithm for real-time communication on slotted shared medium. Journal of Parallel and Distributed Computing, 58(1):1–25, July 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1545/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1545/production/pdf.

Merugu:2005:ASU

[MSZ05] Shashidhar Merugu, Sridhar Srinivasan, and Ellen Zegura. Adding structure to unstructured peer-to-peer networks: the use of small-world graphs. Journal of Parallel and Distributed Computing, 65(2):142–153, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Moldovan:1985:SVA

[MT85] Dan I. Moldovan and Yu-Wen Tung. Snap: a VLSI architecture for artificial intelligence processing. Journal of Parallel and Distributed Computing, 2(2):109–131, May 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ma:1993:EAM

[MT93a] Eva Ma and Lixin X. Tao. Embeddings among meshes and tori. Journal of Parallel and Distributed Computing, 18(1):44– 55, May 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1043/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1043/
production/pdf.

Malhotra:1993:RAR

[MT93b] Manish Malhotra and Kishor S. Trivedi. Reliability analysis of redundant arrays of inexpensive disks. Journal of Parallel and Distributed Computing, 17(1–2):146–151, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1013/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1013/production/
pdf.

Maillet:1995:EIG

[MT95] Eric Maillet and Cecile Tron. On efficiently implementing global time for performance evaluation on multiprocessor systems. Journal of Parallel and Distributed Computing, 28(1): 84–93, July 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1090/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1090/
production/pdf.

Matheson:1996:AMA

[MT96] Lesley R. Matheson and Robert E. Tarjan. Analysis of multigrid algorithms on massively parallel computers: Architectural implications. Journal of Parallel and Distributed Computing, 33(1):33–43, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0022/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0022/production/pdf.

Masuzawa:1997:AFC

[MT97a] Toshimitsu Masuzawa and Nobuki Tokura. An algorithm for finding the causal distributed breakpoint. Journal of Parallel and Distributed Computing, 42(1):60–66, April 10, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1301/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1301/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1301/production/ref.

Murtagh:1997:RBS

[MT97b] Paul J. Murtagh and Ah Chung Tsoi. A reconfigurable bitserial VLSI systolic array neuro-chip. Journal of Parallel and Distributed Computing, 44(1):53–70, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1343/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1343/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1343/production/ref.

Mirchandaney:1990:ALS

[MTS90] R. Mirchandaney, D. Towsley, and J. A. Stankovic. Adaptive load sharing in heterogeneous distributed systems. Journal of Parallel and Distributed Computing, 9(4):331–346, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Martinez:1988:APL

[MV88] Tony R. Martinez and Jacques J. Vidal. Adaptive parallel logic networks. Journal of Parallel and Distributed Computing, 5(1): 26–58, February 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Moonen:1994:JTS

[MV94] M. Moonen and J. Vandewalle. A Jacobi-Type systolic algorithm for Riccati and Lyapunov equations. Journal of Parallel and Distributed Computing, 22(2):314–320, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1089/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1089/production/pdf.

Min:2005:DPE

[MVB05] Wong Han Min, Bharadwaj Veeravalli, and Gerassimos Barlas. Design and performance evaluation of load distribution strategies for multiple divisible loads on heterogeneous linear daisy chain networks. Journal of Parallel and Distributed Computing, 65(12):1558–1577, December 2005. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). See comments [GRV08].

Mohan:2004:RRW

[MVM04] G. Mohan, G. Venkatesan, and C. Siva Ram Murthy. Randomized routing and wavelength requirements in wavelength-routed WDM multistage, hypercube, and de Bruijn networks. Journal of Parallel and Distributed Computing, 64(3):385–399, March 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Moonen:1991:JTA

[MVV91] Marc Moonen, Paul Van Dooren, and Joos Vandewalle. Jacobitype algorithms for LDU and Cholesky factorization. Journal of Parallel and Distributed Computing, 13(1):111–117, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mayr:1995:ORP

[MW95] Ernst W. Mayr and Ralph Werchner. Optimal routing of parentheses on the hypercube. Journal of Parallel and Distributed Computing, 26(2):181–192, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1058/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1058/production/pdf.

Ma:2000:JJE

[MWL00] Matchy J. M. Ma, Cho-Li Wang, and Francis C. M. Lau. JESSICA: Java-enabled single-system-image computing architecture. Journal of Parallel and Distributed Computing, 60(10):1194–1222, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1650;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1650/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1650/ref.

Yuan:1991:FTD

[mYA91] Shyan ming Yuan and Ashok K. Agrawala. Fault-tolerant decentralized commit protocols. Journal of Parallel and Distributed Computing, 13(3):299–311, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mohapatra:1995:LSS

[MYD95] Prasant Mohapatra, Chansu S. Yu, and Chita R. Das. A lazy scheduling scheme for hypercube computers. Journal of Parallel and Distributed Computing, 27(1):26–37, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1069/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1069/production/pdf.

Yeh:1992:ABF

[mYyF92] Yao ming Yeh and Tse yun Feng. Algorithm-based fault tolerance for matrix inversion with maximum pivoting. Journal of Parallel and Distributed Computing, 14(4):373–389, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nesterenko:2002:SPA

[NA02] Mikhail Nesterenko and Anish Arora. Stabilization-preserving atomicity refinement. Journal of Parallel and Distributed Computing, 62(5):766–791, May 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1828;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1828/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1828/ref.

Nagy:2006:CNN

[NA06] Marius Nagy and Selim G. Akl. Computing nearest neighbors in real time. Journal of Parallel and Distributed Computing, 66 (3):359–366, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nakano:1995:OIA

[Nak95] K. Nakano. Optimal initializing algorithms for a reconfigurable mesh. Journal of Parallel and Distributed Computing, 24(2):218–223, February 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1021/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1021/production/pdf.

Nishimura:2004:TDA

[NAK04] Naoki Nishimura, Yasuhiro Awatsuji, and Toshihiro Kubota. Two-dimensional arrangement of spatial patterns representing numerical data in input images for effective use of hardware resources in digital optical computing system based on optical array logic. Journal of Parallel and Distributed Computing, 64(9):1027–1040, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Napolitano:1990:DHP

[Nap90] Leonard M. Napolitano, Jr. The design of a high performance packet-switched network. Journal of Parallel and Distributed Computing, 10(2):103–114, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nassimi:1994:NLT

[Nas94]  David Nassimi.  Nearly logarithmic-time parallel algo-

rithms for the class of ±2^{b} ASCEND computations on a SIMD hypercube. Journal of Parallel and Distributed Computing, 20(3):289–302, March 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1028/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1028/production/pdf.

Nanda:1993:EMA

[NB93] Ashwini K. Nanda and Laxmi N. Bhuyan. Efficient mapping of applications on cache based multiprocessors. Journal of Parallel and Distributed Computing, 19(3):179–191, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1103/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1103/production/
pdf.

Najjar:1993:QAD

[NBM93] Walid A. Najjar, A. P. Wim Bohm, and W. Marcus Miller. A quantitative analysis of dataflow program execution — preliminaries to a hybrid design. Journal of Parallel and Distributed Computing, 18(3):314–326, July 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1067/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1067/production/pdf.

Nardelli:1998:DSM

[NBP98] Enrico Nardelli, Fabio Barillari, and Massimo Pepe. Distributed searching of multi-dimensional data: a performance evaluation study. Journal of Parallel and Distributed Computing, 49(1):111–134, February 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1428/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1428/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1428/production/ref.

Nagar:1999:ACN

[NBSD99] Shailabh Nagar, Ajit Banerjee, Anand Sivasubramaniam, and Chita R. Das. Alternatives to coscheduling a network of workstations. Journal of Parallel and Distributed Computing, 59(2):302–327, November 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1576/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1576/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1576/production/ref.

Nicol:1997:APD

[NC97] David M. Nicol and Gianfranco Ciardo. Automated parallelization of discrete state-space generation. Journal of Parallel and Distributed Computing, 47(2):153–167, December 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1409/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1409/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1409/production/ref.

Nghiem:2009:FBI

[NC09] Thao P. Nghiem and Tae Ho Cho. A fuzzy-based interleaved multi-hop authentication scheme in wireless sensor networks. Journal of Parallel and Distributed Computing, 69(5):441–450, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Natarajan:1993:PVM

[NCA93] Venkat Natarajan, Derek Chiou, and Boon Seong Ang. Performance visualization on Monsoon. Journal of Parallel and Distributed Computing, 18(2):169–180, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1054/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1054/production/pdf.

Navarro:2007:DLC

[NCT^{+}07] Angeles Navarro, Francisco Corbera, Adrian Tineo, Rafael Asenjo, and Emilio L. Zapata. Detecting loop-carried dependences in programs with dynamic data structures. Journal of Parallel and Distributed Computing, 67(1):47–62, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nistor:2009:OPC

[NCTT09] Adrian Nistor, Wei-Ngan Chin, Tiow-Seng Tan, and Nicolae Tapus. Optimizing the parallel computation of linear recurrences using compact matrix representations. Journal of Parallel and Distributed Computing, 69(4):373–381, April 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Noh:1999:HMM

[NDZA99] Sam H. Noh, Klaudia Dussa-Zieger, and Ashok K. Agrawala. A heterogeneous mixed-mode execution model for massively parallel systems. Journal of Parallel and Distributed Computing, 56(1):2–16, January 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1511/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1511/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1511/production/ref.

Nudd:1985:TDV

[NEG85] Graham R. Nudd, R. David Etchells, and Jan Grinberg. Threedimensional VLSI architecture for image understanding. Journal of Parallel and Distributed Computing, 2(1):1–29, February 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nastea:1997:LBS

[NFEG97] Sorin G. Nastea, Ophir Frieder, and Tarek El-Ghazawi. Load-balanced sparse matrix-vector multiplication on parallel computers. Journal of Parallel and Distributed Computing, 46(2):180–193, November 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1361/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1361/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1361/production/ref.

Ngom:2006:PES

[Ngo06] Alioune Ngom. Parallel evolution strategy on Grids for the protein threading problem. Journal of Parallel and Distributed Computing, 66(12):1489–1502, December 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nicol:1993:OPS

[NH93] David M. Nicol and Philip Heidelberger. Optimistic parallel simulation of continuous time Markov chains using uniformization. Journal of Parallel and Distributed Computing, 18(4):395–410, August 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1072/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1072/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1993.1073/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1073/production/pdf.

Nicolau:1988:LQG

[Nic88] A. Nicolau. Loop quantization: a generalized loop unwinding technique. Journal of Parallel and Distributed Computing, 5 (5):568–586, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nicol:1994:RPI

[Nic94] D. M. Nicol. Rectilinear partitioning of irregular data parallel computations. Journal of Parallel and Distributed Computing, 23(2):119–134, November 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1126/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1126/production/pdf.

Niculescu:2007:CEP

[Nic07] Virginia Niculescu. Cost-efficient parallel programs based on set-distributions for polynomial interpolation. Journal of Parallel and Distributed Computing, 67(8):935–946, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nieman:1994:IDC

[Nie94] D. E. Nieman. Issues in the design and control of parallel rule-firing production systems. Journal of Parallel and Distributed Computing, 23(3):338–363, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1145/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1145/production/pdf.

Nikolopoulos:2003:QCB

[Nik03] Dimitrios S. Nikolopoulos. Quantifying contention and balancing memory load on hardware DSM multiprocessors. Journal of Parallel and Distributed Computing, 63(9):866–886, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nikolopoulos:2004:PAH

[Nik04] Stavros D. Nikolopoulos. Parallel algorithms for Hamiltonian problems on quasi-threshold graphs. Journal of Parallel and Distributed Computing, 64(1):48–67, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ngai:1986:RAT

[NIR86] Tin-Fook Ngai, Mary Jane Irwin, and Shishpal Rawat. Regular, area-time efficient carry-lookahead adders. Journal of Parallel and Distributed Computing, 3(1):92–105, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Navaneethan:1991:DNS

[NJ91] P. Navaneethan and Lawrence Jenkins. Design of nonequivalent self-routing networks based on a matrix model. Journal of Parallel and Distributed Computing, 12(1):70–73, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nick:1997:PSS

[NKC^{+}97] Jeffrey M. Nick, Gary M. King, Jen-Yao Chung, Nicholas S. Bowen, and Ching-Shan Peng. Parallel sysplex: a scalable, highly available, high performance commercial system. Journal of Parallel and Distributed Computing, 43(2):179–189, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1342/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1342/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1342/production/ref.

Nicol:1995:APT

[NM95] David M. Nicol and Weizhen Z. Mao. Automated parallelization of timed Petri-Net simulations. Journal of Parallel and Distributed Computing, 29(1):60–74, August 15, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1106/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1106/production/pdf.

Nesterenko:2002:QBS

[NM02] Mikhail Nesterenko and Masaaki Mizuno. A quorum-based self-stabilizing distributed mutual exclusion algorithm. Journal of Parallel and Distributed Computing, 62(2):284–305, February 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1792;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1792/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1792/ref.

Nation:1993:MEC

[NMS93] Wayne G. Nation, Anthony A. Maciejewski, and Howard Jay Siegel. A methodology for exploiting concurrency among independent tasks in partitionable parallel processing systems.

Journal of Parallel and Distributed Computing, 19(3):271–278, November 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1111/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1111/
production/pdf.

Nogueira:2009:TBD

[NP09] Luís Nogueira and Luís Miguel Pinho. Time-bounded distributed QoS-aware service configuration in heterogeneous cooperative environments. Journal of Parallel and Distributed Computing, 69(6):491–507, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ng:1996:EPM

[NPI^{+}96] Cindy K. Y. Ng, Lawrence K. L. Pun, Dixon M. C. Ip, Mounir Hamdi, and Ishfaq Ahmad. Embedding pyramids into 3D meshes. Journal of Parallel and Distributed Computing, 36(2):173–184, August 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0097/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0097/production/pdf.

Nikolopoulos:2002:SAD

[NPP^{+}02] Dimitrios S. Nikolopoulos, Constantine D. Polychronopoulos, Theodore S. Papatheodorou, Jesús Labarta, and Eduard Ayguadé. Scheduler-activated dynamic page migration for multiprogrammed DSM multiprocessors. Journal of Parallel and Distributed Computing, 62(6):1069–1103, June 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1817;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1817/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1817/ref.

Nyland:1997:ASP

[NPY^{+}97] Lars Nyland, Jan Prins, Ru Huai Yun, Jan Hermans, Hye-Chung Kum, and Lei Wang. Achieving scalable parallel molecular dynamics using dynamic spatial domain decomposition techniques. Journal of Parallel and Distributed Computing, 47(2):125–138, December 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1408/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1408/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1408/production/ref.

Noeth:2009:SSC

[NRM^{+}09] Michael Noeth, Prasun Ratn, Frank Mueller, Martin Schulz, and Bronis R. de Supinski. ScalaTrace: Scalable compression and replay of communication traces for high-performance computing. Journal of Parallel and Distributed Computing, 69(8): 696–710, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nakanishi:1995:ESC

[NRS95] H. Nakanishi, V. Rego, and V. Sunderam. On the effectiveness of superconcurrent computations on heterogeneous networks. Journal of Parallel and Distributed Computing, 24(2):177–190, February 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1017/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1017/production/pdf.

Nation:1990:DPP

[NS90] W. G. Nation and H. J. Siegel. Disjoint path properties of the data manipulator network family. Journal of Parallel and Distributed Computing, 9(4):419–423, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nordstrom:1992:UDM

[NS92] Tomas Nordström and Bertil Svensson. Using and designing massively parallel computers for artificial neural networks. Journal of Parallel and Distributed Computing, 14(3):260–285, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nigam:1994:SNR

[NS94] M. Nigam and S. Sahni. Sorting n numbers on n × n reconfigurable meshes with buses. Journal of Parallel and Distributed Computing, 23(1):37–48, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1117/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1117/production/pdf.

Newman:1995:HPW

[NS95] Ilan Newman and Assaf Schuster. Hot potato worm routing via store-and-forward packet routing. Journal of Parallel and Distributed Computing, 30(1):76–84, October 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1127/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1127/production/pdf.

Neves:1997:TRS

[NS97] Richard Neves and Robert B. Schnabel. Threaded runtime support for execution of fine grain parallel code on coarse grain multiprocessors. Journal of Parallel and Distributed Computing, 42(2):128–142, May 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1322/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1322/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1322/production/ref.

Nam:1999:SLB

[NSLK99] Kyungwan Nam, Jaewon Seo, Sunggu Lee, and Jong Kim. Synchronous load balancing in hypercube multicomputers with faulty nodes. Journal of Parallel and Distributed Computing, 58(1):26–43, July 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1547/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1547/production/pdf.

Nastou:1998:ACA

[NSM98] Panagiotis E. Nastou, Dimitrios N. Serpanos, and Dimitrios G. Maritsas. Average case analysis of searching in associative processing. Journal of Parallel and Distributed Computing, 54(2):133–161, November 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1461/production;   http://www.idealibrary.

com/links/doi/10.1006/jpdc.1998.1461/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1461/production/ref.

No:2002:DIP

[NsPPC02] Jaechun No, Sung soon Park, Jesus Carretero Perez, and Alok Choudhary. Design and implementation of a parallel I/O runtime system for irregular applications. Journal of Parallel and Distributed Computing, 62(2):193–220, February 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1788;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1788/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1788/ref.

Naik:1997:PAM

[NSS97] Vijay K. Naik, Sanjeev K. Setia, and Mark S. Squillante. Processor allocation in multiprogrammed distributed-memory parallel computer systems. Journal of Parallel and Distributed Computing, 46(1):28–47, October 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1377/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1377/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1377/production/ref.

Narahari:1999:RST

[NSSS99] Bhagirath Narahari, Sunil Shende, Rahul Simha, and S. R. Subramanya. Routing and scheduling I/O transfers on wormhole-routed mesh networks. Journal of Parallel and Distributed Computing, 57(1):1–13, April 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1506/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1506/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1506/production/ref.

Nanda:1991:RCS

[NSTN91] Arun K. Nanda, Honda Shing, Ten-Hwan Tzen, and Lionel M. Ni. Resource contention in shared-memory multiprocessors: a parameterized performance degradation model. Journal of Parallel and Distributed Computing, 12(4):313–328, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ni:1990:SIS

[NT90] L. M. Ni and K. C. Tai. Special issue on software tools for parallel programming and visualization: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 9(2):101– 102, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nassimi:1993:EIB

[NT93] David Nassimi and Yuh-Dong D. Tsai. An efficient implementation of Batcher’s odd-even merge on a SIMD hypercube. Journal of Parallel and Distributed Computing, 19(1):58–63, September 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1090/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1090/
production/pdf.

Naimi:1996:LDM

[NTA96] Mohamed Naimi, Michel Trehel, and André Arnold. A log (N) distributed mutual exclusion algorithm based on path reversal. Journal of Parallel and Distributed Computing, 34(1):1– 13, April 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0041/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0041/
production/pdf.

No:2003:HPS

[NTC03] Jaechun No, Rajeev Thakur, and Alok Choudhary. Highperformance scientific data management system. Journal of Parallel and Distributed Computing, 63(4):434–447, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Numrich:2007:NSL

[Num07] Robert W. Numrich. A note on scaling the Linpack benchmark. Journal of Parallel and Distributed Computing, 67(4):491–498, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Numrich:2008:CFL

[Num08] Robert W. Numrich. Computational forces in the Linpack benchmark. Journal of Parallel and Distributed Computing, 68(9):1283–1290, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Numrich:2009:CFS

[Num09] Robert W. Numrich. Computational forces in the SAGE benchmark. Journal of Parallel and Distributed Computing, 69(3): 315–325, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nicol:1988:PSP

[NW88] D. M. Nicol and F. H. Willard. Problem size, parallel architecture, and optimal speedup. Journal of Parallel and Distributed Computing, 5(4):404–420, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oklobdzija:1988:IAV

[OB88] Vojin G. Oklobdzija and Earl R. Barnes. On implementing addition in VLSI technology. Journal of Parallel and Distributed Computing, 5(6):716–728, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oliker:1998:PPL

[OB98] Leonid Oliker and Rupak Biswas. PLUM: Parallel load balancing for adaptive unstructured meshes. Journal of Parallel and Distributed Computing, 52(2):150–177, August 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1469/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1469/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1469/production/ref.

Ohn:2007:PCC

[OC07] Kyungoh Ohn and Haengrae Cho. Path conscious caching of B^{+} tree indexes in a shared disks cluster. Journal of Parallel and Distributed Computing, 67(3):286–301, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ohring:1995:IHE

[OD95a] Sabine Ohring and Sajal K. Das. Incomplete hypercubes: Embeddings of tree-related networks. Journal of Parallel and Distributed Computing, 26(1):36–47, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1045/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1045/production/pdf.

OKeefe:1995:SBM

[OD95b] Matthew T. O’Keefe and Henry G. Dietz. Static barrier MIMD: Architecture and performance analysis. Journal of Parallel and Distributed Computing, 25(2):126–132, March 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1035/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1035/production/pdf.

Olariu:2007:AAN

[OEY07] Stephan Olariu, Mohamed Eltoweissy, and Mohamed Younis. ANSWER: AutoNomouS netWorked sEnsoR system. Journal of Parallel and Distributed Computing, 67(1):111–124, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

OBoyle:2003:TGE

[OFS03] M. F. P. O’Boyle, R. W. Ford, and E. A. Stohr. Towards general and exact distributed invalidation. Journal of Parallel and Distributed Computing, 63(11):1123–1137, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oudshoorn:2002:ETO

[OH02] Michael J. Oudshoorn and Lin Huang. Evolving toward an optimal scheduling solution through adaptivity. Journal of Parallel and Distributed Computing, 62(7):1203–1222, July 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1844;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1844/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1844/ref.

Ould-Khaoua:2001:DCS

[OK01] M. Ould-Khaoua. Distributed crossbar switch hypermeshes: Efficient networks for large-scale multicomputers. Journal of Parallel and Distributed Computing, 61(8):999–1012, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1739;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1739/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1739/ref.

OBoyle:2002:ILD

[OK02]  M. F. P. O’Boyle and P. M. W. Knijnenburg.  In-

tegrating loop and data transformations for global optimization. Journal of Parallel and Distributed Computing, 62(4):563–590, April 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1815;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1815/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1815/ref.

Oboyle:1995:SMS

[OKB95] M. F. P. O’Boyle, L. Kervella, and F. Bodin. Synchronization minimization in a SPMD execution model. Journal of Parallel and Distributed Computing, 29(2):196–210, September 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1117/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1117/production/
pdf.

Olariu:2001:PSI

[Ola01] Stephan Olariu. Preface: Special issue on wireless networks. Journal of Parallel and Distributed Computing, 61(7):853–854, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1715;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1715/pdf.

Opper:1984:RAM

[OM84] Eli Opper and Miroslaw Malek. On resource allocation in multistage interconnection network-based systems. Journal of Parallel and Distributed Computing, 1(2):206–220, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Olukotun:1990:HGA

[OM90] O. A. Olukotun and T. N. Mudge. Hierarchical gate-array routing on a hypercube multiprocessor. Journal of Parallel and Distributed Computing, 8(4):313–324, April 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ovalle-Martinez:2005:FMT

[OMSGNSG05] Francisco Javier Ovalle-Martínez, Ivan Stojmenović, Fabián García-Nocetti, and Julio Solano-González. Finding minimum transmission radii for preserving connectivity and constructing minimal spanning trees in ad hoc and sensor networks. Journal of Parallel and Distributed Computing, 65(2):132–141, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oruc:1985:ERA

[OO85] A. Yavuz Oruc and M. Yaman Oruc. Equivalence relations among interconnection networks. Journal of Parallel and Distributed Computing, 2(1):30–49, February 1985. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

OCearbhaill:2005:PIT

[OO05] Eoin A. O’Cearbhaill and Margaret O’Mahony. Parallel implementation of a transportation network model. Journal of Parallel and Distributed Computing, 65(1):1–14, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Olariu:1995:RBT

[OOW95] Stephan Olariu, Michael Overstreet, and Zhaofang F. Wen. Reconstructing a binary tree from its traversals in doubly logarithmic CREW time. Journal of Parallel and Distributed Computing, 27(1):100–105, May 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1075/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1075/production/pdf.

Ouyang:1996:VEC

[OP96] Pei Ouyang and Krishna V. Palem. Very efficient cyclic shifts on hypercubes. Journal of Parallel and Distributed Computing, 39(1):79–86, November 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0156/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0156/production/pdf.

Orlando:1998:CIS

[OP98] Salvatore Orlando and Raffaele Perego. A comparison of implementation strategies for nonuniform data-parallel computations. Journal of Parallel and Distributed Computing, 52(2):132–149, August 1, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1456/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1456/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1456/production/ref.

Oh:2008:OEO

[OPG08] Soon Y. Oh, Joon-Sang Park, and Mario Gerla. E-ODMRP: Enhanced ODMRP with motion adaptive refresh. Journal of Parallel and Distributed Computing, 68(8):1044–1053, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

orPanaite:2000:OBF

[oPP00] Petri or Panaite and Andrzej Pelc. Optimal broadcasting in faulty trees. Journal of Parallel and Distributed Computing, 60(5):566–584, May 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1625;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1625/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1625/ref.

Ou:1997:PRA

[OR97] Chao-Wei Ou and Sanjay Ranka. Parallel remapping of adaptive problems. Journal of Parallel and Distributed Computing, 42(2):109–121, May 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1314/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1314/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1314/production/ref.

Ortigosa:2003:PSP

[ORR03] E. M. Ortigosa, L. F. Romero, and J. I. Ramos. Parallel scheduling of the PCG method for banded matrices rising from FDM/FEM. Journal of Parallel and Distributed Computing, 63(12):1243–1256, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oruc:1987:DCP

[Oru87] A. Yavuz Oruc. Designing cellular permutation networks through coset decompositions of symmetric groups. Journal of Parallel and Distributed Computing, 4(4):404–422, August 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oruc:1994:SPN

[Oru94] A. Yavuz Oruc. A study of permutation networks: New designs and some generalizations. Journal of Parallel and Distributed Computing, 22(2):359–366, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1096/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1096/production/pdf.

Ofek:1993:DAH

[OS93] Yoram Ofek and Moshe Sidi. Design and analysis of a hybrid access control to an optical star using WDM. Journal of Parallel and Distributed Computing, 17(3):259–265, March 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1025/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1025/production/pdf.

Olariu:1996:NDS

[OS96a] Stephan Olariu and James L. Schwing. A novel deterministic sampling scheme with applications to broadcast-efficient sorting on the reconfigurable mesh. Journal of Parallel and Distributed Computing, 32(2):215–222, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0015/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0015/production/pdf.

Olariu:1996:TON

[OS96b] Stephan Olariu and Ivan Stojmenovič. Time-optimal nearestneighbor computations on enhanced meshes. Journal of Parallel and Distributed Computing, 36(2):144–155, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0095/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0095/production/
pdf.

Obenaus:1997:ESG

[OS97] Stefan Thomas Obenaus and Ted H. Szymanski. Embeddings of star graphs into optical meshes without bends. Journal of Parallel and Distributed Computing, 44(2):97–106, August 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1356/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1356/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1356/production/ref.

Ong:1998:ISM

[OS98] Soohwan Ong and Myung H. Sunwoo. Implementation of a sliding memory plane image processor. Journal of Parallel and Distributed Computing, 55(2):236–248, December 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1499/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1499/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1499/production/ref.

Oida:2004:PSA

[OS04] Kazumasa Oida and Shigeru Saito. A packet-size aware adaptive routing algorithm for parallel transmission server systems. Journal of Parallel and Distributed Computing, 64(1):36–47, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Oliveira:2005:PAH

[OSL05] Leonardo B. Oliveira, Isabela G. Siqueira, and Antonio A. F. Loureiro. On the performance of ad hoc routing protocols under a peer-to-peer application. Journal of Parallel and Distributed Computing, 65(11):1337–1347, November 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Olariu:1998:TOP

[OSZ98] Stephan Olariu, Ivan Stojmenovič, and Albert Y. Zomaya. Time-optimal proximity graph computations on enhanced meshes. Journal of Parallel and Distributed Computing, 49 (2):204–217, March 15, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1997.1421/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1421/production/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1421/production/ref.

Onaga:1986:WAL

[OT86] Kenji Onaga and Takahiro Takechi. A wavefront algorithm for LU decomposition of a partitioned matrix on VLSI processor arrays. Journal of Parallel and Distributed Computing, 3 (2):158–182, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Olariu:2001:GEI

[OW01] Stephan Olariu and Jie Wu. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 61(9):1145–1147, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1762;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1762/pdf.

Ofek:2000:CAS

[OY00] Yoram Ofek and Moti Yung. Combined asynchronous/ synchronous packet switching architecture: QoS guarantees for integrated parallel computing and real-time traffic. Journal of Parallel and Distributed Computing, 60(3):275–295, March 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1999.1607;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1607/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1607/ref.

Ozaktas:2004:IFI

[Oza04] Haldun M. Ozaktas. Information flow and interconnections in computing: extensions and applications of Rent’s rule. Journal of Parallel and Distributed Computing, 64(12):1360–1370, December 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Panwar:1994:MPS

[PA94] R. Panwar and G. Agha. A methodology for programming scalable architectures. Journal of Parallel and Distributed Computing, 22(3):479–487, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1105/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1105/production/pdf.

Pande:1996:SIC

[PA96] Santosh Pande and Dharma P. Agrawal. Special issue on compilation techniques for distributed memory systems: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 38(2):107–113, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0134/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0134/production/pdf.

Park:1997:EDF

[PA97] Hyunmin Park and Dharma P. Agrawal. Efficient deadlockfree wormhole routing and virtual-channel reduction in shufflebased networks. Journal of Parallel and Distributed Computing, 46(2):165–179, November 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1380/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1380/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1380/production/ref.

Park:2001:GDM

[PA01] Hyunmin Park and Dharma P. Agrawal. A generic design methodology for deadlock-free routing in multicomputer networks. Journal of Parallel and Distributed Computing, 61(9):1225–1248, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1748;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1748/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1748/ref.

Pinar:2004:FOL

[PA04] Ali Pınar and Cevdet Aykanat. Fast optimal load balancing algorithms for 1D partitioning. Journal of Parallel and Distributed Computing, 64(8):974–996, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Padmanabhan:1991:EAD

[Pad91] Krishnan Padmanabhan. Efficient architectures for data access in a shared memory hierarchy. Journal of Parallel and Distributed Computing, 11(4):314–327, April 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Padmanabhan:1993:SBA

[Pad93] Krishnan Padmanabhan. A shuffle-based alternative to the ADM interconnection architecture. Journal of Parallel and Distributed Computing, 18(1):14–24, May 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1040/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1040/production/pdf.

Plimpton:1998:PTD

[PAH^{+}98] Steve Plimpton, Steve Attaway, Bruce Hendrickson, Jeff Swegle, Courtenay Vaughan, and David Gardner. Parallel transient dynamics simulations: Algorithms for contact detection and smoothed particle hydrodynamics. Journal of Parallel and Distributed Computing, 50(1–2):104–122, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1433/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1433/production/pdf;

 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1433/production/ref.

Ponnuswamy:1997:PBB

[PAJC97] Subburajan Ponnuswamy, Minesh B. Amin, Rakesh Jha, and David A. Castañon. A C^{3}I parallel benchmark based on genetic algorithms — implementation and performance analysis. Journal of Parallel and Distributed Computing, 47(1):23–38, November 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1393/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1393/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1393/production/ref.

Pakzad:1989:FTA

[Pak89] S. Pakzad. Fault tolerance analysis of the class of rearrangeable interconnection networks. Journal of Parallel and Distributed Computing, 7(1):148–164, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pande:1994:TSS

[PAM94] Santoshkumar S. Pande, Dharma P. Agrawal, and Jon Mauney. A threshold scheduling strategy for Sisal on distributed memory machines. Journal of Parallel and Distributed Computing, 21(2):223–236, May 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1054/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1054/production/pdf.

Panda:2009:ICG

[Pan09] Dhabaleswar K. Panda. IPDPS 2007: Comments from the Guest Editor. Journal of Parallel and Distributed Computing, 69(8):679, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Parker:1989:CSS

[Par89] Jeff D. Parker. A concurrent search structure. Journal of Parallel and Distributed Computing, 7(2):256–278, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:1992:MED

[Par92] Haesun Park. On multiple error detection in matrix triangularizations using checksum methods. Journal of Parallel and Distributed Computing, 14(1):90–97, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:1996:PAG

[Par96] Taegeun Park. A parallel algorithm for global routing using an associative processor. Journal of Parallel and Distributed Computing, 38(1):51–62, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0128/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0128/production/pdf.

Parfenoff:1998:EPA

[Par98] I. Parfenoff. An efficient parallel algorithm for maximum matching for some classes of graphs. Journal of Parallel and Distributed Computing, 52(1):96–108, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1470/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1470/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1470/production/ref.

Parhami:2005:SIN

[Par05] Behrooz Parhami. Swapped interconnection networks: Topological, performance, and robustness attributes. Journal of Parallel and Distributed Computing, 65(11):1443–1452, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Patsouris:2001:AMA

[Pat01] Panos A. Patsouris. Algebraic modeling of an ad hoc network for mobile computing. Journal of Parallel and Distributed Computing, 61(7):884–897, July 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1718;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1718/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1718/ref.

Park:1990:RCC

[PB90] Arvin Park and K. Balasubramanian. Reducing communication costs for sorting on mesh-connected and linearly connected parallel computers. Journal of Parallel and Distributed Computing, 9(3):318–322, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pfluegl:1995:NIA

[PB95] Manfred J. Pfluegl and Douglas M. Blough. A new and improved algorithm for fault-tolerant clock synchronization. Journal of Parallel and Distributed Computing, 27(1):1–14, May 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1067/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1067/production/
pdf.

Pande:1999:CLB

[PB99] Santosh Pande and Tareq Bali. A computation+communication load balanced loop partitioning method for distributed memory systems. Journal of Parallel and Distributed Computing, 58(3):515–545, September 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1567/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1567/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1567/production/ref;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1580/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1580/
production/pdf.

Pussente:2009:ACS

[PB09] Rodolfo M. Pussente and Valmir C. Barbosa. An algorithm for clock synchronization with the gradient property in sensor networks. Journal of Parallel and Distributed Computing, 69 (3):261–265, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pedersen:2008:ABA

[PBS08] Jan Bækgaard Pedersen, Alex Brodsky, and Jeffrey Sampson. Approximating the buffer allocation problem using epochs.

Journal of Parallel and Distributed Computing, 68(9):1263– 1282, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Percus:1992:PAC

[PD92] Ora E. Percus and Susan R. Dickey. Performance analysis of clock-regulated queues with output multiplexing in three different 2 × 2 crossbar switch architectures. Journal of Parallel and Distributed Computing, 16(1):27–40, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Panda:2005:PRA

[PD05] B. S. Panda and Sajal K. Das. Parallel recognition algorithms for chordal planar graphs and planar k-trees. Journal of Parallel and Distributed Computing, 65(8):922–926, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:1993:AGS

[PE93] Haesun Park and L. Magnus Ewerbring. An algorithm for the generalized singular value decomposition on massively parallel computers. Journal of Parallel and Distributed Computing, 17(4):267–276, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1026/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1026/production/pdf.

Pinkston:1995:AOI

[PEC95] Timothy Mark Pinkston, Uzi Efron, and Michael Campbell. Applying optical interconnects to the 3-D computer: a performance evaluation. Journal of Parallel and Distributed Computing, 30(1):23–37, October 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1123/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1123/production/pdf.

Peleg:1990:TOL

[Pel90] David Peleg. Time-optimal leader election in general networks. Journal of Parallel and Distributed Computing, 8(1):96– 99, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Peleg:1995:NOT

[Pel95] David Peleg. A note on optimal time broadcast in faulty hypercubes. Journal of Parallel and Distributed Computing, 26(1):132–135, April 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1054/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1054/production/pdf.

Plateau:1991:MSM

[PF91] Brigitte Plateau and Jean-Michel Fourneau. A methodology for solving Markov models of parallel systems. Journal of Parallel and Distributed Computing, 12(4):370–387, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Prodan:2004:ZGM

[PF04] Radu Prodan and Thomas Fahringer. ZENTURIO: a grid middleware-based tool for experiment management of parallel and distributed applications. Journal of Parallel and Distributed Computing, 64(6):693–707, June 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pedone:2008:PHA

[PF08] Fernando Pedone and Svend Frølund. Pronto: High availability for standard off-the-shelf databases. Journal of Parallel and Distributed Computing, 68(2):150–164, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pfeiffer:1990:HLL

[Pfe90] Joseph J. Pfeiffer, Jr. HCL: a language for low-level image analysis. Journal of Parallel and Distributed Computing, 8(3): 231–244, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2004:GFS

[PFJ04] Jae H. Park, Gary Friedman, and Mark Jones. Geographical feature sensitive sensor placement. Journal of Parallel and Distributed Computing, 64(7):815–825, July 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Paul:2006:FSC

[PGS06] Himadri Sekhar Paul, Arobinda Gupta, and Amit Sharma. Finding a suitable checkpoint and recovery protocol for a distributed application. Journal of Parallel and Distributed Computing, 66(5):732–749, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Panda:1991:FDM

[PH91] Dhabaleswar K. Panda and Kai Hwang. Fast data manipulation in multiprocessors using parallel pipelined memories. Journal of Parallel and Distributed Computing, 12(2):130–145, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Parashar:2000:IPP

[PH00] Manish Parashar and Salim Hariri. Interpretive performance prediction for parallel application development. Journal of Parallel and Distributed Computing, 60(1):17–47, January 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1592/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1592/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1592/production/ref.

Palermo:1996:DDP

[PHB96] Daniel J. Palermo, Eugene W. Hodges IV, and Prithviraj Banerjee. Dynamic data partitioning for distributedmemory multicomputers. Journal of Parallel and Distributed Computing, 38(2):158–175, November 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0138/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0138/production/pdf.

Plimpton:2004:PRA

[PHS04] Steven J. Plimpton, Bruce Hendrickson, and James R. Stewart. A parallel rendezvous algorithm for interpolation between multiple grids. Journal of Parallel and Distributed Computing, 64(2):266–276, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Parker:1990:DAM

[PI90] J. R. Parker and T. R. Ingoldsby. Design and analysis of a multiprocessor for image processing. Journal of Parallel and Distributed Computing, 9(3):297–303, July 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Puente:2001:ABR

[PIB^{+}01] V. Puente, C. Izu, R. Beivide, J. A. Gregorio, F. Vallejo, and J. M. Prellezo. The adaptive bubble router. Journal of Parallel and Distributed Computing, 61(9):1180–1208, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1746;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1746/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1746/ref.

Piuri:2001:AFT

[Piu01] Vincenzo Piuri. Analysis of fault tolerance in artificial neural networks. Journal of Parallel and Distributed Computing, 61(1):18–48, January 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1663;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1663/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1663/ref.

Percus:1989:RNG

[PK89] Ora E. Percus and Malvin H. Kalos. Random number generators for MIMD parallel processors. Journal of Parallel and Distributed Computing, 6(3):477–497, June 1989. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Parhami:2004:IAC

[PK04a] Behrooz Parhami and Ding-Ming Kwai. Incomplete k-ary ncube and its derivatives. Journal of Parallel and Distributed Computing, 64(2):183–190, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2004:LPC

[PK04b] Jung-Heum Park and Hee-Chul Kim. Longest paths and cycles in faulty star graphs. Journal of Parallel and Distributed Computing, 64(11):1286–1296, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2005:OTS

[PK05a] Hee-Jun Park and Byung Kook Kim. Optimal task scheduling algorithm for cyclic synchronous tasks in general multiprocessor networks. Journal of Parallel and Distributed Computing, 65 (3):261–274, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2005:DVS

[PK05b] Inho Park and Seon Wook Kim. The distributed virtual sharedmemory system based on the InfiniBand architecture. Journal of Parallel and Distributed Computing, 65(10):1271–1280, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2005:SRD

[PK05c] Sangjoon Park and Byunggi Kim. Self-reproducible DEVS formalism. Journal of Parallel and Distributed Computing, 65(11): 1329–1336, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Plaku:2007:DCG

[PK07] Erion Plaku and Lydia E. Kavraki. Distributed computation of the k nn graph for large high-dimensional point sets. Journal of Parallel and Distributed Computing, 67(3):346–359, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Plank:1997:FTM

[PKD97] James S. Plank, Youngbae Kim, and Jack J. Dongarra. Faulttolerant matrix operations for networks of workstations using diskless checkpointing. Journal of Parallel and Distributed Computing, 43(2):125–138, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1336/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1336/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1336/production/ref.

Psarris:1991:ABT

[PKK91] Kleanthis Psarris, David Klappholz, and Xiangyun Kong. On the accuracy of the Banerjee test. Journal of Parallel and Distributed Computing, 12(2):152–157, June 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Page:2008:SDH

[PKN08] Andrew J. Page, Thomas M. Keane, and Thomas J. Naughton. Scheduling in a dynamic heterogeneous distributed system using estimation error. Journal of Parallel and Distributed Computing, 68(11):1452–1462, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Peir:1993:LAR

[PL93] Jih-Kwon K. Peir and Yann-Hang H. Lee. Look-ahead routing switches for multistage interconnection networks. Journal of Parallel and Distributed Computing, 19(1):1–10, September 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1085/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1085/production/
pdf.

Peng:1994:SOP

[PL94] Shietung T. Peng and Win-Tsung T. Lo. A simple optimal parallel algorithm for a core of a tree. Journal of Parallel and Distributed Computing, 20(3):388–392, March 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1035/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1035/production/pdf.

Petersen:1995:MCC

[PL95] Karin Petersen and Kai Li. Multiprocessor cache coherence based on virtual memory support. Journal of Parallel and Distributed Computing, 29(2):158–178, September 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1115/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1115/production/pdf.

Potter:1998:AAC

[PL98] Jerry L. Potter and Kathy J. Liszka. ACE: An associative calculus data structure. Journal of Parallel and Distributed Computing, 51(1):63–74, May 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1449/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1449/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1449/production/ref.

Peinado:2003:PGW

[PL03a] Marcus Peinado and Thomas Lengauer. Parallel “go with the winners” algorithms in distributed memory models. Journal of Parallel and Distributed Computing, 63(9):801–814, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Price:2003:CAF

[PL03b] Gregory W. Price and David K. Lowenthal. A comparative analysis of fine-grain threads packages. Journal of Parallel and Distributed Computing, 63(11):1050–1063, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2006:AAT

[PL06] Jae-Hyun Park and Heung-Kyu Lee. Algebraic analysis of the topological properties of a banyan network and its application in fault-tolerant switching networks. Journal of Parallel and Distributed Computing, 66(6):773–779, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Plaza:2008:PTI

[Pla08] Antonio J. Plaza. Parallel techniques for information extraction from hyperspectral imagery using heterogeneous networks of workstations. Journal of Parallel and Distributed Computing, 68(1):93–111, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Perrott:1987:DIP

[PLD87] R. H. Perrott, R. W. Lyttle, and P. S. Dhillon. The design and implementation of a Pascal-based language for array processor architectures. Journal of Parallel and Distributed Computing, 4(3):266–287, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Peltier:2003:TPA

[PLL^{+}03] Steven T. Peltier, Abel W. Lin, David Lee, Stephen Mock, Stephan Lamont, Tomas Molina, Mona Wong, Lu Dai, Maryann E. Martone, and Mark H. Ellisman. The Telescience Portal for advanced tomography applications. Journal of Parallel and Distributed Computing, 63(5):539–550, May 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Powell:2007:EOD

[PLR07] Olivier Powell, Pierre Leone, and José Rolim. Energy optimal data propagation in wireless sensor networks. Journal of Parallel and Distributed Computing, 67(3):302–317, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Provost:1992:DAE

[PM92] Foster J. Provost and Rami Melhem. A distributed algorithm for embedding trees in hypercubes with modifications for runtime fault tolerance. Journal of Parallel and Distributed Computing, 14(1):85–89, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Piriyakumar:1996:OCT

[PM96] D. Antony Louis Piriyakumar and C. Siva Ram Murthy. Optimal compile-time multiprocessor scheduling based on the 0–1 linear programming algorithm with the branch and bound technique. Journal of Parallel and Distributed Computing, 35(2):199–204, June 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0081/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0081/production/pdf.

Palmer:2005:OHP

[PM05] J. Palmer and I. Mitrani. Optimal and heuristic policies for dynamic server allocation. Journal of Parallel and Distributed Computing, 65(10):1204–1211, October 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pan:2005:SMK

[PMV05] Linqiang Pan and Carlos Martín-Vide. Solving multidimensional 0–1 knapsack problem by P systems with input and active membranes. Journal of Parallel and Distributed Computing, 65(12):1578–1584, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pan:2006:FRP

[PMV06] Linqiang Pan and Carlos Martín-Vide. Further remark on P systems with active membranes and two polarizations. Journal of Parallel and Distributed Computing, 66(6):867–872, June 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Panda:1997:SIWa

[PN97a] Dhabaleswar K. Panda and Lionel M. Ni. Special issue on workstation clusters and network-based computing: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 40(1):1–3, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1263/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1263/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1263/production/ref.

Panda:1997:SIWb

[PN97b] Dhabaleswar K. Panda and Lionel M. Ni. Special issue on workstation clusters and network-based computing: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 43(2):63–64, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1331/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1331/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1331/production/ref.

Poplawski:1991:SMD

[Pop91] David A. Poplawski. Synthetic models of distributed-memory parallel programs. Journal of Parallel and Distributed Computing, 12(4):423–426, August 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Povitsky:1999:PPA

[Pov99] A. Povitsky. Parallelization of pipelined algorithms for sets of linear banded systems. Journal of Parallel and Distributed Computing, 59(1):68–97, October 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1568/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1568/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1568/production/ref.

Prakash:1992:SSA

[PP92] Shiv Prakash and Alice C. Parker. SOS: Synthesis of application-specific heterogeneous multiprocessor systems. Journal of Parallel and Distributed Computing, 16(4):338–351, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pande:1996:PRV

[PP96] Santosh Pande and Kleanthis Psarris. Program repartitioning on varying communication cost parallel architectures. Journal of Parallel and Distributed Computing, 33(2):205–213, March 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0039/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0039/
production/pdf.

Paquette:2006:ODS

[PP06] Michel Paquette and Andrzej Pelc. Optimal decision strategies in Byzantine environments. Journal of Parallel and Distributed Computing, 66(3):419–427, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2004:DKB

[PPC04] Heejin Park, Kunsoo Park, and Yookun Cho. Deleting keys of B-trees in parallel. Journal of Parallel and Distributed Computing, 64(9):1041–1050, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pani:2006:SAA

[PR06] Danilo Pani and Luigi Raffo. Stigmergic approaches applied to flexible fault-tolerant digital VLSI architectures. Journal of Parallel and Distributed Computing, 66(8):1014–1024, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Prasad:1993:ESP

[Pra93] Sushil Prasad. Efficient and scalable PRAM algorithms for discrete-event simulation of bounded degree networks. Journal of Parallel and Distributed Computing, 18(4):524–530, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1083/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1083/production/
pdf.

Purtilo:1988:EPP

[PRG88] J. M. Purtilo, D. A. Reed, and D. C. Grunwald. Environments for prototyping parallel algorithms. Journal of Parallel and Distributed Computing, 5(4):421–437, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Park:2006:RPR

[PRHB06] Jungkeun Park, Minsoo Ryu, Seongsoo Hong, and Lucia Lo Bello. Rapid performance re-engineering of distributed embedded systems via latency analysis and k-level diagonal search. Journal of Parallel and Distributed Computing, 66(1):19–31, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pagani:2009:OOB

[PRP09] Elena Pagani, Gian Paolo Rossi, and Enrico Pertoso. ORION — Ontology-based queRy routIng in Overlay Networks. Journal of Parallel and Distributed Computing, 69(1):28–38, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Prakash:1997:ACO

[PRS97] Ravi Prakash, Michel Raynal, and Mukesh Singhal. An adaptive causal ordering algorithm suited to mobile computing environments. Journal of Parallel and Distributed Computing, 41(2):190–204, March 15, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1300/production;   http://www.idealibrary.

com/links/doi/10.1006/jpdc.1996.1300/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1300/production/ref.

Palis:1994:PRP

[PRW94] Michael A. Palis, Sanguthevar Rajasekaran, and David S. L. Wei. Packet routing and PRAM emulation on star graphs and leveled networks. Journal of Parallel and Distributed Computing, 20(2):145–157, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1015/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1015/production/pdf.

Purushothaman:1988:RAS

[PS88] S. Purushothaman and P. A. Subrahmanyam. Reasoning about systolic algorithms. Journal of Parallel and Distributed Computing, 5(6):669–699, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Peng:1993:NPM

[PS93] Dar-Tzen T. Peng and Kang G. Shin. A new performance measure for scheduling independent real-time tasks. Journal of Parallel and Distributed Computing, 19(1):11–26, September 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1086/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1086/production/
pdf.

Protopopov:2001:MMP

[PS01] Boris V. Protopopov and Anthony Skjellum. A multithreaded Message Passing Interface (MPI) architecture: Performance and program issues. Journal of Parallel and Distributed Computing, 61(4):449–466, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1674;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1674/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1674/ref.

Psarris:1996:BWG

[Psa96] Kleanthis Psarris. The Banerjee-Wolfe and GCD tests on exact data dependence information. Journal of Parallel and Distributed Computing, 32(2):119–138, February 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0009/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0009/production/pdf.

Prieto:2001:PMA

[PSE^{+}01] M. Prieto, R. Santiago, D. Espadas, I. M. Llorente, and F. Tirado. Parallel multigrid for anisotropic elliptic equations. Journal of Parallel and Distributed Computing, 61(1):96–114, January 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1666;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1666/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1666/ref.

Perez:2005:NFD

[PSPR05] María S. Pérez, Alberto Sánchez, José M. Peña, and Víctor Robles. A new formalism for dynamic reconfiguration of data servers in a cluster. Journal of Parallel and Distributed Computing, 65(10):1134–1145, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Prylli:1997:FRB

[PT97] Loïc Prylli and Bernard Tourancheau. Fast runtime block cyclic data redistribution on multiprocessors. Journal of Parallel and Distributed Computing, 45(1):63–72, August 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1351/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1351/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1351/production/ref.

Plank:2001:PAC

[PT01] James S. Plank and Michael G. Thomason. Processor allocation and checkpoint interval selection in cluster computing systems. Journal of Parallel and Distributed Computing, 61(11):1570–1590, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1757;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1757/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1757/ref.

Pinar:2008:ODP

[PTA08] Ali Pınar, E. Kartal Tabak, and Cevdet Aykanat. Onedimensional partitioning for heterogeneous systems: Theory and practice. Journal of Parallel and Distributed Computing, 68(11):1473–1486, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ponnusamy:1993:EPE

[PTC^{+}93] Ravi Ponnusamy, Rajeev Thakur, Alok Choudhary, Kishore Velamakanni, Zeki Bozkus, and Geoffrey Fox. Experimental performance evaluation of the CM-5. Journal of Parallel and Distributed Computing, 19(3):192–202, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1104/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1104/production/pdf.

Parmentier:2006:LSM

[PTZ06] Gilles Parmentier, Denis Trystram, and Jaroslaw Zola. Large scale multiple sequence alignment with simultaneous phylogeny inference. Journal of Parallel and Distributed Computing, 66 (12):1534–1545, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Peleg:1989:PDR

[PV89] David Peleg and Allen Van Gelder. Packet distribution on a ring. Journal of Parallel and Distributed Computing, 6(3):558– 567, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Petit:2007:OSS

[PV07] Franck Petit and Vincent Villain. Optimal snap-stabilizing depth-first token circulation in tree networks. Journal of Parallel and Distributed Computing, 67(1):1–12, January 2007. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Perez-Vidal:2009:HSF

[PVG09] Carlos Perez-Vidal and Luis Gracia. High speed filtering using reconfigurable hardware. Journal of Parallel and Distributed Computing, 69(11):896–904, November 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pontelli:2006:SST

[PVGG06] Enrico Pontelli, Karen Villaverde, Hai-Feng Guo, and Gopal Gupta. Stack splitting: a technique for efficient exploitation of search parallelism on share-nothing platforms. Journal of Parallel and Distributed Computing, 66(10):1267–1293, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Plaza:2006:CCB

[PVPM06] Antonio Plaza, David Valencia, Javier Plaza, and Pablo Martinez. Commodity cluster-based parallel processing of hyperspectral imagery. Journal of Parallel and Distributed Computing, 66(3):345–358, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Palaniswamy:1996:PTW

[PW96] Avinash C. Palaniswamy and Philip A. Wilsey. Parameterized Time Warp (PTW): An integrated adaptive solution to optimistic PDES. Journal of Parallel and Distributed Computing, 37(2):134–145, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0114/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0114/production/pdf.

Poulsen:1996:IFG

[PY96] David K. Poulsen and Pen-Chung Yew. Integrating finegrained message passing in cache coherent shared memory multiprocessors. Journal of Parallel and Distributed Computing, 33(2):172–188, March 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0036/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0036/production/pdf.

Pan:2009:BGM

[PY09a] Deng Pan and Yuanyuan Yang. Bandwidth guaranteed multicast scheduling for virtual output queued packet switches. Journal of Parallel and Distributed Computing, 69(12):939–949, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pan:2009:BML

[PY09b] Deng Pan and Yuanyuan Yang. Buffer management for lossless service in shared buffer switches. Journal of Parallel and Distributed Computing, 69(11):885–895, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Patarasuk:2009:BOA

[PY09c] Pitch Patarasuk and Xin Yuan. Bandwidth optimal all-reduce algorithms for clusters of workstations. Journal of Parallel and Distributed Computing, 69(2):117–124, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Patarasuk:2008:TPB

[PYF08] Pitch Patarasuk, Xin Yuan, and Ahmad Faraj. Techniques for pipelined broadcast on Ethernet switched clusters. Journal of Parallel and Distributed Computing, 68(6):809–824, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Qiu:2009:LSB

[QGL^{+}09] Meikang Qiu, Minyi Guo, Meiqin Liu, Chun Jason Xue, Laurence T. Yang, and Edwin H.-M. Sha. Loop scheduling and bank type assignment for heterogeneous multi-bank memory. Journal of Parallel and Distributed Computing, 69(6):546–558, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Quinn:1996:UCC

[QH96] Michael J. Quinn and Philip J. Hatcher. On the utility of communication-computation overlap in data-parallel programs. Journal of Parallel and Distributed Computing, 33(2):197–204, March 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0038/production;   http:

//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0038/
production/pdf.

Qiao:1997:TLP

[Qia97] Chunming Qiao. A two-level process for diagnosing crosstalk in photonic dilated Benes networks. Journal of Parallel and Distributed Computing, 41(1):53–66, February 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1284/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1284/production/pdf.

Qin:2005:DRD

[QJ05] Xiao Qin and Hong Jiang. A dynamic and reliability-driven scheduling algorithm for parallel real-time jobs executing on heterogeneous clusters. Journal of Parallel and Distributed Computing, 65(8):885–900, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Qiao:2001:CSC

[QM01] Chunming Qiao and Yousong Mei. A comparative study of cost effective multiplexing approaches for online permutation embedding and scheduling in optical networks. Journal of Parallel and Distributed Computing, 61(1):1–17, January 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1672;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1672/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1672/ref.

Qiao:1994:DRO

[QMCL94] Chunming M. Qiao, R. Melhem, D. Chiarulli, and S. Levitan. Dynamic reconfiguration of optically interconnected networks with time-division multiplexing. Journal of Parallel and Distributed Computing, 22(2):268–278, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1086/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1086/production/pdf.

Quintana-Orti:2001:SPA

[QOvdG01] Enrique S. Quintana-Ortí and Robert van de Geijn. Specialized parallel algorithms for solving Lyapunov and Stein equations. Journal of Parallel and Distributed Computing, 61(10):1489–1504, October 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1732;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1732/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1732/ref.

Quaglia:2005:MON

[QS05] Francesco Quaglia and Andrea Santoro. Modeling and optimization of non-blocking checkpointing for optimistic simulation on Myrinet clusters. Journal of Parallel and Distributed Computing, 65(6):667–677, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Qiu:2008:EML

[QSL^{+}08] Meikang Qiu, Edwin H.-M. Sha, Meilin Liu, Man Lin, Shaoxiong Hua, and Laurence T. Yang. Energy minimization with loop fusion and multi-functional-unit scheduling for multidimensional DSP. Journal of Parallel and Distributed Computing, 68(4):443–455, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Qian:1994:AML

[QY94] Xiaoshu S. Qian and Qing Yang. An analytical model for load balancing on symmetrical multiprocessor systems. Journal of Parallel and Distributed Computing, 20(2):198–211, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1019/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1019/production/
pdf.

Qi:1994:SDP

[QZ94] Ruijin J. Qi and S. A. Zenios. On the scalability of data-parallel decomposition algorithms for stochastic programs. Journal of Parallel and Distributed Computing, 22(3):565–570, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1112/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1112/production/
pdf.

Rahman:2008:RSS

[RAB08] Rashedur M. Rahman, Reda Alhajj, and Ken Barker. Replica selection strategies in Data Grid. Journal of Parallel and Distributed Computing, 68(12):1561–1574, December 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rajasekaran:1996:RSH

[Raj96] Sanguthevar Rajasekaran. Randomized selection on the hypercube. Journal of Parallel and Distributed Computing, 37(2):187–193, September 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0118/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0118/production/pdf.

Rajasekaran:2001:SAP

[Raj01] Sanguthevar Rajasekaran. Selection algorithms for parallel disk systems. Journal of Parallel and Distributed Computing, 61(4):536–544, April 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1682;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1682/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1682/ref.

Rajasekaran:2004:CCM

[Raj04] Sanguthevar Rajasekaran. Out-of-core computing on mesh connected computers. Journal of Parallel and Distributed Computing, 64(11):1311–1317, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rajasekaran:2008:SIP

[Raj08] Sanguthevar Rajasekaran. Special issue on parallel techniques for information extraction. Journal of Parallel and Distributed Computing, 68(1):1–2, January 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramarao:1989:DMI

[Ram89] K. V. S. Ramarao. Detection of mutual inconsistency in distributed databases. Journal of Parallel and Distributed Computing, 6(3):498–514, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramacher:1992:SNS

[Ram92] Ulrich Ramacher. Synapse — a neurocomputer that synthesizes neural algorithms on a parallel systolic engine. Journal of Parallel and Distributed Computing, 14(3):306–318, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ren:1996:MAP

[RAS96] Shangping Ren, Gul A. Agha, and Masahiko Saito. A modular approach to programming distributed real-time systems. Journal of Parallel and Distributed Computing, 36(1):4–12, July 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0085/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0085/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0086/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0086/production/pdf.

Roy:2008:BSP

[RBD08] Sasanka Roy, Debabrata Bardhan, and Sandip Das. Base station placement on boundary of a convex polygon. Journal of Parallel and Distributed Computing, 68(2):265–273, February 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Reddy:1993:DEG

[RCB93] A. L. Narasimha Reddy, John Chandy, and P. Banerjee. Design and evaluation of gracefully degradable disk arrays. Journal of Parallel and Distributed Computing, 17(1–2):28–40, January/February 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1003/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1003/production/pdf.

Rajasekaran:1997:UTS

[RCY97] Sanguthevar Rajasekaran, Wang Chen, and Shibu Yooseph. Unifying themes for selection on any network. Journal of Parallel and Distributed Computing, 46(1):105–111, October 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1382/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1382/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1382/production/ref.

Rajasekaran:2005:PRS

[RD05] Sanguthevar Rajasekaran and Jaime Davila. Packet routing and selection on the POPS network. Journal of Parallel and Distributed Computing, 65(8):927–933, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rhee:1995:MWD

[RDL95] C. Rhee, S. K. Dhall, and S. Lakshmivarahan. The minimum weight dominating set problem for permutation graphs is in NC. Journal of Parallel and Distributed Computing, 28(2):109–112, August 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1093/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1093/production/pdf.

Ravindran:2002:ARM

[RDS02] Binoy Ravindran, Ravi K. Devarasetty, and Behrooz Shirazi. Adaptive resource management algorithms for periodic tasks in dynamic real-time distributed systems. Journal of Parallel and Distributed Computing, 62(10):1527–1547, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1864;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1864/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1864/ref.

Reeves:1984:PPE

[Ree84] Anthony P. Reeves. Parallel Pascal: an extended Pascal for parallel computers. Journal of Parallel and Distributed Computing, 1(1):64–80, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rangarajan:1994:PDW

[RFM94] Sampath Rangarajan, Donald Fussell, and Miroslav Malek. Pipelined diagnosis of wafer-scale linear arrays. Journal of Parallel and Distributed Computing, 20(2):212–223, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1020/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1020/production/
pdf.

Ros:2008:TPI

[RFPAG08] Alberto Ros, Ricardo Fernández-Pascual, Manuel E. Acacio, and José M. García. Two proposals for the inclusion of directory information in the last-level private caches of glueless sharedmemory multiprocessors. Journal of Parallel and Distributed Computing, 68(11):1413–1424, November 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ritter:1987:IAT

[RG87] Gerhard X. Ritter and P. D. Gader. Image algebra techniques for parallel image processing. Journal of Parallel and Distributed Computing, 4(1):7–44, February 1987. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ro:2006:SPE

[RG06] Won W. Ro and Jean-Luc Gaudiot. Speculative pre-execution assisted by compiler (SPEAR). Journal of Parallel and Distributed Computing, 66(8):1076–1089, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramachandran:2003:EBQ

[RGD03] Vijaya Ramachandran, Brian Grayson, and Michael Dahlin. Emulations between QSM, BSP and LogP: a framework for general-purpose parallel algorithm design. Journal of Parallel and Distributed Computing, 63(12):1175–1192, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Reisman:2000:IRA

[RGS00] Amit Reisman, Craig Gotsman, and Assaf Schuster. Interactiverate animation generation by parallel progressive ray-tracing on distributed-memory machines. Journal of Parallel and Distributed Computing, 60(9):1074–1102, September 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1633;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2000.1633/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1633/ref.

Ruiz:2008:RCP

[RGU08] Antonio Ruiz, Nicolás Guil, and Manuel Ujaldón. Recognition of circular patterns on GPUs: Performance analysis and contributions. Journal of Parallel and Distributed Computing, 68(10): 1329–1338, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rodrigues:2000:DLW

[RGVB00] Luís Rodrigues, Katherine Guo, Paulo Veríssimo, and Kenneth P. Birman. A dynamic light-weight group service. Journal of Parallel and Distributed Computing, 60(12):1449–1479, December 1, 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2000.1645;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2000.1645/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2000.1645/ref.

Rahman:2005:RPE

[RH05] M. M. Hafizur Rahman and Susumu Horiguchi. Routing performance enhancement in hierarchical torus network by linkselection algorithm. Journal of Parallel and Distributed Computing, 65(11):1453–1461, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rodriguez:1996:POW

[RHH96] Bernardo Rodriguez, Leslie Hart, and Tom Henderson. Parallelizing operational weather forecast models for portable and fast execution. Journal of Parallel and Distributed Computing, 37(2):159–170, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0116/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0116/production/pdf.

Rostami:2008:SRS

[RHL08] Habib Rostami, Jafar Habibi, and Emad Livani. Semantic routing of search queries in P2P networks. Journal of Parallel and Distributed Computing, 68(12):1590–1602, December 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Richard:1998:EVT

[Ric98] Golden G. Richard III. Efficient vector time with dynamic process creation and termination. Journal of Parallel and Distributed Computing, 55(1):109–120, November 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1493/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1493/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1493/production/ref.

Rivera:1990:PSE

[RIZ90] F. F. Rivera, M. A. Ismail, and E. L. Zapata. Parallel squared error clustering on hypercube arrays. Journal of Parallel and Distributed Computing, 8(3):292–299, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rodriguez:1997:MTC

[RJA97] Bernardo Rodriguez, Harry Jordan, and Gita Alaghband. A metric for the temporal characterization of parallel programs. Journal of Parallel and Distributed Computing, 46(2):113–124, November 1, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1379/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1379/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1379/production/ref.

Robinson:1995:EMA

[RJMC95] David F. Robinson, Dan Judd, Philip K. McKinley, and Betty H. C. Cheng. Efficient multicast in all-port wormholerouted hypercubes. Journal of Parallel and Distributed Computing, 31(2):126–140, December 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1151/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1151/production/pdf.

Ray:1996:CDA

[RJY96] Sibabrata Ray, Hong Jiang, and Qing Yang. A compilerdirected approach to network latency reduction for distributed shared memory multiprocessors. Journal of Parallel and Distributed Computing, 38(2):267–276, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0147/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0147/production/pdf.

Rozier:1997:DEP

[RKK97] Richard G. Rozier, Fouad E. Kiamilev, and Ashok V. Krishnamoorthy. Design and evaluation of a photonic FFT processor. Journal of Parallel and Distributed Computing, 41(1):131–136, February 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1291/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1291/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1291/production/ref.

Rashad:2006:UMO

[RKK06] Sherif Rashad, Mehmed Kantardzic, and Anup Kumar. User mobility oriented predictive call admission control and resource reservation for next-generation mobile networks. Journal of Parallel and Distributed Computing, 66(7):971–988, July 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rotem:1987:ADA

[RKS87] D. Rotem, E. Korach, and N. Santoro. Analysis of a distributed algorithm for extrema finding in a ring. Journal of Parallel and Distributed Computing, 4(6):575–591, December 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Richards:1995:DCP

[RL95] Dana Richards and Arthur L. Liestman. Degree-constrained pyramid spanners. Journal of Parallel and Distributed Computing, 25(1):1–6, February 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1024/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1024/production/pdf.

Ramachandran:1996:CBS

[RL96] Umakishore Ramachandran and Joonwon Lee. Cache-based synchronization in shared memory multiprocessors. Journal of Parallel and Distributed Computing, 32(1):11–27, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0001/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0001/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0002/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0002/production/pdf.

Randall:2002:PIA

[RL02] Marcus Randall and Andrew Lewis. A parallel implementation of ant colony optimization. Journal of Parallel and Distributed Computing, 62(9):1421–1432, September 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1854;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1854/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1854/ref.

Ravindran:2003:PRA

[RLH03] Binoy Ravindran, Peng Li, and Tamir Hegazy. Proactive resource allocation for asynchronous real-time distributed systems in the presence of processor failures. Journal of Parallel and Distributed Computing, 63(12):1219–1242, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rouskov:1996:CFD

[RLS96] Yordan Rouskov, Shahram Latifi, and Pradip K. Srimani. Conditional fault diameter of star graph networks. Journal of Parallel and Distributed Computing, 33(1):91–97, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0028/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0028/production/
pdf.

Rego:1990:CET

[RM90] Vernon Rego and Aditya P. Mathur. Concurrency enhancement through program unification: a performance analysis. Journal of Parallel and Distributed Computing, 8(3):201–217, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Robinson:1997:PBM

[RMC97] David F. Robinson, Philip K. McKinley, and Betty H. C. Cheng. Path-based multicast communication in wormholerouted unidirectional torus networks. Journal of Parallel and Distributed Computing, 45(2):104–121, September 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1372/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1372/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1372/production/ref.

Rotaru:2004:DLB

[RN04] Tiberiu Rotaru and H.-H. Hans-Heinrich Nägeli. Dynamic load balancing by diffusion in heterogeneous systems. Journal of Parallel and Distributed Computing, 64(4):481–497, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Roh:1996:GOE

[RNSB96] Lucas Roh, Walid A. Najjar, Bhanu Shankar, and A. P. Wim Böhm. Generation, optimization, and evaluation of multithreaded code. Journal of Parallel and Distributed Computing, 32(2):188–204, February 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0013/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0013/production/pdf.

Rajasekaran:1992:CQR

[RO92] Sanguthevar Rajasekaran and Richard Overholt. Constant queue routing on a mesh. Journal of Parallel and Distributed Computing, 15(2):160–166, June 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Robert:2009:BPP

[Rob09] Yves Robert. Best papers and panel summary, IPDPS 2008. Journal of Parallel and Distributed Computing, 69(9):749, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rosenfeld:1985:PMA

[Ros85] Azriel Rosenfeld. The prism machine: an alternative to the pyramid. Journal of Parallel and Distributed Computing, 2(4): 404–411, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rosenfeld:1989:ACP

[Ros89] Azriel Rosenfeld. Arc colorings, partial path groups, and parallel graph contractions. Journal of Parallel and Distributed Computing, 7(2):335–354, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rosenberg:1999:GDP

[Ros99] Arnold L. Rosenberg. Guidelines for data-parallel cyclestealing in networks of workstations. I. on maximizing expected output. Journal of Parallel and Distributed Computing, 59(1):31–53, October 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1564/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1564/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1564/production/ref. See erratum [Ano00d].

Rosenberg:2007:BPA

[Ros07] Arnold L. Rosenberg. Best Paper Award recipients: 2006 International Parallel and Distributed Processing Symposium. Journal of Parallel and Distributed Computing, 67(12):1231, December 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ryang:1995:TLD

[RP95] Dal-Soo Ryang and Kyu Ho Park. A two-level distributed detection algorithm of AND/OR deadlocks. Journal of Parallel and Distributed Computing, 28(2):149–161, August 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1096/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1096/production/
pdf.

Raksapatcharawong:1998:MFS

[RP98] Mongkol Raksapatcharawong and Timothy Mark Pinkston. Modeling free-space optical k-ary n-cube wormhole networks. Journal of Parallel and Distributed Computing, 55(1):60– 93, November 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1998.1490/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1490/production/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1490/production/ref.

Reynolds:1993:DPA

[RPS93] Paul F. Reynolds, Jr., Carmen M. Pancerella, and Sudhir Srinivasan. Design and performance analysis of hardware support for parallel simulations. Journal of Parallel and Distributed Computing, 18(4):435–453, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1076/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1076/production/pdf.

Rajasekaran:1995:RAM

[RR95a] Sanguthevar Rajasekaran and Mukund Raghavachari. A randomized algorithm for multipacket routing on the mesh. Journal of Parallel and Distributed Computing, 26(2):257–260, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1064/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1064/
production/pdf.

Rajasekaran:1995:OMA

[RR95b] Sanguthevar Rajasekaran and Suneeta Ramaswami. Optimal mesh algorithms for the Voronoi diagram of line segments and motion planning in the plane. Journal of Parallel and Distributed Computing, 26(1):99–115, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1051/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1051/production/pdf.

Rauber:2005:TLS

[RR05] Thomas Rauber and Gudula Rünger. Tlib—a library to support programming with hierarchical multi-processor tasks. Journal of Parallel and Distributed Computing, 65(3):347–360, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ryoo:2008:POC

[RRS^{+}08] Shane Ryoo, Christopher I. Rodrigues, Sam S. Stone, John A. Stratton, Sain-Zee Ueng, Sara S. Baghsorkhi, and Wen mei W. Hwu. Program optimization carving for GPU computing. Journal of Parallel and Distributed Computing, 68(10): 1389–1401, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ranka:1990:ITM

[RS90a] S. Ranka and S. Sahni. Image template matching on MIMD hypercube multicomputers. Journal of Parallel and Distributed Computing, 10(1):79–84, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ranka:1990:SES

[RS90b] S. Ranka and S. Sahni. String editing on an SIMD hypercube multicomputer. Journal of Parallel and Distributed Computing, 9(4):411–418, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramanujam:1992:TMI

[RS92a] J. Ramanujam and P. Sadayappan. Tiling multidimensional iteration spaces for multicomputers. Journal of Parallel and Distributed Computing, 16(2):108–120, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Raschid:1992:PPS

[RS92b] Louiqa Raschid and Stanley Y. W. Su. A parallel pipelined strategy for evaluating linear recursive predicates in a multiprocessor environment. Journal of Parallel and Distributed Computing, 14(2):146–162, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ravindran:1992:FCB

[RS92c] K. Ravindran and S. Samdarshi. A flexible causal broadcast communication interface for distributed applications. Journal of Parallel and Distributed Computing, 16(2):134–157, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rego:1992:ECS

[RS92d] Vernon J. Rego and V. S. Sunderam. Experiments in concurrent stochastic simulation: The EcliPSe paradigm. Journal of Parallel and Distributed Computing, 14(1):66–84, January 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Raghavendra:1994:RPH

[RS94] C. S. Raghavendra and M. A. Sridhar. Routing permutations on hypercube machines with half-duplex links. Journal of Parallel and Distributed Computing, 20(1):14–19, January 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1002/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1002/production/pdf.

Raghavendra:1996:DOB

[RS96a] C. S. Raghavendra and M. A. Sridhar. Dimension ordering and broadcast algorithms in faulty SIMD hypercubes. Journal of Parallel and Distributed Computing, 35(1):57–66, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0068/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0068/production/
pdf.

Raghavendra:1996:ESD

[RS96b] C. S. Raghavendra and M. A. Sridhar. Exact solutions to diameter and routing problems in PEC networks. Journal of Parallel and Distributed Computing, 34(2):202–210, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0056/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0056/production/pdf.

Rajasekaran:2008:RSI

[RS08] Sanguthevar Rajasekaran and Mingjun Song. A relaxation scheme for increasing the parallelism in Jacobi–SVD. Journal of Parallel and Distributed Computing, 68(6):769–777, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ramaswamy:1996:OEA

[RSB96] Shankar Ramaswamy, Barbara Simons, and Prithviraj Banerjee. Optimizations for efficient array redistribution on distributed memory multicomputers. Journal of Parallel and Distributed Computing, 38(2):217–228, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0142/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0142/production/pdf.

Roh:2001:RMD

[RSBN01] Lucas Roh, Bhanu Shankar, Wim Böhm, and Walid Najjar. Resource management in dataflow-based multithreaded execution. Journal of Parallel and Distributed Computing, 61(5):581–608, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1708;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1708/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1708/ref.

Raschid:1994:SBS

[RSD94]  Louiqa Raschid, Timos Sellis, and Alexios Delis.  A

simulation-based study on the concurrent execution of rules in a database environment. Journal of Parallel and Distributed Computing, 20(1):20–42, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1003/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1003/production/pdf.

Ram:2004:MPP

[RSR04] D. Janaki Ram, A. Vijay Srinivas, and P. Manjula Rani. A model for parallel programming over CORBA. Journal of Parallel and Distributed Computing, 64(11):1256–1269, November 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ram:1996:PSA

[RSS96] D. Janaki Ram, T. H. Sreenivas, and K. Ganapathy Subramaniam. Parallel simulated annealing algorithms. Journal of Parallel and Distributed Computing, 37(2):207–212, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0121/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0121/production/
pdf.

Ramakrishnan:1999:ETN

[RSS99] Vara Ramakrishnan, Isaac D. Scherson, and Raghu Subramanian. Efficient techniques for nested and disjoint barrier synchronization. Journal of Parallel and Distributed Computing, 58(2):333–356, August 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1556/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1556/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1556/production/ref.

Rice:1990:SDP

[RSW90] M. D. Rice, S. B. Seidman, and P. Y. Wang. The specification of data parallel algorithms. Journal of Parallel and Distributed Computing, 8(2):191–195, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rosing:1991:DPP

[RSW91] Matthew Rosing, Robert B. Schnabel, and Robert P. Weaver. The DINO parallel programming language. Journal of Parallel and Distributed Computing, 13(1):30–42, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Rover:1991:SPA

[RTCG91] Diane Rover, Vicki Tsai, Yin-Shan Chow, and John Gustafson. Signal-processing algorithms on parallel architectures: a performance update. Journal of Parallel and Distributed Computing, 13(2):237–245, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Reddy:1999:RTC

[RU99] A. L. Narasimha Reddy and Eli Upfal. Real-time communication scheduling in a multicomputer video server. Journal of Parallel and Distributed Computing, 58(3):425–445, September 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1560/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1560/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1560/production/ref.

Rover:1993:VPS

[RW93] Diane T. Rover and Charles T. Wright, Jr. Visualizing the performance of SPMD and data-parallel programs. Journal of Parallel and Distributed Computing, 18(2):129–146, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1052/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1052/production/
pdf.

Rajasekaran:1997:SRS

[RW97] Sanguthevar Rajasekaran and David S. L. Wei. Selection, routing, and sorting on the star graph. Journal of Parallel and Distributed Computing, 41(2):225–233, March 15, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1297/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1297/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1297/production/ref.

Reisin:2001:INS

[RW01] Tamir G. Reisin and Sabine C. Wurzler. Implementation of a numerical solution of the Multicomponent Kinetic Collection Equation (MKCE) on parallel computers. Journal of Parallel and Distributed Computing, 61(5):641–661, May 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1685;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2000.1685/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1685/ref.

Rao:2002:ULS

[RW02] Dhananjai M. Rao and Philip A. Wilsey. An ultra-large-scale simulation framework. Journal of Parallel and Distributed Computing, 62(11):1670–1693, November 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ranka:1995:IPC

[RWK95] Sanjay Ranka, Jhy-Chun C. Wang, and Manoj Kumar. Irregular personalized communication on distributed memory machines. Journal of Parallel and Distributed Computing, 25(1):58–71, February 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1029/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1029/production/pdf.

Ranka:1993:EMS

[RWY93] Sanjay Ranka, Jhy-Chun C. Wang, and Nangkang Yeh. Embedding meshes on the star graph. Journal of Parallel and Distributed Computing, 19(2):131–135, October 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1098/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1098/production/pdf.

Sunwoo:1990:FCM

[SA90] Myung Hoon Sunwoo and J. K. Aggarwal. Flexibly coupled multiprocessors for image processing. Journal of Parallel and Distributed Computing, 10(2):115–129, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sterling:1993:FGD

[SA93] Thomas Lawrence Sterling and Jeffrey M. Arnold. Fine grain dataflow computation without tokens for balanced execution. Journal of Parallel and Distributed Computing, 18(3):327– 339, July 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1068/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1068/
production/pdf.

Sintorn:2008:FPG

[SA08] Erik Sintorn and Ulf Assarsson. Fast parallel GPU-sorting using a hybrid algorithm. Journal of Parallel and Distributed Computing, 68(10):1381–1388, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Siegel:1992:RPW

[SAB^{+}92] Howard Jay Siegel, Seth Abraham, William L. Bain, Kenneth E. Batcher, Thomas L. Casavant, Doug DeGroot, Jack B. Dennis, David C. Douglas, Tse-Yun Feng, James R. Goodman, Alan Huang, Harry F. Jordan, J. Robert Jump, Yale N. Patt, Alan J. Smith, James E. Smith, Lawrence Snyder, Harold S. Stone, Russ Tuck, and Benjamin W. Wah. Report of the Purdue workshop on grand challenges in computer architecture for the support of high performance computing. Journal of Parallel and Distributed Computing, 16(3):199–211, November 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sabot:1994:OCF

[Sab94] G. W. Sabot. Optimizing CM Fortran compiler for Connection Machine computers. Journal of Parallel and Distributed Computing, 23(2):224–238, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1134/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1134/production/pdf.

Shaw:1998:CIP

[SAC^{+}98] Andrew Shaw, Arvind, Kyoo-Chan Cho, Christopher Hill, R. Paul Johnson, and John Marshall. A comparison of implicitly parallel multithreaded and data-parallel implementations of an ocean model. Journal of Parallel and Distributed Computing, 48(1):1–51, January 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1390/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1390/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1390/production/ref.

Sevilgen:2005:PAT

[SAF05] Fatih E. Sevilgen, Srinivas Aluru, and Natsuhiko Futamura. Parallel algorithms for tree accumulations. Journal of Parallel and Distributed Computing, 65(1):85–93, January 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sandhu:1995:ADS

[San95] Harjinder S. Sandhu. Algorithms for dynamic software cache coherence. Journal of Parallel and Distributed Computing, 29(2):142–157, September 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1114/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1114/production/pdf.

Sanders:1998:RPQ

[San98] Peter Sanders. Randomized priority queues for fast parallel access. Journal of Parallel and Distributed Computing, 49(1):86–97, February 25, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1429/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1429/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1429/production/ref.

Santos:1999:ONO

[San99] Eunice E. Santos. Optimal and near-optimal algorithms for k-item broadcast. Journal of Parallel and Distributed Computing, 57(2):121–139, May 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1529/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1529/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1529/production/ref.

Santos:2002:OEA

[San02] Eunice E. Santos. Optimal and efficient algorithms for summing and prefix summing on parallel machines. Journal of Parallel and Distributed Computing, 62(4):517–543, April 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1698;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1698/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1698/ref.

Sela:1994:MPN

[SANY94] J. G. Sela, P. B. Anderson, D. W. Norton, and M. A. Young. Massive parallelization of NMC’s spectral model. Journal of Parallel and Distributed Computing, 21(1):140–149, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1047/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1047/production/
pdf.

Sarbazi-Azad:2003:AMW

[SAOKM03] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Analytical modelling of wormhole-routed k-ary n-cubes in the presence of matrix-transpose traffic. Journal of Parallel and Distributed Computing, 63(4):396–409, April 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarbazi-Azad:2002:PAL

[SAOKMA02] H. Sarbazi-Azad, M. Ould-Khaoua, L. M. Mackenzie, and S. G. Akl. A parallel algorithm for Lagrange interpolation on the star graph. Journal of Parallel and Distributed Computing, 62(4):605–621, April 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1812;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1812/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1812/ref.

Sarbazi-Azad:2005:DPNa

[SAOKZ05a] H. Sarbazi-Azad, M. Ould-Khaoua, and A. Y. Zomaya. Design and performance of networks for super-, cluster-, and Gridcomputing: Part I. Journal of Parallel and Distributed Computing, 65(10):1119–1122, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarbazi-Azad:2005:DPNb

[SAOKZ05b] H. Sarbazi-Azad, M. Ould-Khaoua, and A. Y. Zomaya. Design and performance of networks for super-, cluster-, and Gridcomputing: Part II. Journal of Parallel and Distributed Computing, 65(11):1301–1304, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Su:1984:DPM

[SB84] Stanley Y. W. Su and Chaitanya K. Baru. Dynamically partitionable multicomputers with switchable memory. Journal of Parallel and Distributed Computing, 1(2):152–184, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Skillicorn:1993:CP

[SB93] D. B. Skillicorn and D. T. Barnard. Compiling in parallel. Journal of Parallel and Distributed Computing, 17(4):337–352, April 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1032/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1032/production/
pdf.

Sohn:1997:SID

[SB97] Andrew Sohn and Rupak Biswas. Special issue on dynamic load balancing. Journal of Parallel and Distributed Computing, 47(2):99–101, December 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1406/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1406/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1406/production/ref.

Schwiebert:2002:PTA

[SB02] Loren Schwiebert and Renelius Bell. Performance tuning of adaptive wormhole routing through selection function choice. Journal of Parallel and Distributed Computing, 62(7):1121–1141, July 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1837;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1837/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1837/ref.

Sun:2004:MKN

[SB04] Xian-He Sun and A. R. Alan R. Blatecky. Middleware: the key to next generation computing. Journal of Parallel and Distributed Computing, 64(6):689–691, June 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stoyenko:1996:LBM

[SBAM96] Alexander D. Stoyenko, Jan Bosch, Mehmet Akšit, and Thomas J. Marlowe. Load balanced mapping of distributed objects to minimize network communication. Journal of Parallel and Distributed Computing, 34(2):117–136, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0050/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0050/production/pdf.

Suhler:1990:TTL

[SBKB90] P. A. Suhler, J. Biswas, K. M. Korner, and J. C. Browne. TDFL: A Task-Level Dataflow Language. Journal of Parallel and Distributed Computing, 9(2):103–115, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherson:1991:COE

[SC91a] Isaac D. Scherson and Peter F. Corbett. Communications overhead and the expected speedup of multidimensional meshconnected parallel processors. Journal of Parallel and Distributed Computing, 11(1):86–96, January 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Strevell:1991:DTT

[SC91b] Michael W. Strevell and Harvey G. Cragon. Data type transformation in heterogeneous shared memory multiprocessors. Journal of Parallel and Distributed Computing, 12(2):164–170, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Skillicorn:1995:CCP

[SC95] D. B. Skillicorn and W. Cai. A cost calculus for parallel functional programming. Journal of Parallel and Distributed Computing, 28(1):65–83, July 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1089/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1089/production/pdf.

Shen:2004:HPD

[SC04] Xiaohui Shen and Alok Choudhary. A high-performance distributed parallel file system for data-intensive computations. Journal of Parallel and Distributed Computing, 64(10):1157– 1167, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schenk:2008:APS

[SCB08] Olaf Schenk, Matthias Christen, and Helmar Burkhart. Algorithmic performance studies on graphics processing units. Journal of Parallel and Distributed Computing, 68(10):1360–1369, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shafi:2009:NPM

[SCB09] Aamir Shafi, Bryan Carpenter, and Mark Baker. Nested parallelism for multi-core HPC systems using Java. Journal of Parallel and Distributed Computing, 69(6):532–545, June 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sheu:1992:FTS

[SCC92] Jang-Ping Sheu, Yuh-Shyan Chen, and Chih-Yung Chang. Fault-tolerant sorting algorithm on hypercube multicomputers. Journal of Parallel and Distributed Computing, 16(2):185–197, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shao:2006:HSO

[SCC^{+}06] Zili Shao, Jiannong Cao, Keith C. C. Chan, Chun Xue, and Edwin H.-M. Sha. Hardware/software optimization for array & pointer boundary checking against buffer overflow attacks. Journal of Parallel and Distributed Computing, 66(9):1129– 1136, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Su:1999:BIW

[SCD99] Ming-Yang Su, Gen-Huey Chen, and Dyi-Rong Duh. Broadcasting on incomplete WK-recursive networks. Journal of Parallel and Distributed Computing, 57(3):271–294, June 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1538/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1538/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1538/production/ref.

Schwandt:1987:IAB

[Sch87] Hartmut Schwandt. Interval arithmetic block cyclic reduction on vector computers. Journal of Parallel and Distributed Computing, 4(5):459–487, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schaeffer:1989:DGT

[Sch89a] J. Schaeffer. Distributed game-tree searching. Journal of Parallel and Distributed Computing, 6(1):90–114, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schroder:1989:TDI

[Sch89b] H. Schroder. Top-down designs of instruction systolic arrays for polynomial interpolation and evaluation. Journal of Parallel and Distributed Computing, 6(3):692–703, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schaefer:1990:SIM

[Sch90] David H. Schaefer. Special issue on massively parallel computation: Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 8(2):101, February 1990. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schmolze:1991:GSR

[Sch91] James G. Schmolze. Guaranteeing serializable results in synchronous parallel production systems. Journal of Parallel and Distributed Computing, 13(4):348–365, December 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shin:2008:TSA

[SCJ^{+}08] KwangSik Shin, MyongJin Cha, MunSuck Jang, JinHa Jung, WanOh Yoon, and SangBang Choi. Task scheduling algorithm using minimized duplications in homogeneous systems. Journal of Parallel and Distributed Computing, 68(8):1146–1156, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:2003:MOC

[SCK03] W. Shi, E. Collins, and V. Karamcheti. Modeling object characteristics of dynamic Web content. Journal of Parallel and Distributed Computing, 63(10):963–980, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Salisbury:1999:MCL

[SCM99] C. Salisbury, Z. Chen, and R. Melhem. Modeling communication locality in multiprocessors. Journal of Parallel and Distributed Computing, 56(2):71–98, February 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1513/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1513/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1513/production/ref.

Saltz:1990:RTS

[SCMB90] Joel Saltz, Kathleen Crowley, Ravi Mirchandaney, and Harry Berryman. Run-time scheduling and execution of loops on message passing machines. Journal of Parallel and Distributed Computing, 8(4):303–312, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shestak:2008:HBB

[SCS^{+}08] Vladimir Shestak, Edwin K. P. Chong, Howard Jay Siegel, Anthony A. Maciejewski, Lotfi Benmohamed, I-Jeng Wang, and Rose Daley. A hybrid branch-and-bound and evolutionary approach for allocating strings of applications to heterogeneous distributed computing systems. Journal of Parallel and Distributed Computing, 68(4):410–426, April 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smith:1988:ASD

[SD88a] S. G. Smith and P. B. Denyer. Advanced serial-data computation. Journal of Parallel and Distributed Computing, 5(3):228– 249, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stewart:1988:SAP

[SD88b] Charles V. Stewart and Charles R. Dyer. Scheduling algorithms for PIPE (Pipelined Image-Processing Engine). Journal of Parallel and Distributed Computing, 5(2):131–153, April 1988. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scheurich:1991:LFC

[SD91] C. Scheurich and M. Dubois. Lockup-free caches in highperformance multiprocessors. Journal of Parallel and Distributed Computing, 11(1):25–36, January 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Skeppstedt:2000:CCP

[SD00]  Jonas Skeppstedt and Michel Dubois.  Compiler con-

trolled prefetching for multiprocessors using low-overhead traps and prefetch engines. Journal of Parallel and Distributed Computing, 60(5):585–615, May 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1623;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1623/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1623/ref.

Stantchev:2008:FPP

[SDG08] George Stantchev, William Dorland, and Nail Gumerov. Fast parallel particle-to-grid interpolation for plasma PIC simulations on the GPU. Journal of Parallel and Distributed Computing, 68(10):1339–1349, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schoneveld:1997:TAP

[SdS97] A. Schoneveld, J. F. de Ronde, and P. M. A. Sloot. Task allocation by parallel evolutionary computing. Journal of Parallel and Distributed Computing, 47(1):91–97, November 25, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1398/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1398/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1398/production/ref.

Skeppstedt:1999:ECC

[SDS99] Jonas Skeppstedt, Fredrik Dahlgren, and Per Stenström. Evaluation of compiler-controlled updating to reduce coherencemiss penalties in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 56(2):122–143, February 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1510/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1510/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1510/production/ref.

Salceda:2004:MAD

[SDTD04] Jesús Salceda, Iván Díaz, Juan Touriño, and R. Ramón Doallo. A middleware architecture for distributed systems management. Journal of Parallel and Distributed Computing, 64(6): 759–766, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Seban:1991:FTP

[Seb91] Robert R. Seban. FTN topology and protocols. Journal of Parallel and Distributed Computing, 11(1):51–62, January 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Seban:1995:DCS

[Seb95] Robert R. Seban. A distributed critical section protocol for general topology. Journal of Parallel and Distributed Computing, 28(1):32–42, July 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1087/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1087/production/pdf.

Seiferas:2005:NSM

[Sei05] Joel Seiferas. Networks for sorting multitonic sequences. Journal of Parallel and Distributed Computing, 65(12):1601–1606, December 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Severance:1996:MOB

[SEP96] Charles Severance, Richard Enbody, and Paul Petersen. Managing the overall balance of operating system threads on a multiprocessor using automatic self-allocating threads (ASAT). Journal of Parallel and Distributed Computing, 37(1):106–112, August 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0111/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0111/production/pdf.

Seredynski:1997:CCM

[Ser97] Franciszek Seredynski. Competitive coevolutionary multiagent systems: The application to mapping and scheduling problems. Journal of Parallel and Distributed Computing, 47(1):39–57, November 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1394/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1394/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1394/production/ref.

Stotts:1990:BPP

[SF90] P. David Stotts and Richard Furuta. Browsing parallel process networks. Journal of Parallel and Distributed Computing, 9 (2):224–235, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shmueli:2005:BLO

[SF05] Edi Shmueli and Dror G. Feitelson. Backfilling with lookahead to optimize the packing of parallel jobs. Journal of Parallel and Distributed Computing, 65(9):1090–1107, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smith:2006:CST

[SFEF06] Matthew Smith, Thomas Friese, Michael Engel, and Bernd Freisleben. Countering security threats in service-oriented ondemand Grid computing using sandboxing and trusted computing techniques. Journal of Parallel and Distributed Computing, 66(9):1189–1204, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smith:2004:PAR

[SFT04] Warren Smith, Ian Foster, and Valerie Taylor. Predicting application run times with historical information. Journal of Parallel and Distributed Computing, 64(9):1007–1016, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarukkai:1993:SPD

[SG93] Sekhar R. Sarukkai and Dennis Gannon. SIEVE: a performance debugging environment for parallel programs. Journal of Parallel and Distributed Computing, 18(2):147–168, June 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1053/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1053/production/pdf.

Sundaresan:1996:COO

[SG96] Neelakantan Sundaresan and Dennis Gannon. Coir: An objectoriented system for control and dynamic data parallelism. Journal of Parallel and Distributed Computing, 37(1):98–105, August 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0110/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0110/
production/pdf.

Sanchez:1999:SDP

[SG99] Jesús Sánchez and Antonio González. Software data prefetching for software pipelined loops. Journal of Parallel and Distributed Computing, 58(2):236–259, August 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1553/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1553/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1553/production/ref.

Schwiegelshohn:1991:OPA

[SGE91] Uwe Schwiegelshohn, Franco Gasperoni, and Kemal Ebcioglu. On optimal parallelization of arbitrary loops. Journal of Parallel and Distributed Computing, 11(2):130–134, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saxena:2003:DOC

[SGR03] P. C. Saxena, Sangita Gupta, and Jagmohan Rai. A delay optimal coterie on the k-dimensional folded Petersen graph. Journal of Parallel and Distributed Computing, 63(11):1026– 1035, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sibeyn:1999:PPL

[SGS99] Jop F. Sibeyn, Frank Guillaume, and Tillmann Seidel. Practical parallel list ranking. Journal of Parallel and Distributed Computing, 56(2):156–180, February 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1508/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1508/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1508/production/ref.

Sinha:2008:FDB

[SGS08] Koushik Sinha, Suranjan Ghose, and Pradip K. Srimani. Fast deterministic broadcast and gossiping algorithms for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 68(7):922–938, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Szymanski:1989:UMM

[SH89] T. H. Szymanski and V. C. Hamacher. On the universality of multipath multistage interconnection networks. Journal of Parallel and Distributed Computing, 7(3):541–569, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sammur:1990:MSP

[SH90] N. M. Sammur and M. T. Hagan. Mapping signal processing algorithms on parallel architectures. Journal of Parallel and Distributed Computing, 8(2):180–185, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shirazi:1992:SIS

[SH92a] Behrooz Shirazi and A. R. Hurson. Special issue on scheduling and load balancing: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 16(4):271–275, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Singh:1992:FEP

[SH92b] Jaswinder Pal Singh and John L. Hennessy. Finding and exploiting parallelism in an ocean simulation program: Experience, results, and implications. Journal of Parallel and Distributed Computing, 15(1):27–48, May 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarnath:1997:PSS

[SH97] R. Sarnath and Xin He. On parallel selection and searching in partial orders: Sorted matrices. Journal of Parallel and Distributed Computing, 40(2):242–247, February 1, 1997. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1258/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1258/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1258/production/ref.

Szymanski:1998:ATF

[SH98] Ted H. Szymanski and H. Scott Hinton. Architecture of a terabit free-space intelligent optical backplane. Journal of Parallel and Distributed Computing, 55(1):1–31, November 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1489/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1489/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1489/production/ref.

Song:1993:DAO

[SHC93] Q. Wang Song, Salim Hariri, and Alok Choudhary. Design and analysis of an optical communications processor. Journal of Parallel and Distributed Computing, 17(3):222–229, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1021/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1021/production/
pdf.

Shen:2006:PSL

[She06] Kai Shen. Parallel sparse LU factorization on different message passing platforms. Journal of Parallel and Distributed Computing, 66(11):1387–1403, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2009:PBI

[She09] Haiying Shen. A P2P-based intelligent resource discovery mechanism in Internet-based distributed systems. Journal of Parallel and Distributed Computing, 69(2):197–209, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:1995:EAC

[SHL95] Xiaojun J. Shen, Qing Hu, and Weifa F. Liang. Embedding K-ary complete trees into hypercubes. Journal of Parallel and Distributed Computing, 24(1):100–106, January 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1010/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1010/production/pdf.

Shi:2009:PAS

[SHLN09] Xiaoqiu Shi, Jinsong Han, Yunhao Liu, and Lionel M. Ni. Popularity adaptive search in hybrid P2P systems. Journal of Parallel and Distributed Computing, 69(2):125–134, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Singh:1995:LBD

[SHT^{+}95] Jaswinder Pal Singh, Chris Holt, Takashi Totsuka, Anoop Gupta, and John Hennessy. Load balancing and data locality in adaptive hierarchical N-body methods: Barnes– Hut, fast multipole, and radiosity. Journal of Parallel and Distributed Computing, 27(2):118–141, June 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1077/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1077/production/pdf.

Stone:2008:AAM

[SHT^{+}08] S. S. Stone, J. P. Haldar, S. C. Tsao, W. m.W. Hwu, B. P. Sutton, and Z.-P. Liang. Accelerating advanced MRI reconstructions on GPUs. Journal of Parallel and Distributed Computing, 68(10):1307–1318, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shu:1995:PIS

[Shu95] Wei Shu. Parallel implementation of a sparse simplex algorithm on MIMD distributed memory computers. Journal of Parallel and Distributed Computing, 31(1):25–40, November 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1142/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1142/production/
pdf.

Sado:1986:SPS

[SI86] Kazuhiro Sado and Yoshihide Igarashi. Some parallel sorts on a mesh-connected processor array and their time efficiency. Journal of Parallel and Distributed Computing, 3(3):398–410, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherson:1989:RFP

[SI89] I. D. Scherson and S. Ilgen. A reconfigurable fully parallel associative processor. Journal of Parallel and Distributed Computing, 6(1):69–89, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smitley:1991:BSS

[SI91] David Smitley and Kent Iobst. Bit-serial SIMD on the CM-2 and the Cray-2. Journal of Parallel and Distributed Computing, 11(2):135–145, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Silverman:1990:PPA

[Sil90] Robert D. Silverman. Parallel polynomial arithmetic over finite rings. Journal of Parallel and Distributed Computing, 10(3): 265–270, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sinclair:1987:ECO

[Sin87] J. B. Sinclair. Efficient computation of optimal assignments for distributed tasks. Journal of Parallel and Distributed Computing, 4(4):342–362, August 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Singhal:1993:TDM

[Sin93] Mukesh Singhal. A taxonomy of distributed mutual exclusion. Journal of Parallel and Distributed Computing, 18(1):94– 101, May 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1048/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1048/
production/pdf.

Singh:1995:FPN

[Sin95] Ambuj K. Singh. A framework for programming with nonatomic memories. Journal of Parallel and Distributed Computing, 26(2):211–224, April 15, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1060/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1060/production/pdf.

Sridhar:1992:RSP

[SIR92] Radhakrishnan Sridhar, Sitharama S. Iyengar, and Subbiah Rajanarayanan. Range search in parallel using distributed data structures. Journal of Parallel and Distributed Computing, 15(1):70–74, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schwiebert:1995:OFA

[SJ95] Loren Schwiebert and D. N. Jayasimha. Optimal fully adaptive minimal wormhole routing for meshes. Journal of Parallel and Distributed Computing, 27(1):56–70, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1071/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1071/production/pdf.

Schwiebert:1996:NSC

[SJ96] Loren Schwiebert and D. N. Jayasimha. A necessary and sufficient condition for deadlock-free wormhole routing. Journal of Parallel and Distributed Computing, 32(1):103–117, January 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0008/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0008/
production/pdf.

Skillicorn:1989:GMN

[SK89a] D. B. Skillicorn and B. Kocay. A global measure of network connectivity. Journal of Parallel and Distributed Computing, 7 (1):165–177, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stark:1989:SMD

[SK89b] W. Richard Stark and Leon Kotin. The social metaphor for distributed processing. Journal of Parallel and Distributed Computing, 7(1):125–147, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sheu:1990:QOD

[SK90] Phillip C-Y. Sheu and R. L. Kashyap. Query optimization in distributed logic-oriented object bases. Journal of Parallel and Distributed Computing, 8(1):60–71, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shu:1991:CKR

[SK91] Wei Shu and L. V. Kale. Chare kernel — a runtime support system for parallel computations. Journal of Parallel and Distributed Computing, 11(3):198–211, March 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stasko:1993:MBA

[SK93] John T. Stasko and Eileen Kraemer. A methodology for building application-specific visualizations of parallel programs. Journal of Parallel and Distributed Computing, 18(2):258– 264, June 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1062/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1062/
production/pdf.

Singh:1994:EGL

[SK94] Suresh Singh and James F. Kurose. Electing “good” leaders. Journal of Parallel and Distributed Computing, 21(2):184– 201, May 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1051/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1051/
production/pdf.

Sleem:2005:HMW

[SK05a] AlaaEldin Sleem and Anup Kumar. Handoff management in wireless data networks using topography-aware mobility prediction. Journal of Parallel and Distributed Computing, 65(8): 963–982, August 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Strumpen:2005:CMR

[SK05b] Volker Strumpen and Arvind Krishnamurthy. A collision model for randomized routing in fat-tree networks. Journal of Parallel and Distributed Computing, 65(9):1007–1021, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saeed:2009:DDS

[SK09] Fahad Saeed and Ashfaq Khokhar. A domain decomposition strategy for alignment of multiple biological sequences on multiprocessor platforms. Journal of Parallel and Distributed Computing, 69(7):666–677, July 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Subramaniam:1996:CLU

[SKH96] Krishnan R. Subramaniam, Suraj C. Kothari, and Don Heller. A communication library using active messages to improve performance of PVM. Journal of Parallel and Distributed Computing, 39(2):146–152, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0162/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0162/production/pdf.

Skillicorn:1996:PIT

[Ski96] D. B. Skillicorn. Parallel implementation of tree skeletons. Journal of Parallel and Distributed Computing, 39(2):115– 125, December 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1996.0160/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.0160/production/pdf.

Sitaram:1991:RDA

[SKK91] Dipak Sitaram, Israel Koren, and C. M. Krishna. A random, distributed algorithm to embed trees in partially faulty processor arrays. Journal of Parallel and Distributed Computing, 12(1):1–11, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schloegel:1997:MDS

[SKK97] Kirk Schloegel, George Karypis, and Vipin Kumar. Multilevel diffusion schemes for repartitioning of adaptive meshes. Journal of Parallel and Distributed Computing, 47(2):109–124, December 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1410/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1410/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1410/production/ref.

Sisodia:2004:NSS

[SKMM04] R. S. Sisodia, I. Karthigeyan, B. S. Manoj, and C. Siva Ram Murthy. A novel scheme for supporting integrated unicast and multicast traffic in ad hoc wireless networks. Journal of Parallel and Distributed Computing, 64(10):1185–1210, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sarukkai:1993:MVP

[SKR93] Sekhar R. Sarukkai, Doug Kimelman, and Larry Rudolph. A methodology for visualizing performance of loosely synchronous programs. Journal of Parallel and Distributed Computing, 18(2):242–251, June 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1060/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1060/production/pdf.

Smitley:1989:SMT

[SL89] David Smitley and Insup Lee. Synthesizing minimum total expansion topologies for reconfigurable interconnection networks. Journal of Parallel and Distributed Computing, 7(1):178–199, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sips:1990:NML

[SL90] Henk J. Sips and Hai Xiang Lin. A new model for on-line arithmetic with an application to the reciprocal calculation. Journal of Parallel and Distributed Computing, 8(3):218–230, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Seznec:1995:OMS

[SL95] Andre Seznec and Jacques Lenfant. Odd memory systems: a new approach. Journal of Parallel and Distributed Computing, 26(2):248–256, April 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1063/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1063/production/pdf.

Song:1997:POP

[SL97] Seok Ho Song and El-Hang Lee. Parallel optical packet address detection using planar optical interconnections. Journal of Parallel and Distributed Computing, 41(1):115–119, February 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1289/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1289/production/
pdf.

Srivatsa:2006:SDR

[SL06] Mudhakar Srivatsa and Ling Liu. Securing decentralized reputation management using TrustGuard. Journal of Parallel and Distributed Computing, 66(9):1217–1232, September 2006. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shin:2006:ADT

[SLG06] Chulho Shin, Seong-Won Lee, and Jean-Luc Gaudiot. Adaptive dynamic thread scheduling for simultaneous multithreaded architectures with a detector thread. Journal of Parallel and Distributed Computing, 66(10):1304–1321, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:1998:PAD

[SLP^{+}98] Hong Shen, K. Li, Y. Pan, G. H. Young, and S. Q. Zheng. Performance analysis for dynamic tree embedding in k-partite networks by a random walk. Journal of Parallel and Distributed Computing, 50(1–2):144–156, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1440/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1440/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1440/production/ref.

Song:2005:DLD

[SLWW05] Wen-Zhan Song, Xiang-Yang Li, Yu Wang, and Weizhao Wang. dBBlue: low diameter and self-routing Bluetooth scatternet. Journal of Parallel and Distributed Computing, 65(2):178–190, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stolfo:1986:DPS

[SM86] Salvatore J. Stolfo and Daniel P. Miranker. The DADO production system machine. Journal of Parallel and Distributed Computing, 3(2):269–296, June 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherson:1989:AAO

[SM89a] Isaac D. Scherson and Yiming Ma. Analysis and applications of the orthogonal access multiprocessor. Journal of Parallel and Distributed Computing, 7(2):232–255, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schieber:1989:PAM

[SM89b] B. Schieber and S. Moran. Parallel algorithms for maximum bipartite matchings and maximum 0–1 flows. Journal of Parallel and Distributed Computing, 6(1):20–38, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sanders:1992:DEU

[SM92a] William H. Sanders and Luai M. Malhis. Dependability evaluation using composed SAN-Based reward models. Journal of Parallel and Distributed Computing, 15(3):238–254, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Simone:1992:AEE

[SM92b] C. Simone and M. Ajmone Marsan. The application of EB-Equivalence rules to the structural reduction of GSPN models. Journal of Parallel and Distributed Computing, 15(3):296–302, July 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Satyanarayanan:1994:MIR

[SM94] R. Satyanarayanan and C. R. Muthukrishnan. Multiple instance resource allocation in distributed computing systems. Journal of Parallel and Distributed Computing, 23(1):94–100, October 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1122/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1122/
production/pdf.

Sinha:2000:PSA

[SM00] Bhabani P. Sinha and Amar Mukherjee. Parallel sorting algorithm using multiway merge and its implementation on a multi-mesh network. Journal of Parallel and Distributed Computing, 60(7):891–907, July 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1636;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1636/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1636/ref.

Stone:2004:LES

[SM04] Christopher Stone and Suresh Menon. Large-eddy simulations on distributed shared memory clusters. Journal of Parallel and Distributed Computing, 64(10):1103–1112, October 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Skillicorn:2008:DPV

[SM08a] D. B. Skillicorn and S. M. McConnell. Distributed prediction from vertically partitioned data. Journal of Parallel and Distributed Computing, 68(1):16–36, January 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Soliman:2008:HEI

[SM08b] Mostafa I. Soliman and Samir A. Mohamed. A highly efficient implementation of a backpropagation learning algorithm using matrix ISA. Journal of Parallel and Distributed Computing, 68 (7):949–961, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Strout:1991:ECS

[SMH91] Robert E. Strout, II, James R. McGraw, and Alan C. Hindmarsh. An examination of the conversion of software to multiprocessors. Journal of Parallel and Distributed Computing, 13 (1):1–16, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saavedra:1994:POD

[SMH94] R. H. Saavedra, Weihua H. Mao, and Kai Hwang. Performance and optimization of data prefetching strategies in scalable multiprocessors. Journal of Parallel and Distributed Computing, 22(3):427–448, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1102/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1102/production/pdf.

Steck:1993:PIR

[SMKL93] James Edward Steck, Bruce McMillin, K. Krishnamurthy, and Gary G. Leininger. Parallel implementation of a recursive least-squares neural network training method on the Intel iPSC/2. Journal of Parallel and Distributed Computing, 18(1): 89–93, May 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1047/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1047/
production/pdf.

Sang:1996:MPB

[SMR96] Janche Sang, Edward Mascarenhas, and Vernon Rego. Mobileprocess-based parallel simulation. Journal of Parallel and Distributed Computing, 33(1):12–23, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0020/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0020/production/pdf.

Sun:1993:SPM

[SN93] Xian-He H. Sun and Lionel M. Ni. Scalable problems and memory-bounded speedup. Journal of Parallel and Distributed Computing, 19(1):27–37, September 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1087/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1087/production/pdf.

Snir:2003:BPI

[Sni03] Marc Snir. Best papers from the 2002 International Parallel and Distributed Processing Symposium. Journal of Parallel and Distributed Computing, 63(9):846, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stichnoth:1994:GCA

[SOG94] James M. Stichnoth, David O’Hallaron, and Thomas R. Gross. Generating communication for array statements: Design, implementation, and evaluation. Journal of Parallel and Distributed Computing, 21(1):150–159, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1048/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1048/production/pdf.

Sohn:1996:PST

[Soh96] Andrew Sohn. Parallel satisfiability test with synchronous simulated annealing on distributed-memory multiprocessor. Journal of Parallel and Distributed Computing, 36(2):195–204, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0100/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0100/

production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0101/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0101/production/
pdf.

Stotts:1990:CGC

[SP90] P. David Stotts and T. W. Pratt. Coverability graphs for a class of synchronously executed unbounded Petri net. Journal of Parallel and Distributed Computing, 10(3):253–260, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sharma:1996:SAM

[SP96] Debendra Das Sharma and Dhiraj K. Pradhan. Submesh allocation in mesh multicomputers using busy-list: a best-fit approach with complete recognition capability. Journal of Parallel and Distributed Computing, 36(2):106–118, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0093/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0093/production/
pdf.

Schmidt:2008:SES

[SP08] Cristina Schmidt and Manish Parashar. Squid: Enabling search in DHT-based systems. Journal of Parallel and Distributed Computing, 68(7):962–975, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Saltz:1991:PEI

[SPBR91] Joel Saltz, Serge Petiton, Harry Berryman, and Adam Rifkin. Performance effects of irregular communication patterns on massively parallel multiprocessors. Journal of Parallel and Distributed Computing, 13(2):202–212, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schmidt:2003:MPP

[SPVvH03] H. A. Schmidt, E. Petzold, M. Vingron, and A. von Haeseler. Molecular phylogenetics: parallelized parameter estimation and quartet puzzling. Journal of Parallel and Distributed Computing, 63(7–8):719–727, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherson:1988:MOA

[SR88a] Isaac D. Scherson and Smil Ruhman. Multi-operand arithmetic in a partitioned associative architecture. Journal of Parallel and Distributed Computing, 5(6):655–668, December 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sridhar:1988:UMF

[SR88b] M. A. Sridhar and C. S. Raghavendra. Uniform minimal fullaccess networks. Journal of Parallel and Distributed Computing, 5(4):383–403, August 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sridhar:1990:MFA

[SR90] M. A. Sridhar and C. S. Raghavendra. Minimal full-access networks: enumeration and characterization. Journal of Parallel and Distributed Computing, 9(4):347–356, August 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Storer:1991:PAH

[SR91] James A. Storer and John H. Reif. A parallel architecture for high-speed data compression. Journal of Parallel and Distributed Computing, 13(2):222–227, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:1994:FAI

[SR94] Hongchi C. Shi and G. X. Ritter. A fast algorithm for image component labeling with local operators on mesh connected computers. Journal of Parallel and Distributed Computing, 23(3):455–461, December 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1156/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1156/production/pdf.

Sridhar:1995:CLS

[SR95] M. A. Sridhar and C. S. Raghavendra. Computing large subcubes in residual hypercubes. Journal of Parallel and Distributed Computing, 24(2):213–217, February 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1020/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1020/production/pdf.

Shankar:1997:RDAa

[SR97a] Ravi V. Shankar and Sanjay Ranka. Random data accesses on a coarse-grained parallel machine. I. one-to-one mappings. Journal of Parallel and Distributed Computing, 44(1):14–23, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1316/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1997.1316/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1316/production/ref.

Shankar:1997:RDAb

[SR97b] Ravi V. Shankar and Sanjay Ranka. Random data accesses on a coarse-grained parallel machine. II. one-to-many and many-to-one mappings. Journal of Parallel and Distributed Computing, 44(1):24–34, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1315/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1315/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1315/production/ref.

Shen:1990:VVE

[SRGB90] V. Y. Shen, C. Richter, M. L. Graf, and J. A. Brumfield. VERDI: a visual environment for designing distributed systems. Journal of Parallel and Distributed Computing, 9(2):128–137, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Siu:1995:TMP

[SRK95] Kai-Yeung Y. Siu, V. Roychowdhury, and T. Kailath. Toward massively parallel design of multipliers. Journal of Parallel and Distributed Computing, 24(1):86–93, January 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1008/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1008/production/pdf.

Saad:1989:DCH

[SS89] Y. Saad and M. H. Schultz. Data communication in hypercubes. Journal of Parallel and Distributed Computing, 6(1):115–135, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:1992:PSR

[SS92] Hanmao Shi and Jonathan Schaeffer. Parallel sorting by regular sampling. Journal of Parallel and Distributed Computing, 14 (4):361–372, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Som:1993:NPP

[SS93] Tapas K. Som and Robert G. Sargent. A new process to processor assignment criterion for reducing rollbacks in optimistic simulation. Journal of Parallel and Distributed Computing, 18(4):509–515, August 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1081/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1081/production/pdf.

Sinharoy:1994:DTA

[SS94a] Balaram Sinharoy and Boleslaw K. Szymanski. Data and task alignment in distributed memory architectures. Journal of Parallel and Distributed Computing, 21(1):61–74, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1042/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1042/production/pdf.

Squier:1994:CTA

[SS94b] R. K. Squier and K. Steiglitz. A comparison of two applicationspecific architectures for 2-d mesh computations. Journal of Parallel and Distributed Computing, 23(3):376–381, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1147/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1147/production/
pdf.

Sur:1995:ARF

[SS95] Sumit Sur and Pradip K. Srimani. An almost regular fault tolerant network with arbitrary number of nodes. Journal of Parallel and Distributed Computing, 26(1):54–60, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1047/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1047/production/pdf.

Sanchis:1996:PAC

[SS96] Laura A. Sanchis and Matthew B. Squire. Parallel algorithms for counting and randomly generating integer partitions. Journal of Parallel and Distributed Computing, 34(1):29–35, April 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1996.0043/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1996.0043/
production/pdf.

Szuba:1997:PEC

[SS97] Tadeusz Szuba and Robert Stras. Parallel evolutionary computing with the random PROLOG processor. Journal of Parallel and Distributed Computing, 47(1):78–85, November 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1396/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1396/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1396/production/ref.

Scheiman:1999:EBC

[SS99] Chris J. Scheiman and Klaus E. Schauser. Evaluating the benefits of communication coprocessors. Journal of Parallel and Distributed Computing, 57(2):236–256, May 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1528/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1528/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1528/production/ref.

Subhlok:2000:APM

[SS00] Jaspal Subhlok and Peter Steenkiste. Airshed pollution modeling in an HPF style environment. Journal of Parallel and Distributed Computing, 60(6):690–715, June 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1598;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.1999.1598/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1598/ref.

Sajith:2003:FPE

[SS03] G. Sajith and S. Saxena. Fast parallel edge colouring of graphs. Journal of Parallel and Distributed Computing, 63(9):775–785, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:2005:LEH

[SS05] Wei Shi and Pradip K. Srimani. Leader election in hierarchical star network. Journal of Parallel and Distributed Computing, 65(11):1435–1442, November 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scheutz:2006:AAD

[SS06] Matthias Scheutz and Paul Schermerhorn. Adaptive algorithms for the dynamic distribution and parallel execution of agentbased models. Journal of Parallel and Distributed Computing, 66(8):1037–1051, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sha:2008:CDD

[SS08] Kewei Sha and Weisong Shi. Consistency-driven data quality management of networked sensor systems. Journal of Parallel and Distributed Computing, 68(9):1207–1221, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sen:1991:GSI

[SSB91] Arunabha Sen, Abhutt Sengupta, and Subir Bandyopadhyay. Generalized supercube: an incrementally expandable interconnection network. Journal of Parallel and Distributed Computing, 13(3):338–344, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Simon:1998:HDS

[SSB98] Horst D. Simon, Andrew Sohn, and Rupak Biswas. HARP: a dynamic spectral partitioner. Journal of Parallel and Distributed Computing, 50(1–2):83–103, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1445/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1445/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1445/production/ref.

Saghi:1993:PPS

[SSG93] Gene Saghi, Howard Jay Siegel, and Jeffrey L. Gray. Predicting performance and selecting modes of parallelism: a case study using cyclic reduction on three parallel machines. Journal of Parallel and Distributed Computing, 19(3):219–233, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1106/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1106/production/
pdf.

Shih:2000:EIG

[SSHC00] Kuei-Ping Shih, Jang-Ping Sheu, Chua-Huang Huang, and Chih-Yung Chang. Efficient index generation for compiling twolevel mappings in data-parallel programs. Journal of Parallel and Distributed Computing, 60(2):189–216, February 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1601/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1601/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1601/production/ref.

Stoica:1996:HMC

[SSK96] Ion Stoica, Florin Sultan, and David Keyes. A hyperbolic model for communication in layered parallel processing environments. Journal of Parallel and Distributed Computing, 39(1):29–45, November 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0152/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0152/production/pdf.

Sim:2004:MLR

[SSL04] Leo Chin Sim, Heiko Schroder, and Graham Leedham. Major line removal morphological Hough transform on a hybrid system. Journal of Parallel and Distributed Computing, 64(9):

1060–1068, September 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scherson:1989:TNO

[SSM89] I. D. Scherson, S. Sen, and Y. Ma. Two nearly optimal sorting algorithms for mesh-connected processor arrays using Shearsort. Journal of Parallel and Distributed Computing, 6(1): 151–165, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shivle:2006:SAR

[SSM^{+}06] Sameer Shivle, H. J. Siegel, Anthony A. Maciejewski, Prasanna Sugavanam, Tarun Banka, Ralph Castain, Kiran Chindam, Steve Dussinger, Prakash Pichumani, Praveen Satyasekaran, William Saylor, David Sendek, J. Sousa, Jayashree Sridharan, and José Velazco. Static allocation of resources to communicating subtasks in a heterogeneous ad hoc Grid environment. Journal of Parallel and Distributed Computing, 66(4):600–611, April 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sugavanam:2007:RSA

[SSM^{+}07] Prasanna Sugavanam, H. J. Siegel, Anthony A. Maciejewski, Mohana Oltikar, Ashish Mehta, Ron Pichel, Aaron Horiuchi, Vladimir Shestak, Mohammad Al-Otaibi, Yogish Krishnamurthy, Syed Ali, Junxing Zhang, Mahir Aydin, Panho Lee, Kumara Guru, Michael Raskey, and Alan Pippin. Robust static allocation of resources for independent tasks under makespan and dollar cost constraints. Journal of Parallel and Distributed Computing, 67(4):400–416, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shestak:2008:SRM

[SSMS08] Vladimir Shestak, Jay Smith, Anthony A. Maciejewski, and Howard Jay Siegel. Stochastic robustness metric and its use for static resource allocations. Journal of Parallel and Distributed Computing, 68(8):1157–1173, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shan:2002:CTP

[SSOB02] Hongzhang Shan, Jaswinder Pal Singh, Leonid Oliker, and Rupak Biswas. A comparison of three programming models for adaptive applications on the Origin2000. Journal of Parallel and Distributed Computing, 62(2):241–266, February 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1777;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1777/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1777/ref.

Sivasubramaniam:1994:SBS

[SSRV94] A. Sivasubramaniam, A. Singla, U. Ramachandran, and H. Venkateswaran. A simulation-based scalability study of parallel systems. Journal of Parallel and Distributed Computing, 22(3):411–426, September 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1101/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1101/production/pdf.

Santoro:1988:ECD

[SSS88] Nicola Santoro, Michael Scheutzow, and Jeffrey B. Sidney. On the expected complexity of distributed selection. Journal of Parallel and Distributed Computing, 5(2):194–203, April 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shacham:2007:SMC

[SSS07] Ohad Shacham, Mooly Sagiv, and Assaf Schuster. Scaling model checking of dataraces using dynamic information. Journal of Parallel and Distributed Computing, 67(5):536–550, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schmidt:2009:SPA

[SSTP09] Matthew C. Schmidt, Nagiza F. Samatova, Kevin Thomas, and Byung-Hoon Park. A scalable, parallel algorithm for maximal clique enumeration. Journal of Parallel and Distributed Computing, 69(4):417–428, April 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sohn:1997:DWD

[SSYG97] Andrew Sohn, Mitsuhisa Sato, Namhoon Yoo, and Jean-Luc Gaudiot. Data and workload distribution in a multithreaded architecture. Journal of Parallel and Distributed Computing, 40(2):256–264, February 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1262/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.1262/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1996.1262/production/ref.

Smith:1985:FIP

[ST85] S. P. Smith and H. C. Torng. A fast inner product processor based on equal alignments. Journal of Parallel and Distributed Computing, 2(4):376–390, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schwiegelshohn:1987:SAC

[ST87] Uwe Schwiegelshohn and Lothar Thiele. A systolic array for cyclic-by-rows Jacobi algorithms. Journal of Parallel and Distributed Computing, 4(3):334–340, June 1987. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schwiegelshohn:1989:LSA

[ST89] Uwe Schwiegelshohn and Lothar Thiele. Linear systolic arrays for matrix computations. Journal of Parallel and Distributed Computing, 7(1):28–39, August 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Salinger:2002:BAO

[ST02] Petr Salinger and Pavel Tvrdík. Broadcasting in all-output-port meshes of trees with distance-insensitive switching. Journal of Parallel and Distributed Computing, 62(8):1272–1294, August 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1848;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1848/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1848/ref.

Sundell:2005:FLF

[ST05] Håkan Sundell and Philippas Tsigas. Fast and lock-free concurrent priority queues for multi-thread systems. Journal of Parallel and Distributed Computing, 65(5):609–627, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Salinger:2006:BAO

[ST06] Petr Salinger and Pavel Tvrdík. Broadcasting in all-output-port cube-connected cycles with distance-insensitive switching. Journal of Parallel and Distributed Computing, 66(8):1103– 1110, August 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sun:2008:ARD

[ST08a] Yang Sun and Teng Tiow Tay. Analysis and reduction of data spikes in thin client computing. Journal of Parallel and Distributed Computing, 68(11):1463–1472, November 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sundell:2008:LFD

[ST08b] Håkan Sundell and Philippas Tsigas. Lock-free deques and doubly linked lists. Journal of Parallel and Distributed Computing, 68(7):1008–1020, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stadtherr:2004:SIO

[Sta04] Hans Stadtherr. Scheduling interval orders with communication delays in parallel. Journal of Parallel and Distributed Computing, 64(1):1–15, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stewart:1995:RAD

[Ste95] A. Stewart. Reasoning about data-parallel array assignment. Journal of Parallel and Distributed Computing, 27(1):79–85, May 1995. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1073/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1995.1073/
production/pdf.

Shimada:1992:RTP

[STN92] Toshio Shimada, Kenji Toda, and Kenji Nishida. Real-time parallel architecture for sensor fusion. Journal of Parallel and Distributed Computing, 15(2):143–152, June 1992. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stout:1987:SDC

[Sto87] Quentin F. Stout. Supporting divide-and-conquer algorithms for image processing. Journal of Parallel and Distributed Computing, 4(1):95–115, February 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stout:1990:SIA

[Sto90] Quentin F. Stout. Special issue on algorithms for hypercube computers: Guest Editor’s introduction. Journal of Parallel and Distributed Computing, 8(4):301–302, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shamir:1987:PAL

[SU87] Eli Shamir and Eli Upfal. A probabilistic approach to the loadsharing problem in distributed systems. Journal of Parallel and Distributed Computing, 4(5):521–530, October 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sun:2002:SVE

[Sun02] Xian-He Sun. Scalability versus execution time in scalable systems. Journal of Parallel and Distributed Computing, 62(2):173–192, February 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1773;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1773/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1773/ref.

Subhlok:2000:OUM

[SV00] Jaspal Subhlok and Gary Vondran. Optimal use of mixed task and data parallelism for pipelined computations. Journal of Parallel and Distributed Computing, 60(3):297–319, March 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1999.1596;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1596/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1596/ref.

Sanjay:2008:PMP

[SV08] H. A. Sanjay and Sathish Vadhiyar. Performance modeling of parallel applications for Grid scheduling. Journal of Parallel and Distributed Computing, 68(8):1135–1145, August 2008.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Stout:1990:IHC

[SW90] Quentin F. Stout and Bruce Wagar. Intensive hypercube communication prearranged communication in link-bound machines. Journal of Parallel and Distributed Computing, 10(2): 167–181, October 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Savage:1991:PGP

[SW91] John E. Savage and Markus G. Wloka. Parallelism in graphpartitioning. Journal of Parallel and Distributed Computing, 13 (3):257–272, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schikarski:1996:EPM

[SW96] Andreas Schikarski and Dorothea Wagner. Efficient parallel matrix inversion on interconnection networks. Journal of Parallel and Distributed Computing, 34(2):196–201, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0055/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0055/production/pdf.

Swarztrauber:1998:TAM

[Swa98] Paul N. Swarztrauber. Transposing arrays on multicomputers using de Bruijn sequences. Journal of Parallel and Distributed Computing, 53(1):63–77, August 25, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1476/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1476/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1476/production/ref.

Stolfo:1991:PPR

[SWC^{+}91] Salvatore J. Stolfo, Ouri Wolfson, Philip K. Chan, Hasanat M. Dewan, Leland Woodbury, Jason S. Glazier, and David A. Ohsie. PARULEL: Parallel rule processing using meta-rules for redaction. Journal of Parallel and Distributed Computing, 13(4):366–382, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shirazi:1990:AEH

[SWP90] Behrooz Shirazi, Mingfang Wang, and Girishi Pathak. Analysis and evaluation of heuristic methods for static task scheduling. Journal of Parallel and Distributed Computing, 10(3):222–232, November 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2008:HBP

[SX08] Haiying Shen and Cheng-Zhong Xu. Hash-based proximity clustering for efficient load balancing in heterogeneous DHT networks. Journal of Parallel and Distributed Computing, 68 (5):686–702, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shi:2006:SIS

[SXZ06] Weisong Shi, Cheng-Zhong Xu, and Xiaobo Zhou. Special issue: Security in Grid and distributed systems. Journal of Parallel and Distributed Computing, 66(9):1113–1115, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sheu:2001:MAH

[SY01] Shih-Hsien Sheu and Chang-Biau Yang. Multicast algorithms for hypercube multiprocessors. Journal of Parallel and Distributed Computing, 61(1):137–149, January 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1668;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1668/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1668/ref.

Shieh:2004:PNU

[SY04] Shiuh-Pyng Shieh and Wen-Her Yang. Protecting network users in mobile code systems. Journal of Parallel and Distributed Computing, 64(2):191–200, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Swami:1992:AHS

[SYG92] Arun Swami, Honesty C. Young, and Ashish Gupta. Algorithms for handling skew in parallel task scheduling. Journal of Parallel and Distributed Computing, 16(4):363–377, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Scheuermann:1994:CBI

[SYO94] P. Scheuermann, A. Yaagoub, and M. Aris Ouksel. Compression of binary images on a hypercube machine. Journal of Parallel and Distributed Computing, 23(1):49–59, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1118/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1118/production/
pdf.

Sendag:2007:IWP

[SYYU07] Resit Sendag, Ayse Yilmazer, Joshua J. Yi, and Augustus K. Uht. The impact of wrong-path memory references in cachecoherent multiprocessor systems. Journal of Parallel and Distributed Computing, 67(12):1256–1269, December 2007. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shavit:2000:CFD

[SZ00a] Nir Shavit and Asaph Zemach. Combining funnels: a dynamic approach to software combining. Journal of Parallel and Distributed Computing, 60(11):1355–1387, November 1, 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1621;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1621/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1621/ref.

Shen:2000:PES

[SZ00b] Xiaojun Shen and Yixin Zhang. Partitionability of k-extra-stage omega networks and an optimal task migration algorithm. Journal of Parallel and Distributed Computing, 60(3):334–348, March 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1999.1603;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1603/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1603/ref.

Subrata:2003:SRC

[SZ03] Riky Subrata and Albert Y. Zomaya. Some results on the computation of Voronoi diagrams on a mesh with multiple broadcasting. Journal of Parallel and Distributed Computing, 63(12): 1300–1314, December 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2009:PLO

[SZ09] Haiying Shen and Yingwu Zhu. A proactive low-overhead file replication scheme for structured P2P content delivery networks. Journal of Parallel and Distributed Computing, 69 (5):429–440, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sengupta:1992:SRA

[SZB92] A. Sengupta, K. Zemoudeh, and S. Bandyopadhyay. Selfrouting algorithms for strongly regular multistage interconnection networks. Journal of Parallel and Distributed Computing, 14(2):187–192, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2007:PLP

[SZD07] Xipeng Shen, Yutao Zhong, and Chen Ding. Predicting locality phases for dynamic memory optimization. Journal of Parallel and Distributed Computing, 67(7):783–796, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shen:2005:DBI

[SZW05] Chi Shen, Jun Zhang, and Kai Wang. Distributed block independent set algorithms and parallel multilevel ILU preconditioners. Journal of Parallel and Distributed Computing, 65 (3):331–346, March 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Szymanski:1995:HOI

[Szy95] Ted Szymanski. “Hypermeshes”: Optical interconnection networks for parallel computing. Journal of Parallel and Distributed Computing, 26(1):1–23, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1043/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1043/production/pdf.

Tanimoto:1984:HCL

[Tan84] Steven L. Tanimoto. Hierarchical cellular logic for pyramid computers. Journal of Parallel and Distributed Computing, 1 (2):105–132, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Theys:2001:WTT

[TAS^{+}01] Mitchell D. Theys, Shoukat Ali, Howard Jay Siegel, Mani Chandy, Kai Hwang, Ken Kennedy, Lui Sha, Kang G. Shin, Marc Snir, Larry Snyder, and Thomas Sterling. What are the top ten most influential parallel and distributed processing concepts of the past millenium? Journal of Parallel and Distributed Computing, 61(12):1827–1841, December 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1767;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1767/pdf.

Taylor:1987:RAI

[Tay87] Fred J. Taylor. A residue arithmetic implementation of the FFT. Journal of Parallel and Distributed Computing, 4(2):191– 208, April 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tay:2002:CPC

[Tay02] Y. C. Tay. A comparison of pixel complexity in composition techniques for sort-last rendering. Journal of Parallel and Distributed Computing, 62(1):152–171, January 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1771;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1771/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1771/ref.

Taylor:2005:SPO

[Tay05] David Taylor. Scrolling partially ordered event displays. Journal of Parallel and Distributed Computing, 65(5):643–653, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Triantafillou:1990:DNM

[TB90] P. Triantafillou and M. Bauer. Distributed name management in Internet systems: a study of design and performance issues. Journal of Parallel and Distributed Computing, 9(4):357–368, August 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tai:2007:QDR

[TBHA07] Sylvia Tai, Robert R. Benkoczi, Hossam Hassanein, and Selim G. Akl. QoS and data relaying for wireless sensor networks. Journal of Parallel and Distributed Computing, 67(6):715–726, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Trahan:2000:OSP

[TBPV00] Jerry L. Trahan, Anu G. Bourgeois, Yi Pan, and Ramachandran Vaidyanathan. Optimally scaling permutation routing on reconfigurable linear arrays with optical buses. Journal of Parallel and Distributed Computing, 60(9):1125–1136, September 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1643;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1643/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1643/ref.

Tari:2005:LFT

[TBZB05] Zahir Tari, James Broberg, Albert Y. Zomaya, and Roberto Baldoni. A least flow-time first load sharing approach for distributed server farm. Journal of Parallel and Distributed Computing, 65(7):832–842, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tzeng:1992:EAE

[TC92] Nian-Feng Tzeng and Hsing-Lung Chen. An effective approach to the enhancement of incomplete hypercube computers. Journal of Parallel and Distributed Computing, 14(2):163–174, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tsay:1996:DLD

[TC96] Jong-Chuang Tsay and Pen-Yuang Chang. Designing lowerdimensional regular arrays for algorithms with uniform dependencies. Journal of Parallel and Distributed Computing, 33(1):24–32, February 25, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0021/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0021/production/pdf.

Tang:2003:CEW

[TC03] Xueyan Tang and Samuel T. Chanson. On caching effectiveness of Web clusters under persistent connections. Journal of Parallel and Distributed Computing, 63(10):981–995, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tang:2004:AHR

[TC04] Xueyan Tang and Samuel T. Chanson. Adaptive hash routing for a cluster of client-side Web proxies. Journal of Parallel and Distributed Computing, 64(10):1168–1184, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tsouloupas:2007:GTI

[TD07] George Tsouloupas and Marios D. Dikaiakos. GridBench: a tool for the interactive performance exploration of Grid infrastructures. Journal of Parallel and Distributed Computing, 67(9): 1029–1045, September 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Thaeler:2005:IIL

[TDC05] Andrew Thaeler, Min Ding, and Xiuzhen Cheng. iTPS: an improved location discovery scheme for sensor networks with long-range beacons. Journal of Parallel and Distributed Computing, 65(2):98–106, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Touzene:2005:EDS

[TDM05] Abderezak Touzene, Khaled Day, and Burkhard Monien. Edgedisjoint spanning trees for the generalized butterfly networks and their applications. Journal of Parallel and Distributed Computing, 65(11):1384–1396, November 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Teng:1990:APA

[Ten90] Shang-Hua Teng. Adaptive parallel algorithms for integral knapsack problems. Journal of Parallel and Distributed Computing, 8(4):400–406, April 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tamir:1992:HSH

[TF92] Yuval Tamir and Gregory L. Frazier. Hardware support for high-priority traffic in VLSI communication switches. Journal of Parallel and Distributed Computing, 14(4):402–416, April 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tufo:2001:FPD

[TF01] H. M. Tufo and P. F. Fischer. Fast parallel direct solvers for coarse grid problems. Journal of Parallel and Distributed Computing, 61(2):151–177, February 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1676;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1676/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1676/ref.

Tomlinson:1997:MFG

[TG97] Alexander I. Tomlinson and Vijay K. Garg. Monitoring functions on global states of distributed programs. Journal of Parallel and Distributed Computing, 41(2):173–189, March 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1298/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1298/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1298/production/ref.

Tang:1999:APT

[TG99] Xinan Tang and Guang R. Gao. Automatically partitioning threads for multithreaded architectures. Journal of Parallel and Distributed Computing, 58(2):159–189, August 1999. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1551/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1551/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1551/production/ref.

Termehchi:2003:POT

[TG03] Arash Termehchi and Mohammad Ghodsi. Pipelined operator tree scheduling in heterogeneous environments. Journal of Parallel and Distributed Computing, 63(6):630–637, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tarafdar:2004:PCS

[TG04] Ashis Tarafdar and Vijay K. Garg. Predicate control: synchronization in distributed computations with look-ahead. Journal of Parallel and Distributed Computing, 64(2):219–237, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Huang:1990:FPM

[tH90] Shing tsaan Huang. A fully pipelined minimum-spanning-tree constructor. Journal of Parallel and Distributed Computing, 9(1):55–62, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tutsch:2002:GSE

[TH02] Dietmar Tutsch and Günter Hommel. Generating systems of equations for performance evaluation of multistage interconnection networks. Journal of Parallel and Distributed Computing, 62(2):228–240, February 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1776;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1776/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1776/ref.

Tikir:2008:HMD

[TH08] Mustafa M. Tikir and Jeffrey K. Hollingsworth. Hardware monitors for dynamic page migration. Journal of Parallel and Distributed Computing, 68(9):1186–1200, September 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tokhi:1997:PEI

[THBF97] M. O. Tokhi, M. A. Hossain, M. J. Baxter, and P. J. Fleming. Performance evaluation issues in real-time parallel signal processing and control. Journal of Parallel and Distributed Computing, 42(1):67–74, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1307/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1307/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1307/production/ref.

Theodoropoulos:2002:DSA

[The02] Georgios K. Theodoropoulos. Distributed simulation of asynchronous hardware: The program driven synchronization protocol. Journal of Parallel and Distributed Computing, 62(4):622–655, April 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1806;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1806/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1806/ref.

Tewksbury:1993:TCO

[THN^{+}93] S. K. Tewksbury, L. A. Hornak, H. E. Nariman, S. M. Langsjoen, N. J. Hall, J. J. Hall, and S. P. McGinnis. Toward cointegration of optical interconnections within silicon microelectronic systems. Journal of Parallel and Distributed Computing, 17(3):188–199, March 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1017/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1017/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1993.1018/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1018/production/pdf.

Taylor:1987:NCS

[THSS87] Stephen Taylor, Lisa Hellerstein, Shmuel Safra, and Ehud Shapiro. Notes on the complexity of systolic programs. Journal of Parallel and Distributed Computing, 4(3):250–265, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tamir:1992:HCM

[TJ92] Yuval Tamir and G. Janakiraman. Hierarchical coherency management for shared virtual memory multicomputers. Journal of Parallel and Distributed Computing, 15(4):408–419, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Totok:2007:MCW

[TK07] Alexander Totok and Vijay Karamcheti. Modeling of concurrent Web sessions with bounded inconsistency in shared data. Journal of Parallel and Distributed Computing, 67(7):830–847, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Trifunovic:2008:PMA

[TK08] Aleksandar Trifunović and William J. Knottenbelt. Parallel multilevel algorithms for hypergraph partitioning. Journal of Parallel and Distributed Computing, 68(5):563–581, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Thulasiraman:2004:FGL

[TKHG04] Parimala Thulasiraman, Ashfaq A. Khokhar, Gerd Heber, and Guang R. Gao. A fine-grain load-adaptive algorithm of the 2D discrete wavelet transform for multithreaded architectures. Journal of Parallel and Distributed Computing, 64(1):68–78, January 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tseng:1996:ECR

[TL96] Yu-Chee Tseng and Ten-Hwang Lai. On the embedding of a class of regular graphs in a faulty hypercube. Journal of Parallel and Distributed Computing, 37(2):200–206, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0120/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0120/production/
pdf.

Tseng:1994:MRG

[TLW94] Yu-Chee C. Tseng, Ten-Hwang H. Lai, and Li-Fen F. Wu. Matrix representation of graph embedding in a hypercube. Journal of Parallel and Distributed Computing, 23(2):215–223, November 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1133/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1133/production/pdf.

Talbi:2006:HPA

[TM06] El-Ghazali Talbi and Hervé Meunier. Hierarchical parallel approach for GSM mobile network design. Journal of Parallel and Distributed Computing, 66(2):274–290, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Trombetti:2006:HPC

[TMM06] G. A. Trombetti, I. Merelli, and L. Milanesi. High performance cDNA sequence analysis using Grid technology. Journal of Parallel and Distributed Computing, 66(12):1482–1488, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

TorrasIGenis:1989:RNL

[Tor89] Carme Torras I. Genis. Relaxation and neural learning: Points of convergence and divergence. Journal of Parallel and Distributed Computing, 6(2):217–244, April 1989. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tagamets:1989:DFI

[TR89] Malle A. Tagamets and James A. Reggia. A data flow implementation of a competition-based connectionist model. Journal of Parallel and Distributed Computing, 6(3):704–714, June 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Thirumalai:1996:ECA

[TR96] Ashwath Thirumalai and J. Ramanujam. Efficient computation of address sequences in data parallel programs using closed forms for basis vectors. Journal of Parallel and Distributed Computing, 38(2):188–203, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0140/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0140/production/pdf.

Traff:2008:OBF

[TR08] Jesper Larsson Träff and Andreas Ripke. Optimal broadcast for fully connected processor-node networks. Journal of Parallel and Distributed Computing, 68(7):887–901, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Traff:2009:WPP

[Trä09] Jesper Larsson Träff. What the parallel-processing community has (failed) to offer the multi/many-core generation. Journal of Parallel and Distributed Computing, 69(9):807–812, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tissot:2006:ORO

[TRSS06] Y. Tissot, G. A. Russell, K. J. Symington, and J. F. Snowdon. Optimization of reconfigurable optically interconnected systems for parallel computing. Journal of Parallel and Distributed Computing, 66(2):238–247, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tong:1991:OFF

[TS91] Charles Tong and Paul N. Swarztrauber. Ordered fast Fourier transforms on a massively parallel hypercube multiprocessor. Journal of Parallel and Distributed Computing, 12(1):50–59, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Taylor:1997:PMD

[TSA97] Valerie E. Taylor, Rick L. Stevens, and Kathryn E. Arnold. Parallel molecular dynamics: Implications for massively parallel machines. Journal of Parallel and Distributed Computing, 45(2):166–175, September 15, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1370/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1370/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1370/production/ref.

Theys:2001:HSD

[TSC01] Mitchell D. Theys, Howard Jay Siegel, and Edwin K. P. Chong. Heuristics for scheduling data requests using collective communications in a distributed communication network. Journal of Parallel and Distributed Computing, 61(9):1337–1366, September 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1753;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1753/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1753/ref.

Tatarchuk:2008:AIM

[TSD08] Natalya Tatarchuk, Jeremy Shopf, and Christopher DeCoro. Advanced interactive medical visualization on the GPU. Journal of Parallel and Distributed Computing, 68(10):1319–1328, October 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tseng:1990:SAP

[Tse90] Ping-Sheng Tseng. A systolic array parallelizing compiler. Journal of Parallel and Distributed Computing, 9(2):116–127, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tseng:1995:DTW

[Tse95] Yu-Chee Tseng. Detecting termination by weight-throwing in a faulty distributed system. Journal of Parallel and Distributed Computing, 25(1):7–15, February 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1025/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1025/production/pdf.

Taylor:2001:BLV

[TSHH01] Valerie E. Taylor, Eric J. Schwabe, Bruce K. Holmer, and Michelle R. Hribar. Balancing load versus decreasing communication: Parameterizing the tradeoff. Journal of Parallel and Distributed Computing, 61(5):567–580, May 1, 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1701;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1701/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1701/ref.

Tsur:2007:ISR

[Tsu07] Dekel Tsur. Improved scheduling in rings. Journal of Parallel and Distributed Computing, 67(5):531–535, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Trdlicka:1998:EAC

[TT98] Jan Trdlička and Pavel Tvrdík. Embedding of k-ary complete trees into hypercubes with uniform load. Journal of Parallel and Distributed Computing, 52(2):120–131, August 1, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1464/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1464/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1464/production/ref.

Turner:2007:DFC

[TT07] Yoshio Turner and Yuval Tamir. Deadlock-free connectionbased adaptive routing with dynamic virtual circuits. Journal of Parallel and Distributed Computing, 67(1):13–32, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Torrellas:1995:EPC

[TTG95] J. Torrellas, A. Tucker, and A. Gupta. Evaluating the performance of cache-affinity scheduling in shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 24(2):139–151, February 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1014/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1014/production/pdf.

Takaoka:1992:EVA

[TU92] Tadao Takaoka and Kiyomi Umehara. An efficient VLSI algorithm for the all pairs shortest path problem. Journal of Parallel and Distributed Computing, 16(3):265–270, November 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tollenaere:1992:PIC

[TVO92] Tom Tollenaere, Marc M. Van Hulle, and Guy A. Orban. Parallel implementation and capabilities of entropy-driven artificial neural networks. Journal of Parallel and Distributed Computing, 14(3):286–305, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Trahan:1997:CTG

[TVS97] Jerry L. Trahan, Ramachandran Vaidyanathan, and Chittur P. Subbaraman. Constant time graph algorithms on the reconfigurable multiple bus machine. Journal of Parallel and Distributed Computing, 46(1):1–14, October 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1385/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1385/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1385/production/ref.

Trahan:1996:PSF

[TVT96] Jerry L. Trahan, Ramachandran Vaidyanathan, and Ratnapuri K. Thiruchelvan. On the power of segmenting and fusing buses. Journal of Parallel and Distributed Computing, 34(1):82–94, April 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0047/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0047/production/pdf.

Teng:1987:PAM

[TW87] Shang-Hua Teng and Bin Wang. Parallel algorithms for message decomposition. Journal of Parallel and Distributed Computing, 4(3):231–249, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tel:1989:HPM

[TW89] Gerard Tel and Harry A. G. Wijshoff. Hierarchical parallel memory systems and multiperiodic skewing schemes. Journal of Parallel and Distributed Computing, 7(2):355–367, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tang:1990:SCA

[TY90a] Peiyi Tang and Pen-Chung Yew. Software combining algorithms for distributing hot-spot addressing. Journal of Parallel and Distributed Computing, 10(2):130–139, October 1990.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tsay:1990:SF

[TY90b] Jong-Chuang Tsay and Sy Yuan. Systolic flow. Journal of Parallel and Distributed Computing, 8(3):286–291, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tzeng:1995:SID

[TY95] N. F. Tzeng and P. C. Yew. Special issue on distributed shared memory systems — Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 29(2):105–107, September 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1111/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1111/production/pdf.

Tian:2009:TCR

[TYH09] Guo-Zhong Tian, Jiong Yu, and Jing-Sha He. Towards critical region reliability support for Grid workflows. Journal of Parallel and Distributed Computing, 69(12):989–995, December 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Traff:2000:SPA

[TZ00] Jesper L. Träff and Christos D. Zaroliagis. A simple parallel algorithm for the single-source shortest path problem on planar digraphs. Journal of Parallel and Distributed Computing, 60(9):1103–1124, September 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1646;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1646/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1646/ref.

Talbi:2006:GBC

[TZ06] El-Ghazali Talbi and Albert Zomaya. Grids in bioinformatics and computational biology. Journal of Parallel and Distributed Computing, 66(12):1481, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Taheri:2007:CTD

[TZ07] Javid Taheri and Albert Y. Zomaya. Clustering techniques for dynamic location management in mobile computing. Journal of Parallel and Distributed Computing, 67(4):430–447, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tzeng:1991:AIT

[Tze91] Nian-Feng Tzeng. Alleviating the impact of tree saturation on multistage interconnection network performance. Journal of Parallel and Distributed Computing, 12(2):107–117, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Tzeng:1993:RCC

[Tze93] Nian-Feng F. Tzeng. A reliable cube-connected cycles structure. Journal of Parallel and Distributed Computing, 19(1): 64–71, September 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1993.1091/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1993.1091/production/pdf.

Tehranian:2006:RFR

[TZH^{+}06] Shahram Tehranian, Yongsheng Zhao, Tony Harvey, Anand Swaroop, and Keith Mckenzie. A robust framework for realtime distributed processing of satellite data. Journal of Parallel and Distributed Computing, 66(3):403–418, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ucar:2006:TAH

[UAKI06] Bora Ucar, Cevdet Aykanat, Kamer Kaya, and Murat Ikinci. Task assignment in heterogeneous computing systems. Journal of Parallel and Distributed Computing, 66(1):32–46, January 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ucar:2007:PIR

[UAPM07] Bora Uçar, Cevdet Aykanat, Mustafa Ç. Pınar, and Tahir Malas. Parallel image restoration using surrogate constraint methods. Journal of Parallel and Distributed Computing, 67 (2):186–204, February 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Uresin:1996:EAC

[ÜD96] Aydin Üresin and Michel Dubois. Effects of asynchronism on the convergence rate of iterative algorithms. Journal of Parallel and Distributed Computing, 34(1):66–81, April 10, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0046/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0046/production/pdf.

Ullman:1984:FSS

[Ull84] Jeffrey D. Ullman. Flux, sorting, and supercomputer organization for AI applications. Journal of Parallel and Distributed Computing, 1(2):133–151, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Umeo:1985:CSM

[Ume85] Hiroshi Umeo. A class of SIMD machines simulated by systolic arrays. Journal of Parallel and Distributed Computing, 2(4): 391–403, November 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Uyar:1994:FRN

[UR94] M. Uemit Uyar and A. P. Reeves. Fault reconfiguration for the near neighbor task in a multistage-network MIMD system. Journal of Parallel and Distributed Computing, 22(1): 1–15, July 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1065/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1065/
production/pdf.

Ujaldon:1996:PTS

[UZSS96] Manuel Ujaldon, Emilio L. Zapata, Shamik D. Sharma, and Joel Saltz. Parallelization techniques for sparse matrix applications. Journal of Parallel and Distributed Computing, 38(2):256–266, November 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0146/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0146/production/pdf.

Venkatraman:2003:SER

[VA03] Lakshmi Venkatraman and Dharma P. Agrawal. Strategies for enhancing routing security in protocols for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 63(2): 214–227, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Verma:2007:GSP

[VA07] Akshat Verma and Ashok Anand. General store placement for response time minimization in parallel disks. Journal of Parallel and Distributed Computing, 67(12):1286–1300, December 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

VandeGeijn:1994:GCO

[Van94] R. A. Van de Geijn. On global combine operations. Journal of Parallel and Distributed Computing, 22(2):324–328, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1091/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1091/production/
pdf.

vanderStok:1996:AOR

[van96] P. D. V. van der Stok. Atomic objects for real-time distributed systems with mixed time constraints. Journal of Parallel and Distributed Computing, 36(1):53–65, July 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0089/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0089/production/pdf.

Varadarajan:1991:ESN

[Var91] Ravi Varadarajan. Embedding shuffle networks in hypercubes. Journal of Parallel and Distributed Computing, 11(3):252–256, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Varvarigos:1994:PMB

[VB94] E. A. Varvarigos and D. P. Bertsekas. Partial multinode broadcast and partial exchange algorithms for d-dimensional meshes. Journal of Parallel and Distributed Computing, 23(2):177–189, November 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1130/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1130/
production/pdf.

Veanes:1996:NCF

[VB96] Margus Veanes and Jonas Barklund. Natural cycletrees: Flexible interconnection graphs. Journal of Parallel and Distributed Computing, 33(1):44–54, February 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0023/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0023/production/pdf.

Veeravalli:2002:ESS

[VB02] Bharadwaj Veeravalli and Gerassimos Barlas. Efficient scheduling strategies for processing multiple divisible loads on bus networks. Journal of Parallel and Distributed Computing, 62(1):132–151, January 1, 2002. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1778;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1778/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1778/ref.

Venugopal:2008:SBH

[VB08] Srikumar Venugopal and Rajkumar Buyya. An SCP-based heuristic approach for scheduling distributed data-intensive applications on global Grids. Journal of Parallel and Distributed Computing, 68(4):471–487, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vin:1990:EDD

[VBM90] H. M. Vin, F. Berman, and J. S. Mattson, Jr. Efficient datadriven evaluation: Theory and implementation. Journal of Parallel and Distributed Computing, 10(4):367–385, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vadhiyar:2004:GGB

[VD04] Sathish S. Vadhiyar and Jack J. Dongarra. GrADSolve — a grid-based RPC system for parallel computing with application-level scheduling. Journal of Parallel and Distributed Computing, 64(6):774–783, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Veldhorst:1989:GEP

[Vel89] M. Veldhorst. Gaussian elimination with partial pivoting on an MIMD computer. Journal of Parallel and Distributed Computing, 6(1):62–68, February 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vidal:2008:PCE

[VGAB08] Antonio M. Vidal, Victor M. Garcia, Pedro Alonso, and Miguel O. Bernabeu. Parallel computation of the eigenvalues of symmetric Toeplitz matrices through iterative methods. Journal of Parallel and Distributed Computing, 68(8):1113–1121, August 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vajapeyam:1993:TES

[VH93] Stiram Vajapeyam and Wei-Chung C. Hsu. Toward effective scalar hardware for highly vectorizable applications. Journal of Parallel and Distributed Computing, 19(3):147–162, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1100/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1100/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1101/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1101/production/pdf.

Vanhatupa:2008:PMI

[VHH08] Timo Vanhatupa, Marko Hännikäinen, and Timo D. Hämäläinen. Performance model for IEEE 802.11s wireless mesh network deployment design. Journal of Parallel and Distributed Computing, 68(3):291–305, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vishkin:1987:RPS

[Vis87] Uzi Vishkin. Randomized parallel speedups for list ranking. Journal of Parallel and Distributed Computing, 4(3):319–333, June 1987. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

VandenBout:1995:TIM

[VM95] David E. Van den Bout and Thomas K. Miller, III. TInMANN: the integer Markovian artificial neural network for performing competitive and Kohonen learning. Journal of Parallel and Distributed Computing, 25(2):107–114, March 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1033/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1033/production/pdf.

Vetter:2003:CCL

[VM03] Jeffrey S. Vetter and Frank Mueller. Communication characteristics of large-scale scientific applications for contemporary cluster architectures. Journal of Parallel and Distributed Computing, 63(9):853–865, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vitter:1993:LSS

[VN93] Jeffrey Scott Vitter and Mark H. Nodine. Large-scale sorting in uniform memory hierarchies. Journal of Parallel and Distributed Computing, 17(1–2):107–114, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1008/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1008/production/
pdf.

VanHulle:1989:EDA

[VO89] M. M. Van Hulle and G. A. Orban. Entropy driven artificial neuronal networks and sensorial representation: a proposal. Journal of Parallel and Distributed Computing, 6(2):264–290, April 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vazquez-Poletti:2006:CHI

[VPHML06] J. L. Vázquez-Poletti, E. Huedo, R. S. Montero, and I. M. Llorente. Coordinated harnessing of the IRISGrid and EGEE testbeds with Grid Way. Journal of Parallel and Distributed Computing, 66(5):763–771, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Varma:1986:PPG

[VR86] A. Varma and C. S. Raghavendra. On permutations passable by the Gamma network. Journal of Parallel and Distributed Computing, 3(1):72–91, March 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Venkatesan:1994:CAF

[VR94] S. Venkatesan and K. V. S. Ramarao. Computing associative functions distributively in spite of link failures. Journal of Parallel and Distributed Computing, 23(3):399–410, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1149/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1149/production/
pdf.

Venkatesan:1995:MFS

[VR95] S. Venkatesan and K. V. S. Ramarao. Making fault-sensitive algorithms tolerate link failures. Journal of Parallel and Distributed Computing, 30(1):64–75, October 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1126/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1126/production/pdf.

vHanxleden:1991:LBM

[vS91] Reinhard v. Hanxleden and L. Ridgeway Scott. Load balancing on message passing architectures. Journal of Parallel and Distributed Computing, 13(3):312–324, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vijaykumar:1999:TSM

[VS99] T. N. Vijaykumar and Gurindar S. Sohi. Task selection for the multiscalar architecture. Journal of Parallel and Distributed Computing, 58(2):132–158, August 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1550/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1550/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1557/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1557/production/pdf;   http:

//www.idealibrary.com/links/doi/10.1006/jpdc.1999.1557/
production/ref.

Varman:1991:MML

[VSIR91] Peter J. Varman, Scott D. Scheufler, Balakrishna R. Iyer, and Gary R. Ricard. Merging multiple lists on hierarchical-memory multiprocessors. Journal of Parallel and Distributed Computing, 12(2):171–177, June 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

VanderWijngaart:1996:AOS

[VSM96] Rob F. Van der Wijngaart, Sekhar R. Sarukkai, and Pankaj Mehra. Analysis and optimization of software pipeline performance on MIMD parallel computers. Journal of Parallel and Distributed Computing, 38(1):37–50, October 10, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0127/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0127/production/pdf.

Veen:1990:RCD

[VV90] A. H. Veen and R. Van Den Born. The RC compiler for the DTN dataflow computer. Journal of Parallel and Distributed Computing, 10(4):319–332, December 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Verhoosel:1996:ITC

[VWHL96] Jack P. C. Verhoosel, Lonnie R. Welch, Dieter K. Hammer, and Erik J. Luit. Incorporating temporal considerations during assignment and pre-run-time scheduling of objects and processes. Journal of Parallel and Distributed Computing, 36(1):13–31, July 10, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0087/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0087/production/pdf.

West:2002:GAA

[WA02] Jack M. West and John K. Antonio. A genetic-algorithm approach to scheduling communications for embedded parallel space-time adaptive processing algorithms. Journal of Parallel and Distributed Computing, 62(9):1386–1406, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1852;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1852/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1852/ref.

Weissman:2003:ISB

[WAE03] Jon B. Weissman, Lakshman Rao Abburi, and Darin England. Integrated scheduling: the best of both worlds. Journal of Parallel and Distributed Computing, 63(6):649–668, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wagner:1989:EAB

[Wag89] Alan Wagner. Embedding arbitrary binary trees in a hypercube. Journal of Parallel and Distributed Computing, 7(3): 503–520, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wagner:1993:EAB

[Wag93] A. S. Wagner. Embedding all binary trees in the hypercube. Journal of Parallel and Distributed Computing, 18(1): 33–43, May 1993. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1993.1042/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1993.1042/
production/pdf.

Wagner:1994:ECT

[Wag94] A. S. Wagner. Embedding the complete tree in the hypercube. Journal of Parallel and Distributed Computing, 20(2):241–247, February 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1023/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1023/production/pdf.

Wang:1996:LBM

[Wan96] Sying-Jyan Wang. Load-balancing in multistage interconnection networks under multiple-pass routing. Journal of Parallel and Distributed Computing, 36(2):189–194, August 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0099/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0099/production/
pdf.

Wang:2001:EHC

[Wan01a] Dajin Wang. Embedding Hamiltonian cycles into folded hypercubes with faulty links. Journal of Parallel and Distributed Computing, 61(4):545–564, April 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1681;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1681/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1681/ref.

Wang:2001:DDM

[Wan01b] Sying-Jyan Wang. Distributed diagnosis in multistage interconnection networks. Journal of Parallel and Distributed Computing, 61(2):254–264, February 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1675;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1675/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1675/ref.

Wang:2006:LWC

[Wan06] Jun Wang. A light-weight, collaborative temporary file system for clustered Web servers. Journal of Parallel and Distributed Computing, 66(5):750–762, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2007:ETA

[Wan07] Yuh-Rau Wang. An efficient O(1) time 3D all nearest neighbor algorithm from image processing perspective. Journal of Parallel and Distributed Computing, 67(10):1082–1091, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wah:1988:SDB

[WAS88] Benjamin W. Wah, Mokhtar Aboelaze, and Weijia Shang. Systematic designs of buffers in macropipelines of systolic arrays. Journal of Parallel and Distributed Computing, 5(1):1–25, February 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

White:1995:PNP

[WAS95] S. White, A. Alund, and V. S. Sunderam. Performance of the NAS parallel benchmarks on PVM-Based networks. Journal of Parallel and Distributed Computing, 26(1):61–71, April 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1048/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1048/production/
pdf.

Walinsky:1994:DPF

[WB94] C. Walinsky and D. Banerjee. A data-parallel FP compiler. Journal of Parallel and Distributed Computing, 22(2):138–153, August 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1077/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1994.1077/
production/pdf.

Wang:1996:PEI

[WB96] X. Wang and E. K. Blum. Parallel execution of iterative computations on workstation clusters. Journal of Parallel and Distributed Computing, 34(2):218–226, May 1, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0058/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0058/production/pdf.

Wang:2001:MWS

[WB01] Honge Wang and Douglas M. Blough. Multicast in wormholeswitched torus networks using edge-disjoint spanning trees. Journal of Parallel and Distributed Computing, 61(9):1278– 1306, September 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2001.1751;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2001.
1751/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2001.1751/ref.

Wang:2009:TPS

[WBTM09] Guoqiang Wang, Ladislau Bölöni, Damla Turgut, and Dan C. Marinescu. Time-parallel simulation of wireless ad hoc networks with compressed history. Journal of Parallel and Distributed Computing, 69(2):168–179, February 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:1991:COP

[WC91] Biing-Feng Wang and Gen-Huey Chen. Cost-optimal parallel algorithms for constructing 2–3 trees. Journal of Parallel and Distributed Computing, 11(3):257–261, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2005:GEI

[WC05] Jie Wu and Mihaela Cardei. Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 65(2):95–97, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2002:DHP

[WCC02] Nen-Chung Wang, Chih-Ping Chu, and Tzung-Shi Chen. A dual-Hamiltonian-path-based multicasting strategy for wormhole-routed star graph interconnection networks. Journal of Parallel and Distributed Computing, 62(12):1747–1762, December 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Walshaw:1997:PDG

[WCE97]  C. Walshaw, M. Cross, and M. G. Everett.  Paral-

lel dynamic graph partitioning for adaptive unstructured meshes. Journal of Parallel and Distributed Computing, 47(2):102–108, December 15, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1407/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1407/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1407/production/ref.

Wang:1994:SPC

[WCF94] Yi-Min M. Wang, Pi-Yu Y. Chung, and W. Kent Fuchs. Scheduling for periodic concurrent error detection in processor arrays. Journal of Parallel and Distributed Computing, 23(3):306–313, December 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1142/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1142/production/pdf.

Wang:2006:PLS

[WCKD06] Kai Wang, Anthony Chang, Laxmikant V. Kale, and Jonathan A. Dantzig. Parallelization of a level set method for simulating dendritic growth. Journal of Parallel and Distributed Computing, 66(11):1379–1386, November 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Williams:2009:OLB

[WCO^{+}09] Samuel Williams, Jonathan Carter, Leonid Oliker, John Shalf, and Katherine Yelick. Optimization of a lattice Boltzmann computation on state-of-the-art multicore platforms. Journal of Parallel and Distributed Computing, 69(9):762–777, September 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2003:DAF

[WCWH03] Fu-Hsing Wang, Jou-Ming Chang, Yue-Li Wang, and Sun-Jen Huang. Distributed algorithms for finding the unique minimum distance dominating set in directed split-stars. Journal of Parallel and Distributed Computing, 63(4):481–487, April 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2008:UAZ

[WCYR08] Weigang Wu, Jiannong Cao, Jin Yang, and Michel Raynal. Using asynchrony and zero degradation to speed up indulgent consensus protocols. Journal of Parallel and Distributed Computing, 68(7):984–996, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wong:1992:TBP

[WD92] Yiwan Wong and Jean-Marc Delosme. Transformation of broadcasts into propagations in systolic algorithms. Journal of Parallel and Distributed Computing, 14(2):121–145, February 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Weemeeuw:1994:FGP

[WD94] Patrick Weemeeuw and Bart Demoen. A fine-grained parallel memory compaction algorithm. Journal of Parallel and Distributed Computing, 20(2):176–186, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1017/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1017/production/pdf.

Wang:2009:CFC

[WDDK09] Zhijun Wang, Anwitaman Datta, Sajal K. Das, and Mohan Kumar. CMV: File consistency maintenance through virtual servers in peer-to-peer systems. Journal of Parallel and Distributed Computing, 69(4):360–372, April 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Weems:2001:SII

[Wee01] Charles C. Weems, Jr. Special issue: International Parallel and Distributed Processing Symposium 2000. Journal of Parallel and Distributed Computing, 61(12):1707–1708, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1785;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1785/pdf.

Weissman:1998:GBW

[Wei98] Jon B. Weissman. Gallop: The benefits of wide-area computing for parallel processing. Journal of Parallel and Distributed Computing, 54(2):183–205, November 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1487/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1487/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1487/production/ref.

Weissman:2002:PCB

[Wei02] Jon B. Weissman. Predicting the cost and benefit of adapting data parallel applications in clusters. Journal of Parallel and Distributed Computing, 62(8):1248–1271, August 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1838;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1838/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1838/ref.

Wong:1989:PAP

[WF89] Ken Wong and Mark A. Franklin. Performance analysis of a parallel logic simulation machine. Journal of Parallel and Distributed Computing, 7(3):416–440, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wise:1990:CQR

[WF90] David S. Wise and John Franco. Costs of quadtree representation of nondense matrices. Journal of Parallel and Distributed Computing, 9(3):282–296, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wolski:1993:PPN

[WF93] Richard M. Wolski and John T. Feo. Program partitioning for NUMA multiprocessor computer systems. Journal of Parallel and Distributed Computing, 19(3):203–218, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1993.1105/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1993.1105/production/
pdf.

Wong:1996:CDC

[WF96] Kenneth F. Wong and Mark Franklin. Checkpointing in distributed computing systems. Journal of Parallel and Distributed Computing, 35(1):67–75, May 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0069/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0069/production/pdf.

Wu:1998:EBT

[WFL98] Jie Wu, Eduardo B. Fernandez, and Yingqiu Luo. Embedding of binomial trees in hypercubes with link faults. Journal of Parallel and Distributed Computing, 54(1):59–74, October 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1482/production;   http://www.

idealibrary.com/links/doi/10.1006/jpdc.1998.1482/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1482/production/ref.

Wing:1993:TVC

[WG93] Jeannette M. Wing and Chun Gong. Testing and verifying concurrent objects. Journal of Parallel and Distributed Computing, 17(1–2):164–182, January/February 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1015/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1015/production/pdf.

Wang:2008:CAG

[WG08] Yong Wang and Qian-Ping Gu. On the complexity and algorithm of grooming regular traffic in WDM optical networks. Journal of Parallel and Distributed Computing, 68(6):877–886, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2009:ABP

[WGC09] Jun Wang, Peng Gu, and Hailong Cai. An advertisement-based peer-to-peer search algorithm. Journal of Parallel and Distributed Computing, 69(7):638–651, July 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2009:PPI

[WGCZ09] Qishi Wu, Jinzhu Gao, Zizhong Chen, and Mengxia Zhu. Pipelining parallel image compositing and delivery for efficient remote visualization. Journal of Parallel and Distributed Computing, 69(3):230–238, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2008:TMR

[WGS08] Yongwei Wang, Venkata C. Giruka, and Mukesh Singhal. Truthful multipath routing for ad hoc networks with selfish nodes. Journal of Parallel and Distributed Computing, 68(6): 778–789, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:1997:EPS

[WH97] Poting Wu and Elias N. Houstis. EPPOD: a problem solving environment for parallel electronic prototyping of physical object design. Journal of Parallel and Distributed Computing, 42(2):157–172, May 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1323/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1323/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1323/production/ref.

Wu:2008:IWM

[WH08] Tsai-Wei Wu and Hung-Yun Hsieh. Interworking wireless mesh networks: Problems, performance characterization, and perspectives. Journal of Parallel and Distributed Computing, 68 (3):348–360, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2000:EPA

[WHT00] Chin-Hsiung Wu, Shi-Jinn Horng, and Horng-Ren Tsai. Efficient parallel algorithms for hierarchical clustering on arrays with reconfigurable optical buses. Journal of Parallel and Distributed Computing, 60(9):1137–1153, September 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1644;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1644/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1644/ref.

Wu:2002:OPA

[WHT02] Chin-Hsiung Wu, Shi-Jinn Horng, and Horng-Ren Tsai. Optimal parallel algorithms for computer vision problems. Journal of Parallel and Distributed Computing, 62(6):1021–1041, June 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1821;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1821/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1821/ref.

Wada:1997:HFT

[WIKC97] Koichi Wada, Takaharu Ikeo, Kimio Kawaguchi, and Wei Chen. Highly fault-tolerant routings and fault-induced diameter for generalized hypercube graphs. Journal of Parallel and Distributed Computing, 43(1):57–62, May 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1326/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1326/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1326/production/ref.

Wilkinson:1990:CRC

[Wil90] B. Wilkinson. Cascaded rhombic crossbar interconnection networks. Journal of Parallel and Distributed Computing, 10(1): 96–101, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wilkinson:1992:OCI

[Wil92] Barry Wilkinson. Overlapping connectivity interconnection networks for shared memory multi-processor systems. Journal of Parallel and Distributed Computing, 15(1):49–61, May 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wing:1985:CAS

[Win85] Omar Wing. Content-addressable systolic array for sparse matrix computation. Journal of Parallel and Distributed Computing, 2(2):170–181, May 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Weil:1991:DIS

[WJD91] Frank Weil, Leah H. Jamifson, and Edward J. Delp. Dynamic intelligent scheduling and control of reconfigurable parallel architectures for computer vision/image processing. Journal of Parallel and Distributed Computing, 13(3):273–285, November 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2007:ESP

[WJV07] Qin Wang, Joseph JaJa, and Amitabh Varshney. An efficient and scalable parallel algorithm for out-of-core isosurface extraction and rendering. Journal of Parallel and Distributed Computing, 67(5):592–603, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wah:1990:OPE

[WL90] B. W. Wah and Guo-Jie Li. Optimal parallel evaluation of AND trees. Journal of Parallel and Distributed Computing, 10 (1):1–17, September 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wilson:1992:HSM

[WL92] Andrew W. Wilson, Jr. and Richard P. LaRowe, Jr. Hiding shared memory reference latency on the Galactica Net distributed shared memory architecture. Journal of Parallel and Distributed Computing, 15(4):351–367, August 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2004:COS

[WL04] Cheng Wang and Zhiyuan Li. A computation offloading scheme on handheld devices. Journal of Parallel and Distributed Computing, 64(6):740–746, June 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2005:WRT

[WL05] Xiao Wu and Victor C. S. Lee. Wireless real-time on-demand data broadcast scheduling with dual deadlines. Journal of Parallel and Distributed Computing, 65(6):714–728, June 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2000:RCC

[WLD00] Fen Lin Wu, S. Lakshmivarahan, and S. K. Dhall. Routing in a class of Cayley graphs of semidirect products of finite groups. Journal of Parallel and Distributed Computing, 60(5):539–565, May 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1624;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1624/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1624/ref.

Wong:2002:GAP

[WLID02] Allan K. Y. Wong, Wilfred W. K. Lin, May T. W. Ip, and Tharam S. Dillon. Genetic algorithm and PID control together for dynamic anticipative marginal buffer management:

An effective approach to enhance dependability and performance for distributed mobile object-based real-time computing over the Internet. Journal of Parallel and Distributed Computing, 62(9):1433–1453, September 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1855;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1855/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1855/ref.

Wu:2008:ORP

[WLL08] Jan-Jan Wu, Yi-Fang Lin, and Pangfeng Liu. Optimal replica placement in hierarchical Data Grids with locality assurance. Journal of Parallel and Distributed Computing, 68(12):1517– 1538, December 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Weatherly:2006:DMS

[WLNL06] D. Brent Weatherly, David K. Lowenthal, Mario Nakazawa, and Franklin Lowenthal. Dyn-MPI: Supporting MPI on mediumscale, non-dedicated clusters. Journal of Parallel and Distributed Computing, 66(6):822–838, June 2006. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Willebeek-LeMair:1990:SNP

[WLR90] M. Willebeek-LeMair and A. P. Reeves. Solving nonuniform problems on SIMD computers: Case study on region growing. Journal of Parallel and Distributed Computing, 8(2):135–149, February 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:2009:OSP

[WLWW09] Jan-Jan Wu, Yi-Fang Lin, Da-Wei Wang, and Chien-Min Wang. Optimizing server placement for parallel I/O in switchbased clusters. Journal of Parallel and Distributed Computing, 69(3):266–281, March 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wan:2001:ORB

[WLY01] Peng-Jun Wan, Liwu Liu, and Yuanyuan Yang. Optimal routing based on super topology in optical parallel interconnect. Journal of Parallel and Distributed Computing, 61(9):1209–1224, September 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1750;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1750/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1750/ref.

Wang:1992:PSN

[WM92] Li-Xin Wang and Jerry M. Mendel. Parallel structured networks for solving a wide variety of matrix algebra problems. Journal of Parallel and Distributed Computing, 14(3):236–247, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Weissman:2001:ORF

[WMG01] Jon B. Weissman, Mahesh Marina, and Michael Gingras. Optimizing remote file access for parallel and distributed network applications. Journal of Parallel and Distributed Computing, 61(11):1591–1608, November 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1744;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1744/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1744/ref.

Wang:2009:PEE

[WMW09] Xue Wang, Junjie Ma, and Sheng Wang. Parallel energyefficient coverage optimization with maximum entropy clustering in wireless sensor networks. Journal of Parallel and Distributed Computing, 69(10):838–847, October 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Woo:1994:ORA

[WN94] Nam-Sung S. Woo and B. Naylor. Optimal routing algorithms for a class of cylindrical Banyan multicomputers. Journal of Parallel and Distributed Computing, 23(2):239–245, November 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1135/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1135/production/
pdf.

Wang:1994:MQF

[WNA^{+}94] Mu-Cheng Wang, Wayne G. Nation, James B. Armstrong, Howard Jay Siegel, Shin Dug Kim, Mark A. Nichols, and Michael Gherrity. Multiple quadratic forms: a case study in the design of data-parallel algorithms. Journal of Parallel and Distributed Computing, 21(1):124–139, April 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1046/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1046/production/pdf.

Wolfe:1988:VOV

[Wol88] M. Wolfe. Vector optimization vs vectorization. Journal of Parallel and Distributed Computing, 5(5):551–567, October 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wong:1999:HMM

[Won99] Raymond K. Wong. Heterogeneous and multifaceted multimedia objects in DOOR/MM: a role-based approach with views. Journal of Parallel and Distributed Computing, 56(3):251–271, March 1999. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1522/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1522/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1522/production/ref.

Worland:1993:PMO

[Wor93] Peter B. Worland. Parallel methods for ODEs with improved absolute stability boundaries. Journal of Parallel and Distributed Computing, 18(1):25–32, May 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1041/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1041/production/pdf.

Warnakulasuriya:2002:CDI

[WP02] Sugath Warnakulasuriya and Timothy Mark Pinkston. Characterization of deadlocks in irregular networks. Journal of Parallel and Distributed Computing, 62(1):61–84, January 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1782;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2001.1782/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1782/ref.

Wang:1994:SDP

[WPKK94] Cho-Li Wang, Viktor K. Prasanna, Hyoung J. Kim, and Ashfaq A. Khokhar. Scalable data parallel implementations of object recognition using geometric hashing. Journal of Parallel and Distributed Computing, 21(1):96–109, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1044/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1044/production/pdf.

Williams:1995:CAA

[WR95] C. C. Williams and P. F. Reynolds, Jr. Combining atomic actions. Journal of Parallel and Distributed Computing, 24(2):152–163, February 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1015/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1015/production/pdf.

Wang:1997:OCA

[WR97] Xiaodong Wang and Vwani P. Roychowdhury. Optimal communication algorithms for heterogeneous computing over ATM networks. Journal of Parallel and Distributed Computing, 46(1):62–79, October 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1388/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1388/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1388/production/ref.

Wei:2002:ESS

[WRC^{+}02] David S. L. Wei, Sanguthevar Rajasekaran, Z. Cheng, K. Naik, and Sy-Yen Kuo. Efficient selection and sorting schemes using coteries for processing large distributed files. Journal of Parallel and Distributed Computing, 62(8):1295–1313, August 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2002.1861;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2002.1861/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2002.1861/ref.

Weems:1991:DIU

[WRHR91] Charles Weems, Edward Riseman, Allen Hanson, and Azriel Rosenfeld. The DARPA image understanding benchmark for parallel computers. Journal of Parallel and Distributed Computing, 11(1):1–24, January 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wright:1991:PAG

[Wri91] William E. Wright. Parallel algorithms for generating the raster representation of straight lines and circles. Journal of Parallel and Distributed Computing, 11(2):170–173, February 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wittenbrink:1995:OPI

[WS95] Craig M. Wittenbrink and Arun K. Somani. 2D and 3D optimal parallel image warping. Journal of Parallel and Distributed Computing, 25(2):197–208, March 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1040/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1040/production/pdf.

Wittenbrink:1997:TSO

[WS97a] Craig M. Wittenbrink and Arun K. Somani. Time and space optimal data parallel volume rendering using permutation warping. Journal of Parallel and Distributed Computing, 46(2):148–164, November 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1386/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1386/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1386/production/ref.

Wu:1997:DMD

[WS97b] Min-You Wu and Wei Shu. DDE: a modified dimension exchange method for load balancing in k-ary n-cubes. Journal of Parallel and Distributed Computing, 44(1):88–96, July 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1345/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1345/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1345/production/ref.

Wu:2006:GHS

[WS06] Ming Wu and Xian-He Sun. Grid harvest service: a performance system of Grid computing. Journal of Parallel and Distributed Computing, 66(10):1322–1337, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Watson:1994:BBM

[WSA^{+}94] D. W. Watson, H. J. Siegel, J. K. Antonio, M. A. Nichols, and M. J. Atallah. A block-based mode selection model for SIMD/SPMD parallel environments. Journal of Parallel and Distributed Computing, 21(3):271–288, June 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1059/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1059/production/pdf.

Wei:1991:DSG

[WSG91] Yi-Hsiu Wei, Alexander D. Stoyenko, and German S. Goldszmidt. The design of a stub generator for heterogeneous RPC systems. Journal of Parallel and Distributed Computing, 11 (3):188–197, March 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2003:CAM

[WSH^{+}03] Ruye Wang, Lucas Scharenbroich, Christopher Hart, Barbara Wold, and Eric Mjolsness. Clustering analysis of microarray gene expression data by splitting algorithm. Journal of Parallel and Distributed Computing, 63(7–8):692–706, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:1997:TMS

[WSRM97] Lee Wang, Howard Jay Siegel, Vwani P. Roychowdhury, and Anthony A. Maciejewski. Task matching and scheduling in heterogeneous computing environments using a geneticalgorithm-based approach. Journal of Parallel and Distributed Computing, 47(1):8–22, November 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1392/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1392/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1392/production/ref.

Weiss:1993:AID

[WSS93] Shlomit Weiss, Ilan Y. Spillinger, and Gabriel M. Silberman. Architectural improvements for a data-driven VLSI processing array. Journal of Parallel and Distributed Computing, 19(4):308–322, December 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1114/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1114/production/pdf.

Wu:1992:CAL

[WT92] Chwan-Hwa (John) Wu and Jyun-Hwei Tsai. Concurrent asynchronous learning algorithms for massively parallel recurrent neural networks. Journal of Parallel and Distributed Computing, 14(3):345–353, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2009:GDC

[WT09] Xianbing Wang and Yong Meng Teo. Global data computation in chordal rings. Journal of Parallel and Distributed Computing, 69(8):725–736, August 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2008:IRP

[WTB^{+}08] Guoqiang Wang, Damla Turgut, Ladislau Bölöni, Yongchang Ji, and Dan C. Marinescu. Improving routing performance through m-limited forwarding in power-constrained wireless ad hoc networks. Journal of Parallel and Distributed Computing, 68(4):501–514, April 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2008:CMT

[WTC08a] Xianbing Wang, Yong Meng Teo, and Jiannong Cao. Corrigendum to “Message and time efficient consensus protocols for synchronous distributed systems” [J. Parallel Distrib. Comput. 68 (2008) 641–654]. Journal of Parallel and Distributed Computing, 68(6):876, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See [WTC08b].

Wang:2008:MTE

[WTC08b] Xianbing Wang, Yong Meng Teo, and Jiannong Cao. Message and time efficient consensus protocols for synchronous distributed systems. Journal of Parallel and Distributed Computing, 68(5):641–654, May 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). See corrigendum [WTC08a].

Wu:2003:SLS

[WTS03] Yun Wu, Terence D. Todd, and Shahram Shirani. SCO link sharing in Bluetooth voice access networks. Journal of Parallel and Distributed Computing, 63(1):45–57, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:1985:ETN

[Wu85] Angela Y. Wu. Embedding of tree networks into hypercubes. Journal of Parallel and Distributed Computing, 2(3):238–249, August 1985. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wu:1994:OFT

[Wu94] Jie Wu. An optimal fault-tolerant nonredundant broadcasting scheme in injured hypercubes. Journal of Parallel and Distributed Computing, 22(2):295–313, August 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1088/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1088/production/pdf.

Wu:2002:DFS

[Wu02] Jie Wu. A distributed formation of smallest faulty orthogonal convex polygons in 2-D meshes. Journal of Parallel and Distributed Computing, 62(7):1168–1185, July 1, 2002. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1842;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1842/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1842/ref.

Wu:2003:OBI

[Wu03] Jie Wu. Optimal broadcasting in injured hypercubes using directed safety levels. Journal of Parallel and Distributed Computing, 63(9):815–826, September 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Woo:1999:MSS

[WUG99] Miae Woo, Reha Uzsoy, and Arif Ghafoor. Media streams scheduling for synchronization in distributed multimedia systems. Journal of Parallel and Distributed Computing, 56 (3):272–295, March 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1998.1526/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1526/production/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1526/production/ref.

Wang:1996:SCB

[WW96] Paul Wang and William E. Weihl. Scalable concurrent Btrees using multi-version memory. Journal of Parallel and Distributed Computing, 32(1):28–48, January 10, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0003/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0003/production/pdf.

Wang:1997:MAT

[WW97] Dajin Wang and Zhongxian Wang. Minimum assignment of test links for hypercubes with lower fault bounds. Journal of Parallel and Distributed Computing, 40(2):185–193, February 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1279/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1279/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1279/production/ref.

Wattenhofer:1998:IBD

[WW98] Roger Wattenhofer and Peter Widmayer. An inherent bottleneck in distributed counting. Journal of Parallel and Distributed Computing, 49(1):135–145, February 25, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1431/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1431/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1431/production/ref.

Weissman:2003:GEI

[WW03] Jon Weissman and Rich Wolski. Guest editor introduction: special issue on Computational Grids. Journal of Parallel and Distributed Computing, 63(5):503–504, May 2003. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wattenhofer:2004:CPA

[WW04] Roger Wattenhofer and Peter Widmayer. The counting pyramid: an adaptive distributed counting scheme. Journal of Parallel and Distributed Computing, 64(4):449–460, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wang:2007:CSP

[WW07] Xue Wang and Sheng Wang. Collaborative signal processing for target tracking in distributed wireless sensor networks. Journal of Parallel and Distributed Computing, 67(5):501–515, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2000:PDC

[wXH00] Chong wei Xu and Bo He. PCB — a distributed computing system in CORBA. Journal of Parallel and Distributed Computing, 60(10):1293–1310, October 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1654;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1654/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1654/ref.

Wang:2005:AER

[WXZ05] Shengquan Wang, Dong Xuan, and Wei Zhao. Analyzing and enhancing the resilience of structured peer-to-peer systems. Journal of Parallel and Distributed Computing, 65(2):207–219, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Wein:1991:MPS

[WZ91] Joel M. Wein and Stavros A. Zenios. On the massively parallel solution of the assignment problem. Journal of Parallel and Distributed Computing, 13(2):228–236, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xiao:2007:DAC

[XBK07] Lin Xiao, Stephen Boyd, and Seung-Jean Kim. Distributed average consensus with least-mean-square deviation. Journal of Parallel and Distributed Computing, 67(1):33–46, January 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xiao:2008:ADA

[XCH08] Bin Xiao, Wei Chen, and Yanxiang He. An autonomous defense against SYN flooding attacks: Detect and throttle attacks at the victim side independently. Journal of Parallel and Distributed Computing, 68(4):456–470, April 2008. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xing:2007:LCS

[XCLR07] Kai Xing, Xiuzhen Cheng, Fang Liu, and Shmuel Rotenstreich. Location-centric storage for safety warning based on roadway sensor networks. Journal of Parallel and Distributed Computing, 67(3):336–345, March 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xiang:2006:FTM

[XCS06] Dong Xiang, Ai Chen, and Jia-Guang Sun. Fault-tolerant multicasting in hypercubes using local safety information. Journal of Parallel and Distributed Computing, 66(2):248–256, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xiao:2003:SLA

[XCZL03] Li Xiao, Xin Chen, Xiaodong Zhang, and Yunhao Liu. On scalable and locality-aware Web document sharing. Journal of Parallel and Distributed Computing, 63(10):945–962, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xue:2003:AHQ

[XG03] Qi Xue and Aura Ganz. Ad hoc QoS on-demand routing (AQOR) in mobile ad hoc networks. Journal of Parallel and Distributed Computing, 63(2):154–165, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1991:MRB

[XH91] Jian Xu and Kai Hwang. Mapping rule-based systems onto multicomputers using simulated annealing. Journal of Parallel and Distributed Computing, 13(4):442–455, December 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1993:HMD

[XH93] Jian Xu and Kai Hwang. Heuristic methods for dynamic load balancing in a message-passing multicomputer. Journal of Parallel and Distributed Computing, 18(1):1–13, May 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1039/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1039/production/pdf.

Xu:2003:LRA

[XHG03] Kaixin Xu, Xiaoyan Hong, and Mario Gerla. Landmark routing in ad hoc networks with mobile backbones. Journal of Parallel and Distributed Computing, 63(2):110–122, February 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xia:2007:DIR

[XHY07] Zhonghang Xia, Wei Hao, and I-Ling Yen. A distributed integrated request processing algorithm for QoS assurance in largescale media-based systems. Journal of Parallel and Distributed Computing, 67(7):863–875, July 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2003:SIS

[XJS03] Cheng-Zhong Xu, Hai Jin, and Pradip K. Srimani. Special issue on scalable Web services and architecture. Journal of Parallel and Distributed Computing, 63(10):913–915, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1994:CCP

[XKMN94] Hong Xu, E. T. Kalns, P. K. McKinley, and L. M. Ni. ComPaSS: a communication package for scalable software design. Journal of Parallel and Distributed Computing, 22(3):449–461, September 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1103/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1103/production/pdf.

Xu:1992:AGD

[XL92] C. Z. Xu and F. C. M. Lau. Analysis of the generalized dimension exchange method for dynamic load balancing. Journal of Parallel and Distributed Computing, 16(4):385–393, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1995:GDE

[XL95] Cheng-Zhong Z. Xu and F. C. M. Lau. The generalized dimension exchange method for load balancing in κary n-cubes and variants. Journal of Parallel and Distributed Computing, 24(1):72–85, January 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1995.1007/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1995.1007/production/pdf.

Xiao:2006:MAO

[XLG^{+}06] Li Xiao, Yunhao Liu, Wenjun Gu, Dong Xuan, and Xiaomei Liu. Mutual anonymous overlay multicast. Journal of Parallel and Distributed Computing, 66(9):1205–1216, September 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1992:EIB

[XMN92] Hong Xu, Philip K. McKinley, and Lionel M. Ni. Efficient implementation of barrier synchronization in wormhole-routed hypercube multicomputers. Journal of Parallel and Distributed Computing, 16(2):172–184, October 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xue:2005:LRL

[XO05] Liyin Xue and Mehmet A. Orgun. Locking without requesting a lock: a consistency maintenance mechanism in Internet-based real-time group editors. Journal of Parallel and Distributed Computing, 65(7):801–814, July 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:2004:DCP

[XQ04] Yingyue Xu and Hairong Qi. Distributed computing paradigms for collaborative signal and information processing in sensor networks. Journal of Parallel and Distributed Computing, 64 (8):945–959, August 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xie:2007:PEN

[XQ07] Tao Xie and Xiao Qin. Performance evaluation of a new scheduling algorithm for distributed systems with security heterogeneity. Journal of Parallel and Distributed Computing, 67 (10):1067–1081, October 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xue:1997:CMT

[Xue97] Jingling Xue. Communication-minimal tiling of uniform dependence loops. Journal of Parallel and Distributed Computing, 42(1):42–59, April 10, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1310/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1310/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1310/production/ref.

Xu:2008:DDM

[XWC^{+}08] WeiQiang Xu, YaMing Wang, JiMing Chen, George Baciu, and Youxian Sun. Dual decomposition method for optimal and fair congestion control in ad hoc networks: Algorithm, implementation and evaluation. Journal of Parallel and Distributed Computing, 68(7):997–1007, July 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xia:2006:IAC

[XYDL06] Zhonghang Xia, I-Ling Yen, Donglei Du, and Peng Li. An integrated admission control scheme for the delivery of streaming media. Journal of Parallel and Distributed Computing, 66 (3):334–344, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xiao:2007:CIS

[XYG07] Bin Xiao, Bo Yu, and Chuanshan Gao. CHEMAS: Identify suspect nodes in selective forwarding attacks. Journal of Parallel and Distributed Computing, 67(11):1218–1230, November 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xie:2008:LBM

[XYKA08] Bin Xie, Yingbing Yu, Anup Kumar, and Dharma P. Agrawal. Load-balanced mesh router migration for wireless mesh networks. Journal of Parallel and Distributed Computing, 68(6): 825–839, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Xu:1996:SEM

[XZS96] Zhichen Xu, Xiaodong Zhang, and Lin Sun. Semi-empirical multiprocessor performance predictions. Journal of Parallel and Distributed Computing, 39(1):14–28, November 25, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0151/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0151/production/
pdf.

Yang:1993:PCM

[Yan93] Qing Yang. Performance of cache memories for vector computers. Journal of Parallel and Distributed Computing, 19(3):163–178, November 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/

jpdc.1993.1102/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1102/production/pdf.

Yang:2000:PMB

[Yan00] Yuanyuan Yang. The performance of multicast Banyan networks. Journal of Parallel and Distributed Computing, 60 (8):909–923, August 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.2000.1641;
http://www.idealibrary.com/links/doi/10.1006/jpdc.2000.
1641/pdf;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.2000.1641/ref.

Yang:2004:FPO

[Yan04] Xiaofan Yang. A fast pessimistic one-step diagnosis algorithm for hypercube multicomputer systems. Journal of Parallel and Distributed Computing, 64(4):546–553, April 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2009:CCA

[Yan09] Haitao Yang. A compensation cost analysis of service-aggregate transaction for DTNs clients. Journal of Parallel and Distributed Computing, 69(11):930–936, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:1998:POC

[YAS98] Qing Yang, Sridar Adina, and T. Sun. Performance of one’s complement caches. Journal of Parallel and Distributed Computing, 48(2):143–164, February 1, 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1415/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1415/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1415/production/ref.

Yang:1990:PMB

[YB90] Qing Yang and Laxmi N. Bhuyan. Performance of multiple-bus interconnections for multiprocessors. Journal of Parallel and Distributed Computing, 8(3):267–273, March 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yen:1995:PHC

[YB95] I-Ling L. Yen and Farokh B. Bastani. Parallel hashing: Collision resolution strategies and performance. Journal of Parallel and Distributed Computing, 31(2):190–198, December 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1995.1158/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1995.1158/production/
pdf.

Yuan:2001:PIF

[YB01] Yanhong Yuan and Prith Banerjee. A parallel implementation of a fast multipole-based 3-D capacitance extraction program on distributed memory multicomputers. Journal of Parallel and Distributed Computing, 61(12):1751–1774, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1725;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1725/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1725/ref.

Yener:1997:FTC

[YBOY97] Bülent Yener, Inderpal Bhandari, Yoram Ofek, and Moti Yung. Fault-tolerant convergence routing. Journal of Parallel and Distributed Computing, 42(2):173–183, May 1, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1320/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1320/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1320/production/ref.

Yen:2004:MRS

[YC04] Li-Hsing Yen and Kuang-Hwei Chi. Maintaining a ring structure for mobile ad hoc computing. Journal of Parallel and Distributed Computing, 64(12):1371–1379, December 2004. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yin:2005:GTD

[YCC05] Liangzhong Yin, Guohong Cao, and Ying Cai. A generalized target-driven cache replacement policy for mobile environments. Journal of Parallel and Distributed Computing, 65 (5):583–594, May 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:1998:PIA

[yCM98] Chung yen Chang and Prasant Mohapatra. Performance improvement of allocation schemes for mesh-connected computers. Journal of Parallel and Distributed Computing, 52(1):40– 68, July 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1998.1459/production;   http:
//www.idealibrary.com/links/doi/10.1006/jpdc.1998.1459/
production/pdf;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1459/production/ref.

Yoo:2000:TRT

[YCY^{+}00] Seong-Moo Yoo, Hyunseung Choo, Hee Yong Youn, Chansu Yu, and Younghee Lee. On task relocation in twodimensional meshes. Journal of Parallel and Distributed Computing, 60(5):616–638, May 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1604;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1604/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1604/ref.

Yoo:1998:FEP

[YD98] Byung S. Yoo and Chita R. Das. A fast and efficient processor management scheme for k-ary n-cubes. Journal of Parallel and Distributed Computing, 55(2):192–214, December 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1498/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1498/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1498/production/ref.

Yen:2001:PUE

[Yen01] Li-Hsing Yen. Precluding useless events for on-line global predicate detections. Journal of Parallel and Distributed Computing, 61(8):1077–1095, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2001.1740;   http://www.idealibrary.com/links/

doi/10.1006/jpdc.2001.1740/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2001.1740/ref.

Yang:2007:CFD

[YF07] Mengkun Yang and Zongming Fei. A cooperative failure detection mechanism for overlay multicast. Journal of Parallel and Distributed Computing, 67(6):635–647, June 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2009:NAI

[YF09] Mengkun Yang and Zongming Fei. A novel approach to improving search efficiency in unstructured peer-to-peer networks. Journal of Parallel and Distributed Computing, 69(11):877–884, November 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yen:1997:RVC

[YH97] Li-Hsing Yen and Ting-Lu Huang. Resetting vector clocks in distributed systems. Journal of Parallel and Distributed Computing, 43(1):15–20, May 25, 1997. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1330/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1330/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1330/production/ref.

Yero:2007:SSA

[YH07] Eduardo Javier Huerta Yero and Marco Aurélio Amaral Henriques. Speedup and scalability analysis of master–slave applications on large heterogeneous clusters. Journal of Parallel and Distributed Computing, 67(11):1155–1167, November 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hsu:1997:PEW

[yHY97] William Tsun yuk Hsu and Pen-Chung Yew. Performance evaluation of wire-limited hierarchical networks. Journal of Parallel and Distributed Computing, 41(2):156–172, March 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.1296/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.1296/production/

pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.1296/production/ref.

Yang:1996:PPS

[YI96] Tao Yang and Oscar H. Ibarra. Performance prediction in symbolic scheduling of partitioned programs with weight variation. Journal of Parallel and Distributed Computing, 34(2):171–182, May 1, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0053/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0053/production/pdf.

Yu:1997:PPR

[YIY97] Yizhou Yu, Oscar H. Ibarra, and Tao Yang. Parallel progressive radiosity with adaptive meshing. Journal of Parallel and Distributed Computing, 42(1):30–41, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1309/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1309/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1309/production/ref.

Yau:1991:PPO

[YJB91] Stephen S. Yau, Xiaoping Jia, and Doo-Hwan Bae. PROOF: a parallel object-oriented functional computation model. Journal of Parallel and Distributed Computing, 12(3):202–212, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yau:2004:CSM

[YK04] Stephen S. Yau and Fariaz Karim. A context-sensitive middleware for dynamic integration of mobile devices with network infrastructures. Journal of Parallel and Distributed Computing, 64(2):301–317, February 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yoon:1989:BDN

[YL89] Hyunsoo Yoon and Kyungsook Y. Lee. B-banyan and B-delta networks for multiprocessor systems. Journal of Parallel and Distributed Computing, 7(3):570–582, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yue:1998:CPA

[YL98] Kelvin K. Yue and David J. Lilja. Comparing processor allocation strategies in multiprogrammed shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 49(2):245–258, March 15, 1998. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1443/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1998.1443/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1998.1443/production/ref.

Yoon:1990:MTP

[YLB90] Hyunsoo Yoon, Kyungsook Y. Lee, and Amos Bahiri. On the modulo M translators for the prime memory system. Journal of Parallel and Distributed Computing, 8(1):72–76, January 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2006:OSP

[YME06] Xiaofan Yang, Graham M. Megson, and David J. Evans. An oblivious shortest-path routing algorithm for fully connected cubic networks. Journal of Parallel and Distributed Computing, 66(10):1294–1303, October 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yuan:2001:PMH

[YMG01] Xin Yuan, Rami Melhem, and Rajiv Gupta. Performance of multi-hop communications using logical topologies on optical torus networks. Journal of Parallel and Distributed Computing, 61(6):748–766, June 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1711;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1711/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1711/ref.

Young-Myers:1993:EST

[YMR93] Helene Young-Myers and Louiqa Raschid. An experimental study of three dataflow paradigms in multithreaded database transitive closure algorithms on shared memory multiprocessors. Journal of Parallel and Distributed Computing, 18(3):371–389, July 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1071/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1071/production/pdf.

Youssef:1992:TPG

[YN92] Abdou Youssef and Bhagirath Narahari. Topological properties of generalized Banyan-Hypercube networks. Journal of Parallel and Distributed Computing, 14(1):98–103, January 1992. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yun:1996:HHN

[YP96] Sang-Kyun Yun and Kyu Ho Park. Hierarchical hypercube networks (HHN) for massively parallel computers. Journal of Parallel and Distributed Computing, 37(2):194–199, September 15, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0119/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0119/production/
pdf.

Yu:2008:ICL

[YSL08] Oliver Yu, Emir Saric, and Anfei Li. Integrated connectionlevel and packet-level QoS controls over wireless mesh networks. Journal of Parallel and Distributed Computing, 68(3):336–347, March 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2005:RRM

[YT05] Ruoyun Yang and Mitchell D. Theys. RMF: Resource monitoring framework for integrating active and passive monitoring tools in Grid environments. Journal of Parallel and Distributed Computing, 65(11):1419–1428, November 2005. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2007:HCL

[YTH07] Ming-Chien Yang, Jimmy J. M. Tan, and Lih-Hsing Hsu. Hamiltonian circuit and linear array embeddings in faulty kary n-cubes. Journal of Parallel and Distributed Computing, 67(4):362–368, April 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:1994:RRM

[YTR94] Pei-Ji J. Yang, Sing-Ban B. Tien, and C. S. Raghavendra. Reconfiguration of rings and meshes in faulty hypercubes. Journal of Parallel and Distributed Computing, 22(1):96–106, July 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1072/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1072/production/
pdf.

Yew:1991:SIS

[YW91] Pen-Chung Yew and Benjamin W. Wah. Special issue on shared-memory multiprocessors. Journal of Parallel and Distributed Computing, 12(2):85–86, June 1991. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2008:SOC

[YWD08] Haitao Yang, Zhenghua Wang, and Qinghua Deng. Scheduling optimization in coupling independent services as a Grid transaction. Journal of Parallel and Distributed Computing, 68 (6):840–854, June 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:2000:PCO

[YWP00] Yuanyuan Yang, Jianchao Wang, and Yi Pan. Permutation capability of optical multistage interconnection networks. Journal of Parallel and Distributed Computing, 60(1):72–91, January 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1595/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1999.1595/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1595/production/ref.

Yu:2009:EPB

[YZLT09] Kun-Ming Yu, Jiayi Zhou, Chun-Yuan Lin, and Chuan Yi Tang. Efficient parallel branch-and-bound algorithm for constructing minimum ultrametric trees. Journal of Parallel and Distributed Computing, 69(11):905–914, November 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Yang:1996:EPP

[YZS96] Yong Yang, Xiaodong Zhang, and Yongsheng Song. An effective and practical performance prediction model for parallel computing on nondedicated heterogeneous NOW. Journal of Parallel and Distributed Computing, 38(1):63–80, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1996.0129/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1996.0129/production/
pdf.

Yan:1996:FTC

[YZY96] Yong Yan, Xiaodong Zhang, and Haixu Yang. A fast token-chasing mutual exclusion algorithm in arbitrary network topologies. Journal of Parallel and Distributed Computing, 35(2):156–172, June 15, 1996. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0078/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0078/production/pdf.

Zwietering:1991:PBM

[ZA91] Patrick Zwietering and Emile Aarts. Parallel Boltzmann machines: a mathematical model. Journal of Parallel and Distributed Computing, 13(1):65–75, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2005:DPB

[ZA05] Qi Zhang and Dharma P. Agrawal. Dynamic probabilistic broadcasting in MANETs. Journal of Parallel and Distributed Computing, 65(2):220–233, February 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zaki:2001:PSM

[Zak01] Mohammed J. Zaki. Parallel sequence mining on sharedmemory machines. Journal of Parallel and Distributed Computing, 61(3):401–426, March 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1695;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1695/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1695/ref.

Zwietering:1994:MNL

[ZAW94] P. J. Zwietering, E. H. L. Aarts, and J. Wessels. The minimal number of layers of a perceptron that sorts. Journal of Parallel and Distributed Computing, 20(3):380–387, March 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1034/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1034/production/pdf.

Zhou:1997:PRO

[ZB97] B. B. Zhou and R. P. Brent. A parallel ring ordering algorithm for efficient one-sided Jacobi SVD computations. Journal of Parallel and Distributed Computing, 42(1):1–10, April 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1997.1304/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1997.1304/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1304/production/ref.

Zhou:2003:EMC

[ZB03] B. B. Zhou and R. P. Brent. An efficient method for computing eigenvalues of a real normal matrix. Journal of Parallel and Distributed Computing, 63(6):638–648, June 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2009:SIV

[ZB09] Xiaoyu Zhang and Chandrajit Bajaj. Scalable isosurface visualization of massive datasets on commodity off-the-shelf clusters. Journal of Parallel and Distributed Computing, 69(1):39– 53, January 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2005:AUM

[ZBF05] Xiao Zhang, Laxmi N. Bhuyan, and Wu-Chun Feng. Anatomy of UDP and M-VIA for cluster communication. Journal of Parallel and Distributed Computing, 65(10):1290–1298, October 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zou:2004:UAC

[ZC04] Y. Zou and K. Chakrabarty. Uncertainty-aware and coverageoriented deployment for sensor networks. Journal of Parallel and Distributed Computing, 64(7):788–798, July 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zeigler:2002:QBF

[ZCK^{+}02] Bernard P. Zeigler, Hyup J. Cho, Jeong G. Kim, Hessam S. Sarjoughian, and Jong. S. Lee. Quantization-based filtering in distributed discrete event simulation. Journal of Parallel and Distributed Computing, 62(11):1629–1647, November 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhou:2006:NFI

[ZDC06] Shuming Zhou, Ni Du, and Baoxing Chen. A new family of interconnection networks of odd fixed degrees. Journal of Parallel and Distributed Computing, 66(5):698–704, May 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zeppenfeld:1991:PSB

[Zep91] Klaus Zeppenfeld. Parallel search in bivalued game trees. Journal of Parallel and Distributed Computing, 12(3):290–296, July 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:1989:NPS

[ZFL89] Chuan-Qi Zhu, Zhixi Fang, and Xiaobo Li. A new parallel sorting approach with sorting memory module. Journal of Parallel and Distributed Computing, 7(3):482–502, December 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2007:SAT

[ZFS07] Xuehai Zhang, Jeffrey L. Freschl, and Jennifer M. Schopf. Scalability analysis of three monitoring and information systems: MDS2, R-GMA, and Hawkeye. Journal of Parallel and Distributed Computing, 67(8):883–902, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:2006:PCA

[ZFWF06] Dianhui Zhu, Yuriy Fofanov, Richard C. Willson, and George E. Fox. A parallel computing algorithm for 16S rRNA probe design. Journal of Parallel and Distributed Computing, 66(12): 1546–1551, December 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:1999:COP

[ZH99] Yingchun Zhu and Laurie J. Hendren. Communication optimizations for parallel C programs. Journal of Parallel and Distributed Computing, 58(2):301–332, August 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1554/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1554/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1999.1554/production/ref.

Zhu:2007:ESS

[ZH07] Yingwu Zhu and Yiming Hu. Efficient semantic search on DHT overlays. Journal of Parallel and Distributed Computing, 67 (5):604–616, May 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zomaya:2003:SOU

[ZHO03] Albert Y. Zomaya, Michael Haydock, and Stephan Olariu. Some observations on using meta-heuristics for efficient location management in mobile computing networks. Journal of Parallel and Distributed Computing, 63(1):33–44, January 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:1992:EPA

[Zhu92] Yahui Zhu. Efficient processor allocation strategies for meshconnected parallel computers. Journal of Parallel and Distributed Computing, 16(4):328–337, December 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhou:2008:RAO

[ZI08] Xiaobo Zhou and Dennis Ippoliti. Resource allocation optimization for quantitative service differentiation on server clusters. Journal of Parallel and Distributed Computing, 68(9): 1250–1262, September 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ziavras:1992:PEH

[Zia92] Sotirios G. Ziavras. On the problem of expanding hypercubebased systems. Journal of Parallel and Distributed Computing, 16(1):41–53, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zimmermann:1990:TAS

[Zim90] Karl-Heinz Zimmermann. The theory of acyclic systolic systems. Journal of Parallel and Distributed Computing, 9(1):26– 41, May 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zimmermann:1996:RAR

[Zim96] Chris Zimmermann. Reflections on an adaptable realtime metalevel architecture. Journal of Parallel and Distributed Computing, 36(1):81–89, July 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1996.0091/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1996.0091/production/pdf.

Zhu:2006:CCE

[ZJ06] Yifeng Zhu and Hong Jiang. CEFT: a cost-effective, faulttolerant parallel virtual file system. Journal of Parallel and Distributed Computing, 66(2):291–306, February 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:1994:PHO

[ZK94] Weixiong X. Zhang and Richard E. Korf. Parallel heap operations on an EREW PRAM. Journal of Parallel and Distributed Computing, 20(2):248–255, February 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1024/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1024/production/pdf.

Zeigler:1993:SEI

[ZL93] Bernard P. Zeigler and Ahmed Louri. A simulation environment for intelligent machine architectures. Journal of Parallel and Distributed Computing, 18(1):77–88, May 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1046/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1046/production/pdf.

Zaki:1997:CDL

[ZLP97] Mohammed Javeed Zaki, Wei Li, and Srinivasan Parthasarathy. Customized dynamic load balancing for a network of workstations. Journal of Parallel and Distributed Computing, 43(2):156–162, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

idealibrary.com/links/doi/10.1006/jpdc.1997.1339/production;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1339/production/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1339/production/ref.

Zheng:2001:GCP

[ZLPP01] S. Q. Zheng, Keqin Li, Yi Pan, and M. Cristina Pinotti. Generalized coincident pulse technique and new addressing schemes for time-division multiplexing optical buses. Journal of Parallel and Distributed Computing, 61(8):1033–1051, August 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1726;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1726/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1726/ref.

Zapata:1991:MGS

[ZLRP91] E. L. Zapata, J. A. Lamas, F. F. Rivera, and O. G. Plata. Modified Gram–Schmidt QR factorization on hypercube SIMD computers. Journal of Parallel and Distributed Computing, 12(1): 60–69, May 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ziavras:1994:AMS

[ZM94a] S. G. Ziavras and P. Meer. Adaptive multiresolution structures for image processing on parallel computers. Journal of Parallel and Distributed Computing, 23(3):475–483, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1159/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1159/production/
pdf.

Znati:1994:UFD

[ZM94b] T. F. Znati and R. G. Melhem. A uniform framework for dynamic load balancing strategies in distributed processing systems. Journal of Parallel and Distributed Computing, 23(2):246–255, November 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1994.1136/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1994.1136/production/pdf.

Zhang:2006:ILM

[ZMC06] Lei Zhang, Jogesh K. Muppala, and Samuel T. Chanson. Integrated location management and location-aided routing system for mobile ad hoc networks. Journal of Parallel and Distributed Computing, 66(3):367–378, March 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zane:2000:SNA

[ZMPE00] Francis Zane, Philippe Marchand, Ramamohan Paturi, and Sadik Esener. Scalable network architectures using the optical transpose interconnection system (OTIS). Journal of Parallel and Distributed Computing, 60(5):521–538, May 2000. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1627;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1627/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1627/ref.

Zambonelli:2001:DFI

[ZN01] Franco Zambonelli and Robert H. B. Netzer. Deadlockfree incremental replay of message-passing programs. Journal of Parallel and Distributed Computing, 61(5):667–678, May 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.2001.1703;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.2001.1703/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
2001.1703/ref.

Zhang:1993:MGT

[ZNQ93] Xiaodong D. Zhang, Naga S. Nalluri, and Xiaohan H. Qin. MIN-Graph: a tool for monitoring and visualizing MIN-Based multiprocessor performance. Journal of Parallel and Distributed Computing, 18(2):231–241, June 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1993.1059/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1993.1059/production/pdf.

Zomaya:1997:SIP

[ZO97] Albert Y. Zomaya and Stephan Olariu. Special issue on parallel evolutionary computing. Journal of Parallel and Distributed Computing, 47(1):1–7, November 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1997.1391/production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1997.1391/production/pdf;
 http://www.idealibrary.com/links/doi/10.1006/jpdc.
1997.1391/production/ref.

Zarrelli:2006:EPE

[ZPI06] Roberto Zarrelli, Mario Petrone, and Angelo Iannaccio. Enabling PVM to exploit the SCTP protocol. Journal of Parallel and Distributed Computing, 66(11):1472–1479, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2000:IMP

[ZR00] Guodong Zhang and Stephen S. Rappaport. Integrated multimedia personal communications with asymmetric services for mobile users in cellular systems. Journal of Parallel and Distributed Computing, 60(4):353–373, April 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.1999.1611;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1611/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.1611/ref.

Zhao:1999:DMM

[ZRC99] J. Leon Zhao, Doron Rotem, and Su-Shing Chen. Data management for multiuser access to digital video libraries. Journal of Parallel and Distributed Computing, 56(3):208–234, March 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1998.1519/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1998.1519/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jpdc.1998.1519/production/ref.

Zeng:2006:DSS

[ZV06] Zeng Zeng and Bharadwaj Veeravalli. Distributed scheduling strategy for divisible loads on arbitrarily configured distributed networks using load balancing via virtual routing. Journal of Parallel and Distributed Computing, 66(11):1404–1418, November 2006. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zeng:2009:NDA

[ZV09a] Zeng Zeng and Bharadwaj Veeravalli. A novel distributed architecture of large-scale multimedia storage system using autonomous object-based storage devices. Journal of Parallel and Distributed Computing, 69(4):349–359, April 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zheng:2009:DCA

[ZV09b] Qin Zheng and Bharadwaj Veeravalli. On the design of communication-aware fault-tolerant scheduling algorithms for precedence constrained tasks in grid computing systems with dedicated communication devices. Journal of Parallel and Distributed Computing, 69(3):282–294, March 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zheng:2000:DCG

[ZW00] S. Q. Zheng and Jie Wu. Dual of a complete graph as an interconnection network. Journal of Parallel and Distributed Computing, 60(8):1028–1046, August 2000. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2000.1640;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2000.1640/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2000.1640/ref.

Zhuo:2003:DRD

[ZWL03] Ling Zhuo, Cho-Li Wang, and Francis C. M. Lau. Document replication and distribution in extensible geographically distributed Web servers. Journal of Parallel and Distributed Computing, 63(10):927–944, October 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:2007:OPD

[ZWRI07] Mengxia Zhu, Qishi Wu, Nageswara S. V. Rao, and Sitharama Iyengar. Optimal pipeline decomposition and adaptive network mapping to support distributed remote visualization. Journal of Parallel and Distributed Computing, 67(8):947–956, August 2007. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:2009:OPR

[ZWS09] Zhao Zhang, Weili Wu, and Shashi Shekhar. Optimal placements of replicas in a ring network with majority voting protocol. Journal of Parallel and Distributed Computing, 69(5):461– 469, May 2009. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhao:2009:LBS

[ZXP09] Chenggui Zhao, Wenjun Xiao, and Behrooz Parhami. Loadbalancing on swapped or OTIS networks. Journal of Parallel and Distributed Computing, 69(4):389–399, April 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhang:1994:LME

[ZYH94] Xiaodong D. Zhang, Yong Yan, and Keqiang Q. He. Latency metric: An experimental method for measuring and evaluating parallel program and architecture scalability. Journal of Parallel and Distributed Computing, 22(3):392–410, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1100/production;   http://www.
idealibrary.com/links/doi/10.1006/jpdc.1994.1100/production/
pdf.

Zomaya:2002:FTR

[ZYO02] Albert Y. Zomaya, Adrian Yates, and Stephan Olariu. Faulttolerant recursive least-squares computations on a meshconnected parallel processor. Journal of Parallel and Distributed Computing, 62(7):1142–1167, July 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.idealibrary.com/links/doi/10.1006/
jpdc.2002.1840;   http://www.idealibrary.com/links/
doi/10.1006/jpdc.2002.1840/pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.2002.1840/ref.

Zeigler:1990:MHD

[ZZ90] Bernard P. Zeigler and Guoqing Zhang. Mapping hierarchical discrete event models to multiprocessor systems: concepts, algorithm, and simulation. Journal of Parallel and Distributed Computing, 9(3):271–281, July 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Zhu:1992:LNN

[ZZC92] Xianing Zhu, Shengwei Zhang, and A. G. Constantinides. Lagrange neural networks for linear programming. Journal of Parallel and Distributed Computing, 14(3):354–360, March 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).